A Bibliography of Books about the Internet and Networking: 1969–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

07 May 2024 Version 5.35

Abstract

This bibliography records books about the Internet and networking. It was started from the contents of Internet RFC 1432 by John S. Quarterman.

Title word cross-reference

#503 [Str95].

3 [Hal99, MBM^{+}95, Wol98]. $tuff [Vin94b, Vin95b, Vin96].

.net [Int95g].

1 [Sta99b]. 10-Minute [Ano97j]. 100VG [Cos94a]. 100VG-AnyLAN [Cos94a]. 10BASE [IEE90, Hel98]. 10BASE-T [IEE90, Hel98]. 10th [Mei97]. 10x [Ano96a]. 1100R [Hub93]. 12 [Ano93h, Ano94-47, BU94b, Bla93c, Car93a, Dav95, Din94, Ess95, Gol93, JM95, JM96a, JM96b, JT96, LJ95, LMLJ95, Mil94b, Mil94a, Mil95a, Mil95b, Mil96a, Moh95, NCS93, Nat93b, Sto95c, Str94, Int93a, Wil94b]. 12-step [Mat95a]. 12th [PE95]. 13 [New95b]. 1394a [And99]. 13th [Int76a, Int76b]. 166MHz [Ano96u]. 16th [Bro96]. 17-hands [Ano94v]. 17th [Int76a, Int76b]. 18-hands [Ano94w]. 18th [IDK98]. 1967 [Ano69]. 1976 [Int76b]. 1981 [Ano81]. 1990s [Gou90]. 1992 [Int92d]. 1993 [Lin93c]. 1993/1994 [Bla93a]. 1994 [Sto95c]. 1995 [Ano95l, Str95]. 19th [IEE94]. 1st [Cai94].

2 [Ano97i, Ive95, LO99, Mor95d, PP91, Pil90, Sal95b, TK96, Tys95, Vac99]. 2.0 [Jam96b]. 2.2t [Pot95]. 2000 [EE99, Web99, YY98]. 20th [IEE95]. 215 [Mil94f]. 21st [Ano94-61, BTC^{+}95, McK94b, McK95, Sac97, Wal95b]. 22nd [Ano99b, IEE97]. 23rd [IEE98]. 2nd [Ano98a, Ano99a, LB95a, LY96].

3 [HB98b, Hen99, Par93]. 3-D [Hen99, Par93]. 31 [KSW99]. 31-August [KSW99]. 3D [McN95].

4.5 [Pre97b, Pre97c]. 4.x [MK98a]. 4.X. [MK98b, MS98, MK98c, MK98d, MK98e, MK98f]. 400 [LS98]. 4th [Ano88a].

5 [ELB99, Mil99a]. 5.5 [Gon99]. 56Kbps [Wil98]. 57th [Ano97a]. 5th [HL99, Pla94b].

6 [Web95]. 63rd [Ano97a]. 6th [Ano79, GRS79].

7 [Byr99]. 7-ilgan [TI95]. 7.0-1.0 [Ano99e]. 77th [Ano96x]. 7th [Ass95b].

802 [BDPS85]. 802.3i [IEE90]. 802.3i-1990 [IEE90]. 8th [VVB85].

90’s [VVB85, OO95]. 94 [Ano94-33]. 95 [Sne95c]. ’96 [ACM96, VT97, Ano97i]. ’97 [Hir97, IEE97, Ano97i]. ’98 [Hir98, KW98, NDM98]. ’99 [Ano98a, Fra99, HL99]. 9th [Ano88a].

= [Edd94b, Edd95a, Edd95b, Hen95b, LGRS93, LB95a, MY94, Rik94, SG94b, WLM90].

A-L [Ano93a]. AALS [Ass95a]. AARP [DHM^{+}96]. ABCs [Cru96a]. Abdominal [EO99]. ABOARD [Des93a, Des93b, Des94]. Abstracts [Ano79, TL93, ACM99b, GMA^{+}98, Int94a]. abuse [PS95a]. AC&NC [Ano99e]. acadamia [Bak94]. academe [OO95]. academia [Lin93a]. Academic [Bre93, KKO94, Ril95, Ano93t, Ehn94, Hic94, Wil94b]. academica [Fre95]. ACCD [Sau92]. Accelerated [Ano96u]. acceptable [Fin95]. acceptance [BG98]. Access [Ano93m, Ano93p, Ano95-59, Ano96v, Bla98f, Bla93c, DGH93, E^{+}93, HSF94, IEE90, Jam94, LH93b, Lic94a, Lic94b, Mel94, Not94a, Not94b, Ano93-29, Osh93, Pea94, Pea95, PM94, Rey93, RS97, TR95, Tus93, Wal93, ANS92, Ano94z, Ano94y, Ano94u, Ano94-55, Ano95-30, Ano95-31, Ano95-29, Ano95-61, Bri94a, BKH95, Cle93a, Eme94, Eme95a, Edm95, Eng93b, Eng95a, FHS94, Gal94, Gil95e, Gra95, Hor98, Hug94, Inf93, Int93d, KK95a, KG98, Koh97, KG97, KP94, LH93a, LE96, Lic95a, Lic95b, LF95, Mad96, Max95a, MW94e, MBC^{+}96, MP95, MS94, Owe93, RM96, Sar94, Sch96a, SSI^{+}98, SC95c, SH95b, TK95a, Wag95b, Wil96a]. Accessibility [LG99]. accessible [Ano90c, Int89a, SL91, SLR91, SL92]. Accessing [Ano93-46, Bri95c, Bra94b, Fei94, Gar93, HKB95a, HKB95b, HK95a, Leg94b, PR91a, PR91b, RP93, Smi94a, US 92, US 93, Ano95w, Cal94b, Cal94a, Hug94, Leo94b, Mil94b, Pet89]. Accessories [Hew86]. accommodations [MBZ94]. accompany [RN97, Str94]. Accomplishments [RAD90]. Account [Not93b]. accounting [MHR91]. accounts [Owe93]. Achieving [SJZ^{+}98]. acht [MW94c]. acquiring [MS94]. acquisition [Sha91b]. acquisitions [SLJ94]. ACRL [Ass95b]. Acrobat [SP95b]. Across [CP99, EST93d, GS95c, San94a, Sch92a]. ACS [Lar94a]. ACSM [Ano97a]. ACSM/ASPRS [Ano97a]. Actes [Ano69, Min94b]. action [U.S93]. Active [BN96]. ActiveState [Ano99d]. activism [Zel95]. activists [Ano95-45, Ste91a].

Activities [Cer89, DGH93, Ash95, BI96, Cer91, Dav95, FK95a, FP95a, Vir97]. Actually [Hug95]. Acu4GL [Ano96v]. Acute [EO99]. adaptation [BK99a, PH97]. adaptations [KB92]. adapted [KB92, Ril95]. adapter [VFH98a, VFH98b]. Adaptive [FK99, Sne95a]. add [KH95]. add-on [KH95]. addictions [KL95]. Adding [Des94]. additions [Pri95]. address [Bar92b, SV99]. Addresses [God94a, Bar93, KK96a, Lea95a, Rou94]. Addressing [FRB93, FA89, FA93, FA94, Mau99b, Fra94c, Hel94, Pet86a]. Administering [Byr99]. Administration [Hun92, Hun98, LG93, ML94, Neo94b, Unixx, Her93]. Administrator [Moh99b, MP92, Mui95]. administrators [BL93, SHN97]. adopter [Eme95b]. adoption [Eme95b, Lee98, Woo95b]. Adressbuch [KK95b]. Ads [Hes95, HH95b, HB95]. ADSL [Sum99]. adult [Ela94, EHPT95]. adults [Bro94a]. Advanced [Ano96v, Bla99c, MW94d, BU94a, Ham99, YM96]. advantage [Int94e, CB95d, Cob95, Cro96a]. Adventure [Aik95, Lei96, Her95b, Woo96]. Adventures [BS96, Ash95, Her95c, LS95, LS96]. advertise [Jan95, SB94]. advertising [SB94]. advice [Ros96]. advisor [Int96b]. Advisory [Ano95f]. Aero [Sar94]. Aeronautics [LG93]. aerospace [Har94b]. aesthetic [Kac96]. Aether [WS94]. affect [Dra95]. afraid [Thi93]. Africa [Hum89, WR89, Dix95]. African [BH95, BH96, BH94a, Din94, Int96e]. After [Edd94c, Woo94, Don95, Edd96, Woo95a]. Aftermath [Spa89a, Ano88b]. Again [Gar97]. Age [Moh99a, AH95b, CRI95, GP95b, Ano93n, Tur95a]. agenda [Ano95-65, U.S93]. Agent [Ano96v, WYLO98, WYLZ99, Bup95, COZ99]. Agent-Mediated [WYLZ99]. Agent-Trudger [Ano96v]. agents [Che96, KW98, KSW99, KG99b, MDW99, Wat96, dAMW97]. Ages [Opp95]. aging [DHM^{+}96, Pos95]. Agony [Bai94]. Agreement [Gar97, Ano94-55]. agreements [Bla99a]. AGRICOLA [San94a]. agriculture [Cam95c, Dre92, Dre93a, Dre93b, Dre94, Dre95]. ahead [BGMP98]. ai [BZ95]. aid [KK96a]. AIDS [Ano95a]. air [Sil95]. Airlines [Hen95a]. AirLinks [Sil95]. al [KMM95, Kro95a]. Alameda [Ano95-33]. Alert [Sto95a]. Algorithm [YN97]. Algorithms [Leo98a]. Alison [KB92]. All-In-One [GM94a]. alla [Pas95]. alle [Pas95]. alleys [Ano95l]. Alliance [Gar97]. allied [Edw95]. Almanac [Rit93]. Almanaque [Ano95b]. Almost [Kno96]. along [BDC^{+}95]. alsu [SG94b]. Alta [Ano99d]. AltaVista [SRR97]. Altera [Ano99e]. Alternative [Cro94b]. amazing [Eng95c, WC95a]. America [BGS94, BGS95, GBS95, WLM90, Ano94-60, Lic94a, Lic94b, Lic95a, Lic95b, SPR98, Tat95a, Wri93]. American [Ano94a, BH95, BH96, Ebi95, WS80, BH94a, Cro94a, Cro95, Mol93a, PU94, Smi95a]. AMIA [Cim96, Mas97]. Amiga [Lar94b]. AMIGOS [Hic94]. among [Hic94, Sch90]. amplify [Shu93, Eng93a]. Amtec [Ano96u]. Analysis [Ano99c, Pau99a, SC93, She95b, Spa89b, sC86, GH94d, Gla94, Hor98, Liu95, NC90, Pau99b, SC92, SW93, Smi95a, Spa88, SC88]. Analyzers [Hel91a]. Analyzing [Mal90, Mil92a, Mil96b]. Anatomy [FPV95, Moy98]. Andrew [Ano96v]. anesthesia [IDK98]. Anfanger [LB94b]. Anguilla [Fra99, Hir97, Hir98]. Animadversions [Pad85]. animal [Bos94]. animal-related [Bos94]. animation [Ham99]. Anixter [Ani93a, Ani93b]. Annotated [Mog89, JE96, Liu95, Liu96, Mog88]. Announcement [Ano98a, Ano99a, Nat96b]. Announcements [Ano98b]. announces [Ass94a]. Annuaire [And95c, HS95c].

Annual [Ano97a, IEE98, JG91, LG93, Mas97, PE95, Pla94b, Ass95a, ACM99b, Ano94a, Cin96, Bro96, Cim96, IEE97, Lei93, TL93]. Anonymizing [Mar98]. anonymous [Fel93, Fel94b, FK95a]. ANSI [CS98, Mad89]. ANSWER [Ano94q]. answers [MM91, MM92, Sav93]. anthology [Fin95, Ano95c]. Anti [Arn95]. Anti-net [Arn95]. anticipate [RRS98]. AnyLAN [Cos94a]. Anything [SRR97]. Anyway [Har93b, OO95]. Anywhere [Gar95a]. AOL [Lic94a, Lic94b, Lic95a, Lic95b, Woo95a]. APIs [Ste98b]. Applet [Ano96v]. Applets [Ano96v]. Appliance [Ano97f]. Appliances [Col99]. Application [Fou98, HB98b, Heg99, Lam96, RG91, Tow95, Bra89a, D^{+}86, Eas94, Gre99]. application-level [Eas94]. Applications [Ack95a, BBH93, Buy99, Car98b, CRH98, CS93b, CS93a, CS97, Dig90b, Fee93c, Fur99, GW99, HHS94, HS98, Leo98g, Leo98f, Leo98b, Lov99, MOD93a, Mad89, MB93, Mes99, Ogb98, Pad93, RV99, SC93, Sei98, Smi93a, WA98, ACM99b, Ele91, Ano94-55, Bas99, BK99a, BEA98, Bou98, CKTV96, Com92, Com94a, CS96a, CB96b, Dig95, ECC94, Gor95b, Gre95c, GMA^{+}98, GP95b, HL99, Int93b, INP95, Leo98a, Leo98d, Leo98e, Leo98c, LC98, Mar95b, MKT98, Mil99b, NGP^{+}99, PG96, PA95, PKL90, SC92, Sha91b, Tho96b, VE96, Ano94-50]. applies [Fol95b]. Applying [DG97, SP94a, Swi96]. apprendimento [Pas95]. Approach [Hel91a, MMLP97, RG93, BBN^{+}93, CKY94, Eas94, Fan95, Kes97, Pet86b, Xie94]. approaches [BDMS94, Sch92b]. Approaching [MKBB95]. Apps [Kuo97]. Aprendiendo [Ran95a]. April [Ass95b, Ano97a, Ano99b, LG93, Int92d, Str94]. Arab [GP95a, GP96a]. Archaeology [Hol94]. Archetypes [Ste96a]. Archie [SBGK94, Gil94a, Gil94b, Ano92h, Har93a, ED92]. Architecture [And99, BW97b, CC83, Com88a, Com91b, Com95c, Com88b, Fei93, Fei96, Fra93c, KP93, KSM98, ML94, MSD90, ME95, Sum99, BM98, CKTV96, CCC^{+}91, Gre99, Gup92, Hal93, HCBO99, Kar96, Mol99, RT98, SR99, SP94a, JGKR96]. Architectures [Bla97b, Heg99, Leo98a, Smy95, Hal97, HGKB99]. archival [FB91]. Archive [Del93, Han96, Sim93, Ano94-38]. archived [HW95]. Archives [Ano96t, IRC95, Sch90, Str92, Str93a, Ano95-55]. Ardent [Ano99e]. ARDI [Ano96u]. Area [BDPS85, Bla98a, Cah98, Dig82, DM95, LAN86, McN85, Sta93a, Sta94a, Wal93, ANS92, Cac89, GS95c, KT92, Mai95, Shu93, TJ97]. areas [Coo93b]. Argument [CES^{+}99]. ARIADNE [AAB^{+}98]. armadillos [MLJ95]. Armchair [CES^{+}99]. ARPA [Abb94, PSC81, Ros84]. ARPA-sponsored [Abb94]. ARPANET [Abb94, Ano93h, Bol94, New81, Sal95c]. Art [GG98, Keh93, Keh94, Pad85, Res97, B^{+}96b, Bur95a, Hei95b, IDK98, Keh92, Keh96, SY95]. Arte [KMM95]. Artecon [Ano99d]. article [OO95]. articles [Hub96]. Artificial [OKP97]. artresources [Fer95]. Arts [Kan94b]. AS/400 [LS98]. ASCILITE [PE95]. Asia [HT95, Mul92]. Asian [Ebi95, SR95b]. Ask [SW97, Thi93]. asked [Iow94, Int95b, MM91, MM92, Rot95b, Sav93]. askew [Ros96]. Asking [Saw93]. Asnuntuck [Ril95]. aspects [Bro96, KRS94, KRK95]. ASPRS [Ano97a]. Assembly [Ano95-61, Ano96w]. Assessing [Dil93, DJB93, Gal94, Ano95-61]. assessment [San92]. assist [Sta98b]. Assistant [KH95, Roa95a, Ros96]. assisted [Bri94a]. assistito [Pas95]. Associates [Ano99e]. Association [Ass94a, Ano99b, Ano94a]. AT&T [ATT93, TT95a, TT95b, CS93b, Com94a]. Atlantis [Ano94b, BW93]. atlas [BL94]. ATLAS2000 [FMS98]. Atlases [FMS98]. ATM [CKMV99, Ano95-51, Ben98, Bla95a, Bla98b, Bla98c, Bla98d, Bla99d, Bla99b, BEA98, Cam97, CX97, CL95, GG98, HHS94, Ibe97, IMI^{+}95, Ker98, Kes97, KSO95, ME97, SKB96, SC92, SC93, Sta95a, TJ97, Wu97]. ATM-based [CX97]. Attachment [IEE90]. Attack [Den90]. attacks [dVdVI98]. attitudes [Rus94, Sch95b]. attorneys [Ano96q]. attractions [BDC^{+}95]. Auctionlike [WYLZ99]. Audio [GGA^{+}97, GAM^{+}97, HSK98, WLKS98]. audio/video [Tho95b]. audiocast [CD92]. Aug [Ano94-33]. August [B^{+}96b, GV90a, TL93, KSW99, Lei93]. Australia [HT95]. Australian [PE95]. Austria [VT97]. Authentication [Opp96, Tsu89, Tun99]. authoring [Pao96, QRV95, Web95]. Authors [Fly95]. Auto [Ano99e]. Auto-Changer [Ano99e]. Automated [Ano96w]. Automating [Mel94, SK98]. Automation [Ano94c, The96c, PN95]. automatizada [Ubi95]. Autonomous [GS95c]. autopista [CS95]. autres [Pla95]. aux [CD94]. Available [Ano92a, Ano93f, Ano93d, Ano93g, Ano93-31, Wol98, Ano92f, Ano95-61, Rya94, Rya95a, Rya95b, Rya95c, Rya95d, Rya95e, Rya96a, SUR92, Uni93a, Uni93b, WoRIGSoO95]. Aviation [Avi96]. avviso [BZ95]. Aware [Kuo97, MW92, MSB^{+}92]. Awareness [Tro94, Pri95]. AX.25 [Rei92].

B [Fol95b]. Back [Amo99, Ano95l, SLB94, TMF97]. Backbone [Ano93-37, Aga99, CNY87]. Background [MHR91, MW92]. Bah [Sto95a]. Balance [CES^{+}99]. Baltics [Int95b]. Bamba [WLKS98]. banche [Pas95]. Bandwidth [ALS96, CP99, Ell99, KKC98]. bank [Wam95, SPR98]. banking [Kol96]. barely [Hof96b]. bargains [Win96b]. Barron [Bar93]. Bartleby [Van95]. BAS [BAS95]. BASCOM [Ano99d]. Baseband [IEE90].

Based [BB99, Cam97, CRH98, EO99, GS99, Rak95, RV99, Ben95c, Bre95a, CX97, CB96b, EW94, Gla94, Har99a, Hug94, KSM98, KTS94, Lee98, LC98, Mac95b, McL96b, Pet86a, PH97, RR94b, SG93b, Ter99, ZCD97]. baseline [Mle95]. Basic [Ano93r, Ano93s, Ano95l, BW94a, BWC95a, Fra96, LS93a, Mar95b, VH94e, Ehn94, Fel93, Fel94b, FK95a, LS93b, McQ94a, dVdVI98]. basico [MPF95]. Basics [Ano94d, LH93b, TE99, Abr98, Ens95, Gre95a, How95a, Int95k, IV95, LH93a, Nof95a, SLB94, Ten92, T^{+}98]. Battle [Ano93-37, DFS98]. battles [Kan94b]. bazaar [Wat93]. bazak [HY94]. BBN [Bol94]. BBS [Ano96s, Ban95, Pla95, Tom93]. BBSs [Lar94b]. Be [RG93, Ano94-48, Bie93, Cla95a, Fol95b, Ken90, Law95, SY95]. Beagle [Dar93b]. beasties [Wil96b]. Become [Ano93g]. becomes [Tur95e]. becoming [New95b]. before [Ano94y, Ano94-32, Ano95l, Bro89, SRS^{+}95]. Beginner [Ano95d, Cas94, Eva95, Keh93, Keh94, LR93c, LaQ94b, BS96, Keh92, Keh96, LR93a, LR93b, LaQ94a, LDH95, Mil96c, Par95, PM95b, Sua94, Sua95]. beginners [Bur96]. Beginning [MB93, SS94b]. Beginnings [Cat89]. behavior [BPS99, Pax97, San92]. Behind [Ano93-43]. being [Mai95]. beliefs [Sch95b]. believe [Dra95]. Belong [Sto99]. belt [Ano97b]. Benchmark [Ano99d]. Beneath [HC95]. Benefits [GS99, Uni95a]. Beowulf [Gum93a]. bereavement [JP94]. Berkeley [Zho84]. besieged [DD98]. Best [Bal93b, Ano95e, Inf95a, Max95b]. bestseller [SG94b]. Better [Lam95, BGMP98]. Between [Ano99d, Haa85, PS92]. Beyond [Ass94b, Abo93, Ano94e, Ano95f, BW97a, Cim96, Fee93a, Fee93b, GMB93, MP91, Opp97, Pow98, Sal95c, Abr98, Ano95-63, Ano95-65, Eas95, Esc98, Flo95, How95a, Kle94, McM94, Nie95, Ott94, Res97, Smi95b, Tex95]. BGP4 [Ste99b]. Bible [Mor95a, Pfa95g]. biblia [QP95]. bibliographic [Bar92a, Bar92b, Bar93, Int94a, MS94]. Bibliography [Bee93, Mog89, Ano94-34, Ano96l, Bol94, Hum89, JE96, Kot93, Liu95, Liu96, Mog88, Str92, WR89]. Bibliotheken [Jun95]. Big [Ano93w, EKG94a, EKG94b, GKC93a, GH93, GKC93b, GH94a, GH94b, KG94, PE94, PG94, HB96, Car99b]. Biggest [Chr98]. biginazu [LGRS93]. Bill [Mil94d]. Billy [Bar93]. Bimodal [BHO^{+}99]. BIND [AL92, Alb93, AL94, Zho84]. biodiversity [MM95]. Biological [Mac93a]. Biologist [Smi93b, Smi93d, Ano94-31, Smi93c, SMM96]. biologists [Ano96a, Fre93]. biomedical [Lea95a]. Biosis [Ano92a]. biotech [Lea95a]. Biotechnology [Bio96]. bird [LeG95b]. Birmingham [Int76a, Int76b]. birthday [ACE^{+}95]. bis [SBGK94]. bit [Kur96a, RT99]. bite [RAWS95]. Bitnet [Noo92a, Noo92b, Noo94, Noo91a, Anoxxa, Bos94, Dar93a, Dre92, Dre93a, Dre93b, Dre94, Dre95, Han92, Han93, Han94a, Han94b]. BITNIS [Chi93]. BitWizard [Ano97h]. bizzare [Woo95a]. Black [Ano97b]. Blackbird [Reh96]. Blacksburg [Wie93]. blanc [Kap96]. BLS [Uni95a]. Board [Bry94, Bos94, Hel94, Ano95-61, Cer89]. Boards [Abo93, Mot94, RWD95]. bodies [Shi96]. body [Ano95e]. bolsillo [Hof95g]. Bomb [YY98]. bonnes [Vin95a]. Book [Ano94f, Ano94g, Ano95g, Ano96b, Ano96c, Ano97c, Ano97d, Bar94, Chr98, Com97, EE94, Hop95, Hum98, Laz97, Pfa95f, Roo97, Ros91, Tem95, AJR95, BT96, Bud96, Com95a, Com95b, EE96, Ive95, Jam96a, Jam96b, KK96a, MMJ95, Pla94a, Ros93, Ros94b, Sch95c, Sch95d, SS95d]. Books [Bee93, Qua93, Bro94a]. Bookshelf [Mor93, Ass99, TA96, Ano96r]. boost [Ano95-65]. boot [PR95]. border [RG91]. borderless [Jar95]. Boston [Ano81, Ano91]. botanical [BOB93]. ’bots [Che96, Wil96b]. bottom [Coo95a]. Boulevard [ACM99a]. Box [Ano94k, Och95, SPR94, KP94, SPR95]. brain [Sun95]. Branding [Gon95]. Brave [Ert95]. Bridges [GD98, Per92a]. Bridging [Int76a, Int76b, Pot93, Pot94a, Pot94b]. Brief [Ano93a, Ano93b, Sou93, HG95]. briefing [Gry94]. Briefs [Gar97]. Brin [MSB^{+}92]. Bringing [Dyr93, Ano96o, Sto95c, Vac96b, War95]. British [Fra99, GV90a, Hir97, Hir98, Ano93m]. broad [Hen95a]. Broadband [Bla95a, Bla98c, Bla98f, Bla99d, CWJT99, Fla91, KSO95, Sta92, Sta95a, Har92, HRB92, Koh97, Kam93]. Broadband-ISDN [Kam93]. broadcast [Swa91]. broadcasting [Bog82, BX83]. Broker [TGL^{+}97]. Brokerages [HWV96]. Brokering [BWWF99]. brokers [Che96]. Brown [KB92]. Browser [Ken95c, CB96b, McN95, Tho95b]. browser-based [CB96b]. Browsers [Rau95]. Browsing [Mui94, Sne95a, Bir93, DG95, Gil94b, SK98]. Bruce [MSB^{+}92]. BSCW [PH97]. BSD [Com92, CS93a, Hub96]. BSD-UNIX [Hub96]. Bucher [Ble95]. Buddhist [SY95]. Buffer [Sfe82]. Build [LE95, Net96a, Atk99, Bea94, HN96, NGP^{+}99]. builder [Ano95-40]. Building [Aga99, Bay99, Bla99a, Buy96, CZ95, Dig90b, Eck95, FPV95, FD94, GL99, Gre95c, HE95a, HAS95, HL93, Kah92, KJGH98, Lov99, Mai96, Mar95b, Row96, ZP98, Res95a, BT96, Bud96, GRLJ94, Har92, HRB92, MMJ95, Rya95f, Rya95g, Rya95h, Rya96b, Wit96, Ano96b]. built [Mar99]. Buku [Rah94]. Bulletin [Abo93, Ano95h, Bry94, Int94d, Bos94, Hel94, Mot94, RWD95]. Bureau [Anoxxc, Ano95-72]. Business [AAS^{+}97, A^{+}98, Ano93c, Ano93f, Ano93v, Ano93o, Ano94h, Phi95, Bay99, BBSH98, Bla96c, CES^{+}99, Dav96, DN99, EE94, GS99, GP95b, Her96a, Ano93n, KRS94, KRK95, Ogb98, PA95, PT95a, PT95b, Pee96b, Shu99, Sta95e, Ame95b, AP95a, All96, AH95b, Int94e, Int94h, Net94d, Int95o, Sma95, Net96b, AT94, Ber95a, B^{+}96a, Bla95c, Boc96, Bro96, Cam94b, Cam95a, CB95d, Cob95, Cro93, Cro94a, Cro95, Cro96b, EE95a, EE96, Eme95c, EV96, ECC94, Foc95, Hof96b, HB96, Int93g, KLW95, Mah95, Mal95b, Mse90, RT94b, RT94a, Ric95a, RR94a, RL95b, RJV95c, RJV95d, RA96b, SSC96, Sch96b, Sme96, WM95, Wes97, Woo96]. Business-to-Business [Bay99, SSC96]. businesses [Jan96, Mad96]. BusinessLink [Bus95]. busy [Cru96b, Mog95]. Buyer [Dig93, Est93b, Bre95a, Est93c]. Byteline [Bur93].

C [AB94, Ano96v, CS98, Fol95b, Gre95c, JCD96, Mas93, Rog97, Vol97b, Vol97a, dG93]. C-Kermit [dG93]. C/C [Vol97b, Vol97a]. CA [NDM98, Uni96a]. Cable [CFL99, Hew86, TE99, Gil95a, Gil95b]. Cabling [Hew88, Vac98]. Cache [YN97]. caching [Hal93]. cafe [Uni95e, Rat93a, Rat93b, Rat93c]. Calamities [Gla99]. calcolatore [Pas95]. Caldera [Ano96s]. California [Ano94-53, Cal94b, DLC92b, JG91, TL93, Uni96b, USE97, Whi95, Cal94a, Lei93, Wan94, Zha94]. call [IBM89a]. called [Ril95]. calls [BA95]. Camelot [Spe87]. camera [AUA94]. camera-ready [AUA94]. camp [PR95]. campus [Sin81]. campus-wide [Sin81]. Can [Bie93, Ano95-45, Cla95a, Jan96, Nov94, Ric95a]. Canada [ACM96, Car95b, ILSB94, Cle93a, Cle93b]. Canadian [CB95d, DOO95, Eis95, CB94a, CB94b, CB95c, Ric95a]. cancer [Ano93y, Kra94, SS95d]. Cansearch [Kra94]. Can’t [Mui95, Ano95-45]. CapAccess [Wal93]. Capacity [BMK88, MA98]. Capital [Wal93].

Capitalizing [CTN95, GCW95]. Capitol [Ano95i]. Capture [RS97]. care [Ano94-52, Coo93b, Hob96, IDK98, Mas97, McC95]. Career [Ins96, KK96a, Mle95, Ril95]. Caribbean [WLM90]. Caribe [WLM90]. Carlos [Mou96]. carnet [Vin95a]. Carrier [ANS92, IEE90]. carriers [Hen95a]. Case [Chr98, GK95, Mog99, RV99, Sti94a, Tri91, Bag95, Cla99, Fos91, FK95b, Gal94, KIP99, Lin93a, SPR98, Sti94b, TW95, Tri93]. cash [RD95]. Casting [Bir90, MPF96, Sal95c]. Catalog [ATT93, Ano92i, Dig90b, Hub93, Kro92, Kro93, Kro94, LeB93, MOD93b, Net93, Ano94q, Ano95-69, Kro95b, KK96b, ML95, Who93, Ani93b]. Cataloging [Nic95, Ols95, Jou97, SLJ94, Sha95a]. Catalogs [Hes95, Ano90c, Bir90, HH95b, HB95, Int89a, SL91, SLR91, SL92]. Catamaran [Ano94-53]. CATRIONA [Nic95]. cats [Hei95a]. causality [SJZ^{+}98]. Cause [Ano93d, EGC^{+}94]. CBI [WMS96]. CCITT [Ano92e]. cd [RCL96, ANS92, Ano94b, Ano95-68, Har96a, IEE90, Ins96, Joh96b, LL95a, Mor95b, Neo94a, NR96, Ass99, Pow93, Wal94b]. cd-rom [RCL96, Ano94b, Ano95-68, Ins96, Joh96b, Mor95b, Pow93, Wal94b]. CD-ROMS [Har96a]. CDMA [Vit95]. CEAL [Ass94a]. celebrities [Ber96b]. cell [dAMW97]. Cells [Fla91]. Censorship [AN96, Her96b, RM96]. census [Cha95a]. Center [ACM99a, Ano94-50, Gro90a, Inf93, Ril95, Ano94-55, Net96a, SY95, Ano97g, Car99b]. centered [Gan95]. centers [Fol95a]. centers/institutes [Fol95a]. Central [Inf93, Gal94, Ram95]. Centre [Int76a, Int76b, Ano93m]. Centric [Col99, Coh99]. Century [McK94b, Sac97, Sta98a, Ano94-61, BTC^{+}95, McK95, Wal95b]. CEO [HH95a]. Cerf [And92, Ano93-43]. CERFnet [Cal90]. CERT [Car99b, FF94]. CGI [IH98, Row96]. cha [LY95b]. chaek [SG94b]. chairman [Uni89, UU89]. challenges [Cal91, McC92, MMR93, MMR94]. Chameleon [Ano94-42]. change [Bup95, WoRIGSoO95]. Changer [Ano99e]. Changes [Ano97e, Sny96, Sch91a, Sch91b]. Changes/Linux [Ano97e]. Changing [SQ93, SP94a]. Channel [Tho99, ER98]. channels [BM98, Hen95a]. chao [AJRmT95]. chaos [Eag93]. Chapter [Ano91, Web95]. Character [Sun96]. Characteristics [Woo95b]. characterization [AW97, Cla94]. Characters [DFG^{+}98]. Chargefree [BGGS95]. Charlotte [Ano99b]. Chart [Ano97h]. chasseur [How95c]. Chat [Kin95b, Rei93, Ros95a, Pyr95, Har95b, Rei93]. Che [Sal95d]. Cheap [LE96]. cheapest [BA95]. Checksum [BBP89]. chemical [IE98]. chemin [Ano69]. Chemistry [Bac95b, Wig93]. chemists [Bac96]. cheng [HSskw95]. chi [AJRmT95, cH95]. Chia [cCrmdh92]. Chicago [BCB99, Hay94a, Nor93]. chieh [Edd95b, kTlC94]. chih [cH95, Neu98a, kTlC95]. Child [PS95a]. childhood [Bla95d, Rot95b]. Children [Bro94a, Ano95l]. China [Cin96, HL99, TMF97, TFG99]. Chinese [FK95b]. Ching [Hen95b]. Chip [Gar97, Ano96u]. chips [Ano95-51]. choice [Her96b]. Choosing [Joh96a, Pow93]. choreograph [MO92]. Christian [Bak95, Bak95, Boy96, GR94, Gal95b]. Christianity [Ano96k]. Christians [Sch95e]. chronic [JP94]. chronicle [Ano94-58]. chronology [Ano93-36]. chu [cCrmdh92, hLcL95]. chun [cCrmdh92]. church [Wri95b]. ci [Sal95d]. CIA [KW98]. CIA’99 [KSW99]. ciag [LY94a]. ciberspazio [Gia94]. cifras [Ano95b]. Cinfolink [Cin96]. circuit [Con89]. Circulation [Min94b]. citations [Ano94-33]. Citing [LC93, LC96]. City [Mei97, Ano93w, IDK98, Sch95b, Sea90, Car99b]. CIX [SI 93]. CJK [Sun96]. Clara [Tee94]. Clarines [Ano93e]. Clarity [VVB85]. Class [Ano96v, SS95f]. classes [ALS96, RM94]. classical [Ehn94]. Classroom [Ben95a, Dyr93, GMDS95a, Gia95b, Sto99, AUC93, Ame95c, Ame95d, And92, Bri95d, Bri95e, BTC^{+}95, Cot95, Dav95, Ess95, GMDS95b, GMDS95c, GMDS95d, LJ95, Moh95, PHW^{+}93, PW95b, NAS93, War95]. classrooms [LMLJ95, Rot95a]. cleans [Ano89b]. Clearinghouse [GBJ94]. Clearly [Har99b]. Clés [SLP94, SLP95]. Click [God94b, Bri95b, Ken95g, KP94]. Client [Ano99e, CS91a, Com92, CS93b, CS93a, Com94a, CS96a, CS97, Fou98, INP95, Ano97b, GGK^{+}99, dAMW97, New95c, Par93, Sne95a]. client-relationships [GGK^{+}99]. Client-Server [CS91a, CS97, Com92, CS93b, CS93a, Com94a, CS96a, dAMW97]. Client/Server [Fou98, INP95, Ano97b, New95c]. clinical [McC95]. CLNP [KF93]. Clock [ASB97, Bla96b]. closing [Ros93]. clothes [Moo95]. Clothing [Lam96]. Club [Bak93b]. Cluster [Ano99d, Buy99, Pet86a]. Cluster-based [Pet86a]. Clusters [SYTM98]. Clustor [Ano97h]. CMIP [Sta93c]. CNN [Com95d]. co [Ste91a]. co-sponsored [Ste91a]. coach [Net95c]. Coalition [Boy96]. coaxial [CNY87]. COBOL [Gar97]. CocoBase [Ano96v]. Code [Zim95b]. codes [IBM91a]. coexistence [Ano92e, Son92]. COGIS [Uni93a, Uni93b]. Cognition [Nil93]. Collaborating [PRY99]. Collaboration [Pet93b]. Collaborative [CTG96, Sky99, NGP^{+}99]. Collaboratory [PRY99]. Collecting [AR98]. Collection [AR98, BD88, Bre94, Hub96, MW92, Tho96d]. collections [Mul92]. College [Ano94m, Bro95a, BL93, Cor96, Glo95, GG96a, ODO95, Ril95]. Colleges [Mei97, RG93, Ale95, BL93, Boy96, Mar96b, MS95, Mle95]. Collision [IEE90, ANS92]. colloque [Min94b]. Colorado [IC90, Ano94j, Sch93b]. Columbia [GV90a]. Columbus [PMB^{+}92]. COM [Lov99]. combinés [Vin95a]. Combustion [PRY99]. Comdex [BR95]. come [GG96b, GP95b, SE96]. comes [CL98]. Comfind [Foc95]. comfortably [WC94]. coming [Ano88b]. command [IBM88]. commands [Ehn94, Fel93, Fel94b, FK95a]. Commentary [LG99, Mue99a, Nov94]. Commerce [A^{+}98, Bro89, Lit93, NI95, NIP96, O’R96, Pyl96, TS98, Ubo95, WYLZ99, DL96, Hor95, KW98, LM98, LL95b, LL96, Net96a, Pel95b, Wri95a, Ano93f]. CommerceNet/Nielsen [CN95b]. Commercial [Ano90a, Ano94i, Bhi96, SI 93, Ano93c, Bos94, ECC94, GP95a, GP96a, Ano93n, Jan95, PA95]. Commercialization [Pre94]. Commission [OD95]. Committee [Ass94a, Ano94y, Ano94-32, Ano95f, Ano95l, Bro89]. commoditization [Jon94a]. Common [Ano96d, Ari93]. commonly [Iow94, MM91, MM92]. commonsense [Sch96a]. comms [Int95c]. communicating [QCM94a]. Communication [Com71, GW95, KPS95, LS89, Vit95, Wil94a, AHS90, Ano94m, BD95a, Bet87, Dec95, Dil93, Fre93, HSK98, JO95, Kon91, Sch92a, SY95]. Communications [Ano93l, Bla97a, Car98b, Car99a, CFL99, Com84, DJB93, Dut98, Eag95a, Far97, Goo97, HL93, Iow94, KW93, Mil91a, Mil95c, Mit91, Osh93, Rea98, Sta95e, Ste99a, Wes98, And95a, Ano96d, Bra89b, Cha95a, Don95, J3 95, New94b, New95c, New96, Ros95a, Tri91, Tri93, Dat74]. Communicators [Fre94, Eme95b]. Communities [Ben95a, RG93, Bry95, Coo96]. Community [Ano98b, FPV95, Rhe94a, Rhe94b, Ril95, Wal93, Ale95, Bri94a, Int95p, Lof95, Nor95, Rik94]. Community-Technical [Ril95]. commutée [Cle93b]. Como [CS95]. Companies [Gla99, Cro94a, Cro95, Gor95a, Lea95a]. companies/organisations [Lea95a]. Companion [Ari93, HA95, LR93c, LaQ94b, AH95b, Bak95, Ben94, Ben95b, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95c, Gia95d, Gia95a, Gia95b, GMDS96, Ken95g, LGRS93, LR93a, LR93b, LaQ94a, LDH95, McK96b, QCM94a, SC94b]. Company [Ano99e, Lan94, McC97, Net96a, Kar96]. Compare [CES^{+}99]. Comparison [Fra94c, Sch92b, ZCD97]. Compatibles [LR93d]. compendium [Dav96, RJV94, RJV95e, RJV95a, RJV95f, RJV95b, RJV95c, RJV95d]. competitive [Inf95a]. competitiveness [Cro96a, Shu93]. Complete [Bra93b, Car95a, Dup95, HS94a, Ken94b, Kin95b, McF95, SPR94, Ben94, Ben95b, Bre95a, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95b, Gol94, Gon98, Gre95b, Gre96, Hah96a, Hel94, Hub96, Ive95, Ken94c, Ken94d, Ken94e, Ken95d, Ken95e, KP94, LR95b, Mal95c, Mor95b, Pan96, RK95c, Sch96b, Str92, Str93a, Wei95a]. completion [Ano88a]. complex [RZ96]. Complexity [Bre95b, TAH99]. Components [ATT93, BLL^{+}96, Dig90b, MOD93b, Ano95-61]. Comprehensive [Uni93a, Uni93b, FP95b, Tay95]. Compression [Del93]. Compromise [Gar97]. CompuServe [Ano94-60, Slu95, Tat95b, Wag95b, Woo95a]. Computation [MDZ92]. Computational [Bac95b, Ols96]. Computations [LD99]. Computer [Ano99e, Bac95a, BBH93, BL97, Coo93b, Che97, Com97, CD99, Com71, Fle94, FH95, GB95, HM91, HL99, Kin95a, LaQ89, Lev84, Per93, RZ96, SBN82, Sto89, Sto99, Tan88, Tan96, Uni89, UU89, Val93b, Yan88, Abb94, Ano95j, BP93, Bos94, Bri94a, Bro89, Bro96, Bur95a, Com95a, Cor96, Dec95, Dil93, Hug94, Ins96, Kes97, Kon82, Lin93a, McC88, Moh95, Mot94, Nor95, New94b, New95c, New96, PW95a, Sta88, Tri91, Tri93, War95, DJB93]. computer-assisted [Bri94a]. computer-mediated [Dec95, Dil93, DJB93]. Computer-Supported [BBH93]. computer/Internet [Sta88, Tri93]. Computers [Ano93-34, Col99, Den90, PE95, Sto99, WA98, Ano96n, MDW99, RN97]. Computing [ACM98, Ano97i, BBP89, Buy99, Col99, CS91a, Den89, Gla99, IDK98, Mei97, Mes99, Net90, Pre93, Pre95a, Pre97a, ZTLV98, Ano93t, BBST96, Bri94a, MKT98, Pas96]. comstockery [CRI95]. comuni [Pet95]. Comunicacion [Fre95]. comunidade [Rib95]. comyunithi [Rik94]. con [Gia94]. conception [Ano88a]. Concepts [Cah98, CDK94, HHS94, Heg99, Kes90, Kes93, KL98, RT98, SCJ96, Str94]. concern [Hic94]. concerned [Hum94]. concerning [JO95]. concerns [Cla99, Fol95b, WLW98]. Concise [Eag95a]. Conectate [Kro95a]. conexion [JE95]. Conference [Ano98c, Ano99b, FS95, Fra99, GMA^{+}98, Hay95, Hir97, Hir98, HL99, IEE94, IEE95, IEE97, IEE98, JG91, Lei93, LG93, MKT98, NDM98, Pla94b, TL93, Uni96b, ACM99b, Ass95b, Ano95-64, Cai94, FBGP98, KT95, LM98, Mei97, PE95, Uni96a, Bac95b]. Conferences [CKPT99, Ano96x, Car95c, HCBO99, Str92, Str93a]. Conferencing [Bac95b, MKBB95, Ben95c, HCBO99, Jac94, Moh95, Som94]. Confidence [Saw93]. Configuration [QCM94b, Rau95]. confirm [Cla99]. Conflict [FPV95, Int94a]. confluence [Ass95b]. Confman [FBGP98]. Congestion [VSN86, Dow86, FF99, KG99a]. Congestion-Control [VSN86]. Congram [MP89]. Congram-oriented [MP89]. congrès [Ano69]. Congress [Ano74, Int76a, Ano79, Ano88a, Ano94y, Ano94-32, Ano95f, Gaa76, GRS79, Int76b, Rii82, VVB85, Ano95l]. Connect [Ano97j, GRRM95, LaQ94b, Lar94b, Mad96, Wig95a, Ano96g, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95b, LR93b]. Connected [Col99, Con94, Con95b, Ger91, GO95]. Connecticut [BD95b]. Connecting [Ano94j, Ano94k, Cas93, CS94b, D’E95, Est93b, Est93c, Est93a, Fie93, Gil95a, Gil95b, MRTR95, Och95, PM94, SFL95, War99, Ano94-54, McC94a, McC94b, MK95, NYS94, Sal95b]. Connection [Ayr94, BL97, Des93a, Fow99, Har94g, Hil97, IMI^{+}95, QCM94b, AJR95, Int95d, Ano95-52, App95, Cra92, Cru95c, DS96, Fon94, Gil95e, HGKM98, Med94a, Med94b, Mor95d, Pub95, PHW^{+}93, SD96b]. connection-oriented [Cra92]. connectionless [TJ97]. Connections [Ano93p, E^{+}93, Kam93, Luk96, Nat96b, Not94b, VH94f, Wil98, Ano90a, Ano93-40, Ano94r, BU94b, Eng93b, Eng95a, Par95, Ros84, Sch96a, SC95d, Yan93, Yan94]. Connectivity [QCM94b, War99, Ano94-37, Cro96a, FB91, Grz94, McQ94b, SBM96, Sto93, Tex95, WNS^{+}94, Xie94]. connexion [And94, And95d, SS94c]. Connexions [Con87]. Conquering [Rut96]. conscience [Fre94]. conscious [AIS^{+}99]. Considerations [IEE90, Sch91a, Sch91b]. Consortium [Mei97, Ano90e, Hic94]. conspiracies [Mam96]. Constructing [AAB^{+}98, HPS95, CB96b]. Construction [LaM97, Bea95, Bow90, FD95, Joh96b]. Constructivism [Bra97]. Consultant [Ano95g, Sav94]. Consumer [RAWS95, WLW98]. contact [SLP94, SLP95]. contacts [Bar92a]. containing [MM95]. Content [Fry95, Aik95, BM98, DFS98, Smi95a]. contented [Fol95b]. Contents [Ano93-35]. Continues [Ell99]. continuity [Ass95b]. contractors [INT88, INT89b, Vic88]. Contrarian [Sto99]. Contributing [Eng93a]. Control [Haa85, IBM89c, IBM89a, IBM89b, IBM89f, IBM89d, IBM89e, IBM92b, Leo98b, VSN86, Dow86, ET91, FF99, FBGP98, IBM88, IBM90, IBM91b, IBM91a, IBM92a, Pil90, Ram95]. controlled [Sne95a, SV99]. controller [Rei92]. controls [RM96]. Convention [ACM99a, Ano97a]. Conventional [LG95]. convergence [SJZ^{+}98]. converging [Hor95]. conversations [Ros95a]. Cook [Coo92]. Cookbook [HB98a, Bou97, Sch96a]. coolest [Ano95k, Kro95c, Net94c, Woo95a]. cooperation [PH97, ST89]. Cooperative [BBH93, KW98, KSW99, Ano94-55, SJZ^{+}98]. coordinate [CKTV96]. coordinated [TFG99]. Coordinating [Ano90d, NAS90, Hun93]. Coordination [Car99b, COZ99, RV99, Ram95]. Copy [LeB93]. copyright [Ano94z, Ano95-30, Int96a]. Copyrights [Gar97]. CORBA [ZTLV98]. Core [Atk99, Keo99, Rya93, TK99, Del98]. Corner [Sch95g]. Corp [Ano99d, Ano99e]. Corporate [Cor95, HPS95, Ber96a, Cro96a, Mai96]. Corporation [Ano99e, Bre95b, Ano99e]. Correlation [Amo99]. COS [Ano96u, Ano96v]. COS/Print [Ano96u]. COS/Relay [Ano96v]. cosa [Sal95d]. COSINE [BK91]. Cosmos [Ano96s]. Cost [Bol96b, Les95, Gil95a, Gil95b, Mad96, MBB95, Mot94, Sch96a]. Cost-Effective [Les95, Mad96]. Cost-quality [Bol96b]. costs [Hen95a, Lan94, MBB95]. Could [Gar97]. Council [NAS90, Her96a, Ano90d]. countering [DD98]. countries [ACE^{+}95, PK99]. County [Ano95-33]. coup [Int92d]. Course [LG95, Roc93, Ker95, McQ94a, McQ94b, Sar94, SG93b]. courses [Cor96, Mat93]. court [Sta98b]. Courte [Sch94]. courts [Ano96q, Sta98b]. Cover [Ayr94, Bar95a, GL94]. coverage [WMS96]. CoWeb [JGKR96]. CoWeb-architecture [JGKR96]. cows [Dre92, Dre93a, Dre93b, Dre95]. CPAs [Int94d]. CPS [GBJ94]. Crackdown [Ste92]. cracked [Ano94-48]. Create [Kuo97, Ber96a, HH95b, HB95, HH98]. Creating [BU94b, Ben95a, Bry94, ER98, Fry95, Hes95, Jar95, Nof96a, Nof96b, Sky99, Ano94v, Int95g, Joh96b, Tho96b]. Creation [Pel95a]. creative [FP95b, Jon94a]. Créez [Pig95]. Crime [BK99b, Bol96a]. criminal [Luc94]. Crisis [Spa89a, YY98]. CRiSP [Ano96v]. criteria [Hin91]. critique [Ano69]. critters [MLJ95]. Cross [Ano96v]. Cross-Platform [Ano96v]. Crossing [TOLF93, TOL94a, Ano94-61]. crossroads [HOO95]. Crucial [Gar97]. cruise [Ano95-46, Kel92a, Mer92, Mer93]. Cruisin [LJ95]. Cruising [Ano91, BH94b, Fra93b, Fra94b, FAW95, FJ95, Fra93a, Fra94a, Mor95d, Sch95e]. Cruzando [TOL94b]. Cryptography [Hir98, Smi97, Sta99a, Fra99, Hir97]. CSCW [BBH93]. CSM [Ano99d]. CSM-USA [Ano99d]. CSMA [ANS92, IEE90]. CSMA/CD [ANS92, IEE90]. CTiTEK [Ano99d]. CU [Sat95]. CU-SeeMe [Sat95]. Cuba [Kle95]. Cuckoo [Sto89]. cults [Mam96]. Cultural [Cla95d, Bry95, GP95a, GP96a]. Culture [AAS^{+}97, Inf95b, Ano96e, Moo95]. Cultures [Shi96, BBN^{+}93]. CUNYVM [Lar94a]. Curious [Sne95b]. Current [Ano93e, Har94b, Tro94, FAR93a, Pri95]. curriculum [Ass94b, Ano94e, Bur95a, Gan95, Moh95, Pie95, Str94, Int93a]. curso [Fre95]. custom [ER98]. Customer [EH95]. customers [GG96b]. customization [IBM92b]. Customizing [RS97, Bla98e]. cutting [CB96a]. cutting-edge [CB96a]. CWRU [Tri91, Tri93]. CWRU-PC [Tri91, Tri93]. CWRU-PC/IP [Tri93, Tri91]. Cyber [May96, Sta98b, Tsu95, Ano95-34, Gue96]. Cybercafé [McC97]. Cyberdiritto [Pas95]. cyberespace [Gue96, Gui95]. Cyberhood [Bar95a]. CyberHound [Ano95k, Gal96].

Cybermanners [Ros94a]. Cybernetic [Bre93]. Cyberporn [Ano95l]. Cyberpunk [HM91, MW92, MSB^{+}92]. CyberSafe [Ano95m]. cyberskeptic [Cyb96]. Cybersmut [Coo96]. Cyberspace [Bar95a, Ben91, Des94, Gar97, Ros95c, Vie93, Aik95, Ale96, Bak95, Cla95a, DD98, GH93, GH94a, GH94b, GH94c, Hel94, Hen95a, Kug95, LaS95a, LaS95b]. Cyberstrategy [BBSH98]. CyberSurfer [ANT96]. Cyclades [Ano99d, Ano99d]. Cyclades-PR4000 [Ano99d]. Cypress [CNY87]. Cyrix [Ano96u]. Czech [Pla94b].

D [Fol95b, Hal99, Hen99, MBM^{+}95, Par93, Wol98]. d’adresses [Vin95a]. dai [Ban95, CRH98]. Daily [Ano93f]. dalszy [LY94a]. dan [Rah94]. Dance [Ano95n, Kan95a]. Danger [BM96a]. dangers [Ano95-65]. Danish [Gaa76]. dans [Gui95]. Dark [Woo94, Woo95a]. DARPA [Bet87, Lei85, Mil90, RAD90, SDUU85]. DARPA/NSF [Mil90]. Data [Car99a, CP94a, CFL99, GK95, Hac98, Ros89, Sta95e, TGL^{+}97, Wis90, Net94d, Ano94-55, BOB93, sC86, FP95a, Haa85, IE98, New94b, New95c, New96, Rou94, Sfe82, Sha91b, Sin99a, TJ97, TAH99, YmtEC99, Dat74]. Data-driven [IE98]. DataBase [Ano96v, ACM96, Ano93d, Ano97h, Ano99d, Ano99e, Den96, Gut99, Han96, KS97, LS98, Not93c, Vol97b, Ano94-33, Ano94-55, Bar92a, Bar92b, Hof95b, Nram94, Row96, Tom93]. Databases [Les95, Ano90c, Hea94, Bus95, Bar93, Eag93, Gal95a, Gal96, Int89a, LL99, Mot94, SL91, SLR91, SL92, ZB95, Zil95a, Zil95b, Zil95c, Zil95d]. Datacomm [Ano96u]. Datagram [DC90, Tsu89, Con89]. datasets [WoRIGSoO95]. Datenverarbeitung [AB94]. dati [Pas95]. David [MSB^{+}92]. Davis [Mei92]. Dawn [Ano94-51]. Days [Gar95a, Par94b, Ran94]. DB2 [CSJ98]. DBA [Lon94]. DC [Cim96, DLC92a, Mot94]. DCOM [RB99].

Dear [Vie93]. debate [Bla93b, BSK^{+}96]. Debian [Ano96u]. Debugging [Car98a]. Debut [Not93a]. Dec [PE95, Dig82]. DEC-Intel-Xerox [Dig82]. Decade [New81]. DECconnect [Dig90b]. December [Ano95f, HL99, IC90, USE97, McN95]. Decision [Com84, EO99, McC95]. Decision-Support [EO99]. decisions [BK99a]. decoder [Tho95b]. Dedicated [Ano95-59, JO95]. default [Hus94]. defense [GH94d, Pil90]. definition [Hal94]. Definitive [MK96b, GP96b, Jam96a, Jam96b, Nug95]. degeneration [FK95b]. degree [Rou94]. del [CS95, KMM95]. delivering [Mat95a]. delivery [Ano93z, Ben95c, McL96b, Mil95e]. della [Sal95d]. DELPHI [BR94a, Woo95a]. Deluxe [SG94a]. Demand [Gar97]. demographic [CN95b]. demographics [Ano95-57]. demonstration [SK98]. Demythologizing [BB93]. Department [Ano95w, Ano95-61, Ano96w, Pil90]. Deploying [CP99, Fra93c, Mau99a]. derechos [WLM90]. descriptive [Bow90]. Descubra [Bog94]. desde [JE95]. Design [Ano97h, Bir94, Cah98, Car98a, sC86, Com87a, CS91c, Com91a, CS91b, CS96b, CS98, CS91e, Com87b, CDK94, Dev96, Haa85, MWB95, Nie98, Per98, PT99, Pow98, Pun91, SC93, Sun96, TR94b, Yan88, BLF99, Cha95a, Com95d, Gre99, KL98, PMTJ98, RT98, SC92, Sch91a, Sch91b, Sin81, SZ99, TB90, ZDE93, Zho84, CS91d]. Designer [AFZ95, Ins96]. Designing [B^{+}95, Cos94a, Kin95a, Smy95, TS98, Wit96]. Desk [Mel94, Ano95-38, Hee93a, Hee95a, Hee95b]. Desktop [RS98]. Desperately [Hus94]. Destination [Hof95d, Hof95e, Sea90]. destiny [BLF99]. detect [Bol96a]. Detection [Amo99, IEE90, ANS92, Esc98, Sha91a]. determine [Mai95]. Developer [Ano96u, HB98b, IH98, Tow95, Int96c, Ins96]. Developers [Pet93a]. Developing [Fuo95, KW93, LG95, Rob95, ACE^{+}95, Lin93a, PK99]. Development [BBH93, DGH93, Fou98, Gut99, Kah92, SS95c, Uni96b, Wal93, Ano94-55, Ano96w, Ano97b, Bar95b, Dav95, FB91, INT88, INT89b, Lof95, MW92, Mit91, Mit92, NS94, Peh99, Pel95b, Tho96d, Uni96a, Vic88, Wes95]. developments [Har94b]. Device [Dha94]. devil [Bie94]. d’experts [LO99]. d’exploration [Soh95b, Soh96]. DFUe [RS93b]. Dial [Ano93p, Cle93a, E^{+}93, Gil93b, Gil93a, Gil94c, Not93b, Eng93b, Eng95a, Gen83, SC95c, Rey93]. Dial-Up [Ano93p, E^{+}93, Cle93a, Gil93b, Gil93a, Gil94c, Not93b, Eng93b, Eng95a, SC95c, Rey93]. Diario [Ano95o]. dias [Ran95a]. Diccionario [FPAR95]. Dictionary [Fah94, Godxx, RS91, RS93a, Bie94, Cru95b, FP94, New94b, New95c, New96, PW95a, Ano94g]. Diego [Ano94-53, Ano97g, DLC92b]. Dienste [SBGK94]. Diesel [PRY99]. differentiated [FK99]. differentiated-services [FK99]. Diffusion [DS97, GPRR94, Lin93a, Pre97a, PBF^{+}98, RRS98]. digest [Ten92]. Digital [Bla97b, BW97b, FMS98, Gar97, IE98, KP95, LL95b, LL96, Owe93, PE99, RD95, TE99, Ano94-55, Com95d, Nor95, Tho95b, Lit93, SCS99]. digital-audio [Tho95b]. digital-audio/video-interactive-decoder [Tho95b]. Dikon [Kon91]. d’images [How95c]. Dimension [Ano96u]. Dimensions [BW97a, GP95a, GP96a, Mil95f]. Dinos [Rya95i]. d’introduction [CD94]. Direct [Har94g, Gal94, HW95, MK95, RAWS95, SC95d]. directed [Rus94]. Directing [BW97a]. directions [Man96]. Director [Bro89]. Directories [New93, New94a]. Directorio [WLM90]. directors [Mai95]. Directory [Ano94l, Bra93b, ED92, End94, FA89, FA93, FA94, KKO94, LaQ89, Mec94, New93, New94a, Not94a, Rou94, TT95a, TT95b, Leg94a, Hea95, Avi96, Int96k, Ano96r, Bla93a, Bra94c, BL96, Ela94, EHPT95, Foc95, FMB95, Int87, Hum89, HS94b, HS95a, HK91a, HK91b, Har94c, Hel94, Int94f, Int94g, LJN94, Lev95, Max95d, Mil94a, Mil95a, Mil95b, Mil96a, NN94, NC95a, SSC96, Ste95b, Str93a, Wal95a, WS80, WMRS87, WR89, WLM90]. disabilities [SSI^{+}98]. disabled [PH97]. Discarding [CX97]. discourse [Gla94]. discourses [Bau95]. Discovering [Ano95p, SW93, Wen95]. discovers [Mou96]. Discovery [HPS95, ODL93, Sch93b, BDS93, BDMS94, Deu92, GGMT99, KW98, Sch90, Sch91c, Sch92b]. Discusses [Gar97]. Discussion [Ano99c, Ben95a, Mag94, KKO94, Ano94-41, Bry95, EGC^{+}94, HOO95]. Disk [Ano96s, Inf95c, Joh96b, JE96]. Disorder [Ste92]. Dispatcher [HGKM98]. Disseminating [GK95]. dissemination [Wes95]. distance [Ano95-66, Ben95c]. Distributed [AR98, Ano94-53, BM98, BN96, CDK94, HS98, KvRvST92, KvRvST93, KP93, MWB95, Ros89, SBN84, Sin99b, Slo94, TB90, ZM93, BAG94, CKTV96, HK91a, Hof95b, LM98, LL99, MM95, MO92, dAMW97, Smi95a, Spe87, VJ99]. Distributing [Hen99, Wit96]. Distribution [BAG94, Vie93, Ass94a, All94, Hen95a]. District [Tee94]. Diverse [Bra93a, Ste91a]. division [Wes95, Ano96w]. DIX [Dig82]. dla [LY94a]. DNS [AL92, Alb93, AL94, BAG94, Tan93]. do [Ano95-45, Che97, Max95e]. Docs [Hub96]. Document [Ano93m, RV99]. Document-Based [RV99]. Documentacion [Ubi95]. documentation [Ano93z, CCN94]. Documents [Hes95, BDKK96, HH95b, HB95, Hub96, MO92, RR94b, Str92, Str93a, Wit96]. DOD [Pos80, CC83, She86]. Doelen [VVB85].

Does [SW97, KKM^{+}98, RAWS95]. Doing [Boc96, CC95, Cro94a, Cro95, HH95a, RL95b, DKF^{+}95, Har96b, HB96, Rot96, Sme96]. dollars [DFS98]. Domain [Moc94, Neo94b, Ste96b, Int93c, Ste99b, Zho84]. Domestico [JE95]. Don’t [CES^{+}99, SW97, Sto99]. donuts [DKF^{+}95]. d’or [GS95b]. DOS/Windows [Int93d]. double [RR95a, RR95b]. double-edged [RR95a, RR95b]. down [Coo95a]. downloading [DG95]. Dr [Ran96a]. dragon [Ree93]. Dragons [HT95]. dream [BK96a, GCS96]. Dreams [Ste96a]. drink [GS96b]. Drinking [WoRIGSoO95]. Drive [Gar95a]. driven [IE98]. Driver [Ano94-35, Bra94a, GFF^{+}95, LSBM95]. Drivers [Dha94]. Driving [Neu94]. Drug [Dev96]. DSI [Son92]. Dumme [LB94b]. Dummies [AH95a, LB93, CC96b, LB94a, LY94b, LY94c, LBY95, LY95a, LB95a, LBY96, Lev96, LY96, Sei94c, Sei94b, Sei94a, SP95a, Syd96, You94]. dummy [EKG94a, EKG94b, GKC93a, GH93, GKC93b, GH94a, GH94b, KG94, PE94, PG94]. Dynamic [Leo98b, LD99, Sav96, Tsu89, Atk99, FHU99, LC98, NM98, Tho96b, YmtEC99]. dynamics [Pax99].

E-Business [Shu99]. e-mail [BL95a, Bau93, BW94a, BWC95a, Ehn94, Int95b, Kw95, Ban94, RAWS95, SB95, BAS95, AH94, Bac95a, BL95d, God94a, McF95, VH94d, Als95, Ano95-68, DC94, Gar94, Inn96, JEK92, Lev96, Pos95, QCM94a, Shi94, SG93b, TR94a, War95, Wri95a, Ken90, Mul99]. E-meil [cYlC95]. early [Bla95d, Ehn94, Eme95b, GBJ94, Rot95b, Woo95b]. EARN [PS92]. EARN/INTERNET [PS92]. earning [LR95c]. Earth [KL94, Rit93, Koh97, Ano94-55]. Earthquake [HM97]. ease [Ano93c]. easiest [DG95, Int95q]. easily [Too95].

East [Ass94a, Gal94, GP95a, GP96a, Mar93]. Eastern [WMRS87]. Easy [Fra93b, Fra94b, FAW95, FJ95, Lic94a, Lic94b, Mil94c, Mil95d, Min95, MNT94, Tus93, Bar95b, Fra93a, Fra94a, Ken95g, LE96, Lic95a, Lic95b, PA94, Slu95, Wag95b, Wil96a, Wol96]. EcoLinking [Rit92]. Economic [GS99, MMV94, AT94]. economics [MMV93, MKM94, Pel95b, Roo97]. economists [Ano93-30]. economy [Bro91]. Ed [Ano92i, Ano95-41, LB95a]. EDGAR [GK95, KG98, Ano93g]. edge [CB96a, WM95]. edged [RR95a, RR95b]. EDI [Lee98, Wri95a]. Edinburgh [FS95]. editing [SJZ^{+}98]. Edition [Ano99e, BGS94, BC95d, HK95b, SG94a, TPBH94, AH95a, BW93, BGS95, Bra94b, CB94b, Dar93b, Edd96, GBS95, GS95a, Gum93a, Jam96b, LY96, Lic94a, Lic94b, TA96]. Editor [Ano96v, CB96b, LJN94]. Education [AAS^{+}97, ACE^{+}95, And93, Ano95q, Ell94, Gou90, HWV96, Kah92, MBDR91, Mei92, Par90, SS95c, SW97, AHH88, Res92, Ano94-35, Int95m, Ano95-66, Bra94a, Ben95c, Bur95b, Ela94, EHPT95, EGC^{+}94, Har96b, LSBM95, Mai95, McL96b, O’H93, Wag95a, Wal94a, Wil94b, Wol98, Ano95w, PE95]. Educational [Edu95, Har94d, Pla94a, Fuo95, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95b, Har94e, Har95a, KK96a, KJTS96, LS95, LS96, Moh95, Zha94]. educativa [Fre95]. Educator [BH93, Con95a, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95c, Gia95d, Gia95a, GMDS96, PDP96, Ano95-54, Bla95d, Dav95, Gia95b, GVW94, GVW95, Vir97, Wol96]. Educators [Ano93s, Ano94-47, Ano96j, Gla94, Jos95a, Jos95b, Mac95a, PS95b, Rot95b, Rou94, RH97, WMS96, Yas95]. Edutainment [Ras94]. EFF [GH94c]. effect [Gal94]. Effective [Ham99, Les95, RCL96, Mad96, Str94]. Effectively [LP95, QCM94a, RK95c]. effects [Rus94]. efficiency [FL95a]. Efficient [TGL^{+}97]. EG1A [Ano97h]. Egg [Sto89]. eight [MW94c]. eighteenth [GV90a]. eighth [Ano90d, NAS90]. Einblicke [BGGS95]. Eisenhower [GBJ94]. ekonomiku [Spi92]. electric [Str92, Str93a]. Electrical [Hel91a]. electro [Ano95e]. electro-industrial [Ano95e]. Electronic [A^{+}98, Ele91, Ano93j, Ano94m, Bac95b, Ben95a, Buc95a, Car98b, Chu93, Cro94b, ED92, End94, FA89, FA93, FA94, GMB93, HWV96, Mag94, Joh96b, Kan94a, KKO94, LH93b, LC93, LC96, Mee95, Nil93, Pyl96, RWD95, Rei93, Res97, Rhe94a, Rhe94b, RBM94a, RBM94b, SM93, VJ98, Wie93, WC95d, Zor95, AH95b, Ano91, Ano94-41, Ano95-52, Ano96e, Bak98, Bos94, Cro94a, Cro95, Din94, Fel93, Fel94b, FK95a, Fol95a, GG96b, HOO95, Har94b, Hf96, Hel94, Hen95a, Hor95, Soc95a, Ken95g, KW98, LH93a, LM98, Lar94a, Mar96a, Mot94, Nat93a, Pel95b, Per95a, PS92, Pod95, RK95a, Ros93, Rus94, Ste92, Str92, Str93a, TL96, Vic89, Wat93, Wis90, Wri95a]. Electronics [Net93]. elektronische [Ble95]. elementary [Int93a]. Elements [AH94, Pad85, Hor95]. elettronico [Gio95]. Elmhurst [Ano94m]. elsewhere [Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b]. Email [Hf96, SG98, WC95b, JO95, Kar96, Win96a, LP95]. Emergence [Col99, Bro91, Mas97]. emergencia [Rib95]. Emerging [Bla97a, BL95d, Mue99b, Val93b, Bak98, BBK^{+}98, RA96b]. Emotional [JP94]. emperor [Moo95]. emphasis [Wil94b]. Empirical [CES^{+}99]. Employee [Uni95a]. Employment [Ril95]. empowerment [AJR95]. Empress [Ano99d]. Emulation [Ano96t, HKR91]. Emulus [Ano96t]. Enabling [BK99a, LF95]. Encapsulation [SCT81]. Encryption [Bac95a, Gar97, LS89, Ken95g]. Encyclopedia [Abo93]. End [Gol92, Pax97, Pax99, FK99, FF99, KIP99, Mor94, Not93b, San94a, San92, SP94b, LS89]. End-to-end [Gol92, Pax97, Pax99, FK99, FF99, San92, LS89]. end-users [Mor94]. Energy [Bro89, SS95b, She95b]. Engineering [Ano96u, CTG96, DLC92b, DLC92a, GV90a, GV90b, Gro90b, Gro90a, Pow93, Pow98, Ano94r, Gil95a, Gil95b, Kes97, TW95]. engineers [He98, Tho95a, Tho96a]. Engines [DH97]. English [BGGS95, Pfa96a, SBGK94, Int95g, Pfa94a, War95]. Enhanced [Ken93, Kol96, Asa91, Wis90]. Enhancing [SSI^{+}98]. enjeux [Kap96]. Enterprise [CP99, Har99a, KJGH98, Mau99a, Nej94, RS98, Sky99, EI 95, Bro93]. Enterprising [Sav96]. enters [Opp95]. Entertainment [RBM94a, RBM94b]. entrepreneur [LR95c]. entreprise [AL95]. Entwicklung [Jun95]. Environment [Bre93, KSM98, Yan88, Ben95c, BP91, Deu92, Dow86, Eas94, Ger91, McL96b, Mil95e, PH97, PJ95, QRV95, San91, San94b, She85, SC88, Tho95b, Wil94b]. Environmental [BEM96, FB95, Per95b, Rit92, Sch95f, She77]. Environments [LS89, RV99, MW92, SZ99]. equal [Str95]. era [LL95b, LL96]. ERIC [Ten92]. erotique [CD95]. Erotisme [Cau95]. ERS [Edu95]. Espana [Par96b]. espanol [MPF95]. especially [NCS93, Nat93b]. Espionage [Sto89]. Essays [Pad85, MW92]. Essence [Rea98]. Essential [Ano95s, BLB99, Dod99, ELB99, Gil93b, Gil93a, Gil94a, Gil94c, Man95a, Ber95a, Gil94b, Gil95c, Gil95e, IV95, Str94]. Essentials [Arm94b, Ibe97, Keo99, TR94a, TR94b, TR95, Arm94a, Cla95b]. establish [Ano95-61]. establishment [EST93d]. Estimating [LCP91]. etc [Ano95-68]. Ethel [GK96]. Ethernet [HL93, BMK88, Dig82, Haa85, Har99b, Hel98, Hel99, KKC98, Maz90, Sei98, SL95]. Etherphone [TS88]. Ethnic [Int94a]. ethnomusicology [Ehn94]. Europe [AP95a, SEDK94, WMRS87, Ano93z]. European [LDH95, Pla94b, SEDK94, Ewe95, Gar97, IC94, Mar93]. Evaluating [Rey93, Cha95a, RK95c]. Evaluation [Gha92, Mit92, MMLP97, McC94b, PH97, PM94]. events [Ano93-36, Kac96]. Every [Pit95a, Pit95b, Pit95c]. Everybody [Gaf94, Rub94]. Everyone [Fri94, Mau99b, Rit92, Wig95b, CS94a, JD95a, JD95b, Los95]. Everything [Com97, Sny96, Thi93, Com95a, GH93, GH94a, GH94b, GH94c, HH98, Sch95e]. eVO [Ano96e]. Evolution [Kah94b, Int91]. Evolving [BL95d]. Examining [RG93]. example [Hal94, Shu93]. Examples [SBN82, Ack95b]. excellence [Col95]. Exchange [Gon99, Sto93, Net95d, Bir93, Int85, Kol96, Nem96]. Exchanger [She95b]. executive [BL95b]. Executor [Ano96u]. exercises [Ack95b]. Exhibition [Int76a, Int76b, Ano95-64]. Existing [SS99, Wes97]. exits [RHS95]. Exon [CRI95]. Expanding [Car99b, Dee95, Ger91, Ano95-63, Mal95b]. expansion [SV99]. expedition [HU95]. Experience [Mei92, SBN84, CG95]. Experiences [DH97, SKB96]. Experiment [MKBB95]. Experimental [Mog89, San92, Yan88, Mog88, Sch91a, Sch91b, Som94]. experiments [Sha91a]. Expert [Leo98c]. experts [Dra95]. Explained [Har99b, Mil97, Pad95]. exploit [Hof96a]. Exploiting [ST89, Cim96]. Exploration [Bri94a, Gil93b, Gil93a, Gil94c, EGC^{+}94, FP95a, Rus94]. explorations [Hal99]. exploratory [Mai95]. Explore [Cob96, Kro95c, Net94c, PE99, FTP95a]. Explorer [ED94, Sam94, Cob96, Cat96, FTP95b, Ken95h, Lov99, Roa95b]. explorer/MSN [Cob96]. Exploring [Ano94n, Ble93, BBST96, Cro94b, D. 93, De 94, ECC94, EE95c, FTP95a, FTP95b, GNU95a, GNU95b, Har94a, Hem93, HM96, Jon90, Lan93, LNH^{+}93, Mal92, MG96, MB93, PGS93, PGS94, San95, SLB94, Tha95, Aik95, BC95b, Hal95b, Mar96a, Pfa95b, Pfa95c, Pfa95d, Pfa95e, Pfa96c, Wat93]. Explosion [KS97]. Exposed [MSK99]. Exposition [Ano97a]. Express [Gun94, Rae95a, Rae95b, Rae95c, Rae95d, Rae95e]. expression [FP95b]. Extended [DC90, GH94c]. Extending [COZ99, Pao96]. extensions [HK91a, Mol93b]. external [Bro91]. Extraction [Sim93]. Extranets [Bay99, BBSH98]. extremists [Hof96a].

fėvrier [Min94b]. F1 [Ass94b, Ano94e]. faccio [Sal95d]. Faces [KS97]. facil [Ken95f]. facilitated [SB94]. facilitating [JO95]. facilities [Kes90, Kes93]. facility [Spe87]. facsimile [Pun91]. Fact [Ano94o]. Factors [Dav95]. facts [Con94, Con95b]. faculty [Ale95, Bar95b, Byr95, Sta94b, Zha94]. Failed [Gla99]. FairCom [Ano96v]. Fall [Cim96, Mas97]. Famous [God94a]. FAQ [Int95b]. FAQs [MMV94, Sav94, YL95]. Farmer [JFJ96]. Farmers [Pet93a]. Farming [Hac98]. Fast [Bak93a, MD91, SV99, DeR96]. fatica [LB94d]. Faulkner [CDEF94]. Faxes [Bak93c, BA95]. faxing [Win96a]. FC [Fra99, Hir97, Hir98]. FDA [Tom93]. FDDI [Jai94, Ros86, Ros89]. Fears [Gar97]. Feasibility [SI 93, Nic95]. features [ZDE93]. featuring [Nat96a]. February [Ano94-53, Fra99, Gro90b, Hir97, Hir98, JG91]. FEC [CX97]. Federal [LJN94, Par94a, Ano95-27, Can94, Lev95, LN95, Max95a, Max95b, NN94, Zaz95, Gou90]. fee [Mac95b]. fee-based [Mac95b]. Feminism [Kug95]. Ferret [Har94d, Har94e, Har95a]. Fiber [Jai94, Ros89]. Fibre [Tho99]. Field [KSM^{+}96, Nel95, Str94, Vie93]. Fifteenth [ACM96, GV90b]. Fifth [Int76a, Int76b, Sav96, DLC92a, Moh95, Yas95]. File [BH93, Del93, Gar94, PR91b, RP93, Saw93, Sim93, VH94b, Zha96, Ano95-35, EHS92, Hal93, IBM89a, Int92a, Kas93, Kie95, Moj93, PR91a]. Files [RP93, Rob94, VH94b, Ano95-61, Ehn94, PR91a, PR91b]. Filling [JGKR96]. Film [DH96, Gar97, WG94b]. Filtering [MMLP97, Mul99]. Final [BU94a, MBZ94, MBB95, McN95]. Finance [Bro89, Uni89, UU89]. finances [New95b]. Financial [Ano96f, Fra99, GK95, Hir97, Hir98, Gor95a, KK96a, Lee98]. Find [Mui95, SRR97, Con95a, LJN94, Lev95, NN94]. finder [AP96]. Finding [Ano94p, Ano95t, Gil94a, Gil94b, Gil96, GG95a, Har94d, LeB93, Max95e, New95a, Pfa95f, DS96, Fel94b, FK95a, GRRM95, Har94e, Har95a, LS95, LS96, RK95c]. findings [LT92]. FindLaw [Ano95u]. FINDsomeone.com [Gra98]. FINFS [Moj93]. Finland [Ano96x]. firehose [WoRIGSoO95]. Firehunter [Ano99e]. Firewall [Ano96t, Esc98, GRLJ94]. Firewalls [BCRL95, CB94c, Gon98, Opp97, SH95a, CZ95, INP95, SS95e, WC94, Ano94f]. FireWire [And99]. Firing [BSK^{+}96]. firm [Ano96o]. First [New81, Ben91, CD92, Hir97, Ano95l, Ano95v, Ano95-46, Bac95b]. Fish [LaM97]. fitness [NC95a, NC95b, Rye95]. five [Int81, Mai95]. Fixed [Bla98f]. FL [GMA^{+}98]. Flame [FPV95]. Flaw [Gar97]. flexible [Moj93, ZDE93]. FLIP [KvRvST92, KvRvST93]. Florida [Gro90b, Gal94, Mac95a, SH95b]. Flow [Dow86]. flu [Ano93y]. fly [LeG95b]. Focus [Ell99]. Fodor [Goo96]. folk [Bry95]. folks [Bar95b]. follow [Bar95b]. follow-up [Bar95b]. food [GS96b]. For-profit [Bal93a]. Force [DLC92b, DLC92a, GV90a, GV90b, Gro90b, Gro90a, IC90, ILSB94, Ano89b]. Forest [Bri95c]. Formal [Roc93, Sev95]. Formats [Sim93, Ano95-35, Kie95, Zha96]. Formed [Gar97]. formerly [Cim96, Mas97]. forms [HH95b, JGKR96, Kol96, Nof96a, YmtEC99].

Forrestal [Lin93c]. forskning [Dam93]. fortnightly [Pri95]. fortuna [CS95]. fortune [CS94a]. Forum [CES^{+}99, Can94, GRLJ94, Soc95a]. Forums [Hes95, Car95c, HB95]. Forward [Bak93a, TMF97, Xie94]. forwarding [EST93d]. Found [Bel93, Eng96a, Kug95]. Foundation [Bla95a, Bla99d]. foundations [Wil99]. four [Coo93b, HT95]. Fournisseurs [Cle93b]. Fourth [Ano74, Ano95l, AHH88, Yas95, Int94b]. fracasos [Ano95b]. Fragmentation [Tsu89]. Frame [MD91, Smi93a, Sta95a]. FrameMaker [RR94b]. Frames [Fla91]. Framework [Bla93c, Pet86b, PBF^{+}98]. frameworks [Gre99]. France [KW98, Kap96, KW94]. Francisco [Hay94c, TL93, Ano94-58, Lei93]. Free [EV96, Lof95, Vin94a, Vin94b, Vin96, TT95b, Mac95b, Mot94, Vin95b, Win96b]. Free-nets [Lof95]. Freebies [Wol98]. Freedom [CES^{+}99, Her96b, Ano95-34]. frequency [Wes95]. frequently [Int95b, Rot95b, Sav93]. friendly [Kro95c, Net94c, Ano90f]. friends [CDEF94]. Frontier [BL95d, Chu93, HM91, Mee95, Rhe94a, Rhe94b, Cro96b, MBS95, Ste92]. Frostburg [Owe93]. FSU [Owe93]. ftp [BH93, Fel93, Ano92g, Bos94, BW94a, BWC95a, FTP95a, FTP95b, Fel94b, FK95a, LPJ^{+}94, Nof95c, Rob94, SG93a, VH94b]. Full [Hub93, Bre94]. full-text [Bre94]. Fun [Mor95c, Hof95a]. functionalities [KJTS96]. functionality [JGKR96]. Functions [Car98b]. Fundamentals [Ber95b, Mau99b, Sha95b, Inf95c, Kur96b]. Funding [Hal94, Con94, Con95b]. funniest [May96]. Future [AAS^{+}97, Far97, GGA^{+}97, Gol93, KS97, LT93, LCC^{+}97, Man96, MSD90, Sac97, Bre95b, Bro95a, CCC^{+}91, Dra95, FMS98, GAM^{+}97, HOO95, Har92, HRB92, Kle94, KG99b, LT92, MDB^{+}94]. Futurekids [HU95]. Fuzzy [Leo98c]. FX [Ano97h]. FYI [MM91, MM92].

G0 [Ano99d]. G0-Joe [Ano99d]. gaido [KM95, LGRS93]. gain [Wam95, Ano94-54, McC94a, McC94b, NYS94]. gains [BG98]. Gale [Gal95a, ZB95]. gallery [DOB95, SY95]. gambling [Neu98b]. Game [Tan93, Bel95]. gamer [Pan96]. games [Bel95, Car95c, Gre96, Tav95]. Gap [Int76a, Int76b, Pot93, Pot94a, Pot94b]. Garbage [AR98]. gardeners [Pet93a]. Gas [Uni93a, Uni93b, Cro96c, SR95b]. gatekeeping [All94]. Gates [Mil94d]. Gateway [VSN86, BAG94, Gen83, Haa85, Mei82, PS92, RG91, Ros84, Sfe82, US 92, US 93, Zu 82]. Gateways [Tsu89, BP87, Fol95a, Gas96]. Gathering [WYLO98, SP94a]. Gebührenfreies [BGGS95]. Gen [Hof95b]. Gen-Web [Hof95b]. gene [SS95d]. genealogist [Arc95]. genealogy [Haw96a, Hof95b, oS94]. General [Ano95-61, Ano96w]. Generating [CO99]. Generation [AK97, Bla99e, Sav96, Ano93x, BPR88, BM96b, DFG^{+}98, dAMW97, Sal95d, Tho96c]. Generic [Ano96u]. Geneva [Cai94]. genocide [Int85]. geographic [FB91]. geographical [Fra94c]. geography [Pie95]. Geospatial [Ano99b]. Germany [LM98]. Gestion [RJ95]. Get [Car95b, MPG94b, MPG94a, Whi95, Win96b, BK96a, Con94, Con95b, DS96, Ent95, Ger91, Hf96, Ken95b, Ric95a, SD96b]. Getting [AL96, ACE^{+}95, Arm94b, Bax95, GA95, LD95, MKNW92, McK94b, McQ94a, PA94, SMLD95, Web99, Wri95b, Arn96, Ehn94, GO95, Ive95, MK92, M^{+}93, MKNW94, McK95, MLJ95, SC95b]. Ghana [Int94b]. Ghosh [HH95a]. GI [LM98]. GIF [AIS^{+}99]. Gigabit [Hel98, KKC98, Par94c, Sei98, Hel99]. giornale [Gio95]. Gist [Mil94f]. GIX [Sto93]. gizmos [DHS95]. Glance [Fel94a, Fel93, Fel94b, FK95a]. Glasgow [Ano88a]. Global [AUC93, Ame95c, Ame95d, Ano99d, BBH93, CWJT99, Cro96a, Dyr93, GPRR94, HA94, Kam93, LR93c, LaQ94b, LH93b, MKBB95, Nil93, Pre97a, SQ93, Sto93, Uni96b, NAS93, WC95d, Ano94m, Hea94, Ano94-37, Bus95, Bri95d, Bri95e, Cam94b, FB91, GH93, GH94a, GH94b, GH94c, Ger91, J3 95, Koh97, LR93a, LR93b, LaQ94a, LDH95, LH93a, PHW^{+}93, QCM94a, RRS98, ST89, Sch91a, Sch91b, Sch91c, Shu93, Uni96a, WNS^{+}94, WoRIGSoO95, Zil95a, Zil95b, Zil95c, Zil95d, PM94, Sto93]. Globalizing [AAS^{+}97]. Glosario [MPF95]. GlOSS [GGMT99]. Glossary [Ano93a, Ano93b, Ano93-39, Mac93b, MP93]. glue [Bro93]. GO [Ano99d, Mil99a, Bal93b, SG93b]. GO-Between [Ano99d]. GO-Global [Ano99d]. Goals [Sch91b, Sch91a, Sto95c]. God [Kel96]. goes [Win93]. Going [Uni96b, Ano95-48, Uni96a]. Gold [Lit93, MZ93, US 92, US 93]. golden [HS94b, HS95a, NC95a]. Goldmine [Ano97g]. Good [Gar95b, Gen95, Hal94]. GoodNet [Gol93]. Gopher [Ano95w, CCN94, Fra95, Lin93b, Mor94, Nof95d, Ano93-41, Ano94-56, Ano95x, BW94a, BWC95a, Leg93, Sne95a, Ano92h, Ano93-42, Fel94b, FK95a, Gil94a, Gil94b, LPJ^{+}94]. Gopher/Veronica [Fel94b]. gopherin [SG93b]. Gophers [MLJ95, Bos94, Bab94, Har93a]. Gopherspace [Mar96a]. governed [Cla95a]. governing [Ram95]. Government [AT94, Bro89, Kah94b, Nef94, Ano95-27, Can94, Gum93b, Gum93c, Gum94, LJN94, Lin93a, Max95a, Max95b, Max95e, NN94, New95b, Pub95, Par94a, Rya94, Rya95a, Rya95b, Rya95c, Rya95d, Rya95e, Rya96a, TL96, Zaz95]. governmental [Ste91a]. Governor [Ano95-61, Ano96w]. Governs [FRG97]. GPM [D^{+}86].

GPM/INTERNET [D^{+}86]. Grace [BBH93]. grade [DG97, Moh95, RK96, Yas95]. grades [Mai95]. gradient [TAH99]. grant [Ano95w, Mle95]. grantees [INT88, INT89b, Vic88]. Grapevine [SBN84]. graph [SW93, ZCD97]. graph-based [ZCD97]. Graphical [Lic94a, Lic94b, Tes96, Ano93-45, BDKK96, Gil95e, Kas93, Lic95a, Lic95b, Wen94, Wag95b]. Graphics [FMS98, IE98, DOB95, MBM^{+}95, Par93, Rim96]. GraphOn [Ano99d, Ano99e]. grassroots [RMU96]. gratifications [But95a, Rya95i]. Great [Hes95, HH95b, HB95, Tur95d, Wat93]. Great-looking [Hes95, HH95b, HB95]. Greenbelt [LG93]. Grif [Pao96]. Group [And95a, Ano98b, JG91, LG93, Ano92f, Bry95, Hen94]. groupware [Hor95]. grow [Eme95c]. growing [AH95b, MW92, SP94a]. Grown [Gar97]. Growth [HT95, SBN84, Coo93b, Lot92]. GSM [GW99]. GSQL [MWB95]. guaranteed [AHS90]. guaranteed-performance [AHS90]. guarantees [Low90a, Low90b, Low91]. Guerres [Gui95]. Guerrilla [LR95b, LR95c, CS94a]. guerrillero [CS95]. GUI [Ano96v, Ano93j]. Guia [CS95, LR94]. guida [Pas95, Pet95]. Guide [Ano92i, Ano93i, Ano93p, Ano93r, Ano93s, Ano95x, Ano95-55, Ano97j, Ari93, Bac95a, BH94a, Bla98a, Blo94, Bra93a, Cal94b, Cal94a, CC95, Cas94, CSJ98, Dig90a, Dig93, Der94, Des93b, Dod99, DC94, Dup95, EAR94, Eag95a, EE99, E^{+}93, Est93b, Fla90, Fly95, Fra93b, Fra94b, FAW95, FJ95, Gaf94, GKM^{+}99, Gal95b, Gar94, Gil93b, Gil93a, Gil94a, Gil94c, Hew88, Har95b, Har95c, HB98b, HA94, Mag94, Kan94a, Keh93, Keh94, Ken94b, Ken94a, Kic94, Kin95b, Kro92, Kro93, Kro94, Leg94b, LR93c, LS93a, LC93, Lin93b, Luc94, Mah95, Man95a, Mat93, McF95, MB93, Mes99, MD99, Mil91a, Mil95c, MP92, MD91, MK96b, NSF91, Neu94, New93, Nor92, Per93, Pfa94b, Rau95, RdG95, Rit92, RBM94a, Rya95a, Rya95b, Rya95c, Rya95d, Rya95e, SS94b, Seg95, SL95, Smi93d, SS95c]. Guide [Soh96, Sta93b, Sta93c, Ste98a, Swi96, Tow95, Twe94, VM95, War99, Web99, Wig95b, Wil93, WS94, Wol98, WG94b, X/O88, Zel99, Zim95a, AJR95, ANT96, AFZ95, Ale96, All95, And92, And94, And95d, Ano89a, Ano91, Ano92c, Ano93u, Ano94i, Ano94x, Ano94-27, Ano94-37, Ano94-41, Net94b, Ano95d, Ano95k, Ano95m, Ano95j, Ano95s, Ano95q, Ano95-27, Ano95-39, Ano95-48, Ano95-49, Ano95-58, Cyb96, Ano96d, Ano96e, Ano96j, Ano96q, Med96, Arc95, AP95b, Bri94b, Bac96, Bak95, BR94a, Bar93, BH95, BH96, Bea94, BC95c, Bla95d, Blo95, BD88, Bra95, BU95, Bre95a, BEM96, Bro94a, BS96, Con95a, Car96, Cal90, Cam95c, CS94a, CD94, Chi95, Cla95c, Cla96, CT95a, CT95b, CT96a, CN95a, Com95b, Dav94, DeR96, Den96, DOO95, Din94, DHM^{+}96, DG94, DG95]. guide [Dre92, Dre93a, Dre93b, Dre94, Dre95, EKG94a, EKG94b, Ebi95, Ehn94, Eis95, Eng93b, Eng95a, Est93c, Eva95, Fra93a, Fra94a, FP95b, oIT94, Gal95a, Gal96, Gac96, GKC93a, GH93, GKC93b, GH94a, GH94b, GH94c, GR94, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95b, GK96, Gil94b, Gil95c, Gil95e, Gil96, Glo95, GG96a, GVW94, GVW95, Gre95b, Gre96, Gri95, GO95, GP96b, HTW94, Han95, Har96b, HE95b, HV95, Hel94, HKR91, HA95, Hob96, HH96, How95c, Hug94, HB96, Inf93, Inf95a, Inf96, IBM89f, IBM92a, Int95g, Int95q, Ive95, JW95, JFJ96, Jam96a, Jam96b, JEK92, JE96, Jos95a, Jos95b, JP94, JM96b, JT96, JD95a, JD95b, Kan95b, Kan95c, KG94, Keh92, Keh96, Ken95a, Ken94c, Ken94d, Ken95d, Ken95e, Ker98, KH95, KM96b, KHJG91, KHJ91, KB92, KHJ92]. guide [Koc93, KL95, KL94, Kov95, Kra94, Kro87, Kro89, KCG95, Kro95b, KK96b, LR93a, LaQ94a, LDH95, LaS95b, Lar94a, Lar94b, LoI94, LS93b, LS95, LS96, LN94, LR95b, LR95c, LC96, Mac95b, Mai96, Mal95c, Mar96a, Mar96b, Mil96c, Mor94, MBS95, Nat93a, NCS93, Nat93b, NSF89, NSF92, NYS92, Nel95, Nem96, Neo95, Net94a, New94a, Noo91a, Noo92a, Noo92b, Noo94, NM94, NM98, Nug95, Ols95, Oti95, Ove94, ODO95, Pee96a, PE94, Pre97b, Pub95, Par95, PM95b, PN91, Per91c, Per91a, Per91b, Per92d, Per92e, Per92b, Per92c, Per95b, PU94, Pie95, Pit95a, Pit95b, Pit95c, PMTJ98, PG94, PS95b, Ran96a, Ree93, RT94b, RT94a, Ric95a, Ril95, RRO96, Roa95a, Roa95b, RK95b, RO94a, RO94b, Rom96, Ros96, RK95c, RMU96, RA96b, Rya94, Rya96a, Sua94, SUR90, Cen93, SRI82a, Net83, SUR93]. guide [SUR92, SBM96, SC94b, Sch96a, SS95b, Sch95f, She95a, She77, Sil95, Siy94, Slu95, Sme96, Smi93b, Smi93c, Smi94a, Soh94, Soh95b, Sta95d, Str92, Str93a, Str94, Sua95, Tay95, Tee94, Tho96b, Tho95c, TB93, Til95, TG95, TPH96, Tur95d, Uni93c, Vir97, Van94, Ano94-30, Wis96, Wag95b, Wat93, Wei95a, Whi95, WNS^{+}94, Wil96a, Win96b, Wol96, Woo95a, Woo96, YM96, ZB95, Zaz95, Zef96, Zha96, RBM94b, Ani93a, Ano93-31]. Guidebook [CTN95, Haw96b]. guided [Cru95a, FP95a]. Guidelines [Cer91, PCF91, Ano95-61, HC94]. Guides [Bar94, Tus93, VH94a, VH94b, VH94c, VH94d, VH94e, VH94f, RJV94, RJV95e, RJV95a, RJV95f, RJV95b, RJV95c, RJV95d, VH94g]. Guiding [Pet93a]. Guidon [Ano93j]. Gulf [BG98]. Gurus [B^{+}94].

ha-ishi [PEH95]. ha-Leksikon [PEH95]. ha-Yisreeli [Mal95a]. Habermasian [Fan95]. hacer [CS95]. Hacker [CP94a, CB94c, RS91, RS93a, Ste92]. Hackers [Ano93k, CES^{+}99, HM91, Lev84]. Hacking [MSK99]. hadash [PEH95]. hagi [cYlC95]. Hahn [Hah96a]. Haiti [Peh99].

Hamamatsu [IDK98]. Hamburg [LM98]. hand [Bro91]. Handbook [Bac95a, CB94a, CB94b, CP94a, DK94a, DK94b, DKF94, Fly95, Fra95, Fur99, Har94f, Har95c, He98, Jai94, Mil89, Mil93a, MNT94, RS93b, TOLF93, Vac98, Buy96, CB95c, Cro96b, GLSW89, Grz94, Jan95, Kro95c, Lon94, LR93e, Mil93c, MRJ96, MNT93, Nat93a, Net94c, PWK94, RBK^{+}94, RMU96, SDUU85, Sta95b, TOL94a, Uta94, Vir97, Wen94]. Handbuch [RS93b, SK95]. Handler [Caw95]. Handouts [Ass94a, Wor94, AUA94]. Hands [HA94, Rau95, SS94b, SS94a, Sal95a, Sal96, Zna94, Ano94v, Ano94w, Ill91, Owe93, PT95a, PT95b, Pee96b]. Hands-on [HA94, Rau95, SS94b, SS94a, Sal95a, Sal96, Zna94, Ill91, Owe93, PT95a, PT95b, Pee96b]. hangouts [Woo95a]. Hans [MSB^{+}92]. Happy [ACE^{+}95]. Harbor [Ano94p]. Hard [Ano96s, Kan94b]. hardware [SMLD95]. Harley [Hah96a]. harnessing [Ano95-65]. Harp [LL99]. Harvard [Ste91a]. Hastings [Ano95y]. hate [Ano95-34, Hof96a]. having [Web95]. Hawaii [GV90b]. he-hadash [PEH95]. Headliner [Cer93]. Headquarters [Ano94s, Nat93a]. Health [Lib95, MD93, Anoxxa, Blo95, Bro95a, Coo93b, Edw95, Han92, Han93, Han94a, Han94b, Hob96, HH96, HM96, KT95, Mas97, NC95a, NC95b, PS95b, RJV95e, RJV95a, Rye95]. Healthcare [Hea95, Ste95b, Int95a, Mil96c, Nat95, SHB95]. HealthLink [Hea94]. hearing [Ano94-32, Ano95l]. hearings [Ano94y]. Heartland [Eve95]. Heat [She95b]. Heights [Wil94a]. held [Int76a, Ano91, BCB99, FS95, Int76b, IDK98, JG91, LG93]. Heller [Int95m]. helminthiasis [Rey91]. help [Ano96j, JV94]. Helpful [Tus93]. Helps [Car99b, NGP^{+}99]. Herausforderungen [Jun95]. here [Mam96]. Heretic [Sto99]. Heroes [Lev84]. Heterogeneous [MWB95, Pet86b]. Hewlett [Ano99e]. HFC [Bla98f]. hidden [FS95]. Hierarchical [LS89]. High [ACM98, Buy99, Cos94a, HL89a, HL89b, Jai94, Luk96, MP89, Maz90, Mei92, Sei98, Sto99, Wol98, Chi95, Gal94, KKC98, ODO95, Sch95b, SD96b, Int93a]. high-bandwidth [KKC98]. High-Performance [ACM98, Luk96]. High-Speed [Cos94a, Sei98, HL89a, HL89b, Jai94, SD96b]. higher [EGC^{+}94]. highlights [Uni89, UU89]. Highway [Ano93k, Edm95, Fis93, FD94, Kan94a, RBM94a, RBM94b, Twe94, Ano93u, Ano95-31, Cro94a, Cro95, DOO95, DG94, Eme94, Eme95a, FVNT94, LJ95, ODO95, Sta94b, Tur95e]. Highways [MS95, Ano91]. Hill [Ano95i, Dat74, WE96]. hired [Ken95b]. historians [Oti95]. Historical [RAD90]. histories [Shi96]. History [New81, Laz97, LCC^{+}97, NS94, Abb94, HG95, HH97, Win98]. Hitachi [Ano97h]. Hitchhiker [Kan94a, JEK92]. hitchhikers [Kro87, Kro89]. HJR [Vir94]. Holes [Hof96b]. Holocaust [Int85, Cen95]. Home [Hes95, Ano94v, Cha95a, HN96, HH95b, HB95, McA95, Tim94, Unixx, Kra96]. HomeNet [KSM^{+}96]. homepage [Wes95]. homes [Gil95a, Gil95b]. homesteader [MBS95, The94a]. Homesteading [Rhe94a, Rhe94b]. Hong [HL99]. Hook [Ken95b]. Hop [Wei95b]. Host [LDG96, Cho95, LMG95, PMJ94]. hosts [Bra89b, Bra89a, Eas94, KIP99, LCP91, Sri93]. Hot [Wei95c, Pow93]. Hotel [Int76a, Ano94-53, Cim96, Int76b, Mas97]. Hours [Edd94c, Edd96]. House [Ano94y, Ano94-32, Ano95l, Bro89]. Houston [GMB93]. How-to [Hen94, Ano94x, DG94, Hen95b, Mac95b]. how-to-do-it [Max95e]. HP [Ano99e, Hew88]. HPCC [Bir93]. hsiao [cH95]. hsueh [Neu98a]. HTML [Ano94v, Bri98, Che97, CC96a, Fly95, FD95, FHU99, HH95b, HB95, HH98, Int95g, JGKR96, KH95, LO99, Lam96, Lem95, LG95, LC98, MK98a, MK98b, MS98, MK98c, MK98d, MK98e, MK98f, MP98, Mor95c, MK96b, NR96, Neu98a, PMTJ98, Rus98, YmtEC99]. HTTP [Ste96b]. hui [Neu98a]. Human [BL97, Hum76, Hum80, HSHF95, Hil97, Nil93, WMRS87, WLM90, Ano89c, Fre94, Hum89, Mas89, Hum94, MW92, Ste91a, WS80, WR89]. humanitarian [Fre94]. humanities [RJV95f, RJV95b]. humanos [WLM90]. Hundred [Ano94y, Ano94-32, Ano95l]. hybrids [Kac96]. hype [PA95]. Hyper [MKBB95, Sto95a]. Hyper-Media [MKBB95]. Hypermedia [KP93, HA95, MO92]. HyperNet [MO92]. Hyperspatial [Haw96b]. Hypertext [Lam96, Nie95, Seg95]. Hypocenters [HM97].

I-Way [BK99b]. i.e [Ame95b]. I/O [Ano96v]. IBM [Dis93, IBM88, IBM89c, IBM90, IBM91b, IBM91a, IBM92a, PS92]. IBR [Ano93o]. ICANN [Mue99a]. ICCL’98 [BCB99]. ICE [Ano94r]. ICE.PPN [Ano99e]. ICS [Ano99d]. ICSC [HL99]. Idea [Ano93-34]. ideas [Ano94-47, BI96, Coo96, LMLJ95, Int93a]. identifies [Pow93]. identity [Tur95a]. Idiot [Dup95, Ken94b, Kin95b, McF95, Gol94, Ken94c, Ken94d, Ken94e, Ken95d, Ken95e, Wei95a]. IE5 [FHU99, Hom99]. IEEE [And99, BDPS85, IEE90, Mad89]. IEEE/ANSI [Mad89]. IETF [CD92, Int91, SR99]. if [Ric95a]. IFIP [LM98]. IFIP/GI [LM98]. IFLA [Ano94s]. IHR [Seg95]. IHR-info [Seg95]. iHTML [Ano99e]. II [Ano99e, Ano93b, Bla98b, Bla98c, Col95, CS98, KW98]. III [Bla98b, Bla98d, CS97, Iow94, KSW99]. ilgan [TI95]. ilko [SG94b]. ILL [SLJ94]. ILLINET [ZM93]. Illinois [ZM93]. illness [JP94]. Illusion [Hel99]. Illustrated [Eag94a, Ste94, Ste96b, WS95, CA95]. im [Ble95]. image [AIS^{+}99, Nof96a]. imaginada [Rib95]. IMHP [PMJ94]. Impact [Far97, Laz97, Nil93, AP95a, EGC^{+}94, Int92d, HH97, HGKB99, INP95, Kin94, KKM^{+}98, McC95, MS94]. implement [Lin93c]. Implementation [Int76a, CS91c, Com91a, CS91d, CS91b, CS98, CS91e, Fei93, Fei96, Int76b, Leo98d, MWB95, Sum99, Sun96, TGL^{+}97, WS95, Asa91, CS96b, Cra92, Dav95, D^{+}86, PP91, Pil90, Pun91, SRI82a, Net83, Sch91a, Sch91b, SZ99, Wis90, Zho84]. Implemented [Kon82, Ken90]. Implementing [Ben92, Cam95a, Coo95b, Gon99, Hel91b, Lei96, Vac99, Tho96c]. Implications [Gac98, Tsu89, Ano93x, AW97, Har92, HRB92, LT92, RD95]. imprescindible [Gru95]. Improve [DS97]. improved [Uni89, UU89]. Improvement [Wan93]. improvements [Web95]. Improving [FRB93]. in-state [Ano95-61]. in/for [FMS98, IE98]. Inc. [Ano99d, PM94]. Incident [Car99b, FF94]. Includes [LaQ94b, LR93b, Pot94a, WMS96]. including [Gil94b, Kw95, Kan95b, Kan95c]. Incomplete [NCS93, Nat93b]. Incorporating [Bre94]. Independent [BL93, RMU96]. index [Ano94r, Vic89]. indexer [LoI94]. indexes [Ano94d]. Indexing [EFIJ99]. India [Aga99, BGMP98]. Indiana [Ano96h]. Indies [Fra99, Hir97, Hir98]. Individual [Gil93b, Gil93a, Gil94c]. indlucing [Har94b]. Inductive [Ano99c, Pau99a, Pau99b]. Industrial [Leo98e, Ano95e]. industry [Ano96i, Cru95c, IC94, Lin93a]. INET [Lei93, TL93]. inexpertos [LB95b]. Inference [Pea98]. INFNet [BAG94, Ghi89]. Info [Rit93, Ano95-37, Wal94b, Seg95]. Infobahn [Inf95b]. Infohighway [Neu94]. InfoMagic [Ano96v]. informacion [CS95]. informatica [Ano95b]. informatics [Cim96]. Information [ANS92, Ano93r, Ano94-40, Ano95f, Phi95, Ano95-61, Ano96w, Ano99b, Bre93, Can94, Coo95a, Dec95, Den99, DJB93, Eag94a, Eag95a, Edm95, End94, FS95, Fie93, For95, FL95a, FD94, GMB93, HSF94, Har99a, Har93a, Hen99, Inf93, Kah92, Kah94a, KW93, KW94, KS97, Leg94b, LS93a, LC93, LPJ^{+}94, Man95a, MMLP97, MWL93, New93, OD95, PK94a, Pet93a, PM94, Rit92, RS97, RBM94a, RBM94b, Sau92, SYTM98, Sto95b, SMB94, Twe94, Uni93b, Wal93, WYLO98, Wes95, AP95a, Aik95, Ano93u, Ano94z, Ano94t, Ano94-35, Ano95w, Ano95-30, Ano95-31, Ano95-38, Cin96, AT94, Bra94a, Bag95, BR94b, Bla96c, BDC^{+}95, Bro95a, Con95a, Cal94b, Cal94a, CS94a, CHR^{+}96, D’E95, Dil93, DKF^{+}95, DG94, Eme94, Eme95a, For94, Fel94b, FK95a, FB91, FHS94, GP95a, GP96a, Gie94, GP95b, Gum93b]. information [Gum93c, Gum94, HSHF95, Har94b, Int85, Kah90, KG98, Ken95g, Kle94, KW98, KSW99, KL95, KT95, LJ95, LS93b, LG99, Leo94b, LJN94, LSBM95, LC96, Lin93a, Lin93c, Liu95, Liu96, Min94b, Mac95b, MM95, MMJ95, Mat95a, Max95e, MP91, McM94, MW92, Mil94b, Mle95, Med94b, MRJ96, Mot94, NN94, Nat95, New94a, New95b, Nov94, Ott94, Par94a, Per95b, Pri95, RT94b, RHK95, RK95c, Rud95, Rya94, Rya95a, Rya95b, Rya95c, Rya95d, Rya95e, Rya96a, SUR92, Sch94, Sch95b, Sch90, SP94a, SBP^{+}95, Sil95, SM93, Sne95a, Sta94b, SHB95, TW95, TL96, Thi94, Tho95b, TPH96, Tur95e, U.S93, US 92, US 93, Wal95b, Wan93, Wat93, WSA^{+}94, Wil95a, YM96, Bro89, RHS95]. informational [RM94]. informazione [Gio95]. InfoScavenger [Int96d]. Infoserver [Bag95]. Infotrac [Inf93]. InfoWorld [Inf80]. Infrastructure [Ano95f, HL93, HPS95, Kah92, Kah94a, KW93, Mes99, SQ93, SS99, Wei93, Coo95a, Gro95a, Kah90, TFG99, U.S93, BM98]. infrastructures [Mue99b, Wes97]. ing [Dam97]. ingles [MPF95]. ingles-espanol [MPF95]. Ingres [Ano99e]. initiated [MW94d]. initiation [LSLS94]. Initiative [Byr95, VJ95]. Injury [PS95b]. Inline [Ano99e]. Inner [Sch95b]. Innovative [Ano96u, Ano94-55]. innovator [Eme95b]. Inquiry [SW97]. Inroads [Cho96]. ins [BGGS95]. inservice [De 94]. insider [The94b, Pre95b, She95a]. Insight [BGGS95]. INSPEC [Ano94-33]. instability [LMJ98]. Installation [Hew86, Hew88, IBM89d]. installing [DS96, Hf96]. Instant [Ano94u, HA94, Hof94b, Jam94, SS95a, HY94, Hof95f, KP94]. Institute [Ano96x, Lam96, LR96, Bro96]. institutes [Fol95a]. Institution [Ano93d]. institutions [Ale95, Rot95a, Rou94]. instruction [BI96]. Instructional [TOLF93, CT95a, CT95b, CT96a, JM96b, JT96, TOL94a]. instructivo [TOL94b]. instructor [CT96b]. insurance [Cru95c, Gor95a]. Insurers [Hal95a]. Insurgent [Dix95]. Intanetto [Hof94a, KM95, MY94, Nih95, Nag94, Oma95, Rik94, Edd95a]. Integra4 [Ano97h]. Integrate [LG95]. Integrated [Heg99, ME95, ME97]. Integrating [Bur95a, Int93a, DO96, Moh95, Ste95a, Str94]. Integration [BB99, Wil94b, Mil95e]. Integrity [Wu97, Wis90]. Intel [Dig82]. Intellectual [Sim95a, OO95]. Intelligence [OKP97, Dav96, Inf95a, Mar99]. Intelligent [CWJT99, MMLP97, Pea98, SS99, Bla98e, KL98, PJ95, Wat96]. intensive [IDK98, Ker95]. intention [SJZ^{+}98]. intentions [Sch95b]. inter [Ste99b]. inter-domain [Ste99b]. Interacting [Eng96b, WWBD94]. interaction [Rya95i]. Interactive [Ano93l, Moh99b, Par93, Pre93, Ras94, She95a, Sin99a, Car95c, Chi95, GP95b, IE98, Kac96, Kum96, LG95, Tho96b, Tho95b, Ano95-55]. Intercat [JO95]. Interchange [BU94a]. Intercommunication [CK74].

Intercomputer [Pad85]. interconnected [Kon91]. Interconnecting [Bra93a, McC88]. Interconnection [BDPS85, Yan88]. Interconnections [Per92a]. interdisciplinary [B^{+}96b]. Interest [Ano93c, Coo93a, Ano92f]. interests [SW93]. Interface [Ano96v, PCD92, Ros89, Ano93-45, EW94, Fuo95, Gil95c, Kas93, PH97]. Interfaces [McQ94b]. interim [Spe87]. Interlink [Mar99]. internacional [Fre95]. internal [Ber96a, PS92]. Internals [CS91c, Com91a, CS91d, CS91b, CS98, CS91e, Zim95b, CS96b]. International [Ano99e, Bre95b, BG98, BGMP98, Cen95, FS95, FRG97, Fra99, GPRR94, Hir97, Hir98, HL99, Mag94, IDK98, KW98, KSW99, Lei93, MKT98, Not94a, TL93, Uni96a, Uni96b, Ano94-41, Net96b, Coh99, Cro96a, D^{+}86, Int94g, Int85, Soc95a, Soc95b, INT88, INT89b, LM98, Sea90, She77, SG94b, Vic88, VT97, Wal95a]. Internationalization [Hiu96]. internationell [Dam93]. internaute [Soh94, Soh95b, Soh96]. INTERNET [CKMV99, Ano96v, AB94, All94, Amo99, ACE^{+}95, Ano69, Int92b, Ano92b, Ano92i, Ano93-33, Ano94b, Ano94f, Ano94-42, Ano94r, Ano95g, Ano95z, Ano95-55, Ano95-68, Ano96b, Ano96c, Ano96f, Ano96s, Ano96t, Ano96u, Ano97d, Ano97h, Ano98a, Ano98c, Ano99a, Ano99d, Ano99e, BCB99, Bee93, Bil99, BGGS95, Bla99c, Ble95, BJS94, Bos94, Bud96, Com93, Coo92, Car94a, Cat96, Che93a, Che94, Che95, CB94c, Chr98, Cim96, CO99, CP94b, Duf95, Duf96, EL95, EM95, Est93a, Gil93a, Gos95, HS94c, Har96a, HJ93, HH98, Hor95, HL99, Int89a, IBM89a, IBM89b, Ins96, Kan95b, Kan95c, Ken95h, KF94, KW98, Koe95, Kur96a, Int92c, LaS95a, Lap96, Lar94b, LJN94, Lin93a, Low90b, McM94, McN95, MW92, Mog99, Mol93b, Mou96, Nat94a, Nej94, NDM98, Noo91a, Noo92a, Noo92b, Noo94, OO96].

Internet [Ols96, Pea95, Per92e, Pla95, PSC81, Qua93, RM96, Roa95b, Roo97, Rot96, SBGK94, SR99, She85, Sne95b, Soh95a, SLR91, Sta94b, SB94, Thi93, Uni96a, USE97, VP95, VFH98b, VT97, Wol98, BP94, Cal91, YmtEC99, Abb94, Ano95b, Bol94, Cha92, Des93a, D^{+}86, Hof95d, Mag94, INT88, INT89b, Kas93, LB95d, McC92, PS92, Pol92, Pos80, Soh94, Soh95b, Vic88, Vic89, VVB85, cH95, kTlC94, Ame95a, Ame95b, And95b, AUA94, AL96, AUC93, Ame95c, Ame95d, Ass94b, Abb99, AR98, AJR95, Acc95, Ack95b, AAS^{+}97, AP95a, Aga99, ANT96, Ahu96, Aik95, AFZ95, Ale95, AP96, Ale96, All95, All96, Als95, AAB^{+}98, AIS^{+}99, AHS90, And95a, And92, And94, And95c, And95d, AL95, AN96, AH95b, Ano74, Int76a, Ano79, Ano81, Ano88a, Ano88b, Ano89a, Ano89b]. Internet [Ano90b, Ano90c, Ano90d, Ano90e, Ano91, Ano92a, Ano92e, Ano92c, Ano92d, Int92f, Ano92f, Ano92g, Ano92h, Ano93h, Ano93c, Ano93a, Ano93b, Ano93f, Ano93e, Ano93d, Ano93g, Ano93j, Ano93i, Ano93-27, Ano93m, Ano93w, Ano93v, Ano93o, Ano93p, Int93h, The93, Ano93q, Ano93r, Ano93s, Int93i, Ano93t, Ano93u, Ano93x, Ano93y, Ano93z, Ano93-30, Ano93-31, Ano93-32, Ano93-35, Ano93-40, Ano93-39, Ano93-38, Ano93-43, Ano93-42, Ano93-45, Ano93-46, Ano94e, Ano94h, Ano94d, Ano94j, Ano94i, Ano94k, Ano94l, Ano94n, Ano94m, Ano94o, Ano94p, Hea94, Ano94q, Ano94-62, Ano94z, Ano94y, Int94e, Int94d, Int94h, Ano94v, Ano94x, Ano94s, Ano94-43, Ano94u, Ano94-27, The94b, Ano94w, Ano94-44, Ano94-45, Int94j, Ano94-28, Ano94-29, Ano94-31, Ano94-34, Ano94-35, Ano94-32, Ano94-33]. Internet [Ano94-36, Ano94-37, Ano94-39, Ano94-38, Int94k, Ano94-40, Ano94-41, Ano94-47, Leg94a, Ano94-49, Ano94-48, Net94d, Net94b, Ano94-54, Ano94-53, Ano94-55, Ano94-57, Thr94, Ano94-60, Ano94-61, Ano95a, Ano95c, Ano95d, Bus95, Ano95f, Ano95e, Ano95k, Cor95, Ano95m, Ano95l, Ano95j, Ano95n, Ano95p, Ano95s, Ano95q, Ano95t, Ano95v, Ano95u, Ano95w, Ano95x, Hea95, Ano95-30, Ano95-31, Ano95-29, Ano95-33, Int95d, Ano95-34, Int95c, Ano95-35, Ano95-27, Ano95-47, Ano95-37, Ano95-28, Ano95-38, Int95e, Ano95-36, Int95f, Inf95b, Int95h, Int95j, Int95i, Ano95-39, Int95l, Ano95-32, Ano95-40, Ano95-42, Ano95-41, Int95m, Ano95-57, Ano95-43, Ano95-44, Int95n, Ano95-45, Int95o, Ano95-46, Ano95-51, Ano95-48, Ano95-49, Leg95a, Ano95-50, Ano95-52, Mar95c, Mec95b, Ano95-53, Net95d, Net95a, Phi95, Ano95-58]. Internet [Res95a, Ano95-59, Ano95-61, Sma95, Ano95-62, Ano95-64, Ano95-63, Ano95-65, Ano95-66, Of 95, Ano95-69, Ano95-70, Ano95-72, ZD 95, Avi96, Bio96, Cin96, Cyb96, Ano96d, Ano96e, Ano96g, Int96b, Int96c, Ano96j, Int96d, Int96g, Int96f, Int96h, Int96i, The96a, Ano96a, Ano96i, Int96j, Ano96k, Ano96l, Int96k, Int96l, The96b, Ano96m, Ano96h, The96c, Ano96o, Ano96n, Ano96q, Ano96p, Ano96r, Med96, Net96b, Ano96w, Yah96, Ano97c, Ano97g, Jou97, Ano97i, Ano97j, Ano97e, Ano97f, Ano98b, Ano99c, Anoxxa, Anoxxb, Anoxxc, App95, Arc95, AW97, Arm94a, Arm94b, Arm96, Arn95, AA94, Arn96, AADF96, Asa91, AK97, Ash95, AP95b, Aug96, AT94, Ayr94, BN90, Bri95b, Bri95c, BB95, Bri95d, Bri95e, Bri94b, BU94a, Bra94a, BU94b, Bac95b, Bac96]. Internet [BW93, Bac94, B^{+}94, BS93b, BS93a, BL94, BL95a, BL95b, BL95c, Bad95, Bag95, Bak93a, Bak93b, Bak93c, Bak94, Bak95, Bak98, Bal93a, Bal93b, Bal91, Ban95, BMG^{+}94, Bar94, BL93, BK91, BD95a, Bar95b, Bar92a, Bar92b, Bar93, BI96, BH94a, BH95, BH96, Bau95, Bau93, Bax95, BOB93, Bea94, Bea95, BGS94, BGS95, BM98, BC95a, Bel95, BT96, Bel93, BCRL95, BBK^{+}98, Ben92, Ben95a, BC95b, BWWF99, BR94b, Ben94, Ben95b, Ben95c, Ben98, BR95, BA95, Ber95a, Ber96a, BBN^{+}93, Ber95b, B^{+}96a, BZ95, Ber95c, Ber96b, BP93, Bet87, BK99a, Bhi96, BBSH98, BC95c, BM96a, Bir90, Bir93, Bla93a, Bla95c, Bla95d, Bla93b, Bla96c, Bla93c, Ble93, Blo95, Boc96, Bog94, Bog82, BX83, BH94b, Bol96a, Bol96b, BLL^{+}96, BAG94, BK99b, BEA98, BB99, BK96a, Bou97]. Internet [Bow95, Bow96, BM94, BPR88, Bow90, BP91, BDS93, BDMS94, BBST96, Boy96, BP87, BD88, BBP89, Bra89b, Bra89a, Bra95, BU95, BM96b, BW94a, BW94b, BWC95a, BWC95b, Bra97, Bra94b, BDC^{+}95, Bra93b, Bra94c, BL96, Bre91, BB93, Bre95a, Bw96, Bre94, BD95b, Bre95b, Bri95a, BEM96, Bri94a, Bro89, Bro94a, Bro95a, Bro96, Bro94b, Bro95b, Bro91, Bry95, Buc95b, BS94, Bup95, Bur95a, Bur96, BG98, BGMP98, Bur93, Bur95b, BS96, BKH95, But94, But95c, But95b, But95a, Buy96, Byr95, CAC94, Con94, Con95a, Con95b, CCN94, Car96, Can94, Cal94b, Car93a, CN95b, Can95, ST95, CTN95, Col93, CKPT99, Cac89, CM95, Cai95, Cal94a, Cam94a, Cam94b, Cam95a, Cam96, CB95a, CC95, Cam95b, Cam95c, CS94a, CS95, Car94b, CA95, CHHL91, Car95a, CB95b, Car99b, CB94a, CB94b]. Internet [CB95c, Car95b, CB95d, CG95, Car95c, CB96a, CD92, Cas94, Cat89, Cau95, CD95, Ced95, CC83, Cer89, Cer91, Cer93, Cha95a, Cha95b, nChCiY94, CWJT99, CDEF94, CZ95, Che93b, CNO94, cCrmdh92, CP94a, Che97, Che96, Che87, Chi94, CD94, sC86, Chi95, Cho95, Cho96, CS94b, CC96a, Cla94, CCC^{+}91, Cla95d, Cla95b, Cla95a, Cla95c, Cla96, CRH98, CL98, Cla99, CT95a, CT95b, CT96a, CT96b, Cle93a, Cle93b, Clo93, Cly93, CN95a, Cob94, Cob95, Cob96, Coc95, CP98, Cog94, Cog95, Coh99, Col99, Com95a, Com95b, Com97, Com95d, Com94c, Con89, Coo96, CC96b, Coo95b, CRI95, Cor96, Cot95, Cra92, Cra99, CES^{+}99, COZ99, CP82, Cro93, Cro94a, Cro95, Cro96a, Cro96b, CB91, Cro96c, Cru95a, Cru95b, Cru96a, Cru96b, CKY94, Cru95c, Cul95, CW93, CTG96, Dar93a]. Internet [Dis93, Dig95, D’E95, D. 93, Dat94, Dav96, DL96, DO96, DG97, Dam97, Dam93, Dar92, Dar93b, Dav95, DLC92b, DLC92a, Dav94, DOB95, DeR96, De 94, Dea96, Dec95, DN99, Dee95, Dee93, DH96, Del93, Den89, DD98, Den96, Der94, Des93b, Des94, DS97, Deu92, DFS98, DFG^{+}98, DHS95, DOO95, Dil93, DJB93, Din94, Dix95, DKF^{+}95, Don96, Don95, DH99, DHM^{+}96, DS95, Dow86, Dra95, DG94, DG95, Dre92, Dre93a, Dre93b, Dre94, Dre95, DS96, DGH93, Dyr93, EKG94a, EKG94b, Eve95, EH95, Edu95, Eme94, Eme95a, Ent95, Eag93, Eag94c, Eas95, Eas94, EFIJ99, Ebi95, Eck95, Edd94a, Edd94b, Edd95a, Edd95b, Edd94c, Edd96, Edm95, Edw95, Ehn94, EO99, Eis95, Ela94, EHPT95, Ell94, EE94, EE95a, EE95b, EE96, EB^{+}97, ER98, Eme95c, Eme95b, EV96, EGC^{+}94, ED92, Eng96b]. Internet [Eng93a, E^{+}93, Eng93b, Eng95a, Eng93c, ED94, Eng94a, Eng94b, Eng95b, ELS95, ELO95, Eng95c, Ens95, ECC94, EE95c, Ess95, Est93b, Est93c, EW94, Eva95, Ewe95, EHS92, FAR93a, Fir95, Fre94, Flo95, Foc95, FTP95a, FTP95b, Fer95, FP94, Fah94, FPAR95, FS95, Fai95a, Fai95b, Fal94a, Fal94b, Fal96, Fan95, FE91, FE92, For94, Fee93c, Fee93a, Fee93b, Feh93, Fei94, Fel93, Fel94b, Fel94a, FK95a, Fel94c, FK99, Fie93, Fin95, Fis93, FVNT94, Fis95b, FF94, Fle94, FB95, FF99, Fol95b, Fol95a, Fon94, For95, FL95a, FL95b, Fos91, FB91, FRG97, Fra93a, Fra93b, FHS94, FH94a, FH94b, Fra94a, Fra94b, FAW95, FJ95, Fra94c, FPV95, Fra96, Fra95, FKA95, Fre95, FMB95, FH95, Fre93, FMS98, FBGP98, FP95a, FP95b, Fry95, Fum98, FK95b, Fuo95, Fur99, Gru95, oIT94, GNU95a, GNU95b]. Internet [Gal95a, Gal96, Gaa76, GP95a, GP96a, Gac96, Gac98, Gad94, GKC93a, GH93, GKC93b, GH94a, GH94b, GH94c, Gaf94, Gag94a, Gag94b, GR94, Gal95b, Gal94, Gan95, Gar97, GLSW89, Gar89, Gar94, GRRM95, Gar95a, GBS95, GS95a, GS96a, Gar93, GRLJ94, Gas96, Gen95, Gen83, Ger91, GFF^{+}95, Gha92, GM94a, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95c, Gia95d, Gia95a, Gia95b, GMDS96, Gia94, GCW95, GS95b, GS99, Gie94, GH94d, GK96, GB95, Gil95a, Gil95b, Gil93b, Gil94a, Gil94b, Gil94c, Gil95c, GVJ95, Gil95d, Gil95e, Gil96, GK95, Gio95, Gla94, GG94, GG95a, Glo95, GG95b, GG96a, GM94b, God94b, GGA^{+}97, GAM^{+}97, GVW94, GVW95, Gol92, Gol94, GG96b, Gon95, GBJ94, GPRR94, Goo96, GCS96, Gor95a, GS96b, Gou90, Gra95, GCGMP97]. Internet [GGMT99, GA95, Gre95a, Gre95b, Gre96, Gre95c, Gri95, GV90a, GV90b, Gro90b, Gro90a, GP95b, Gro95a, Gro95b, GO95, Gru94, Gry94, Grz94, Gue96, GB96, Gui95, Gum93a, Gum93b, Gum93c, Gum94, Gun94, Gup92, GP96b, GRS79, HH95a, Hel91b, Hum76, Hum80, Int81, Int86, Int87, Her94, Haa85, HMG^{+}95, HL96, HS94a, HS94b, HSF94, HS94d, HS95a, HSskw95, HS95b, HS95c, Hah96a, Hah96b, Hal97, Hal93, Hal95a, Hal99, Hal94, Hal95b, HWV96, HC94, HTW94, Han92, Han93, Han94a, Han94b, Han95, Han96, HCBO99, HSHF95, HK91a, HK91b, HN92a, HN92b, HN93, HSK98, HE95a, HOO95, HAS95, Har94b, Har93a, Har94c, Har94e, Har94d, HU95, Har95b, HKB95a, HKB95b, HK95a, Har95a, Har96b, HS98, Har94g, Har92, HRB92, HG95, HLB^{+}95, Has96, HP96, HE95b, HH97, Haw96a, Haw96b]. Internet [Hay94a, Hay94b, Hay94c, Hay95, HN96, Hf96, HV95, He98, Hed87a, Hed87b, Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b, Hei95a, Hei95b, Hel94, HL89a, HL89b, Hem93, Hen82, Hen94, Hen95a, Hen95b, HM97, Hen99, HT95, HKR91, Her96a, Her96b, Her95a, Her95b, Her95c, HA94, HH95b, HB95, Hes95, HA95, Hic94, Hil96, Hin91, Hin93, Hob96, Hoe95, Hof94b, HY94, HL94a, HL94b, Hof95e, Hof95a, Hof95b, Hof95f, Hof95g, Hof95c, Hof96b, HKN96, Hof96a, HH96, Hol94, Hon95, HG97, Hor98, How95a, How95b, How95c, HGKB99, Hub96, Hud94, Hug94, Hug95, Hui95, Hui96, Hui99, HGKM98, HB96, HM96, Hus94, HPS95, Inf93, Inf95a, Inf96, Ano93n, IBM88, IBM89c, IBM89f, IBM89d, IBM89e, IBM90, IBM91b, IBM91a, IBM92a, IBM92b, Int93g, Int91, Inn96, Int85, Int96e]. Internet [Ill91, IC94, Int76b, Int96a, Int95a, IRC95, Int95b, Soc95a, Soc95b, Inf95c, Mec96, Int94l, IE98, INP95, IV95, Int95p, Int95q, Ive95, J3 95, JG91, Jac94, Jaf94, JT94a, JT94b, JT94c, JW95, JFJ96, Jam96a, Jam96b, Jam94, JC95, JCD96, Jan95, Jan96, JN95, JE95, JV94, Joh94, Joh95, Joh96a, Joh96b, JE96, Jon90, Jon94b, Jon94a, Jor94, Jos95a, Jos95b, JP94, JO95, Jun95, JM95, JM96a, JM96b, JT96, JD95a, JD95b, Kac96, Kah90, KK95a, Kah94b, Kal94a, Kal94b, KIP99, KG98, Kan95a, Kw95, Kan95b, Kan95c, Kap96, KG94, KM96a, KLW95, Keh92, Keh93, Keh94, KMM95, Keh96, Kel92b, Kel92a, Kel96, Ken95b, Ken95a, Ken90, Ken93, Ken94b, Ken94c, Ken94d, Ken94e, Ken94a, Ken95d, Ken95e, Ken95f, Ker95, Kes97, KW94, KP95, Kic94, KH95, Kie95, KM96b]. Internet [Kin94, KRS94, KRK95, KKB95, Kla94, KK95b, Kle95, Kle94, KKF95, Klu95, KK96a, Kno96, KHJG91, KHJ91, KB92, KHJ92, Koc93, KSM98, Koh97, Kol96, KL95, KG97, KG99a, Kon82, KT95, Kot93, KG99b, KL94, Kov95, Kra94, KSM^{+}96, Kra96, KKM^{+}98, KS95, KEdT96, Kro87, Kro89, Kro92, Kro93, KP94, Kro94, Kro95a, KCG95, Kro95b, KK96b, Kro95c, Kug95, Kum96, Kuo97, KL98, Kur96b, Leg94b, LaM97, LGRS93, LR93a, LR93c, LR93b, LaQ94a, LaQ94b, LR94, LDH95, LR95a, LaS95b, LMJ98, LO99, LT92, LT93, Lam96, LJ95, LH93b, LH93a, LDG96, Lan94, LS93a, Lan93, LNH^{+}93, LSLS94, Lar94a, LoI94, LS93b, LeB93, LL95a, Le 96, LeG95b, LE96, Lea95a, Lee98, Lei85, Lei93, Lei94, LCC^{+}97, Leo94a, Leo94b, Les95, LS95, LS96, LG93, LJN94, LN94, Lev95, LN95, LB93, LB94a]. Internet [LB94b, LY94a, LY94b, LB94c, LY94c, LB94d, LBY95, LY95a, LB95b, LB95a, LB95c, LY95b, LSBM95, LBP95, LBY96, Lev96, LY96, LR95b, LR95c, Li90, LD99, Lic94a, Lic94b, Lic95a, Lic95b, Lin93b, Lin93c, LMLJ95, Lit93, LPJ^{+}94, Liu95, hLcL95, Liu96, LR96, Lof95, LCP90, LCP91, LF95, LMG95, Lot92, LG95, Lov99, Low90a, Low91, LS89, Luc94, Luk96, LD95, Lun94, LR93e, LL95b, LL96, Min94b, Mer92, Mer93, Mas95, Mas96, Mac95a, MMV93, MMV94, Mac95b, Mac93b, MBM^{+}95, MM95, MRTR95, MS96, Mad96, Mad98, MMJ95, Mah95, MW94a, MW94b, MD94, Mai95, Mai96, Mak93, Mal95a, Mal92, MDB^{+}94, MM91, MM92, MP93, MPF95, Mal95b, Mal95c, Mam96, Man95a, Mar95a, Mar95b, Mar94, MK92, MKNW92, M^{+}93, MKNW94, Mar93, Mar96a, Mar98, ME96, Mar96b, MG96, Mat93]. Internet [Mat95a, Mat96a, Mat96b, Mat95b, ML95, MG94a, MG94b, Max95a, Max95d, Max95b, Max95e, MKBB95, May96, MP89, Maz90, McA95, MG94c, McB95, McB96, McC95, MMR93, MMR94, McC94a, McC94b, MBZ94, MBB95, McC97, McF95, McK94a, McK96a, McK94b, McK95, McK96b, MB93, McL96a, McL96b, McQ94a, Mec95a, MD96, Mee95, Mei82, Mei92, Mel94, MS95, MPF96, Mey95, MD93, Mil92a, MZ93, Mil94d, Mil94c, Mil94b, Mil94a, Mil95d, MK95, Mil95a, Mil95b, Mil96a, Mil96c, Mil96b, Mil92b, Mil93c, Mil93d, MW94d, MW94e, Mil94e, Mil95e, MK96a, MSD90, Mil90, MHR91, Mil95f, Mil94f, Min94a, MPG94b, MPG94a, Mis95, Mit92, Mle95, Moc94, Mog88, Mog89, Mog95, Moh99a, Moh95, Ban94, Moj93, Mol93a, Mol99, Moo96, Moo95, Mor95a, MSB^{+}92, MKM94, Mor94, Mor95b, Med94a]. Internet [Med94b, Mor95d, MRJ96, MNT93, MNT94, Mot94, Moy98, Mse90, Mue99a, Mue99b, Mui94, Mul97a, MP95, MLJ95, Mul92, MS94, MY94, MBS95, Nat93a, Net94c, Net95c, NCS93, Nat93b, Nat96a, Nor95, NN94, New95a, Nat94b, Nat95, NSF89, NSF91, NSF92, Nat96b, NI95, NIP96, NYS92, NYS94, NC95a, NC95b, NS94, Nram94, Nef94, Nel95, Neo94a, Neo94b, Neo95, NR96, Net96a, Net95b, Neu98b, New93, New94a, NST96, New95b, New94b, New95c, New96, Ngu95, Nic95, NGP^{+}99, Nie95, NC90, NDM98, Nof95a, Nof95b, Nof96c, Nol97, Noo91b, NM94, Nor92, Not93c, Not93a, Not93b, Not94a, Not94b, Nov94, NM98, Nug95, O’H93, Who93, O’R96, Obl92, ODL93, Och95, OM95, Ogb98, Ols95, Ano93-29, Opp95, Opp97, Opp98, OD95, Osh93, Oti95, Ott94, Ove94, ODO95, Owe93, PA94, Pee96a, PE94, Pub94]. Internet [Pub95, PG96, Pad95, PE99, PRY99, Pan95, Pan96, Par96a, Par94a, Par93, PA95, Par95, Par96b, PP91, Pas95, Pas96, Pau99a, Pau99b, Pax97, Pax99, Pea94, PM95b, PG95, PT95a, PT95b, Pee96b, Peh99, Pel95a, Pel95b, PK94b, PK94a, PMJ94, Per95a, PN91, Per91c, Per91a, Per91b, Per92d, Per92b, Per92c, Per95b, PR95, Pet93a, Pet86a, Pet86b, Pet87, Pet89, PR91a, PR91b, Pet93b, PCF91, PK99, Pet95, Pfa94a, Pfa95a, Pfa95b, Pfa95c, Pfa95d, Pfa95e, PW95a, Pfa95f, Pfa96a, Pfa96b, Pfa96c, PU94, PHW^{+}93, PH97, PS95a, Pie95, Pig95, PP94, Pik95, PB95, Pik96, Pil90, Pil91, Pit95a, Pit95b, Pit95c, Pla94a, PDP96, Pla94b, Pod95, PMB^{+}92, PGS93, PGS94, PC95, Pol96, Pop94, Pos95, Pos83, Pot93, Pot94a, Pot94b, Pot95]. Internet [Pow93, PG93, PG94, PN95, Pre93, Pre94, Pre95a, Pre97a, PBF^{+}98, PS95b, Pre95b, Pri95, Pro96, PM94, PWK94, PW95b, PKL90, Pun91, Pyl96, Pyr95, Que95, QP95, QCM94b, QCM94a, Que94, Ric94a, Rae95a, Rae95b, Rae95c, Rae95d, Rae95e, Rah94, RRS98, Rak95, RR95a, Ram95, Ran94, Ran95a, RRF95, RL95a, Ran95b, Ran96a, RP93, RSD98, Ras94, Rat93a, Rat93b, Rat93c, Rav95, RZ96, RK95a, RK96, Ree93, Reg93, Reh96, Rei93, Rei92, RB96, RBK^{+}94, RG91, RT94b, RT94a, RA96a, Res95b, Res97, Rey91, Rey93, Rib95, Ric95a, Ric94b, Ric95b, RdG95, Rii82, Rik94, RR94a, Ril95, RRO96, Rim94, RdB95, Rim95a, Rim95b, Rim96, Roa95a, RL95b, RML^{+}94, Rob95, RN93, Rob94, RK95b, Roc93, RO94a, RO94b, RCL96, Rom96, Ros84, Ros93, Ros94a, Ros94b, Ros95a, RS98, RHK95]. Internet [RJV94, RJV95e, RJV95a, RJV95f, RJV95b, RJV95c, RJV95d, Ros95b, RHS95, Ros96, RN97, RV99, RAWS95, Rot95a, Rot95b, Rou94, RM94, Row96, RK95c, Rub94, Rub95, Rud95, RMU96, RA96b, Rus98, Rus94, Rut96, Rya94, RNW^{+}94, Rya95a, Rya95b, Rya95c, Rya95d, Rya95e, Rya95f, Rya95g, Rya95h, Rya95i, Rya96a, Rya96b, RH97, Rye95, Sua94, SI 93, Spe93, Sil96, SUR90, Cen93, Int92e, SPR95, SRI82a, SRI82b, SDUU85, SRI82c, Net83, The94a, SUR93, SUR92, SS94b, SS95a, SB95, SLP94, SLP95, SS94c, Sal95a, Sal96, Sal95b, Sal95c, SBM96, Sal95d, Sam94, Sam95, SC94a, SC95a, San94a, San95, San92, Sar94, Sat95, Sau92, SG93a, SMLD95, Sav93, Sav94, SRL96, Saw93, SBGK94, SC94b, SC92, SC93, Sch95a, Sch96a, Sch95b, SSC96, SK95, Sch95c, Sch95d, Sch95e, Sch92a]. Internet [SS95b, Sch95f, Sch93a, Sch96b, ST89, Sch90, Sch91a, Sch91b, Sch91c, Sch92b, SQ93, Sch93b, SP94a, SBP^{+}95, Sea90, Seg95, SPR98, SSI^{+}98, Sei94c, Sei94b, Sei94a, SP95a, SRR97, Sfe82, SLJ94, Sha95a, Sha91a, SR95a, SR96, Sha91b, Sha95b, She95a, SCJ96, She86, She77, SYTM98, Shi96, SY95, Shi94, Shu93, SG95, Sil95, Sim93, Sim95a, Sim95b, Sin81, SM93, SFL95, SE96, SH95a, Slu95, Sme96, Smi93b, Smi93c, Smi93d, SG93c, Smi94a, SLB94, Smi94b, SG94b, SG94c, SG94a, Smi95a, SS95c, Smi95b, SGM95, Smi97, SR95b, Sne95a, Sne95c, Sny95, SS95d, SS99, SW97, Som94, Son92, Sou93, Spa88, Spa89a, Spa89b, SRV95, Spe87, Spi92, SPR94, Sri93, SL91, SL92, Sta95b, Sta98a, Sta98b, Sta88, SS95e, Sta94b, Sta95f, SS96, oS94, SRS^{+}95, Ste96a]. Internet [Ste91a, SC95b, SC95c, SC95d, Ste95a, Ste95b, Ste99b, SHB95, Sti94a, Sti94b, SH95b, Sto93, Sto95a, SMB94, Sto95c, SD96a, Str92, Str93a, SB94, Str94, SSA95, SC88, Sua95, SK98, SD96b, Sun95, SS95f, Swa91, SMM96, Syd96, Tex95, Tsu95, TRHR96, TW95, TMF97, TFG99, TL96, Tat95a, Tat95b, TK95a, TK95b, TK96, Tav95, Tay95, Tee94, Tem95, Ten92, TOLF93, TOL94a, TOL94b, Tes96, Tha95, Thi94, Tho95a, Tho96a, Tho96c, Tho96b, Tho95b, Tho95c, Tho96d, TB93, Til95, Tim94, TR95, TG95, TL93, Tol94c, TPBH94, Tom93, Too95, TS98, Tre95, Tri87, TB90, Tri91, Tri93, Tro94, yTZ94, TPH96, Tsu89, kTlC95, Tür94, Tur95a, TH96, Tur95c, Tur95d, Tur95e, T^{+}98, Tus93, Tys95, Uni96b, Uni95d, Uni95c, Uni89, Uni93a, NAS90, NAS93, U.S83]. Internet [Uni93b, UU89, Uni95b, Uta94, US 92, US 93, Uni93c, Uni95e, Ubi95, Upj95, Vir94, Lea95b, Vir97, Vac96a, Vac96b, Val93a, VJ95, VE96, Van95, VFH98a, VJ98, Van94, VB95a, VB95b, VH94g, VH94a, VH94b, VH94c, VH94d, VH94e, VH94f, Vel95, Vc94, Via94, VM95, VE98, Vin94a, Vin94b, Vin95b, Vin96, VJ99, VRFJ97, Vol97b, Vol97a, Wal94b, Int94c, Ano94-30, BAS95, Int93a, Wor94, Str95, Wes94, Wen94, WC95a, Wis96, WC94, Wag95b, Wag95a, WE96, Wal94a, WG94a, Wal95a, Wal95b, Wam95, Wan93, Wan94, Wan96, WLW98, WYLO98, WYLZ99, War99, War95, WM95, Wat93, Wat96, WvRA96, WK95, vdW95, Wei95a, WWBD94, Wes95, Wes97, Whi95, WMS96, Wie96, Wig93, Wig95b, Wig95a, WSA^{+}94, WLKS98, Wil95a, Wil98, Wil94b, WNS^{+}94, Wil95b, Wil96b, Wil96a, Wil95c, Win96a]. Internet [Wn95, Win93, Win98, Win96b, WS80, WMRS87, Wis90, Wit96, Wol96, WC95d, WG94b, Woo95a, WoRIGSoO95, Woo95b, Woo96, Wri95b, Wri95a, Wya94a, Wya94b, Wya95a, Wya95b, Wya95c, Xer81a, Xer81b, Xie94, Yan93, Yan94, Yas95, YN97, You94, YL95, You95, YA95, Yu93, ZB95, Zaz95, Zef96, ZCD97, Zel95, ZDE93, Zha94, Zha96, Zho84, ZTLV98, Zil95a, Zil95b, Zil95c, Zil95d, Zu 82, dVdVI98, Hum98, Laz97]. Internet-accessible [Int89a, SLR91, Ano90c]. Internet-Aware [Kuo97]. Internet-Based [EO99, GS99, RV99, Gla94, Lee98, PH97]. Internet-facilitated [SB94]. InTerNet-LINK [Ano96v]. Internet-wide [Bow90]. Internet/Bitnet [Noo92a, Noo92b, Noo94, Noo91a, Bos94, Anoxxa, Dre92, Dre93a, Dre93b, Dre94, Dre95, Han92, Han93, Han94a, Han94b]. Internet/Intranet [Ano97h, Bud96]. Internet/NREN [Lin93a, McC92, Ano93v, MMR93, MMR94]. Internet/School [Des93a]. Internet2 [BM98]. internetable [Mas97]. Internets [CD99, EST93d]. Internetting [TJ97]. InternetUser [Int96m]. Internetwork [KvRvST93, SCT81, Sta95c, VSN86, ET91, Kon91, KvRvST92, RT99]. Internetworked [ME97]. Internetworking [EI 95, Bla98d, Cly88, CS91d, CS91b, CS91a, CS96b, CS96a, CS97, CS98, CS91e, Com88b, Com87b, CHW99, GD98, Ghi89, Hun93, IL93, KSO95, ME95, SKB96, Smy95, DeR99, Ker98, Pab96, San91, San94b, Bra93a, Cam97, CMQ93, Com87a, Com88a, CS91c, Com91a, Com91b, Com92, CS93b, CS93a, Com94a, Com95c, IMI^{+}95, JL94, McC88, Mil91a, Mil95c, Ano90a]. Internetworks [DC90, ALS96, GS95c, B^{+}95, CX97, Mil91b, Mil93b]. Interoperability [Con87]. Interpersonal [Gac98]. interpretation [Bau95]. Interprocess [Ste99a]. InterRamp [Ano94-42]. intervention [Cla99, Int85]. interview [HH95a]. interworking [Ano92e]. Intonet [Edd94b, SG94b, SG94b, TI95, cYlC95]. Intranet [Ano96u, Ano97h, Bud96, Opp98, Mar99]. Intranet/Web [Ano96u]. Intranets [DH99, BBSH98]. Intrapersonal [Gac98]. introduce [Bar95b]. Introducing [IBM89c, Jaf94, JT94a, JT94b, JT94c, MD93, McK96a]. Introduction [Amo99, Ano93w, Ano94-43, Ano94-44, Ano94-45, Ano95-47, Ano96v, BR94b, BM94, Dav88, FL95b, Han96, Hed87a, Hed87b, How95a, IL93, Mil93d, PK94b, PG93, Res95b, Sil96, SC94a, SC95a, Sta89, WC95d, Zor95, AJR95, Acc95, Ack95b, Ano94w, BL95b, Bla95c, BD88, BD95b, Cha92, Cly93, Ent95, Fel94c, Joh94, Kw95, McN85, MS95, Mil92b, MW94e, Mil94e, Pol92, RN93, Ros94b, SB94, WC94, Wan96, RN97]. introductory [Ano94-27, Ano95q, SCJ96]. Intruders [Den90]. Intrusion [Amo99, Esc98]. inunga [Edd94b]. Invades [Sch93a]. Invasion [Chr98]. Inventing [Abb99]. inventor [BLF99]. Inverted [Saw93]. investigation [Hei95b]. investment [TRHR96]. Investor [Kic94]. Invisible [RG93, Bro91]. InvisibleWeb [Ano96u]. Invited [Ano98b, VT97]. involving [Ano94-55]. Iowa [Iow94]. IP [Ano89b, Ano93-37, Ari93, Arn94, Bar92b, B^{+}95, Bla92, Bla95b, Bla99a, Bla99f, Bre95a, Bro93, CMQ93, Com88a, CS91c, Com91a, CS91d, CS91b, CS91a, Com91b, Com92, CS93b, CS93a, Com94a, Com95c, CS96b, CS96a, CS97, CS98, CS91e, Com88b, CB91, Dav88, Dav98, Fei93, Fei96, Fel94c, FRB93, Gor95b, Hun92, Hun98, IBM92a, IBM92b, Int92a, Int93b, Int93d, Int93c, IBM93, Int93e, Int93f, KF93, Ker98, Kon91, LDG96, Los95, ML94, Mau99a, Mau99b, Mil92a, Mil96b, Mil97, MSD90, Mur98, PG96, Par94b, Per98, Pil90, PC93, Res95b, San91, San94b, SKB96, Sch91a, Sch91b, Siy94, Son92, Sta88, Ste94, Ste96b, Tho96c, Tri91, Tri93, WE93, Wil93, WS95, Zna94]. IP/IPng [SKB96]. IPCC [Wil94a]. IPL [Uni95c]. IPng [BM96b, SKB96, Tho96c]. IPSec [DH99]. IPv6 [Ano96c, Hui96, Hui99]. Irasuto [Edd95a]. IRC [Har95b, She95a]. Ireland [AP95b, Nug95, Smi94a]. IRIS [GMB93]. IRMSP [Bet87]. ISDN [Bla97b, Gil95a, Gil95b, Hop95, Kam93, Kes90, Kes93, Sta89, Sta92, Sta93b, Sta95a, SC88, SD96b]. ishi [PEH95]. isn’t [JEK92]. ISO [Ano92e]. ISO/CCITT [Ano92e]. ISOC [Soc95a]. Issue [IL93, Ano93-29]. Issued [Gar97]. Issues [A^{+}98, DM95, Gar97, Kah92, Rei93, Tri87, Cam96, FAR93a, Gre95a, Int96a, Liu96, Lun94, MMR93, MMR94, MBZ94, OO95, RA96b, SMLD95, TB90, VJ99]. italiane [Pas95]. itta [SG94b].

J [Ano99e]. Jack [Bro89]. jacketed [Xie94]. Jade [PR91a, PR91b, RP93]. Jan [Ano94-33]. JANET [CB91]. Japan [IDK98, MKT98, SI 93]. Japanese [TW95]. Java [ACM98, Ano96u, Ano96v, Hum98, TG95, CP98, LO99, LY97, Rus98, TK99, Tho96b, T^{+}98, ZTLV98]. JavaScript [BLB99]. JClass [Ano96v, Ano97h]. Jeopardy [Gar97]. Jet [Ano94-46]. Jet-setters [Ano94-46]. JetStor [Ano99e]. Jewish [Rom96, Ano90e]. job [AL96, BK96a, GRRM95, GG95a, GCS96, JN95, Ken95b, New95a, Ril95, RRO96]. jobs [The94b]. jobsurfing [Car96]. Joe [Ano99d]. Johnson [Mei97]. Join [Ano98b]. Joining [Bak93b]. Joint [Ano81, Pla94b, Ano95l]. Joke [Tem95]. jokes [May96]. Jonanthan [KB92]. Jose [Uni96a, Uni96b]. Journal [Jou97, Ano93n, EI 95, Bio96, Int96c, Dis93, Int96e, Ano93v]. Journalist [Reg93, Bri94a, RK95a]. Journals [Ano93j, Mag94, KKO94, Ano94-41, HOO95]. journey [Ano94-40]. Joy [Bai94]. Jr [ACM99a]. JTR [MNT93]. ju [cCrmdh92, Edd95b, cH95]. Judaism [Ano96k]. Jughead [Har93a]. July [GMA^{+}98, GV90a, ILSB94, KW98, KSW99, VT97, Ano95l, Sav93, Sto95c]. Jump [Arm94b, Wei95b]. Jump-Start [Arm94b]. June [ACM96, LM98, Pla94b, Ano90d, McN95, Str94, NAS90]. jungle [BB95]. junior [Chi95]. junk [RAWS95]. junkies [Has96]. juristen [vdW95]. just [Dre92, Dre93a, Dre93b, Dre95, JEK92, Nov94]. Justice [LR96, Hum94, Luc94, LR96]. JWA [Don95].

K-12 [Bla93c, Gol93]. K12Link [Zil95a]. kakumei [Oma95]. kalastusta [Hin93]. kan [kTlC94]. Kansas [MP95]. kao [HSskw95]. kayo [LB95a]. Keep [Car99b]. Keeping [Far93b, WC94]. Kerberos [Tun99]. Kermit [dG93]. kernel [Pet89]. Key [BH93, LS89, Ass94b, Ano94e, Mor95d]. keystroke [Pot94a]. keystroke-by-keystroke [Pot94a]. Kid [Jor94]. Kids [Ano94-28, Tur95c, Wie96, ANT96, FKA95, PM95b, Pol96]. kidstuff [Arm96]. kill [New95b]. King [ACM99a]. Kiosk [McC97, Rak95, RR95a]. Kit [Ano96u, Eng93c, ED94, Eng94a, FHS94, GM94a, HSF94, Har94g, LaQ94b, Sam94, Ame95a, Ano95-68, App95, Bri95b, Bea95, Ced95, EE95a, Eng94b, Eng95b, ELS95, ELO95, FD95, FH94a, FH94b, Int92a, Int93b, Int93d, Int93e, Int93f, Joh96b, LR93b, Pfa95a, Pfa96b, RRF95, SS94c, Sei94c, You94, LaM97]. kkaji [cYlC95]. Know [Che97, Com97, Mau99b, BM96a, Com95a, Ent95, Kur96a, Thi93, Wn95]. knowing [MW92]. Knowledge [BWWF99, DS97, Sav96, Sky99, Ano95-28, Dra95]. Kochmer [KB92]. kokusai [Rik94].

Kommunikation [MW94a, MW94b]. Kompyuserv [Mal95a]. Kong [HL99]. kooks [Mam96]. kot [SG94b]. Krol [Ano92i]. kua [LB95d]. kuikku [Hof94a].

L [Bro89, Ano93a, Fee93c]. la-mahshev [PEH95]. la-mishtamesh [Mal95a]. laboratory [Mil95e]. LAIR [PN95]. LambdaMOO [Rya95i]. LAN [Cam97, Cly88, DC94, GD98, Hew86, LAN86, Mad96, Mil89, Mil91a, Mil93a, Mil95c, ME97, Net93, Sta93b, War99, Yan88, Ano90a]. LAN-ATM [Cam97]. LAN-Internet [War99]. Lan-to-LAN [Ano90a]. LAN/WAN [Cly88]. land [GCS96, Mle95]. land-grant [Mle95]. landscape [BBK^{+}98]. Lane [Gun94]. Language [Lam96, KTS94, Mar94, War95]. languages [BCB99]. LANs [DC90, KKC98, Kon91, Lar94b, Mad89, ME95, ME97, Sei98, The94c]. Lansing [Str95]. Large [Chr98, RG93, HCBO99]. Largest [Fle94, Mar99]. Last [Ell99]. Last-Mile [Ell99]. Late [HL96]. Latin [WLM90, Mol93a]. Latina [WLM90]. Launches [Gar97]. Law [Ano94a, Ano96n, Ste91a, Ste92, Ano94-61, Ano95s, Ano95u, Int95e, Ano95-32, Int96f, Ano96o, Anoxxb, Bro96, Fol95b, Fol95a, Gro95b, Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b, Kur96b, Luc94, Max95b, Mil93c, RZ96, RWD95, RJV95c, RJV95d, Sme96, Wri95a, Mil94e]. law-related [Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b]. lawless [Ano96n]. LawLink [Zil95b]. Lawoffice.com [Ano96o]. lawyer [All95, Int95f, Ano95-48, Eis95, RA96b]. Lawyers [Ano95-49, Car96, Ker95, Ano95-36, Ano95-48, Car95a, Uni95d]. Layer [BDPS85, IEE90, MD99, ANS92]. layers [Bra89b]. Layout [BDKK96]. LB452 [Ano93h]. LC [Not93a]. LCN [IEE97]. LCN’98 [IEE98]. lead [ZDE93]. leaf [dAMW97]. Learn [Ano95-50, Int95k, Lea95b, Win96a, Bal93b, Int95q]. Learned [Gla99, PCD92, Mou96]. Learning [Ack95b, Bal93b, Bur96, HWV96, PE95, Rob94, Wol96, DOO95, Gru94, Hal99, KW98, MS95, Mil94b, ODO95, Wam95, War95]. leased [SC95d]. Left [Sch95g]. legacy [LL99]. Legal [A^{+}98, Leg94a, Leg95a, Ano96q, Ano96p, Bla93c, Hee94a, Hee94b, Hee94c, Jon94b, Ame95a, Ano94-61, Ano95-38, The96c, Net96b, Bla93a, Bro96, BS94, Coo96, Hee93a, Hee93b, Hee94d, Hee95a, Hee95b, Ker95, KRS94, KRK95, Kur96b, Mac95b, PN95, RD95, RCL96, RA96b, SLJ94, Sme96]. LEGI [Leg93]. Legislative [Ano96w, Leg94b, Cal94b, Wil95a]. Leksikon [PEH95]. lesson [Bar95b, DG94, Int93a]. Lessons [BB93, Gla99, KG98, PCD92, Peh99, Cro96b, KW94, Mou96]. Let [SG93b]. letter [The93, Ano94-52]. Level [MK98b, MS98, MK98c, MK98d, MK98e, MK98f, Bla99a, Eas94, Sch91a, Sch91b]. levels [dVdVI98]. liability [Wri95a]. LIBCAT [NM94]. Liberian [Sch95a]. Librarian [Ano93p, E^{+}93, Ano91, Int93h, Eng93b, Eng95a, NDM98, Sch96a, TB93]. Librarians [Ano93s, Kin94, LT93, MD93, Tro94, Ben94, Ben95b, Chi94, Cly93, LT92, Max95b, MW92, Mil94a, Mil95a, Mil95b, Mil96a, Mil93c, MSB^{+}92, Mor94, Spe93, Til95]. Libraries [Ass94a, Ano95-69, Ano96v, Cay93, FMS98, GMB93, IE98, MMR93, MMR94, McK94a, SEDK94, Tex95, Wal94a, Ano94-54, Ano95-29, FAR93a, Gre95a, KP95, LL99, MRTR95, McC92, McC94a, McC94b, MBZ94, MBB95, NYS94, Noo91b, PM94, SH95b, VJ95, Ano93-41]. Library [Ano93m, Ano93z, Ano94q, Ano94-56, Ano95f, Ano95-61, Ano96u, Bla98b, Bla99b, Bre93, Cam95b, DJB93, FE91, FE92, Gai93, KKB95, Nic93, OD95, Par90, RBK^{+}94, Sti94a, Wig95a, ZM93, Ame95b, Ano90c, Ano94x, Ano94-55, Bir90, Bir93, BW94a, BW94b, Bre94, BS94, Dil93, Hic94, Int89a, LT92, Liu95, Liu96, NM94, SLJ94, Sim95b, SM93, SL91, SLR91, SL92, Sti94b, TA96, TPH96, Uni95c, Uta94, VE96, Van95, Wan94, Ano95x, Ano95-33, MBC^{+}96]. library-wide [BS94, SLJ94]. libro [Fre95]. Life [Ano92a, LaS95a, LaS95b, Pet93b, Sav94, Tur95a, ZD 95, Yah96, GH93, GH94a, GH94b, GH94c, MM95, SP94b]. light [Int95g]. Lighthouse [Hal99]. Lightweight [HS98]. ligne [Cle93b]. Lincoln [Ano93-41, Ano95-69]. L’indispensable [Pla95]. Line [Hub93, Ano94-40, Ano95-48, Ano95-49, Ano95-58, Bar92a, CS94a, CS95, Cot95, Ehn94, Gen95, Her95b, Her95c, Int95q, LR95b, MBC^{+}96, Pot93, Pot94a, Pot94b, RCL96, Sta98a, SC95d, Whi95, And93, SCS99]. LINK [Ano96v, Jos95a, Jos95b]. linking [Lan94]. links [Ano93c]. LinkScan [Ano97h]. LinkWay [WMS96]. Linux [Ano95-55, Ano96b, Ano96s, Ano96t, Ano96u, Ano96v, Ano97h, Ano99d, Ano99e, HE95a, Le 96, Ano94k, Ano96s, Ano96t, Ano97j, Ano97e, Ano97f, DN99, Don96, McC97, Och95, Ogb98, VRFJ97]. LinuxCare [Ano99d]. Linz [VT97]. Liquid [Ano96u]. List [Ano92f, Ano95-51, Bar92a, Caw95, Ano92b, Bar93, Bos94, Boy96, Ehn94, Mas89, Han92, Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b, Hol94, JO95, KK96a, MP95]. list-server [Ehn94]. Listing [Hum94]. Lists [Ben95a, HN92a, HN93, Mag94, KKO94, All94, Ano92b, Ano92f, Ano94-41, Mar95c, Bos94, Ehn94, Fel93, Fel94b, FK95a, HN92b, Neo95]. Listserv [Fol95b, Gla94]. listServers [Dar93a]. listservs [Ale95, Fol95b]. LITA [MW92, MSB^{+}92]. Lite [SS93]. Literacy [Hop95, Ela94, EHPT95]. Literature [Mog89, Ano94d, Bri94b, Bro94a, Mog88]. litigation [Int96g]. little [Kur96a]. live [RML^{+}94]. LiveTable [Ano96v]. living [Shi96]. Livre [GS95b, Kap96]. LJ [PC95]. LLC [Ano99e]. Local [BDPS85, Bla98a, Dig82, DM95, IEE94, IEE95, IEE97, IEE98, LAN86, McN85, Sta90, Sta93a, Sta94a, ANS92, New95b, Pub95]. Locating [Hen82, Mac95b, MRJ96]. Location [Cho95, Har93a]. Log [SW97]. logged [JV94]. logged-on [JV94]. logging [Mil94f]. Logic [Leo98c]. logical [Bir93]. login [Kon82]. logros [Ano95b]. longitudinal [LMG95]. look [Ano95v, Coo96, GK96]. Looking [Gol93, HH95b, HB95, Hes95, HH98]. lookups [SV99]. loss [JP94]. lost [JV94]. Lotus [DO96, SY95]. Louis [Mil92b, Mil94e]. Louisiana [Ben92]. Low [Not93b, Ros84, Koh97, Mot94, Sch96a]. low-cost [Sch96a]. low-earth-orbit [Koh97]. Low-end [Not93b]. Lowell [IEE98]. Ltd [Ano99e]. LTER [FB91]. lu [nChCiY94, cH95, kTlC95]. lurking [Ano95-65]. Luther [ACM99a]. Lycos [ML95]. LynxArray [Ano99d]. LynxNSS [Ano99d].

M [Ano99d, Ano93b]. M-Cluster [Ano99d]. M-Z [Ano93b]. MA [Ano81]. Mac [BT96, PB95, Sta95f, SS96, BDPS85, Bra94b, Fra93b, Gia95c, HB98a, Val93a]. Mac-edition [Bra94b]. Mach [Spe87]. Machine [LY97]. machines [Kan94b]. Macintosh [DK94b, DS95, Eng94b, Eng95b, Eng95c, HSF94, MPG94a, Bra93a, Bri94a, Eng93c, ED94, FAW95, HKB95a, HH95b, Lar94a, LE95, Lic94a, Lic95a, Pfa95d, Pit95b, She86, SG95, Val93a]. Macs [Sei94c, Sei94b, SP95a, Syd96, Sei94a]. Macsyma [Ano99e]. made [McB95, Slu95]. Madefast [CTG96]. Madness [WS94]. madrikh [HY94]. Mae [Fon94]. Magazine [Dat74, LAN86, Int94h, Inf95b, Net95a, Int96h, Com95d, Eve95]. magia [Wya95a].

Magical [GL94]. Magnate [Ano99d]. mahshev [PEH95]. Mail [FA89, FA93, FA94, Ken90, Ken93, Mol93b, VJ98, Ano94m, Asa91, BL95a, Bau93, Bos94, BW94a, BWC95a, Buc95a, CP82, Ehn94, Fel93, Fel94b, FK95a, Hf96, Hel94, Int95b, JEK92, Kw95, Lar94a, Mar96a, Ban94, Mul99, Nat93a, PS92, Pod95, Ros93, RAWS95, Rus94, SB95, BAS95, Wis90, Yu93, AH94, Bac95a, BL95d, God94a, Kol96, McF95, VH94d, Als95, Ano95-68, DC94, Gar94, Inn96, JEK92, Lev96, Pos95, QCM94a, Shi94, SG93b, TR94a, War95, Wri95a]. Mailing [Caw95, HN92a, HN93, Ano92f, Bos94, Ehn94, Fel93, Fel94b, FK95a, HN92b, Neo95]. Main [Ano99d]. Mainframe [Bra93a, EE99]. mainstream [Leo94a]. Maintain [Ste97]. maintaining [FK99, NM98]. maintenance [IBM89d]. major [Hof96b]. Majordomo [Caw95]. Majors [CES^{+}99]. make [BA95, CS94a, Int95g, Pot93, Pot94b]. Makers [Com84]. Making [Ano95-52, Ayr94, BW97a, CDEF94, Clo93, Coo93a, Eag95a, Fon94, Fow99, GG95b, Bab94, LR95b, McC95, RK96]. malcontents [Mad98]. Man [Ano93-43, Bau93, Sta93b]. manage [Ber96a, Net96a, VJ95]. Management [Int76a, Ano88a, Ano96u, Ano99e, CKPT99, GG98, Har94a, Har99a, Heg99, Int76b, IMI^{+}95, LS89, Moh99a, PT99, Sav96, Slo94, Sta93c, Sti94a, Ter99, Vac99, Zel99, Ano74, Ano81, Ano92e, D^{+}86, FBGP98, Gha92, Li90, Maz90, Nat95, Pub94, Rii82, Ros94b, Rot95a, SHB95, Sti94b, Tri87, TB90, Uni89, UU89, VVB85, ZP98]. Manager [Dav94, Eag95a, BN90, Mai96, Per95b]. Managing [Bla98a, CP99, LPJ^{+}94, Mil93b, OT88, Ste91b, TS88, TO92, Als95, Res95a, Rya95f, Rya95g, Rya95h, Rya96b]. Manes [Ano94-48]. manifesto [Sal95d]. manipulation [Bri94a]. manos [Car94a, Car94b]. MANs [KSO95]. Manual [Gru95, Hew86, MOD93a, Ano95-62, Of 95, CB95b, Cly93, Gru94, HS94c, Hof95g, IBM88, IBM89a, IBM89b, KKB95, Max95e, Ols95, Owe93, Pre97c, PW95b, TOL94b, Thi94, Tri91, Tri93, Ubi95, WE96, Yas95]. manuals [Bri95c]. Manufacturing [GS99, Leo98e, Mat96b]. Map [RBM94a, RBM94b, Ano94-57, Goo96, Pil91]. Mapping [BC95a, BAG94]. maps [Nof96a]. March [Ass95b, Ano94-32, Can94, DLC92b, IDK98, LG93, MKT98]. Market [Les95, Int95a, Jon94a, Ngu95, Wes97, HH95a]. Marketing [Mar95c, But95c, EE95b, Mat95a, Mat96a, RSD98, Ano94l, Int94j, Int95h, BD95a, CS94a, EH95, Jan95, Ker95, LaS95a, LaS95b, LR95b, LR95c, PE99, RAWS95, Rot95a, Ste95a, SB94, Uni95b, WLW98]. Marketplace [Val93b, Coo96]. marketplaces [Bak98]. Markets [BL95d, HWV96, Int95m]. marking [FK99]. Markup [Lam96]. Marriott [Lin93c]. Martin [ACM99a]. Maryland [LG93]. Mass [Fry95, Tho99, ER98, OM95]. mass-production [ER98]. Massachusetts [IEE98]. Massively [KP93]. Master [Mas89, Ano97b, Lea95b]. Mastering [Bre98, CM95, Par95, Ros96]. Masterlist [Ano89c, Mas94, Hum94]. Mastery [BS93b, BS93a]. Mateo [JG91]. Materials [LG95, Ano95-68, JE96, MBC^{+}96, Str94, TW95]. math [Bri95d]. mathematics [Gla94, Hal95b, O’H93]. Matrix [Qua90, Sav95]. Matter [Saw93]. MAU [IEE90]. May [Ano95a, Ano96x, BCB99, Cai94, Gro90a, Smi93c, VVB85]. Maze [Sto89]. Mb [IEE90]. Mb/s [IEE90]. MBone [Kum96, SRL96]. MCC [Ano93c]. McGraw [Dat74, WE96]. McGraw-Hill [Dat74, WE96]. McKinley [Max95d]. MCSE [Byr99, Del98, Gon99, Keo99, Vac99]. MDA [FS95]. mean [KKM^{+}98]. Means [DN99, Ogb98, YY98, Rot96]. Measured [BMK88, Mil90]. measurement [Cer91, EHS92, Sch91a, Sch91b]. Measurements [Cac89]. measuring [PKL90]. Mechanics [Mac93a]. Mechanism [DGH93, Jon94a]. Mechanisms [Cul95, Sha91a]. Mecklermedia [Mec95b, Bw96, HE95b, Mec96, Kan95b, Kan95c, Sta95b, TG95]. Med [Med96]. meda [Mal95a]. Media [Jai94, MKBB95, Win98, Ame95b, PA95, Sim95b, VE96]. Mediated [WYLZ99, Dec95, Dil93, DJB93]. Mediators [AAB^{+}98]. Medical [MB93, Chi94, Cim96, HM96]. medicina [Par96a]. Medicine [DS97, McK96a, Int96i, Lib95, NC95a, NC95b]. Medium [Fry95, IEE90, Sav95, AADF96, Gon95, OM95]. Meeting [Ano91, Ass95a, Ame95b, Ano90d, Ano94a, Don95, NAS90, Bla93b]. meets [GP95a, GP96a, New95b]. Meganet [Kon91]. meil [cYlC95]. meilleurs [RdG95]. member [Win93]. members [BU94a, Slu95]. Membership [FHS94, HSF94, FH94a, FH94b]. Memory [Ano99d, HPS95, Sfe82]. men [Edd95b, cH95]. mercadeo [CS95]. Merchant [Ano99e]. Merit [Mer92, Mer93, Kel92a]. mes [Bel95]. message [Ros93]. messages [Als95, IBM91a]. Messaging [RS98, Hor95]. meta [GCGMP97]. meta-searching [GCGMP97]. MetaCard [Ano99e]. Metadata [BB99]. Metadata-Based [BB99]. Metaphor [BW97a]. Metaphors [Ste96a]. Metasearch [DH97]. Method [IEE90, ANS92, TAH99]. Méthodes [Ano69]. methodologies [Rud95]. Methodology [Fou98, PKL90]. Methods [Hor98, Ano69, CN95a, Gha92, Gon95, MA98, Wan94]. metrics [MA98]. metro [Mai95]. Metropole [Int76a, Int76b]. Metropolitan [KT92, Sta93a]. Mexico [Don95]. Michigan [Ano94q, Ano95x]. micro [Bre95a]. micro-based [Bre95a]. Microsoft [KH95, Ano96r, Byr99, Cat96, DK94a, GS95a, Gon99, Hud94, Ken95h, Nem96, Pfa95a, Pfa96b, Pil90, Roa95b, Ros96, Vac99, Win96a]. middle [Pie95, Int93a, Woo95b, GP95a, GP96a, Opp95]. MIDI [IRC95]. might [Thi93]. migrating [KKC98]. Migration [LDG96]. migratory [Hen82]. Mile [Ell99]. military [DG97, U.S83]. millennium [AADF96]. millions [GG96b, Mat95a]. mind [Ano95e]. mind/body [Ano95e]. Minding [Ros94a]. mine [US 92, US 93]. Mining [Har93a, RNW^{+}94, WSA^{+}94, DG94]. Minitel [KW94]. Minneapolis [IEE94, IEE95, IEE97, Mai95]. Minneapolis/St. [Mai95]. Minnesota [IEE94, IEE95, IEE97, Gra95, Wil94b]. Minute [Ano97j, Ken94a, Gri95, HE95b, Kan95b, Kan95c, Nem96, Roa95a, Roa95b, TG95]. minutes [Car95b, MPG94b, MPG94a]. MIS [Mai96, Wig95a]. mishtamesh [Mal95a]. Misinformation [Vie93]. missing [CDEF94]. Missouri [Sta94b]. Mitigating [KG99a]. Mobile [Bla96a, Bla99e, CKMV99, Eas94, KG99b, LDG96, Per98, VT97, Xie94, PMJ94, VJ99]. Mobility [MDW99, Cho95, KW98]. MOD [MOD93a, MOD93b]. MOD-TAP [MOD93a, MOD93b]. moda [Ban95]. Model [CC83, MMLP97, Cra92, Gil95a, Gil95b]. Models [O’R96, MBB95, MA98, ZCD97]. Modem [Ban95, Mot94]. Modems [Wil98]. Modern [CFL99, RCL96]. moderne [AB94]. modun [SG94b]. molecular [SMM96]. Monde [KCG95]. Mondo [Rim94, Sal95d]. Monetary [Gar97]. Money [GG95b, Sny96, Fis95b, LR95b, LL95b, LL96]. Monitor [Del93, Val93a]. Monterey [NDM98, USE97]. monthly [Dat94, Int84]. Montréal [ACM96]. MOO [Car95c].

MOOs [She95a]. Moravec [MSB^{+}92]. Morocco [CL98]. mortali [Pet95]. MOS [VT97]. Mosaic [AH95a, Ano94w, Bow95, Bro94b, Bro95b, DK94b, DKF94, Gau94, Gil94d, Gil95c, Pfa94b, Pik94, Rau95, SS94a, TK95a, Bra94b, BC95d, Chi95, DK94a, Eag95b, EE95c, Gil94b, HA95, KLW95, MWB95, Man95b, Rak95, Val93a, VSW94]. Mosaic-Based [Rak95]. most [Hf96, Hub96, Jam96a, Jam96b, Mar99, McQ94a, Bab94, Woo95a]. Mousetrap [Lam95]. movement [Ste91a]. movie [Nram94]. Moving [Coh99, Tex95]. MP [Ano97h]. MP-EG1A [Ano97h]. Mr [Vie93]. Mr. [Gol93]. MS [Tri91, Tri93]. MS-DOS [Tri91, Tri93]. MSHA [McA95]. MSN [Cob96]. MUD [Car95c]. MUDS [BC95b, SR95a, She95a]. mujakchong [cYlC95]. mujer [Ano95b]. mukete [SI 93]. Multi [COZ99, Sin99b, BH93]. Multi-agent [COZ99]. Multi-Platform [BH93]. Multi-processing [Sin99b]. Multicast [DC90, Mau99a, Mil99b, BHO^{+}99, Bla99a]. multicasting [SRL96]. Multicomputer [ASB97]. multidatabase [KTS94]. multidomain [EST93d]. multilayer [Bla99a]. Multilevel [MMLP97]. Multimedia [CKPT99, CHW99, Des94, EFIJ99, Fur99, IMI^{+}95, Jac94, MWB95, Nie95, Wes98, Ben95c, Dea96, EE95b, HCBO99, Hor95, Ins96, Joh96b, Kum96, McM94, NGP^{+}99, RT98, Rus98, TA96]. multiparty [HSK98]. Multiplatform [Ano96v]. Multiple [IEE90, ANS92, GGK^{+}99]. Multipoint [Ano94-49, MP89]. Multipoint/Internet [Ano94-49]. Multiport [Sfe82]. Multipurpose [Mol93b]. Multisegment [IEE90]. mundial [Ano95b]. mundiales [LR94]. mundo [Ano95b, Bog94, Car94a, Car94b, JE95, Kro95a]. Museum [Rak95]. museums [FS95]. MUSHes [Car95c]. Music [Tro94, Ano94-38, Ano95e, Ber95c, Bry95, Ehn94, FMB95, GP96b]. Musical [GL94]. musicians [Ehn94]. musicology [Ehn94]. Must [RG93]. Mutual [SCT81]. MVS [IBM89f, IBM89d, IBM89e, IBM92a, IBM92b]. myriad [Ano95-61]. mysteries [Mam96]. Mystery [BS93b, BS93a]. mystic [Str92, Str93a]. Myths [BMK88, Ste96a].

’n [And92, Hud94, RL95a]. NAFTA [Don95]. Naked [Jan89, Moo95]. Name [Moc94, Tan93, BPR88, Int93c, Tri87, TB90, Zho84]. Naming [BP91, Pet86b, Pet88, Bow90, Pet87]. namyon [SG94b]. nan [cH95]. nao [Edd95b]. NASA [Ano90d, DGH93, HJ93, JG91, LG93, Nat93a, NC90]. Nashville [Mas97]. National [Int76a, Ano95f, GBJ94, Int76b, Kah92, KW93, LG93, Par90, U.S93, Wei93, Ass95b, Aga99, Bro89, Ano94-50, Bla93b, Cat89, Gou90, LR96, MBDR91, Wal93]. Nations [Ano94-27]. nationwide [Fol95a]. native [BEA98]. navegando [Ano95b]. navegante [GVJ95]. naviganti [BZ95]. navigate [Yas95]. Navigating [Ano95-53, BB95, Bow95, Bra94b, DS95, HKB95a, HKB95b, HK95a, Joh95, Kel92b, Kel92a, PMB^{+}92, Ric94b, SG93c, SG94c, SG94a, SGM95, Sne95c, Tat95a, Tat95b, Tys95, VH94c, Wya94a, DG94, Ens95, RO94a, RO94b, WK95]. navigation [Pot95, Ric94a]. Navigational [Ano93-42]. Navigator [Gil93b, Gil93a, Gil94c, Gil94d, Gil95d, Jam96b, Net95c, TK96, Gil95c, Gil96, Jam96a, Neo95, Pfa95d, Ano93q, Jam96a, Kro95c, Net94c, Pfa95b, Pfa95c, Pfa95e, Pfa96c]. NC [Ano99b]. NCSA [Val93a]. Near [Fra93c]. Near-Term [Fra93c]. nearly [LJN94, Lev95, NN94]. Nebraska [Ano93-41, Ano93h, Ano95-69, Hei95b]. Nebraska-Lincoln [Ano93-41, Ano95-69]. Need [Che97, Com97, Gar95a, Ano95-61, Con95a, Com95a, HH98, Kur96a, Sch95e, Uni89, UU89, Wn95]. Needs [Mau99b, Jon94a, KLW95, San94a]. Negotiation [WYLZ99]. Neighborhood [Bar95a]. Nel [Gia94, Sal95d]. nell’era [Gio95]. NEMICRO [BH94b]. ’net [Pub95, Net95a, Bak93a, Bak93b, Bak93c, Col99, Eng96a, GRRM95, Kur96a, Mal95c, Man95c, PM95b, RBM94a, RNW^{+}94, Sal95c, Sch95e, Wam95, WSA^{+}94, Woo94, AUA94, Arn95, Bir90, Buy96, Cob96, Goo96, Hol94, LR96, Net94a, Ric95a, Woo95a, RBM94b, BW97a, FTP95b]. Net-Centric [Col99]. Net.10 [Net96b]. Net.Commerce [Shu99]. Net.profit [Mal95b]. net.sex [RT95]. Net.speak [FP94, Fah94]. netchat [Mal94]. Netcruiser [Net95b, DeR96]. Netfind [SP94a]. netgames [Wol94, Gre96]. NetGuide [Arc95, Net94b]. nethead [Her95b, Her95c]. Netherlands [VVB85]. Netiquette [Fol95b, She94]. Netizens [HH97, Laz97]. Netjobs [GCS96]. netmoney [Wol95]. Netmusic [Gre95b]. Netnews [Fri94, Gar94, Har95c]. NetPAC [Dat96]. NetPages [Com94b]. NetPower [Per93]. NetPractice [Mil96c]. Nets [Not93c, Not93a, Not93b, Not94c, RS93b, JEK92, Lof95, Tom93]. NetScape [Wen95, Bas99, DG95, Dup95, HKB95a, HKB95b, HK95a, HK95b, Jam96a, Jam96b, Kro95c, Min95, Net94c, Pfa95b, Pfa95c, Pfa95d, Pfa95e, Pfa96c, TK95b, TK96]. Netstation [VFH98a, VFH98b]. nettowaku [Rik94]. NetTracker [Ano99e]. Netware [Net95d, Gas96, SFL95, Siy94, The94c]. Netware/Internet [Net95d]. Netwise [Net94d]. Network [ACM98, Afs95, Ahu96, AIS^{+}99, Ano69, Ano94-53, Ano95f, Ano99d, Bir94, Blo94, Bre98, Cah98, Cam97, CK74, Col99, Com84, Dig82, Dig90a, EAR94, Fle94, Gil93b, Gil93a, Gil94c, Gre99, Har94a, Hel91a, Hun92, HGKM98, Hun98, Iow94, Jan89, Kah92, KPS95, KKB95, KSM98, LAN86, Leo98g, Leo98f, Leo98b, Mas93, MSK99, MH94, Mil91a, Mil95c, Net90, Net93, Par90, PT99, Rag93, RG93, Rya93, She85, She95b, SH95a, Slo94, SS99, Sta93c, Sta95c, Ste90, Ste98b, Ste99a, Tan93, Ter99, TR94b, Tun99, Wal93, WE93, Wu97, ZM93, Zel99, AA94, BPS99, Ber96a, BP93, Bla99a, Cen95, Con89, Esc98, FB91, GS95a, Haa85, IBM89a, Int92a, Kel92a, Kes97, Leo98a, Leo98d, Leo98e, Leo98c, Lin93a, Mak93, Mar99, Mit91, Mit92, Moj93, Nat94a, Owe93, PG96, PS92]. network [Sch92a, She86, She77, Sta99a, TB93, Tri91, Tri93, TJ97, Uta94, ZP98, Ano92a, Gou90, MBDR91, Mil90, Net86, PM94, RBK^{+}94]. Network-conscious [AIS^{+}99]. Network-Management [Sta93c]. Networked [Ano95-54, Bre93, GMB93, Har93a, Heg99, Mes99, RT98, SZ99, WA98, Ano95-68, KT95, MP91, Str92, Str93a]. Networking [Ani93a, Ano93r, Ano93s, Ano93x, Ano93-39, Ano96v, Bee93, Cas93, Com97, Dat74, Dha94, Fla90, Fla91, Har99b, HA94, Jai94, Keo99, KW89, LR93c, LaQ94b, LS93a, Lei93, ML94, MD91, Nat92, Ass99, Pad85, Pad93, Par94c, Pil91, PC93, Pla94b, Sav96, Sky99, Sta93b, Ste98b, TL93, NAS92, Val93b, X/O88, Ele91, Res92, Bar95b, Com95a, Ano93n, Kes97, LR93a, LR93b, LaQ94a, LDH95, LS93b, Mac93b, Mil96c, Mil99b, New94b, New95c, New96, RHK95, SEDK94]. Networks [BDPS85, Bla95a, Bla96a, Bla97b, BW97b, Bla98c, Bla98a, Bla98f, Bla99d, Bla99e, Blo94, Bra93a, CKMV99, CD99, Cos94a, CP99, Dig93, Dav98, DM95, Dev96, DH99, Far93b, Fow99, FA89, FA93, FA94, GKM^{+}99, GG98, HHS94, Hel98, IEE90, IEE94, IEE95, IEE97, IEE98, Ibe97, Jar95, Kin95a, LaQ89, Mac93a, Mal90, McM94, McN85, Mer87, Nil93, OD98, Pea98, Rog97, SWE98, Slo93, Sta90, Sta93a, Sta94a, Sto95c, Swi96, Tan88, Tan96, Zor95, ANS92, Abb94, Bau93, Bla98e, Bla99a, Bro89, Bro93, Coo93b, GH93, GH94a, GH94b, GH94c, Gro95a, Har92, HRB92, HL89a, HL89b, Kes97, KT92, KW94, LF95, Lun94, Mak93, McC88, MZ93, Nat94a, Opp96, QCM94a, SEDK94, TJ97, TAH99, War95, Wis90]. Networld [Rot96]. Netzen [RS93b]. Netzwerktechniken [Ano69]. Neural [Dev96, Leo98g, Leo98a, Leo98f, Leo98d, Leo98e, Leo98c, Leo98b, Mac93a, Mas93, OD98, Rog97, Swi96, Zor95]. Neuron [MDZ92]. newbies [Wil96a]. newest [Woo95a]. newly [RMU96]. News [Ano93e, Ano95-56, Bla93b, Gar97, Has96, Soc92a, Soc92b, Kw95, Ric95b, SG98, Net94d, Int95i, Int96g, Com95d, Nor95, Int92e, Ano97i]. newsgroup [FK95b]. Newsgroups [Pfa95f, Bos94, Ehn94, FK95a, Kug95]. Newsletter [Ano93y, Sma95, The96a, Edu95, Hum76, Her94, Mat93]. Newsletters [Hes95, Mag94, KKO94, Ano94-41, HH95b, HB95]. Newspaper [Com84, Ano96h, ST95]. newspapers [PE99, Smi95a]. Newsroom [Com95d, Int95j]. Newton [New94b, New95c, New96]. Newtons [Ano94g]. NExS [Ano96v]. next [BPR88, BM96b, DFG^{+}98, HMG^{+}95, Ken94e, Sal95d, Tho96c, Tür94, Fos91]. NFS [Her93, San91, San94b, Siy94, Ste91b]. NFSNET [Cat89]. NGOs [IV95]. ni [SI 93]. Nineteenth [Sta98a]. Ninteenth [IC90]. Ninth [Uni96a, Uni96b]. NIS [Int95b, Ste91b, Upj95]. NIST [Wes95]. NJ [Lin93c]. NLM [Nat94b]. NNStat [BD88]. NNTP [Ste96b]. No [Hub93, Ker98]. no-nonsense [Ker98]. non [Coh99, Ste91a]. non-governmental [Ste91a]. non-U.S.-centric [Coh99]. nonprofit [SMLD95, Zef96]. nonsense [Ker98]. norte [Ano95r]. North [WS80]. Northcoast [Nor95]. Northeast [Ben92]. NorthWestNet [KHJG91, KHJ91, KB92, KHJ92, Koc93, RK95b]. Norton [RN97]. notebook [Ano96a, CC93]. Notes [DO96, Gaa76, Vie93, Sar94]. notice [Ano94-55]. NOV*IX [Fir95]. Novel [ZDE93]. Novell [Mal90]. November [ACM99a, DLC92a, FS95, GV90b, IEE97, Mei97, NDM98, Ano91, Bro89]. Novice [Ano97j, Com94c, MW94e, Pee96a]. NREN [Ano93a, Ano93b, Ano93v, Ano93x, Coo92, Lin93a, MBDR91, McC92, MMR93, MMR94, Par90, Wei93]. NREN/ [Ano93x]. NSF [Luk96, Mil90]. NSFnet [Coo93a]. NT [HAS95, Bud96, Buy96, CSJ98, DeR99, New96]. nugu [SG94b]. nuls [LB94c]. Numeric [Bar92b]. nurses [Edw95]. nursing [Ano94d]. NUSIRG [KHJG91, KHJ91, KB92, KHJ92]. Nutshell [AL92, Nie98, Alb93, AL94]. NYSERNet [NYS92, PN91, Per91c, Per92d, Per92e, PM94]. Nyumon [Nih95, Edd95a, Nag94].

O [Ano96v]. O2 [Ano99e]. OARnet [KB92]. Object [Ano99e, HS98, Rog97, VT97]. Object-Oriented [HS98, Rog97]. Objects [BN96, BP91, Hen82, VJ99]. ObjectSpace [Ano96u, Ano96v, Ano96u]. OCLC [Ano92g, Ano93j, Ano93z, Dil93, LeB93]. Oct [Ano74]. October [Ano94y, Cim96, GV90b, IEE94, IEE95, IEE98, Mas97, Sau92]. ODBC [Ano96v]. odyssey [Ano95-28]. Off [GGA^{+}97, GAM^{+}97, RAWS95, Wal95b]. Offbeat [WS94]. offered [Ano93z]. Offers [Ano92g, Ano93-35]. Office [Ano96s, Ano95-39]. Official [BCW95, Jam96a, Jam96b, MG94a, Zim95a, Mec95b, BR94a, Bw96, HE95b, Mec96, Kan95b, Kan95c, MG94b, New94b, New95c, New96, Sta95b, TG95]. Offline [Ano96u]. Oil [Cro96c, Sto95b, Uni93a, Uni93b, SR95b]. on-a-disk [Inf95c]. on-disk [JE96]. On-line [Ehn94, Gen95, Sta98a, Ano94-40, Ano95-48, Ano95-49, Ano95-58, Bar92a, CS94a, CS95, Cot95, Her95b, Her95c, Int95q, LR95b, MBC^{+}96, Pot93, Pot94a, Pot94b, RCL96, Whi95, And93]. On-Net [FTP95b]. On-ramp [HSF94, FHS94]. On-ramps [RHS95]. on/off [Wal95b]. ONC [San94b]. ONC/NFS [San94b]. One [Ano94y, Ano94-32, Ano95l, Cul95, GM94a, Bar95b, TMF97, Wen94, Val93a]. One-Stop [Val93a]. One-Time [Cul95]. Online [Abo93, AP95a, Bal93b, BGS94, BGS95, Bla93b, Byr95, Cro94b, Edm95, Gol93, Hes95, Ive95, Jan95, Kin95b, LH93b, Les95, Ano93-29, Per93, dPW94, Rit92, Rit93, Sav94, Sme96, Tus93, WS94, Als95, Ano94q, Int94e, Net94b, Ano96i, Bak95, Bar92b, Bar93, BH96, Bos94, Con95a, Dar93b, DS96, GBS95, Goo96, Gor95a, Gre95b, Gre96, Han92, Har96a, HH95b, HB95, KRS94, KRK95, Koc93, LH93a, Lar94b, LR95c, Mal95c, Max95c, McK96a, Mot94, RK95a, RK96, RK95b, Ros95a, Smi95a, Str92, Str93a, Tho95a, Tho96a, Wan94, Win93, Win96b, Wol96, Ano94-60, Ble95, Lic94a, Lic94b, Lic95a, Lic95b, Tat95a]. only [Bau93]. OnTheInternet [Soc95b]. Onto [Hal95a, Ano96o, Mil95e, Wri95b]. oo [Sim93]. OPAC [Mec94]. OPACs [HKR91, HKR91]. Open [Ano96u, CHR^{+}96, Cas93, PC93, QW93, GG96b, Wis90, Dig90b, CKMV99, HH95a, QW93]. OpenGL [Ano96u]. OPENSIG’98 [CKMV99]. Operating [Com87a, Com87b, Mog95, Pil90, Tri91, Tri93]. operation [PCF91]. Operational [Heg99]. Operations [GS99, HK91a]. Opinion [Coo93a, Har93b, Nol97]. opornych [LY94a]. Opportunities [OM95, Ano95-65, CB95d, Gor95a, Ano93n, McC92, NCS93, Ngu95, Ril95]. opportunity [Cam94b, PA95]. Opryland [Mas97]. Optical [Dut98]. Optimal [GD98, YN97, TAH99]. Optimization [Cah98, Leo98f, She95b, Cho95]. Option [Roc93]. Options [Not94b, Rey93, BU94a, CB95a, Pos83]. Oracle [HB98b, Lon94]. orbit [Koh97, Ano92a]. ordenadores [JE95]. Order [Eag93, Kol96]. Oregon [ACM99a]. organisation [SMLD95]. organisations [Ano96f, Lea95a]. organisé [Min94b]. organizaciones [WLM90]. Organization [Nil93, Fre94, Jar95, Rud95]. organizational [Ano96d, Bre95b, Lee98]. Organizations [OKP97, Osh93, Ano89c, Ano94-27, Bup95, Mas89, Hum94, Mad96, RMU96, Ste91a]. organized [Can94]. organizing [RK95c]. Oriented [HS98, Rog97, Cra92, MP89, Wes97]. original [BLF99]. Origins [HL96]. Orion [Ano96t]. Orlando [GMA^{+}98]. Orleans [B^{+}96b]. OS/ [HW95, Tha95]. OS/2 [Ive95, Mor95d, Pil90, Sal95b, Tys95]. OSI [NC90, PC93, Sta93b]. OSPF [Moy98]. Other [Gil94a, Mag94, Jai94, Pad85, Sto99, Ano94-41, BU94a, CS94a, Ehn94, FK95a, Gil94b, KL95, Mad96, MLJ95, NCS93, QCM94a, RZ96, RK95a, SMLD95, TL96, Wil96b]. Ottawa [Can94]. ottoke [Edd94b]. Our [Har93b, Ano95l, Koc93, RK95b]. Outlaws [HM91]. ouverts [RJ95]. Overview [Mur98, Ros89, MS96]. OWL [ANT96]. Own [LE95, HN96, Int95g, MMJ95]. oyster [Pot93, Pot94b].

Pace [Car99b]. Pacific [Hus94]. Package [Ano99d, BD88, She86]. Packard [Ano99e]. Packet [BPS99, CK74, MD91, CNY87, EST93d, FK99, Fos91, Gen83, Low90a, Low90b, Low91, Pax99, Sfe82]. packet-switching [Low90a, Low90b]. Packets [Bel93, Fla91, Rei92]. Page [Pow98, Ano95-40, HN96, McA95, Unixx]. Pages [BCW95, CD95, GM94b, HS94d, Hes95, MG94a, Aik95, Ano94v, Mec95b, Bw96, Cha95a, HS94b, HS95b, HS95c, Hah96b, HH95b, HB95, HH98, Mec96, MG94b, MG94c, Mse90, Nat96a, NC95b, NST96, Nof96b, Pig95, PDP96, Pol96, ST89, SP94a, SD96a]. PageSpace [CKTV96]. Pain [EO99]. Painless [Sim93, Ran96a].

Pair [IEE90, Hew88]. pak [Wen94]. Palliative [Ano94-52]. Panel [BLL^{+}96, DFG^{+}98, GFF^{+}95, GAM^{+}97, MBM^{+}95]. paper [Gry94, RR94b]. papers [Ano97a, Bol94, Cai94, Gaa76, Har96a, Pla94b, VT97]. para [CS95, Fre95, LB95b, MPF95, ME96]. paradigm [KG97]. Parallel [WA98]. ParaSoft [Ano99d]. parejas [Spi92]. parent [Ano95m]. parents [BM96a, Jos95a, Jos95b, SSA95]. Paris [Ano74, KW98, Min94b]. Parliamentaire [Can95]. Parliamentary [Can95]. Part [ANS92, Ano93a, Ano93b, Iow94, RHS95, Edm95]. partners [Ste91a]. partnerships [KT95]. Passcode [Cul95]. passport [Koc93, RK95b]. Past [LCC^{+}97, Web99]. path [AA94]. Pathfinder [Rya93, Tim94]. pathological [BPS99]. patient [McC95]. Pattern [OD98]. Patterns [GPRR94]. Paul [Mai95]. pay [Per95a, VJ95]. Payoff [Eag95a]. PC [Ano99e, Ano95-58, Col99, Del93, Eag95a, Fra94a, Gar93, Int95q, JE95, LR93d, Lar94a, LN94, New96, Pee96a, SS94b, SG95, Zha96]. PC-Compatibles [LR93d]. PC-DOS [SG95]. PC/DOS [Lar94a]. PC/IP [Sta88, Tri93, Tri91]. PCs [Ano96v]. PDQ [Ano93-31]. PDS [ATT93]. pegangan [Rah94]. PEN [Ano95c]. Pennsylvania [Ass95b, Wan94, Zha94]. penny [Win96b]. penny-pincher [Win96b]. People [KJGH98, Cru96b, MRJ96, Rot96, SSI^{+}98]. perceptions [Hei95b]. Perfect [Cha95c, Le 96]. Performance [ACM98, Buy99, CP98, Hum98, Kil98, Luk96, MP89, MA98, Mil90, Mil95f, ME97, Mog99, The94c, AHS90, AW97, Dra95, Gol92, Gup92, Hor98, Low90a, Low90b, Low91, TB90]. performance-analysis [Hor98]. peril [Jon94b]. perils [Wil95c]. perioda [Spi92]. Perl [Mul97a, Mul97b, IH98]. PerlDirect [Ano99d]. Persian [BG98]. Person [RB96]. Personal [GW95, Goo97, Kam93, Pre93, Pre95a, Pre97a, Wei93, ZTLV98, KLW95, Kon82, Mot94, Sta88, Tri91, Tri93, Tur95d]. personnel [Rou94]. Perspective [AN96, Hus94, KIP99, Pub94, SR99, TRHR96, TB93]. Perspectives [FRG97, GPRR94, MBDR91, Par90, PT99, BG98, BGMP98, MMR93, MMR94]. Peru [Int92d]. Peter [RN97]. PGP [Gar95b, Kar96, Ken95g, Sta95d, Zim95a, Zim95b]. PGPmail [Pre97b, Pre97c]. pharmaceutical [Smi94b]. Phillips [Phi95]. Philosophy [Ale96]. philosophy-related [Ale96]. phone [BA95]. phonebook [BN90]. Photoshop [Mil99a, ELB99]. PHP [Atk99, CRS^{+}99]. phrases [Str94]. phylogeny [MM95]. Physical [IEE90, ANS92, JP94]. Physician [Han95, McC95, Med96]. Physics [Ano97i]. PICS [RM96]. Piecing [Bow95]. pien [cCrmdh92]. pilot [CHHL91]. pincher [Win96b]. Pine [Buc95a]. pinpointing [Ano95-65]. Pioneers [Sta98a]. Pip [Fra93c]. Pirates [PG95]. piryohan [SG94b]. pit [Nor93, RHS95]. pitfalls [Gry94, Wil95c]. Pittsburgh [Ass95b, Gro90a]. Pixel [Ano97h]. Plain [Pfa94a, Int95g, Pfa96a]. Plainsboro [Lin93c]. plan [Int91, Mat95a, Net96a, VJ95]. plane [LeG95b]. Planet [Ano99e, Rim95a, Rim95b]. planetaire [LSLS94]. planete [Gue96]. PlanetUplink [Ano99e]. Planners [Pet93a, Gen95]. Planning [Cos94a, MA98, Gaa76, IBM92b, Mit92]. plans [Int93a]. Platform [Ano96v, Pet89, Web95, BH93]. Platforms [Gut99, Lei96]. Plausible [Pea98]. Play [Tan93, Hud94, RL95a]. Playing [SR95a, SR96]. pleasures [Wil95c]. Plug [GP96b, Hud94, RL95a, RBK^{+}94, Mas95]. Plug-n-play [RL95a]. Plundering [Spe93]. Plus [Ano99d, LaQ94b, Jaf94, JT96, LR93b, PT95b, CT96b, Pee96b, Ano96u]. PMI [Ano81]. PMI/Internet [Ano81]. Pocket [SL95, VH94g, VH94a, VH94b, VH94c, VH94d, VH94e, VH94f, Arm96, Ber95c, Ber96b, Dea96, Fai95a, Fis95b, Gac96, Gol94, GS96b, Gro95b, RA96a, Rye95, Tav95, VB95a]. PODS [ACM96]. Point [Bri95b, God94b, Ken95g, KP94]. point-&-click [Ken95g]. point-and-click [KP94]. points [Hor98]. policies [Fin95, Inn96]. Policy [BW97a, Coo93a, Ell99, MBDR91, VSN86, Ele91, Ano93x, Ano95-41, Bla99a, Hus94, Kah90, MBZ94, Mil95f, Sch91a, Sch91b, Wis96, Wal94a]. Policy-Making [BW97a]. political [Dix95, GP95a, GP96a, Lev95]. politicians [ACE^{+}95]. Politics [Man95c]. pollution [Sil95]. Polyhedra [Ano99d]. Ponders [Mil94d]. Pooh [Web95]. poor [Bau93]. Popular [CES^{+}99, Rau95, Jam96a, Jam96b]. pornography [Her96b]. Portability [X/O88]. Portland [ACM99a]. Portrait [Pet93b]. positioning [FB91, Shu93]. POSIX [QW93]. Possibility [LeB93]. Post [Kal94a, Ale95, RMU96]. Post-Soviet [Kal94a, RMU96]. post-teleconference [Ale95]. Postings [SG98]. postsecondary [Rot95a]. Potential [Hel99, Bro95a, Hof96b, MW92]. potpourri [Fol95b]. pour [Hoe95, Kap96, LSLS94, LB94c, Pla95, SLP94, SLP95]. Power [Not93b, CB96a, McQ94b, Ros95b]. PPP [Ano97j, Car98a, Gil95e, Kno96, MK95, Par95, SC95c]. PR4000 [Ano99d]. Practical [Bac95a, Bla98a, CMQ93, GS96a, He98, Hel91a, Mas93, Sta93c, Swi96, War99, Zel99, Ano97b, Cha92, CKY94, HC94, LMLJ95, Ols95, Woo96, Ano97d]. Practice [AV95, Ano94-61, Ano95-32, Ano95-65, KM96a, Sta99a, Str94, Wal94a]. Practices [Per98]. Prague [Pla94b]. Praxis [Jun95]. Pre [Uni96a, Ano93h, Rus94]. Pre-conference [Uni96a]. pre-service [Rus94]. precursor [Gie94]. predict [Jon94a]. prediction [Gol92]. prefix [SV99]. prepared [MW92, OD95, PM94]. presentation [Bar95b, Sau92]. presentations [MSB^{+}92, VT97]. presented [Sau92]. presents [Int95g, Mer92, Mer93, BAS95]. preservation [SJZ^{+}98]. President [MW92, MSB^{+}92, Vie93]. Press [CES^{+}99, ACE^{+}95]. Pretty [Gar95b]. prevention [Int85, PS95a, PS95b]. Preview [Ano93-35]. Pricing [CG95, Nol97, Per95a]. Primer [Ano93r, Ano93s, Car99a, LS93a, MH94, Tro94, All96, Ano94l, Ano94-29, BL93, LS93b, MD94, MP95, WK95]. Princeton [Lin93c]. principiante [LR94]. Principles [ACM96, Afs95, CKPT99, Com88a, Com91b, Com95c, Com88b, GW99, Per98, SBN82, Smi93a, Vit95, Sta99a]. Print [Ano96u]. printers [Ngu95]. Priority [Cam97]. Priority-Based [Cam97]. Privacy [Asa91, Bac95a, Fol95b, Gar95b, Ken90, Ken93, Tür94, Cla99, Cra99, Sta95d, WLW98, Wis90, Kol96]. privacy-enhanced [Wis90]. Private [DH99, Fow99, GKM^{+}99, KPS95, SWE98, Bak94]. Privatization [Coo93a]. Pro [Ano96v, Ano95-40]. Probabilistic [Pea98]. Probabilities [Car98b]. Problem [BWWF99, Mog99]. Problem-Solving [BWWF99]. problemas [Spi92]. Problems [GPRR94, MKBB95, Moh95, BDS93, BDMS94, Cal91, Hof96b, Man96, SLJ94]. procedure [IBM89a]. procedures [Ken90]. Proceedings [ACM96, ACM99b, Ano94-53, DLC92b, DLC92a, GV90a, GV90b, Gro90b, Gro90a, GRS79, IEE94, IEE95, IEE97, IC90, IC92, ILSB94, Int76b, JG91, Lei93, NDM98, USE97, Wil94a, Int76a, Ano81, Ano88a, Ano95f, BCB99, B^{+}96b, Cim96, FS95, Fra99, Hir97, Hir98, HL99, IEE98, IDK98, KW98, KSW99, LM98, LG93, MKT98, Mas97, Rii82, Uni96a, VVB85]. Process [She95b, Sev95]. Processes [KJGH98, MDW99, Sta98b]. Processing [Ack95a, Mul99, ANS92, New94b, New95c, New96, Sin99b]. Prodigy [Joh95, Ano94-60]. Product [Ani93b, Jon94a, Mat93, Mil95e]. production [ER98]. Products [Ani93a, Ano95-55, Ano96s, Ano96t, Ano96u, Ano96v, Ano97h, Ano99d, Ano99e, Gla99, IBM89c, Int95a, Jan95, Mat95a]. profession [Til95]. Professional [Ano96v, CRS^{+}99, GL99, Int95l, Dav95, HH98, HH96, LaS95b, Pri95, TPH96]. professional-looking [HH98]. Professionals [Ano93r, BLB99, ELB99, LS93a, MD99, Blo95, Edw95, LS93b, PS95b, Pri95]. professions [BBN^{+}93]. Profile [Pet87, Net96b, Pet88]. profiles [INT88, INT89b, Sne95a, Vic88, Vic89]. Profit [Mor95c, Bal93a, RT94b]. profits [LR95c]. Program [Ano90d, Ano93z, Luk96, MW92, Spa89b, Ste91a, NAS90, Ame95b, Col93, Cer93, Eng93a, Joh94, MSB^{+}92, Nat96b, Net96a, Pre97b, Pre97c, Spa88]. programmable [VT97]. Programmation [JCD96, LO99]. programmer [FHU99, Hom99, IBM89b, IBM89e, IBM90, IBM91b, IBM93]. Programmers [Fly99, Gar97, Ano97b]. Programming [Ano96v, Ano97c, Atk99, Bas99, Buy99, CRS^{+}99, CRH98, CS93b, CS93a, CS97, LS98, ML94, Rag93, Ste90, Ste98b, Ste99a, Wat96, WA98, BCB99, Com92, Com94a, CS96a, Fra96, IH98, JC95, SK98, Tho96b, T^{+}98, VJ99, WvRA96]. programming-by-demonstration [SK98]. programs [Ame95b, B^{+}96b, HW95]. Progress [Wal93, Ano93x, Gol92]. Progressing [Sin99b]. Progressive [Zel95]. Project [Ano74, Int76a, Ano88a, Ano94-54, D^{+}86, GK95, GMB93, Int76b, JO95, McC94a, NYS94, Rii82, Van95, VVB85, Ano81, BM98, CHHL91, Che87, Dil93, Gaa76, KG98, Low90a, Low90b, MM95, McC94b, PM94, SH95b, Int93a, Zho84, McC94b]. projection [TAH99]. Projects [Gla99, RAD90, FAR93a, Str94]. Projektu [Spi92]. Prolog [CRH98]. promise [Ass95b, Jon94b]. Promises [Gry94]. promote [Jan95]. promotes [LR96]. Promoting [FF99, Mat95a]. proof [Wri95a]. Prop [New95b]. propects [VE96]. Property [Sim95a, OO95]. proposal [GCGMP97, HOO95]. proposals [Ano94-55]. proposes [Web95]. proprietary [DFS98]. prospects [Gou90]. Prostate [Ano93y]. Protect [Sta95d]. protecting [Ano95l, Wei95a]. Protocol [Ano99c, Ben98, Bet87, CK74, Dee93, Haa85, Har94a, IBM89c, IBM89a, IBM89b, IBM89f, IBM89d, IBM89e, IBM92b, KvRvST93, MP89, Moy98, Pau99a, SRI82b, SC93, VH94b, WYLZ99, Wil93, AHS90, BM96b, Che87, EST93d, Fel94c, GLSW89, Hal93, Hin91, Hui96, Hui99, IBM88, IBM90, IBM91b, IBM91a, IBM92a, Kas93, Kol96, Lei85, MDB^{+}94, PP91, Pau99b, PMJ94, Pil90, Pos80, Pos83, Rei92, RG91, SRI82a, SDUU85, SRI82c, Net83, SC92, Sha91b, Sta88, Swa91, Tri91, Tri93, U.S83, Wan93, Yu93, ZDE93, Ano96c, Haa85, IBM89a, IBM89b, KvRvST92, Mil90, PSC81, ZDE93]. Protocol/Internet [IBM89a, IBM89b, IBM89c, IBM89f, IBM89d, IBM89e, IBM92b, IBM88, IBM90, IBM91b, IBM91a, IBM92a]. Protocols [Bla92, Bla95b, Com88a, Com91b, Com95c, Com88b, Fei93, Fei96, Fum98, HHS94, Hed87a, Low90a, Low90b, Low91, Mil91b, Mil92a, MSD90, SCT81, Ste94, Ano95-61, Cho95, Cra92, CP82, CB91, Gor95b, Hed87b, Leo94a, Maz90, Mil96b, NS94, PG96, SDUU85, SR99, She86, Tho96c, Xer81a, Xer81b, Ste96b]. Prototype [BB99, Ano96w, Wis90]. prototypes [Bre95b]. proven [Mat95a]. Provide [For95, Ano96w, HK91a, HK91b, Hen95a, ST89]. provided [Int95p]. Provider [Joh96a, Ano94p, CB95b, Den96, Fra94c]. provider-rooted [Fra94c]. Providers [Not94a, Wig95b, Int96k, Cle93a, DFS98, MP95, SC95c, SC95d, YM96]. Providing [DJB93, Koh97, Osh93, SS96, Dil93, Low90a, Low90b, Low91, MS94, VJ95, Wal95b]. providors [SMLD95]. provision [Har94b]. Proxy [Ano96u, Ano99d]. Pseudo [KG97]. Pseudo-serving [KG97]. pseudonymous [GGK^{+}99]. pseudoserving [KG99a]. Psychology [Gac98, Feh93]. Public [Ano94-55, Coo93a, Ell99, KK95a, KG98, KPS95, KSM98, Lun94, McC92, MBZ94, Ros95c, Wal93, Ano94q, Ano94-54, Ben92, Dra95, Fan95, FK95b, Haa85, Hal94, Hug94, JO95, Kah90, KL95, KT95, LaS95a, LaS95b, LL99, Mar96b, McC94a, McC94b, MBB95, McC97, NYS94, OD95, PS95b, PM94, Sch90, SH95b, TJ97, Uni95c, Van95]. Publication [Com71, Soc95a, Soc95b]. publications [Ano89c, Mas89]. publicize [Jan95]. Publish [Law95, FP95b, HN96]. published [HS94b]. Publisher [Ano97e, Ano97f, FD95]. Publishers [Fly95, Ngu95]. Publishing [AADF96, Fra95, Fry95, Hes95, HH98, Lem95, Ngu95, Nof96c, O’R96, Pre95a, Res97, WC95d, Ano95-52, Bri98, CC96a, Fol95a, Gry94, HOO95, HH95b, HB95, Joh96b, KH95, NR96, PJ95, Reh96]. publising [Wit96]. publizieren [Ble95]. punishment [Bol96a]. Purpose [CN95a]. purposes [KJTS96]. puto [cYlC95]. Putting [RR94a, GFF^{+}95, KP95, SRS^{+}95]. Puzzle [QW93]. pyolgon [LB95a]. Python [Ano97c, WvRA96].

QOS [CP99, BK99a, Bla99a]. QSAR [Dev96]. Qu [CRH98]. Qu-Prolog [CRH98]. qualitative [Dra95]. Quality [GG98, SBP^{+}95, Wol98, Bol96b, Cha95a, Sch90]. Quant [Ano99e]. Quant-X [Ano99e]. quantitative [ZCD97]. Quarterly [Com71, Int96j, Int96k]. Que [LR94, LR95a, PW95a]. Queens [Ano94-56]. query [LL99]. quest [AUC93, Ame95c, Ame95d, NAS93]. Questions [Saw93, Iow94, Int95b, MM91, MM92, Rot95b, Sav93]. Quick [Bra94b, BC95d, HKB95a, HKB95b, HK95a, HK95b, Que94, SP95b, Tol94c, Als95, Bra95, BU95, Car95c, DeR96, LY94b, LY95a, Pre97b, Ros95a, RHS95, Sei94b, Syd96]. Quicken [HP96]. QuicKIT [Mis95]. quickstart [PP94].

R [LO99, AHH88]. racial [Int94a]. radio [Pan95]. RAID [Ano99e]. Railroad [Bow96]. Railroad-related [Bow96]. ramifications [Dix95]. ramp [FHS94, HSF94, Med94b, Wal95b]. ramps [Ano95-29, RHS95]. Range [Mit92]. Rapidly [Car99b]. rate [RT99]. rätt [Dam93]. RawPacket [Ros84]. RBOCs [Ano95-59]. rc [Sim93]. RDBMS [Ano97h]. reach [Nov94, RBK^{+}94]. reachability [Sch91a, Sch91b]. reaching [Bau93]. read [Pri95]. Reader [Tha95, Uni95a]. Reader/2 [Tha95]. Ready [McK94b, AUA94, McK95]. Real [BB93, Fly99, Sha91b, ALS96, Bet87, Ros95a, Shi96, SJZ^{+}98]. real-time [ALS96, Bet87, Ros95a, SJZ^{+}98]. Reality [Ano96u, BMK88, CO99, Nol97, Tur95e, Vac96b]. Realizing [Kle94]. Really [LE95, Kur96a, LT92, Mas97, RAWS95, Rot96]. Reasoning [Pea98]. Reasons [SMB94]. REBMS [Ano99d]. Rec.g [Bel95]. receiving [Als95]. Recherchen [RS93b]. Recipes [Mas93]. recliners [Ano93y]. Recognition [OD98]. Records [Ano92g]. red [Cha95b, Ubi95]. redes [Cha95b, LR94]. reduce [Hen95a]. reevaluation [Lun94]. referatu [Spi92]. Reference [HS94a, Hof94b, Mel94, Nic93, Que94, Saw93, SL95, Ste98a, Tho96d, Tol94c, Ano95-38, Ano95-62, Int96j, Bre94, CC93, Car95a, Dee95, FHU99, Gol94, Hah96a, Hee93a, Hee95a, Hee95b, Hel94, HY94, Hof95f, Hom99, IBM88, IBM89e, IBM90, IBM91b, Jon94b, Kin94, LY94b, LY95a, Pre97c, Ree93, RNW^{+}94, Sei94b, SLJ94, Smi94b, Syd96, TA96, WSA^{+}94]. referencia [HS94c]. Reflection [KJTS96]. Reflections [Sto99]. RefLink [Zil95c]. reform [Hun93, Sto95c].

Reforma [Ano95-60]. regarding [Ano95-61]. regional [Shu93, Whi95]. registration [Ken90]. Registry [Lea95a]. Regulating [Cla95a]. regulation [ALS96, Coo96, Mue99a]. regulations [TMF97]. Related [Bla92, Bla95b, Dav98, Ale96, Ano95q, Bos94, Bow96, Bro94a, Dav95, DHM^{+}96, Dre92, Dre93a, Dre93b, Dre94, Dre95, Fol95a, Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b, Hub96, Lea95a, SLJ94]. Relations [Ros95c, Dra95, LaS95a, LaS95b, Mar96b]. relationships [GGK^{+}99]. Relay [Ano96v, Har95b, MD91, Rei93, Smi93a, Sta95a, Pyr95]. Release [Ano96u, Ano96v, MP92, GBJ94, Hud94, IBM91b, IBM91a, IBM92a, IBM92b]. Relevance [SS99]. relevant [Str92, Str93a]. Reliability [KIP99, LCP90, LCP91, LMG95, Sri93]. Reliable [Bro93, Kol96]. religious [Int94a, Str92, Str93a]. remain [Hof96b]. Remarkable [Sta98a]. remote [Ano94-55, FB91, IBM89a, Kon82]. Removing [Mul99]. Renater [Min94b]. reordering [BPS99]. Repelling [CB94c]. Replacement [YN97]. Replacing [KF93]. replay [Sha91a]. Replication [HK91a, HK91b, GS95c]. Report [Ano93o, Ano95-61, Ano96w, Ano98c, Con87, Ano93x, Ano94-54, Int95e, Int95h, Int95m, BU94a, Can94, Coo92, FB91, Int81, Her96a, Int93g, Int95p, McC94a, McC94b, MBZ94, MBB95, McN95, MD96, New95b, OD95, PM94, Spe87, Ste91a, SH95b, Uni89, UU89]. reporter [Net96b, Hum80, Int86]. Reporting [Car99b, Bri94a]. reports [Int84]. Representatives [Ano94y, Ano94-32, Ano95l, Uni89, UU89, Bro89]. Republic [Pla94b, Int94b]. Requirements [BP87, Bra89b, Bra89a, HK91b]. Research [Res92, Ano95l, BDS93, CC95, Cat89, DGH93, Gou90, Kah92, KS97, Les95, MBDR91, Mer87, Par90, RS93b, Sny95, Ele91, Int93i, Ano95-41, Ano95-52, Cyb96, Ano96q, Ano96p, Bak94, BDMS94, BW94b, BWC95b, BS94, Che87, Dil93, Har96a, Har96b, KM96a, Ker95, Low90a, Low90b, Man96, McK96b, RCL96, SD96a, Zho84]. researcher [Leg94a, Int96f, Bla93a, Bla95c, Mac95b]. Researchers [PRY99]. Researching [RK95c]. réseau [LSLS94]. réseaux [Min94b, Pla95, RJ95, Sch94]. Reserch [SI 93]. reservation [AHS90, ZDE93]. Reserve [Tri91, Tri93]. Residential [Bla98f, KSM^{+}96]. Resource [BH94a, Bla98b, Bla99b, Des93b, Deu92, DC94, EAR94, IH98, NSF91, Nor92, Not94a, ODL93, Per93, Pet93a, Que94, Rya93, Sch91c, Sch93b, She95b, Slo93, Tol94c, ZM93, ZDE93, AHS90, Ano89a, Ano92c, BH95, BH96, BDS93, BDMS94, BS94, Con95a, FS95, Fol95a, Han92, Han93, Har94c, Hol94, Int85, KHJG91, KHJ91, KB92, KHJ92, Maz90, Mil94a, Mil95a, Mil95b, Mil96a, NSF89, NSF92, Pie95, Sch90, Sch92b, SLJ94, Thi93, Uni93c, Ano94-30]. Resources [Ano92h, Ano93-30, Res95a, Dyr93, FH95, Gar93, Har94d, Har93b, He98, Mag94, Kam93, LG95, Mel94, Rya93, Rya95f, Rya95g, Rya95h, Rya96b, Smi93d, Vol97b, Ame95a, Ale96, Ano89a, Ano94l, Hea94, Ano94-31, Ano94-40, Ano94-41, Ano95a, Ano95c, Bus95, Ano95q, Leg95a, Ano96k, Anoxxa, Bri95d, Bri95e, Bak95, Bel95, BBN^{+}93, Bos94, Bow96, Bre94, Bro94a, Bur95b, Cro93, Din94, DHM^{+}96, Dre92, Dre93a, Dre93b, Dre94, Dre95, Ehn94, Ela94, EHPT95, FE91, FE92, Feh93, FB95, Gen95, GMDS95a, GMDS95b, GMDS95c, GMDS95d, Gia95b, Goo96, GB96, Her94, Han94a, Han94b, Har94e, Har95a, Hee93a, Hee93b, Hee94a, Hee94b, Hee94c, Hee94d, Hee95a, Hee95b, Hob96, HM96, JP94, JM95, JM96a, JM96b, JT96, Kal94a, Kan95a, KM96b, Kle95, Kra94, LS95, LS96, Luc94, Mac95a]. resources [Mah95, McK96a, Mil94b, Mol93a, MKM94, Mul92, MS94, Nat95, NYS92, NM94, O’H93, Ols96, Ols95, PT95a, PT95b, Pee96b, PN91, Per91c, Per91a, Per91b, Per92d, Per92e, Per92b, Per92c, Pet87, Pet89, PU94, PS95a, Pla94a, Pos95, PS95b, RK95a, Ril95, RJV94, RJV95e, RJV95a, RJV95f, RJV95b, RJV95c, RJV95d, Rud95, RMU96, SUR90, Cen93, SUR93, Sha95a, She77, Smi93b, Smi93c, Smi94b, SR95b, Sne95a, SHB95, TL96, Tee94, Tho95a, Tho96a, Tho95c, Tur95d, Upj95, Vir97, Wig93, WG94b, Woo96, Zil95a, Zil95b, Zil95c, Zil95d]. Response [Amo99, FF94]. responsible [KG97]. ressources [CD94, KCG95, LO99]. rest [Ano95-29, Tsu95]. Restructuring [Sar94]. Results [Pow93, Dil93, MBZ94]. Retail [WYLZ99]. Retailing [Rub95]. rete [Pas95]. reti [Sal95d]. retreat [Ste91a]. Retrieval [Bre93, Rud95, Sch95b]. retrieving [Mac95b]. Return [Rim95b]. Reveals [Gar97]. Review [Ano92i, Ano93r, Ano94f, Ano95g, Ano96b, Ano96c, Ano97c, Ano97d, Chr98, Com71, Hum98, Laz97, Not93c, RAD90, Roo97, Rya93, The94b, Cin96, The96c, Den96, Liu95, PN95, Tom93]. Reviews [Ano94g, Bar94]. Revised [Ano93-31, PP91]. Revolution [Lev84, MWL93, SRR97, Fre93, GP95a, GP96a, Ken95b]. REXX [Ano95-55]. RFC [BAG94]. Rich [God94a, MZ93]. ride [Net94a]. Rider [MG94b, MG94a]. Riderlink [CB95a]. Riders [BCW95]. Riding [Fis93, FVNT94, MWL93, RT94b]. rifarensu [Hof94a]. Right [Fow99, GL99, Smy95]. Righteous [CES^{+}99]. Rights [Hum76, Hum80, WMRS87, Ano89c, Fre94, Mas89, HSHF95, RAWS95, Ste91a, WS80, WR89, WLM90]. Riley [Ril95]. Ring [Bir94, Cam97]. rishtot [Mal95a]. Risks [Sny95, Neu98b]. River [Ano99e, ACE^{+}95]. RLIN [LeB93]. RMON [Sta99b]. Road [Ano94-57, New95b, Pil91]. Roadmap [Cri94, Ano95-72, Anoxxc, Fal94a, Fal94b, Fal96]. roadside [BDC^{+}95, Uni95e]. robbed [Law95]. Robbery [BK99b]. robots [MW92, MSB^{+}92]. rock [Gre95b]. Role [Kah94b, Val93b, Bak98, Gan95, Har94b, Lin93a, You95]. rom [RCL96, Ano94b, Ano95-68, Ins96, Joh96b, Mor95b, Pow93, Wal94b]. ROMS [Har96a]. rooted [Fra94c]. roses [CD95]. Rotterdam [VVB85]. rough [Ken95a, PMB^{+}92, Wat93]. round [GH93, GH94a, GH94b, GH94c]. route [EST93d, Gie94]. router [Con89, HGKM98]. Routers [Per92a]. routes [Wes97]. Routing [Ano94-36, CWJT99, DC90, FRB93, GD98, Hui95, Mau99b, Moy98, Sin81, Tsu89, Cho95, Gup92, Hal97, Hin91, LMJ98, Mil95f, NS94, Pax97, Rei92, Ste99b, TAH99, Wan93]. RSVP [SKB96, ZDE93]. rules [Ano94l, BAG94, BDKK96]. Running [Cha95c, Le 96, WE93]. Rural [Wal95b, Ano94-54, McC94a, McC94b, NYS94, PM94]. Rush [Lit93]. Russian [Mar93].

S [Ano95-55, IEE90]. S/REXX [Ano95-55]. safari [And92]. safe [Kug95]. safety [Blo95, Rou94]. said [Gar89, Web95]. Sailing [Ano93-32]. Sailor [Ano95-62]. Saint [Mil92b, Mil94e]. saizensen [Rik94]. SALBIN [Gar93]. Sallie [Fon94]. Samba [DeR99]. Sampler [Rak95, EU95]. San [Ano94-53, DLC92b, Hay94c, JG91, Mou96, TL93, Uni96a, Uni96b, SC94b, Ano94-58, Ano97g, Lei93]. Sane [Ano99e]. Santa [Tee94]. santé [CD94]. Satan [RR95b, RR95a]. Satellite [CWJT99]. Satellites [Bak93c, Koh97]. SATURN [Cas94]. Say [SE96, Wig95a]. sayong [SG94b]. Says [SMB94, For94]. SC’99 [ACM99a]. Scalable [BDMS94, HG97, BDS93, HGKM98, ZDE93]. Scaling [Wil94a]. SCAMC [Cim96, Mas97]. scenes [BDC^{+}95]. scenic [Gie94]. schema [BK91]. Scheme [LS89]. Schemes [IMI^{+}95, Gha92]. Scholarly [HOO95]. Scholarship [Byr95]. School [Ano93z, Des93a, Mei92, Mil94e, Ste91a, Gal94, Mas96, MD94, ODO95, Pie95, PWK94, Sch95b, VE96, Int93a, Woo95b, Mei92]. schools [Ben92, Gra95, Hil96, MS95, Mil94b, Sto95c, Wal95b, Ano94a]. Science [AAS^{+}97, Ano92a, Ano94y, Ano94-32, Che97, CT95a, CT95b, CT96a, CT96b, Den89, Eag93, Far97, FH95, Han96, HL99, JG91, KS97, LCC^{+}97, LG93, Nat92, Nat93a, Sac97, NAS92, And95c, Ano94-55, KM96a, NC90, O’H93, RJV95e, RJV95a, Rou94, SD96a, TW95, Yas95, Ano95l]. Sciences [MD93, Anoxxa, Bri95d, CD94, Dre92, Dre93a, Dre93b, Dre94, Dre95, Han92, Han93, Han94a, Han94b, KL94, RJV95c, RJV95d]. scientific [FP95a, HOO95]. Scientist [Ano93-33, Van96]. scientists [Tho95a, Tho96a]. scofflaws [DD98]. Scope [COZ99]. Scotland [FS95]. scout [Int95p]. Scrapbook [SK98]. Screen [Tha95, Tur95a]. SCSI [VFH98b, VFH98a]. SDK [Kuo97]. Sea [MWL93]. Search [AUA94, DH97, Gil94a, HKR91, SRR97, WC95c, Buc95b, Gil94b, Gil96, JN95, Ken95b, Rud95, Wan94]. searchable [JE96]. searcher [MRJ96]. Searching [LG98, LeB93, Les95, TB93, Ano94d, Bir90, D’E95, DG95, GCGMP97, RRO96]. seat [GFF^{+}95]. Seattle [Ano97a, Sea90]. Second [Bla99e, Hir98, JG91, KW98, MKT98, Sto95b, Ano94y, Edd96, NDM98, VT97]. secondary [Gan95]. seconds [MW94c]. secondsession [Ano94-32]. secret [Mar99]. secretos [LBP95]. Secrets [MSK99, Gui95, LB95c, Vac96a]. Section [Ame95b, IEE90]. Secure [CP94a, ET91, Kar96, Kol96, LS89, VJ99, WYLO98, WYLZ99, GGK^{+}99, Mar99, Opp96, PCF91, WC94]. secured [Kol96]. Securing [Bhi96, Bla93c, VJ98, PG96]. securité [BCRL95]. Security [Ano94-53, Bre98, Cam96, CB94c, COZ99, DH99, Fum98, Gar97, GS96a, KPS95, MSK99, Opp97, Rav95, SH95a, Sou93, Sta95c, Ste98a, Ahu96, Ano88b, Ano94-32, Ano94-33, Ber95b, B^{+}96a, Coo95b, DG97, Dav94, Den99, Esc98, Fol95b, GRLJ94, Hof96b, How95a, Hug95, INP95, Lun94, Mol99, Opp95, Opp98, PG96, Res95b, SPR98, Sta95b, Sta99a, SS95e, Uni89, UU89, Vac96a, VJ99, Wis90, YA95, dVdVI98, Ano94f, Ano97d]. SEDIT [Ano95-55]. seeking [Hus94]. seem [Nov94]. SeeMe [Sat95]. seen [Web95]. Sekunden [MW94c, MW94a, MW94b]. select [Ano95q, HJ93]. Selected [Cai94, Pla94b, RAD90, Ano94-34, Ano96l, Dav96, SUR90, Cen93, SUR93, SUR92, VT97]. Selecting [DH97, Den96]. Selection [Ame95b, FE91, FE92]. Selective [CX97, HKR91]. self [FP95a]. self-guided [FP95a]. sell [Hen95a, Jan95]. Selling [GG96b, Mat95a]. Semantic [SYTM98]. semantics [ST89]. Semi [AK97]. Semi-structured [AK97]. Semiconductor [Afs95, Gar97]. Seminar [Inf95c, De 94]. send [BA95]. sending [Als95]. Sendmail [AV95, Cos94b]. Senior [Mei92]. Sense [IEE90, ANS92, Ano96d, Clo93]. sensing [Ano94-55, FB91]. senza [LB94d]. Sept [Ano74, Ano81]. Sept. [Str95]. September [Int76a, Ano88a, Int76b, Lin93c, Sto95c, Uni96a, Uni96b]. Serial [Ano96v, Ano89c, Mas89]. serials [SLJ94, SM93, Str92, Str93a]. Series [Ano94-36, Moc94]. Server [Ano96s, Ano96u, Ano96v, Ano99d, Ano99e, Byr99, CS91a, CS97, Don96, EO99, Fou98, Gon99, HB98a, HB98b, INP95, LE95, Mog99, Moh99b, Seg95, TK99, Vac99, VRFJ97, Ano97b, BOB93, Bea95, BT96, BPR88, Bud96, Che87, Com92, CS93b, CS93a, Com94a, CS96a, Eck95, Ehn94, HE95a, HAS95, Int93c, Int93f, KG99a, MMJ95, dAMW97, New95c, Nof95b, Par93, YM96, Ano96b, Ano96v, Ano97h]. server-side [KG99a]. Servers [Mor94, Mul97a, AW97, Bas99, Bea94, Bos94, GGK^{+}99, Mog95, Row96, SBGK94, Zho84]. serveurs [Pel95a]. Service [Bri95c, Edu95, ED92, GG98, Joh96a, KSM98, ME97, Pet93a, Pet88, SQ93, Sin81, Wal93, Int96k, Bla99a, CB95b, Dee95, DFS98, EH95, Int95p, Leg93, Pun91, Rus94, SMLD95, Sch91a, Sch91b, Str95]. service-level [Sch91a, Sch91b]. Services [Ano99d, Bil99, Blo94, DJB93, Ibe97, KSO95, KSM98, KSM^{+}96, LPJ^{+}94, ME95, Nor92, ODL93, Per93, RBM94a, RBM94b, Tan93, X/O88, AP95a, Ano90a, Net94b, Ano95-58, Res95a, Cin96, Ano96f, Int96j, Ano96h, BH96, Bla98e, Bos94, Buy96, CS94a, CHR^{+}96, Cle93b, Dil93, FK99, Goo96, Gor95a, Gre95b, Gre96, GP95b, Gui95, Har96a, HGKM98, Int95a, Int95q, Jan95, Jon94b, Kes90, Kes93, Kin94, KRS94, KRK95, KHJG91, KHJ91, KB92, KHJ92, Lar94b, Leo94a, Lin93c, Liu95, Liu96, LR96, Mal95c, Mat93, Mat95a, Ban94, MS94, Oti95, RO94a, RO94b, RCL96, Rya95f, Rya95g, Rya95h, Rya96b, Sav93, SS96, VJ95, Win93]. servicios [CS95]. serving [Bou97, KLW95, KG97]. Sesame [LSLS94]. session [Ano94y, Ano95l, GFF^{+}95, MBM^{+}95]. sessions [SG93b]. Set [Leo98g, Rau95, Ste97, Sun96, McC97, Pao96]. Set-up [Rau95]. setters [Ano94-46]. Setting [CC96b, Nof95b]. Seven [Leo98g]. Seven-Volume [Leo98g]. Seventeenth [Gro90a]. Seventh [FS95]. SGML [Pao96, Wit96]. sha [LB95d]. Shadow [Ano94-60]. Shamash [Ano90e]. shang [hLcL95]. Shannon [CDEF94]. Shared [HS98, Har93b, SW93]. shares [Mou96]. shareware [Par95]. Sharing [Bag95, Slo93, ZM93]. sheet [Ano94o]. sheets [HKR91, PMTJ98]. shelves [Spe93]. Sheraton [Cim96]. shih [But95b, nChCiY94, yTZ94, kTlC94, kTlC95].

Shikhar [HH95a]. Shipping [Ano96s]. shoot [MBM^{+}95]. shoot-out [MBM^{+}95]. Shootout [Ken95c]. Shop [Ano97h, Bla96b]. Shopping [Eas95, Val93a, RA96a]. short [Mat93, Pad95]. shots [Ano93y]. shou [AJRmT95, HSskw95, LY95b, yTZ94]. should [BM96a]. shouyouka [SI 93]. show [Ano94-35, LSBM95]. Shows [BR95]. shu [LB95d]. si [SG94b]. Siberia [BBK^{+}98]. sic [KB92]. Side [Ric95c, KG99a]. SIGACT [ACM96]. SIGACT-SIGMOD-SIGART [ACM96]. SIGART [ACM96]. SIGCOMM [Com71]. SIGGRAPH [ACM99b, Ano98b, B^{+}96b, GMA^{+}98]. siglo [Fre95]. SIGMOD [ACM96, Win93]. Signal [Ack95a]. Signaling [Bla97b, Bla98c, CKMV99]. significant [Bak94]. Silicon [Sto95b]. Simple [Cog94, Cog95, Dee93, Har94a, Ros91, Ano96q, DeR96, McB95, Ros94b, PCD92]. simplified [Mar95a]. simplifying [YmtEC99]. Simply [Eng95c]. simpozija [Spi92]. simulation [Wan93]. simulator [Pot95, SG95]. simultaneously [JGKR96]. Singapore [AN96]. Single [MDZ92]. SIP [Dee93]. sistem [Rah94]. Site [Cha95c, Pow98, Ste97, AP96, BT96, Bud96, CL98, CC96b, Int95g, KKM^{+}98, Le 96, MMJ95, Pri95, Rya95i, WC94, Pri95]. Sites [GL99, Wei95c, Ano95-41, Atk99, Ber95a, Bos94, BDC^{+}95, Buy96, Dav96, DH96, Iow94, KK96a, KL95, Lev95, LCP90, Max95b, NM98, RdG95]. SiteWire [Hew88]. six [Gaa76, Mai95]. Sixteenth [Gro90b]. Skills [Ril95, Ano95-63, Bar95b, BBST96, RHK95]. Skip [Wei95b]. SLA [Ano91]. Slander [Fol95b]. SLATE [Leg93]. Slick [GG94]. SlickEdit [Ano96t, Ano96v]. slightly [Ros96]. SLIP [Bai94, Gil95e, MK95, Par95, SC95c]. SLIP/PPP [Gil95e, SC95c]. SLiRP [Kno96]. Slotted [Cam97]. Slotted-Ring [Cam97]. Slouching [Nov94]. SM [Per92e].

Small [Mei97, Ano95-29, Sma95, HB96, Mad96]. Smiley [Godxx]. smileys [Str93b]. Smoothing [RT99]. snail [JEK92]. Snake [Sto95b]. SNMP [Har94a, Mil93b, Sta93c, Sta99b, Tow95]. SNMPv2 [RJ95, Sta93c, Sta99b]. SNMPv3 [Sta99b, Zel99]. Soaring [Col95]. Social [Han96, Rei93, Bri95e, Hum94, KM96a, KKM^{+}98, RJV95c, RJV95d, Rya95i]. Societal [BW97a]. Society [Ano94-53, Kan94b, PE95, Pla94b, TL93, Win98, Lei93]. Society/5th [Pla94b]. Socket [CS93a]. Sockets [Com92, CS97, Mul97b, Ste98b]. Soft [Kan94b]. softbot [EW94]. softbot-based [EW94]. Software [Ano94-59, Ano96t, Ano96u, Ano99d, Ano99e, Gar93, LaQ94b, Nej94, Sam94, Uni96a, Uni96b, Wol98, YY98, Bio96, Ano97b, Bir93, Bre95a, DS96, D^{+}86, Eng93a, GB95, Gro95a, Ken90, LR93b, Lar94a, LN94, MRJ96, Net95b, SS93, SMLD95, Str92, Str93a, Tri91, Tri93, Wen94, FTP95a, FTP95b]. Software/Systems [Ano94-59]. soliciting [Ano94-55]. SOLID [Ano96v]. Solution [Ano96u, JL94, SPR94, KP94]. Solutions [Ano96v, Ano99e, MSK99, Int94e, Cal91]. Solving [BWWF99]. Some [MMV93, Max95b, Wig93, Hof96b, Rot95b]. sometimes [Ril95]. Sonet [BW97b]. songs [Jor94]. Sound [Ano93-34, GGA^{+}97, GAM^{+}97]. Sounding [GGA^{+}97, GAM^{+}97]. sounds [BDC^{+}95]. Source [Uni93a, Uni93b, Zim95b, Ano95u, GGMT99, Gup92, LJN94, Lev95, LN95, NN94]. Sourcebook [Kah94a, Ano96i]. Sources [AAB^{+}98, AK97, BB99, End94, New93, Tro94, Vol97b, Vol97a, AT94, Dec95, Gum93b, Gum93c, Gum94, HSHF95, Har94b, LJN94, Mot94, NN94, New94a, SUR92]. Southeast [Mul92, Sta94b]. Southeastern [Mei97]. Southern [Whi95, BD95b].

Southwest [Vel95]. Soviet

[Kal94a, RMU96]. SP3D [Hal99]. SPA [Sme96]. Space [Ano94y, Ano94-32, Gar97, LG93, SYTM98, Ano94-55, Kug95, Mil95e]. SpaceMath [Hal95b]. spaces [Shi96]. Spam [Mul99, SG98]. Spatio [HM97]. Spatio-Temporal [HM97]. Special [BC95d, HK95b, IL93, LT93, TPBH94, Yan93, Yan94, Ano92f, LT92, Ano93-29, Spe93, Til95]. specialists [Sim95b, SM93, VE96]. Specification [Dig82, LY97, MP89, Ano89b]. Specifications [IEE90, ANS92]. specimen [BOB93]. Spectrum [Hel91a, Vit95, Sch90]. Speculations [Nil93]. speech [Ano95-34]. Speed [Cos94a, Hel99, Sei98, HL89a, HL89b, Jai94, Maz90, Pot94a, SD96b]. Speedier [Mog99]. speeds [Ano93z]. sphere [Fan95, FK95b]. spiders [Che96]. Spinnin [Ano95-63]. Spinning [Fis95a, For95, Sac97]. Spire [Ano96u]. split [Ros84]. split-gateway [Ros84]. sponsored [Abb94, LG93, Ste91a]. Sports [Fai95b, Fai95a]. Spread [Vit95]. Spreading [ZM93]. Spring [Ano95-64, Mil95b, BR95]. Springer [Ano93-35]. Springer-Verlag [Ano93-35]. Sprint [Ano93c]. Sprint-MCC [Ano93c]. Sprite [Che87]. Spy [Sto89, Pro96]. SQL [Ano99e, Byr99]. Squid [Mog99]. SRP [AHS90]. SS7 [Bla97b]. SSL [LC98]. ST [PP91]. ST-2 [PP91]. St. [Mai95]. staff [BW94a, BW94b, Hic94, Sta94b, Tee94]. stages [Hic94]. Stamping [SG98]. Standalone [ME97]. Standard [Dig82, DH99, Pos80, U.S83, Wis90]. standardization [Hin91]. Standards [Mad89, QW93, Sev95, Sta93b, Sta93c, Sum99, SC88]. Stanford [GCGMP97]. Starlink [PHW^{+}93]. starptautiska [Spi92]. Start [Arm94b, Des93b, LaQ94b, CB95b, Jan96, LR93b, Sch95e]. Start-up [LaQ94b, LR93b]. Started [MKNW92, Arn96, Bax95, GA95, LD95, MK92, M^{+}93, MKNW94, MLJ95, SMLD95, SC95b]. Starter [Eng93c, Eng94a, Eng94b, Eng95b, ELS95, ELO95, HTW94, Pfa96b, RRF95, Sei94c, You94]. STARTS [GCGMP97]. State [BD95b, GG98, Gro90b, IDK98, Owe93, Sta94b, Ano95-61, Lev95, Mle95, New95b, TFG99, Ano95-61]. state-coordinated [TFG99]. State-of-the-Art [GG98, IDK98]. statement [Bro89]. States [HS94b, Mot94, RMU96]. stations [Pan95]. Statistics [Ano95-72, Anoxxc, BD88]. status [FB91]. Stay [HL96]. Step [Neu98a, Ste98a, Ano94-40, Ano96j, Bea94, Cat96, FP95b, GO95, Ken94e, LS95, LS96, Mat95a, TMF97, Tür94, Int93a]. Step-by-Step [Ste98a, Ano94-40, Ano96j, Bea94, FP95b, GO95, LS95, LS96, Int93a]. Steps [Ben91, TMF97]. Sterling [MSB^{+}92]. STI [Ano90d, NAS90]. Still [CDEF94]. Stones [ACE^{+}95]. Stop [Val93a]. Stopping [SG98]. stops [RHS95]. Storage [BM98, Tho99, FB91]. Store [Ano99d, Ano94-60]. Stored [TS88]. storefront [GG96b]. Stories [MWL93]. storms [Swa91]. Story [Ayr94, Bar95a, GL94, Sta98a, Vie93, Ano93h, Nor93]. Strange [Chr98]. straniere [Pas95]. Strategic [Jar95, Nej94, Inf95a]. Strategies [DM95, Cor95, Int95m, EE95b, FE91, FE92, McQ94b]. strategist [KM96b]. strategy [Ano92e, BBSH98, Cro96b, Mai96, RR94b, Son92, Ste95a]. stream [PP91]. streaming [Rus98, WLKS98]. StreetWatch [Str95]. strict [Ken90]. Striking [MZ93]. striving [SY95]. structure [Bre95b, Ram95]. structured [AK97, LL99, QRV95]. Structures [SBN82, TGL^{+}97]. Strut [Eng96a]. Student [CC95, Cla95c, SW97, Of 95, Bla93b, Cla96, DOO95, Glo95, Mar96a, ODO95, Pit95a, Pit95b, Pit95c, RK96, Sch95b]. Students [Che97, He98, Klu95, Ano95s, Chi95, GG96a, Gru94, Jos95a, Jos95b, NCS93, Nat93b, PS95b, Sta94b]. Studies [CES^{+}99, Mar93, Sti94a, Bri95e, Int96e, Mol93a, PU94, Sti94b, Str92, Str93a]. Study [Com95b, GK95, Mog99, Pow93, RV99, Ale95, Ano96k, Bag95, But95a, CN95b, Dra95, EHS92, Fos91, FK95b, Gal94, Gup92, Kal94a, KIP99, Li90, Lin93a, LCP90, Mai95, MBZ94, Mle95, Moh95, Ngu95, Nic95, Rya95i, Sch91a, Sch91b, SEDK94, Som94, Sri93, TW95, TB90, Wan94, Zha94]. Stuff [Eng96a, Vin94a, Ano95k, EV96, May96]. Stumpers [Fee93c]. Stumpers-L [Fee93c]. Style [AH94, LC93, Pad85, LC96, PMTJ98]. SU-PC [Sta88]. SU-PC/IP [Sta88]. Subcommittee [Ano94y, Ano94-32, Ano95l, Bro89]. Subject [Mag94, HV95, RJV94, RJV95e, RJV95a, RJV95f, RJV95b, RJV95c, RJV95d, Tho95c]. Suboptimal [GD98]. Suboptimal-Optimal [GD98]. Subscriber [SCS99]. subscribers [Edu95]. subscribing [Ehn94]. substitute [JEK92]. subversive [HOO95]. succeeding [Ano95-70]. Success [Wya94b, Wya95b, Int94e, KKF95]. Successful [Bry94, HSK98, SE96, Uni95b, WE93]. such [Bie93]. Sudan [Ano93-36]. SUIT [PCD92]. Suite [Ano96s, Wil93, BP94, Lei85]. summary [INT89b, LT92]. Summer [BH94b]. Sun [Hum98, Ano94-51, CP98]. Super [Edm95, Ano93u, DG94]. Supercomputer [Ano97g, Gro90a]. Supercomputing [Ano94-50, Ano97i]. Superhighway [Ano94-40, Eag94a, HSF94, LH93b, Sto95b, SMB94, Ano94t, Ano94-35, Net94d, Ash95, Bra94a, BR94b, BDC^{+}95, CS94a, Cim96, DKF^{+}95, For94, FHS94, Gie94, LH93a, LSBM95, Mat95a, McM94, Mil94b, Med94b, New95b, Ott94, PK94a, Per95a, Per95b, RT94b, Thi94, Wal95b, RHS95]. Supernet [Ano94j]. superstitions [Mam96]. Supplement [IEE90, LeB93, CT96b]. supplies [ACE^{+}95]. Supply [Ano93m].

Support [Ano92g, BBH93, DGH93, EO99, KSO95, Lei96, SW97, Ale95, Net94d, Bra89a, JP94, Mat93, McL96b, Mog95]. Supported [BBH93]. Supporting [Gon99, KvRvST92, KvRvST93, Ran96b, Sch90, Vac99, ZTLV98, Cho95]. SURAnet [Cen93, SUR93, SUR92]. Surf [HC95, Bol96a, DG95]. surfboard [Wya95c]. Surfer [Ano95-42]. Surfergrrrls [GK96]. Surfez [KS95, KEdT96]. Surfin [LMLJ95, Net94a, Tsu95]. Surfing [Her95b, Her95c, Hof95c, MK96a, Min94a, MWL93, Pub95, Pol92, Que95, TK95b, TK96, McN95, PM95b, Pfa95b, Pfa95c, Pfa95d, Pfa95e, Pfa96c]. surfisti [Her95a]. Surveillance [Amo99]. Survey [AR98, Gar97, Pow93, Sou93, Coo93b, Jon94a, LMG95, OD95, Uni95a]. Surveying [Jon94a]. survie [Soh94, Soh96]. Survival [EE99, Har95b, Web99, Par95]. Surviving [Pfa95f]. Sverige [BJS94]. SVR4 [Com94a]. Sweden [KSW99]. sweet [Pad95]. Switch [Ano96v, SC93, CNY87, SC92]. Switched [Bla98a, Bla99a]. switcher [Gen83]. Switching [CL95, Dav98, MD99, Bla99a, Low90a, Low90b, Low91]. Switzerland [Cai94]. sword [RR95a, RR95b]. syllabus [Chi94]. Symposia [Pao96]. Symposium [ACM96, Ano81, Ano94-53, Ano98a, Ano98c, Ano99a, Cim96, IDK98, Mas97, USE97, D^{+}86]. Synchronization [ASB97]. syndrome [Ano93w]. Synthesis [She95b]. System [AAB^{+}98, And99, Ano94-53, Ano96v, Ano99e, BBH93, CN95a, DKF94, Her93, Hiu96, IEE90, KP93, LS98, MWB95, MSD90, Mil90, Mit92, Moc94, MMLP97, Mui95, Pad93, PR91b, QCM94b, Rag93, Ran96b, RP93, SBN84, Sun96, TS88, Tun99, Vac99, BAG94, Bow90, Bri94a, Bry95, sC86, FBGP98, Hof96b, IBM89a, Int92a, Int93f, KTS94, LL99, LR93e, Moj93, dAMW97, Pet87, PR91a, Pil90, Rud95, SHN97, Tri91, Tri93, Ano96t, MP92]. système [Sch94]. systemic [Hun93]. Systems [ACM96, Ani93b, Ano94-59, Ano96u, Ano98a, Ano99e, Bry94, Cas93, CL95, Com87a, Com87b, CDK94, COZ99, Cro94b, Fur99, GW95, Goo97, GMB93, Heg99, KvRvST92, KvRvST93, Leo98g, Leo98f, Leo98b, Pea98, PC93, Sau92, Slo94, Ter99, TE99, TS98, USE97, YN97, ANS92, Ano95-68, Bos94, Bro95a, FB91, Hel94, Hug94, Jan95, KG98, LM98, Leo98a, Leo98d, Leo98e, Leo98c, Mas97, McC88, Mil95e, Mog95, Opp96, RT98, SJZ^{+}98, Tri87, TB90, VT97, Ano96w, Ano98c, QW93, Ano99a]. Systimax [ATT93].

T [Hel98, IEE90]. T1 [BW97b, Fla90]. Table [Ano93-35]. tables [Nof96a]. Tactician [Ano96u]. Tag [Dav98, Pao96]. tags [Pao96]. Taking [Gie94]. Talk [Har95b]. talking [Rya95i]. taming [Ree93]. Tampere [Ano96x]. tao [cCrmdh92]. tap [Vel95, FH95, MOD93a, MOD93b]. tapped [Hof96b]. Tapping [FH95]. Target [Mee95]. Task [DLC92b, DLC92a, GV90a, GV90b, Gro90b, Gro90a, Ano89b]. tasks [SK98]. Taxing [Gar97]. Tcl [Ano95g, Fly99, IH98, ZP98]. Tcl/Tk [Fly99, ZP98]. TCP [Ano89b, Ano93-37, Ari93, Arn94, B^{+}95, Bla92, Bla95b, Bre95a, Bro93, CMQ93, Com88a, CS91c, Com91a, CS91d, CS91b, CS91a, Com91b, Com92, CS93b, CS93a, Com94a, Com95c, CS96b, CS96a, CS97, CS98, CS91e, Com88b, CB91, Dav88, Fei93, Fei96, Gor95b, Hun92, Hun98, IBM92a, IBM92b, Int92a, Int93b, Int93d, Int93c, IBM93, Int93e, Int93f, Ker98, Kon91, Los95, ML94, Mil92a, Mil96b, Mil97, MSD90, Mur98, PG96, Par94b, Pil90, PC93, Res95b, San91, San94b, Sch91a, Sch91b, Siy94, Ste94, Ste96b, Tho96c, WE93, Wil93, WS95, Zna94]. TCP/IP [Ano89b, Ano93-37, Ari93, Arn94, B^{+}95, Bla92, Bla95b, Bre95a, Bro93, CMQ93, Com88a, CS91c, Com91a, CS91d, CS91b, CS91a, Com91b, Com92, CS93b, CS93a, Com94a, Com95c, CS96b, CS96a, CS97, CS98, CS91e, Com88b, CB91, Dav88, Fei93, Fei96, Gor95b, Hun92, Hun98, IBM92a, IBM92b, Int92a, Int93b, Int93d, Int93c, IBM93, Int93e, Int93f, Ker98, Kon91, Los95, ML94, Mil92a, Mil96b, Mil97, MSD90, Mur98, PG96, Par94b, Pil90, PC93, Res95b, San91, San94b, Sch91a, Sch91b, Siy94, Ste94, Ste96b, Tho96c, WE93, Wil93, WS95, Zna94]. Teach [Lem95, Par94b, Ran94, Ran95b, Via94]. Teacher [Mey95, The96b, oIT94, Gan95, Hei95b, Pla94a, She77, Van94, Ano90f]. teacher-centered [Gan95]. Teacher-Friendly [Ano90f]. teachers [BU94b, Gal94, Klu95, Mai95, Mar94, Mil94a, Mil95a, Mil95b, Mil96a, NCS93, Nat93b, Pie95, Pla94a, Rus94, SRS^{+}95, VE96, Wil95b]. Teaching [Rob94, Roc93, SRS^{+}95, Ass94b, Ano94e, Bra97, Cot95, War95]. TEAMate [Ano96s]. teambuilding [SE96]. Teaming [Sav96]. Tech [Sto99, Ros84]. Technical [A^{+}98, Ano98b, Mal92, Mur98, Nol97, PT99, RAD90, Ril95, Wil94a, Ano97a, Mat93, MS94, SLJ94, Wil99]. Technik [AB94]. Techniques [Leo98g, Leo98f, Leo98b, Mar98, Mil91b, WA98, Ano97b, BT96, Bud96, CB96a, Gil96, Ham99, Hug95, Leo98a, Leo98d, Leo98e, Leo98c, MMJ95, SCJ96]. technologie [And95c]. Technologies [Ano96u, Ano98a, Ano98c, Ano99a, Bla97a, Bla99c, BL95d, Dav98, GW95, HPS95, KJGH98, USE97, RZ96]. Technology [Ack95a, AAS^{+}97, Ano90f, Ano93r, Ano93s, Ano93-38, Ano94y, Ano94-32, Ano94-36, Ano94-60, Ano94-61, Ano95l, Ano99b, Ano99d, CAC94, CFL99, Dat74, DM95, Far97, FB91, KS97, LS93a, Lei94, LCC^{+}97, Moc94, Sei98, Sac97, SS95c, SCS99, Ano94-55, Int95h, Ano95-65, CT95a, CT95b, CT96a, CT96b, Dix95, Hei95b, Her96a, Hun93, IDK98, LS93b, Lin93a, MRTR95, Mai95, McN85, MW92, Moh95, Rus98, SRS^{+}95, SD96a, Win98, Wri95a, YM96, Ano95-61, Ano96w]. tecnologia [Fre95]. Tecplot [Ano96u]. Teens [Cay93]. Telcom [Ano95-66]. Tele [MKBB95]. Tele-Conferencing [MKBB95]. Telecom [Ano94g, New94b, New95c, New96]. Telecommunication [CN95a, Bla98e, HGKB99]. Telecommunications [Ano93-39, Car98b, Car99a, Dig93, Dod99, Mac93b, Rus94, Ano94-61, BU94a, Bal93b, Bro89, NCS93, New94b, New95c, New96]. Teleconference [KT95, Ale95, BTC^{+}95, Ble93, PHW^{+}93]. teleconferencing [Rus98]. Telegraph [Sta98a, Win98]. Telenet [Pos83]. telephone [Kes97]. Telephony [ME97, Fos91, New94b, New95c, New96, SR99]. Television [CFL99, Pre93]. teleworkers [PH97]. telex [Mei82, Zu 82]. Telix [SS93]. Telnet [Nof95e, BW94a, BWC95a, Fel93, FK95a, Bar93, Fel94b, SG93a]. Telnetting [VH94a]. Temporal [HM97]. tendences [Spi92]. Tennessee [Dav95, Eme95b]. TERIS [Mit92]. term [Har96a, Fra93c]. Terminal [Ano96t, VH94f, Sch92a]. terminals [McC97]. terminology [McK96a]. Tertiary [PE95]. Test [Gar95a, Mit92]. Testing [Afs95]. tether [SP94b]. Texas [FMB95]. text [Bre94, GGMT99]. text-source [GGMT99]. Textile [Ano96x]. Textiles [Lam96]. Texts [Mag94, Ano94-41]. tezes [Spi92]. thar [MZ93]. their [HOO95, Heg99]. them [Bea94, LJN94, Lev95, MZ93, NN94]. TheoLink [Zil95d]. Theoretical [Mac93a]. Theory [AV95, Ehn94, Lee98]. therapeutics [SS95d]. therapy [SS95d]. There [Bar95a, Ano95k, MZ93]. thesis [Sha91b]. Thin [Ano99e]. things [Bie93]. Thinking [MW92, MSB^{+}92, Ass94b, Ano94e]. Thinternet [SP94b]. Third [Ano94y, Ano94-32, Fra99, KSW99, LG93, DLC92b]. thirtieth [ILSB94]. THOMAS [Wil95a]. those [Owe93]. Thought [Ano96u, PMB^{+}92]. Thoughts [Sto95b]. thousands [HW95]. Threading [Thr94]. Threaten [Gar97]. threats [SS95e]. Three [CW93]. Threshold [TOLF93, Ano94-61, TOL94a]. throughput [FK99]. thru [McN95]. ti [kTlC94]. ticketless [Vel95]. tickets [Hen95a]. Tien [Edd95b]. tietoverkoilla [Hin93]. tighter [Ano88b]. Time [Cul95, Hil97, Mil90, Tim94, YY98, ALS96, Bet87, Ros95a, Sha91b, SJZ^{+}98, Wes95]. timely [JE96]. Times [DC94, BGMP98]. TimesFax [Ano95-67]. Timken [Kar96]. TINA [CKPT99, KSM98]. TINA-based [KSM98]. tip [Wes94]. Tips [CP94a, Ano94l]. tirgus [Spi92]. Title [Mag94]. Titles [Pow93]. Tk [Ano95g, Fly99, ZP98]. TLI [CS93b, Com94a]. TLS [Ano99c, Pau99a, Pau99b]. TN [Mas97, Mei97]. today [Int95c, Int96l, BC95a, Pow93, Sch95e, VE96, BK96b]. todo [JE95]. together [Bow95]. Token [Bir94]. toll [TT95b, Per95a]. toll-free [TT95b]. tomorrow [BC95a, BK96b, Pow93, SRL96, VE96]. Tool [Ano93-42, Ano99d, MKBB95, Nej94, Reg93, Bak94, BD95a, Bir93, Ced95, Lof95, Mai95, MO92, Mar96b, SP94a]. toolchest [Mor95b]. Toolkit [Ano96u, Ano96v, Del93, McK94b, CB96a, IBM93, McK95, NR96, Rim96, Sch96b, Ano95g, PCD92]. Tools [Bre93, Cah98, EAR94, Gil94a, GL99, HSF94, Har93a, Mil91b, MW92, SBGK94, SHN97, Tro94, Ano94l, BL95c, BT96, BW94a, BW94b, BWC95a, BWC95b, Buc95b, Bud96, CB96a, FHS94, Gil94b, Gil96, Gro95a, MMJ95, McQ94a, McQ94b, MLJ95, Ric94a, Rii82, Ros95b, RM94, Rus98, SLJ94, Str94, Tho95a, Tho96a, Til95, WC95c, Web95, ZP98]. Top [Ano93-28, Mar99, Sch95g, Ber95a, Coo95a, DH96]. Topics [Car98b]. Topology [COZ99, ZCD97]. Toronto [ILSB94]. torpes [ME96]. Total [Har94a, AA94, Int93e]. Tou [nChCiY94]. Tour [Bar94, Bra94b, BC95d, Bre91, Fra93b, Fra94b, FAW95, FJ95, HKB95a, HKB95b, HK95a, HK95b, SP95b, Als95, Arm96, Ber95c, Ber96b, Car95c, Cru95a, Dea96, Fai95a, Fis95b, Fra93a, Fra94a, GS96b, Gro95b, HLB^{+}95, RA96a, Ros95a, RHS95, Rye95, Tav95, VB95a, Wag95b]. tourguide [Wen94]. Tourism [End94, Cha95a]. tourists [Cha95a]. Tours [GL94]. tous [Hoe95]. tout [BP94]. TowerEiffel [Ano96u]. TP4 [Son92]. TP4/IP [Son92]. TPR [Hay95]. TPR-NE [Hay95]. Trace [Amo99]. Tracking [Pre97a, Sto89]. Trade [BR95]. tradeoffs [Bol96b]. Trading [WYLO98]. Traffic [GG98, IMI^{+}95, SC93, ALS96, Cac89, Cla94, ET91, EHS92, SC92]. Train8 [Sch93a]. Trainer [JT94a, Jaf94, JT94b, JT94c, Kov95]. trainers [Ano95-52]. Training [Ame95a, Ano95-68, BGGS95, Hic94, LT93, Wil95c, Ano95-52, Bla95c, Bla93b, BW94a, BW94b, BWC95a, BWC95b, Col93, Gra95, Int94i, Joh94, KKB95, RBK^{+}94, Siy94, TB93, Uta94, Wes94, WE96]. transaction [Spe87]. Transactional [Bil99]. Transactions [Ste96b]. Transfer [Ano92g, Gar94, Rob94, VH94b, EHS92, Hal93, Kas93, YmtEC99]. Transferring [VH94b]. Transfers [BH93]. transformation [Ass95b]. transforming [Cro94a, Cro95]. transit [ET91]. Transition [Hil97, SRI82b, Ano95f, SRI82c, Son92]. Transmission [IBM89c, IBM89a, IBM89b, IBM89f, IBM89d, IBM89e, IBM92b, AIS^{+}99, HG97, IBM88, IBM90, IBM91b, IBM91a, IBM92a, Pil90, Haa85]. transnacional [Rib95]. transparency [She85]. Transparent [GD98, LDG96, Dow86]. Transpersonal [Gac98]. Transport [BW97b, Wu97, sC86, Sin81, Xer81a, Xer81b]. Transportation [Anoxxc, GB96, CB95a, Ano95-72]. Traps [Amo99]. travees [CS95]. travel [BC95c, Goo96, Haw96b, VB95a, Woo96]. Traveler [End94, Twe94, SC94b]. travelers [Vel95]. Traveling [Dar92, Thi94]. Traveller [Cro94b]. Travelogue [Mal92]. travers [Min94b]. traves [Bro95b, Fre95]. treasures [Cla95d, DG94, Tes96, TB93]. TREC’98 [LM98]. tree [MM95]. Trends [LM98, Ano95-57, Bro95a, D^{+}86, Liu96]. Trial [KSM^{+}96]. trib [Med96]. Trib.com [ST95]. Tribune [Ano95y]. Tricks [B^{+}94, GG94]. trip [BB95, GH93, GH94a, GH94b, GH94c]. troubleshooter [JV94]. Troubleshooting [Ano93-40, Dig90a, Mil89, Mil91b, Mil92a, Mil93a, Mil96b, The94c]. Trudger [Ano96v]. truth [Mam96, Moo95]. Trying [LE95]. tsan [HSskw95]. tse [AJRmT95, HSskw95, LY95b, yTZ94]. Tsen [But95b]. tso [Neu98a]. Tsou [hLcL95]. Tsukuba [MKT98]. Tsung [kTlC94]. ttara [cYlC95]. tu [Edd95b]. TUBA [KF93]. tung [Hen95b]. Tuning [CP98, Hum98, Kil98]. tus [Car94a, Car94b]. tutor [LL95a]. Tutorial [Mol93b, Mur98, Ros86, SS94a, RN97]. tutorials [Pot94a, Uni96a]. TV [Ano94-37, Gar97, Pre95a, Sat95, TE99, WNS^{+}94]. Twenty [CP94a, DLC92b, DLC92a]. twenty-fifth [DLC92a]. twenty-third [DLC92b]. Twisted [Hew88, IEE90]. Twisted-Pair [IEE90, Hew88]. Two [BBN^{+}93, Ale95, Bau95, TMF97]. two-year [Ale95]. Type [IEE90].

U.S. [Ano94y, Ano94-32, Ano95l, Coh99, HKR91, HKN96, Mar99, Mer87, Pan95, TL96]. U.S.A [Ano81]. ui [SG94b, TI95]. UK [Int76b, Int76a, MD94, Pop94, Sch95c, Sch95d, Smi94a]. UKOLN [OD95]. UKOLUG [BU95]. ul [SG94b]. ultimate [Ano94x, BD95a, BLF99, HLB^{+}95]. Ultrix [SC94b]. umbral [TOL94b]. umjik [Edd94b]. Uncover [Ano94d, Not93c]. Underground [RdB95, Ano94-38, Int95n, Ros96, Woo95a]. understand [Ano96j, BKH95]. Understanding [And95b, Dut98, Kin95a, LC98, McL96a, Nof95c, Nof95d, Nof95e, Nof95f, Not94b, Obl92, RB99, SCS99, Wen95, Bra97, Bro96, Int85]. unexplored [OM95]. Unicode [Uni96a, Hiu96, Lei96, Ran96b]. Unified [Sun96]. Union [Gar97]. unique [NYS92, PN91, Per91c, Per91a, Per91b, Per92d, Per92e, Per92b, Per92c]. Unit [IEE90]. unite [Mad98]. United [HS94b, Ano94-27]. UniTrends [Ano99e]. Univers [BPR88, Bow90]. Universal [CSJ98, ZTLV98]. Universe [Fie93]. universities [Boy96, Mar96b, Mle95]. University [GV90a, GV90b, Gro90b, IC90, ILSB94, Mil92b, Mil94e, Owe93, Tri91, Wan94, Yan88, Zha94, Cor96, Ano93-41, Ano95-69, BD95b, Cam95b, GMB93, Sch93b, Sta88, Sta94b, Tri93]. Unix [Ano95-55, Ano97d, Hug94, Rah94, Fel93, And95a, Bou97, Bra93a, CMQ93, Cas93, Eck95, Gar97, GS96a, Her93, Hiu96, Hub96, KW89, Lei96, Mat93, Moh99b, Mui95, Pab96, Pad93, QW93, Rag93, Ran96b, San91, San94b, Sch94, SCJ96, SG95, SHN97, Ste90, Ste96b, Ste98b, Ste99a, Sun96, Vol97b, Wan96, AB94, DeR99, Fel94b, FK95a, Tan93, Web95]. Unix-based [Hug94]. Unleashed [BMG^{+}94, DR94, DR95, DR96, Sam95, Ano94-39, Ano95-43, Ano95-44, Ano96m, Bac94, EB^{+}97]. unrestricted [Gal94]. unsolved [Hof96b]. Unwanted [SG98]. Update [CNO94, Che93b, Edu95, Ewe95, Int94b, Hus94, Mil95b]. updates [Mat93]. updating [NM98]. Upgrading [DN99]. Uppsala [KSW99]. Urban [Bur95b, Gen95]. USA [Ano99d, Ano96h, BCB99, Lei93, Mot94, Sea90, TL93, Uni96b]. usability [PH97]. usage [Gra95, Hei95b].

Usare [LB94d]. USD [RO94a, RO94b]. USDA [Ano93i]. Use [Ano93p, BBH93, But94, CKPT99, E^{+}93, HKN96, LT93, Pan95, Sti94a, SMB94, Too95, Ack95b, Ano94q, Ano94-55, Ano95-27, Ano95-50, Ano96j, BA95, BKH95, Cla95a, Eme95b, Eng93b, Eng95a, Fre94, FE91, FE92, For94, Fin95, FF99, GH94d, GCS96, Gru94, HC94, Hf96, Inn96, Int95q, KB92, LT92, LJN94, Lev95, McC95, MW94e, Moh95, NN94, PW95b, Rot95a, SG93a, Sti94b, Tho95b, Tre95, Zaz95, Zha94]. used [Mai95]. useful [Hug95, NYS92, PN91, Per91c, Per91a, Per91b, Per92d, Per92e, Per92b, Per92c, Tee94]. Usenet [Har94f, Laz97, Har95c, Bos94, FK95a, Fri94, Gai93, HH97, Kw95, OT88, Pfa95f, TD91, TO92, VH94c]. USENIX [Ano98a, Ano99a, Ano98c]. User [Abo93, Ano92i, Ano93-31, Ano96v, Ari93, Blo94, Gar94, Gil93b, Gil93a, Gil94c, Har95c, Inf93, JG91, JL94, Kro92, Kro93, Kro94, LaQ89, LG93, Lin93b, Nor92, Not93b, Oti95, PCD92, Pfa94b, Tri91, Tri93, Zim95a, And92, Ano94i, Ano95-39, Ano95-57, Blo95, BD88, Cal90, CN95a, DFS98, Fol95b, GFF^{+}95, IBM89f, IBM92a, JD95a, JD95b, KIP99, Kas93, KHJG91, KHJ91, KB92, KHJ92, KG97, Kro95b, KK96b, Kur96b, MM91, MM92, MP93, NYS92, Owe93, PN91, Per91c, Per91a, Per91b, Per92d, Per92e, Per92b, Per92c, San94a, Sne95a, Sta95d]. user-responsible [KG97]. Users [Der94, Fly95, MB93, Neu94, RG93, Rya93, SS94b, SS94a, Wig95b, Ano95j, CB96a, Mat93, Mor94, Neo95, Pet86b, PWK94, SC94b, Thi93, Wil95c, Zha96]. uses [But95a, McC97, Mey95, Rya95i]. Using [Of 95, Ano97j, Bal91, Buc95b, Cas94, Chi95, Cos94a, Dig95, DS97, DH97, Eag94c, Eag94b, GD98, Gau94, Har96a, HM97, Hen99, Hon95, Ill91, JN95, KM96a, Ken95h, LP95, LeB93, LG95, Nat96a, Not93b, Pfa95f, Pik94, Pik95, PB95, Pik96, Pyr95, SB95, SRV95, Sta95f, oS94, TD91, TPBH94, VH94c, Wag95a, Wes97, WA98, dG93, Ano90d, Ano95-49, Atk99, Ben95c, BK96a, BD95b, BS94, DeR96, Dee95, DS96, FP95a, Haa85, HK91a, HK91b, Har95a, Hf96, Jai94, Jon94a, Kar96, KH95, Koh97, LS95, LS96, Lof95, LCP91, MWB95, Mal95b, McL96a, Mil92b, Mil93d, Mil94e, Nof95c, Nof95d, Nof95e, Nof95f, ODO95, Pao96, Par93, PHW^{+}93, RK95a, Ric95a, RCL96, Rud95, Rus94, RNW^{+}94, Sch95b, Sch90, SW93, She86, Sne95a, SV99, Sta98b, NAS90, WSA^{+}94]. USITS [Ano98a]. uso [Ubi95]. USSR [WMRS87]. usuarios [MPF95]. Utah [KKB95, RBK^{+}94, Uta94]. Utilities [Del93, LS98, VH94e]. utilization [Hel94]. Utilizing [Hil96, KSM98, HKR91]. utlandsk [Dam93]. UUCP [Gar94, OT88, TD91, TO92]. UVa [Byr95]. uz [Spi92].

V [Her93, Pad93, Rag93]. v.2 [SDUU85]. v5 [Ano96v]. v5-Multiplatform [Ano96v]. v8.10 [Ano99d]. Vacuum [Coo93a]. vadisanas [Spi92]. Valley [Tee94]. value [Sev95]. VanillaSearch [Ano96u]. variable [RT99]. variable-bit-rate [RT99]. Variables [HS98]. VariCAD [Ano99e]. vast [Ano94-40]. Vatican [MBC^{+}96]. VAX [Neu94, Owe93]. VAX/VMS [Neu94]. ve [Mal95a]. ve-rishtot [Mal95a]. Velocis [Ano97h]. Vendors [Osh93]. Venture [Hal95a]. verion [LS96]. verktyg [Dam93]. Verlag [Ano93-35]. vernetzt [Kla94]. Veronica [Ano93-42, Fel94b, FK95a, Gil94a, Gil94b, Har93a]. Version [Ano96v, Ano97h, Ano99d, Com92, CS93a, Com94a, CS97, CS98, Dig82, Pfa95d, Pfa95e, Sav93, SG95, Ano99e, Bou97, BL96, Gia95c, Gia95d, HSF94, IBM90, IBM91b, IBM91a, IBM92a, IBM92b, Kah94a, Per92e, Pit95b, Pit95c, Pot95, Smi93c]. Very [HCBO99]. vi [Fel93, Fel94b]. Via [Ano95-72, DGH93, Not93b, Ano94d, Ano94j, Leg95a, Ano96w, Anoxxc, Bag95, Bir90, BAG94, Bri94a, Bro94b, Mot94, QCM94a, Sin99a, SS96, Wes95, Win93, Wol98]. viaggio [Sal95d]. Victorian [Sta98a]. vid [Dam93]. Video [Car99a, CFL99, Ove94, Som94, Ano94-37, Ano94-40, Com95d, DH96, DKF^{+}95, HG97, Moh95, RT99, WLKS98, WNS^{+}94, WG94b]. video-conferencing [Moh95]. video-interactive-decoder [Tho95b]. Videoconferencing [TH96]. Vienne [Ano69]. View [Mer87, Sch95g]. Viewer [CTN95]. Viewpoint [Kah94b, Luk96]. views [CW93]. Village [Lin93c, Wie93]. Vinton [Ano93-43]. Virginia [Cam95b, Ano95c, Ano95-61, Ano96w]. Virtual [Bry95, CO99, DH99, Fow99, GKM^{+}99, LY97, Nic93, OKP97, Pel95b, Rhe94a, Rhe94b, Sav96, SWE98, Val93b, VE98, BC95b, Car95c, Con89, Coo96, Moo95, Spe93, SC94b, Shi96, SZ99, Vac96b, VFH98a, VFH98b]. virus [Ano88b, Bro89, Uni89, UU89]. Viruses [Den90, Bro89]. Visa [Bre95b, VFH98a, VFH98b]. Vision [Wei93]. visions [MW92, Rii82]. Visual [Ano96t, B^{+}96b, Fra96, Gre95c, Mar95b, Ano96v]. VisualExpresso [CO99]. Visualization [HM97, LD99]. Visualizing [GFF^{+}95, IE98, SYTM98]. VM [IBM88, IBM89a, IBM89b, IBM91b, IBM91a]. VMS [Mat93, Neu94, SC94b]. vocational [Wag95a]. Voice [Bla99f, Car99a, CFL99, Sch92a, TS88, sC86, New94b, New95c, New96, Sfe82, Yu93, Wri93]. voice/data [sC86, Sfe82]. Vol [CS97, CS98]. Volcano [Ano94-63]. Volume [Ano94-36, Bla95a, Bla98c, Bla98d, Bla99d, Buy99, Leo98g, Ste94, Ste96b, VH94a, VH94b, VH94c, VH94d, VH94e, VH94f, WS95, Ano94n]. Volumes [Bla98b]. voor [vdW95]. vos [Pig95]. voyage [Dar93b]. voyager [Int94k, HA95]. Voyages [VB95b]. voyeur [How95b, How95c]. VRML [BK96b, HE95b, HM97, Hen99, Vac96b]. vs [Bar95a, CRI95, CES^{+}99, Gil95a, Gil95b, Nol97]. VT100 [HKR91]. Vulnerabilities [Cul95]. Vulnerability [Gar97, Bro89].

W3D [McN95]. WAIS

[Fra95, Gil94a, Gil94b, Mor94]. Waite [And95a, Hen94]. Walk [Ric95c]. Walking [Tur95b, Tur95d]. walls [MP91]. wan [HSskw95, Cly88, Mil91a, Mil95c]. wanderers [Che96]. wang [nChCiY94, Edd95b, cH95, Neu98a]. want [Thi93]. Warehouse [Hac98]. warehousing [Sin99a]. warfare [Den99]. wargames [SR96]. Warner [Tim94]. Warning [Gar97]. Warp [Ive95, Mor95d, Sal95b, Tys95]. Warping [Sal95b]. was [Gar89]. Washington [Ano97a, Cim96, DLC92a, Max95c, Mot94]. Watch [Ano95i]. Water [Tee94]. wave [HMG^{+}95]. Way [BK99b, Fra93b, Fra94b, FAW95, FJ95, Har94e, Har94d, Har95a, Lic94a, Lic94b, Vin95b, Bal93b, BA95, DG95, Fra93a, Fra94a, Int95q, Kro95c, Lic95a, Lic95b, Net94c, Pot94a, Wag95b]. ways [Hof95a]. Weather [Noa96]. Weaves [Van96]. Weaving [BLF99]. Web [AJR95, Aik95, Mec95b, Ano96s, Ano96u, Ano96v, Aug96, BT96, Ber95a, BGGS95, Bow95, BWC95a, Bw96, Bro94b, Bro95b, Bud96, Cai94, Cam96, DO96, DR95, DR96, Eag94b, Fis95a, For95, Fra95, Gag95, GO95, HC95, HB98b, HN96, Hof95e, Hof95b, Kan95b, Kan95c, Kur96a, Lem95, Leo94b, MMJ95, Mar95a, McN95, Mor95d, Net96a, Nof96b, Pel95a, Pfa95g, Pfa96c, Ros96, SBGK94, Ste95a, Tur95b, Whi95, Med96, Hof95d, Ken95a, KM96b, LG99, PU94, AJR95, Abr98, Ano93-46, Ano95-40, Ano95-63, AW97, Atk99, BLB99, Bay99, Ber96a, BLF99, BDKK96, Bos94, Bra94b, Bri98, Bro94a, BCW95, Cha95c, CKTV96, CB96b, Dav96, DR94, Des94, EU95, Eag95b, ELB99, EE95b, EE95c, FK95a, Fol95a, Fou98, FD95, FP95b, GL99, Gut99, Hac98, Hal95a, Ham99, Har99a]. Web [HKB95a, HKB95b, HK95a, HB98a, Hen95a, HA95, HH98, HKN96, Hof96a, Int95g, IH98, KG98, Kan95a, Kw95, Kil98, LG98, LPJ^{+}94, LC98, Man96, Man95b, Mar96a, MA98, Mil99a, Min95, Moh99b, Nat96a, Ngu95, Nie98, NM98, Pao96, Pfa95b, Pfa95c, Pfa95d, Pfa95e, Pig95, Pow98, Pre95a, QRV95, Rau95, Sin99a, Ste97, Ste98a, SK98, TK99, Ter99, Tho94, Tur95d, Van96, VSW94, Wen95, Web95, Wes97, Wil99, YM96, YmtEC99]. Web-based [Har99a, LC98, Ter99]. Web/BBS [Ano96s]. WebMaker [RR94b]. Webmaster [Ano99e, LE95, SBM96, WC95d, You95]. Webs [Sac97]. WebSite [PM95a]. Websurfer [SS95a]. WebWriter [CB96b]. Week [Com84, Lem95, Phi95, Ran95b]. Weekend [Cer93]. Weird [Ric95c]. Welcome [BS93b, BS93a, Bad95, Car99b, SG95]. Welt [MW94a, MW94b, MW94c]. Weltweit [Kla94]. weltweiten [RS93b]. were [Mou96]. Werkzeuge [SBGK94]. West [Fra99, Hir97, Hir98, McL96b]. Western [Tri91, Tri93, Eve95]. Wheels [Ubo95]. Where [Des93b, HL96, Con95a]. which [Pow93, Web95]. White [GM94b, MG94c, ST89, SP94a]. Who [Ano95-70, FRG97]. Whole [Ano92i, Kro92, Kro93, Kro94, Who93, Rit93, Kro95b, KK96b]. Whose [OO95]. Wide [AJR95, Abr98, Aik95, Ano93-46, Mec95b, Aug96, Ber96a, BL97, BLF99, BGGS95, Bos94, Bow95, BWC95a, Bra94b, Bw96, Bri98, Bro94b, Bro95b, BCW95, Cah98, Cai94, Cam96, Dav96, DR95, DR96, EU95, Eag94b, Eag95b, EE95b, EE95c, FK95a, Fol95a, Fra95, FP95b, GO95, HC95, HKB95a, HKB95b, HK95a, HN96, Hen95a, Hof95e, Kan95b, Kan95c, LG98, Leo94b, LPJ^{+}94, Man96, Man95b, Mar95a, Mar96a, McN95, Min95, Mor95d, Nat96a, Ngu95, Nof96b, Pel95a, Pfa95g, SBGK94, Ste95a, Tho94, Tur95b, Tur95d, VSW94, Wen95, Whi95, Wil99, YM96, Bow90, BS94, Cac89, GS95c, Hof95d, Hof95b, Ken95a, LE95, Net96a, SLJ94, Sin81, TJ97, DR94, Des94, HA95, Kw95, QRV95]. wide-area [GS95c]. Wild [Ano93-38]. Wilde [Wil99]. wilderness [BBN^{+}93]. wilds [Ehn94]. Will [Ano93g, Dra95, Ken90, LeG95b, Shu93, SE96]. Wily [CB94c]. Window [DKF94, MP92, Int93f, Sha91a, Hiu96]. Windows [AH95a, Bea95, BGS94, BGS95, Bud96, Buy96, Dat96, Eme95a, ELS95, ELO95, GBS95, Gau94, Gri95, HB95, Hes95, Hon95, JW95, Ken95d, Ken95e, KKB95, Kro95b, McB96, MPG94b, Mis95, Moo96, Neo95, NR96, Par95, Pfa95a, Pik96, Rau95, Sua94, SS94a, Sal96, Wya95c, You94, DHS95, LE96, Ano93-44, Ano94-42, CS97, CSJ98, DK94a, Eng94a, Fra94b, FJ95, GS95a, Gia95d, Gut99, HAS95, HKB95b, HK95a, Hud94, Int93d, Jam96b, Ken95g, Lic94b, Lic95b, Mor95b, Net95c, New96, Pfa95e, Pit95c, Sne95c, SRV95, kTlC94, WMS96, Win96a]. WinGopher [Ano93-45]. Winning [Ano93-37]. Winsock [Rob95]. WinWord [Ros96]. WIPO [Gar97]. Wired [Ano95-71, Car93b]. Wireless [Bla96a, Bla98f, Bla99e, DM95, Goo97, Wes98]. Wiretap [Ano92d, Dar93b, Gum93a]. Wiring [Ani93b, Dig90b]. Wisconsin [Eve95]. Wissenschaft [AB94]. within [Ass94b, Ano94e]. Without [LE95, Ram95, RM96, ST89]. Wizards [HL96]. WLN [Wes94]. WLN.COM [Sch93a]. Women [Bau95, FH95, ACE^{+}95, Kug95]. Won’t [Wig95a]. Word [Roa95a]. Wordlink [LR93b, LaQ94b]. words [Str94]. Work [BBH93, Eag95a, Wal93, Gol92, KP95, Mas97, She85]. Workbook [Moh99b, SRI82b, Bla95c, BW94a, BW94b, BWC95a, BWC95b, Bur96, Gad94, Int94i, SRI82c].

Workflow [Moh99a]. Workgroup [Ano96v, DO96]. Working [JG91, LG93, Int95o, LM98, MY94]. workload [AW97]. Works [Com97, Edd94a, Com95a, Edd94b, Edd95a, Edd95b, WMS96]. Workshop [ACM98, BCB99, JT94a, KW98, KSW99, MSD90, Wor94, Ano93t, BBST96, Cly93, Ill91, Jaf94, JT94b, JT94c, PT95a, PT95b, Pee96b, VT97]. Workshops [BH94b, Str94]. Workstation [Cal91]. Workstations [WA98]. Worl [Bro95b]. Worl-Wide [Bro95b]. World [AJR95, Aik95, Int76a, Ano88a, Ano93-28, Ano93-38, Ano93-44, Ano94-41, Ano95-64, Ano96x, Aug96, BL97, BGGS95, Bow95, BWC95a, BB93, Bro94b, Cai94, Cam96, Car93b, Chr98, DR95, DR96, Eag94b, Ert95, Far93b, Fle94, Fra95, GO95, HC95, Har95b, HKB95a, HKB95b, HK95a, HN96, Hof95e, Int76b, Jos95a, Jos95b, Kan95b, Kan95c, KPS95, Leo94b, LE95, Mar95a, Mat95b, Mor95d, Nof96b, Pel95a, Pfa95g, dPW94, Ran94, Rii82, SBGK94, Ste95a, Tur95b, Van96, VVB85, Whi95, WS94, Ano81, Int92f, Ano94-63, Mec95b, GP95a, GP96a, HE95b, Hof95d, Hof95b, Mec96, Int94l, Jam96a, Jam96b, Ken95a, KKB95, Koc93, Leo94a, Mad98, MW94c, Mar99, MP91, Mec95a, Net96a, NST96, Nov94, NM98, Pot93, Pot94a, Pot94b, Pri95, RBK^{+}94, RML^{+}94, RK95b, SC94b, SG95, Sta95b]. world [TG95, Abr98, Ano93-46, BR95, Ber96a, BLF99, Bos94, Bra94b, Bri98, BCW95, Dav96, DR94, Des94, EU95, Eag95b, EE95b, EE95c, FK95a, Fol95a, FP95b, Hen95a, HA95, Kw95, LG98, LPJ^{+}94, Man96, Man95b, Mar96a, McN95, Min95, Nat96a, Net86, Ngu95, QRV95, Tho94, Tur95d, VSW94, Wen95, Wil99, YM96]. World-Wide [AJR95, BL97, Bro94b, Cai94, DR95, DR96, Fra95, Kan95b, Kan95c, Pel95a, LE95, DR94, Des94, HA95, Kw95, QRV95]. WorldLink [LR93d]. worlds [BC95b, BBN^{+}93, Car95c, VE98]. Worldwide [MKT98, RS93b, Hum94, Kac96, MO92, MBC^{+}96, Mue99b]. WorldWideWeb [Fly95]. Worm [Spa89a, Spa89b, Den89, Gar89, Spa88]. Worms [Den90]. Worth [Bak93a, Bak93b, Bak93c, Lan94]. Wrapper [AK97]. wrecker [SO94]. Writing [RR94b, Har96a]. wu [yTZ94]. WWCA’98 [MKT98]. WWW [BGGS95, AJRmT95, Cha95a, Fol95a, Gil94a, Gil94b, KJTS96, Nof95f, RR94b, Wil99, cYlC95]. WWWhither [Dam97]. Wyoming [Gre95a].

X [Ano96u, Ano99e, Pet89, DKF94, Hiu96, Int93f, MP92, Wan96]. x-kernel [Pet89]. X-Window [Hiu96]. X.400 [CHHL91]. X.500 [BK91, HK91a, HK91b]. X11 [MP92]. XDSL [Bla98f]. Xerox [Dig82]. XINU [Com87a, Com87b]. XML [Bou98, Hom99, LO99]. XTI [Ste98b]. XXI [Fre95].

Y2K [Web99]. Yahoo [Yah96]. yang [But95b]. yao [cH95]. year [Ale95, Int81, Ken90, EE99, Van96, Web99, YY98]. yeh [Neu98a]. Yellow [BCW95, HS94d, HS95c, MG94a, Mec95b, Bw96, HS94b, HS95b, Hah96b, Mec96, MG94b, NC95b, NST96, PDP96, Pol96, SD96a]. yesterday [BC95a]. Yisreeli [Mal95a]. yogoshu [MY94]. yohaeng [TI95]. yomu [Edd95a]. York [Hay94b]. Young [Ano95-71, Bro94a]. you’re [Ric95a]. Yourself [Lem95, Par94b, Ran94, Ran95b, Via94, Wei95a]. yu [cH95]. Yucatan [Ano95o]. yung [But95b, kTlC95]. yuzazu [KM95].

Z [Ano93b, Bla95d, Rit93, Sim93]. Z39.50 [MH94]. ZD [ZD 95]. Zealand [Anoxxb, CG95]. Zeitschriften [Ble95]. Zen [Keh92, Keh93, Keh94, KMM95, Keh96]. zones [BM96a]. zoo [Bos94].

References

Adam:1998:ECT

[A^{+}98] Nabil Adam et al. Electronic Commerce: Technical, Legal and Business Issues. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-949082-5. xiv + 209 pp. LCCN HF5548.32.E374

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0139490825.html.

Arnold:1994:IPT

[AA94] Stephen E. Arnold and Erik S. Arnold. Internet 2000: the path to the total network. Infonortics in-depth briefings. Infonortics, Calne, Wiltshire (England), 1994. ISBN 1-873699-08-5. vi + 197 pp. LCCN ????

Ambite:1998:ASC

[AAB^{+}98] José Luis Ambite, Naveen Ashish, Greg Barish, Craig A. Knoblock, Steven Minton, Pragnesh J. Modi, Ion Muslea, Andrew Philpot, and Sheila Tejada. ARIADNE: a system for constructing mediators for Internet sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2): 561–563, June 1998. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arnold:1996:PIN

[AADF96] Stephen E. Arnold, Erik S. Arnold, Ulla De Stricker, and Richard D. Fiust. Publishing on the Internet: a new medium for a new millennium. Infonortics in-depth briefings. Infonortics, Calne, Wiltshire, England, 1996. ISBN 1-873699-28-X. iv + 159 pp. LCCN 100 A76 1996.

Adam:1997:SFT

[AAS^{+}97] Nabil Adam, Baruch Awerbuch, Jacob Slonim, Peter Wegner, and Yelena Yesha. The science of future technology: Globalizing business, education, culture through the Internet. Communications of the ACM, 40(2):115– 121, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-2/
p115-adam/.

Alex:1994:UCI

[AB94] Wulf Alex and Gerhard Bernor. Unix, C und Internet: moderne Datenverarbeitung in Wissenschaft und Technik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57881-1 (Berlin), 0-387-57881-1 (New York). xix + 535 pp. LCCN TA403.N39 1995.

Abbate:1994:AIH

[Abb94] Janet Abbate. From ARPA-NET to INTERNET: a history of ARPA-sponsored computer networks, 1966–1988. Thesis (Ph.D.), University of Pennsylvania, Philadelphia, PA, USA, 1994. vii + 178 pp.

Abbate:1999:II

[Abb99] Janet Abbate. Inventing the Internet. Inside technology. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-01172-7 (hardcover). viii + 264 pp. LCCN TK5105.875.I57 A23 1999.

URL http://lcweb.loc.
gov/catdir/toc/98047647.
html.

Aboba:1993:OUE

[Abo93] Bernard Aboba. The Online User’s Encyclopedia: Bulletin Boards and Beyond. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62214-9. xxiv + 806 pp. LCCN QA76.9.B84 A26 1993. US$32.95.

Abrams:1998:WWW

[Abr98] Marc Abrams. World Wide Web: beyond the basics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-954785-1. xi + 483 pp. LCCN TK5105.888.A254

Accardi:1995:II

[Acc95] Joe Accardi. An introduction to the Internet, 1995. 1 videocassette (85 min.).

Anderson:1995:EWD

[ACE^{+}95] Linda Anderson, commentator, Sara S. Chapman, Beth Engeler, Karen Kelly, commentator, and Nancy Woodhull. Education for women in developing countries; stones from the river; getting supplies on the internet; women politicians and the press; happy birthday, 51%, 1995. 1 sound cassette (25 min.).

Ackenhusen:1995:SPT

[Ack95a] John Ackenhusen. Signal Processing Technology and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2469-2. 530 pp. LCCN TK5102.9 .S547 1995. US$59.95.

Ackermann:1995:LUI

[Ack95b] Ernest C. Ackermann. Learning to use the Internet: an introduction with examples and exercises. Franklin, Beedle, and Associates, Wilsonville, OR, USA, 1995. ISBN 0-938661-92-2. x + 358 pp. LCCN TK 5105.875 I57 A25 1995.

ACM:1996:PPF

[ACM96] ACM, editor. PODS ’96. Proceedings of the Fifteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, PODS 1996, Montréal, Canada, June 3–5, 1996, volume 15. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996. ACM order number: 475960.

ACM:1998:AWJ

[ACM98] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing. ACM Press, New York, NY 10036, USA, 1998. ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
program.html. Possibly un-

published, except electronically.

ACM:1999:SPO

[ACM99a] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. LCCN ????

ACM:1999:SPCd

[ACM99b] ACM, editor. SIGGRAPH 99. Proceedings of the 1999 SIGGRAPH annual conference: Conference abstracts and applications, Computer Graphics. ACM Press, New York, NY 10036, USA, 1999. ISBN 0-201-48560-5. ISSN 1069-529X. LCCN T385 .S54

                                        1999. URL http://info.
                                        acm.org/pubs/contents/proceedings/

graph/. ACM order number

428990.

Afshar:1995:PSN

[Afs95]  Amir Afshar.  Princi-

ples of Semiconductor Network Testing. Butterworth-Heinemann, Boston, MA, USA, 1995. ISBN 0-7506-9472-6. xiv + 213 pp. LCCN TK7874 .A339 1995. US$85.95.

Albertson:1995:DGI

[AFZ95] Rick Albertson, Jeffrey Fine, and Mike Zender. Designer’s guide to the Internet. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-229-0 (paperback). xv + 2–417 pp. LCCN TK5105.875.I57 A4 1995 Ref.

Agarwal:1999:BIN

[Aga99]  P. K. Agarwal.  Build-

ing India’s national Internet backbone. Communications of the ACM, 42(6): 53–58, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-6/
                                        p53-agarwal/.

Angell:1994:EEM

[AH94] David Angell and Brent Heslop. The Elements of E-Mail Style. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-62709-4. xii + 157 pp. LCCN HE6239.E54 A53 1994. US$12.95. On how to write e-mail effectively ….

Angell:1995:MDW

[AH95a] David Angell and Brent Heslop. Mosaic for Dummies: Windows edition. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-242-2. 356 pp. US$19.99. Uses surfing jargon to teach users to catch a wave (fundamentals), customize your surfboard (change the software configurations), power surf (using advanced techniques), creating your own home page, and surfing after dark (entertaining WWW sites).

Angell:1995:IBC

[AH95b] David Angell and Brent D.

Heslop. The Internet business companion: growing your business in the electronic age. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40850-3. xiii + 242 pp. LCCN HD30.335 .A54 1995.

AmnestyInternational:1988:HRE

[AHH88] Human rights education: the fourth R, page various, 1988. ISSN 1063-8768. Amnesty International USA’s National Steering Committee on Human Rights Education, Jonesboro, AR, USA.

Anderson:1990:SRR

[AHS90] David P. Anderson, Ralf Guido Herrtwich, and Carl Schaefer. SRP: a resource reservation protocol for guaranteedperformance communication in the Internet. Report UCB/CSD 90-562, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, February 1990. 26 pp.

Ahuja:1996:NIS

[Ahu96] Vijay Ahuja. Network and Internet security. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-045595-1. xix + 324 pp. LCCN TK5105.59.A39 1996.

Aikat:1995:ACE

[Aik95] Debashis Aikat. Adventure in cyberspace: exploring the information content of the World Wide Web pages on the Internet. Thesis (Ph.D.), Ohio University, Athens, OH, USA, August 1995. xii + 203 pp.

Amer:1999:NCG

[AIS^{+}99] Paul D. Amer, Sami Iren, Gul E. Sezen, Phillip T. Conrad, Mason Taube, and Armando Caro. Networkconscious GIF image transmission over the Internet.

Computer Networks (Amsterdam, Netherlands: 1999), 31 (7):693–708, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/7/2101.pdf.

Abraham:1995:WEB

[AJR95] Ralph Abraham, Frank Arnoud Jas, and Willard Russell. The WEB empowerment book: an introduction and connection guide to the Internet and the World-Wide Web. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94431-1. xxxi + 192 pp. LCCN TK5105.888.A25 1995. US$24.95.

Abraham:1995:WCC

[AJRmT95] Ralph Abraham, Frank Arnoud Jas, Willard Russell, and Chien ming Tseng. WWW chao chi shou tse. Po shuo ku wen “ching tien hsi lieh”. Po shuo ku wen yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-8948-59-X. x + 214 pp. LCCN ISBN 0-471-11629-7.

Ashish:1997:WGS

[AK97] Naveen Ashish and Craig A.

Knoblock. Wrapper generation for semi-structured Internet sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):8–15, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Albitz:1992:DBN

[AL92] Paul Albitz and Cricket Liu. DNS and BIND in a Nutshell. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-010-4. xviii + 381 pp. LCCN TK5105.875.I57 A43 1992. US$29.95. Discusses one of the Internet’s fundamental building blocks: the distributed host information database that’s responsible for translating host names into addresses, routing mail and other services. Intended for system administrators, covers the technical stuff but is readable.

Albitz:1994:DBN

[AL94] Paul Albitz and Cricket Liu. DNS and BIND in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-010-4. xxviii + 381 pp. LCCN TK5105.875.I57A43 1992.

Andrieu:1995:IL

[AL95] Olivier Andrieu and Denis Lafont. Internet et l’entreprise. Eyrolles, Paris, France, 1995. ISBN 2-212-08906-6. xviii + 395 pp. LCCN ????

AHML:1996:GJT

[AL96] Arlington Heights Memorial Library and Library Cable Network. Getting a job through the Internet, 1996. 1 videocassette (60 min.).

Albitz:1993:DBN

[Alb93] Paul Albitz. DNS and BIND in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, corrected edition, March 1993. ISBN ???? xxviii + 381 pp. LCCN ????

Alexander:1995:SIL

[Ale95] Linda H. (Linda Hackney)

Alexander. A study of Internet listservs as postteleconference support to faculty at community colleges and two-year institutions. Thesis (Ph.D.), University of North Texas, Denton, TX, USA, May 1995. ix + 250 pp.

Alexander:1996:PCG

[Ale96] Dey Alexander. Philosophy in cyberspace: a guide to philosophy-related resources on the Internet. Philosophy Documentation Center, Bowling Green State University, Bowling Green, OH, USA, 1996. ISBN 0-912632-59-3 (spiral-bound). ???? pp. LCCN ????

Allen:1994:IDL

[All94] Chris Allen. The internet, distribution lists, and gatekeeping. Technical report, ????, ????, 1994. 21 pp.

Allison:1995:LGI

[All95] G. Burgess Allison. The lawyer’s guide to the Internet. Section of Law Practice Management, American Bar Association, Chicago, IL, USA, 1995. ISBN 1-57073-149-7 (paperback). xvi + 347 pp. LCCN TK5105.875.I57 A55 1995. US$29.95.

Allison:1996:IBP

[All96] Wayne Allison. The Internet business primer. Small business sourcebooks. Sourcebooks, Naperville, IL, USA, 1996. ISBN 1-57071-065-1 (hardcover), 1-57071-064-3 (paperback). viii + 116 pp. LCCN HD30.37.A38 1996.

Alspach:1995:IEQ

[Als95] Ted Alspach. Internet Email quick tour: sending, receiving and managing your messages online. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-220-8. xviii + 122 pp. LCCN TK5105.875.I57 A438 1995. US$14.00.

Akyildiz:1996:BRR

[ALS96] Ian F. Akyildiz, Jörg Liebeherr, and Debapriya Sarkar. Bandwidth regulation of realtime traffic classes in internetworks. Computer Networks and ISDN Systems, 28 (6):855–872, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=6&aid=1487.

AALLRIC:1995:TKL

[Ame95a] American Association of Law Libraries. Research Instruction and Caucus. Training kit: legal resources on the Internet. West Publishing, St. Paul, MN, USA, 1995. 194 pp.

AASLISS:1995:ISS

[Ame95b] American Association of School Librarians. Independent Schools and Section. Independent schools selection [i.e. section] program and business meeting Internet in school library media programs, 1995. 2 sound cassettes.

ARCITB:1995:GQIa

[Ame95c] Ames Research Center. Imaging Technology Branch. Global quest: the Internet in the classroom, 1995. 1 videocassette (10 min., 40 sec.).

ARCITB:1995:GQIb

[Ame95d] Ames Research Center. Imaging Technology Branch. Global quest: the Internet in the classroom, 1995. 1 videocassette (16 min.).

Amoroso:1999:IDI

[Amo99] Edward G. Amoroso. Intrusion Detection: An Introduction to Internet Surveillance, Correlation, Trace Back, Traps, and Response. Intrusion.Net Books, Sparta, NJ, USA, 1999. ISBN 0-9666700-7-8. 218 pp. LCCN TK5105.59 .A465 1999. US$49.95.

Ang:1996:CIS

[AN96] Peng Hwa Ang and Berlinda Nadarajan. Censorship and the Internet: A Singapore perspective. Communications of the ACM, 39(6): 72–78, June 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228520.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228520.html.

Andres:1992:CSC

[And92] Yvonne Marie Andres. Cerf ’n safari: a classroom users’ guide to the Internet. California Education and Research Federation Network, San Diego, CA, USA, 1992. iv + 36 pp.

Andres:1993:EO

[And93] Yvonne Marie Andres. Education On-Line. The Executive educator, 15(6):21–??, June 1, 1993. CODEN EX-EREA. ISSN 0161-9500.

Andrieu:1994:IGC

[And94] Olivier Andrieu. Internet:

guide de connexion. Eyrolles, Paris, France, 1994. ISBN 2-212-08900-7. vi + 239 pp. LCCN ????

Anderson:1995:WGU

[And95a] Bart Anderson. The Waite Group’s UNIX communications and the Internet. SAMS Publishing, Indianapolis, Ind., third edition, 1995. ISBN 0-672-30537-2. xxv + 817 pp. LCCN QA76.76.O63 A5 1995. Partial Contents: UNIX communications fundamentals — UNIX mail — UNIX news: USENET — UNIX file transfer: UUCP — Internet communications — Appendixes.

ACA:1995:UI

[And95b] Andrew Cochran Associates.

Understanding the Internet, 1995. 1 videocassette (ca. 50 min.).

Andrieu:1995:AST

[And95c] Olivier Andrieu. Annuaire de la science et de la technologie sur Internet [1995]. Technical report, ADIT, Strasbourg, France, 1995. 340 pp.

Andrieu:1995:IGC

[And95d] Olivier Andrieu. Internet:

guide de connexion. Eyrolles, Paris, France, 2e mise à jour. edition, 1995. ISBN 2-212-08900-7. 256 pp. LCCN ????

Anderson:1999:FSA

[And99] Don Anderson. FireWire System Architecture: IEEE 1394a. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-48535-4. 544 pp. LCCN TK7895.B87 A52 1999. US$39.95.

Anixter:1993:NPG

[Ani93a] Anixter Bros., Inc., 4711 Golf Road, Skokie, IL 60076, USA, Tel: (312) 677-2600. Anixter Networking Products Guide, 1993.

Anixter:1993:WSP

[Ani93b] Anixter Bros., Inc., 4711 Golf Road, Skokie, IL 60076, USA, Tel: (312) 677-2600. Anixter Wiring Systems Product Catalog, 1993.

Anonymous:1969:MCC

[Ano69] Anonymous. Méthodes ä chemin critique. Network methods. Netzwerktechniken. Actes du congrès Internet I, Vienne, 1967. Dunod, Paris, France, 1969. xxvi + 598 pp. LCCN ????

Anonymous:1974:PMF

[Ano74] Anonymous, editor. Project management: Fourth Internet Congress, Paris, Sept. 30–Oct. 3, 1974. AFCET, Paris, France, 1974. ISBN ???? LCCN ????

Anonymous:1979:AIC

[Ano79]  Anonymous, editor.  Ab-

stracts of the 6th Internet Congress, 1979. VDI-Verlag, Dusseldorf, Germany, 1979. ISBN ???? LCCN ???? Anonymous:1981:WPM

[Ano81] Anonymous, editor. The world of project management: PMI/Internet Joint Symposium, Boston, MA., U.S.A., Sept. 28–30, 1981: 1981 proceedings. Project Management Institute, Drexel Hill, PA, USA, 1981. ISBN ???? LCCN ????

Anonymous:1988:ICC

[Ano88a] Anonymous, editor. Internet 88: from conception to completion; proceedings of the 9th World Congress on Project Management, September 4th – 9th 1988, Glasgow. NCC Publications, Manchester, UK, 1988. ISBN ???? LCCN ????

Anonymous:1988:IVA

[Ano88b] Anonymous. Internet virus aftermath: Is tighter security coming? Data communications, 17(14):52–??, December 1, 1988. CODEN DA-CODM. ISSN 0363-6399.

Anonymous:1989:IRR

[Ano89a] Anonymous. Internet resources (or resource??) guide. Technical Report 1175, NSF Network Service Center, Cambridge, MA, 1989. various pp.

Anonymous:1989:ITF

[Ano89b] Anonymous. Internet task force cleans up TCP/IP specification. Data communications, 18(12):17–??, September 21, 1989. CODEN DA-CODM. ISSN 0363-6399.

Anonymous:1989:MHR

[Ano89c] Anonymous. Masterlist of human rights organizations and serial publications. Technical report, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1989. 107 pp. Human Rights Internet reporter. Supplement; v. 13 (1989–1990).

Anonymous:1990:CIS

[Ano90a]  Anonymous.  Commer-

cial Internetworking services for Lan-to-LAN connections. Data communications, 19(2): 46–??, February 1, 1990. CO-DEN DACODM. ISSN 0363-6399.

Anonymous:1990:I

[Ano90b] Anonymous. Internet 101, 1990. 1 videocassette (21 min.).

Anonymous:1990:ILC

[Ano90c]  Anonymous.  Internet-

accessible library catalogs and databases. Technical report, ??, ??, 1990. 44 pp.

Anonymous:1990:NSP

[Ano90d]  Anonymous.  NASA STI

Program Coordinating Council, eighth meeting, June 2, 1992: using the Internet. NASA technical memorandum NASA-TM-108019, National Aeronautics and Space Administration, Washington, DC, USA, 1990. various pp.

Anonymous:1990:SJI

[Ano90e] Anonymous. Shamash: the Jewish Internet consortium, 1990.

Anonymous:1990:TT

[Ano90f] Anonymous. Teacher-Friendly technology. Learning (Palo Alto, CA), 19(2):82–??, September 2, 1990. ISSN 0090-3167.

Anonymous:1991:CEH

[Ano91] Anonymous. Cruising the electronic highways: librarian’s guide to the Internet, Boston Chapter, SLA Internet meeting [held] November 21, 1991, 1991. 1 videocassette (190 min.) sd. + b and w + + instruction materials.

Anonymous:1992:BLS

[Ano92a] Anonymous. Biosis Life Science Network and Orbit now available on the Internet. Information intelligence online newsletter, 13(10):4–??, October 1, 1992. CODEN IIONDK. ISSN 0194-0694.

Anonymous:1992:ILL

[Ano92b] Anonymous. Internet, list of lists. Technical report, ????, ????, 1992. 62 pp.

Anonymous:1992:IRG

[Ano92c] Anonymous. Internet resource guide. Technical report, NSF Network Service Center, Cambridge, MA, 1992. various pp.

Anonymous:1992:IWa

[Ano92d] Anonymous. Internet wiretap, 1992.

Anonymous:1992:ICI

[Ano92e] Anonymous. ISO/CCITT and Internet management: coexistence and interworking strategy. X/Open guide. X/Open Company, Reading, UK, 1992. ISBN 1-872630-67-7. xvi + 62 pp. LCCN ????

Anonymous:1992:LLS

[Ano92f] Anonymous. List of lists: special interest group mailing lists available on the Internet. Technical report, ????, ????, 1992. 94 pp.

Anonymous:1992:OOF

[Ano92g] Anonymous. OCLC offers FTP support for transfer of records over the Internet. Information intelligence, online libraries, and microcomputers, 10(11):5–??, November 1, 1992. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1992:RIA

[Ano92h] Anonymous. Resources on the Internet: Archie and Gopher. Information intelligence, online libraries, and microcomputers, 10(10):1–??, October 1, 1992. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1992:RWI

[Ano92i] Anonymous. [review] The Whole Internet User’s Guide & Catalog by Ed Krol. Information intelligence, online libraries, and microcomputers, 10(10):10–??, October 1, 1992. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:BINa

[Ano93a] Anonymous. Brief Internet & NREN glossary: Part I (A-L). Information intelligence, online libraries, and microcomputers, 11(5):1–??, May 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:BINb

[Ano93b] Anonymous. Brief Internet & NREN glossary: Part II (M-Z). Information intelligence, online libraries, and microcomputers, 11(6 / 7):1– ??, June 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:BII

[Ano93c] Anonymous. A business interest in Internet: Sprint-MCC ease commercial links. Information Week, 430:71– ??, June 21, 1993. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1993:CID

[Ano93d] Anonymous. Cause institution database available on the Internet. Information intelligence, online libraries, and microcomputers, 11(3):8– ??, March 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:CCN

[Ano93e] Anonymous. Clarines: Current news over the Internet. Information intelligence, online libraries, and microcomputers, 11(8 / 9):4–??, September 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:CBD

[Ano93f]  Anonymous.  Commerce

Business Daily available on the Internet. Information intelligence, online libraries, and microcomputers, 11(6 / 7):11–??, June 1, 1993. CO-DEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:EWB

[Ano93g] Anonymous. Edgar will become available over the Internet. Information intelligence, online libraries, and microcomputers, 11(11):12– ??, November 1, 1993. CO-DEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:ALN

[Ano93h] Anonymous. From ARPA-NET to LB452: Nebraska’s pre K–12 Internet story. Technical report, Nebraska Dept. of Education, Division of Education Services, Lincoln, NE, USA, May 3, 1993. 6 pp.

Anonymous:1993:GUI

[Ano93i] Anonymous. Guide to the USDA Internet. U.S. Dept. of Agriculture, Office of Information Resources Management, National Computer Center, Telecommunications and Applications Services Center, Washington, DC, USA, version 1.0 edition, August 1993. viii + 63 + 96 pp.

Anonymous:1993:GOG

[Ano93j] Anonymous. Guidon: OCLC’s GUI for electronic journals on the Internet. Information intelligence, online libraries, and microcomputers, 11(3):8– ??, March 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:HH

[Ano93k] Anonymous. Hackers highway? Mediaweek, 3(45):10– ??, November 8, 1993. ISSN 1055-176X.

Anonymous:1993:IC

[Ano93l] Anonymous. Interactive communications. Editor & publisher, 126(48):14–??, November 27, 1993. ISSN 0013-094X.

Anonymous:1993:IAB

[Ano93m] Anonymous. Internet access to the British Library Document Supply Centre. Information intelligence, online libraries, and microcomputers, 11(2):8–??, February 1, 1993. CODEN IIOMEI. ISSN 0737-

IBJ:1993:IBJ

[Ano93n] The Internet Business Journal: commercial opportunities in the networking age, 1993. ISSN 1192-8646. Strangelove Press, Ottawa, ON, Canada.

Anonymous:1993:IBR

[Ano93o] Anonymous. Internet Business Report (IBR). Information intelligence, online libraries, and microcomputers, 11(10):10–??, October 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:ICL

[Ano93p] Anonymous. Internet connections: a librarian’s guide to dial-up access and use. Information intelligence, online libraries, and microcomputers, 11(8 / 9):10–??, September 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:IN

[Ano93q] Anonymous. The Internet Navigator. Information intelligence online newsletter, 14(11):13–??, November 1, 1993. CODEN IIONDK. ISSN 0194-0694.

Anonymous:1993:IPI

[Ano93r] Anonymous. Internet primer for information professionals: a basic guide to Internet networking technology [review]. Information intelligence, online libraries, and microcomputers, 11(5):11–??, May 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:IPL

[Ano93s] Anonymous. An Internet primer for librarians and educators: a basic guide to Internet networking technology. Information intelligence, online libraries, and microcomputers, 11(1):10–??, January 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:IWA

[Ano93t] Anonymous. Internet workshop: academic computing. Technical report, Norfolk State University (VA). Academic Computing Center and Academic computing, Norfolk, VA, USA, 1993. 21 pp.

Anonymous:1993:IYG

[Ano93u] Anonymous. The Internet:

your guide to the information super highway. Online access. special issue, v. 8, no. 6, Online Access Pub. Group, Chicago, IL, USA, October 1993. 102 pp.

Anonymous:1993:IBJ

[Ano93v] Anonymous. The Internet/ NREN Business Journal. Information intelligence, online libraries, and microcomputers, 11(3):10–??, March 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:IBC

[Ano93w] Anonymous. Introduction:

Big city syndrome and the Internet. UNIX/world, 10(12): 64–??, December 1, 1993. ISSN 0739-5922.

Anonymous:1993:NNG

[Ano93x] Anonymous. Networking for the new generation NREN/ Internet progress report and policy implications, 1993. 1 sound cassette.

Anonymous:1993:NPC

[Ano93y] Anonymous. The newsletter: Prostate cancer: Internet: flu shots: recliners. U.S. news and world report, 115(15):79– ??, October 18, 1993. CO-DEN XNWRAV. ISSN 0041-5537.

Anonymous:1993:OLS

[Ano93z] Anonymous. OCLC library school program offered in Europe. Internet speeds documentation delivery. OCLC Newsletter, 203:15–??, May 1, 1993. ISSN 0163-898X.

Anonymous:1993:I

[Ano93-27] Anonymous. On Internet 1993. Information intelligence, online libraries, and microcomputers, 11(5):12–??, May 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:TW

[Ano93-28] Anonymous. On top of the world. Contract design, 35 (9):54–??, September 1, 1993. ISSN 1053-5632.

OnlineAccess:1993:OAS

[Ano93-29] Anonymous. Online access special issue, the Internet, 1993.

Anonymous:1993:REI

[Ano93-30] Anonymous. Resources for economists on the Internet, 1993. ISSN 1081-4248.

Anonymous:1993:RPU

[Ano93-31] Anonymous. Revised PDQ User Guide available on Internet. Information intelligence online newsletter, 14(10):11– ??, October 1, 1993. CODEN IIONDK. ISSN 0194-0694.

Anonymous:1993:SIa

[Ano93-32] Anonymous. Sailing on the Internet. Information intelligence online newsletter, 14 (8 / 9):1–??, September 1, 1993. CODEN IIONDK. ISSN 0194-0694.

Anonymous:1993:SIb

[Ano93-33] Anonymous. The scientist — (internet). Information intelligence online newsletter, 14(8 / 9):9–??, September 1, 1993. CODEN IIONDK. ISSN 0194-0694.

Anonymous:1993:SIC

[Ano93-34] Anonymous. Sound idea for computers. Broadcasting & cable, 123(12):27–??, March 22, 1993. ISSN 0007-2028.

Anonymous:1993:SOT

[Ano93-35] Anonymous. Springer-Verlag offers table of contents preview on Internet. Information intelligence, online libraries, and microcomputers, 11(5):6– ??, May 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:SCE

[Ano93-36] Anonymous. Sudan: chronology of events, 1955–1993. Occasional papers, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, May 1993. 23 pp.

Anonymous:1993:TWB

[Ano93-37] Anonymous. TCP/IP winning the backbone battle. Datamation, 39(19):26–??, October 1, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TWW

[Ano93-38]  Anonymous.  Technology:

The wild world of Internet. Time, 142(24):62–??, December 6, 1993. CODEN TYMEA9. ISSN 0040-781X.

Anonymous:1993:TNI

[Ano93-39] Anonymous. Telecommunications, networking and Internet glossary. Information intelligence, online libraries, and microcomputers, 11(10): 10–??, October 1, 1993. CO-DEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:TIC

[Ano93-40] Anonymous. Troubleshooting Internet connections. OCLC Newsletter, 204:31–??, July 1, 1993. ISSN 0163-898X.

Anonymous:1993:UNL

[Ano93-41] Anonymous. The University of Nebraska-Lincoln Libraries gopher. Technical report, the University of Nebraska Libraries and Schmid Law Library in collaboration with the UNL Computing Resource Center, Lincoln, NE, USA, 1993.

Anonymous:1993:VGN

[Ano93-42] Anonymous. Veronica: a Gopher navigational tool on the Internet. Information intelligence, online libraries, and microcomputers, 11(10):1–??, October 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1993:VCM

[Ano93-43] Anonymous. Vinton Cerf:

the man behind the Internet. UNIX/world, 10(12):62– ??, December 1, 1993. ISSN 0739-5922.

Anonymous:1993:WW

[Ano93-44] Anonymous. Windows on the world. Information Week, 449:17–??, November 1, 1993. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1993:WGI

[Ano93-45]  Anonymous.  WinGopher

graphical interface for the Internet, 1993. 1 computer disk.

Anonymous:1993:WWW

[Ano93-46] Anonymous. World Wide Web: Accessing the Internet.

Information intelligence, online libraries, and microcomputers, 11(1):1–??, January 1, 1993. CODEN IIOMEI. ISSN 0737-7770.

Anonymous:1994:AAL

[Ano94a] Anonymous. Association of American Law Schools annual meeting, 1994. sound cassettes.

Anonymous:1994:AIC

[Ano94b] Anonymous. Atlantis Internet CD-ROM, 1994. ISSN 1076-9323. Computer laser optical disks.

Anonymous:1994:A

[Ano94c] Anonymous. Automation. Legal management: the journal of the Association of Legal Administrators, 13(5):12– ??, September 1, 1994. CO-DEN LEMAEB. ISSN 1043-7355.

Anonymous:1994:BSI

[Ano94d] Anonymous. Basics of searching indexes; searching the nursing literature; searching uncover via the Internet, 1994. 1 videocassette (17 min.).

Anonymous:1994:BFK

[Ano94e] Anonymous. Beyond the F1 key thinking and teaching the Internet within the curriculum, 1994. 2 sound cassettes.

Anonymous:1994:BRF

[Ano94f]  Anonymous.  Book re-

view: Firewalls and Internet Security. Linux Journal, 6:??, October 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue6/firewall.
xtg.html.

Anonymous:1994:BRN

[Ano94g] Anonymous. Book reviews:

Newtons Telecom Dictionary. Linux Journal, 3:??, July 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1994:BI

[Ano94h] Anonymous. Business and the Internet, 1994. 2 videocassettes (210 min.).

Anonymous:1994:CUG

[Ano94i]  Anonymous.  Commercial

user’s guide to the Internet. Technical report, Thompson Pub. Group, Washington, DC, USA, 1994. various pp.

Anonymous:1994:CIC

[Ano94j] Anonymous. Connecting to the Internet via Colorado Supernet, 1994. 2 videocassettes (120 and 100 min.).

Anonymous:1994:CYL

[Ano94k]  Anonymous.  Connecting

your Linux box to the Internet. Linux Journal, 8: ??, December 1994. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1994:DMI

[Ano94l] Anonymous. Directory of marketing on the Internet: a primer: tips, tools, rules and resources. Technical report, Thompson Publishing Group, Washington, DC, USA?, 1994. 22 pp.

Anonymous:1994:EMA

[Ano94m] Anonymous. Electronic mail at elmhurst college global communication on the Internet, 1994. 1 videocassette (VHS) (30 min.).

Anonymous:1994:EIV

[Ano94n] Anonymous. Exploring the Internet: volume 1. Technical report, The University, Albany, NY, USA, 1994. 28 pp.

Anonymous:1994:FSI

[Ano94o] Anonymous. Fact sheet on the Internet. Technical Report 94-074, Library of Michigan, Lansing, MI, USA, 1994. 1 pp.

Anonymous:1994:HFI

[Ano94p] Anonymous. Harbor (finding an Internet provider). Linux Journal, 6:??, October 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1994:HUL

[Ano94q] How to use the Library of Michigan ANSWER online public catalog from the Internet, page various, 1994. Library of Michigan, Lansing, MI, USA.

Anonymous:1994:III

[Ano94r] Anonymous. The ICE index (internet connections for engineering), 1994.

Anonymous:1994:IHI

[Ano94s] Anonymous. IFLA headquarters and the Internet. IFLA journal / International Federation of Library Associations, 20(3):369–??, 1994. ISSN 0340-0352.

Anonymous:1994:IS

[Ano94t] Anonymous. The information superhighway. Transportation & distribution, 35 (9):83–??, September 1, 1994. ISSN 0895-8548.

Anonymous:1994:IIA

[Ano94u] Anonymous. Instant Internet access, 1994. ISBN 1-884133-06-1. Includes CD-ROM.

Anonymous:1994:ICH

[Ano94v] Anonymous. Internet 17hands on creating home pages with HTML, 1994. 1 sound cassette.

Anonymous:1994:IIM

[Ano94w] Anonymous. Internet 18hands on introduction to mosaic, 1994. 1 sound cassette.

Anonymous:1994:IHG

[Ano94x] Anonymous. Internet: a howto guide to the ultimate library. Asiaweek, 20(37):39– ??, September 14, 1994. ISSN 1012-6244.

Anonymous:1994:IAH

[Ano94y] Anonymous. Internet access: hearings before the Subcommittee on Science of the Committee on Science, Space, and Technology, U.S. House of Representatives, One Hundred Third Congress, second session, October 4, 1994. United States Government Printing Office, Washington, DC, USA, 1994. ISBN 0-16-046876-0. iii + 201 pp. LCCN KF27.S399 1994f.

Anonymous:1994:IAC

[Ano94z] Anonymous. Internet access to copyright information. Technical report, Copyright Office, Library of Congress, Washington, DC, USA, 1994. 1 pp.

Anonymous:1994:IIG

[Ano94-27]  Anonymous.  The Inter-

net: an introductory guide for United Nations organizations. United Nations. Advisory Committee for the Co-ordination of Information Systems, Geneva, Switzerland, 1994. ISBN 92-1-100678-3. viii + 140 pp. LCCN ????

Anonymous:1994:IMK

[Ano94-28] Anonymous. The Internet and More for Kids. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 0-7821-1517-9. US$16.99.

Anonymous:1994:IP

[Ano94-29] Anonymous. The Internet primer, 1994. 1 videocassette (36 min.) card.

WCUP:1994:IRG

[Ano94-30] Anonymous. An Internet resource guide. West Chester, PA, USA, 1994.

Anonymous:1994:IRB

[Ano94-31] Anonymous. Internet resources for the biologist, 1994.

Anonymous:1994:ISH

[Ano94-32] Anonymous. Internet security: hearing before the Subcommittee on Science of the Committee on Science, Space, and Technology, U.S. House of Representatives, One Hundred Third Congress, secondsession, March 22, 1994. United States Government Printing Office, Washington, DC, USA, 1994. ISBN 0-16-044854-9. iii + 138 pp. LCCN KF27.S399 1994d.

Anonymous:1994:ISJ

[Ano94-33] Anonymous. Internet security: (Jan 86 — Aug 94), citations from the IN-SPEC database. Published search, National Technical Information Service, Washington, DC, USA, 1994. various pp.

Anonymous:1994:ISB

[Ano94-34] Anonymous. The Internet: selected bibliography. Technical Report 94-075, Library of Michigan, Lansing, MI, USA, 1994. 1 pp.

Anonymous:1994:ISD

[Ano94-35] Anonymous. The Internet show: Driver’s education for the information superhighway, 1994. 1 videocasssette (VHS) (60 min.).

Anonymous:1994:ITS

[Ano94-36] Anonymous. Internet Technology Series Volume I: Routing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-106857-1. US$40.50.

Anonymous:1994:ITV

[Ano94-37] Anonymous. Internet TV video guide to global connectivity, 1994. 1 videocassette (120 min.).

Anonymous:1994:IUM

[Ano94-38] Anonymous. Internet underground music archive, 1994.

Anonymous:1994:IU

[Ano94-39] Anonymous. The Internet unleashed, 1994. 1 computer disk.

Anonymous:1994:IVS

[Ano94-40] Anonymous. The Internet video: a step-by-step on-line journey through the vast resources of the information superhighway, 1994. 1 videocassette (91 min.).

Anonymous:1994:IWI

[Ano94-41] Anonymous. Internet World’s on Internet 94: an international guide to electronic journals, newsletters, texts, discussion lists, and other resources on the Internet. Mecklermedia, Westport, CT, USA, 1994. ISBN 0-88736-929-4. xxv + 453 pp. LCCN TK 5105.875 I57 I59 1994 Reference Desk. Contents: Discussion lists and special interest mailing lists — Electronic journals and newsletters — Electronic texts, text archives, selected FTP sites, and Internet resource guides — Freenets and other community-based information services — Campuswide information systems — Commercial services on the Internet — Usenet newsgroups and other mailing lists — WAIS-accessible databases.

Anonymous:1994:IIC

[Ano94-42] Anonymous. InterRamp Internet Chameleon for Windows, 1994. 3 computer disks + instruction card.

Anonymous:1994:II

[Ano94-43] Anonymous. Introduction to the Internet, 1994. 1 videocassette (27 min.).

Anonymous:1994:IId

[Ano94-44] Anonymous. Introduction to the Internet, 1994. 1 videocassette (36 min.).

Anonymous:1994:IIe

[Ano94-45] Anonymous. Introduction to the Internet, 1994. 1 videocassette (110 min.).

Anonymous:1994:J

[Ano94-46]  Anonymous.  Jet-setters,

1994. ISSN 1079-0403.

Anonymous:1994:KII

[Ano94-47] Anonymous. K–12 and the Internet: ideas from educators. Technical report, North-WestNet: Northwest Academic Computing Consortium, Inc. and the University of Washington, Seattle, WA, USA, March 1994. 34 pp.

Anonymous:1994:MIA

[Ano94-48] Anonymous. Manes: Internet is not all it’s cracked up to be. Information Week, 493:86– ??, September 19, 1994. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1994:MI

[Ano94-49]  Anonymous.  Multipoint/

Internet, 1994. 1 videocassette (105 min.).

Anonymous:1994:NCS

[Ano94-50] Anonymous. National Center for Supercomputing Applications. Industry Week, 243 (23):56–??, December 1994.

CODEN IWEEA4. ISSN 0039-0895.

Anonymous:1994:NDS

[Ano94-51] Anonymous. A new dawn for Sun. Information Week, 493: 15–??, September 19, 1994. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1994:PCL

[Ano94-52] Anonymous. Palliative care letter, 1994.

Anonymous:1994:PIS

[Ano94-53] Anonymous, editor. Proceedings — Internet Society: Symposium on Network and Distributed System Security: February 3–4, 1994, Catamaran Hotel, San Diego, California. Internet Society, Reston, VA, USA, 1994. ISBN ???? LCCN ????

Anonymous:1994:PGR

[Ano94-54] Anonymous. The project GAIN report connecting rural public libraries to the Internet, 1994. 1 videocassette (27 min.).

Anonymous:1994:PUE

[Ano94-55]  Anonymous.  Public use

of Earth and space science data over the Internet: a cooperative agreement notice soliciting proposals for (1) innovative remote sensing database applications, (2) digital library technology development, and (3) a remote sensing public access center involving public use Earth and space science data over the Internet. Technical Report CAN-OA-94-1, Office of Aeronautics, National Aeronautics and Space Administration, Washington, DC, USA, March 9, 1994. various pp.

Anonymous:1994:QLG

[Ano94-56] Anonymous. Queens library gopher, 1994.

Anonymous:1994:RMI

[Ano94-57]  Anonymous.  Road map

to the Internet. Technical report, Ziff-Davis Press, Emeryville, CA, USA, 1994. 1 pp.

Anonymous:1994:SFC

[Ano94-58] Anonymous. San Francisco chronicle, 1994.

Anonymous:1994:SS

[Ano94-59]  Anonymous.  Software/

systems. Government computer news, 13(16):56–??, July 1994. ISSN 0738-4300.

Anonymous:1994:TSI

[Ano94-60] Anonymous. Technology: In the shadow of the Internet: What’s in store for CompuServe, Prodigy and America Online? Time, 144(13): 58–??, September 26, 1994. CODEN TYMEA9. ISSN 0040-781X.

Anonymous:1994:TTL

[Ano94-61] Anonymous, editor. Technology, telecommunications and the law: crossing the Internet threshold: law and legal practice for the 21st century. State Law Library Montana, Helena, MT, USA, 1994. ISBN ???? LCCN ????

Anonymous:1994:I

[Ano94-62] Anonymous. This is Internet, 1994. ISBN 1-56641-019-3. 1 videocassette (20 min.) in.

Anonymous:1994:VW

[Ano94-63] Anonymous. Volcano world.

Technical report, Center for Aerospace Sciences, University of North Dakota, Grand Forks, ND, USA, 1994.

Anonymous:1995:ARI

[Ano95a] Anonymous. AIDS resources on the Internet as of 1 may 1995, 1995.

Anonymous:1995:AMM

[Ano95b]  Anonymous.  Almanaque

mundial 1996: la mujer en el mundo: cifras, logros, fracasos, informatica: navegando en la INTERNET. SAMRA Televisa, Mexico DF, Mexico, 1995. ISBN 1-56259-032-4. 608 pp. LCCN ????

Anonymous:1995:AVP

[Ano95c] Anonymous. Anthology Virginia’s PEN K–12 Internet resources, 1995.

Anonymous:1995:BGIa

[Ano95d]  Anonymous.  Beginner’s

guide to the Internet. Technical report, Micro Age Corporation, Pierre, SD, USA, 1995. 30 + 12 pp.

Anonymous:1995:BMB

[Ano95e] Anonymous. The best of mind/body electro-industrial music from the Internet, 1995. 1 sound disc.

Anonymous:1995:BIT

[Ano95f]  Anonymous, editor.  Be-

yond the Internet the transition to the National Information Infrastructure: proceedings of the Library of Congress Network Advisory Committee, December 4–5, 1994, Network planning paper; no. 28. Network Development and MARC Standards Office, Library of Congress, Washington, DC, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:BRY

[Ano95g] Anonymous. Book review: Your Internet Consultant; Tcl and the Tk Toolkit. Linux Journal, 11:??, March 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue11/br11ic.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue11/br11tk.html.

Anonymous:1995:B

[Ano95h] Anonymous. Bulletin. Stereo review, 60(7):8–??, July 1, 1995. ISSN 0039-1220.

Anonymous:1995:CHW

[Ano95i] Anonymous. Capitol Hill watch. National journal, 27 (23):1404–??, June 10, 1995. ISSN 0360-4217.

Anonymous:1995:CUG

[Ano95j] Anonymous. The computer users guide to the Internet. Technical report, Internet Voyager, Clarion, IA, USA, 1995. 29 pp.

Anonymous:1995:CIG

[Ano95k] Anonymous. CyberHound’s Internet guide to the coolest stuff out there. Visible Ink Press, Detroit, MI, USA, 1995. ISBN 0-7876-0688-X. ???? pp. LCCN TK5105.875.I57C93 1995.

Anonymous:1995:CPO

[Ano95l] Anonymous. Cyberporn — protecting our children from the back alleys of the Internet: joint hearing before the Subcommittee on Basic Research and the Subcommittee on Technology of the Committee on Science, U.S. House of Representatives, One Hundred Fourth Congress, first session, July 26, 1995. United States Government Printing Office, Washington, DC, USA, 1995. ISBN 0-16-047717-4. v + 130 pp. LCCN DOC Y 4.SCI 2:104/16. Distributed to some depository libraries in microfiche. Shipping list no.: 96-0043-P. “No. 16.”.

Anonymous:1995:CPG

[Ano95m] Anonymous. CyberSafe a parent’s guide to the Internet, 1995. 1 videocassette (5 p.).

Anonymous:1995:DIa

[Ano95n] Anonymous. Dance on the Internet. Afterimages, 2(2–3): 1, 4–9, Spring–Summer 1995. ISSN 1073-2101.

Anonymous:1995:DY

[Ano95o] Anonymous. Diario de Yucatan, 1995.

Anonymous:1995:DIb

[Ano95p] Anonymous. Discovering the Internet, 1995. 1 videocassette (60 min.).

Anonymous:1995:ERI

[Ano95q] Anonymous. Education resources on the Internet: an introductory guide to select Internet education and related resources. Technical report, University of the State of New York, State Education Dept., Albany, NY (Education Bldg., Rm. 120, 12234), August 1995. vi + 43 pp.

Anonymous:1995:Nb

[Ano95r] Anonymous. El norte, 1995.

Anonymous:1995:EIG

[Ano95s] Anonymous. The essential Internet guide for law students. Technical report, State University of New York at Buffalo. School of Law and Law and Technology Issues Society, Buffalo, NY, USA, 1995. 44 + 2 pp.

Anonymous:1995:FII

[Ano95t] Anonymous. Finding it on the Internet, 1995. 1 videocassette (ca. 30 min.).

Anonymous:1995:FYS

[Ano95u] Anonymous. FindLaw your source to law on the Internet, 1995.

Anonymous:1995:FLA

[Ano95v] Anonymous. A first look at the Internet. Technical report, CCI Computer Courseware International Inc., ??, BC, Canada, 1995. 94 pp.

Anonymous:1995:GIA

[Ano95w] Anonymous. Gopher it!: accessing Department of Education grant information on the Internet. Consumer guide; no. 11 consumer guide (United States Dept. of Education); no. 11, United States. Dept. of Education, Washington, DC, USA, 1995. 2 pp.

Anonymous:1995:GLM

[Ano95x] Anonymous. Guide to the Library of Michigan Internet gopher. Technical report, Library of Michigan, Lansing, MI, USA, 1995. 2 pp.

Anonymous:1995:HT

[Ano95y] Anonymous. The Hastings Tribune, 1995.

Anonymous:1995:I

[Ano95z] Anonymous. Internet, 1995. ISSN 1081-2482.

Anonymous:1995:IGI

[Ano95-27] Anonymous. The Internet: a guide to Internet use in the federal government. Treasury Board of Canada Secretariat, Financial and Information Management Branch, Information, Communications and Security Policy Branch, Ottawa, ON, Canada, 1995. ISBN 0-662-23277-1. 50 pp. LCCN ????

Anonymous:1995:IKO

[Ano95-28] Anonymous. Internet: a knowledge odyssey, 1995. ISBN 1-57559-002-6. Includes CD-ROM.

Anonymous:1995:IAS

[Ano95-29] Anonymous. Internet access for small libraries on ramps for the rest of us, 1995. 2 sound cassettes.

Anonymous:1995:IAC

[Ano95-30] Anonymous. Internet access to copyright information. Technical report, Copyright Office, Library of Congress, Washington, DC, USA, 1995. 1 pp.

Anonymous:1995:IAI

[Ano95-31] Anonymous. Internet access to the information highway, 1995. 1 computer disk.

Anonymous:1995:IPL

[Ano95-32] Anonymous. The Internet and the practice of law. Technical report, Minnesota Continuing Legal Education, St. Paul, MN, USA, September 1995. various pp.

Anonymous:1995:IAY

[Ano95-33] Anonymous. The Internet at your Alameda County Library, 1995. 1 videorecording (24 min.).

Anonymous:1995:ICH

[Ano95-34] Anonymous. The Internet cyber hate and freedom of speech, 1995. 1 videocassette (20 min.).

Anonymous:1995:IFF

[Ano95-35] Anonymous. Internet file formats, 1995. Includes CD-ROM.

Anonymous:1995:ILa

[Ano95-36] Anonymous. The Internet for lawyers. Pbi; no. pbi (series); no. 1995-1034, Pennsylvania Bar Institute, Harrisburg, PA, USA, 1995. various pp.

Anonymous:1995:IIb

[Ano95-37] Anonymous. Internet info, 1995. ISSN 1085-8776. Includes CD-ROM.

Anonymous:1995:ILI

[Ano95-38] Anonymous. Internet legal information reference desk,

Anonymous:1995:IOU

[Ano95-39] Anonymous. Internet office: user guide. SPRY, Inc., Seattle, WA, USA, 1995. ISBN ???? various pp. LCCN ????

Anonymous:1995:IPW

[Ano95-40] Anonymous. Internet pro Web page builder, 1995. ISBN 1-56737-149-3. Includes CD-ROM.

Anonymous:1995:ISEa

[Ano95-41] Anonymous. Internet sites ed policy and research, 1995.

Anonymous:1995:IS

[Ano95-42] Anonymous. Internet surfer, 1995. Includes CD-ROM.

Anonymous:1995:IUa

[Ano95-43] Anonymous. The Internet unleashed. Sams.net Pub., Indianapolis, IN, USA, second edition, 1995. ISBN 0-672-30714-6 (paperback). lii + 1384 pp. LCCN TK5105.875.I57I586 1995.

Anonymous:1995:IUb

[Ano95-44]  Anonymous.  The Inter-

net unleashed. Sams.net Pub., Indianapolis, IN, USA, third edition, 1995. ISBN 0-672-30714-6 (paperback). li + 1398 pp. LCCN TK5105.875.I57I586 1995.

Anonymous:1995:IWI

[Ano95-45] Anonymous. The Internet: what it can and can’t do for activists. Technical report, Social Justice Connections, Arlington, VA, USA, 1995. 11 pp.

Anonymous:1995:IYF

[Ano95-46] Anonymous. The Internet your first cruise, 1995. 1 videocassette (25 min.).

Anonymous:1995:IIa

[Ano95-47] Anonymous. Introduction to the Internet. Technical report, South Dakota. Information Services, Pierre, SD, USA, 1995. various pp.

Anonymous:1995:LGI

[Ano95-48] Anonymous. A lawyer’s guide to the Internet: how and why lawyers are going on-line. Technical report, State Bar of Wisconsin CLE, Madison, WI, USA, July 1995. various pp.

Anonymous:1995:LOG

[Ano95-49] Anonymous. Lawyers on-line:

a guide to using the Internet. Technical report, Virginia Law Foundation. Committee on Continuing Legal Education, Charlottesville, VA, USA, 1995. various pp.

Anonymous:1995:LUI

[Ano95-50] Anonymous. Learn to use the Internet, 1995. ISBN 1-57142-006-1. 1 videocassette (VHS) (45 min.).

Anonymous:1995:LAC

[Ano95-51] Anonymous. List of ATM chips now on Internet. Electronic Design, 43(12):34–??, June 12, 1995. CODEN ELO-DAW. ISSN 0013-4872.

Anonymous:1995:MIC

[Ano95-52] Anonymous. Making the Internet connection: electronic publishing and research, training the trainers. Technical report, Case Western University, Cleveland, OH, USA, July 12–14, 1995. various pp.

Anonymous:1995:NI

[Ano95-53] Anonymous. Navigating the Internet, 1995. 2 videocassettes (95 min.).

Anonymous:1995:NE

[Ano95-54] Anonymous. Networked educator, 1995.

Anonymous:1995:NPI

[Ano95-55]  Anonymous.  New prod-

ucts: Interactive Unix Guide; SEDIT and S/REXX for Linux; Linux Internet Archives. Linux Journal, 16:??, August 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue16/np16a.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue16/np16b.html;   http:
//noframes.linuxjournal.
com/lj-issues/issue16/np16c.
html.

Anonymous:1995:Na

[Ano95-56] Anonymous. News. Computers in Physics, 9(3):251– ??, May 1, 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1995:IUD

[Ano95-57] Anonymous. On the Internet: user demographics and trends. SIMBA Information Inc., Wilton, CT, USA, 1995. ISBN 0-88709-089-3. ix + 123 pp. LCCN HE7581.O5 1995.

Anonymous:1995:PGI

[Ano95-58] Anonymous. PC guide for Internet and on-line services, 1995. ISBN 1-881979-20-2 (book/video), 1-881979-22-9 (video), 1-881979-21-0 (book). 1 videocassette (1 hr.).

Anonymous:1995:RDI

[Ano95-59] Anonymous. RBOCs dedicated to Internet access. America’s network: technology for the information highway, 99(11):46–??, June 1, 1995. CODEN ANETE4. ISSN 1075-5292.

Anonymous:1995:R

[Ano95-60] Anonymous. Reforma, 1995.

Anonymous:1995:RSL

[Ano95-61] Anonymous. Report of the State Library Board and the Department of Information Technology on assessing the need to establish protocols and guidelines regarding instate access to the myriad files and components available through the Internet to the Governor and the General Assembly of Virginia. Technical Report 15, Virginia. Division of Legislative Automated Systems and Virginia. Dept. of Information Technology, Commonwealth of Virginia, Richmond, VA, USA, 1995. various pp.

Anonymous:1995:SIR

[Ano95-62] Anonymous. Sailor Internet: reference manual. Technical report, Maryland. Division of Library Development and Services, Baltimore, MD, USA, 1995. 50 pp.

Anonymous:1995:SWB

[Ano95-63] Anonymous. Spinnin’ the Web and beyond expanding your Internet skills, 1995. 1 videocassette (20 min.).

Anonymous:1995:SIW

[Ano95-64] Anonymous, editor. Spring Internet World 95 conference and exhibition, Virtual conference. Mecklermedia, Westport, CT, USA, 1995. ISBN ???? LCCN ???? Includes CD-ROM.

Anonymous:1995:TAB

[Ano95-65] Anonymous. A technology agenda for 1996 and beyond: harnessing technology to boost your practice: pinpointing dangers and opportunities lurking on the Internet. Technical report, American Lawyer Media, New York, NY, USA, 1995. 50 pp.

Anonymous:1995:TID

[Ano95-66] Anonymous. Telcom the Internet distance education, 1995. 1 videocassette.

Anonymous:1995:T

[Ano95-67] Anonymous. TimesFax, 1995.

Anonymous:1995:TKT

[Ano95-68] Anonymous. Training kit:

training materials for networked systems: CD-ROM, internet, E-mail, etc. Technical report, West Publishing, St. Paul, MN, USA, 1995. 157 pp.

Anonymous:1995:UNL

[Ano95-69] Anonymous. University of Nebraska-Lincoln libraries Internet catalog. Technical report, University of Nebraska-Lincoln Libraries, Lincoln, NE, USA, 1995.

Anonymous:1995:WSI

[Ano95-70] Anonymous. Who’s succeeding on the Internet and how? Technical report, ActivMedia, Peterborough, NH, USA, 1995. viii + 80 pp.

Anonymous:1995:YW

[Ano95-71]  Anonymous.  Young and

wired. American photo, 6(4): 89–??, July 1, 1995. ISSN 1046-8986.

Anonymous:1995:YRI

[Ano95-72] Anonymous. Your roadmap to the Internet: Via the Bureau of Transportation statistics, 1995. 1 videocassette (18 min.) sd. + col. kit (73 p., 28 cm.).

Anonymous:1996:INB

[Ano96a] Anonymous. 10x: Internet notebook for biologists, 1996. ISSN 1088-2766.

Anonymous:1996:BRB

[Ano96b]  Anonymous.  Book re-

view: Building a Linux Internet Server. Linux Journal, 23:??, March 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue23/blisrev.
html.

Anonymous:1996:BRIa

[Ano96c] Anonymous. Book review: IPv6: The New Internet Protocol. Linux Journal, 25:??, May 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue25/1249.
html.

Anonymous:1996:CSG

[Ano96d] Anonymous. The common sense guide to organizational communications on the Internet. Technical report, Lawrence Ragan Communications, Inc., Chicago, IL, USA, 1996. 24 pp.

Anonymous:1996:EEG

[Ano96e] Anonymous. eVO: electronic guide to the Internet culture, 1996. ISSN 1088-2758.

Anonymous:1996:FSO

[Ano96f] Anonymous. Financial services organisations on the Internet, 1996. Market Assessment Publications, London, UK, 1996. ISBN 1-86111-033-2 (spiral). 89 pp. LCCN ????

Anonymous:1996:HCI

[Ano96g] Anonymous. How to connect to the Internet. Iconnect publication series. ALA Editions, Chicago, IL, USA, 1996. ISBN 0-8389-3457-9 (paperback). 16 pp. LCCN ????

Anonymous:1996:IUN

[Ano96h] Anonymous. Indiana, USA newspaper services on the Internet, 1996.

Anonymous:1996:IOI

[Ano96i] Anonymous. The Internet and online industry sourcebook. Gateway Publishing, North Adams, MA, USA, 1996. ISBN ???? 474 pp. LCCN ????

Anonymous:1996:IES

[Ano96j] Anonymous. Internet for educators a step-by-step guide to help educators understand and use the Internet, 1996.

ISBN 0-9648717-1-8. 1 videocassette (66 min.).

Anonymous:1996:IRSb

[Ano96k] Anonymous. Internet resources for the study of Judaism and Christianity. Technical report, University of Pennsylvania, Philadelphia, PA, USA, 1996.

Anonymous:1996:ISB

[Ano96l] Anonymous. The Internet: selected bibliography. Technical Report 96-032, Library of Michigan, Lansing, MI, USA, 1996. 1 pp.

Anonymous:1996:IU

[Ano96m] Anonymous. The Internet unleashed 1996. Sams.net Pub., Indianapolis, IN, USA, third edition, 1996. ISBN 1-57521-041-X. li + 1398 pp. LCCN TK5105.875.I57I586 1995b.

Anonymous:1996:LCL

[Ano96n] Anonymous. Law and computers: the lawless Internet?, 1996. Association of American Law Schools. Meeting (San Antonio, TX, USA). 2 sound cassettes.

Anonymous:1996:LCB

[Ano96o] Anonymous. Lawoffice.com: bringing your law firm onto the Internet. Technical report, Minnesota Continuing Legal Education, St. Paul, MN, USA, January 1996. various pp.

Anonymous:1996:LRIb

[Ano96p]  Anonymous, editor.  Le-

gal research on the Internet. Kansas Bar Association, Topeka, KS, USA, 1996. ISBN ???? LCCN ????

Anonymous:1996:LRIa

[Ano96q] Anonymous. Legal research on the Internet: a simple, “how to” guide for courts and attorneys. Technical report, Washington (State). Office of the Administrator of the Courts, Olympia, WA, USA, February 15, 1996.

Anonymous:1996:MBI

[Ano96r] Anonymous. Microsoft bookshelf Internet directory. Microsoft Press, Bellevue, WA, USA, 1996–1997 edition, 1996. ISBN 1-55615-947-1. xiii + 699 pp. LCCN QA76.57.M52M52 1996.

Anonymous:1996:NPC

[Ano96s] Anonymous. New products:

Caldera Internet Office Suite; Cosmos Shipping Linux on a Hard Disk; TEAMate Web/ BBS Server for Linux. Linux Journal, 28:??, August 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1996:NPL

[Ano96t] Anonymous. New products:

Linux Internet Archives; Visual SlickEdit for Linux; Emulus — Terminal Emulation Software; Orion Firewall System. Linux Journal, 25:??, May 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue25/0111.
html.

Anonymous:1996:NPO

[Ano96u] Anonymous. New products: ObjectSpace Java Generic Library; ObjectSpace Web Toolkit; Innovative Software InvisibleWeb & Offline Proxy Server; Datacomm Internet and Intranet/Web Server with Cyrix 166MHz chip; ARDI Executor 2; TowerEiffel Release 2.0; Debian Linux 1.1; Open Systems Management COS/ Print; Amtec Engineering Tecplot 7.0; Thought, Inc. VanillaSearch; Dimension X Liquid Reality Developer’s Kit; X Inside, Inc. Accelerated OpenGL Solution for Linux; Spire Technologies Tactician Plus. Linux Journal, 30:??, October 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue30/newprod.
html.

Anonymous:1996:NPV

[Ano96v]  Anonymous.  New prod-

ucts: Visual SlickEdit Version 2.0; C++ Version of Andrew User Interface; JClass LiveTable Pro & LiveTable Applet; FairCom Web Server; New Release of ObjectSpace C++ Class Libraries; CocoBase Java DataBase Access System; InTerNet-LINK; Linux Advanced GUI Toolkit; Introduction to Programming Java Applets; SOLID Server 2.1 for Linux; COS/Relay; Linux Serial I/O Solutions; CRiSP v5-Multiplatform Professional Editor; Cross-Platform Web Agent-Trudger; NExS version 1.2 for Linux; Networking Switch for Access 8 PCs; Acu4GL for ODBC; InfoMagic Workgroup Server. Linux Journal, 32:??, December 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1996:RDL

[Ano96w] Anonymous. Report of the Division of Legislative Automated Systems and the Department of Information Technology on development of a prototype to provide legislative information via the Internet to the Governor and the General Assembly of Virginia. Technical Report 25, Virginia. Division of Legislative Automated Systems and Virginia. Dept. of Information Technology, Commonwealth of Virginia, Richmond, VA, USA, 1996. 6 pp.

Anonymous:1996:TIW

[Ano96x] Anonymous, editor. Textile Institute: World conferences; 77th — May 1996, Tampere, Finland. Institute of Fiber, Textile and Clothing Science, Tampere University of Technology, Tampere, Finland, 1996. ISBN 951-722-545-8. LCCN ????

Anonymous:1997:AAA

[Ano97a] Anonymous, editor. 1997 ACSM/ASPRS Annual Convention & Exposition: technical papers: ACSM 57th Annual Convention, ASPRS 63rd Annual Convention, Seattle, Washington, April 7– 10, 1997, volume 2 of ACSM ASPRS Annual Convention and Exposition. American Congress on Surveying and Mapping and American Society for Photogrammetry and Remote Sensing, Bethesda, MD, USA, 1997. ISBN 1-57083-043-6 (set), 1-57083-044-4 (vol. 1), 1-57083-045-2 (vol. 2), 1-57083-047-9 (vol. 4), 1-57083-048-7 (vol. 5). LCCN ???? Five volumes.

Anonymous:1997:BBC

[Ano97b]  Anonymous.  Black belt

client/server software development: practical techniques from master programmers. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 1997. ISBN 0-87930-498-7. xii + 323 pp. LCCN ???? US$34.95. Includes floppy disk.

Anonymous:1997:BRI

[Ano97c]  Anonymous.  Book re-

view: Internet Programming with Python. Linux Journal, 42:??, October 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/2152.html.

Anonymous:1997:BRPa

[Ano97d] Anonymous. Book review:

Practical Unix and Internet Security. Linux Journal, 34:??, February 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/0182.html.

Anonymous:1997:PIC

[Ano97e] Anonymous. From the publisher: Internet changes/ Linux changes. Linux Journal, 42:??, October 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:PLI

[Ano97f] Anonymous. From the publisher: Linux — the Internet appliance? Linux Journal, 36:??, April 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:IGS

[Ano97g] Anonymous. Internet goldmine: San Diego Supercomputer Center. Computers in Physics, 11(3):229–??, ????

1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:NPI

[Ano97h] Anonymous. New products: Integra4 RDBMS for Linux; Clustor 1.0; BitWizard; Hitachi MP-EG1A; LinkScan; Pixel!FX Version 5.1; JClass Chart; Internet/Intranet Design Shop; Velocis Database Server. Linux Journal, 38:??, June 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:NIP

[Ano97i] Anonymous. NEWS — Internet 2, Physics Computing ’97, Supercomputing ’96. Computers in Physics, 11(1): 8–??, ???? 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:NNM

[Ano97j]  Anonymous.  Novice to

novice: a 10-minute guide for using PPP to connect Linux to the Internet. Linux Journal, 36:??, April 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //noframes.linuxjournal.
                                        com/lj-issues/issue36/ppp.
                                        html.

Anonymous:1998:AUS

[Ano98a] Anonymous. Announcement: 2nd USENIX Symposium on Internet Technologies and Systems (USITS ’99). ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:ASC

[Ano98b]  Anonymous.  Announce-

ments: SIGGRAPH community invited to join Internet technical group. Computer Graphics, 32(2):74, May 1998. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Anonymous:1998:CRUa

[Ano98c] Anonymous. Conference report: USENIX Symposium on Internet Technologies and Systems. ;login: the USENIX Association newsletter, 23 (2):??, April 1998. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/usits97/
summaries.html.

Anonymous:1999:AUSa

[Ano99a] Anonymous. Announcement:

2nd USENIX Symposium on Internet Technologies and Systems. ;login: the USENIX Association newsletter, 24 (3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/usits99/

.

Anonymous:1999:GIT

[Ano99b] Anonymous, editor. Geospatial Information and Technology Association: Conference; 22nd — April 1999, Charlotte, NC, GITA - PROCEEDINGS- 1999; 22nd. Geospatial Information and Technology Association, ????, 1999. ISBN ???? LCCN ????

Anonymous:1999:IAI

[Ano99c] Anonymous. Inductive analysis of the Internet Protocol TLS: Discussion. Lecture Notes in Computer Science, 1550:13–23, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1999:NPCa

[Ano99d]  Anonymous.  New prod-

ucts: Cyclades-PR4000, Cyclades Corp.; PerlDirect, ActiveState Tool Corp.; ICS, BASCOM Global Internet Services, Inc.; Linux Network Server Package, CTiTEK; CSM Proxy Plus for Linux Version 4.1, CSM-USA, Inc.; Empress REBMS v8.10, Empress Software; M-Cluster, Alta Technology Corp.; LynxArray and LynxNSS, Artecon; GO-Global, GO-Between, G0-Joe, GraphOn Corp.; Linux Main Memory Database Benchmark, Polyhedra, Inc.; Magnate Internet Store, ParaSoft Corp.; LinuxCare, LinuxCare, Inc. Linux Journal, 60:??, April 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1999:NPI

[Ano99e] Anonymous. New products:

iHTML Merchant 2.0, Inline Internet Systems, Inc.; Auto-Changer, UniTrends Software Corporation; JetStor RAID, AC&NC; VariCAD 7.0-1.0, VariCAD; MetaCard 2.2, MetaCard Corp.; Macsyma 421 for Linux, Macsyma Inc.; HP Firehunter, Hewlett–Packard Company; PlanetUplink, Planet Computer; NetTracker 3.5 and Webmaster 3.0, Sane Solutions LLC; ICE.PPN, J. River, Inc.; O2 Object Database Management System, Ardent Software, Inc.; Quant-X and Altera SQL Server version 2.0, Altera Ltd.; Linux Edition of Ingres II, Computer Associates International, Inc.; Thin Client PC X Server, GraphOn Corporation. Linux Journal, 58:??, February 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:19xx:IBH

[Anoxxa] Anonymous. Internet/Bitnet health sciences resources, 19xx.

Anonymous:19xx:NZL

[Anoxxb] Anonymous. New Zealand law on the Internet. Technical report, University of Waikato, Law Library School of Law, Hamilton, New Zealand, 19xx.

Anonymous:19xx:YRI

[Anoxxc] Anonymous. Your roadmap to the Internet via the bureau of transportation statistics, 19xx. 1 videocassette (18 min.).

ANSI:1992:IPS

[ANS92] ANSI. Information processing systems: local area networks — Part 3. Carrier sense multiple access with collision detection (CSMA/CD) access method and physical layer specifications. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1992. ISBN 1-55937-049-1. 220 pp. LCCN TK5105.7.I54 1990. ANSI/ IEEE Std 802.3-1990 Edition. International standard ISO/ IEC 8802-3, IEEE product number: SH13482.

Ahmad:1996:COI

[ANT96] Nyla Ahmad, Martha Newbigging, and Keltie Thomas. CyberSurfer: the OWL Internet guide for kids. Owl Books: Greey de Pencier (distributed in the U.S. by Firefly Books), Toronto, ON, Canada, 1996. ISBN 1-895688-50-7. 72 pp. LCCN ISBN 1-56604-358-1.

Adamson:1995:OSE

[AP95a] Mark Adamson and Fraser Pearce. Online services in Europe: the impact of the Internet on business information services. Financial Times management reports. FT Telecoms and Media Publishing, London, UK, 1995. ISBN 1-85334-398-6 (spiral). viii + 137 pp. LCCN ISBN 1-56884-658-4.

Aughney:1995:IGI

[AP95b] Jim Aughney and Jimmy Plenderleith. The Ireland guide to the Internet. Urban Dynamix, ??, Ireland, 1995. ISBN 0-9526847-0-5 (paperback). 215 pp. LCCN ????

Alexander:1996:ISF

[AP96] David Alexander and Jerome Preisler. The Internet site finder. Avon Books, New York, NY, USA, 1996. ISBN 0-380-78269-3 (paperback). 227 pp. LCCN ????

AppleComputer:1995:AIC

[App95] Apple Computer, Inc. Apple Internet connection kit, 1995. 10 computer disks.

Abdullahi:1998:GCI

[AR98] Saleh E. Abdullahi and Graem A. Ringwood. Garbage collecting the Internet: a survey of distributed garbage collection. ACM Computing Surveys, 30(3):330–373, September 1998. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/surveys/1998-30-
3/p330-abdullahi/.

Archer:1995:NGG

[Arc95] George W. (George Warren) Archer. Netguide: genealogist’s guide to Internet. Technical report, G.W. Archer, McLean, VA (P.O. Box 6233, McLean 22106), 1995. various pp.

Arick:1993:TIC

[Ari93] Martin R. Arick. TCP/IP Companion—A Guide for the Common User. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1993. ISBN 0-89435-466-3. xxv + 240 pp. LCCN TK5105.55 .A75 1993. US$32.95.

Armbruster:1994:IE

[Arm94a] Lynda Armbruster. Internet essentials. Que College, Indianapolis, IN, USA, 1994. ISBN 1-56529-771-7 (paperback). x + 95 pp. LCCN TK5105.875.I57A76 1994.

Armbruster:1994:IEJ

[Arm94b] Lynda Armbruster. Internet Essentials: a Jump-Start to Getting on the Internet. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 1-56529-771-7. x + 95 pp. LCCN ???? US$25.50.

Armstrong:1996:PTK

[Arm96] Sara Armstrong. A pocket tour of kidstuff on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1803-

8. xii + 185 pp. LCCN TK5105.875.I57A78 1996.

Arnett:1994:ITI

[Arn94] Matthew Flint Arnett. Inside TCP/IP. New Riders Publishing, Carmel, IN, USA, 1994. ISBN 1-56205-354-X. xx + 535 pp. US$39.99.

Arnett:1995:IAN

[Arn95] Nick Arnett. The Internet and the anti-net. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

????  URL http://www.
ncsa.uiuc.edu/SDG/IT94/
Proceedings/Overviews/arnett/
Antinet.html. Includes CD-

ROM.

Arnold:1996:IGS

[Arn96] David Arnold. Internet getting started, 1996. 1 videocassette (40 min.).

Asar:1991:PEM

[Asa91] Ravin Asar. Privacy enhanced mail: an implementation for the Internet. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1991. vii + 83 pp.

Abali:1997:CSM

[ASB97] Bulent Abali, Craig B. Stunkel, and Caroline Benveniste. Clock synchronization on a multicomputer. Journal of Parallel and Distributed Computing, 40(1):118–130, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1268/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1268/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1268/production/ref.

Ashton:1995:IAA

[Ash95]  Gary L. Ashton.  Inter-

net activities: adventures on the superhighway. South-Western Educational Pub., Cincinnati, OH, USA, 1995. ISBN 0-538-65147-4. ???? pp. LCCN HE7570.I57 1995.

AAS:1994:SLT

[Ass94a] Associates for Asian Studies. Subcommittee on Library Technology and Workshop on Handling Vernacular on Personal Computers and over the Internet. The Subcommittee on Library Technology, The Committee on East Asian Libraries (CEAL), The Association for East Asian Studies announces the distribution of Handouts of Workshop on Handling Vernacular on Personal Computers and over the Internet. CEAL Subcommittee on Library Technology, ????, 1994. various pp.

ARCL:1994:BFK

[Ass94b] Association of College and Research Libraries. Bibliographic and Instruction Section. Beyond the F1 key thinking and teaching the Internet within the curriculum, 1994. 2 sound cassettes.

AALS:1995:AAM

[Ass95a] Association of American Law Schools. 1995 AALS annual meeting, 1995. sound cassettes.

ACRL:1995:ANC

[Ass95b] Association of College and Research Libraries. ACRL 7th national conference continuity and transformation: the promise of confluence (March 29–April 1, 1995, Pittsburgh, Pennsylvania), 1995. sound cassettes.

OReilly:1999:NCB

[Ass99] O’Reilly and Inc. Associates.

The Networking CD Bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-523-8. 456 pp. LCCN ???? US$79.95. URL http://

www.oreilly.com/catalog/

netcd/. Six O’Reilly networking books on CD-ROM.

Issued in container and accompanied by a print ed. of DNS and BIND (3rd ed.). Contents: DNS and BIND (3rd ed.) – TCP/IP network administration (2nd ed.) – Building Internet firewalls – Practical UNIX and Internet security (2nd ed.) – Sendmail desktop reference – Sendmail (2nd ed.).

Austin:1994:GSB

[AT94] Terese Austin and Kim Tsang. Government sources of business and economic information on the Internet. Technical report, University of Michigan Library, Ann Arbor, MI, USA, 1994. 50 pp.

Atkinson:1999:CPP

[Atk99] Leon Atkinson. Core PHP Programming: using PHP to build dynamic Web sites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-020787-X. 450 pp. LCCN QA76.73.P224A85

1999. URL http://www.
phptr.com/ptrbooks/ptr_
013020787X.html.

ATT:1993:SPC

[ATT93] AT&T, Tel: (800) 344-0223 ext 1102 (USA), 32-2-676-3623 (Europe) (416) 756-5118 (Canada). AT&T Systimax PDS Components Catalog, 1993.

ACRL:1994:SNI

[AUA94] Association of College and Research Libraries, University of Texas at Austin. General Libraries. Reference and Information Services Department, and American Library Association. Search the net Internet camera-ready handouts, 1994. 15 posters.

ARCITB:1993:GQI

[AUC93] Ames Research Center. Imaging Technology Branch, United States. National Aeronautics and Space Administration, and Central Operation of Resources for Educators. Global quest the Internet in the classroom, 1993. 1 videocassette (11 min. + 45 sec.) Internet background information booklet (40 p.; 28 cm.).

Augspurger:1996:IWW

[Aug96] Lynn L. Augspurger. Internet — world wide web. Technical report, IBM Corporation, San Jose, CA, USA, 1996. various pp. Six volumes.

Avolio:1995:STP

[AV95] Frederick M. Avolio and Paul A. Vixie. Sendmail: Theory and Practice. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1995. ISBN 1-55558-127-7. xv + 262 pp. LCCN HE6239.E54 A96 1995. US$29.95.

Anonymous:1996:AID

[Avi96] Aviation Internet directory, page various, 1996. ISSN 1088-0526. SRAS Pub., Lehighton, PA, USA.

Arlitt:1997:IWS

[AW97] Martin F. Arlitt and Carey L. Williamson. Internet Web servers: workload characterization and performance implications. IEEE/ACM Transactions on Networking, 5(5):631–645, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1997-5-5/p631-arlitt/.

Ayre:1994:CSM

[Ayr94] Rick Ayre. Cover story: Making the Internet connection. PC Magazine, 13(17):118–??, October 11, 1994. ISSN 0888-8507.

Baczewski:1994:TIG

[B^{+}94] Philip Baczewski et al. Tricks of the Internet Gurus. Howard W. Sams, Indianapolis, IN 46268, USA, December 1994. ISBN 0-672-30599-2. xxxiii + 809 pp. LCCN TK5105.875.I57 T75 1994. US$35.00, CDN$47.95. A collection of chapters written by 21 well known Internet user/authors, each of whom is an Internet “expert” in his or her own field. This book is not for the newbie or novice Internaut, but is filled with Internet tricks and secrets that will help turn a moderately experienced person into a seasoned veteran. These are not mere tutorials, but detailed high-level essays written by visible, online personalities. Other chapters cover how to send a fax from the Internet, discussion forums, educational uses of the Internet, research, business and commerce. Well-written and presented.

Bennet:1995:DTI

[B^{+}95] Geoff Bennet et al. Designing TCP/IP Internetworks. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01880-0. xii + 601 pp. LCCN TK5105.585.B46 1995. US$49.95. Includes two diskettes.

Bernstein:1996:ISB

[B^{+}96a] Terry Bernstein et al. Internet security for business. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-13752-9 (paperback). various pp. LCCN HD30.38.I57 1996.

Blau:1996:VPA

[B^{+}96b] Brian Blau et al., editors.

Visual proceedings: the art and interdisciplinary programs of SIGGRAPH 96: SIGGRAPH 96, August 4– 9, 1996, New Orleans, LA, Computer Graphics. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-784-7. ISSN 1069-5419. LCCN T385 .S54

                                        1996b. URL http://info.
                                        acm.org/pubs/contents/proceedings/
                                        graph/.

Berger:1995:CWM

[BA95] David A. Berger and Charles R. Azer. The cheapest way to make phone calls, send faxes, or use the Internet. Technical report, Value Added Services Marketing Group, Ann Arbor, MI (P.O. Box 3271, Ann Arbor 48106), 1995. xiv + 121 pp.

Tillman:1994:MMG

[Bab94] Hope N. Babson. Making the most of gophers. Technical report, Special Libraries Association, Washington, DC, USA, 1994. 7 pp.

Baczewski:1994:IU

[Bac94] Philip Baczewski. The Internet unleashed. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30466-X (paperback). lviii + 1387 pp. LCCN TK5105.875.I57I586 1994.

Bacard:1995:CPH

[Bac95a] André Bacard. Computer Privacy Handbook: a Practical Guide to E-Mail Encryption. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1995. ISBN 1-56609-171-3. 274 pp. LCCN ???? US$24.95.

Bachrach:1995:ECI

[Bac95b] Steven M. Bachrach. Electronic conferencing on the Internet: The First Electronic Computational Chemistry Conference. Journal of Chemical Information and Computer Sciences, 35(3): 431–??, May 1, 1995. CO-DEN JCISD8. ISSN 0095-2338.

Bachrach:1996:IGC

[Bac96] Steven M. Bachrach. The Internet: a guide for chemists. Computer applications in chemistry books. American Chemical Society, Washington, DC, USA, 1996. ISBN 0-8412-3223-7 (clothbound), 0-8412-3224-5 (paperback). xv + 344 pp. LCCN QD8.3 .I57 1996.

Badgett:1995:WI

[Bad95] Tom Badgett. Welcome to — the Internet. MIS-Press, New York, NY, USA, second edition, 1995. ISBN 1-55828-424-9. xiii + 461 pp. LCCN TK5105.875.I57B33 1995.

Bonetti:1994:DRM

[BAG94] Pierluigi Bonetti, Claudio Allocchio, and Antonia Ghiselli. Distribution of RFC 1327 mapping rules via the Internet DNS: the INFNet distributed gateway system. Computer Networks and ISDN Systems, 27(3): 461–469, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/

                                        comnet/cas_sub/browse/browse.
                                        cgi?year=1994&volume=27&
                                        issue=3&aid=1366.

Bagwill:1995:SII

[Bag95] Robert H. Bagwill. Sharing information via the Internet: an infoserver case study. NISTIR; 5757 NIS-TIR 5757, U.S. Dept. of Commerce, Technology Administration, National Institute of Standards and Technology, Computer Systems Laboratory; National Technical Information Service, distributor, Gaithersburg, MD, USA, 1995. 23 pp.

Baird:1994:JAS

[Bai94] Warren Baird. The joy (and agony) of SLIP. Linux Journal, 6:??, October 1994. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Baker:1993:NWF

[Bak93a] Steven Baker. Net worth: Fast forward to the Internet. UNIX review, 11(11):15–??, November 1, 1993. CODEN UNRED5. ISSN 0742-3136.

Baker:1993:NWJ

[Bak93b] Steven Baker. Net worth: Joining the Internet club. UNIX review, 11(10):23–??, October 1, 1993. CODEN UNRED5. ISSN 0742-3136.

Baker:1993:NWS

[Bak93c] Steven Baker. Net worth: Satellites and faxes on the Internet. UNIX review, 11(12): 23–??, December 1, 1993. CODEN UNRED5. ISSN 0742-3136.

Baker:1994:ISR

[Bak94] Susan Ratzlaff Baker. Internet: a significant research tool for private acadamia. Thesis (M.S.), Oklahoma State University, Stillwater, OK, USA, 1994. v + 127 pp.

Baker:1995:CCC

[Bak95] Jason D. Baker. Christian cyberspace companion: a guide to the Internet and Christian online resources. Baker Books, Grand Rapids, MI, USA, 1995. ISBN 0-8010-5248-3 (paper). 209 pp. LCCN BR99.B35 1995.

Bakos:1998:ERE

[Bak98] Yannis Bakos. The emerging role of electronic marketplaces on the Internet. Communications of the ACM, 41 (8):35–42, August 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-8/
p35-bakos/.

Balkits:1991:UI

[Bal91] Ivars Balkits. Using the Internet. Technical report, ????, ????, April 3, 1991. 33 pp.

Bal:1993:FI

[Bal93a] S. Bal. A for-profit Internet. Telecommunications (International ed.), 27(8):61–??, August 1, 1993. CODEN TLCOAY. ISSN 0040-2494, 0278-484X.

Balas:1993:LAI

[Bal93b] Janet Balas. Learning about the Internet online: Best way to learn about telecommunications is to go online. Computers in Libraries, 13(9):26– ??, October 1, 1993. CODEN CPLIE8. ISSN 1041-7915.

MohsenBanan:1994:IES

[Ban94] Mohsen Banan. Internet email services. Technical report, Electronic Messaging Association, Arlington, VA, USA, 1994. ii + 59 pp.

Banaudi:1995:EMM

[Ban95]  Giorgio Banaudi.  E’ di

moda il Modem: dai BBS a Internet. Muzzio societa. F.Muzzio, Padova, Italy, 1995. ISBN 88-7021-713-2. v + 143 pp. LCCN ????

Barron:1992:LCI

[Bar92a] List of contacts for Internet on-line bibliographic database, August 10, 1992. University of North Texas, Denton, TX, USA.

Barron:1992:NIA

[Bar92b] Numeric IP address of Internet online bibliographic database, August 10, 1992. University of North Texas, Denton, TX, USA.

Barron:1993:IBB

[Bar93] Billy Barron. Internet: Billy Barron’s guide to online bibliographic databases: a list of Telnet addresses. Technical report, University of North Texas, Denton, TX, USA, August 31, 1993. 252 pp.

Barber:1994:BRI

[Bar94] Putnam Barber. Book reviews: (Internet tour guides). Linux Journal, 5:??, September 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Barlow:1995:CSC

[Bar95a] John Perry Barlow. Cover story: Cyberhood vs. neighborhood. is there A there in cyberspace? The Utne reader, 68:52–??, March 1, 1995. ISSN 8750-0256.

Barrett:1995:HWI

[Bar95b] Rosemary Barrett. How we introduce folks to the Internet in one easy lesson, or A follow-up presentation on: networking skills for faculty development. Cause information resources library, CAUSE, Boulder, CO, USA, 1995. 177–185 pp.

WDTBAS:1995:BPI

[BAS95] BAS presents Internet and email, 1995. 1 videocassette quick reference card + notes.

Bassiri:1999:PAN

[Bas99] Kaveh Gh. Bassiri. Programming applications for Netscape servers. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-41970-X. xxx + 975 pp. LCCN TK5105.888.B375 1999. US$49.95.

Bauwens:1993:PMI

[Bau93] Michel Bauwens. The poor man’s Internet: reaching the networks with e-mail only. ASLIB Proceedings, 45 (7 / 8):201–??, July 1, 1993. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Baum:1995:WII

[Bau95] Alaina M. Baum. Women and the Internet: an interpretation of two discourses. Thesis (M.S.), University of Oregon, Eugene, OR, USA, 1995. vi + 149 pp.

Baxter:1995:GSI

[Bax95]  R. Earl Baxter.  Get-

ting started on the Internet. Parker Publishing, Alexandria, CA, USA, 1995. ISBN 1-900369-00-1 (paperback). 59 pp. LCCN ????

Bayles:1999:EBB

[Bay99] Deborah L. Bayles. Extranets: Building the Businessto-Business Web. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999.

ISBN 0-13-650912-6. xvi + 347 pp. LCCN HD30.382.B39

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0136509126.html.

Breeden:1993:DIL

[BB93] Laura Breeden and Lawrence Bouman. Demythologizing the Internet: Lessons from the real world. Educom review, 28(5):29–??, September 1, 1993. CODEN EDREEW. ISSN 1045-9146.

BCR:1995:NIT

[BB95] Bell Communications Research and Bellcore solutions. Navigating the Internet a trip through the jungle, 1995. 3 videocassettes (102 min.).

Bornhoevd:1999:PMB

[BB99] C. Bornhoevd and A. P.

Buchmann. A prototype for metadata-based integration of Internet sources. Lecture Notes in Computer Science, 1626:439–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benford:1993:GSS

[BBH93] Steve Benford, Adrian Bullock, and Paul Harvey. Grace: a system to support the development and use of global computer-supported cooperative work (CSCW) applications. Internet research, 3(1): 25–??, Spring 1993. CODEN IRESEF. ISSN 1066-2243.

Belov:1998:EIL

[BBK^{+}98] Serge D. Belov, Sergey V. Bredikhin, Serge P. Kovalyov, Serge A. Kulagin, Sam L. Musher, Natalie G. Scherbakova, and Igor V. Shabalnikov. The emerging Internet landscape in Siberia. Computer Networks and ISDN Systems, 30(16– 18):1657–1662, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/16-18/2004.pdf.

Bernbom:1993:TWT

[BBN^{+}93] Gerry Bernbom, Ann Bristow, James G. Neal, Howard Rosenbaum, Mark Sheehan, and Sherrill Weaver. Two worlds, two cultures how two professions approach the wilderness of Internet resources, 1993. 1 videocassette (ca. 84 min.) notes ([2] leaves; 28 cm.).

Braden:1989:CIC

[BBP89] R. Braden, D. Borman, and C. Partridge. Computing the Internet checksum. Computer Communication Review, 19 (2):86–??, April 1, 1989. CO-DEN CCRED2. ISSN 0146-4833.

Bickerton:1998:CBS

[BBSH98] Pauline Bickerton, Matthew Bickerton, and Kate Simpson-Holley. Cyberstrategy: Business strategy for extranets, intranets and the Internet. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 0-7506-4203-3. vi + 182 pp. LCCN ???? US$54.95.

Boyce:1996:EIC

[BBST96] Jennifer Boyce, Joan Bewley, Cindy Stell, and John Tofanelli. Exploring the Internet: computing skills workshop, 1995–1996. College custom series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-064684-8. 98 pp. LCCN ????

Bell:1995:MIY

[BC95a] Gordon Bell and University Video Communications. Mapping the Internet yesterday, today and tomorrow, 1995. 1 videocassette (38 min.).

Benedikt:1995:MEV

[BC95b] Claire Lisette Benedikt and Dave Ciskowski. MUDS: exploring virtual worlds on the Internet. BradyGames, Indianapolis, IN, USA, 1995. ISBN 1-56686-246-9 (paperback). vi + 282 pp. LCCN ????

Bickford:1995:ITG

[BC95c] Jeffrey R. Bickford and Salem State College. The Internet travel guide. College custom series. McGraw Hill, Inc., New York, NY, USA, 1995. ISBN 0-07-005780-X. 356 pp. LCCN ????

Branwyn:1995:MQT

[BC95d] Gareth Branwyn and Sean Carton. Mosaic Quick Tour, Special Edition. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-215-1 (Macintosh), 1-56604-214-3 (Windows). 232 pp. LCCN TK5105.875.I57 B72 1994. US$19.95 (Macintosh), US$24.95 (Windows). Includes disks.

Bal:1999:IPL

[BCB99] Henri Bal, Luca Cardelli, and Boumediene Belkhouche, editors. Internet programming languages: ICCL’98 Workshop, held in Chicago, IL, USA, May 13, 1998; proceedings, volume 1686 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66673-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.625 I23 1998.

Bellovin:1995:FSI

[BCRL95] Stephen M. Bellovin, William R. Cheswick, Catherine Reclus, and Philippe Langlois. Firewalls et securité Internet. Collection Brian W. Kernigham. Addison-Wesley France, Paris, France, 1995. ISBN 2-87908-099-1. xvi + 342 pp. LCCN ????

Bussey:1995:NRO

[BCW95] Andrew Busey, Larry Colker, and Hank Weghorst. New Riders Official World Wide Web Yellow Pages. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-449-X. 672 pp. LCCN TK5105.888 .B87 1995. US$29.95. Includes disk.

Braden:1988:NIS

[BD88] Robert T. Braden and Annette L. DeSchon. NN-Stat: Internet statistics collection package, introduction and user guide. ISI research report ISI/RR-88-206, University of Southern California, Information Sciences Institute, Marina del Rey, CA, USA, 1988. 38 pp.

Barnett:1995:IUM

[BD95a] Paul Barnett and Steve Ducharme. The Internet the ultimate marketing and communication tool!, 1995. ISBN 1-887119-05-1. 1 videocassette (ca. 45 min.).

Bretthauer:1995:IUI

[BD95b] David Bretthauer and Christine Drew. An introduction to using the Internet at Southern Connecticut State University. Technical report, Southern Connecticut State University, New Haven, CT 06515, USA, 1995. 94 pp.

Branwyn:1995:IRA

[BDC^{+}95] Gareth Branwyn, Luke Duncan, Sean Carton, Donald Rose, Tom Lichty, Shannon R. Turlingon, and Jan Weingarten. Internet roadside attractions: sites, sounds and scenes along the information superhighway. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-193-7. xxxiv + 320 pp. LCCN TK5105.875.I57 I573 1995.

Borchers:1996:LRG

[BDKK96] Jan Borchers, Oliver Deussen, Arnold Klingert, and Clemens Knörzer. Layout rules for graphical Web documents. Computers and Graphics, 20 (3):415–426, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.cgi?
year=1996&volume=20&issue=
3&aid=9600011.

Bowman:1994:SIR

[BDMS94] C. Mic Bowman, Peter B. Danzig, Udi Manber, and Michael F. Schwartz. Scalable Internet resource discovery: research problems and approaches. Communications of the ACM, 37 (8):98–107, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/

toc/Abstracts/0001-0782/
179704.html.

Berntsen:1985:MLI

[BDPS85] Janet A. Berntsen, James R.

Davin, Daniel A. Pitt, and Neil G. Sullivan. MAC layer interconnection of IEEE 802 local area networks. Computer Networks and ISDN Systems, 10(5):259–273, December 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bowman:1993:RPS

[BDS93] C. Mic Bowman, Peter Danzig, and Michael P. Schwartz. Research problems for scalable Internet resource discovery. Technical report CU-CS-643-93, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, March 1993. 12 pp.

Bean:1994:ISS

[Bea94] Greg Bean. Internet servers:

a step-by-step guide on how to build them. CyberGroup, Inc., Baltimore, MD, USA, third edition, 1994. ISBN 0-9642041-1-8. 9 + 159 pp. LCCN ????

Bean:1995:ISC

[Bea95] Greg Bean. Internet server construction kit for Windows. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-12696-9 (paperback). xviii + 412 pp. LCCN TK5105.875.I57B42 1995.

Bonjour:1998:IAN

[BEA98] Dominique Bonjour, Omar Elloumi, and Hossam Afifi. Internet applications over native ATM. Computer Networks and ISDN Systems, 30(12):1097–1110, July 13, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/12/1815.pdf.

Beebe:1993:BBA

[Bee93] Nelson H. F. Beebe. A Bibliography of Books about the Internet and Networking, 1993. This bibliography in BibTEX format includes the Quarterman list [Qua93], and many others. It, and many others maintained by the same author, are updated frequently, and available for anonymous

ftp on ftp.math.utah.edu
in /pub/tex/bib. Via e-

mail, send mail to tuglib@

math.utah.edu with the text
help and send index from
tex/bib.

Bellovin:1993:PFI

[Bel93] S. Bellovin. Packets found on an Internet. Computer Communication Review, 23(3):26– ??, July 1, 1993. CODEN CCRED2. ISSN 0146-4833.

Bell:1995:RGG

[Bel95] Joseph Grant Bell. Rec.g@mes:

games and game resources on the Internet. Brady Games, Indianapolis, IN, USA, 1995. ISBN 1-56686-311-2. 414 pp. LCCN ????

Briggs-Erickson:1996:EGI

[BEM96] Carol Briggs-Erickson and Toni Murphy. Environmental guide to the Internet. Government Institutes, Inc., Rockville, MD, USA, second edition, 1996. ISBN 0-86587-517-0. xxiv + 210 pp. LCCN ????

Benedikt:1991:FS

[Ben91] Michael Benedikt, ed. Cyberspace: First Steps. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-02327-X. viii + 436 pp. LCCN QC173.59.S65 C93 1991. US$15.95. An anthology of sociological examinations of networks and related topics, by writers, scholars, and public figures. Edited by a professor of Architecture at the University of Texas.

Benavides:1992:IIP

[Ben92] Alba E. Benavides. Implementing Internet in public schools in Northeast Louisiana. Thesis (Ed.S.), Northeast Louisiana University, Monroe, LA, USA, 1992. xii + 217 pp.

Benson:1994:CIC

[Ben94] Allen C. Benson. The complete Internet companion for librarians. Neal-Schuman Publishers, New York, NY, USA, 1994. ISBN 1-55570-178-7. ???? pp. LCCN TK5105.875.I57 B46 1994.

Bender:1995:CCI

[Ben95a] Robert M. Bender. Creating communities on the Internet: Electronic discussion lists in the classroom. Computers in Libraries, 15(5):38– ??, May 1, 1995. CODEN CPLIE8. ISSN 1041-7915.

Benson:1995:CIC

[Ben95b] Allen C. Benson. The complete Internet companion for librarians. Neal-Schuman Publishers, New York, NY, USA, 1995. ISBN 1-55570-178-7. xxiv + 405 pp. LCCN TK5105.875.I57 B46 1995.

Benson:1995:DEE

[Ben95c] Michael L. Benson. A distance education environment using multimedia conferencing and Internet based delivery. Thesis (M.S.), George Mason University, Fairfax, VA, USA, 1995. xi + 153 pp.

Bentall:1998:AIP

[Ben98] M. Bentall. ATM and Internet Protocol. New English Library, London, UK, 1998. ISBN 0-340-71921-4. viii + 168 pp. LCCN TK5103.75 .B46 1998. US$54.95.

Bernard:1995:IBT

[Ber95a] Ryan Bernard. Internet business 500: the top 500 essential Web sites for business. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-287-9 (paperback). xxvi + 450 pp. LCCN HF 54.56 B47 1995. “Fully hyperlinked text on CD-ROM for Mac and Windows”–Cover.

Bernstein:1995:FIS

[Ber95b]  Terry Bernstein.  Funda-

mentals of Internet security. Technical Report D95-1961, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, November/December 1995. 10 pp.

Berry:1995:PTM

[Ber95c] Colin Berry. A pocket tour of music on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1695-7 (paperback). xvii + 206 pp. LCCN TK5105.875.I57 B47 1995. US$12.99.

Bernard:1996:CIC

[Ber96a] Ryan Bernard. The corporate Internet: create and manage an internal World Wide Web network. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-14929-2 (paperback). various pp. LCCN TK5105.888.B47 1996.

Berry:1996:PTC

[Ber96b] Colin Berry. A pocket tour of celebrities on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1811-9. xv + 191 pp. LCCN ????

Betten:1987:IPR

[Bet87] John R. Betten. IRMSP: a protocol for real-time communication in the DARPA Internet. Master of science, plan ii., Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1987. 19 pp.

Burkhart:1998:IPIa

[BG98] Grey E. Burkhart and Seymour E. Goodman. International perspectives: the Internet gains acceptance in the Persian Gulf. Communications of the ACM, 41(3): 19–25, March 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-3/
                                        p19-burkhart/.

Bischof:1995:EII

[BGGS95] Hans-Peter Bischof, Jutta Goeers, Petra Gremeyer, and Axel T. Schreiner. Einblicke ins Internet. Gebührenfreies Training für das World Wide Web (WWW) (English: Insight in Internet. Chargefree training for World Wide Web (WWW)). Carl Hanser, München, Germany, 1995. ISBN 3-446-18512-7. 20 pp. LCCN ???? Includes CD-ROM.

Burkhart:1998:IPIb

[BGMP98] Grey E. Burkhart, Seymour E. Goodman, Arun Mehta, and Larry Press. International perspectives: the Internet in India: better times ahead? Communications of the ACM, 41(11): 21–26, November 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-11/
p21-burkhart/.

Beatty:1994:IWA

[BGS94] Grace Joely Beatty, David C.

Gardner, and David A. Sauer. Internet for Windows: America Online Edition. Prima Publishing, Rocklin, CA, USA, December 1994. ISBN 1-55958-554-4. vii + 275 pp. LCCN TK5105.875.I57 B43 1995. US$19.95, CDN$27.95.

Beatty:1995:IWA

[BGS95] Grace Joely Beatty, David C.

Gardner, and David A. Sauer. Internet for Windows: America Online edition. Visual learning guides. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 1-55958-554-4. vii + 275 pp. LCCN TK 5105.875 I57 B43 1995. US$19.95, UK£18.49, CDN$27.95.

Brumbaugh:1993:FEK

[BH93] Ken Brumbaugh and Nick Hart. ftp: The educator’s key to Multi-Platform file transfers. T H E journal (Technological Horizons in Education), 20(10):65–??, May 1, 1993. CODEN THEJD4. ISSN 0192-592X.

Battle:1994:AAR

[BH94a] Stafford L. Battle and Rey O. Harris. African American Resource Guide to the Internet. OnDemand Press, Columbia, MD, USA, 1994. ISBN 1-886758-00-X. ???? pp. LCCN ????

Boland:1994:CIN

[BH94b] Mary Jo Boland and Carolyn House. Cruising the Internet: NEMICRO workshops, summer 1994. Technical report, NELINET, Newton, MA, USA, 1994. 30 pp.

Battle:1995:AAR

[BH95] Stafford L. Battle and Rey O. Harris. The African American resource guide to the Internet. OnDemand Press, Columbia, MD, USA, 1995. ISBN 1-886758-00-X. x + 85 pp. LCCN TK5105.875.I57 B38 1995.

Battle:1996:AAR

[BH96] Stafford L. Battle and Rey O.

Harris. The African American resource guide to the Internet and online services. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-005499-1 (paperback). various pp. LCCN E185.B33 1996.

Bhimani:1996:SCI

[Bhi96]  Anish Bhimani.  Secur-

ing the commercial Internet. Communications of the ACM, 39(6):29–35, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228509.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228509.html.

Birman:1999:BM

[BHO^{+}99] Kenneth P. Birman, Mark Hayden, Oznur Ozkasap, Zhen Xiao, Mihai Budiu, and Yaron Minsky. Bimodal multicast. ACM Transactions on Computer Systems, 17 (2):41–88, May 1999. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocs/1999-17-2/
p41-birman/.

Barron:1996:IIA

[BI96] Ann E. Barron and Karen S.

Ivers. The Internet and instruction: activities and ideas. Libraries Unlimited, Englewood, CO, USA, 1996. ISBN 1-56308-331-0. xiii + 159 pp. LCCN LB1044.87 .B37 1996.

Bierce:1993:CST

[Bie93] Ambrose Bierce. Can such things be?, 1993. Includes CD-ROM.

Bierce:1994:DD

[Bie94] Ambrose Bierce. The devil’s dictionary, 1994.

Billard:1999:TSI

[Bil99] D. Billard. Transactional services for the internet. Lecture Notes in Computer Science, 1590:14–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1996:BSI

[Bio96] Biotechnology software and Internet journal, page various, 1996. ISSN 1088-4270. Mary Ann Liebert, Inc., Larchmont, NY, USA.

Birchfield:1990:CNN

[Bir90] Marilee Birchfield, editor. Casting a new net: searching library catalogs via the Internet. ????, ????, 1990. ISBN ???? LCCN ????

Biro:1993:BTI

[Bir93] Ross Biro. A browsing tool for the Internet logical library of the HPCC software exchange. NASA contractor report NASA CR-194545 4004510452, National Aeronautics and Space Administration, Washington, DC, USA, 1993. Distributed to depository libraries in microfiche. Shipping list no.: 94-0085-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1993] 1 microfiche. RIACS technical report TR 93-09.

Bird:1994:TRN

[Bir94] David Bird. Token Ring Network Design. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-62760-4. x + 220 pp. LCCN TK5105.8.I24 B57 1994. US$29.25.

Borg:1994:I

[BJS94] Tage Borg, Jens Jonason, and Christer Sturmark. Internet@Sverige. Bonnier Data-Media, ????, 1994. ISBN 91-644-0010-7. 344 pp. 298 SEK. The Windows’ users guide Internet in Sweden.

Barker:1991:CIX

[BK91] P. Barker and S. Kille. The COSINE and Internet X.500 schema, November 1991. Available via FTP.

Bounds:1996:HGY

[BK96a] Shannon Bounds and Arthur Karl. How to get your dream job using the Internet. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-883577-68-3. 336 pp. LCCN ????

Broll:1996:VTT

[BK96b] Wolfgang Broll and Tanja Koop. VRML: Today and tomorrow. Computers and Graphics, 20(3): 427–434, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.cgi?
year=1996&volume=20&issue=
3&aid=9500012.

Bhatti:1999:EQA

[BK99a] Saleem N. Bhatti and Graham Knight. Enabling QoS adaptation decisions for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 31(7): 669–692, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/7/2100.pdf.

Boni:1999:WRC

[BK99b] William C. Boni and Gerald L. Kovacich. I-Way Robbery: Crime on the Internet. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 0-7506-7029-0. xv + 240 pp. LCCN HV6773 .B66 1999. US$41.95.

Bush:1995:HUA

[BKH95] David Bush, Dave Kuenstle, and Alan Hoffman. How to understand, access and use the Internet, 1995. ISBN 1-56414-176-4 (book). 2 videocassettes.

Barboni:1993:ICI

[BL93] Edward J. Barboni and Karen Lukas. Independent colleges and the Internet: a primer for college administrators. Series of papers on information technologies 2, Council of Independent Colleges, Washington, DC, USA, September 1993. 20 pp.

Badgett:1994:IA

[BL94] Tom Badgett and LearnKey, Inc. Internet atlas, 1994. 1 videocassette (146 min.).

Badgett:1995:IE

[BL95a] Tom Badgett and LearnKey, Inc. Internet e-mail, 1995. 1 videocassette (86 min.).

Badgett:1995:IEI

[BL95b] Tom Badgett and LearnKey, Inc. Internet executive introduction, 1995. 1 videocassette (37 min.).

Badgett:1995:IT

[BL95c] Tom Badgett and LearnKey, Inc. Internet tools, 1995. 1 videocassette (101 min.).

Blum:1995:EMF

[BL95d] Daniel J. Blum and David M.

Litwack. The E-Mail Frontier: Emerging Markets and Evolving Technologies. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-56860-8. xx + 476 pp. LCCN HE6239.E54 B58 1995. US$46.25.

Braun:1996:IDV

[BL96] Eric (Eric Eugene) Braun and Mark Leger. The Internet directory version 2.0. Fawcett Columbine, New York, NY, USA, 1996. ISBN 0-449-98370-6. various pp. LCCN ???? Berners-Lee:1997:HCW

[BL97] Tim Berners-Lee. The human connection: Worldwide computer. Communications of the ACM, 40(2): 57–58, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-2/
p57-berners-lee/.

Black:1992:TIR

[Bla92] Uyless D. Black. TCP/ IP and Related Protocols. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-005553-X. xvii + 372 pp. LCCN TK5105.5.B5663 1992. US$44.95. Explains the TCP/IP suite gateway standards (GGP. EGP RIP etc.).

Blackman:1993:LRI

[Bla93a] Josh Blackman. The legal researcher’s Internet directory:

1993/1994. Technical report, Legal Research of New York, Inc., Brooklyn, NY, USA, 1993. viii + 137 pp.

Blake:1993:NIS

[Bla93b] Paul Blake. News, the Internet and student training: debate at the National Online Meeting. Online & CDROM review: the international journal of online & optical information systems, 17 (3):186–??, June 1, 1993. CO-DEN ONCDEW. ISSN 0309-314X.

Blau:1993:KAI

[Bla93c] Andrew Blau. K-12 access to the Internet: Securing the legal framework. Journal of science education and technology, 2(3):497–??, September 1, 1993. CODEN JSEEEP. ISSN 1059-0145.

Black:1995:AVF

[Bla95a]  Uyless D. Black.  ATM

Volume I: Foundation for Broadband Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-297178-X. 448 pp. LCCN TK5105.35.B53 1995. US$58.00. URL http:/

/www.phptr.com/ptrbooks/
ptr_013297178x.html.

Black:1995:TIR

[Bla95b] Uyless D. Black. TCP/IP and Related Protocols. McGraw-Hill, New York, NY, USA, second edition, 1995. ISBN 0-07-005560-2. xxiii + 375 pp. LCCN TK5105.55.B53 1995. US$50.00. Explains the TCP/IP suite gateway standards (GGP. EGP RIP etc.).

Blackman:1995:BRI

[Bla95c] Josh Blackman. The business researcher’s introduction to the Internet: training workbook. FIND/SVP, Inc., New York, NY, USA, 1995. ISBN 1-56241-222-1. LCCN ????

Blagojevic:1995:ZEC

[Bla95d] Bonnie Blagojevic. A to Z: the early childhood educator’s guide to the Internet. Technical report, ERIC Clearinghouse on Elementary and Early Childhood Education University of Illinois at Urbana-Champaign, Urbana, IL, USA, June 1995. various pp.

Black:1996:MWN

[Bla96a] Uyless D. Black. Mobile and Wireless Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-440546-3. xx + 348 pp. LCCN TK5103.2.B53 1996. US$58.00. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0134405463.html.

Blake:1996:SAC

[Bla96b] Carol Blake. Shop around the clock. Internet World, 7(1): 70–??, January 1, 1996. CO-DEN IERNE8. ISSN 1064-3923.

Blakeman:1996:BII

[Bla96c] Karen Blakeman. Business information on the Internet. RBA Information Services, Caversham, New Zealand (??), 1996. ISBN 0-9527191-0-X (paperback). 123 pp. LCCN ????

Black:1997:ECT

[Bla97a] Uyless D. Black. Emerging Communications Technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1997. ISBN 0-13-742834-0. xxi + 457 pp. LCCN TK5101.B553 1997. US$51.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0137428340.html.

Black:1997:ISA

[Bla97b]  Uyless D. Black.  ISDN

and SS7: Architectures for Digital Signaling Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-259193-6. xxi + 345 pp. LCCN TK5103.75.B53 1997. US$61.00. URL http://www.

phptr.com/ptrbooks/ptr_
0132591936.html.

Black:1998:MSL

[Bla98a] Darryl P. Black. Managing Switched Local Area Networks: a Practical Guide. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-18554-7. xix + 356 pp. LCCN TK5105.7.B57 1998. US$44.95.

Black:1998:ARL

[Bla98b] Uyless D. Black. ATM Resource Library (Volumes I, II, III). P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-793779-2. LCCN ???? US$149.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0137937792.html. Three

volumes.

Black:1998:AVIa

[Bla98c] Uyless D. Black. ATM Volume II: Signaling in Broadband Networks. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-571837-6. 224 pp. LCCN TK5105.35.B53 1995. US$69.00. URL http://www.

                                        phptr.com/ptrbooks/ptr_
                                        0135718376.html.

Black:1998:AVIb

[Bla98d] Uyless D. Black. ATM, Volume III: Internetworking With ATM. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-784182-5. 256 pp. LCCN TK5105.35.B53 1995. US$55.00. URL http://www.

                                        phptr.com/ptrbooks/ptr_
                                        0137841825.html.

Black:1998:INC

[Bla98e] Uyless D. Black. The intelligent networks: customizing telecommunication networks and services. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-793019-4. xvi + 208 pp. LCCN TK5101.B5528

1998. URL http://www.
prenhall.com/allbooks/ptr_
0137930194.html.

Black:1998:RBN

[Bla98f] Uyless D. Black. Residential Broadband Networks: XDSL, HFC and Fixed Wireless Access. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-956442-X. 400 pp. LCCN TK5103.4.B54 1998. US$61.00. URL http://www.

phptr.com/ptrbooks/ptr_
0132591936.html.

Black:1999:BSN

[Bla99a] Darryl P. Black. Building switched networks: multilayer switching, QoS, IP multicast, network policy, and service level agreements. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37953-8. xix + 298 pp. LCCN TK5105.5.B555 1999. US$44.95.

Black:1999:ARL

[Bla99b] Uyless Black. ATM Resource Library. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-083786-5. LCCN

????  URL http://www.
phptr.com/ptrbooks/ptr_
0130837865.html. Boxed

set of ATM, Volume I: Foundations for Broadband Networks, Second Edition; ATM, Volume II: Signaling in Broadband Networks and ATM, Volume III: Internetworking.

Black:1999:AIT

[Bla99c] Uyless D. Black. Advanced Internet Technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-759515-8. 380 pp. LCCN TK5105.875.I57B55 1998. US$49.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0137595158.html.

Black:1999:AFB

[Bla99d] Uyless D. Black. ATM: Foundation for Broadband Networks, Volume 1. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-083218-9. ???? pp. LCCN TK5105.35 .B53

                                        1999. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0130832189.html.

Black:1999:SGM

[Bla99e]  Uyless D. Black.  Sec-

ond Generation Mobile and Wireless Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-621277-8. 400 pp. LCCN TK5103.2.B5323 1999. US$64.00. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0136212778.html.

Black:1999:VI

[Bla99f]  Uyless D. Black.  Voice

over IP. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-022463-4. ???? pp. LCCN TK5105.8865.B53

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130224634.html.

Barrett:1999:EJW

[BLB99] Dan Barrett, Dan Livingston, and Micah Brown. Essential JavaScript for Web Professionals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-013056-7. ???? pp. LCCN ???? URL http:/

/www.phptr.com/ptrbooks/
ptr_0130130567.html.

Blegen:1993:EIT

[Ble93] John C. Blegen. Exploring Internet teleconference, 1993. 1 videocassette (ca. 1 hr. 35 min.).

Bleuel:1995:OPI

[Ble95] Jens Bleuel. Online publizieren im Internet: elektronische Zeitschriften und Bucher. Edition Ergon, Pfungstadt, Germany, 1995. ISBN 3-927960-28-4. 183 pp. LCCN ????

Berners-Lee:1999:WWO

[BLF99] Tim Berners-Lee and Mark Fischetti. Weaving the Web: the original design and ultimate destiny of the World Wide Web by its inventor. HarperCollins Publishers, New York, NY, USA, 1999. ISBN 0-06-251586-1 (hardcover), 0-06-251587-X (paperback). xi + 226 pp. LCCN K5105.888 .B46 1999.

Bonar:1996:PCI

[BLL^{+}96] Jeffrey Bonar, Nancy Lehrer, Kelly Looney, Randy Schnier, James Russel, Ted Selker, and Stewart Nickolas. Panel 6: Components on the Internet. ACM SIGPLAN Notices, 31(10):444–448, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blomeley:1994:NNS

[Blo94] Frances Blomeley. Networks and Network Services — A User’s Guide. Immediate Publishing, ??, May 1994. ISBN 1-898931-00-3. 246 pp. UK£14.95, US$25.95. A broad overview of most aspects of networks and network services from a U.K. user’s perspective.

Blotzer:1995:IUG

[Blo95] Michael J. Blotzer. Internet user’s guide for safety and health professionals. Technical report, Genium Publishing Corp., Schenectady, NY, USA, 1995. viii + 152 pp.

Bowerman:1994:II

[BM94] Jane Bowerman and Vicki McWilliams. Introduction to the Internet. Technical report, University of New Mexico, Computer and Information Resources and Technology (CIRT), Albuquerque, NM, USA, 1994. 100 pp.

Biggar:1996:DZW

[BM96a] Bill Biggar and Joe Myers.

Danger zones: what parents should know about the Internet. Andrews and McMeel, Kansas City, MO, USA, 1996. ISBN 0-8362-1317-3 (paperback). xiv + 129 pp. LCCN HQ784.I58B54 1996.

Bradner:1996:IIP

[BM96b] Scott O. Bradner and Allison Mankin. IPng, Internet protocol next generation. Addison-Wesley IPng series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63395-7. xxxi + 307 pp. LCCN TK5105.875.I57 I67 1996.

Beck:1998:IDS

[BM98] Micah Beck and Terry Moore. The Internet2 Distributed Storage Infrastructure project: an architecture for Internet content channels. Computer Networks and ISDN Systems, 30(22– 23):2141–2148, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/22-23/2051.pdf.

Bang:1994:IU

[BMG^{+}94] Steve Bang, Martin Moore, Rick Gates, et al. The Internet Unleashed. SAMS Publishing, 1994. ISBN 0-672-30466-X. lviii + 1387 pp. US$44.95. Probably the thickest book (much over 1000 pages) covering almost all topics you want to know about the Internet …. The disk contains PC Windows software for connecting with and getting around on the Internet.

Boggs:1988:MCE

[BMK88] David R. Boggs, Jeffrey C. Mogul, and Christopher A. Kent. Measured capacity of an Ethernet: Myths and reality. ACM SIGCOMM, page 31, August 1988.

BBN:1990:IMP

[BN90] Beranek Bolt Newman, Inc. and NSF Network Service Center. Internet manager’s phonebook. Technical report, NSF Network Center, Bolt Beranek and Newman, Cambridge, MA, USA, 1990. 219 pp.

Brown:1996:DAO

[BN96] Marc H. Brown and Marc A. Najork. Distributed active objects. Computer Networks and ISDN Systems, 28(7–11):1037–1052, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.

cgi?year=1996&volume=28&
issue=7-11&aid=1586.

Beach:1993:IBS

[BOB93] J. H. Beach, S. J. Ozminski, and D. E. Boufford. An Internet botanical specimen data server. Taxon, 42(3):627–??, August 1, 1993. CODEN TAXNAP. ISSN 0040-0262.

Bock:1996:DBI

[Boc96] Wally Bock. Doing business on the Internet, 1996. ISBN 1-55852-203-4. 6 sound cassette.

Boggs:1982:IB

[Bog82] David Reeves Boggs. Internet broadcasting. Thesis (Ph.D.), Stanford University, Stanford, CA, USA, 1982. viii + 86 pp.

BogarinNavarro:1994:DMI

[Bog94] Rodrigo Bogarin Navarro.

Descubra el mundo de Internet. Editorial Tecnologica de Costa Rica, Cartago, Costa Rica, 1994. ISBN 9977-66-074-3. 171 pp. LCCN TK5105.875.I57 B64 1994.

BBN:1994:AIB

[Bol94] Bolt, Beranek and Newman, Cambridge, MA, USA. ARPANET and INTERNET: a bibliography of BBN papers, 1994. 52 pp.

Bolen:1996:SDC

[Bol96a] Dennis Edward Bolen. To surf and detect: crime and punishment on the Internet. Go!Guides series. Motion Works Publishing, San Francisco, CA, USA, 1996. ISBN 1-57712-004-3 (paperback). ???? pp. LCCN HV6024.5.B65 1996.

Bolot:1996:CQT

[Bol96b]  Jean-Chrysostome  Bolot.

Cost-quality tradeoffs in the Internet. Computer Networks and ISDN Systems, 28(5):645–651, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=5&aid=1496.

Boschert:1994:EZL

[Bos94] Ken Boschert. The electronic zoo: a list of animal-related computer resources (internet/Bitnet mailing lists, gophers, World Wide Web sites, mail servers, Usenet newsgroups, FTP sites, commercial online services, and bulletin board systems). Technical report, ????, ????, 1994. 196 pp.

Bourne:1997:CSI

[Bou97] Philip E. Bourne. A cookbook for serving the Internet: UNIX version. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-519992-1. xxix + 303 pp. LCCN QA76.9.C55 B68 1997.

Boumphrey:1998:XA

[Bou98] Frank Boumphrey. XML applications. Wrox Press, Chicago, IL, USA, 1998. ISBN 1-86100-152-5. xix + 649 pp. LCCN ????

Bowman:1990:UCI

[Bow90] C. Mic Bowman. Univers:

the construction of an Internetwide descriptive naming system. Washington, DC, USA, 1990. Distributed to depository libraries in microfiche. Shipping list no.: 92-1583-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1992] 1 microfiche.

Bowbrick:1995:PTM

[Bow95] Steve Bowbrick. Piecing together Mosaic: Navigating the Internet and the World Wide Web. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-142-6. 300 pp. LCCN ????

Bowdidge:1996:RIR

[Bow96] Robert Bowdidge. Railroadrelated Internet resources, 1996.

Boyer:1996:CCC

[Boy96] Bill Boyer. Coalition for Christian colleges and universities Internet list. Technical report, ????, ????, 1996. various pp.

Braden:1987:RIG

[BP87] R. Braden and J. Postel. Requirements for Internet gateways. Technical Report RFC 1009, Internet Engineering Task Force. Network Working Group, ????, June 1987. 55 pp.

Bowman:1991:NOI

[BP91] Mic Bowman and Larry L. Peterson. Naming objects in an Internet environment. Technical report CS-91-30, Pennsylvania State University, Dept. of Computer Science, University Park, PA, USA, December 1991. 7 pp.

Besser:1993:ICN

[BP93] Howard Besser and Paul Evan Peters. Internet computer network, 1993. Videocassette. Lecture sponsored by the School of Library and Information Studies, University of California, Berkeley. This event took place during a lecture to the Library and Information Title supplied by cataloger. Performers: Howard Besser, presenter; Paul Peters, lecturer. Paul Peters, founder of the Coalition for Networked Information, speaks on the Internet computer network. He surveys the current features of Internet and introduces areas where he feels fundamental shifts in thinking and design are possible by creators and users of network information sources.

Butler:1994:ITS

[BP94] Mark Butler and Cecile Pierchon. Internet… tout de suite. Dunod, Paris, France, 1994. ISBN 2-10-002463-9. xi + 145 pp. LCCN ????

Bowman:1988:UNS

[BPR88] Mic Bowman, Larry L. Peterson, and Herman Rao. Univers: a name server for the next generation Internet. Technical report TR 88-17, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, 1988. 19 pp.

Bennett:1999:PRP

[BPS99] Jon C. R. Bennett, Craig Partridge, and Nicholas Shectman. Packet reordering is not pathological network behavior. IEEE/ACM Transactions on Networking, 7 (6):789–798, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/1999-
7-6/p789-bennett/p789-bennett.
pdf;   http://www.acm.org/
pubs/citations/journals/
ton/1999-7-6/p789-bennett/

.

Banks:1994:DOG

[BR94a] Michael A. Banks and Eric Romanik. DELPHI: the official guide. Delphi Internet Services, Cambridge, MA, USA, 1994. ISBN 0-9625623-2-7. 289 pp. LCCN ????

Benneman:1994:III

[BR94b] Tom Benneman and Bob Rieger. Introduction to Internet information superhighway, 1994. 1 videocassette.

Bentson:1995:TSC

[BR95] Randolph Bentson and Arnold Robbins. The trade shows: Comdex Spring and Internet World. Linux Journal, 16:??, August 1995. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Braden:1989:RIHb

[Bra89a] R. Braden. Requirements for Internet hosts: application and support. Technical report, Internet Engineering Task Force, ????, 1989. 98 pp.

Braden:1989:RIHa

[Bra89b] R. Braden. Requirements for Internet hosts: communications layers. Technical report, Internet Engineering Task Force, ????, 1989. 116 pp.

Brambert:1993:GI

[Bra93a] Dave Brambert, ed. Guide to Internetworking: Interconnecting Diverse Networks: DOS, Macintosh, UNIX, Mainframe. Miller Freeman Publications, San Francisco, CA, USA, 1993. ISBN 0-87930-262-3. 186 pp.

LCCN TK5105.7 .G85 1993. US$26.95.

Braun:1993:CID

[Bra93b] Eric Braun. The Complete Internet Directory. Fawcett Columbine, New York, NY, USA, 1993. ISBN ?? 325 pp. LCCN ?? US$25. A directory of newsgroups, discussion lists, FTP sites and so on, with just a few pages on how to use these resources.

BP:1994:IDE

[Bra94a] Brandenburg Productions.

The Internet driver’s education for the information superhighway, 1994. 1 videocassette (66 min.).

Branwyn:1994:MQT

[Bra94b] Gareth Branwyn. Mosaic Quick Tour. Mac-edition: Accessing and Navigating the Internet’s World Wide Web. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-195-3 (Mac edition), 1-56604-194-5 (Windows edition). 175 pp. LCCN TK5105.875.I57 B72 1994. US$10.95.

Braun:1994:ID

[Bra94c] Eric (Eric Eugene) Braun.

The Internet directory. Fawcett Columbine, New York, NY, USA, 1994. ISBN 0-449-90898-4. xxii + 704 pp. LCCN TK5105.875.I57B73 1994. US$25.00.

Bradley:1995:QGI

[Bra95]  Julia Case Bradley.  A

quick guide to the Internet. Wadsworth Publishing Company, Pacific Grove, CA, USA, 1995. ISBN 0-534-26076-4. 109 pp. LCCN TK5105.875.I57B68 1995. US$9.95.

Brandt:1997:CTU

[Bra97] D. Scott Brandt. Constructivism: teaching for understanding of the Internet. Communications of the ACM, 40(10):112–117, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        1997-40-10/p112-brandt/.

Breeden:1991:TI

[Bre91] Laura Breeden. Tour of the Internet, 1991. 1 computer disk.

BrettIi:1993:NIR

[Bre93] George H. Brett II. Networked information retrieval tools in the academic environment: Towards a cybernetic library. Internet research, 3 (3):26–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

Bremer:1994:IIF

[Bre94] Marian Bremer. Incorporating Internet full-text and reference resources into a library collection. Technical report, Special Libraries Association, Washington, DC, USA, 1994. 7 + 21 pp.

Breeding:1995:TII

[Bre95a] Marshall Breeding. TCP/IP for the Internet: the complete buyer’s guide to micro-based TCP/IP software. Mecklermedia, Westport, CT, USA, 1995. ISBN 0-88736-980-4. xii + 305 pp. LCCN TK5105.585 .B74 1995.

Breuner:1995:COS

[Bre95b] Emily F. (Emily Florence)

Breuner. Complexity and organizational structure: Internet and visa international as prototypes for the corporation of the future. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1995. 86 pp.

Brenton:1998:MNS

[Bre98] Chris Brenton. Mastering Network Security. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 0-7821-2343-0. xxvi + 672 pp. LCCN QA76.9.A25 B757 1998. US$39.99.

Briggs:1994:ECR

[Bri94a] York M. Briggs. An exploration of computer-assisted reporting via the Internet for the community journalist: access and manipulation of the Internet with the Macintosh computing system. Thesis (m.a.), Central Missouri State University, Warrensburg, MO, USA, 1994. 134 pp.

BLISS:1994:ILG

[Bri94b] British Library Information Sciences Service. Internet: a literature guide. British Library Information Sciences Service, London, UK, 1994. ISBN 0-7123-4200-1 (paperback). 31 pp. LCCN ????

Bride:1995:I

[Bri95a] Mac Bride. The Internet. Teach yourself books. NTC Pub. Group, Lincolnwood, IL, USA, 1995. ISBN 0-8442-3700-0 (paperback). ix + 214 pp. LCCN ????

BCIT:1995:PCI

[Bri95b] British Columbia Institute of Technology. Media Studio. Point and click Internet kit. Reed Books, Markham, ON, Canada, 1995. ISBN 0-433-39614-8 (PC), 0-433-39615-6 (Mac). ???? pp. LCCN ????

BCMFTA:1995:AFS

[Bri95c] Accessing Forest Service manuals through the Internet, page various, 1995. The Ministry, Victoria, BC, Canada.

BCTDEB:1995:TGCa

[Bri95d] British Columbia. Technology and Distance Education Branch. Towards a global classroom: Internet resources for math and the sciences. Technical report, Queen’s Printer for British Columbia, Victoria, BC, Canada, 1995. 108 pp.

BCTDEB:1995:TGCb

[Bri95e] British Columbia. Technology and Distance Education Branch. Towards a global classroom: Internet resources for social studies. Technical report, Queen’s Printer for British Columbia, Victoria, BC, Canada, 1995. 158 pp.

Bride:1998:HPWa

[Bri98] Mac Bride. HTML publishing on the World Wide Web. Teach yourself Teach yourself books. Hodder and Stoughton, London, UK, second edition, 1998. ISBN 0-340-71220-1. vii + 216 pp. LCCN ????

Brock:1989:NIC

[Bro89]  Jack L. Brock.  Novem-

ber 1988 Internet computer virus and the vulnerability of national telecommunications networks to computer viruses: statement of Jack L. Brock, Director, Government Information … before the Subcommittee on Telecommunications and Finance, Committee on Energy and Commerce, House of Representatives. Washington, DC, USA, 1989. 8 pp. Shipping list no.: 90-146-P.

Brownrigg:1991:IEE

[Bro91] Edwin B. (Edwin Blake) Brownrigg. The Internet as an external economy: the emergence of the invisible hand. Technical report, Memex Research Institute, Roseville, CA., Spring 1991. 3 pp.

Brown:1993:TRG

[Bro93] Ronald O. Brown. TCP/IP: Reliable glue for enterprise networks. Networking management, 11(6):76–??, June 1, 1993. CODEN NETME2. ISSN 1052-049X.

Brown:1994:CLW

[Bro94a] David K. Brown. Children’s literature Web guide Internet resources related to books for children and young adults, 1994.

Browne:1994:IMW

[Bro94b] Steve Browne. The Internet via Mosaic and World-Wide Web. Ziff-Davis Press, Emeryville, CA, USA, September 1994. ISBN 1-56276-259-1. xiii + 156 pp. LCCN TK5105.875.I57 B76 1994. US$29.95.

Brown:1995:CHI

[Bro95a] S. Scott Brown. College health information systems on the Internet: trends and future potential. Master of public health. thesis (M.S.), Department of Health Behavior and Health Education, University of North Carolina at Chapel Hill, Chapel Hill, NC, USA, 1995. 56 pp.

Browne:1995:ITM

[Bro95b] Steve Browne. Internet a traves de Mosaic y Worl-Wide Web. Anaya Multimedia, Madrid, Spain, edicion espanola edition, 1995. ISBN 84-7614-715-5. 208 pp. LCCN ????

Brown:1996:AIC

[Bro96] Peter Brown. 16th annual institute on computer law: understanding the business and legal aspects of the Internet. Patents, copyrights, trademarks, and literary property course handbook series G-430, Practising Law Institute, New York, NY, USA, 1996. 360 pp.

Bryant:1994:CSB

[Bry94] Alan D. Bryant. Creating Successful Bulletin Board Systems. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-62668-3. xviii + 393 pp. LCCN QA76.9.B84 B79 1994. US$39.95.

Bryant:1995:VMC

[Bry95] Wanda Bryant. Virtual music communities: the folk music Internet discussion group as a cultural system. Thesis (Ph.D.), UCLA, Los Angeles, CA, USA, 1995. xvi + 610 pp. Two volumes.

Badgett:1993:WIM

[BS93a] Tom Badgett and Corey Sandler. Welcome to — Internet: from mystery to mastery. MIS Press, New York, NY, USA, 1993. ISBN 1-55828-308-0 (paperback). xi + 324 pp. LCCN TK 5105.875 I57 B33 1993. US$19.95.

Badgett:1993:WI

[BS93b] Tom Badgett and Corey Sandler. Welcome to … Internet: From Mystery to Mastery. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1993. ISBN 1-55828-308-0. xi + 324 pp. LCCN TK 5105.875 I57 B33 1993. US$19.95.

Bufalo:1994:ILR

[BS94] Frances Bufalo and Keith Ann Stiverson. Internet as a library-wide resource using the Internet for legal research, 1994. 1 sound cassette.

Busche:1996:AIB

[BS96] Don Busche and Angela Satterlee. Adventures on the Internet: a beginner’s guide. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1996. ISBN 0-7895-0118-X. xii + 244 pp. LCCN TK5105.875.I57B88 1996.

Buckley:1996:FLD

[BSK^{+}96] William F. (Willam Frank)

Buckley, Warren Steibel, Michael Kinsley, South Carolina Educational Television Network, and Firing Line (Television program). A firing line debate, 1996. 1 videocassette (ca. 120 min.) (57 p.; 22 cm.).

Bell:1996:MWS

[BT96] Mark Bell and Rob Terrell.

The Mac Web server book: tools and techniques for building your Internet site. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-341-7. ???? pp. LCCN TK5105.888.B45 1996.

Bender:1995:CCT

[BTC^{+}95] William N. Bender, Pat Tilson, Randy Cooper, Stephen Barkley, Lynn Regan, and Lucy Levy. 21st century classroom teleconference, 1995. 1 videocassette (120 min.).

BMI:1994:FRI

[BU94a] Battelle Memorial Institute and United States. Federal Highway Administration. Final report for Internet and other advanced telecommunications options for interchange members. ???? #BAT-94-020, Battelle, U.S. Dept. of Transportation, Federal Highway Administration, 1994. 2 + 68 + A118 pp.

BVSD:1994:CCK

[BU94b] Boulder Valley School District, Annenberg/CPB Math and Science Project and U. S. West Foundation. Creating connections: K–12 teachers and the Internet. Technical report, Boulder Valley School District, Boulder, Col., 1994. various pp.

Bradley:1995:UQG

[BU95] Phil Bradley and UK Online User Group. UKOLUG quick guide to the Internet. UK Online User Group, London, UK, 1995. ISBN 1-870254-08-2 (paperback). 74 pp. LCCN ????

Buck:1995:EMP

[Buc95a] Tina Buck. Electronic mail (pine), 1995. 1 videocassette (80 min.).

Buck:1995:UIS

[Buc95b] Tina Buck. Using Internet search tools, 1995. 1 videocassette (56 min.).

Budnick:1996:WNW

[Bud96] Larry Budnick. The Windows NT Web server book: tools and techniques for building an Internet/Intranet site. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-342-5. ???? pp. LCCN TK5105.888.W56 1996.

Bupp:1995:IAC

[Bup95] Jonathan Bupp. The Internet as an agent of change for organizations. Thesis (Honors in Computer Science), Millersville University of Pennsylvania, Millersville, PA 17551, USA, 1995. 30 pp.

Burnett:1993:BI

[Bur93] Ingrid Schierling Burnett.

Byteline: The Internet. Colorado libraries, 19(2):38–??, Summer 1993. ISSN 0147-9733.

Burish:1995:IIC

[Bur95a] Helen T. Burish. Integrating the Internet into the computer art curriculum. Thesis (M.A.), University of Wisconsin–Madison, Madison, WI, USA, 1995. ii + 69 pp.

Burnett:1995:UER

[Bur95b] Gary Dean Burnett. Urban education resources on the Internet. Digest; no. 106 ERIC/CUE digest; no. 106. EDO-UD-95-3, ERIC Clearinghouse on Urban Education, New York, NY, USA, 1995.

Burke:1996:LIW

[Bur96] John (John J.) Burke. Learning the Internet: a workbook for beginners. Neal-Schuman net-guide series. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-248-1. xi + 138 pp. LCCN TK5105.875.I57B87

Anonymous:1995:BGIb

[Bus95] BusinessLink: global Internet resources and databases, page various, 1995. ISSN 1077-9760. Marcus P. Zillman, Marco Island, FL, USA.

Butler:1994:HUI

[But94] Mark Butler. How to Use the Internet. Ziff-Davis Press, Emeryville, CA, USA, 1994. ISBN 1-56276-222-2. xi + 145 pp. LCCN TK5105.875 .I57 B874 1994. US$17.95.

Butler:1995:UGS

[But95a] Kristin Deann Butler. A uses and gratifications study of the Internet. Thesis (M.S.), Journalism and Mass Communications Dept., San Jose State University, San Jose, CA, USA, 1995. x + 151 pp.

Butler:1995:TYS

[But95b] Mark Butler. Tsen yang shih yung Internet. Tsai se tien nao tsung shu; 7. Pai tung tu shu ku fen yu hsien kung ssu, Tai-pei shih, Taiwan, ti 1 pan. edition, 1995. ISBN 957-802-020-1 (paperback). 140 pp. LCCN ????

Butler:1995:MI

[But95c] Michelle Lynn Butler. Marketing and the Internet. Thesis (M.A.), University of Texas at Austin, Austin, TX, USA, 1995. x + 123 pp.

Buyens:1996:BNS

[Buy96] Jim Buyens. Building net sites with Windows NT: an Internet services handbook. Addison-Wesley Developers Press, Reading, MA, USA, 1996. ISBN 0-201-47949-4. ???? pp. LCCN TK5105.888.B89 1996.

Buyya:1999:HPC

[Buy99] Rajkumar Buyya. High Performance Cluster Computing, Volume 2: Programming and Applications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-013785-5. ???? pp. LCCN QA76.88 .H489 1999. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130137855.html.

Bacon:1993:IWE

[BW93] Francis Bacon and Internet Wiretap. The Internet wiretap edition of The new Atlantis, 1993.

Bradshaw:1994:BIT

[BW94a] Jacqueline Bradshaw and Maureen Witney. Basic Internet tools: telnet, email, FTP, and gopher: a training workbook for library staff. Electronic Library Network, Open Learning Agency, Burnaby, BC, Canada, v. 1.3. edition, 1994. ISBN 1-55162-028-6 (spiral bound), 1-55162-030-8 (perfect bound). 90 + 40 pp. LCCN ????

Bradshaw:1994:IRT

[BW94b] Jacqueline Bradshaw and Maureen Witney. Internet research tools: a training workbook for library staff. Electronic Library Network, Open Learning Agency, Burnaby, BC, Canada, v. 1.1. edition, October 1994. ISBN 1-55162-007-3. 70 + 84 pp. LCCN ????

Breeding:1996:MOI

[Bw96] Marshall Breeding and Internet world. Mecklermedia’s official Internet World World Wide Web yellow pages. I D G Books Worldwide, Indianapolis, IN, USA, 1996 edition, 1996. ISBN 1-56884-344-5. xli + 862 pp. LCCN ????

Berman:1997:SDD

[BW97a] Jerry Berman and Daniel J. Weitzner. Societal dimensions: Directing policymaking beyond the Net’s metaphor. Communications of the ACM, 40(2):83– 84, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-2/
p83-berman/.

Black:1997:STA

[BW97b] Uyless D. Black and Sharleen Waters. Sonet & T1: Architecture for Digital Transport Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-447590-9. xix + 332 pp. LCCN TK5105.415.B58 1997. US$61.00. URL http:/

/www.phptr.com/ptrbooks/
ptr_0134475909.html.

Bradshaw:1995:BIT

[BWC95a] Jacqueline Bradshaw, Maureen Witney, and Sheila Comeau. Basic Internet tools: telnet, e-mail, FTP, gopher and the World Wide Web: a training workbook. Electronic Library Network, Open Learning Agency, Burnaby, BC, Canada, second edition, 1995. ISBN 1-55162-195-9. v + 136 pp. LCCN ????

Bradshaw:1995:IRT

[BWC95b] Jacqueline Bradshaw, Maureen Witney, and Sheila Comeau. Internet research tools: a training workbook. Electronic Library Network, Open Learning Agency, Burnaby, BC, Canada, second edition, 1995. ISBN 1-55162-196-7. v + 123 pp. LCCN ????

Benjamins:1999:TBP

[BWWF99] V. R. Benjamins, B. Wielinga, J. Wielemaker, and D. Fensel. Towards brokering problemsolving knowledge on the Internet. Lecture Notes in Computer Science, 1621:33– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boggs:1983:IB

[BX83] David Reeves Boggs and Xerox Corporation. Internet broadcasting. Technical Report CSL 83-3, Xerox Corporation, Palo Alto Research Center, Palo Alto, CA, USA, 1983. vi + 84 pp.

Byrnes:1995:OSI

[Byr95]  Peter-John Byrnes.  The

online scholarship initiative UVa faculty scholarship on the Internet, 1995.

Byrne:1999:MAM

[Byr99] Jeffrey Byrne. MCSE: Administering Microsoft SQL Server 7. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010795-6. xxi + 567 pp. LCCN QA76.3 .B88

                                        1998. URL http://www.
                                        phptr.com/bookseri/mcse.
                                        html.

Berretti:1995:IAA

[BZ95] Alberto Berretti and Vittorio Zambardino. Internet: avviso ai naviganti. Interventi 20. Donzelli, Roma, Italy, 1995. ISBN 88-7989-150-2. 109 pp. LCCN ????

Carey:1995:II

[CA95] Patrick Carey and Angela Ambrosia. The Internet: illustrated. International Thomson Pub. Co., Cambridge, MA, USA, 1995. ISBN 1-56527-528-4. xvi + 176 pp. LCCN ????

Caceres:1989:MWA

[Cac89] Ramon Caceres. Measurements of wide area Internet traffic. Report UCB/CSD 89-550, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, December 1989. 12 pp.

CACM:1994:IT

[CAC94] Internet technology. Special Issue of Communications of the ACM, vol. 37, no. 8, August 1994.

Cahn:1998:WAN

[Cah98]  Robert S. Cahn.  Wide

Area Network Design: Concepts and Tools for Optimization. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-458-8. xx + 441 pp. LCCN TK5105.87 .C34 1998. US$77.95.

Cailliau:1994:SPW

[Cai94] R. Cailliau, editor. Selected papers of the 1st World-Wide Web conference, 25–27 May 1994, Geneva, Switzerland, volume 27(2) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, November 1994. CODEN CNISE9. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ???? Published in Computer Networks and ISDN Systems, 27, No. 2.

Caines:1995:I

[Cai95] Richard Caines. The Internet. A Key Note market briefing. Key Note, Hampton, Middlesex England, 1995. ISBN 1-85765-499-4. iii + 20 pp. LCCN ????

CERFnet:1990:CUG

[Cal90] California Education and Research Federation Network, San Diego, CA, USA. CERFnet user’s guide, 1990. various pp.

Caldwell:1991:WIP

[Cal91]  L. G. Caldwell.  Work-

station to internet; problems, solutions, and challenges. Computer Networks and ISDN Systems, 23(1–3): 25–??, November 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

California-LCB:1994:GAC

[Cal94a] California. Legislative Counsel Bureau. A Guide for accessing California legislative information over Internet. Sacramento, CA, USA, 1994. 30 pp. Cover title. Running title: Public access guide. “01/94” — Cover verso. Free to California residents. This pamphlet tells how you can find California legislative information online. Explains what legislation info is available, what assistance is available and how the information is organized. The majority of the book is spent explaining the Internet, how to get access, how to use electronic mail and where to go for more detailed information. There’s also a simple glossary of legislative terms. This pamphlet is also available online, in PostScript format. Send e-mail: To: ftpmail@leginfo.public.ca.gov Body: connect leginfo.public.ca.gov get README public access guide ps quit.

CLCB:1994:GAC

[Cal94b] California. Legislative Counsel Bureau, Sacramento, CA, USA. A Guide for accessing California legislative information over Internet, 1994. 30 pp.

Cameron:1994:I

[Cam94a] Debra Cameron. The Internet. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1994. ISBN 1-56607-035-X. ???? pp. LCCN TK5105.875.I57 C36 1994. US$285.00.

Cameron:1994:IGB

[Cam94b]  Debra Cameron.  The

Internet: a global business opportunity. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1994. ISBN 1-56607-035-X. vi + 229 pp. LCCN TK5105.875.I57C36 1994.

Cameron:1995:IIB

[Cam95a] Debra Cameron. Implementing the Internet for business. Computer Technology Research Corporation, Charleston, SC, USA, 1995. ISBN 1-56607-952-7. xii + 301 pp. LCCN HD30.335.C35 1995.

Campbell:1995:UVI

[Cam95b] James Campbell. The University of Virginia Internet library, 1995.

Campidonica:1995:GAI

[Cam95c] Mark Campidonica. A guide to agriculture on Internet. Technical report, agAccess, Davis, CA, USA, 1995. ii + 101 pp.

Cameron:1996:SII

[Cam96] Debra Cameron. Security issues for the Internet and the World Wide Web. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1996. ISBN 1-56607-973-X. ???? pp. LCCN QA76.9.A25C363 1996.

Cam:1997:LIP

[Cam97] H. Cam. LAN-ATM Internetworking over a prioritybased slotted-ring network. In IEEE Computer Society. Technical Committee on Computer Communications [IEE97], pages 54–63. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

CISTI:1994:IFG

[Can94] Canada Institute for Scientific and Technical Information. The Internet in the federal government: report of a forum organized by the Canada Institute for Scientific and Technical Information, March 16, 1994, Ottawa. Technical Report NRCC 38274, Canada Institute for Scientific and Technical Information (CISTI), National Research Council Canada, Ottawa, ON, Canada, March 16, 1994. 19 pp.

CP:1995:PIP

[Can95] Canada. Parliament. Parliamentary Internet parliamentaire, 1995.

CMU:1993:IKE

[Car93a] Carnegie-Mellon University.

Internet in K–12 Education Project and Team. The Internet in K–12 education. Carnegie Mellon University, Internet in K–12 Education Project Team, Pittsburgh, PA, USA, December 2, 1993. vi + 165 pp.

Carroll:1993:WW

[Car93b] J. A. Carroll. Wired to the world. Ca magazine, 126(7): 28–??, August 1, 1993.

Carballar:1994:IMT

[Car94a]  Jose A. Carballar.  In-

ternet. El mundo en tus manos. Ra-Ma, Libreria y Editorial Microinformatica, Madrid, Spain, November 1994. ISBN 84-7897-158-0. 385 pp. 3500 pesetas, US$27.00.

CarballarFalcon:1994:IMT

[Car94b] Jose Antonio Carballar Falcon. Internet: el mundo en tus manos. Ra-Ma, Libreria y Editorial Microinformatica, Madrid, Spain, 1994. ISBN 84-7897-158-0. XVIII + 372 pp. LCCN TK5105.875.I57C368 1995.

Carp:1995:CIR

[Car95a] Robert H. Carp. The Complete Internet reference for lawyers. Internet Research, Inc., Chestnut Hill, MA, USA, 1995. 228 pp.

Carroll:1995:GIC

[Car95b] Jim Carroll. Get on the Internet in Canada in 5 minutes. Prentice-Hall Canada, 1870 Birchmont Road, Scarborough, ON, Canada M1P 2J7, 1995. ISBN 0-13-371204-4. ???? pp. LCCN ????

Carton:1995:IVW

[Car95c] Sean Carton. Internet virtual worlds quick tour: MUD’s, MOO’s and MUSHes: interactive games, conferences and forums. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-222-4. xxii + 192 pp. LCCN GV1469.15 .C37 1995. US$14.00.

CEI:1996:LGJ

[Car96] Career Education Institutes, Winston-Salem, NC, USA. The Lawyers guide to jobsurfing on the Internet, 1996. 27 pp.

Carlson:1998:PDD

[Car98a] James D. Carlson. PPP Design and Debugging. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-18539-3. xi + 228 pp. LCCN TK5105.582C37 1998. US$26.95.

Carne:1998:TTA

[Car98b] E. Bryan Carne. Telecommunications Topics: Applications of Functions and Probabilities in Electronic Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-645565-4. xxii + 417 pp. LCCN TK5102.9.C35

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0136455654.html.

Carne:1999:TPD

[Car99a] E. Bryan Carne. Telecommunications Primer: Data, Voice, and Video Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-022155-4. 832 pp. LCCN TK5101.C29863

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130221554.html.

Carpenter:1999:WBC

[Car99b] Jeffrey J. Carpenter. Welcome to the Big City: Incident reporting helps the CERT Coordination Center keep pace with a rapidly expanding Internet. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/1999-11/features/

cert.html. Special issue on security.

Cashin:1993:NUC

[Cas93]  Jerry Cashin.  Network-

ing with UNIX—Connecting Open Systems. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1993. ISBN 1-56607-003-1. iv + 171 pp. LCCN TK5105.5 .C37 1992. US$220.00.

Cassidy:1994:SBG

[Cas94] Kyle Cassidy. SATURN: a Beginner’s Guide To Using the Internet. RCNJ Press, ????, 1994. ISBN ?? 110 pp. This is the only Internet book that I know of that is specifically for VAX/VMS users (you poor things.) The manual is somewhat specific for our VAX 4000, but it can be used by anyone who is netting with VMS. It is also available online for free, FT-

Pable from: gboro.rowan.
edu:/pub/Saturn_Guide.

Catlett:1989:NBN

[Cat89] Charles E. Catlett. The NFSNET: Beginnings of a National Research Internet. Academic computing (McKinney, TX), 3(5):18–??, January 1, 1989. ISSN 0892-4694.

Catapult:1996:MIE

[Cat96] Catapult, Inc. Microsoft Internet Explorer step by step. Microsoft Press, Bellevue, WA, USA, 1996. ISBN 1-57231-300-5. ???? pp. LCCN TK5105.875.I57M53 1996.

Caumont:1995:IE

[Cau95] Marc Caumont. Internet Erotisme. Carnet d’adresses. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1576-7. xiii + 157 pp. LCCN ????

Cawley:1995:MML

[Caw95] Piers Cawley. Majordomo: a mailing list handler. Linux Journal, 13:??, May 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue13/majordomo.
html.

Caywood:1993:TL

[Cay93] Carolyn Caywood. Teens & libraries. School library journal: SLJ, 39(10):58–??, October 1, 1993. ISSN 0362-8930.

Crowcroft:1991:TSI

[CB91] J. Crowcroft and R. Bradshaw. TCP/IP Internet protocols and JANET. Computer Networks and ISDN Systems, 23(1–3):177– ??, November 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Carroll:1994:CIHa

[CB94a] Jim Carroll and Rick Broadhead. Canadian Internet Handbook. Prentice-Hall Canada, 1870 Birchmont Road, Scarborough, ON, Canada M1P 2J7, 1994 edition, 1994. ISBN 0-13-304395-9. xvii + 415 pp. LCCN Prentice-Hall Canada. US$14.95. Highly recommended (for Canadian readers). With pretty complete info on Internet providers and resources in Canada ….

Carroll:1994:CIHb

[CB94b] Jim Carroll and Rick Broadhead. Canadian Internet Handbook 1995 edition. Prentice-Hall Canada, 1870 Birchmont Road, Scarborough, ON, Canada M1P 2J7, second edition, November 1994. ISBN 0-13-329350-5. xx + 798 pp. LCCN ????

US$21.95. URL http://

www.csi.nb.ca/handbook/
handbook.html. Notes from

1st ed: If you live in Canada get this book. It contains sections about getting Internet access in Canada, growth of ’net use there, short basic sections about how to use of some of the most popular Internet tools, a huge directory of Canadian Internet service providers, and even more huge list of gopher servers and campus-wide information systems in Canada and to top it off, lists of Canadianbased Usenet groups, WWW, Archie, IRC servers and online catalogs. Notes from the second edition: really, really get this book!

Cheswick:1994:FIS

[CB94c] William R. Cheswick and Steven M. Bellovin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63357-4. xiv + 306 pp. LCCN TK5105.875.I57C44 1994. US$24.95.

Campbell:1995:RTO

[CB95a] Darwin Campbell and Catherine Bradshaw. Riderlink: transportation options on the Internet. Technical report, ????, ????, 1995. 10 + 2 pp.

Carp:1995:ISP

[CB95b] Robert H. Carp and Sandy Bendremer. Internet service provider start up manual. Internet Research, Inc., Chesnut Hill, MA, USA, 1995. 247 pp.

Carroll:1995:CIH

[CB95c] Jim Carroll and Rick Broadhead. Canadian Internet handbook. Prentice-Hall Canada, 1870 Birchmont Road, Scarborough, ON, Canada M1P 2J7, revised 1995 edition, 1995. ISBN 0-13-373614-8. xxiii + 903 pp. LCCN ????

Carroll:1995:IAO

[CB95d] Jim Carroll and Rick Broadhead. The Internet advantage: opportunities for Canadian business. Prentice-Hall Canada, 1870 Birchmont Road, Scarborough, ON, Canada M1P 2J7, 1995. ISBN 0-13-226598-2. xx + 522 pp. LCCN HD30.335 .C38 1995. US$29.95.

Carton:1996:IPT

[CB96a] Sean Carton and Gareth Branwyn. Internet power toolkit: cutting-edge tools and techniques for power users. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-329-8. ???? pp. LCCN TK5105.875.I57C37 1996.

Crespo:1996:WBB

[CB96b] Arturo Crespo and Eric A. Bier. WebWriter: a browserbased editor for constructing Web applications. Computer Networks and ISDN Systems, 28(7–11):1291–1306, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1616.

Cerf:1983:DIA

[CC83] Vinton G. Cerf and Edward A. Cain. The DoD Internet architecture model. Computer Networks: The International Journal of Distributed Informatique, 7(5): 307–318, October 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

CLASS:1993:CII

[CC93] Cooperative Library Agency for Systems and Services (Calif.) and California Internet Institute. California Internet Institute: reference notebook. CLASS, San Jose, CA, USA, 1993. various pp.

Campbell:1995:SGD

[CC95] David R. Campbell and Mary V. Campbell. The Student’s Guide to Doing Research on the Internet. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-48916-3 (paperback). ix + 349 pp. LCCN LB2369 .C28 1995. US$14.95.

Chu:1996:HPI

[CC96a] Kenny Chu and Francis Chin. HTML publishing on the Internet. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912100-4 (paperback). xx + 444 pp. LCCN QA76.76.H94 C58 1996.

Coombs:1996:SIS

[CC96b] Jason Coombs and Ted Coombs. Setting up an Internet site for dummies. – for dummies. IDG Books, San Mateo, CA, USA, 1996. ISBN 1-56884-335-6. xxii + 362 pp. LCCN ????

Clark:1991:TFI

[CCC^{+}91] D. Clark, L. Chapin, V. Cerf, R. Braden, and R. Hobby. Towards the future Internet architecture, December 1991. 1 data file 5 1/4 in. Available via FTP.

CCNY:1994:IGD

[CCN94] CCNY. Internet Gopher documentation. New York, NY, USA, 1994. 10 pp.

Chen:1992:IJM

[cCrmdh92] Chia chun Ch’en and Internet ru men dao hang. Internet ju men tao hang / Ch’en Chia-chun pien chu. Tien nao yu t’ung hsun ts’ung shu; 5. Tzu hsun yu tien nao ch’u pan she, T’ai-pei shih, Taiwan, ch’u pan. edition, 1992. ISBN 957-9076-46-4 (paperback). 303 pp. LCCN ????

Casner:1992:FII

[CD92] Stephen H. Casner and Stephen E. Deering. First IETF Internet audiocast. ISI reprint series ISI/RS-92-293, University of Southern California, Information Sciences Institute, Marina del Rey, CA, USA, July 1992. 6 pp.

Chicoine:1994:ISS

[CD94] Pierre Chicoine and Pierre Duchesneau. Internet et les sciences de la santé: guide d’introduction aux ressources. Hôpital Saint-Luc, Bibliothèque, Montréal, PQ, Canada, 1994. ISBN 2-9801006-2-5. ii + 77 pp. LCCN ????

Caumont:1995:PRI

[CD95] Marc Caumont and Marc Dandreau. Pages roses Internet erotique. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1669-0. 64 pp. LCCN ????

Comer:1999:CNI

[CD99] Douglas E. Comer and Ralph Droms. Computer Networks and Internets. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-083617-6. xxiii + 583 pp. LCCN TK5105.5.C5897

1999. URL http://www.
phptr.com/bookseri/comer.
html;   http://www.prenhall.
com/allbooks/esm_0130836176.

html. CD-ROM and Website by Ralph Droms.

Chapman:1994:SMS

[CDEF94] Sara S. Chapman, Helen Desfosses, Beth Engeler, and Lee Farbman. Still missing; Shannon Faulkner; making friends through the Internet, 1994. 1 sound cassette (25 min.).

Coulouris:1994:DSC

[CDK94] George Coulouris, Jean Dollimore, and Tim Kindberg. Distributed Systems: Concepts and Design. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-62433-8. xix + 644 pp. LCCN QA76.9.D5 C68 1994. US$49.50.

Cedeno:1995:ITK

[Ced95] Nancy Cedeno. The Internet tool kit. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1688-4 (paperback). xxvi + 418 pp. LCCN TK5105.875.I57C42 1995. US$16.99.

SNIC:1993:SGS

[Cen93] SURAnet Network Information Center. SURAnet guide to selected Internet resources. Technical report, SURAnet Network Information Center, College Park, MD, USA, 1993. various pp.

CCGS:1995:INH

[Cen95] International network on Holocaust and genocide, page various, 1995. Macquarie University, NSW, Australia.

Cerf:1989:IAB

[Cer89] V. Cerf. The Internet Activities Board. Computer Networks and ISDN Systems, 17(4–5):337–343, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cerf:1991:GIM

[Cer91] Vinton G. Cerf. Guidelines for Internet measurement activities, October 1991. 1 data file 5 1/4 in. Available via FTP.

Cerf:1993:WHI

[Cer93] Vinton G. Cerf. Weekend headliner Internet program, 1993. 1 sound cassette (ca. 30 min.).

Crawford:1999:FRB

[CES^{+}99] Diane Crawford, John Ehrt, Joe Saltiel, Robert Glass, John Wubbel, John Willmore, and Tom Moran. Forum: a righteous balance of Internet freedom; don’t compare business majors to hackers; “popular” press; empirical studies vs. armchair argument. Communications of the ACM, 42(4):13–16, April 1999.

URL http://www.acm.org:
80/pubs/citations/journals/
cacm/1999-42-4/p13-crawford/

.

Ciciora:1999:MCT

[CFL99] Walter Ciciora, James Farmer, and David Large. Modern Cable Television Technology: Video, Voice, and Data Communications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-416-2. xxxii + 873 pp. LCCN TK6675 .C53 1999. US$84.95.

Carter:1995:PIN

[CG95] Michael Carter and Graeme Guthrie. Pricing Internet: the New Zealand experience. Discussion paper 9501, Dept. of Economics and Operations Research, University of Canterbury, Christchurch, New Zealand, January 1995. 17 pp.

Huang:1995:IHY

[cH95] Yen chung Huang. InterNet hsiao yao yu: wang chi wang lu ju men chih nan. Po shuo ku wen “ching tien hsi lieh”. Po shuo ku wen yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-8948-45-X. 508 pp. LCCN ????

Chandler:1992:IPI

[Cha92] Connee Chandler. INTER-NET: a practical introduction, 1992. 1 videocassette (3 hr. 45 min.).

Chamberlain:1995:CWT

[Cha95a]  Arthur Chamberlain.  A

census of WWW tourism home pages on the Internet: evaluating the quality of design and communications to tourists. Thesis (M.A.), University of Florida, Gainesville, FL, USA, 1995. vi + 160 pp.

Chamorro:1995:RRI

[Cha95b] Rafael Chamorro. La red de redes Internet, 1995.

Chandler:1995:RPW

[Cha95c] David M. Chandler. Running a Perfect Web Site. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0210-X. xviii + 457 pp. LCCN ???? US$39.95. Includes CD ROM.

Cherenson:1987:SIP

[Che87] Andrew Richard Cherenson.

The Sprite Internet protocol server: research project. Master of science, plan ii, University of California, Berkeley. Department of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1987. 29 pp.

Cheifet:1993:I

[Che93a] Stewart Cheifet. Internet, 1993. 1 videocassette (28 min.).

Cheifet:1993:IU

[Che93b] Stewart Cheifet. Internet update, 1993. 1 videocassette (27 min.).

Cheifet:1994:I

[Che94] Stewart Cheifet. Internet, 1994. 1 videocassette (26 min.).

Cheifet:1995:I

[Che95] Stewart Cheifet. Internet, 1995. 1 videocassette (28 min.).

Cheong:1996:IAS

[Che96] Fah-Chun Cheong. Internet agents: spiders, wanderers, brokers, and ’bots. New Riders, Indianapolis, Ind., 1996. ISBN 1-56205-463-5. xiv + 413 pp. LCCN TK5105.875.I57 C43 1996. US$32.00, CDN$43.95.

Chenoweth:1997:WDC

[Che97] J. D. Chenoweth. What do computer science students need to know about HTML and the Internet? In Meinke [Mei97], pages 10–14. ISBN ???? LCCN ????

Cargille:1991:IXP

[CHHL91] C. A. Cargille, R. A. Hagens, A. Hansen, and L. H. Landweber. The Internet X.400 pilot project. Computer Networks and ISDN Systems, 23(1–3):97– ??, November 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

CCNICT:1993:MB

[Chi93] Chile. Comision Nacional de Investigacion Cientifica y. and Tecnologica. Departamento de Informacion and Manual Batch Internet to NLM Intercommunication System. Manual BITNIS. El Departamento, Santiago de Chile, Chile, 1993. 55 pp.

Chiang:1994:IML

[Chi94] Dudee Chiang. Internet for medical librarians: a syllabus. Technical report, Norris Medical Library, University of Southern California, Los Angeles, CA, USA, 1994. 83 pp.

Chiu:1995:UMI

[Chi95] Wanda Chiu. Using Mosaic and the Internet: an interactive guide for junior high students. Technical report TR 95-06, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, July 7, 1995. 74 pp.

Cho:1995:LRO

[Cho95] Gi Hwan Cho. Location and routing optimization protocols supporting Internet host mobility. Thesis (Ph.D.), Dept. of Computing Science, University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1995. ix + 148 pp.

Chon:1996:II

[Cho96] Kilnam Chon. Internet inroads. Communications of the ACM, 39(6):59–60, June 1996. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228515.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228515.html.

Carr:1996:OIS

[CHR^{+}96] Leslie Carr, Gary Hill, David De Roure, Wendy Hall, and Hugh Davis. Open information services. Computer Networks and ISDN Systems, 28(7–11):1027–1036, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1603.

Christenson:1998:BRL

[Chr98]  Nick Christenson.  Book

review: At Large: The Strange Case of the World’s Biggest Internet Invasion. ;login: the USENIX Association newsletter, 23(4): ??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
1998-6/freedman.html.

Churbuck:1993:EF

[Chu93] David C. Churbuck. On the electronic frontier. Forbes, 152(12):170–??, November 22, 1993. CODEN FORBA5. ISSN 0015-6914.

Crowcroft:1999:IM

[CHW99] Jon Crowcroft, Mark Handley, and Ian Wakeman. Internetworking Multimedia. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-584-3. xx + 290 pp. LCCN TK5105.5 .C78 1999. US$59.95.

Cimino:1996:BSE

[Cim96] J. J. Cimino, editor. Beyond the superhighway: exploiting the Internet with medical informatics: proceedings, 1996 AMIA annual Fall Symposium, formerly SCAMC, October 26–30, 1996, Sheraton Washington Hotel, Washington, DC. Hanley and Belfus, Philadelphia, PA, USA, 1996. ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Anonymous:1996:CAR

[Cin96] Cinfolink annual review of information services and the Internet in China, page various, 1996. Cinfolink Services, Toronto, ON, Canada.

Cerf:1974:PPN

[CK74] Vint Cerf and Robert Kahn.

A protocol for packet network intercommunication. IEEE Transactions on Communications, 22(5):637–648, May 1974. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Campbell:1999:OSA

[CKMV99] Andrew T. Campbell, Irene Katzela, Kazuho Miki, and John Vicente. Open Signaling for ATM, INTERNET and Mobile Networks (OPEN-SIG’98). Operating Systems Review, 33(2):15–28, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cabell:1999:UTP

[CKPT99] Q. Cabell, P. Kirstein, T. Pagtzis, and T. Tiropanis. The use of TINA principles in the management of Internet multimedia conferences. Lecture Notes in Computer Science, 1597:482–498, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:PAC

[CKTV96] Paolo Ciancarini, Andreas Knoche, Robert Tolksdorf, and Fabio Vitali. PageSpace: an architecture to coordinate distributed applications on the Web. Computer Networks and ISDN Systems, 28(7–11):941–952, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1598.

Crutchfield:1994:IPA

[CKY94] Dean Crutchfield, Ed Krol, and Scott Yanoff. Internet: a practical approach, 1994. 1 videocassette (180 min.). Video tape of a satellite seminar held May 5, 1994, and recorded with permission from PBS Adult Learning Satellite Service. Performers: Host, Dean Crutchfield; panelists, Ed Krol, Scott Yanoff. Internet services are addressed in this seminar, including demonstrations, aimed at both new and experienced users of the emerging superinformation highway. Goal is to assist academicians in utilizing Internet network to its fullest potential.

Chen:1995:ASS

[CL95] Thomas M. Chen and Stephen S.

Liu. ATM Switching Systems. Artech House Inc., Boston, MA, USA, 1995. ISBN 0-89006-682-5. xvii + 261 pp. LCCN TK5105.35 .C47 1995. US$65.00.

Clark:1998:SIC

[CL98] Jan Guynes Clark and Vincent S. Lai. On site: Internet comes to Morocco. Communications of the ACM, 41(2): 21–26, February 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/

                                        journals/cacm/1998-41-2/
                                        p21-guynes_clark/.

Claffy:1994:ITC

[Cla94] Kimberly C. Claffy. Internet traffic characterization. Thesis (Ph.D.), University of California, San Diego, Department of Computer Science and Engineering, San Diego, CA, USA, 1994. xv + 158 pp. Vita. Thesis (Ph. D.).

Clark:1995:RIC

[Cla95a] Charles S. Clark. Regulating the Internet: can the use of cyberspace be governed? Congressional quarterly researcher, 5(24):561– 584, June 30, 1995. ISSN 1056-2036.

Clark:1995:IE

[Cla95b] David Clark. Internet essentials. Que College, Indianapolis, IN, USA, second edition, 1995. ISBN 0-7897-0246-0 (paperback). 6 + 174 pp. LCCN TK5105.875.I57C548 1995.

Clark:1995:SGI

[Cla95c] David Clark. Student’s guide to the Internet. Alpha Books, Indianapolis, IN, USA, 1995. ISBN 1-56761-545-7. xx + 314 pp. LCCN TK 5105.875 I57 C55 1995. US$14.99. Covers IBM and compatibles, Macintosh and UNIX.

Clark:1995:CTI

[Cla95d] Michael Clark. Cultural treasures of the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-209669-2. xxix + 313 pp. LCCN QA76.55 .C58 1995 Ref.

Clark:1996:SGI

[Cla96] Carol Lea Clark. A student’s guide to the Internet. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-442310-0. vii + 184 pp. LCCN TK5105.875 .I57 C53 1996.

Clarke:1999:IPC

[Cla99] Roger Clarke. Internet privacy concerns confirm the case for intervention. Communications of the ACM, 42(2):60–67, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org:80/pubs/citations/
journals/cacm/1999-42-2/
p60-clarke/.

Cleveland:1993:DIA

[Cle93a] Gary Cleveland. Dial-up Internet access providers in Canada. Technical report, National Library of Canada, Ottawa, ON, Canada, 1993. 7 pp.

Cleveland:1993:FSI

[Cle93b] Gary Cleveland. Fournisseurs de services Internet par ligne commutée au Canada. Technical report, Bibliothèque nationale du Canada, Ottawa, ON, Canada, 1993. 6 pp.

Clockedile:1993:MSI

[Clo93] Sandra L. Clockedile. Making sense of the Internet. Technical report, Graduate School of Library and Information Science, Simmons College, Boston, MA, USA, 1993.

Clyne:1988:LWI

[Cly88] Les Clyne. LAN/WAN internetworking. Computer Networks and ISDN Systems, 16 (1–2):34–39, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Clyde:1993:IIL

[Cly93] Laurel Clyde. The Internet: an introduction for librarians: workshop manual. Technical report, Lindin, Reykjavik, Iceland, 1993. 38 + 10 pp.

Cady:1995:MI

[CM95] Glee Harrah Cady and Pat McGregor. Mastering the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1645-0 (paperback). xliv + 1258 pp. LCCN TK5105.875.I57 C34 1995. US$39.95. System requirements for accompanying disks: IBM PC or compatible; Microsoft Windows.

Carl-Mitchell:1993:PIT

[CMQ93] Smoot Carl-Mitchell and John S. Quarterman. Practical Internetworking with TCP/IP and UNIX. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-58629-0. xviii + 476 pp. LCCN TK5105.5 .C36 1993. US$43.25.

Co:1995:PMU

[CN95a] Van Kiem Co and Oanh Hoang Nguyen. Purpose, methods and user guide of the Internet telecommunication system. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1995. i + 23 pp.

CNC:1995:CNI

[CN95b] CommerceNet Consortium and Nielsen Media Research. CommerceNet/ Nielsen: Internet demographic study. Technical report, CommerceNet Consortium/Nielsen Media Research, New York, NY, USA, November 1995.

Cheifet:1994:IU

[CNO94] Stewart Cheifet, Peter Nichols, and Clementine Oliver. Internet (update), 1994. 1 videocassette (27 min.).

Comer:1987:CCB

[CNY87] Douglas Comer, Thomas Narten, and Rajendra Yavatkar. Cypress coaxial backbone packet switch. Computer Networks and ISDN Systems, 14(2–5):383–388, 1987. CODEN CNISE9.

ISSN 0169-7552 (print), 1879-2324 (electronic).

Cleary:1999:VGV

[CO99] D. Cleary and D. O’Donoghue. VisualExpresso: Generating a virtual reality internet. Lecture Notes in Computer Science, 1593:797–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CobbGroup:1994:II

[Cob94] Inside the Internet, page various, 1994. ISSN 1075-7902. Cobb Group, Louisville, KY, USA.

CobbGroup:1995:IBA

[Cob95] Internet business advantage, page various, 1995. ISSN 1082-393X. Cobb Group, Louisville, KY, USA.

CobbGroup:1996:ENI

[Cob96] Explore the net with Internet explorer/MSN, page various, 1996. ISSN 1087-2965. Cobb Group, Louisville, KY, USA.

Cochrane:1995:I

[Coc95] Kerry Cochrane. The Internet. A first book. Franklin Watts, New York, NY, USA, 1995. ISBN 0-531-20200-3. 63 pp. LCCN TK5105.875.I57C62 1995.

Cogswell:1994:SI

[Cog94] Jeffrey M. Cogswell. Simple Internet. Waite Group Press, Corte Madera, CA, USA, 1994. ISBN 1-878739-79-4. xiii + 162 pp. LCCN TK5105.875.I57 C63 1994.

Cogswell:1995:SI

[Cog95] Jeffrey M. Cogswell. Simple Internet. ????, Madrid, Spain, 1a edition, 1995. ISBN 84-7614-740-6 (paperback). 191 pp. LCCN ????

Cohen:1999:MTN

[Coh99] Robert B. Cohen. Moving toward a non-U.S.-centric international Internet. Communications of the ACM, 42 (6):37–40, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p37-cohen/.

CU:1993:ITP

[Col93] Columbia University. Libraries and Internet training. Internet training program. Technical report, Columbia University, New York, NY, USA, 1993. various pp.

CDLRC:1995:SEI

[Col95] College of DuPage. Learning Resource Center. Soaring to excellence II, 1995. 1 videocassette (75 min.) guide.

Cole:1999:ENC

[Col99] Bernard Cole. The Emergence of Net-Centric Computing: Network Computers, Internet Appliances, and Connected PC’s. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-897869-7. xviii + 362 pp. LCCN QA76.527.C65

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0138978697.html.

CompCommRev:1971:CCR

[Com71] Computer Communication Review—The SIGCOMM Quarterly Publication, 1971. CODEN CCRED2. ISSN 0146-4833. US$37. ACM Press,New York, NY 10036, USA. Quarterly.

CommWeek:1984:CWN

[Com84] Communications Week—The Newspaper For Network Decision Makers, 1984. ISSN 0746-8121. P.O. Box 2070, Manhasset, NY 11030, USA. Weekly; available free to qualified subscribers.

Comer:1987:OSD

[Com87a] Douglas E. Comer. Operating Systems Design. Internetworking with XINU, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1987. ISBN 0-13-637414-X. xxi + 567 pp. LCCN QA76.76.O63 C65 1984.

Comer:OSD87

[Com87b] Douglas E. Comer. Operating Systems Design. Internetworking with XINU, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1987. ISBN 0-13-637414-X. LCCN QA76.6 .C6275 1984.

Comer:1988:ITP

[Com88a] Douglas E. Comer. Internetworking with TCP/IP. Principles, Protocols, and Architecture, volume 1. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-470154-2. xvii + 382 pp. LCCN TK5105.5 .C59 1988. US$27.00.

Comer:ITP88

[Com88b] Douglas E. Comer. Internetworking with TCP/IP. Principles, Protocols, and Architecture. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-470154-2. LCCN TK5105.5 .C59 1988. US$27.00.

Comer:1991:ITDb

[Com91a] Douglas E. Comer. Internetworking with TCP/IP. Design, Implementation and Internals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1991. ISBN 0-13-465378-5. xvi + 524 pp. LCCN ????

Comer:1991:ITP

[Com91b] Douglas E. Comer. Internetworking with TCP/IP. Principles, Protocols, and Architecture. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1991. ISBN 0-13-474321-0, 0-13-468505-9. 534 pp. LCCN TK5105.585 .C66 1991.

Comer:1992:ITC

[Com92]  Douglas E. Comer.  In-

ternetworking with TCP/IP. Client-server programming and applications: BSD Sockets Version. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-474222-2. LCCN TK5105.5 .C59 1991. US$51.00. See [CS93a].

CC:1993:I

[Com93] Computer chronicles (Television program). Internet, 1993. 1 videocassette (27 min.).

Comer:1994:ITC

[Com94a]  Douglas E. Comer.  In-

ternetworking with TCP/IP. Client-server programming and applications: AT&T SVR4 TLI Version. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-474230-3. xx + 508 pp. LCCN TK5105.5 .C59 1994. US$50.00.

Aldea:1994:N

[Com94b] Aldea Communications. Net-Pages. Aldea Communications, P.O. Box 188125, Carlsbad, CA 92009, USA, 1994. ISBN ???? 200 pp. LCCN ???? Of the various Internet White and Yellow pages books out there, this one looks and acts the most like a ‘real’ telephone book: personal listings of your name and e-mail address are free, advertising space in the yellow pages section is available for a fee, and the book itself it distributed for free. With only a few thousand names in the first edition, this listing is hardly the most complete, but it only contains information about people who choose to be listed.

CompuTate:1994:IN

[Com94c] The Internet novice, page various, 1994. US$17.50. CompuTate, Arlington, VA, USA.

Comer:1995:IBE

[Com95a] Douglas Comer. The Internet book: everything you need to know about computer networking and how the Internet works. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-183237-9 (paperback), 0-13-151565-9 (recycled paper). xxii + 312 pp. LCCN TK5105.875.I57 C65 1995.

Comer:1995:SGI

[Com95b] Douglas Comer. Study guide for the Internet book. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-188012-8. iii + 63 pp. LCCN ????

Comer:1995:ITI

[Com95c] Douglas E. Comer. Internetworking with TCP/IP: Principles, Protocols and Architecture. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1995. ISBN 0-13-216987-8. 613 pp. LCCN TK5105.5 .C59 1994. US$52.00.

Compton:1995:ICN

[Com95d] Charles L. (Charles Lacey) Compton. Internet CNN newsroom: the design of a digital video news magazine. Thesis (M. Eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1995. 57 pp.

Comer:1997:IBE

[Com97] Douglas E. Comer. The Internet Book: Everything You Need to Know About Computer Networking and How the Internet Works. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-890161-9. xxiv + 327 pp. LCCN TK5105.875.I57C65 1997.

URL http://www.phptr.
com/bookseri/comer.html;
 http://www.prenhall.com/
allbooks/esm_0138901619.
html.

Connexions:1987:CIR

[Con87] Connexions—The Interoperability Report, 1987. CO-DEN CONNFR. ISSN 0894-5926. 480 San Antonio Road, Mountain View, CA 94040-1219, USA, Tel: (415) 941-3399, E-mail: connexions@ interop.com. U.S./Canada subscriptions US$150 for 12 issues/year.

Conroy:1989:IDR

[Con89]  Patrick T. Conroy.  An

Internet datagram router for a virtual circuit network. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1989. v + 24 pp.

CC:1994:HGC

[Con94] Classroom Connect. How to get connected to the Internet: facts and funding. Technical report, Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1994. 14 pp.

CC:1995:EOR

[Con95a] Classroom Connect. The educator’s online resource guide: where to find the information you need on the Internet. Technical report, Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1995. 11 pp.

CC:1995:HGC

[Con95b] Classroom Connect. How to get connected to the Internet: facts and funding. Technical report, Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1995. 14 pp.

CNC:1992:CRI

[Coo92] The Cook report on Internet, NREN, page various, 1992. ISSN 1071-6327. Cook Network Consultants, Ewing, NJ, USA.

Cook:1993:ONP

[Coo93a]  Gordon Cook.  Opinion:

NSFnet privatization: Policy making in a public interest vacuum. Internet research, 3 (1):3–??, Spring 1993. CO-DEN IRESEF. ISSN 1066-2243.

CND:1993:CNH

[Coo93b] Cook Network Consultants. Computer networks and health care: a survey of four growth areas. The Cook report on Internet — NREN, Cook Network Consultants, Ewing, NJ, USA, October 1993. 24 pp.

CNC:1995:IIT

[Coo95a] Cook Network Consultants. Information infrastructure: top down or bottom up? Technical report, Cook Network Consultants, Ewing, NJ, USA, 1995. 80 pp.

Cooper:1995:IIS

[Coo95b] Frederic J. Cooper. Implementing Internet security. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-471-6. xv + 378 pp. LCCN TK5105.59.I44 1995.

Cook:1996:CVC

[Coo96] Zachary Cook. Cybersmut, virtual communities, and the marketplace of ideas: a look at the legal regulation of the Internet. Thesis (B.A.), Dept. of Political Science, Williams College, Williamstown, MA, USA, 1996. 104 pp.

Anonymous:1995:CIS

[Cor95] Corporate Internet strategies, page various, 1995. ISSN 1087-6251. Cutter Information Corp., Arlington, MA, USA.

Corrigan:1996:IUC

[Cor96] Dan Corrigan. The Internet university: college courses by computer. Cape Software, Harwich, MA, USA, 1996. ISBN 0-9648112-0-0 (paperback). xx + 488 pp. LCCN LC5951.U6C67 1996.

Costa:1994:PDH

[Cos94a] Janis Furtek Costa. Planning and Designing High-Speed Networks Using 100VG-AnyLAN. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-168685-2. viii + 121 pp. LCCN TK5105.7 .C689 1994. US$24.00.

Costales:1994:S

[Cos94b] Bryan Costales. sendmail.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1994. ISBN 1-56592-056-2. 830 pp. US$32.95. While not strictly an Internet book this tome focuses on one thing: the UNIX program sendmail, which is a huge part of how electronic mail moves around on the Internet. Mainly for system administrators, the book shows how to use every function, mode and mood of sendmail to get your e-mail where it’s going. A great, if single-minded, book.

Cotton:1995:OCT

[Cot95] Eileen Giuffre Cotton. The on-line classroom: teaching with the Internet. ERIC Clearinghouse on Reading, English, and Communication, Bloomington, IN, USA, 1995. ISBN 1-883790-18-2 (paperback). ???? pp. LCCN LB1044.87.C68 1995.

Cremonini:1999:MAS

[COZ99] M. Cremonini, A. Omicini, and F. Zambonelli. Multiagent systems on the Internet. extending the scope of coordination towards security and topology. Lecture Notes in Computer Science, 1647:77–88, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crocker:1982:IMP

[CP82] David H. Crocker and Jonathan Bruce Postel. Internet mail protocols. Network Information Center, SRI International, Menlo Park, CA, USA, 1982. 125 pp.

Chen:1994:IHH

[CP94a] Min Chen and William Popper. Internet Hacker’s Handbook: Twenty Tips for Secure Data. ????, ????, 1994. ISBN 0-9625476-1-1. 129 pp. LCCN ???? US$11.95.

Condat:1994:I

[CP94b] Jean-Bernard Condat and Nicolas Pioch. Internet. Collection Informatique Collection Informatique (Montréal, Quebec). Editions J.C.I., 2700 Joliette #201, Montréal, Québec, Canada H1W 3G9, October 1994. ISBN 2-921599-06-6. 214 pp. LCCN ???? FF 195, US$29.95. From the authors: “Written by two well known French Internet pioneers, this book presents all the French use of Internet for company chiefs, students, schools and gives some tips for a total and comprehensive adaptation of Internet in France.”.

Cockcroft:1998:SPT

[CP98] Adrian Cockcroft and Richard Pettit. Sun Performance and Tuning: Java and the Internet. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, February 1998. ISBN 0-13-095249-4. xxxvi + 587 pp. LCCN QA76.8.S86C63 1998.

US$45. URL http://www.
fdds.com/books/books/Cockcroft/
Cockcroft.html;   http:/

                                        /www.prenhall.com/ptrbooks/
                                        ptr_0130952494.html.

Croll:1999:MBD

[CP99] Alistair Croll and Eric Packman. Managing Bandwidth: Deploying QOS Across Enterprise Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-011391-3. ???? pp. LCCN TK5105.5.C76

                                        1999. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0130113913.html.

Cranor:1992:IMC

[Cra92]  Charles D. Cranor.  An

implementation model for connection-oriented Internet protocols. Thesis (M.S.), Department of Computer Science, Washington University, St. Louis, MO, USA, 1992. ix + 92 pp.

Cranor:1999:IP

[Cra99] Lorrie Faith Cranor. Internet privacy. Communications of the ACM, 42(2): 28–38, February 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-2/
                                        p28-cranor/.

Clark:1998:PIB

[CRH98] K. Clark, P. J. Robinson, and R. Hagen. Programming Internet based DAI applications in Qu-Prolog. Lecture Notes in Computer Science, 1544:137–151, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crispen:1994:R

[Cri94] Patrick Douglas Crispen.

Roadmap, 1994.

Corn-Revere:1995:NAC

[CRI95] Robert Corn-Revere and Cato Institute. New age comstockery: Exon vs. the Internet. Policy analysis; no. 233, Cato Institute, Washington, DC, USA, 1995. 27 pp.

Cronin:1993:IBR

[Cro93] Mary J. Cronin. Internet business resources. Database, 16(6):47–??, December 1, 1993. CODEN DTBSDQ. ISSN 0162-4105.

Cronin:1994:DBI

[Cro94a] Mary J. Cronin. Doing business on the Internet: how the electronic highway is transforming American companies. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01770-7. xi + 308 pp. LCCN HD30.335.C76 1994. US$29.95.

Crowe:1994:ETE

[Cro94b] Elizabeth Powell Crowe. The Electronic Traveller: Exploring Alternative Online Systems. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1994. ISBN 0-8306-4498-9. xxiv + 226 pp.

LCCN TK5105.87 .C76 1994. US$16.95.

Cronin:1995:DMB

[Cro95]  Mary J. Cronin.  Doing

more business on the Internet: how the electronic highway is transforming American companies. Van Nostrand Reinhold, New York, NY, USA, second edition, 1995. ISBN 0-442-02047-3. xiii + 368 pp. LCCN HD30.335 .C76 1995. Rev. ed. of: Doing business on the Internet. c1994.

Cronin:1996:GAI

[Cro96a] Mary J. Cronin. Global advantage on the Internet: from corporate connectivity to international competitiveness. Van Nostrand Reinhold, New York, NY, USA, 1996. ISBN 0-442-01938-6. x + 358 pp. LCCN HF54.56 .C76 1996.

Cronin:1996:ISH

[Cro96b] Mary J. Cronin. The Internet strategy handbook: lessons from the new frontier of business. Harvard Business School Press, Boston, Mass., 1996. ISBN 0-87584-720-X, 0-87584-631-9. xi + 296 pp. LCCN HD30.37 .I57 1996.

Crowley:1996:OGI

[Cro96c] William R. Crowley. Oil and gas on the Internet. Gulf Pub. Co., Houston, TX, USA, 1996. ISBN 0-88415-155-7. x + 166 pp. LCCN TN867.C76 1996.

Castagnetto:1999:PPP

[CRS^{+}99] Jesus Castagnetto, Harish Rawat, Sascha Schumann, Chris Scollo, and Deepak Veliath. Professional PHP Programming. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-296-3. xxi + 909 pp. LCCN QA76.73.P23 C37 2000. US$49.99, UK£35.99, CDN$74.95. URL http://

www.wrox.com/Books/Book_
Details.asp?ISBN=1861002963.

Crumlish:1995:GTI

[Cru95a] Christian Crumlish. A guided tour of the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1619-1. xxiii + 276 pp. LCCN TK5105.875.I57C78 1995.

Crumlish:1995:ID

[Cru95b] Christian Crumlish. The Internet dictionary. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1675-2 (paperback). x + 226 pp. LCCN TK 5105.875 I57 C79 1995.

Cruz:1995:III

[Cru95c] Raul A. Cruz. The Internet and the insurance industry connection. Technical report, RAC Associates, Thousand Oaks, CA, USA, 1995. viii + 47 pp.

Crumlish:1996:AI

[Cru96a] Christian Crumlish. The ABCs of the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1887-9 (paperback). xxi + 362 pp. LCCN ????

Crumlish:1996:IBP

[Cru96b] Christian Crumlish. The Internet for busy people. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882108-8. xxiii + 275 pp. LCCN TK5105.875.I57C792 1996.

Comer:1991:ITIc

[CS91a] Douglas Comer and David Stevens. Internetworking with TCP/IP: Client-Server Computing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-474222-2. LCCN TK5105.5 .C59 1991.

Comer:1991:ITIb

[CS91b] Douglas Comer and David Stevens. Internetworking with TCP/IP: Design, Implementation and Internals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-472242-6. LCCN TK5105.5 .C59 1991.

Comer:1991:ITDa

[CS91c] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP. Design, Implementation, and Internals, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-472242-6. xx + 532 pp. LCCN TK5105.5 .C59 1991.

Comer:1991:ITI

[CS91d] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP. Design, Implementation, and Internals, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-465378-5, 0-13-472242-6. LCCN TK5105.5 .C59 1991-92.

Comer:ITD91

[CS91e] Douglas E. Comer and David L. Stevens. Internetworking with TCP/ IP. Design, Implementation, and Internals, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-472242-6. LCCN TK5105.5 .C59 1991-92.

Comer:1993:ITCb

[CS93a] Douglas E. Comer and David L. Stevens. Internetworking with TCP/ IP. Client-server Programming and Applications, BSD Socket Version. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-020272-X. xxiv + 498 pp. LCCN TK5105.585 .C66 1993. US$45.00.

Comer:1993:ITCa

[CS93b] Douglas E. Comer and David L. Stevens. Internetworking with TCP/ IP. Client-server Programming and Applications for the AT&T TLI. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-097627-X. xx + 508 pp. LCCN ???? US$45.00.

Canter:1994:HMF

[CS94a] Laurence A. Canter and Martha S. Siegel. How to make a fortune on the information superhighway: everyone’s guerrilla guide to marketing on the Internet and other on-line services. HarperCollins College Publishers, New York, NY, USA, 1994. ISBN 0-06-270131-2, 0-06-272065-1. 234 pp. LCCN HF6146.I5 C36 1994.

Christian:1994:CI

[CS94b] Eliot J. Christian and Geological Survey (U. S.). Connecting to the Internet, 1994. 1 videocassette (45 min.).

Canter:1995:CHF

[CS95] Laurence A. Canter and Martha S. Siegel. Como hacer una fortuna en la autopista de la informacion: su “Guia del guerrillero” para hacer mercadeo a travees de la Internet y — servicios online. Lasser Press Mexicana, Mexico, DF, Mexico, 1995.

ISBN 968-458-477-6 (paperback). 326 pp. LCCN ????

Comer:1996:ITIb

[CS96a] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP: Client-server programming and applications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-260969-X. ???? pp. LCCN TK5105.585 .C66 1991.

Comer:1996:ITIa

[CS96b] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP: Design, implementation, and internals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-125527-4. ???? pp. LCCN TK5105.585 .C66 1991.

Comer:1997:ITI

[CS97] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP. Vol. III. Client-Server Programming and Applications — Windows Sockets Version. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-848714-6. xxvi + 513 pp. LCCN TK5105.585 .C66 1995. URL http:/

/www.phptr.com/ptrbooks/
esm_0138487146.html.

Comer:1998:ITI

[CS98] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP. Vol.

II, ANSI C Version: Design, Implementation, and Internals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, June 1998. ISBN 0-13-973843-6. 608 pp. LCCN ???? US$62.00. URL

http://www.prenhall.com/
allbooks/esm_0139738436.
html.

Cook:1998:UGD

[CSJ98] Jonathan Cook, Dwaine Snow, and Calene Janacek. The Universal Guide to DB2 for Windows NT. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-099723-4. 528 pp. LCCN QA76.9.D3C6725

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0130997234.html. Includes

CD-ROM.

Clement:1995:STIa

[CT95a] Gail P. Clement and Roy Tennant. Science and technology on the Internet: an instructional guide. Internet workshop series; no. 4. Library Solutions Press, Berkeley, CA, USA, 1995. ISBN 1-882208-12-9. xx + 228 pp. LCCN Q179.96 .C54 1995. US$60.00. Supplement to Crossing the Internet threshold.

Clement:1995:STIb

[CT95b] Gail P. Clement and Roy Tennant. Science and technology on the Internet: an instructional guide. Internet workshop series; no. 4. Library Solutions Press, Berkeley, CA, USA, plus edition, 1995. ISBN 1-882208-13-7. xx + 264 pp. LCCN Q179.97 .C54 1995. US$60.00. “Supplement to Crossing the Internet threshold.” Title on computer disk labels: Science and technology on the Internet PLUS. Diskettes include presentation slides in Power-Point for Macintosh or Windows.

Clement:1996:STIa

[CT96a] Gail P. Clement and Roy Tennant. Science and technology on the Internet: an instructional guide. Internet workshop series; no. 4. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, plus rev. edition, 1996. ISBN 1-882208-13-7. xx + 228 pp. LCCN Tk 5105.875.I57T46 1993 Suppl.

3. Includes 2 computer disks.

Clement:1996:STIb

[CT96b] Gail P. Clement and Roy Tennant. Science and technology on the Internet PLUS: instructor’s supplement. Number 4 in Internet workshop series. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 2nd printing, minor revisions. edition, 1996. ISBN ???? various pp. LCCN ????

Cutkosky:1996:MCE

[CTG96] Mark R. Cutkosky, Jay M.

Tenenbaum, and Jay Glicksman. Madefast: Collaborative engineering over the Internet. Communications of the ACM, 39 (9):78–87, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org/pubs/toc/
                                        Abstracts/cacm/234474.html.

CTN:1995:CIV

[CTN95] CTN Computer Television Network. Capitalizing on the Internet Viewer Guidebook. Network World, Farmingham, MA, USA, 1995. ISBN ???? 182 pp. LCCN ???? Includes CD-ROM.

Cullen:1995:IVO

[Cul95] C. Cullen. Internet vulnerabilities of one-time passcode mechanisms. In IEEE Computer Society. Technical Committee on Computer Communications [IEE95], pages 184–186. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Cunningham:1993:TVI

[CW93] Ann Marie Cunningham and Wendy Wicks. Three views of the Internet. NFAIS report series; 1993/3. National Federation of Abstracting and Information Services, Philadelphia, PA, USA, 1993. ISBN 0-942308-42-5. vi + 105 pp. LCCN TK5105.875.I57 T49 1993.

Chang:1999:IRG

[CWJT99] C.-H. Chang, H.-K. Wu, M.- H. Jin, and Y.-O. Tseng. Intelligent routing for global broadband satellite Internet. Lecture Notes in Computer Science, 1593:80–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatterjee:1997:SDF

[CX97] S. Chatterjee and W. Xiao.

Selective discarding with FEC for ATM-based Internetworks. In IEEE Computer Society. Technical Committee on Computer Communications [IEE97], pages 284–293. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Anonymous:1996:CGI

[Cyb96] The cyberskeptic’s guide to Internet research, page various, 1996. ISSN 1085-2417. BiblioData, Needham Heights, MA, USA.

Yu:1995:IMT

[cYlC95] Chi chang Yu and Kil lak Chong. Intonet mujakchong ttara hagi: E-meil puto WWW kkaji. Kommaeng chamnun kaeguri siriju; 3. Kilbot, Soul-si, Korea, kaejong chungbopan. edition, 1995. ISBN 89-7560-203-6. 424 pp. LCCN ????

Chapman:1995:BIF

[CZ95] D. Brent Chapman and Elizabeth D. Zwicky. Building Internet firewalls. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, minor corrections, nov. 1995 edition, 1995. ISBN 1-56592-124-0. xxvi + 517 pp. LCCN TK5105.875 .I57 C43 1995. On cover: Internet security.

DFA:1993:EI

[D. 93] D. and F. Associates. Exploring Internet, 1993. 1 videocassette (120 min.).

Dworatschek:1986:IGI

[D^{+}86] Sebastian Dworatschek et al., editors. The international GPM/INTERNET symposium, 1986: Project management software; application, implementation, trends. Gesellschaft fur Projektmanagement INTERNET, Munich, Germany, 1986. ISBN 3-924841-05-5. LCCN T56.8.G65 1986.

Damvik:1993:IVV

[Dam93] Mats Damvik. Internet: ett verktyg vid forskning i utlandsk och internationell rätt. IRI-rapport 1993:2, Institutet för rattsinformatik, Stockholms universitet, Stockholm, Sweden, 1993. 78 pp.

Damon:1997:WII

[Dam97] Lee Damon. WWWhither(ing)

Internet. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/1997-12/imho.html.

Moreira:1997:ADC

[dAMW97] Dilvan de Abreu Moreira and Les T. Walczowski. AGENTS a distributed client-server system for leaf cell generation. ACM Transactions on Design Automation of Electronic Systems, 2(1):42–61, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-1/p42-moreira/p42-
moreira.pdf;   http://www.
acm.org/pubs/citations/
journals/todaes/1997-2-
1/p42-moreira/.

Darby:1992:TI

[Dar92] Christine M. Darby. Traveling on the Internet. ERIC reports ED 350 007, ?ERIC Document Reproduction Service, operated by CBIS Federal, Inc., Alexandria, VA, USA, 1992. 48 pp.

DC:1993:IBL

[Dar93a] Dartmouth College. Internet and Bitnet listServers. Technical report, Dartmouth College, Hanover, NH, USA, 1993. various pp.

Darwin:1993:IWO

[Dar93b] Charles Darwin. The Internet wiretap online edition of The voyage of the Beagle, 1993.

DataComm:1974:DCM

[Dat74] Data Communications: McGraw-Hill’s Networking Technology Magazine, 1974. CO-DEN DACODM. ISSN 0363-6399. New York, NY, USA. Monthly; available free to qualified subscribers.

DNSHI:1994:ISM

[Dat94] Internet security monthly, page various, 1994. ISSN 1079-5669. DataNet Security Holding Inc., Washington, DC, USA.

DataTrek:1996:NW

[Dat96] Data Trek, Inc. NetPAC for windows, 1996. Includes 3 computer disks.

Davidson:1988:ITI

[Dav88] John Davidson. An Introduction to TCP/IP. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96651-X (paperback), 3-540-96651-X (paperback). xii + 100 pp. LCCN TK5105.5 .D379 1988. US$28.00.

Davis:1994:MGI

[Dav94] Peter T. Davis. Manager’s guide to Internet security. Technical report, Computer Security Institute, San Francisco, CA, USA, 1994. 21 pp.

Davenport:1995:FRT

[Dav95] Martha K. Davenport. Factors related to the Tennessee K–12 educators’ implementation of the Internet into classroom activities and professional development. Thesis (Ed. D.), Dept. of Educational Leadership and Policy Analysism, East Tennessee State University, Johnson City, TN, USA, 1995. xi + 135 pp.

DVA:1996:BII

[Dav96] David Vine Associates. Business intelligence on Internet: a compendium of selected World Wide Web sites. Technical report, David Vine Associates, Princeton, NJ, USA, 1996. various pp.

Davie:1998:SIN

[Dav98] Bruce Davie. Switching in IP Networks: IP Switching, Tag Switching, and Related Technologies. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-505-3. xv + 255 pp. LCCN TK5105.875.I57 D38 1998. US$50.95.

Deering:1990:MRD

[DC90] Stephen E. Deering and David R. Cheriton. Multicast routing in datagram internetworks and extended LANs. ACM Transactions on Computer Systems, 8(2): 85–110, May 1990. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1990-8-2/
p85-deering/.

Drummond:1994:LTM

[DC94] Rik Drummond and Nancy Cox. The LAN Times E-mail Resource Guide. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882052-9. xxv + 334 pp. LCCN TK5105.73 .D78 1994. US$29.95.

Denning:1998:IBC

[DD98] Dorothy E. Denning and Peter J. Denning, editors. Internet besieged: countering cyberspace scofflaws. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30820-7. xii + 547 pp. LCCN HV6773.I57 1998.

DeVries:1994:EII

[De 94] Henry De Vries. Exploring Internet inservice seminar. Technical report, Electronic Technology Group, Cornell Cooperative Extension, Ithaca, NY, USA, January 17–18, 1994. various pp.

DEmanuele:1995:ICS

[D’E95] A. D’Emanuele. The Internet: (2) connecting and searching for information. Pharmaceutical journal, 254 (6823):58–??, January 14, 1995. CODEN PHJOAV. ISSN 0031-6873.

Dean:1996:PTM

[Dea96] Damon A. Dean. A pocket tour of multimedia on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1821-6. xii + 148 pp. LCCN QA76.575.D42 1996.

December:1995:ISI

[Dec95] John December. Information sources: the Internet and computer-mediated communication. Technical report, ????, ????, January 26, 1995. 93 pp.

Deering:1993:SSI

[Dee93] Stephen E. Deering. SIP:

Simple Internet protocol. IEEE network, 7(3):16–??, May 1, 1993. ISSN 0890-8044.

Deel:1995:ERS

[Dee95] Nancy Deel. Expanding reference service using the Internet, 1995. 1 sound cassette.

Delfino:1993:PMI

[Del93] Erik Delfino. PC monitor — the Internet toolkit: File compression and archive utilities. Online, 17(6):90–??, November 1, 1993. CODEN ONLIDN. ISSN 0146-5422.

Dell:1998:CM

[Del98] Thomas Dell. CORE MCSE. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. LCCN ????

URL http://www.phptr.
com/bookseri/mcse.html.

Denning:1989:SCI

[Den89] Peter J. Denning. The science of computing: The Internet worm. American Scientist, 77(2):126–??, March 1, 1989. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Denning:1990:CUA

[Den90] Peter J. Denning, ed. Computers Under Attack: Intruders, Worms, and Viruses. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-201-53067-8. xx + 554 pp. LCCN QA76.9.A25 C667 1990. US$23.95. Details of celebrated network security cases, including those described in the previous two books. Includes Stoll’s original article about the Wiley Hacker, and responses and articles by others on the same subject. Has extensive coverage of the 1988 Internet Worm. Also includes information on viruses. Has quite a bit of material on the cultures of the networks, and on social, legal, and ethical matters. Starts with the standard historical network papers, including “Notable Computer Networks” by Quarterman and Hoskins.

Dennis:1996:DRG

[Den96] Richard Dennis. Database review: guide to selecting an Internet provider. Technical report, American Bar Association, Washington, DC, USA, 1996. 571–596 pp.

Denning:1999:IWS

[Den99] Dorothy Elizabeth Robling Denning. Information warfare and security. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-43303-6. xvii + 522 pp. LCCN U163 .D46 1999. US$31.45. ACM order number 704982.

Dern:1994:IGN

[Der94]  Daniel P. Dern.  The

Internet Guide for New Users. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-016510-6 (hardcover), 0-07-016511-4 (paperback). xxvii + 570 pp. LCCN TK5105.875.I57 D47 1994. US$40.00 (hardcover), US$27.95 (paperback). Contents: Ramping up, getting started — Internet history and technology: a brief introduction — Internet on a dollar a day: how to get a user account and plug in — Internet naming and addressing — Enough UNIX to survive as an Internet user — Electronic mail, usenet, remote login, and file transfer: the four basic Internet food groups — Electronic mail: how to exchange messages with other Internet users and other exciting things you can do — Usenet: the bulletin board of the Internet — Remote login with telnet — Transferring files with FTP — Navigating the Internet — Internet dashboard: navigating the Internet rapids — Gopher, archie, WAIS, and others: meet the navigators — Help, problems, security, and other aspects of being an Internet citizen — Being an Internet citizen — Help! — Commercial services, archives, communities, and miscellany — Communities of interest: a look at who’s using the Internet — Commercial and other information services on the Internet: databases, libraries, and other info-mongers — Anonymous-FTP and other archives on the Internet: gigabytes for the taking. Contents: (cont) Electronic mailing lists: how to join, read, and participate in them — BITNET: another network worth knowing about — And away you go: tips, toys, and suggestions — Third-party TCP/IP software for your computer — Common editors on UNIX systems and how to exit them — InterNIC access information — Free! a simple Internet front-end UNIX shellscript — Internet-related organizations worth knowing about.

DeRouchey:1996:IQF

[DeR96]  Bill DeRouchey.  Inter-

net quick: a fast and simple guide using NetCruiser. Franklin, Beadle and Associates, Wilsonville, OR, USA, 1996. ISBN 1-887902-10-4. ???? pp. LCCN TK5105.882.D47 1996.

DeRoest:1999:SUN

[DeR99] James DeRoest. Samba: Unix and NT internetworking. Mc-Graw-Hill, New York, NY, USA, 1999. ISBN 0-07-135104-3. ???? pp. LCCN QA76.76.O63 D4727 1999.

Descy:1993:AAIa

[Des93a] Don Descy. ALL ABOARD THE INTERNET, the Internet/school connection. TechTrends: for leaders in education and training, 38(4): 15–??, September 1, 1993. CODEN TETREF. ISSN 8756-3894.

Descy:1993:AAIb

[Des93b] Don E. Descy. All aboard the Internet, where to start: An Internet resource guide. TechTrends: for leaders in education and training, 38(5): 39–??, October 1, 1993. CO-DEN TETREF. ISSN 8756-

Descy:1994:AAI

[Des94]  Don Descy.  All aboard

the Internet. World-Wide Web: Adding multimedia to cyberspace. TechTrends: for leaders in education and training, 39(4):15–??, September 1, 1994. CODEN TETREF. ISSN 8756-3894.

Deutsch:1992:RDI

[Deu92] Peter Deutsch. Resource discovery in an Internet environment. Thesis (M.Sc.), McGill University, Montréal, PQ, Canada, 1992. 1 microfiche; University Microfilms order no. UMI00414690.

Devillers:1996:NNQ

[Dev96] James Devillers. Neural Networks in QSAR and Drug Design. Academic Press, New York, NY, USA, 1996. ISBN 0-12-213815-5. x + 284 pp. LCCN ???? US$141.95.

DiPaola:1998:CIP

[DFG^{+}98] Steve DiPaola, Barnett Fox, A. Thomas Goldberg, Mark S. Meadows, and Celia Pearce. Characters on the Internet (panel): the next generation. In Grisson et al. [GMA^{+}98], pages 225–

227. ISBN 1-58113-046-5, 1-58113-048-1 (CD-ROM). LCCN T385 .S54 1998b.

URL http://www.acm.org:
80/pubs/citations/proceedings/
graph/280953/p225-dipaola/

. ACM order number for CD-ROM: 435984. ACM order number: 435983.

Dewan:1998:ISP

[DFS98] Rajiv M. Dewan, Marshall L.

Freimer, and Abraham Seidmann. Internet service providers, proprietary content, and the battle for users’ dollars. Communications of the ACM, 41(8): 43–48, August 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-8/
p43-dewan/.

daCruz:1993:UCK

[dG93] Frank da Cruz and Christine M. Gianone. Using C-Kermit. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Englewood Cliffs, NJ 07632, USA, 1993. ISBN 1-55558-108-0 (DP), 0-13-037490-3 (PH). xxi + 514 pp. LCCN TK5105.9.D33 1993. Available in German translation [?].

Drake:1994:IYL

[DG94] Chuck Drake and Educational Reform Group. Internet your lesson in navigating the information super highway: a how-to guide to mining its treasures, 1994. 1 videocassette (ca. 70 min.).

Drake:1995:NEW

[DG95] Chuck Drake and Educational Reform Group. Netscape: the easiest way to surf the Internet: your guide to downloading, searching and browsing, 1995. 1 videocassette (40 min.).

Dalton:1997:AMG

[DG97] C. I. Dalton and J. F. Griffin. Applying military grade security to the Internet. Computer Networks and ISDN Systems, 29(15):1799–1808, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        comnet/cas_sub/browse/browse.
                                        cgi?year=1997&volume=29&
                                        issue=15&aid=1775.

Duncan:1993:SRD

[DGH93] Denise Duncan, Curtis Generous, and Judy F. Hunter. Support of research and development activities via the Internet: NASA’s access mechanism. Internet research, 3(3):37–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

Deivert:1996:FVI

[DH96] Bert Deivert and Dan Harries. Film and video on the Internet: the top 500 sites. Michael Wiese Productions, Studio City, CA, USA, 1996. ISBN 0-941188-54-X. vii + 252 pp. LCCN PN1998.A1D36 1996.

Dreilinger:1997:ESS

[DH97] Daniel Dreilinger and Adele E.

Howe. Experiences with selecting search engines using metasearch. ACM Transactions on Information Systems, 15(3):195– 222, July 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/dreilinger.
html.

Doraswamy:1999:INS

[DH99] Naganand Doraswamy and Dan Harkins. IPSec: The New Security Standard for the Internet, Intranets, and Virtual Private Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-011898-2. ???? pp. LCCN TK5105.567.D67

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130118982.html.

Dhawan:1994:NDD

[Dha94] Sanjay Dhawan. Networking Device Drivers. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01943-2. 398 pp. LCCN TK5105.9 .D49 1995. US$44.95.

Doty:1996:AGI

[DHM^{+}96] Gloria M. Doty, Mary K.

Hartz, Leslie Anderson Morales, Charlotte Nusberg, Hugh O’Connor, and Helen M. White. The AARP guide to Internet resources related to aging. Technical report, Research Information Center, American Association of Retired Persons, Washington, DC, USA, 1996. x + 71 pp.

Diamond:1995:IGW

[DHS95] Joel Diamond, Valda Hilley, and Howard Sobel. Internet gizmos for windows. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-451-4. xxxiv + 913 pp. LCCN ????

DEC:1982:DIX

[Dig82] Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA. Version 2.0, DEC-Intel-Xerox (DIX) Ethernet Standard: Ethernet Local Area Network Specification Version 2.0, DEC part number: AA-K759B-TK edition, 1982. 103 pp. US$31.00.

DEC:1990:NTG

[Dig90a] Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA. Network Troubleshooting Guide, digital part number ek-339ab-gd-002 edition, August 1990. 278 pp. US$95.00.

DEC:1990:ODB

[Dig90b] Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA. OPEN DECconnect Building Wiring Components and Applications Catalog, 1990.

DEC:1993:TNB

[Dig93] Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA. Telecommunications and Networks Buyer’s Guide, 1993. Published twice a year.

DEC:1995:UIA

[Dig95] Digital Equipment Corporation. Using Internet applications, 1995. Includes CD-ROM.

Dillon:1993:AIIa

[Dil93] Martin Dillon. Assessing information on the Internet: toward providing library services for computermediated communication: results of an OCLC research project. Technical Report OCLC/OR/RR-93/1, OCLC Online Computer Library Center, Inc., Office of Research, Dublin, OH, USA, January 31, 1993. 33 pp.

Dinar:1994:KEG

[Din94] Ali B. Dinar. K–12 electronic guide for African resources on the Internet. Technical report, University of Pennsylvania, African Studies Program, Philadelphia, PA, USA, 1994. 15 pp.

DCS:1993:IIJ

[Dis93] IBM Internet journal, 1993.

ISSN 1068-1396. Cardinal Business Media, Dallas, TX, USA.

Dixon:1995:ITP

[Dix95] David N. Dixon. Insurgent technology: the political ramifications of the Internet in Africa. Technical report, ????, ????, 1995. 26 pp.

Dillon:1993:AIIb

[DJB93] Martin Dillon, Erik Jul, and Mark Burge. Assessing information on the Internet: Toward providing library services for Computer-Mediated communications. Internet research, 3(1):54–??, Spring 1993. CODEN IRESEF. ISSN 1066-2243.

Dougherty:1994:MHMa

[DK94a] Dale Dougherty and Richard Koman. The Mosaic Handbook for Microsoft Windows. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1994. ISBN 1-56592-094-5. 200 pp. LCCN TK5105.875.I57 D686 1994. US$29.95.

Dougherty:1994:MHMb

[DK94b] Dale Dougherty and Richard Koman. The Mosaic Handbook for the Macintosh. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1994. ISBN 1-56592-096-1. 170 pp. LCCN TK5105.875.I57 D68 1994. US$29.95.

Dougherty:1994:MHX

[DKF94] Dale Dougherty, Richard Koman, and Paula Ferguson. The Mosaic Handbook for the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1994. ISBN 1-56592-095-3. 200 pp. LCCN TK5105.875.I57 D68 1994. US$29.95.

Dobson:1995:IVD

[DKF^{+}95] Jim Dobson, Dawn E.

Kuczwara, Aaron Freeman, Matt Ferraro, Matthew B. Brandabur, Stephen John Stofflet, Peter Zelchenko, and Eric D. Wanger. The Internet video doing donuts on the information superhighway, 1995. 1 videocassette.

Dahl:1996:IC

[DL96] Andrew Dahl and Leslie Lesnick. Internet commerce. New Riders, Indianapolis, Ind., 1996. ISBN 1-56205-496-1. xii + 382 pp. LCCN HF5415.1265.L47 1996.

Davies:1992:PTIb

[DLC92a] Megan Davies, Debra Legare, and Cynthia Clark, editors. Proceedings of the twentyfifth Internet Engineering Task Force, Washington, DC, November 16–20, 1992. Corporation for National Research Initiatives, Reston, VA, USA, 1992. ISBN ???? LCCN ????

Davies:1992:PTIa

[DLC92b] Megan Davies, Debra Legare, and Cynthia Clark, editors. Proceedings of the twentythird Internet Engineering Task Force, San Diego, California, March 16–20, 1992. Corporation for National Research Initiatives, Reston, VA, USA, 1992. ISBN ???? LCCN ????

Davis:1995:WLA

[DM95] Peter T. Davis and Craig R. McGuffin. Wireless Local Area Networks: Technology Issues and Strategies. Mc-Graw-Hill, New York, NY, USA, 1995. ISBN 0-07-015839-8. xv + 250 pp. LCCN TK5105.7.D388 1995. US$40.00.

Dee:1999:LMB

[DN99] Daniel Dee and Dale Nielsen. Linux means business: Upgrading Linux over the Internet. Linux Journal, 61:??, May 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dale:1996:WWI

[DO96] Bob Dale and Barbara Opyt. The Web workgroup: integrating Lotus Notes and the Internet. OnWord Press, Santa Fe, NM, USA, 1996. ISBN 1-56690-110-3. ???? pp. LCCN HF5548.4.L692D35 1996.

DeGroot:1995:IGG

[DOB95] Paul De Groot, Dick Oliver, and Lane Boyd. Internet graphics gallery. Que Corp., Indianapolis, IN, USA, 1995. ISBN 0-7897-0137-5. 593 + 16 pp. LCCN T 385 D44 1995. US$39.99, CDN$53.99. Includes CD-ROM.

Dodd:1999:EGT

[Dod99] Annabel Z. Dodd. The Essential Guide to Telecommunications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-014295-6. ???? pp. LCCN TK5101.D54

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130142956.html.

Donohoe:1995:JMM

[Don95] Lisa Donohoe. JWA meeting on Mexico after NAFTA and Internet communications. Technical Report D95-1897, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, February 1995. 16 pp.

Donkers:1996:LIS

[Don96] Arthur Donkers. Linux as an Internet server. Sys Admin: The Journal for UNIX Systems Administrators, 5(1): 52–??, January 1996. CO-DEN SYADE7. ISSN 1061-

Dickie:1995:LHC

[DOO95] Cheryl Dickie, Trevor Owen, and Ronald Davis Owston. The learning highway: the Canadian student guide to the Internet. Key Porter Books, Toronto, ON, Canada, 1995. ISBN 1-55013-615-1. ???? pp. LCCN ????

Downing:1986:FCC

[Dow86] Alan Robert Downing. Flow control and congestion control for a transparent Internet environment. Thesis (M.S.), Department of Computer Science, University of California, Los Angeles, Los Angeles, CA, USA, 1986. ix + 107 pp. Typescript (photocopy).

Presno:1994:OW

[dPW94] Odd de Presno and Mike Weaver. The Online World. Productive Publications, Toronto, ON, Canada, 1994. ISBN 0-920847-89-7. 300 pp. US$47.00.

December:1994:WU

[DR94] John December and Neil Randall. World-Wide Web Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, November 1994. ISBN 0-672-30617-4. 1058 pp. LCCN TK 5105.875 I57 D43 1994. US$35.00, CDN$47.95. Computer Literacy Bookshops says: Not only the largest book on the Web, but also the most complete. 300 pages cover fundamental concepts and individual wab browsers for different platforms. Covers navigation and searching techniques and list Web sites by subject. Written by a team of WWW experts, the best part is the inclusion of details on technical information. Loads of information on designing and creating a Web site, working with HTML editors and filters, creating forms and image mapping.

December:1995:WU

[DR95] John December and Neil Randall. World-Wide Web Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1995. ISBN 0-672-30737-5. 1346 pp. LCCN TK 5105.875 I57 D43 1994. US$39.99, CDN$53.99, UK£30.95. URL http:/

/www.december.com/works/
wwwu/book.html.

December:1996:WU

[DR96] John December and Neil Randall. World-Wide Web Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1996. ISBN 1-57521-040-1. 1392 pp. LCCN TK 5105.875 I57 D43 1994. US$49.99, CDN$67.99, UK£46.95. URL http:/

/www.december.com/works/
wwwu/book.html. Includes

CD ROM.

Drager:1995:QSH

[Dra95] Philip John Drager. A qualitative study of how experts with knowledge of public relations and the Internet believe that the Internet will affect future public relations performance. Thesis (M.A.), Ball State University, Muncie, IN, USA, 1995. 35 pp.

Drew:1992:JCG

[Dre92] Wilfred Drew. Not just cows: a guide to Internet/Bitnet resources in agriculture and related sciences. Technical report, ????, ????, 1992. 57 pp.

Drew:1993:JCGa

[Dre93a] Wilfred Drew. Not just cows: a guide to Internet/Bitnet resources in agriculture and related sciences. Technical report, ????, ????, 1993. 49 pp.

Drew:1993:JCGb

[Dre93b] Wilfred Drew. Not just cows: a guide to Internet/Bitnet resources in agriculture and related sciences. Technical report, ????, ????, 1993. 86 pp.

Drew:1994:GIB

[Dre94] Wilfred Drew. A guide to Internet/Bitnet resources in agriculture and related sciences. Technical report, ????, ????, 1994. 50 pp.

Drew:1995:JCG

[Dre95] Wilfred Drew. Not just cows a guide to Internet/Bitnet resources in agriculture and related sciences, 1995.

Douglas:1995:NIY

[DS95] Tamsin Douglas and Ned Snell. Navigating the Internet with your Macintosh. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 1-57521-043-6. xviii + 337 + 3 pp. LCCN TK5105.875.I57D68 1995.

Drukier:1996:HGI

[DS96] Laurie Drukier and Kim Schrader. How to get on the Internet: finding, installing, and using Internet connection software to get online. Go!Tools. Motion Works Pub., San Francisco, CA, USA, 1996. ISBN 1-57712-000-0 (paperback). ???? pp. LCCN TK5105.875.I57D78 1996.

Detmer:1997:UII

[DS97] William M. Detmer and Edward H. Shortliffe. Using the Internet to improve knowledge diffusion in medicine. Communications of the ACM, 40(8): 101–108, August 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-8/
p101-detmer/.

Dufour:1995:I

[Duf95] Arnaud Dufour. Internet.

Que sais-je?; 3073. Presses Universitaires de France, Paris, France, 1995. ISBN 2-13-047469-1. 127 pp. LCCN ????

Dufour:1996:I

[Duf96] Arnaud Dufour. Internet. Que sais-je?; 3073. Presses Universitaires de France, Paris, France, 2e edition corrigée edition, 1996. ISBN 2-13-047469-1. 127 pp. LCCN ????

Dupuy:1995:CIG

[Dup95] John Dupuy. The Complete Idiot’s Guide to Netscape. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-612-7. 300 pp. LCCN ???? US$16.99.

Dutton:1998:UOC

[Dut98] Harry J. R. Dutton. Understanding Optical Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-020141-3. xxxvi + 760 pp. LCCN TK5103.59.D88

                                        1998. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0130201413.html.

deVivo:1998:ISA

[dVdVI98] Marco de Vivo, Gabriela O. de Vivo, and Germinal Isern. Internet security attacks at the basic levels. Operating Systems Review, 32(2): 4–15, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dyrli:1993:IBG

[Dyr93] Odvard Egil Dyrli. The Internet: Bringing global resources to the classroom. Technology and learning, 14 (2):50–??, October 1, 1993. ISSN 1053-6728.

Engle:1993:IC

[E^{+}93] Mary E. Engle et al. Internet Connections: a Librarian’s Guide to Dial-Up Access and Use. LITA monographs;

3. American Library Association, Chicago, IL, USA, 1993. ISBN 0-8389-7677-8. x + 166 pp. LCCN Z674.82.I59 I58 1993. US$22.00.

Eagan:1993:OCS

[Eag93] Ann. Eagan. Order out of chaos: Science databases on the Internet. Database, 16 (6):62–??, December 1, 1993. CODEN DTBSDQ. ISSN 0162-4105.

Eager:1994:ISI

[Eag94a] Bill Eager. The Information Superhighway Illustrated. Que Corporation, Indianapolis, IN, USA, October 1994. ISBN 1-56529-892-6. 184 pp. US$24.99. A vividly illustrated book that explains how the technology of the information superhighway (the networks, hardware and software) works and overviews current and future applications.

Eager:1994:UWW

[Eag94b] Bill Eager. Using the World Wide Web. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 0-7897-0016-6. 648 pp. LCCN TK5105.888 .E34 1994. US$27.99, CDN$38.99. Computer Literacy Bookshops says: User friendly tutorial for the fastest-growing part of the Internet. Discusses WWW connections with Cello, Mosaic for Windows and Mac, and the new upcoming versions of Moasic. Most of it is dedicated to finding resources on the Web in areas of art, music, business, education, government and so on.

Eager:1994:UI

[Eag94c] William Eager. Using the Internet. Que Corporation, Indianapolis, IN, USA, December 1994. ISBN 0-7897-0096-4 (paperback). xxiv + 369 pp. LCCN TK5105.875.I57E22 1994. US$19.99. Explains in step-by-step fashion how to get on the Internet and how to use graphical, point-and-click computer programs to navigate through the vast resources of the Internet.

Eager:1995:IP

[Eag95a]  Bill Eager.  The Infor-

mation Payoff: the Manager’s Concise Guide to Making PC Communications Work. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-158296-8. 246 pp. LCCN HD30.3 .E17 1995. US$19.95. How companies around the world are using electronic communications and the Internet to improve productivity and enhance employee participation. Interviews with seniorlevel managers at Fortune 500 companies provide insight into the latest applications and benefits of electronic mail, interactive communications, video teleconferencing, and electronic publications.

Eager:1995:WWW

[Eag95b] William Eager. World Wide Web and Mosaic, 1995. 1 videocassette (VHS) (120 min.) in envelope.

EARN:1994:GNR

[EAR94] EARN Association. Guide to network resource tools. Discussing many different resources and services available on the Internet …. This RFC (rfc1580.txt) is obtainable

from nic.ddn.mil in /rfc

…., March 1994.

Easwaran:1994:MHI

[Eas94] Srividhya Easwaran. Mobile hosts in an Internet environment: an applicationlevel approach. Thesis (M.S.), University of North Carolina at Charlotte, Charlotte, NC, USA, 1994. vii + 113 pp.

Easton:1995:SIB

[Eas95] Jaclyn Easton. Shopping on the Internet and beyond! Coriolis Group Books, Scottsdale, AZ, USA, 1995. ISBN 1-883577-33-0 (paperback). xviii + 382 pp. LCCN ????

Ellsworth:1997:IU

[EB^{+}97] Jill Ellsworth, Billy Barron, et al. The Internet unleashed. Sams.net Pub., Indianapolis, IN, USA, 1997. ISBN 1-57521-346-X. ???? pp. LCCN ????

Ebihara:1995:AAI

[Ebi95] Wataru Ebihara. An Asian American Internet guide. Technical report, APNet, Los Angeles, CA, USA, 1995. 48 pp.

Erbes:1994:EIB

[ECC94] William Erbes, Mary J. Cronin, and Paul Christy. Exploring Internet 3 business and commercial applications, 1994. 1 videocassette (118 min.) guide.

Eckel:1995:BUI

[Eck95] George Eckel. Building a UNIX Internet server. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-494-5. xvii + 325 pp. LCCN QA76.76.O63 E235 1995.

Emtage:1992:AED

[ED92] Alan Emtage and Peter Deutsch. archie — an electronic directory service for the Internet. Technical report, McGill University, Montréal, Québec, Canada, 1992.

Engst:1994:IEK

[ED94] Adam C. Engst and William Dickson. Internet Explorer’s Kit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-089-1 (paperback). xvi + 446 pp. LCCN TK5105.875.I57 E53 1994. US$34.50.

Eddings:1994:HIW

[Edd94a] Joshua Eddings. How the Internet Works. How it works series How it works series (Emeryville, Calif.). Ziff-Davis Press, Emeryville, CA, USA, 1994. ISBN 1-56276-192-7. xi + 218 pp. LCCN TK5105.875.I57 E33 1994. US$24.95. Probably for someone who loves an Internet guide with lots of colourful illustrations/pictures ….

Eddings:1994:IOU

[Edd94b] Joshua Eddings. Intonet un ottoke umjik inunga = How the Internet works. Kanamsa, Soul, Korea, 1-pan. edition, 1994. ISBN 89-7354-108-0. 211 pp. LCCN ????

Eddy:1994:IAH

[Edd94c] Andy Eddy. Internet After Hours. Prima Publishing, Rocklin, CA, USA, 1994. ISBN 1-55958-513-7 (paperback). xxi + 378 pp. LCCN TK 5105.875 I57 E34 1994. US$19.95, CDN$27.95. About games and diversions on the Internet.

Eddings:1995:IYI

[Edd95a] Joshua Eddings. Irasuto de yomu Intanetto nyumon = How the Internet works. Inpuresu, Tokyo, Japan, dai 1han. edition, 1995. ISBN 4-8443-4673-3. viii + 181 pp. LCCN ????

Eddings:1995:TNW

[Edd95b] Joshua Eddings. Tien nao wang lo Internet ju men tu chieh = How the Internet works. Tsai se tien nao tsung shu; 3. Tien tzu kung yeh chu pan she, Wan li chi kou, wan li shu tien, 1995. ISBN 962-14-0925-X (paperback). 173 pp. LCCN ????

Eddy:1996:IAH

[Edd96] Andy Eddy. Internet after hours: second edition. Prima Publishing, Rocklin, CA, USA, second edition, 1996. ISBN 0-7615-0386-2. xviii + 382 pp. LCCN ????

Edmondson:1995:OAI

[Edm95] Dorothy M. Edmondson. Online access to the Internet: part of the Information Super Highway. D. M. Edmondson, ??, ??, USA, 1995. ISBN ???? various pp. LCCN TK5105.875.I57 E36 1995.

ERS:1995:EIU

[Edu95] ERS Internet update: a newsletter for subscribers to Educational Research Service, page various, 1995. Educational Research Service, Arlington, VA, USA.

Edwards:1995:INA

[Edw95] Margaret J. A. Edwards. The Internet for nurses and allied health professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94554-7 (softcover). xiii + 122 pp. LCCN W 26.5 E26i 1995.

Ellsworth:1994:IBB

[EE94] Jill H. Ellsworth and Matthew V.

Ellsworth. The Internet Business Book. Wiley-Intersci-ence, New York, NY, USA, 1994. ISBN 0-471-05809-2 (paperback). xxiv + 376 pp. LCCN HD30.335 .E44 1994. US$22.95. Precisely written full of substantial information motivates the business reader to begin work on the Internet. A very practical guide, offering exceptionally solid business advice to the reader wanting to get on the Internet.

Ellsworth:1995:IBK

[EE95a] Jill H. Ellsworth and Matthew V.

Ellsworth. The Internet business kit, 1995. ISBN 0-471-13327-2. Includes CD-ROM.

Ellsworth:1995:MIM

[EE95b] Jill H. Ellsworth and Matthew V. Ellsworth. Marketing on the Internet: multimedia strategies for the World Wide Web. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11850-8. xxvii + 404 pp. LCCN HF6146.I58 E44 1995.

Erbes:1995:EIW

[EE95c] William Erbes and William Eager. Exploring Internet 4 World Wide Web and Mosaic, 1995. 1 videocassette (120 min.) guide.

Ellsworth:1996:NIB

[EE96] Jill H. Ellsworth and Matthew V. Ellsworth. The new Internet business book. John Wiley, New York, NY, USA, 1996. ISBN 0-471-14160-7. xxviii + 499 pp. LCCN HD30.37 .E44 1996. Rev. ed of: The Internet business book / Jill H. Ellsworth, Matthew V. Ellsworth. c1994.

Easwaran:1999:YMS

[EE99] Sunny Easwaran and Mary Easwaran. The Year 2000 Mainframe Survival Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010481-7. xvii + 562 pp. LCCN QA76.76.S64E27

                                        1999. URL http://www.
                                        phptr.com/year2000;   http:

//www.y2kinfo.com.  In-

cludes CD-ROM.

Eberman:1999:IMI

[EFIJ99] B. Eberman, B. Fidler, R. Iannucci, and C. Joerg. Indexing multimedia for the Internet. Lecture Notes in Computer Science, 1614:195– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmi:1994:CDE

[EGC^{+}94] Michael J. Emmi, Stephen P.

Gardner, Nancy M. Cline, Louise Velazquez, Michael Olenski, and Ted Schilowitz. CAUSE for discussion: an exploration of the impact of the Internet on higher education, 1994. 1 videocassette (53 min.).

EPRI:1995:CSM

[EH95] Electric Power Research Institute and Hummel Consulting. Customer service and marketing on the Internet. Technical Report EPRI TR-105664, Electric Power Research Institute, Palo Alto, CA, USA, November 1995. various pp.

Ehn:1994:ORC

[Ehn94]  Hope Ehn.  On-line re-

sources for classical and academic musicians: a guide through the wilds of the Internet: newsgroups, mailing lists, and other resources for early music, classical music, musicology, music theory, and ethnomusicology: with a guide to basic list-server commands for subscribing to mailing lists and getting files by e-mail. Technical report, H. Ehn, Newton Centre, MA, USA, 1994. i-13 + ii-14 + iii-11 + 2 pp.

Eland:1995:IDL

[EHPT95] Thomas W. Eland, Virginia M. Heinrich, Tim Ponder, and Sarah Tourville. Internet directory of literacy and adult education resources. Technical report, Minnesota/South Dakota Regional Adult Literacy Resource Center, Saint Paul, MN, USA, June 1995. 144 pp.

Ewing:1992:MSI

[EHS92] David J. Ewing, Richard S. Hall, and Michael F. Schwartz. A measurement study of Internet file transfer traffic. Technical report CU-CS-571-92, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, January 1992. 13 pp.

Anonymous:1995:JEI

[EI 95] EI journal: enterprise Internetworking, page various, 1995. ISSN 1082-8761; 1068-1396. Cardinal Business Media, Dallas, TX, USA.

Eisen:1995:CLI

[Eis95] Lewis S. Eisen. The Canadian lawyer’s Internet guide.

Pixley Press, Toronto, ON, Canada, 1995. ISBN 0-9680154-0-9. xviii + 186 pp. LCCN ????

EFF:1994:BDGa

[EKG94a] Electronic Frontier Foundation, Mitchell Kapor, and Adam Gaffin. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. 156 pp.

EFF:1994:BDGb

[EKG94b] Electronic Frontier Foundation, Mitchell Kapor, and Adam Gaffin. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. various pp.

EsebbagBenchimol:1995:Ia

[EL95] Carlos Esebbag Benchimol and Juan Lloret Verdugo. Internet. Guia de iniciacion. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-763-5. 198 pp. LCCN ????

Eland:1994:IDL

[Ela94] Thomas Eland. Internet directory of literacy and adult education resources. Technical report, Minnesota/South Dakota Regional Adult Literacy Resource Center, Saint Paul, MN, USA, 1994. v + 40 pp.

Eigen:1999:EPW

[ELB99] Brad J. Eigen, Dan Livingston, and Micah Brown. Essential Photoshop 5 for Web Professionals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-012833-3. ???? pp. LCCN T385.E395

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130128333.html.

Anonymous:1991:ENR

[Ele91] Electronic networking: research, applications, and policy, 1991. ISSN 1051-4805. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA.

Ellsworth:1994:EI

[Ell94] Jill H. Ellsworth. Education on the Internet. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30595-X. xxvi + 591 pp. LCCN TK5105.875.I57 E45 1994. US$25. Computer Literacy Bookshops says: Packed with ideas resources projects and advice. Teachers and professors will find out how to develop curriculum plans, share teaching ideas and techniques and to start sister school programs. Students and parents will discover research topics, correspond with pen pals, learn and improve foreign language skills and explore museums of art. Sections for K-12, elementary, high school, college and graduate levels.

EBG:1995:EBG

[Ell95] Elliott Brown Gallery. Elliott Brown Gallery, 1995.

Ellis:1999:PPF

[Ell99] Bob Ellis. Public policy:

Focus on “last-mile” bandwidth continues. Computer Graphics, 33(3):29– ??, August 1999. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Engst:1995:ISKc

[ELO95] Adam C. Engst, Corwin S.

Low, and Stanley K. Orchard. Internet starter kit for Windows 95. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-260-6. xxv + 514 pp. LCCN ????

Engst:1995:ISKb

[ELS95] Adam C. Engst, Corwin S.

Low, and Michael A. Simon. Internet starter kit for Windows. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, second edition, 1995. ISBN 1-56830-177-4 (paperback). xxvi + 646 pp. LCCN TK5105.875.I57E532 1995.

EsebbagBenchimol:1995:Ib

[EM95] Carlos Esebbag Benchimol and Julian Martinez Valero. Internet. Guias practicas.

Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-785-6. 345 pp. LCCN ????

ES:1994:IAI

[Eme94] Emerald Software. Internet access to the information highway, 1994. Includes CD-ROM.

ES:1995:IAI

[Eme95a] Emerald Software. Internet access to the information highway for windows, 1995. Includes CD-ROM.

Emery:1995:IAU

[Eme95b]  Margot Emery.  Internet

adoption and use by innovator and early adopter communicators in Tennessee. Thesis (M.S.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1995. vii + 130 pp.

Emery:1995:HGY

[Eme95c] Vince Emery. How to grow your business on the Internet. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-883577-29-2. xxii + 490 pp. LCCN TK5105.875.I57 E45 1995.

Endicott:1994:ETD

[End94] Marcus L. Endicott. The electronic traveler: Directory of tourism information sources. Technical report, ????, ????, 1994. US$50.00. From the publisher’s info: As the title of this new book says, The Electronic Traveler is a directory of tourism information available to everyone with a computer and a telephone: it is a guidebook to online travel information sources. It explains in plain English exactly what travel and tourism information is available on the information highway and how to access it. It covers the Internet, popular proprietary interactive systems, Computer Reservation Systems (CRSs), independent bulletin board systems (BBSs), and commercial fulltext databases.

Engels:1993:CSI

[Eng93a] William R. Engels. Contributing software to the Internet: the Amplify program. Trends in biochemical sciences, 18(11):448–??, November 1, 1993. CODEN TBSCDB, TBSRDM. ISSN 0968-0004, 0167-7640, 0376-5067.

Engle:1993:ICL

[Eng93b] Mary Engle. Internet connections: a librarian’s guide to dial-up access and use. LITA monographs; 3. Library and Information Technology Association, 50 East Huron Street, Chicago, IL 60602, USA, 1993. ISBN 0-8389-7677-8. x + 166 pp. LCCN Z674.82.I59I58 1993.

Engst:1993:ISK

[Eng93c]  Adam Engst.  Internet

Starter Kit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1993. ISBN 1-56830-064-6. xxiii + 641 pp. LCCN TK5105.875.I57 E6 1993. US$29.95. System requirements for computer disk: Macintosh PC; about 25M of hard drive space for software archived in FTP site

                                        ftp.tidbits.com; modem.

“The software in accompanying disk is compressed (using StuffIt Deluxe) so that it can fit on a single 800K Macintosh disk.” From Samuel Ko: For more experienced MAC users …. Software for setting up Internet connection via SLIP/PPP (like Mac-TCP) included ….

Engst:1994:ISKa

[Eng94a]  Adam Engst.  Internet

Starter Kit for Windows. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-094-8. xxi + 608 pp. LCCN TK5105.875.I57E532. US$34.50.

Engst:1994:ISKb

[Eng94b]  Adam C. Engst.  Inter-

net starter kit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, second edition, 1994. ISBN 1-56830-111-1 (paperback). xxviii + 990 pp. LCCN TK5105.875.I57E5317 1994. US$29.95, CDN$39.99.

Engle:1995:ICL

[Eng95a] Mary Engle. Internet connections: a librarian’s guide to dial-up access and use. LITA monographs; 6. American Library Association, Chicago, IL, USA, second edition, 1995. ISBN 0-8389-7793-6. xi + 268 pp. LCCN Z674.82.I59 I58 1995.

Engst:1995:ISKa

[Eng95b]  Adam C. Engst.  Inter-

net starter kit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, third edition, 1995. ISBN 1-56830-197-9 (paperback). xxviii + 713 pp. LCCN TK5105.875.I57 E5317 1995. US$35.00. System requirements for computer disk: Macintosh computer; System 7.0 or higher; 1.4 MB disk drive; modem. Disk includes: MacTCP 2.0.6, MacPPP 2.0.1, InterSLIP 1.0.1, Anarchie 1.5, Eudora 1.5.1, MacWeb 1.00A3.2, Internet Config 1.1; StuffIt Expander 3.5.2.

Engst:1995:SAI

[Eng95c] Adam C. Engst. Simply amazing Internet for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-230-4 (paperback). xxii + 249 pp. LCCN ????

Engelman:1996:FIN

[Eng96a] Linda J. Engelman. Found it on the net: Strut your stuff. Internet World, 7(1):

97–??, January 1, 1996. CO-DEN IERNE8. ISSN 1064-3923.

Engelman:1996:II

[Eng96b] Linda J. Engelman. Interacting on the Internet. Irwin, Chicago, IL, USA, 1996. ISBN 0-256-20353-9. x + 134 pp. LCCN TK5105.875.I57E48 1996.

Ensman:1995:BNI

[Ens95] Richard G. Ensman. The basics of navigating the Internet. Dance teacher now, 17(1):47–50, 52–53, 57, January 1995. ISSN 0199-1795. Raleigh, NC, USA.

ET:1995:IIW

[Ent95] Entrepreneur Television. An introduction to the Internet what to know and how to get on, 1995. 1 videocassette (45 min.).

Eich:1999:IBD

[EO99] H. P. Eich and C. Ohmann. Internet-based decision-support server for acute abdominal pain. Lecture Notes in Computer Science, 1620:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elofson:1998:CCM

[ER98] Greg Elofson and William N. Robinson. Creating a custom mass-production channel on the Internet. Communications of the ACM, 41 (3):56–62, March 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-3/
p56-elofson/.

Ertel:1995:BNW

[Ert95] Monica Ertel. Brave new world. Searcher: the magazine for database professionals, 3(3):28–??, March 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Escamilla:1998:IDN

[Esc98] Terry Escamilla. Intrusion detection: network security beyond the firewall. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-29000-9. xx + 348 pp. LCCN TK5105.59 .E83 1998.

Essex:1995:IKC

[Ess95] Christopher Essex. The Internet and the K–12 classroom. Hot topic guide; 58. EDINFO Press, Bloomington, IN, USA, 1995. ISBN ???? various pp. LCCN ????

Estrada:CI93

[Est93a] Susan Estrada. Connecting to the Internet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK 5105.875 I57 E82c 1993. US$15.95.

Estrada:1993:CI

[Est93b] Susan Estrada. Connecting to the Internet: a Buyer’s Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK 5105.875 I57 E82c 1993. US$15.95. This small book focuses on choosing the best type of network connection for your personal, school or business needs, and how to get the best price for the type of access you require. Explains the differences between SLIP, PPP, ISDN, X.25 and other options. Includes an extensive list of Internet service providers. This is a single- purpose book, telling how to choose a connection and get online. It doesn’t try to teach you how to use the ’net once you’re there. That is graceful in its simplicity.

Estrada:1993:CIB

[Est93c] Susan Estrada. Connecting to the Internet: a buyer’s guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK5105.875.I57E84 1993.

Estrin:1993:PRE

[EST93d] Deborah Estrin, Martha Steenstrup, and Gene Tsudik. A protocol for route establishment and packet forwarding across multidomain internets. IEEE/ACM Transactions on Networking, 1 (1):56–70, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-1/p56-estrin/.

Estrin:1991:SCT

[ET91] D. Estrin and G. Tsudik. Secure control of transit internetwork traffic. Computer Networks and ISDN Systems, 22(5):363–??, 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

EMWAC:1995:WWW

[EU95] European Microsoft Windows NT Academic Centre and University of Edinburgh. Computing Services. World Wide Web sampler, 1995. Includes CD-ROM.

Emery:1996:FBS

[EV96] Vince Emery and Patrick Vincent. Free business stuff from the Internet. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-883577-82-9 (paperback). xvi + 495 pp. LCCN ????

Evanoff:1995:BGI

[Eva95] Joanne Evanoff. Beginner’s guide to the Internet. Technical report, Chester Fritz Library, Grand Forks, ND, USA, 1995. 27 pp.

EG:1995:HWW

[Eve95] Heartland: western Wisconsin’s Internet magazine, page various, 1995. Everest Group, Eau Claire, WI, USA.

Etzioni:1994:SBI

[EW94] Oren Etzioni and Daniel Weld. A softbot-based interface to the Internet. Communications of the ACM, 37 (7):72–76, July 1994. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        176797.html.

Ewers:1995:EIU

[Ewe95] Andy Ewers. The European Internet: an update. British Library R and D report 6198, British Library Research and Development Department, London, UK, 1995. iii + 24 pp.

Frey:1989:DEM

[FA89] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1989. ISBN 0-937175-39-0. xv + 284 pp. LCCN HE6239.E54 F73 1989. US$26.95. A quick desk reference to many networks, with two pages on each one.

Frey:1993:DEM

[FA93] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1993. ISBN 1-56592-031-7. xvi + 443 pp. LCCN HE6239.E54 F78 1993. US$24.95. This book is a reference work. It details the various computer networks with gateways to the Internet. It is common to cite such works as “indispensable”: in fact, most users, and even site managers, muddle along quite happily without it. Quick reference “electronic” versions exist of very similar documents, which provide the addressing schemes for the more common network and commercial service gateways. Frey and Adams have, however, put together a very complete and interesting reference, and I do suggest it to anyone managing, or using, extensive email correspondence. As a user of electronic mail, or the manager of a small Internet node or UUCP site, it would be hard to say that you “need” this book. If, however, you are at all interested in the topic of email, you will find this fascinating and useful.

Frey:1994:DEM

[FA94] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, June 1994. ISBN 1-56592-046-5. 640 pp. US$9.95. This book is a reference work. It details the various computer networks with gateways to the Internet. It is common to cite such works as “indispensable”: in fact, most users, and even site managers, muddle along quite happily without it. Quick reference “electronic” versions exist of very similar documents, which provide the addressing schemes for the more common network and commercial service gateways. Frey and Adams have, however, put together a very complete and interesting reference, and I do suggest it to anyone managing, or using, extensive e-mail correspondence. As a user of electronic mail, or the manager of a small Internet node or UUCP site, it would be hard to say that you “need” this book. If, however, you are at all interested in the topic of e-mail, you will find this fascinating and useful.

Fahey:1994:NSI

[Fah94] Tom Fahey. net.speak: The Internet Dictionary. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-095-6. xii + 203 pp. LCCN TK5105.875.I57 F34 1994. US$15.00.

Fain:1995:PTS

[Fai95a] Terry Fain. A pocket tour of sports on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1693-0 (paperback). xii + 211 pp. LCCN GV568.F35 1995.

Fain:1995:SI

[Fai95b] Terry Fain. Sports sur Internet. Pocket tour Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1632-1. xxii + 203 pp. LCCN ????

Falk:1994:IRa

[Fal94a] Bennett Falk. The Internet roadmap. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 0-7821-1365-6. xvii + 263 pp. LCCN TK5105.875.I57 F3 1994. US$12.99.

Falk:1994:IRb

[Fal94b] Bennett Falk. The Internet roadmap. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1994. ISBN 0-7821-1586-1. xviii + 325 pp. LCCN TK5105.875.I57 F35 1994b. US$14.99.

Falk:1996:IR

[Fal96] Bennett Falk. The Internet roadmap. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, third edition, 1996. ISBN 0-7821-1890-9. xvi + 320 pp. LCCN ????

Fang:1995:IPS

[Fan95] Nien-Hsuan Fang. The Internet as a public sphere: a Habermasian approach. Thesis (Ph.D.), State University of New York at Buffalo, Buffalo, NY, USA, 1995. 214 pp.

FARNET:1993:ILC

[FAR93a] FARNET, Denver, CO, USA. The Internet and libraries: current projects and issues, 1993. various pp.

Farrow:1993:KWY

[Far93b] Rik Farrow. Keeping the world out of your networks. UNIX/world, 10(12):56–??, December 1, 1993. ISSN 0739-5922.

Farber:1997:SFT

[Far97] David J. Farber. The science of future technology: Communications technology and its impact by 2010. Communications of the ACM, 40(2): 135–138, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.

acm.org/pubs/citations/
journals/cacm/1997-40-2/
p135-farber/.

Fraase:1995:MIT

[FAW95] Michael Fraase, Ted Alspach, and Jan Weingarten. The Macintosh Internet Tour Guide—Cruising the Internet the Easy Way. Ventana Press, Chapel Hill, NC, USA, second edition, 1995. ISBN 1-56604-173-2. xxii + 391 pp. LCCN TK5105.875.I57A44 1995. US$29.95.

Foster:1991:TDL

[FB91] David Foster and Emery Boose. Technology development in the LTER network: status report on geographic information systems, remote sensing, Internet connectivity, archival storage and global positioning systems. LTER publication 12, Long-Term Ecological Research Network Office, University of Washington, College of Forest Resources, Seattle, WA, USA, March 1991. 23 pp.

Flood:1995:ERI

[FB95] Michael Flood and John Button. Environmental resources on the Internet. Environmental briefing / Powerful Information; no. 6. 2 Environmental briefing (Powerful Information); no. 6. 2. Powerful Information, Milton Keynes, UK, 1995. ISBN 1-899950-05-2 (paperback). 27 pp. LCCN ????

Fromme:1998:CCM

[FBGP98] Michael Fromme, Bernd Böker, Lutz Grüneberg, and Helmut Pralle. A conference control management system for the Internet: Confman 2.0. Computer Networks and ISDN Systems, 30(16–18): 1457–1465, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/comnet/sub/1998/
                                        30/16-18/1991.pdf.

Freed:1994:BIH

[FD94] Les Freed and Frank Derfler. Building the Information Highway. Ziff-Davis Press, Emeryville, CA, USA, July 1994. ISBN 1-56276-126-9. 400 pp. LCCN TK5105.5 .F74 1994. US$24.95.

Fox:1995:HWP

[FD95] David Fox and Troy Downing. HTML Web publisher’s construction kit. Waite Group Press, Corte Madera, CA, USA, 1995. ISBN 1-57169-018-2. ???? pp. LCCN QA76.76.H94 F69 1995.

Farley:1991:LRI

[FE91] Laine Farley and Mary Engle. Library resources on the Internet: strategies for selection and use. Technical report, American Library Association, Chicago, IL, USA, 1991. 50 pp.

Farley:1992:LRI

[FE92] Laine Farley and Mary Engle. Library resources on the Internet: strategies for selection and use. RASD occasional papers; no. 12. American Library Association, Chicago, IL, USA, 1992. ISBN 0-8389-7576-3. 43 pp. LCCN Z687 .L54 1992. A project of Direct Patron Access to Computer-based Reference Systems Committee, Machine-assisted Reference Section, Reference and Adult Services Division, American Library Association.

Feeney:1993:IBa

[Fee93a] Ann Feeney. The Internet and beyond. Campus-wide information systems: CWIS, 10(3):38–??, May 1, 1993. CODEN CISYF9. ISSN 1065-0741.

Feeney:1993:IBb

[Fee93b] Ann Feeney. The Internet and beyond. Campus-wide information systems: CWIS, 10(5 / 6):36–??, September 1, 1993. CODEN CISYF9. ISSN 1065-0741.

Feeney:1993:IAS

[Fee93c] Ann Feeney. Internet applications: Stumpers-L. Computers in Libraries, 13(5):40– ??, May 1, 1993. CODEN CPLIE8. ISSN 1041-7915.

Fehrmann:1993:IRP

[Feh93] Paul Fehrmann. Internet resources for psychology. College and Research Libraries News, 54(9):510–??, October 1, 1993. ISSN 0099-0086.

Feit:1993:TIA

[Fei93] Sidnie Feit. TCP/IP: Architecture, Protocols and Implementation. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-020346-6. xxiii + 466 pp. LCCN TK5105.5 .F423 1993. US$44.95. Covers protocols plus additional services and products: NFS NIS, BIND, ARP, RIP, KER-BEROS, SNMP, etc. Discusses hoe to invoke network services, plan name/address structure, troubleshoot, connect via bridges and routers.

Feit:1994:AI

[Fei94] Sidnie Feit. Accessing the Internet, 1994. 1 videocassette (2 hr.) and 1 computer disk.

Feit:1996:TIA

[Fei96] Sidnie Feit. TCP/IP: Architecture, Protocols and Implementation. McGraw-Hill, New York, NY, USA, second edition, 1996. ISBN 0-07-021389-5. ???? pp. LCCN TK5105.585 .F45 1996. Covers protocols plus additional services and products: NFS NIS, BIND, ARP, RIP, KER-BEROS, SNMP, etc. Discusses hoe to invoke network services, plan name/address structure, troubleshoot, connect via bridges and routers.

Feldman:1993:IAG

[Fel93] Susan E. Feldman. The Internet at a glance: basic commands for electronic mail, anonymous ftp, telnet, mailing lists, unix, vi. Technical report, Susan E. Feldman, Ithaca, NY, USA, 1993. 11 pp.

Feldman:1994:IG

[Fel94a] Susan E. Feldman. The Internet at a Glance. Datasearch, 1994. 10 pp. US$7.00. A collection of “cheatsheets” for the Internet. It’s very short (only 10 thin pages) but it covers the very basics for using the Internet and UNIX. Topics include Finding Resources on the Internet, Tools, electronic mail, anonymous FTP, telnet, mailing lists and newsgroups, basic UNIX commands and the vi editor.

Feldman:1994:IAG

[Fel94b] Susan E. Feldman. The Internet at a glance: finding information on the Internet: Gopher/Veronica, basic commands for electronic mail, anonymous FTP, Telnet, mailing lists, Unix, vi. Technical report, Datasearch, Ithaca, NY, USA, 1994. 9 pp.

Feldmeier:1994:IIP

[Fel94c] David Feldmeier. An introduction to the Internet protocol (IP), 1994. 1 videocassette (39 min.).

FTG:1995:IA

[Fer95] Ferguson-Taylor Group, Inc. Internet artresources, 1995.

Fithen:1994:CIR

[FF94] Katherine Fithen and Barbara Fraser. CERT incident response and the Internet. Communications of the ACM, 37(8):108–114, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        179721.html.

Floyd:1999:PUE

[FF99] Sally Floyd and Kevin Fall. Promoting the use of endto-end congestion control in the Internet. IEEE/ACM Transactions on Networking, 7(4):458–472, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1999-7-4/p458-floyd/.

Fraase:1994:IMKb

[FH94a] Michael Fraase and Harley Hahn. Internet membership kit, 1994. ISBN 1-56604-171-6, 1-56604-062-0 (The Mac Internet tour guide), 0-07-882023-5. Includes CD-ROM.

Fraase:1994:IMKc

[FH94b] Michael Fraase and Harley Hahn. Internet membership kit, 1994. ISBN 1-56604-172-4, 1-56604-081-7 (The Windows Internet tour guide), 0-07-882023-5. Includes CD-ROM.

Freeman:1995:TTI

[FH95] Elisabeth Freeman and Susanne Hupfer. TAP: Tapping Internet resources for women in computer science. Communications of the ACM, 38 (1):44, January 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fraase:1994:IMKa

[FHS94] Michael Fraase, Harley Hahn, and Rick Stout. Internet Membership Kit: access and tools! Your onramp to the information superhighway! Ventana Media, Chapel Hill, NC, USA, Macintosh version. edition, 1994. ISBN 1-56604-171-6. LCCN TK5105.875.I57F72 1994.

Francis:1999:IDH

[FHU99] Brian Francis, Alex Homer, and Chris Ullman. IE5 dynamic HTML programmer’s reference. Programmer’s reference. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-174-6. xiii + 551 pp. LCCN ????

Field:1993:ICU

[Fie93] Cynthia E. Field. Internet: Connecting to a universe of information. InCider, 11(7): 36–??, July 1, 1993. CODEN INCIDE. ISSN 0740-0101.

Finney:1995:AIA

[Fin95] Peter G. Finney. An anthology of Internet acceptable use policies. Technical report, National Association of Regional Media Centers, ????, 1995. various pp.

FC:1995:NI

[Fir95]  Firefox  Communications.

NOV*IX for Internet. Firefox, San Jose, CA, USA, second edition, 1995. various pp.

Fisher:1993:RIH

[Fis93] Sharon Fisher. Riding the Internet Highway. New Riders Publishing, Carmel, IN, USA, 1993. ISBN 1-56205-192-X. xii + 266 pp. LCCN TK5105.875.I57 F57 1993. US$16.95.

Fisher:1995:SW

[Fis95a] Yuval Fisher. Spinning the Web. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94539-3. 400 pp. LCCN ???? US$29.00.

Fister:1995:PTM

[Fis95b] Mark Fister. A pocket tour of money on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1696-5. x + 165 pp. LCCN HG4515.5.F57 1995.

Fraase:1995:WIT

[FJ95] Michael Fraase and Phil James. The Windows Internet Tour Guide—Cruising the Internet the Easy Way. Ventana Press, Chapel Hill, NC, USA, second edition, 1995. ISBN 1-56604-174-0 (paperback). xxii + 380 pp. LCCN TK5105.875.I57F76 1995. US$29.95.

Feldman:1995:IAG

[FK95a] Susan E. Feldman and Larry Krumenaker. The Internet at a glance: finding information on the Internet: World Wide Web, Gopher, and Veronica: basic commands for electronic mail, Usenet newsgroups, anonymous FTP, telnet, mailing lists, Unix, and other Internet activities. Information Today, Medford, NJ, USA, third edition, 1995. ISBN 1-57387-024-2. 24 pp. LCCN ????

Fung:1995:DPS

[FK95b] Anthony Y. H. Fung and Kent D. Kedl. The degeneration of the public sphere: a case study of a Chinese Internet newsgroup. Technical report, ????, ????, 1995. 20 pp.

Feng:1999:APM

[FK99] Wu-Chang Feng and Dilip D.

Kandlur. Adaptive packet marking for maintaining end-to-end throughput in a differentiated-services Internet. IEEE/ACM Transactions on Networking, 7 (5):685–697, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1999-7-5/p685-feng/.

Frazier:1995:IK

[FKA95] Deneen Frazier, Barbara L. Kurshan, and Sara Armstrong. Internet for kids. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1741-4. xxix + 314 pp. LCCN TK5105.875 .I57 F77 1995. US$22.99.

Forschler:1995:IEI

[FL95a] Scott Forschler and Leonard H. Axe Library. Information efficiency and the Internet. Technical report, Leonard H. Axe Library, Pittsburg State University, Pittsburg, KS, USA, March 1995. 4 pp.

Forschler:1995:II

[FL95b] Scott Forschler and Leonard H. Axe Library. Introduction to the Internet. Technical report, Pittsburg State University, Leonard H. Axe Library, Pittsburg, KS, USA, March 1995. 4 pp.

Flanagan:1990:GTN

[Fla90] William A. Flanagan. The Guide to T1 Networking. Telecom Library Inc., 12 West 21 Street, New York NY 10010, USA, Tel: +1 212 691 8215, +1 800 LIBRARY, FAX: +1 212 691 1191, fourth edition, 1990. ISBN 0-936648-26-0. US$29.95. The T-1 trunk line technology is a standard for voice and data transmission in the US and Canada. The author introduces the subject in clear simple language and explains the various issues and terminology. Most terms are clearly explained as they are introduced; no prior understanding of telephone technology is required. The higher-speed technologies and the corresponding global standards are discussed briefly. Reference material is well organized. I recommend this book as an introduction to the subject. Table of Contents 1 What T1 Does For a Network 2 T-1 Defined Historically 3 Digital Voice 4 T-1 Circuits 5 What Equipment You Will Need At Your Office 6 How To Build a T-1 Network 7 Beyond T-1: Desktop to DS-3 8 Network Management And Control A T-1 Case Examples B Networking Acronyms C Basic Multiplexing D Global Framing Formats Index.

Flanagan:1991:FPC

[Fla91] William A. Flanagan. Frames, Packets and Cells in Broadband Networking. Telecom Library Inc., 12 West 21 Street, New York NY 10010, USA, Tel: +1 212 691 8215, +1 800 LIBRARY, FAX: +1 212 691 1191, 1991. ISBN 0-03-664831-7 (??checksum error??). xi + 269 pp. The author takes on the task of teaching the current state of the new data communication technologies, and does a generally excellent job of it. Chapter 1 introduces the basic concepts of packet networking, explaining that technology in detail and drawing contrasts to the T-1 TDM technology explained in The Guide to T-1 Networking by the same author. ASCII and X.25 are referred to in this chapter, with only the briefest description, so the reader should have some rudimentary knowledge of data communication. However the basic differences between packet switching and circuit switching, and between connectionoriented and connectionless transmission, are clearly explained so the reader can skip over confusing terminology without loss. Chapter 2 does an excellent job of explaining the basic concepts underlying the various types of packet, frame and cell technologies, the relative advantages and disadvantages of each, layered protocols, the OSI reference model and says a few words about the functions of SNA, the Internet protocols, the IEEE LAN standards and Signaling System 7. Chapter 3 uses T-1 as an example to explain the SONET/STS and SDH framing technology. The reader should approach this chapter with a good understanding of T-1 synchronization. HDLC is also described briefly in this chapter. Chapter 4 explains frame relay, including the Local Management Interface, and again makes a comparison to X.25 (which is never described in detail). This is followed by an detailed explanation of the basics of FDDI (but SMT isn’t covered). The MAN standards DQDB, IEEE 802.6 and SMDS, which are all closely related, are explained as a unit. ATM is described, with its relationship (or lack thereof) to SONET. The StrataCom proprietary FastPacket protocol is mentioned briefly, followed by some general design principles for light-weight protocols, with brief descriptions of XTP and HSTP. If you have the necessary background to read it, as explained above, I can recommend this book as an introduction to the new networking technologies that it covers. Table of Contents 1 Packet Overview 2 What All Packets Have In Common 3 Broadband Framing 4 Emerging Packet Technologies A Networking Acronyms B Annotated Bibliography C Time Division Multiplexing Index.

Fleck:1994:WLC

[Fle94] Kathleen Fleck. The world’s largest computer network: The Internet. Agricultural education magazine, 67(2):9–??, August 1, 1994. ISSN 0732-4677.

FMI:1995:IB

[Flo95] Florida Marketing International. The Internet and beyond, 1995. 1 videocassette.

Flynn:1995:WHH

[Fly95] Peter Flynn. The World-WideWeb Handbook: An HTML Guide for Users, Authors and Publishers. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-205-8. xix + 351 pp. LCCN TK5105.888 .F56 1995. Foreword by Tim Berners-Lee. Three sections deal with (1) Getting connected to the Internet and using Internet software; (2) Writing HTML (2.0) files for the World-WideWeb; (3) Running a HTTP server and providing a Web service. Author is a member of the IETF Working Group on HTML. Text includes additional material on SGML; choice of editors, browsers and servers; copyright and intellectual property; and advance details of HTML3.

Flynt:1999:TTR

[Fly99]  Clif Flynt.  Tcl/Tk for

Real Programmers. AP Professional, Boston, MA, USA, 1999. ISBN 0-12-261205-1. 656 pp. LCCN QA76.73.T44F59 1999. US$49.95. Includes CD-ROM.

Freeman:1995:TMI

[FMB95] Deb Freeman, Casey Monahan, and Dave Brichler. Texas music Internet directory, 1995.

Friedrich:1998:GID

[FMS98] M. Friedrich, M. Melle, and D. Saupe. Graphics in/ for digital libraries — AT-LAS2000 — atlases of the future on the Internet. Computers and Graphics, 22(6): 697–701, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/6/624.pdf.

FP:1995:CIB

[Foc95] Focused Presence. Comfind Internet business directory,

Folmsbee:1995:LRE

[Fol95a] Mark Folmsbee. Internet publishing law related electronic publishing on the World Wide Web (WWW); gateways to nationwide electronic resource centers/ institutes, 1995. 1 sound cassette.

Folmsbee:1995:LPA

[Fol95b] Mark Folmsbee. Listserv potpourri (A) how to be a contented listserv user; (B) listserv “netiquette”; (C) privacy and security concerns on listservs; (D) slander law as it applies to the Internet, 1995. 1 sound cassette.

Fontana:1994:MIC

[Fon94] Matteo Fontana. Making the Internet connection at sallie mae. Technical Report IM 599, Marymount University, Tarrytown, NJ 10591, USA, 1994. 52 pp. Final Project (MSIM).

Farnet:1994:RHW

[For94] Farnet, Inc and Fifty one reasons. 51 reasons: how we use the Internet and what it says about the information superhighway. Technical report, FARNET Inc., Lexington, MA, USA, 1994. 124 pp.

Ford:1995:SWH

[For95]  Andrew Ford.  Spinning

the Web: How to Provide Information on the Internet. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01996-3 (Van Nostrand Reinhold), 1-85032-141-8 (ITP UK). xviii + 227 + 8 pp. LCCN TK5105.888 .F67 1995.

Foster:1991:IPT

[Fos91] Steven L. Foster. Internet packet telephony: a NeXT case study. Thesis (M.S.), University of Nevada, Reno, Reno, NV, USA, November 1991. iii + 78 pp.

Fournier:1998:MCS

[Fou98] Roger Fournier. A Methodology for Client/Server and Web Application Development. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-598426-2. xxi + 648 pp. LCCN QA76.9.C55F68

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0135984262.html.

Fowler:1999:VPN

[Fow99] Dennis Fowler. Virtual Private Networks: Making the Right Connection. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-575-4. xviii + 222 pp. LCCN TK5105.875.E87 F69 1999. US$38.95.

Fahey:1994:NID

[FP94] Tom Fahey and Ruffin Prevost. Net.speak: the Internet dictionary. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-095-6 (paperback). xii + 203 pp. LCCN TK5105.875.I57F34 1994.

Froseth:1995:IAU

[FP95a] Stan Froseth and Barbara B. Poppe. Internet activities using scientific data: a selfguided exploration. Technical report, Dept. of Commerce, National Oceanic and Atmospheric Administration, Office of Educational Affairs, Space Environment Laboratory; For sale by the Supt. of Docs., U.S. G.P.O., Boulder, CO, USA, 1995. iv + 124 pp.

Fry:1995:HPI

[FP95b] Andrew Fry and David W. Paul. How to publish on the Internet: a comprehensive step-by-step guide to creative expression on the World Wide Web. Warner Books, New York, NY, USA, 1995. ISBN 0-446-67179-7 (trade paperback). xii + 275 pp. LCCN TK5105.888.F79 1995. Accompanying disk contains SPRY Mosaic PC software.

Fahey:1995:DI

[FPAR95] Tom Fahey, Ruffin Prevost, Hugo A. Acuna Soto, tr, and Raymundo Hugo Rangel, rev. Diccionario de Internet. Prentice-Hall Hispanoamericana, Mexico, DF, Mexico, 1995. ISBN 968-880-512-2. 220 pp. LCCN ????

Franco:1995:AFC

[FPV95] V. Franco, R. Pirto, and N. K. Vidal. Anatomy of a flame: Conflict and community building on the Internet. IEEE technology & society magazine, 14(2):12–??, Summer 1995. ISSN 0278-0097.

Fraase:1993:ITG

[Fra93a] Michael Fraase. The Internet tour guide cruising the Internet the easy way, 1993. 1 computer disk.

Fraase:1993:MIT

[Fra93b] Michael Fraase. The Mac Internet Tour Guide—Cruising the Internet the Easy Way. Ventana Press, Chapel Hill, NC, USA, 1993. ISBN 1-56604-062-0. xxii + 288 pp. LCCN TK5105.875.I57 F73 1993. US$27.95 (includes Macintosh diskette).

Francis:1993:NAD

[Fra93c]  Paul Francis.  A Near-

Term architecture for deploying pip. IEEE network, 7(3): 30–??, May 1, 1993. ISSN 0890-8044.

Fraase:1994:PIT

[Fra94a] Michael Fraase. The PC Internet tour guide: cruising the Internet the easy way. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-084-1. xxiv + 284 pp. LCCN TK5105.875.I57 F57 1994. US$24.95.

Fraase:1994:WIT

[Fra94b] Michael Fraase. The Windows Internet Tour Guide —Cruising the Internet the Easy Way. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-081-

7. xxiv + 344 pp. LCCN TK5105.875.I57 F76 1994. US$24.95. System requirements for computer disk (Chameleon sampler v. 3.11): IBM PC or compatible (at least 80386 processor); at least 4MB RAM; DOS 5.0 or later; Microsoft Windows 3.1 or later. Also requires Internet account and appropriate network connection (SLIP, PPP, or direct connection).

Francis:1994:CGP

[Fra94c]  Paul Francis.  Compar-

ison of geographical and provider-rooted Internet addressing. Computer Networks and ISDN Systems, 27 (3):437–448, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1994&volume=27&
issue=3&aid=1364.

Franks:1995:IPH

[Fra95] Mike Franks. The Internet Publishing Handbook: For World-Wide Web, Gopher, and WAIS. Addison-Wes-ley, Reading, MA, USA, 1995.

ISBN 0-201-48317-3 (paperback). xviii + 380 pp. LCCN TK5105.875.I57 F769 1995. US$22.95. URL http://www.

aw.com/devpress/titles/
48317.html.

Franklin:1996:VBI

[Fra96] Carl Franklin. Visual Basic 4.0 Internet programming. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-13420-1 (paperback). xiii + 480 pp. LCCN QA76.73.B3F72 1996. Includes CD-ROM.

Franklin:1999:FCT

[Fra99] Matthew Franklin, editor. Financial cryptography: Third International Conference, FC ’99, Anguilla, British West Indies, February 22– 25, 1999: proceedings, volume 1648 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66362-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1999.

Ford:1993:IRA

[FRB93] Peter S. Ford, Yakov Rekhter, and Hans-Werner Braun. Improving the routing and addressing of IP. IEEE network, 7(3):10–??, May 1, 1993. ISSN 0890-8044.

Frey:1993:RCB

[Fre93] A. H. Frey. The revolution in communication for biologists — the Internet. The FASEB journal: official publication of the Federation of American Societies for Experimental Biology, 7(13):1206–??, October 1, 1993. CODEN FAJOEC. ISSN 0892-6638.

FFMSC:1994:CCH

[Fre94] Freedom Forum Media Studies Center, New York, NY, USA. Communicators of conscience: humanitarian and human rights organizations’ use of the Internet, 1994. 33 pp.

Frederick:1995:CIA

[Fre95] Howard H. Frederick. La comunicacion internacional academica a traves de Internet: libro de curso: tecnologia educativa para el siglo XXI. Technical report, Emerson College, Boston, MA, USA, 1995. various pp.

Foster:1997:IPW

[FRG97] Will A. Foster, Anthony M. Rutkowski, and Seymour E. Goodman. International perspectives: Who governs the Internet? Communications of the ACM, 40(8): 15–20, August 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/cacm/1997-40-8/
p15-foster/.

Fristrup:1994:UNE

[Fri94] Jenny A. Fristrup. Usenet:

Netnews for Everyone. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-123167-7. xvii + 386 pp. LCCN QA76.76.O63 F79 1994.

Fry:1995:PNM

[Fry95] Andrew L. Fry. Publishing in the new mass medium: Creating content on the Internet. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

????  URL http://www.
ncsa.uiuc.edu/SDG/IT94/
Proceedings/Overviews/fry/

fry.html. Includes CD-ROM.

Fahy:1995:IHR

[FS95] Anne Fahy and Wendy Sudbury, editors. Information: the hidden resource: museums and the Internet: proceedings of the Seventh International Conference of the MDA, held in Edinburgh, Scotland, 6–7 November 1995. Museum Documentation Association, Cambridge, England, 1995. ISBN 0-905963-96-2 (paperback). LCCN ????

FS:1995:EIEa

[FTP95a] FTP Software, Andover, MA, USA. Exploring on the Internet: with Explore from FTP Software, September 2, 1995. vi + 128 pp.

FS:1995:EIEb

[FTP95b] FTP Software, Inc, Andover, MA, USA. Exploring on the Internet: with Explorer On-Net from FTP Software, 1995. 232 pp.

Fumy:1998:ISP

[Fum98] W. Fumy. Internet security protocols. Lecture Notes in Computer Science, 1528:186– 208, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuoco:1995:DEI

[Fuo95] Marguerite M. Fuoco. Developing an educational interface to the Internet. Thesis (M.S.), University of Alabama at Birmingham, Birmingham, AL, USA, 1995. vii + 92 pp.

Furht:1999:HIM

[Fur99]  Borko Furht.  Handbook

of Internet and Multimedia Systems and Applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-1858-0. xvi + 872 pp. LCCN QA76.575 .H353 1999. US$94.95.

Fisher:1994:RIH

[FVNT94] Sharon Fisher, Steven J. Vaughan-Nicholas, and Rob Tidrow. Riding the Internet highway. New Riders Publishing, Carmel, IN, USA, deluxe edition, 1994. ISBN 1-56205-315-9. xv + 381 pp. LCCN TK5105.875.I57F57 1994.

Graves:1995:GSI

[GA95] John Graves and Special Libraries Association. Getting started on the Internet. Special Libraries Association, Washington, DC, USA, 1995. ISBN 0-87111-444-5. ii + 33 + 10 pp. LCCN ????

Gaarslev:1976:NPP

[Gaa76] Axel Gaarslev, editor. Notes on project planning: six Danish papers at the Internet Congress 1976. Danish Project Management Society, Copenhagen, Denmark, 1976. ISBN 87-980516-0-1. 77 pp. LCCN ????

Gach:1996:PGI

[Gac96]  Gary Gach.  The pocket

guide to the Internet. Pocket Books, New York, NY, USA, 1996. ISBN 0-671-56850-7 (paperback). xviii + 347 pp. LCCN ????

Gackenbach:1998:PII

[Gac98] Jayne Gackenbach. Psychology and the Internet: Intrapersonal, Interpersonal, and Transpersonal Implications. Academic Press, New York, NY, USA, 1998. ISBN 0-12-271950-6. xix + 369 pp. LCCN BF637.C45 P79 1998. US$62.95.

Gadzikowski:1994:IW

[Gad94] Claire Gadzikowski. Internet workbook. Technical report, National Network of Libraries of Medicine, Omaha, Neb.?, February 1994. various pp.

Gaffin:1994:EGI

[Gaf94] Adam Gaffin. Everybody’s Guide to the Internet. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-57105-6. xii + 211 pp. LCCN TK5105.875.I57 G34 1994. US$14.95. This is basically a printed version of version 2.2 of the EFF’s online guide The Big Dummy’s Guide to the Internet, plus an index. Litigious lawyers from another publishing house prevented the printed version from sharing the same name.

Gagnon:1994:WIa

[Gag94a] Eric Gagnon. What’s on the Internet. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1994. ISBN 1-56609-162-4. 256 pp. US$19.95. This book is an index of what’s available on the Internet’s newsgroups. It doesn’t delve into any other Internet services. A more appropriate title for this book would be “What’s on Usenet”, because that’s just what it is about. the book is arranged by subject matter and reviews about 1,600 of the most popular, worldwide newsgroups, giving honest opinions about what’s on each, what’s good and bad about them, and how popular each one is.

Gagnon:1994:WIb

[Gag94b] Eric Gagnon. What’s on the Internet. Internet Media Corp., Fairfax, VA, USA, netcom special edition, 1994. ISBN ???? 260 pp. LCCN ????

Gagnon:1995:WW

[Gag95] Eric Gagnon. What’s on the Web. Internet Media Corp., Fairfax, VA, USA, fall/winter 1995/1996 edition, 1995. ISBN 1-884640-18-4. 462 pp. LCCN TK5105.888 .W45 1995.

Gainor:1993:UL

[Gai93] Larry Gainor. Usenet and the library. RSR: Reference Services Review, 21(3):7–??, 1993. ISSN 0090-7324.

Gallo:1994:AEH

[Gal94] Michael Anthony Gallo. Assessing the effect on high school teachers of direct and unrestricted access to the Internet: a case study of an east central Florida high school. Thesis (Ph. D. in Science Education), Florida Institute of Technology, Melbourne, FL, USA, 1994. xi + 218 pp.

GRI:1995:GGI

[Gal95a] Gale guide to Internet databases, 1995. ISSN 1081-2385. Gale Research Inc., Detroit, MI, USA.

Gallagher:1995:CGI

[Gal95b]  Jim Gallagher.  Chris-

tian Guide to the Internet. Wisdom Research, PO Box 292786, Lewisville, TX 75029, 1995. ISBN none listed. US$9.95. From the publisher: “Covers everything you need for step-by-step navigating on the ‘Information Superhighway’ as well as a complete directory of Christian resources and domains on the Internet.”.

GRI:1996:CGI

[Gal96] CyberHound’s guide to Internet databases, page various, 1996. ISSN 1081-2385. Gale Research, Detroit, MI, USA.

Godwin:1997:SAP

[GAM^{+}97] Paul Godwin, Eythor Arnolds, William L. Martens, Tim Cole, and James Grunke. Sounding off on audio (panel): the future of Internet sound. Computer Graphics, 31(Annual Conference Series):465–467, August 1997. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://

www.acm.org:80/pubs/citations/
proceedings/graph/258734/
p465-godwin/.

Gangloff:1995:RIT

[Gan95] Maralea K. Gangloff. The role of Internet on a teachercentered secondary curriculum. Master of Arts in Educational Processes, thesis (M.A.), Maryville University of Saint Louis, St. Louis, MO, USA, May 1995. v + 81 pp.

Gardner:1989:IWW

[Gar89] Phillip E. Gardner. The Internet worm: What was said and when. Computers and Security, 8(4):291–??, June 1, 1989. CODEN CPSEDU. ISSN 0167-4048.

Gartner:1993:SPS

[Gar93] Richard Gartner. SALBIN:

PC software for accessing Internet resources. Information Technology and Libraries, 12 (2):267–??, June 1, 1993. CO-DEN ITLBDC. ISSN 0730-9295.

Gardner:1994:UGI

[Gar94] James Gardner. A DOS User’s Guide to the Internet: E-mail, Netnews, and File Transfer with UUCP. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-106873-3. xi + 302 pp. LCCN TK5105.875.I57G37 1994. Includes IBM PC diskette.

Gardner:1995:IAA

[Gar95a] James Gardner. Internet Anywhere: All You Need to Test Drive the Internet For 30 Days. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-185612-X. xix + 444 pp.

LCCN TK5105.875.I57 I54 1995. US$29.95, CDN$39.95.

Garfinkel:1995:PPG

[Gar95b]  Simson Garfinkel.  PGP:

Pretty Good Privacy. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-098-8. xxxiii + 393 pp. LCCN QA76.9.A25G36 1995. US$24.95.

Garber:1997:NBC

[Gar97] Lee Garber. News briefs: Crucial compromise launches digital TV. US encryption agreement in jeopardy. warning issued about UNIX flaw. WIPO discusses cyberspace copyrights. IT issues could threaten European Monetary Union. COBOL programmers in demand again. chip alliance formed. semiconductor film grown in space. survey reveals security fears and vulnerability. taxing the Internet. Computer, 30(2):18, 19, 22, February 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaskin:1996:NIG

[Gas96] James E. Gaskin. Netware to Internet gateways. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-521774-1 (paperback). ???? pp. LCCN QA76.76.O63G388 1996.

Gauer:1994:UMW

[Gau94] Stephen Gauer. Using Mosaic for Windows. Electric Avenue Press, Vancouver, BC, Canada, 1994. ISBN 0-9698853-0-X. 104 pp. LCCN TK 5105.875 I57G29 1994. US$19.95. Includes diskette.

Gilbreth:1995:CSI

[GB95] William J. Gilbreth and David Bender. Computer software and the Internet. Patents, copyrights, trademarks, and literary property course handbook series; no. G-415. Practising Law Institute, New York, NY, USA, 1995. ISBN 0-87224-200-5. 392 pp. LCCN KF3024.C6 C659 1995. “Prepared for distribution at the Computer software and the Internet program, September 1995, New York City”–p. 5. “G4-3953.”.

Guensler:1996:TRI

[GB96] Randall Guensler and David Bernstein. Transportation resources on the Internet. Institute of Transportation Engineers journal, 66(4):42–48, April 1996. CODEN ITE-JDZ. ISSN 0162-8178.

Gonzalez:1994:CER

[GBJ94] Brenda Gonzalez, Sue Burke, and Linda C. Joseph. CPS early release Internet and Eisenhower National Clearinghouse, 1994. 1 videocassette (45 min.).

Gardner:1995:IWA

[GBS95] David C. Gardner, Grace Joely Beatty, and David A. Sauer. Internet for Windows: America online 2.5 edition. Visual learning guides. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 0-7615-0303-X. vi + 314 pp. LCCN QA76.57.A43 G37 1995.

Gravano:1997:SSP

[GCGMP97] Luis Gravano, Chen-Chuan K. Chang, Héctor García-Molina, and Andreas Paepcke. STARTS: Stanford proposal for Internet meta-searching. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26(2): 207–218, June 1997. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goodwin:1996:NUI

[GCS96] Mary Goodwin, Deborah Cohn, and Donna Spivey. Netjobs: use the Internet to land your dream job. Michael Wolff and Company, New York, NY, USA, 1996. ISBN 0-679-77032-1. xvi + 184 pp. LCCN ????

Gibbs:1995:CI

[GCW95] Mark Gibbs, Dean Crutchfield, and Jeffrey Weiss. Capitalizing on the Internet, 1995. 3 videocassettes (120 min. each).

Garcia:1998:SOR

[GD98] Roman Garcia and Jose Duato. Suboptimal-optimal routing for LAN internetworking using transparent bridges. International Journal of Foundations of Computer Science (IJFCS), 9(2): 139–??, June 1998. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Genka:1983:DPS

[Gen83] Jon K. T. Genka. A dial up packet switcher for an Internet gateway. Thesis (B.S.), M.I.T., Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1983. 74 pp. Supervised by David K. Gifford.

Genaway:1995:ORU

[Gen95] David C. Genaway. On-line resources for urban planners: what good is the Internet? Thesis (M.U.P.), State University of New York at Buffalo, Buffalo, NY, USA, 1995. ix + 69 pp.

Gerich:1991:EIG

[Ger91] E. Gerich. Expanding the Internet to a global environment but … how to get connected? Computer Networks and ISDN Systems, 23(1–3): 43–??, November 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gershon:1995:VIP

[GFF^{+}95] Nahum D. Gershon, Bran Ferren, James Foley, Joseph Hardin, Frank Kappe, and William A. Ruh. Visualizing the Internet (panel session): putting the user in the driver’s seat. Computer Graphics, 29(Annual Conference Series):492–494, November 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/graph/218380/
p492-gershon/.

Glossbrenner:1994:IST

[GG94] Alfred Glossbrenner and Emily Glossbrenner. Internet Slick Tricks. Random House Electronic Pub., New York, NY, USA, 1994. ISBN 0-679-75611-6 (paperback). xii + 271 pp. LCCN TK5105.875.I57 G59 1994. US$16.00. If you’ve been on the Internet forever (or at least for a couple months) and you think you know it all, pick up Internet Slick tricks. I bet you’ll learn something new. This slim guide will show you hundreds of tips and tricks on using the Internet. It’s readable and honest and full of fascinating facts on decoding graphic images, using Archie, finding info online and much more. One of my favorites.

Glossbrenner:1995:FJI

[GG95a] Alfred Glossbrenner and Emily Glossbrenner. Finding a job on the Internet. Mc-Graw, New York, NY, USA, 1995. ISBN 0-07-024055-8. xiv + 290 pp. LCCN HF5382.7 .G56 1995 Ref. US$16.95.

Glossbrenner:1995:MMI

[GG95b] Alfred Glossbrenner and Emily Glossbrenner. Making Money on the Internet. McGraw-Hill, New York, NY, USA, March 1995. ISBN 0-07-024049-3 (hardcover), 0-07-024050-7 (paperback). xiv + 306 pp. LCCN HD30.335 .G55 1995. US$29.95 (hardcover), US$19.95 (paperback). The publisher says: “Businesspeople get the straight scoop on how to take best advantage of the Internet to sell goods and services. You’ll see what other businesses are doing, and learn how to profit from their mistakes.”.

Glossbrenner:1996:ICS

[GG96a] Alfred Glossbrenner and Emily Glossbrenner. Internet 101: a college students guide. McGraw-Hill, New York, NY, USA, third edition, 1996. ISBN 0-07-024448-0 (paperback). various pp. LCCN TK5105.875.I57G58 1996.

Gonyea:1996:SIH

[GG96b] James C. Gonyea and Wayne M.

Gonyea. Selling on the Internet: how to open an electronic storefront and have millions of customers come to you! McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-024187-2 (paperback/ disk). xx + 224 pp. LCCN HD30.37.G66 1996.

Giroux:1998:QSA

[GG98] Natalie Giroux and Sudhakar Ganti. Quality of Service in ATM Networks: Stateof-the-Art Traffic Management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-095387-3. xii + 252 pp. LCCN TK5105.35.G58

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0130953873.html.

Goduin:1997:SAF

[GGA^{+}97] Paul Goduin, James Grunke, Eythor Arnolds, William L. Martens, and Tim Cole. Sounding off on audio: The future of Internet sound. Computer Graphics, 31(Annual Conference Series):465–467, August 1997. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Gabber:1999:SPC

[GGK^{+}99] Eran Gabber, Phillip B. Gibbons, David M. Kristol, Yossi Matias, and Alain Mayer. On secure and pseudonymous client-relationships with multiple servers. ACM Transactions on Information and System Security, 2(4):390– 415, November 1999. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tissec/1999-2-
4/p390-gabber/.

Gravano:1999:GTS

[GGMT99] Luis Gravano, Héctor García-Molina, and Anthony Tomasic. GlOSS: text-source discovery over the Internet. ACM Transactions on Database Systems, 24(2): 229–264, June 1999. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tods/1999-24-2/
p229-gravano/.

Gaffin:1993:BDGb

[GH93] Adam Gaffin and Jorg Heitkotter. Big dummy’s guide to the Internet: a round trip through global networks, life in cyberspace, and everything. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1993. x + 210 pp.

Gaffin:1994:BDGa

[GH94a] Adam Gaffin and Jorg Heitkotter. Big dummy’s guide to the Internet: a round trip through global networks, life in cyberspace, and everything. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. xi + 278 pp.

Gaffin:1994:BDGb

[GH94b] Adam Gaffin and Jorg Heitkotter. Big dummy’s guide to the Internet: a round trip through global networks, life in cyberspace, and everything. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. xii + 302 pp.

Gaffin:1994:EEG

[GH94c] Adam Gaffin and Jorg Heitkotter. EFF’s (extended) guide to the Internet: a round trip through global networks, life in cyberspace and everything. Technical report, Electronic Frontier Foundation, Washington, DC, USA (1001 G. St., N.W., Suite 950 East, Washington 20001), 1994. xii + 304 pp.

Gilbert:1994:UID

[GH94d] Gerald Gilbert and David H. Hsu. The use of the Internet in defense analysis. Technical report, Institute for Defense Analyses, Science and Technology Division, Alexandria, VA, USA, 1994. 163 pp.

Ghatak:1992:EMI

[Gha92] Subhendu Ghatak. Evaluation methods for Internet management schemes. Thesis (Ph.D.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1992. xiii + 169 pp.

Ghiselli:1989:II

[Ghi89] A. Ghiselli. Internetworking in INFNet. Computer Networks and ISDN Systems, 17(4–5):371–375, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gagnon:1995:WI

[GI95] Eric Gagnon and Internet Media Corporation. What’s on the Internet. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, summer/fall 1995. edition, 1995. ISBN 1-56609-184-5. 366 pp. LCCN ????

Gianini:1994:NCC

[Gia94]  Gabriele Gianini.  Nel

ciberspazio con Internet. Telematica. Apogeo, Milano, Italy, 1994. ISBN 88-7303-075-0. x + 237 pp. LCCN ????

Giagnocavo:1995:EICg

[Gia95a] Gregory Giagnocavo. Educator’s Internet companion, 1995. Includes CD-ROM.

Giagnocavo:1995:EICh

[Gia95b] Gregory Giagnocavo. The educator’s Internet companion classroom connect’s complete guide to educational resources on the Internet, 1995. Includes Windows and Macintosh versions on two floppy disks.

Giagnocavo:1995:EICe

[Gia95c] Gregory Giagnocavo. Educator’s Internet companion Mac version, 1995. Includes CD-ROM.

Giagnocavo:1995:EICf

[Gia95d] Gregory Giagnocavo. Educator’s Internet companion Windows version, 1995. Includes CD-ROM.

Giese:1994:TSR

[Gie94] Mark Giese. Taking the scenic route: the Internet as precursor to the information superhighway. Technical report, ????, ????, 1994. 22 pp. A paper written for the Communications Technology and Policy Division, Association for Education in Journalism and Mass Communication Annual Conference, Atlanta, Georgia, August 10–13, [1994].

Gilster:1993:INE

[Gil93a] Paul Gilster. The Internet Navigator — The Essential Guide to Network Exploration for the Individual Dial-up User. Wiley-Intersci-ence, New York, NY, USA, 1993. ISBN 0-471-59782-1. xxiv + 470 pp. LCCN TK5105.875.I57 G55 1993. US$24.95.

Gilster:1993:IN

[Gil93b] Paul Gilster. The Internet Navigator—The Essential Guide to Network Exploration for the Individual Dial-up User. Wiley-Intersci-ence, New York, NY, USA, 1993. ISBN 0-471-59782-1. xxiv + 470 pp. LCCN TK5105.875.I57 G55 1993. US$24.95.

Gilster:1994:FIE

[Gil94a] Paul Gilster. Finding it on the Internet: The Essential Guide to archie, Veronica, Gopher, WAIS, WWW, and Other Search Tools. Wiley-Interscience, New York, NY, USA, 1994. ISBN 0-471-03857-1. xviii + 302 pp. LCCN TK5105.875.I57 G53 1994. US$19.95.

Gilster:1994:FII

[Gil94b] Paul Gilster. Finding it on the Internet: the essential guide to archie, Veronica, Gopher, WAIS, WWW (including Mosaic), and other search and browsing tools. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1994. ISBN 0-471-03857-1 (paperback). xviii + 302 pp. LCCN TK5105.875 I57 G53 1994 Reserves. US$19.95.

Gilster:1994:IN

[Gil94c] Paul Gilster. The Internet Navigator—The Essential Guide to Network Exploration for the Individual Dialup User. Wiley-Interscience, New York, NY, USA, second edition, 1994. ISBN 0-471-05260-4. xxxii + 590 pp.

LCCN TK5105.875.I57 G55 1994.

Gilster:1994:MN

[Gil94d] Paul Gilster. The Mosaic Navigator. John Wiley, New York, NY, USA, 1994. ISBN 0-471-11336-0. x + 243 pp. LCCN TK5105.875.I57 G56 1995. US$16.95. While Gilster provides cogent instructions in obtaining Mosaic from the NSCA, using its tools to browse via telnet, FTP, and e-mail, customizing it for individual comfort, and lists the best WWW sites, at times he rises above the “beginner” level and leaves the reader confused and frustrated.

Gillett:1995:CHIa

[Gil95a] Sharon Eisner Gillett. Connecting homes to the Internet: an engineering cost model of cable vs. ISDN. Thesis (M.B.A.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1995. 282 pp.

Gillett:1995:CHIb

[Gil95b] Sharon Eisner Gillett. Connecting homes to the Internet: an engineering cost model of cable vs. ISDN. Technical Report MIT/LCS/TR-654, Massachusetts Institute of Technology, Cambridge, MA, USA, 1995. 197 pp.

Based on thesis (M.B.A.)– Massachusetts Institute of Technology, Sloan School of Management, 1995.

Gilster:1995:MNE

[Gil95c]  Paul Gilster.  The Mo-

saic navigator: the essential guide to Internet interface. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11336-0 (paperback). x + 243 pp. LCCN TK5105.875.I57 G56 1995. US$16.95.

Gilster:1995:NIN

[Gil95d] Paul Gilster. The new Internet Navigator. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-12694-2 (paperback). xxvii + 735 pp. LCCN TK 5105.875 I57 G55 1995.

Gilster:1995:SPC

[Gil95e] Paul Gilster. The SLIP/ PPP connection: the essential guide to graphical Internet access. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11712-9 (paperback). xvi + 442 pp. LCCN TK5105.875.I57 G57 1995.

Gilster:1996:FII

[Gil96] Paul Gilster. Finding it on the Internet: the Internet navigator’s guide to search tools and techniques. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, revised and expanded, 2nd edition, 1996. ISBN 0-471-12695-

0. xviii + 379 pp. LCCN TK5105.875.I57G54 1996.

Giovanetti:1995:GEL

[Gio95] Pierangelo Giovanetti. Il giornale elettronico: l’informazione nell’era di Internet. Il pensiero moderno. Comunicazione e immagine. Vallecchi, Firenze, Italy, 1995. ISBN 88-8252-047-1. 195 pp. LCCN ????

Ginsburg:1995:EPC

[GK95] Mark Ginsburg and Ajit Kambil. The EDGAR project: a case study in disseminating financial data on the Internet. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

                                        ????  URL http://www.
                                        ncsa.uiuc.edu/SDG/IT94/
                                        Proceedings/Fin/ginsburg/

edgar.html. Includes CD-

ROM.

Gilbert:1996:SLE

[GK96] Laurel Gilbert and Crystal Kile. Surfergrrrls: look Ethel! an Internet guide for us. Seal Press, Seattle, WA, USA, 1996. ISBN 1-878067-79-6. ???? pp. LCCN HQ1180.G55 1996.

Gaffin:1993:BDGa

[GKC93a] Adam Gaffin, Mitch Kapor, and Steve Cisler. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1993. 116 pp.

Gaffin:1993:BDGc

[GKC93b] Adam Gaffin, Mitch Kapor, and Steve Cisler. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1993. 116 pp.

Gaidosch:1999:GVP

[GKM^{+}99] Tamas Gaidosch, Charles Kunzinger, Martin Murhammer, Laura Rademacher, and Andreas Weinfurter. A Guide to Virtual Private Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-083964-7. 200 pp. LCCN TK5105.875.E87G85

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130839647.html.

Gresham-Lancaster:1994:CSM

[GL94]  Scot  Gresham-Lancaster.

Cover story: Magical musical tours. Electronic Musician, 10(10):46–??, October 1, 1994. ISSN 0884-4720.

Greenberg:1999:BPW

[GL99] Jeff Greenberg and J. R.

Lakeland. Building Professional Web Sites with the Right Tools. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-084317-2. ???? pp. LCCN TK5105.888.G72

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130843172.html.

Glazer:1994:ADI

[Gla94] Evan Michael Glazer. An analysis of discourse on an Internet-based listserv for mathematics educators. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1994. xii + 207 pp.

Glass:1999:CCL

[Gla99] Robert Glass. Computing Calamities: Lessons Learned From Products, Projects, and Companies that Failed. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-082862-9. xiv + 302 pp. LCCN HD9696.2.U62 G55 1999. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130828629.html.

Glossbrenner:1995:ICSa

[Glo95] Alfred Glossbrenner. Internet 101: a college student’s guide. McGraw-Hill, New York, NY, USA, second edition, 1995. ISBN 0-07-024207-0 (paperback), 0-07-024054-X. xviii + 334 (or xviii + 350??) pp. LCCN TK 5105.875 I57 G58 1995. US$19.95. Ad copy says: “Ideal for the MTV generation this book tells students how to use the Net for all their academic and personal needs.”.

Garcia-Luna:1989:IPH

[GLSW89] Jose Garcia-Luna, Mary K.

Stahl, and Carol A. Ward. Internet protocol handbook. Technical Report NIC 50007, Network Information Systems Center, SRI International, Menlo Park, CA (333 Ravenswood Ave., Room EJ291, Menlo Park, CA 94025), August 1989. v + 214 pp.

Giagnocavo:1994:AOI

[GM94a] Gregory Giagnocavo and Tim McLain. All-In-One Internet Kit. Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1994. ISBN 0-932577-01-6. 110 pp. LCCN ???? US$49. Ad says: Includes all of the software and instructions both Mac and IBM PC users need to logon to the Internet using a SLIP or PPP connection. This is the first book/disk kit offering all of the software (e-mail, ftp, gopher, telnet, newsreader) needed to access the Net plus a Web browser.

Godin:1994:IWP

[GM94b] Seth Godin and James S.

McBride. 1994 Internet White Pages. IDG Books, San Mateo, CA, USA, 1994. ISBN 1-56884-300-3. LCCN CT120 .G65 1994. US$29.95. A thick book listing about 100000 e-mail addresses. (Although that may seem like a lot, it’s not very comprehensive.) Addresses are listed by last name and are also indexed by Internet domain name. This book may be useful for finding associates’ email addresses, but is probably no more useful than using one of the many net-based email-address search utilities. The book is unevenly edited (I’m in there three times!) and was woefully out of date before the ink was dry.

Grisson:1998:CAA

[GMA^{+}98] Scott Grisson, Janet McAndless, Omar Ahmad, Christopher Stapleton, Adele Newton, Celia Pearce, Ryan Ulyate, and Rick Parent, editors. Conference abstracts and applications: SIGGRAPH 98, July 14– 21, 1998, Orlando, FL, Computer Graphics. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-046-5, 1-58113-048-1 (CD-ROM). LCCN T385 .S54

1998b. URL http://info.
acm.org/pubs/contents/proceedings/

graph/. ACM order number for CD-ROM: 435984. ACM order number: 435983.

Gunning:1993:NEI

[GMB93] Kathleen Gunning, Judy E. Myers, and Charles W. Bailey. Networked electronic information systems at the University of Houston libraries:

The IRIS project and beyond. Library Hi Tech, 11(4):49– ??, 1993. CODEN LIHTD2. ISSN 0737-8831.

Giagnocavo:1995:EICa

[GMDS95a] Gregory Giagnocavo, Tim McLain, Vince DiStefano, and Chris Noonan Sturm. Educator’s Internet companion: Classroom connect’s complete guide to educational resources on the Internet. Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1995. ISBN 0-932577-10-5 (paperback). xvi + 271 pp. LCCN LB1044.87.E37 1995.

Giagnocavo:1995:EICb

[GMDS95b] Gregory Giagnocavo, Tim McLain, Vince DiStefano, and Chris Noonan Sturm. Educator’s Internet companion: classroom connect’s complete guide to educational resources on the Internet. Classroom Connect, Lancaster, PA, USA, second printing. Mac version. edition, 1995. ISBN 0-932577-10-5. xvi + 271 pp. LCCN LB1044.87.E37 1995.

Giagnocavo:1995:EICc

[GMDS95c] Gregory Giagnocavo, Tim McLain, Vince DiStefano, and Chris Noonan Sturm. Educator’s Internet companion: classroom connect’s complete guide to educational resources on the Internet. Classroom Connect, Lancaster, PA, USA, second printing. Windows and Mac version. edition, 1995. ISBN 0-932577-10-5. xvi + 271 pp. LCCN LB1044.87 .E37 1995.

Giagnocavo:1995:EICd

[GMDS95d] Gregory Giagnocavo, Tim McLain, Vince DiStefano, and Chris Noonan Sturm. Educator’s Internet companion: classroom connect’s complete guide to educational resources on the Internet. Classroom Connect, Lancaster, PA, USA, second printing. Windows version edition, 1995. ISBN 0-932577-10-5. xvi + 271 pp. LCCN LB1044.87.E37 1995.

Giagnocavo:1996:EIC

[GMDS96] Gregory Giagnocavo, Tim McLain, Vince DiStefano, and Chris Noonan Sturm. Educator’s Internet companion. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-569484-1. various pp. LCCN LB1044.87.E37 1996.

GPNITL:1995:EIa

[GNU95a] Great Plains National Instructional Television Library, Nebraska Educational Telecommunication Center, and University of Nebraska (Lincoln Campus) Interactive Media Group. Exploring the Internet, 1995. 1 videocassette (ca. 50 min.), plus teacher’s guide.

GPNITL:1995:EIb

[GNU95b] Great Plains National Instructional Television Library, Nebraska Educational Telecommunication Center, and University of Nebraska (Lincoln Campus) Interactive Media Group. Exploring the Internet, 1995. 5 videocassettes (ca. 40 min.), plus teacher’s guide.

Groves:1995:WWW

[GO95] Dawn Groves and John Overbaugh. World Wide Web 101: a step-by-step guide for getting connected to the Internet. SPRY, Inc., Seattle, WA, USA, 1995. ISBN ???? 100 pp. LCCN ????

Godin:1994:EMA

[God94a] Seth Godin. E-Mail Addresses of the Rich and Famous. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-40893-7. LCCN CT120 .G65 1994. US$7.95. A directory of more than 1000 of e-mail addresses of famous folks: politicians, authors, reporters and other personalities. Included are Ed Asner, Tom Brokaw, Bill Clinton, Ross Perot, Dave Barry and others.

Godin:1994:PCI

[God94b] Seth Godin. Point and Click Internet. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, Macintosh edition, 1994.

ISBN 1-56609-161-6. vii + 103 pp. LCCN ???? US$12.95, CDN$17.95. A thin volume that shows how to access the Internet via America Online. Covers electronic mail (including mailing lists), gopher, newsgroups and netiquette. Despite its trim size, it is an interesting and valuable for the beginner, filled with lots of interesting tidbits.

Godin:19xx:SD

[Godxx] Seth Godin. The Smiley Dictionary. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 19xx. ISBN 1-56609-008-3. 96 pp. US$6.95. Illustrates more than 200 smileys :-).

Golding:1992:EPP

[Gol92] Richard Golding. End-to-end performance prediction for the Internet: work in progress. Technical report UCSC-CRL-92-26, University of California, Santa Cruz, Computer Research Laboratory, Santa Cruz, CA 95064, USA, June 19, 1992. 28 pp.

Goldberg:1993:KOF

[Gol93] Fred S. Goldberg. K-12’s online future: Looking for Mr. GoodNet. T H E journal (Technological Horizons in Education), 20(10): 75–??, May 1, 1993. CODEN THEJD4. ISSN 0192-592X.

Goldman:1994:CIP

[Gol94] Neal Goldman. The complete idiot’s pocket reference to the Internet. Alpha Books, Indianapolis, IN, USA, 1994. ISBN 1-56761-528-7. ix + 214 pp. LCCN TK5105.875.I57G65 1994.

Gonzales:1995:BIN

[Gon95] Frank M. Gonzales. Branding on the Internet: new methods for a new medium. Technical report, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, 1995. 5 pp.

Goncalves:1998:FC

[Gon98] Marcus Goncalves. Firewalls complete. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-024645-9. xxi + 632 pp. LCCN TK5105.59 .G65 1998.

Goncalves:1999:MIS

[Gon99] Vinicius A. Goncalves. MCSE:

Implementing and Supporting Microsoft Exchange Server 5.5. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-923525-6. xxviii + 425 pp. LCCN QA76.3.G66

1998. URL http://www.
phptr.com/bookseri/mcse.
html.

Goodwin:1996:FNT

[Goo96] Mary Goodwin. Fodor’s net travel: your map to travel resources on the Internet and online services. Michael Wolff and Co., Inc., New York, NY, USA, 1996. ISBN 0-679-77033-X. xv + 389 pp. LCCN ????

Goodman:1997:WPC

[Goo97] Daniel J. Goodman. Wireless Personal Communications Systems. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63470-8. xxix + 417 pp. LCCN TK5103.485.G66 1997. US$51.95.

Gora:1995:IOS

[Gor95a] Jean Crooks Gora. The Internet and online services: opportunities for insurance and financial services companies. Technical report, LOMA Research Division, Atlanta, GA, USA, 1995. viii + 187 pp.

Goralski:1995:TIA

[Gor95b]  Walter Goralski.  TCP/

IP applications and protocols. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1995. ISBN 1-56607-951-9. iv + 248 pp. LCCN TK5105.585.G67 1995.

Gosselin:1995:I

[Gos95] David Gosselin. Internet. Quick reference guide. DDC Pub., New York, NY, USA, 1995. ISBN 1-56243-254-0. vi + 214 pp. LCCN TK5105.875.I57G67 1995.

Gould:1990:FRI

[Gou90] Stephen Gould. The Federal Research Internet and the National Research and Education Network: prospects for the 1990s. CRS report for Congress; 90-362 SPR Library of Congress. Congressional Research Service. CRS report for Congress; 90-362 SPR. 90-362 SPR, Congressional Research Service, Library of Congress, Washington, DC, USA, 1990. iii + 15 + 2 pp.

Gabbard:1995:IRA

[GP95a] C. Bryan Gabbard and George S. Park. The information revolution in the Arab world: commercial, cultural and political dimensions: the Middle East meets the Internet. ???? P-7920, RAND, Santa Monica, CA, USA, 1995. 29 pp.

Gross:1995:BAN

[GP95b] Adam C. Gross and Kermit M. Patton. Business applications of the new Internet: interactive information services come of age. Report; no. 834 (spring 1995) report (business intelligence program (sri international)); no. 834, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, 1995. 36 pp.

Gabbard:1996:IRA

[GP96a] C. Bryan (Claybourne Bryan) Gabbard and George S. (George Shin-Kye) Park. The information revolution in the Arab world: commercial, cultural and political dimensions: the Middle East meets the Internet. ???? P-7920-1, RAND, Santa Monica, CA, USA, 1996. 32 pp.

Gurley:1996:PDG

[GP96b] Ted Gurley and W. T. Pfefferle. Plug in: the definitive guide to music on the Internet. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-241050-8. xxvi + 320 pp. LCCN ML74.7.G87 1996.

Goodman:1994:IPG

[GPRR94] S. E. Goodman, L. I. Press, S. R. Ruth, and A. M. Rutkowski. International perspectives: The global diffusion of the Internet: Patterns and problems. Communications of the ACM, 37(8):27–31, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        179733.html.

Gallagher:1994:CGI

[GR94] Jim Gallagher and Wisdom Research. Christian guide to the Internet. Technical report, Wisdom Research, Lewisville, TX, USA, 1994. 47 pp.

Grant:1995:IAU

[Gra95] Karen Grant. The Internet:

access, usage and training in Minnesota schools. Graduate paper (m.s. in i.m.), St. Cloud State University, St. Cloud, MN, USA, 1995. 37 pp.

Grasser:1998:FC

[Gra98] R. Scott Grasser. FINDsomeone.com. Butterworth-Heinemann, Boston, MA, USA, 1998. ISBN 0-7506-7020-7. xvii + 137 pp. LCCN HV6762.U5 G73 1998. US$39.95.

Greene:1995:IWL

[Gre95a] Brian A. Greene. Internet for Wyoming libraries: basics and issues. Technical report, Wyoming State Library, division of Dept. of Administration and Information, Cheyenne, WY, USA, 1995. 16 pp.

Greenman:1995:NYC

[Gre95b] Ben Greenman. Netmusic:

your complete guide to rock and more on the Internet and online services. Random House Electronic Publishing: Michael Wolff and Co., New York, NY, USA, 1995. ISBN 0-679-76385-6. xii + 393 pp. LCCN ML12.G74 1995.

Gregory:1995:BIA

[Gre95c] Kate Gregory. Building Internet applications with Visual C++. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0213-4. xvi + 459 pp. LCCN TK5105.875.I57B85 1995.

Greenman:1996:NYC

[Gre96] Ben Greenman. NetGames 2: your complete guide to games on the Internet and online services. Michael Wolff and Co. Pub., New York, NY, USA, 1996. ISBN 0-679-77034-8 (paperback). xii + 359 pp. LCCN ????

Greenberg:1999:NAF

[Gre99] Eric Greenberg. Network application frameworks: design and architecture. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30950-5. xxviii + 378 pp. LCCN TK5105.5.G73 1999. US$49.95.

Grimes:1995:MGI

[Gri95] Galem Grimes. 10 minute guide to the Internet with Windows 95. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0663-6. xi + 188 pp. LCCN TK5105.875.I57G75 1995.

Gary:1994:ISB

[GRLJ94] Dane Gary, Michael Ressler, Steven Lipner, and Mark Jenkins. Internet security: building a firewall, forum, 1994. 1 videocassette.

Gross:1990:PSIb

[Gro90a] Phillip G. Gross, editor. Proceedings of the Seventeenth Internet Engineering Task Force, Pittsburgh Supercomputer Center, May 1–4, 1990. Corporation for National Research Initiatives, Reston, VA, USA, 1990. ISBN ???? LCCN ????

Gross:1990:PSIa

[Gro90b] Phillip G. Gross, editor.

Proceedings of the Sixteenth Internet Engineering Task Force, Florida State University, February 6–9, 1990. Corporation for National Research Initiatives, Reston, VA, USA, 1990. ISBN ???? LCCN ????

Gross:1995:IIS

[Gro95a]  Adam C. Gross.  Inter-

net infrastructure: software tools and networks. Technical report, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, 1995. 16 pp.

Gross:1995:PTL

[Gro95b] Michael Gross. A pocket tour of law on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1792-9 (paperback). xii + 176 pp. LCCN KF242.A1G77 1995.

Gardner:1995:CNF

[GRRM95] E. Ty Gardner, Kevin Rhoades, Randy Ratliff, and Joe McCullough. Connect on the net finding a job on the Internet, 1995. 1 videocassette (25 min.).

Gutsch:1979:PIC

[GRS79] Roland W. Gutsch, Hasso Reschke, and Heinz Schelle, editors. Proceedings of the 6th Internet Congress 1979. VDI-Verlag, Dusseldorf, Federal Republic of Germany, 1979. ISBN 3-18-419061-7. LCCN ????

Grundeen:1994:MSL

[Gru94] Brenda Grundeen. A manual for students learning to use the Internet. Technical report, ????, ????, 1994. iii + 49 pp.

GAI:1995:MII

[Gru95] Grupo Anonimi Internautae. Manual imprescindible de Internet. Manuales imprescindibles. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-811-9. 371 pp. LCCN ????

Grycz:1994:PPB

[Gry94] Czeslaw Jan Grycz. Promises and pitfalls: a briefing paper on Internet publishing. Professional/Scholarly Publishing Division, Association of American Publishers, New York, NY, USA, 1994. ISBN 0-933636-28-8. 76 pp. LCCN ????

Grzybowski:1994:ICH

[Grz94] Carl Grzybowski. Internet:

connectivity handbook. Technical report, ????, ????, 1994. 48 pp.

Gardner:1995:IWM

[GS95a] David C. Gardner and David A. Sauer. Internet for Windows: the Microsoft network edition. Visual learning guides. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 0-7615-0301-3. vii + 280 pp. LCCN TK5105.875.I57G368 1995.

Gibbs:1995:LDI

[GS95b] M. (Mark) Gibbs and Richard J. (Richard James) Smith. Livre d’or Internet. Livre d’or. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1478-7. xxii + 763 pp. LCCN ????

Gwertzman:1995:ARA

[GS95c] James S. Gwertzman and Margo Seltzer. Autonomous replication across wide-area internetworks. Operating Systems Review, 29(5):234, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garfinkel:1996:PUI

[GS96a] Simson Garfinkel and Gene Spafford. Practical UNIX & Internet Security. Computer security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1996. ISBN 1-56592-148-8. xxix + 971 pp. LCCN QA76.76.O63 G38 1996. US$39.95. URL

http://www.ora.com/gnn/
bus/ora/item/pus2.html;
 http://www.oreilly.com/
catalog/puis.

Gordon:1996:PTF

[GS96b] Ellen Gordon and Peter Stokes. A pocket tour of food and drink on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1806-2. xiii + 160 pp. LCCN TX357.G5757 1996.

Gide:1999:EBI

[GS99] E. Gide and F. Soliman. The economic benefits of Internet-based business operations in manufacturing. Lecture Notes in Computer Science, 1611:830–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guedon:1996:PCI

[Gue96] Jean-Claude Guedon. La planete cyber: Internet et cyberespace. Decouvertes Gallimard Techniques; 280. Gallimard, Paris, France, 1996. ISBN 2-07-053349-2. 128 pp. LCCN ????

Guisnel:1995:GCS

[Gui95] Jean Guisnel. Guerres dans le cyberespace: services secrets et Internet. Enquetes. La Decouverte, Paris, France, 1995. ISBN 2-7071-2502-4 (paperback). 251 pp. LCCN HV6773 .G85 1995. 140FF.

Gummere:1993:IWE

[Gum93a] Francis Barton Gummere.

The Internet Wiretap edition of Beowulf, 1993.

Gumprecht:1993:ISGa

[Gum93b] Blake Gumprecht. Internet sources of government information. Government documents notes (Temple University. Paley Library), Temple University, Paley Library, Philadelphia, PA, USA, October 7, 1993. 8 pp.

Gumprecht:1993:ISGb

[Gum93c] Blake Gumprecht. Internet sources of government information. Technical report, Temple University, Paley Library, Philadelphia, PA, USA, 1993. 14 pp.

Gumprecht:1994:ISG

[Gum94] Blake Gumprecht. Internet sources of government information. Government documents notes (Temple University. Paley Library), Temple University, Paley Library, Philadelphia, PA, USA, February 26, 1994. various pp.

Gunn:1994:IEL

[Gun94] Angela Gunn. Internet Express Lane. Robert J. Brady Co., Bowie, MD 20715, USA, 1994. ISBN 1-56686-190-X. US$29.95.

Gupta:1992:PSI

[Gup92] Manish Gupta. A performance study of an Internet source routing architecture. Thesis (M.S.), University of Kentucky, Lexington, KY, USA, 1992. x + 64 pp.

Gutierrez:1999:WDD

[Gut99] Dan D. Gutierrez. Web Database Development for Windows Platforms. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-013985-8. ???? pp. LCCN QA76.9.D26G88

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130139858.html.

Gross:1990:PEI

[GV90a] Phillip G. Gross and Gregory M. Vaudreuil, editors. Proceedings of the eighteenth Internet Engineering Task Force, University of British Columbia, July 30–August 3, 1990. Corporation for National Research Initiatives, Reston, VA, USA, 1990. ISBN ???? LCCN ????

Gross:1990:PFI

[GV90b] Phillip G. Gross and Gregory M. Vaudreuil, editors. Proceedings of the fifteenth Internet Engineering Task Force, University of Hawaii, October 31 – November 3, 1989. Corporation for National Research Initiatives, Reston, VA, USA, 1990. ISBN ???? LCCN ????

Gilster:1995:NI

[GVJ95] Paul Gilster, Jorge Rodriguez Vega, and Alberto Belinchon Jimenez. El navegante de Internet. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-695-7. 645 pp. LCCN ????

Golden:1994:EGI

[GVW94] Patrick S. Golden, Ronald Vasaturo, and Katherine B. Wallet. The educator’s guide to the Internet. Technical report, Virginia Space Grant Consortium, Hampton, VA, USA, 1994. iv + 168 pp. US$16.00. 1 videocassette (11 min., 45 sec.). Written and designed by educators for educators this is a user-friendly guide to make the Internet work for you in the classroom. It includes on-line projects for your students to collaborate with other classes around the world, lesson plans, graphics and software on almost every subject imaginable, etc.

Golden:1995:EGI

[GVW95] Patrick S. Golden, Ronald Vasaturo, and Katherine B. Wallet. The educator’s guide to the Internet. Technical report, Virginia Space Grant Consortium, Hampton, VA, USA, 1995. 191 pp.

Gardiner:1995:PCS

[GW95] John Gardiner and Barry West. Personal Communication Systems and Technologies. Artech House Inc., Boston, MA, USA, 1995. ISBN 0-89006-588-8. xiii + 238 pp. LCCN TK6570.M6 P48 1995. US$59.00.

Garg:1999:PAG

[GW99] Vijay K. Garg and Joseph E. Wilkes. Principles and Applications of GSM. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-949124-4. ???? pp. LCCN TK5103.483.G37

                                        1999. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0139491244.html.

Heslop:1994:IIG

[HA94] Brent D. Heslop and David Angell. The Instant Internet Guide—Hands-on Global Networking. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-62707-8. x + 209 pp. LCCN TK5105.875.I57 H48 1994. US$14.95, CDN$18.95. Contents: Merging onto the Internet — Communicating via electronic mail — Turning on network news — Interactive Internet — Mining files on the Internet — Finding what you want on the Internet — UNIX in about an hour — Getting a dial-up connection to the Internet.

Heslop:1995:WWC

[HA95] Brent D. Heslop and David Angell. The World-Wide Web Companion: a Mosaic voyager’s guide to Internet hypermedia. Addison-Wes-ley, Reading, MA, USA, 1995. ISBN 0-201-40914-3. ???? pp. LCCN ????

Haas:1985:DGB

[Haa85] Walter Oskar Haas. Design of a gateway between a public data network and an Ethernet using the Transmission Control Protocol and Internet Protocol. Thesis (M.S.), Department of Computer Science, University of Utah, Salt Lake City, UT 84112, USA, 1985. vi + 84 + 1 pp.

Hackathorn:1998:WFD

[Hac98] Richard D. Hackathorn. Web Farming for the Data Warehouse. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-503-7. 384 pp. LCCN QA76.9.D37H32 1998. US$44.95.

Hahn:1996:HHI

[Hah96a] Harley Hahn. Harley Hahn’s The Internet complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1996. ISBN 0-07-882138-X (paperback). xxix + 802 pp. LCCN TK5105.875.I57H34 1996.

Hahn:1996:IYP

[Hah96b] Harley Hahn. The Internet yellow pages. Osborne/ McGraw-Hill, Berkeley, CA, USA, third edition, 1996. ISBN 0-07-882182-7 (paperback). l + 830 pp. LCCN ????

Hall:1993:CAI

[Hal93] Richard Scott Hall. A caching architecture for the Internet file transfer protocol. Thesis (M.S.), University of Colorado, Boulder, CO, USA, 1993. viii + 42 pp.

Hallgren:1994:FIP

[Hal94] Martyne M. Hallgren. Funding an Internet public good: definition and example. Computer Networks and ISDN Systems, 27(3):403–409, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        comnet/cas_sub/browse/browse.
                                        cgi?year=1994&volume=27&
                                        issue=3&aid=1360.

Hall:1995:IVI

[Hal95a] Evelyn Hall. Insurers venture onto the Internet Web. Best’s review (Life/health insurance ed.), 96(2):40–??, June 1, 1995. ISSN 0005-9706.

Halter:1995:SEM

[Hal95b] Christopher P. Halter. Spacemath: exploring mathematics through the Internet. Thesis (M.A.), University of California, San Diego, San Diego, CA, USA, 1995. xii + 163 pp.

Halabi:1997:IRA

[Hal97] Bassam Halabi. Internet routing architectures. Cisco Press, Indianapolis, IN, USA, 1997. ISBN 1-56205-652-2. xvii + 477 pp. LCCN TK5105.875.I57 H35 1997.

Hall:1999:SLE

[Hal99] Clint Hall. SP3D and the Lighthouse: explorations in 3D Internet learning. In ACM [ACM99b], pages 95–96. ISBN 0-201-48560-5. ISSN 1069-529X. LCCN T385 .S54

1999. URL http://www.
acm.org/pubs/citations/
proceedings/graph/311625/

p95-hall/. ACM order number 428990.

Hamlin:1999:EWA

[Ham99]  J. Scott Hamlin.  Effec-

tive Web animation: advanced animation techniques for the Web. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-60600-3. 336 pp. LCCN TR897.7.H36 1999. US$44.95.

Hancock:1992:IBO

[Han92] Lee Hancock. Internet/Bitnet online health sciences resource list. Technical report, ????, ????, June 12, 1992. 17 pp. Version 3.0.1. Computer generated printed list of electronic databases, discussion groups, journals, newsletters, etc. Some entries include descriptive information.

Hancock:1993:IBH

[Han93]  Lee Hancock.  Internet/

Bitnet health sciences resource. Technical report, ????, ????, 1993. 67 pp.

Hancock:1994:IBH

[Han94a]  Lee Hancock.  Internet/

Bitnet health sciences resources. Technical report, ????, ????, 1994. various pp.

Hancock:1994:IHS

[Han94b] Lee Hancock. Internet/Bitnet health sciences resources. Technical report, ????, ????, 1994. Previous editions titled: Internet/Bitnet online health sciences resource list. Often referred to as “The Hancock list.” Description based on: version 3-26-94 (copyright 1993) Frequently updated computer generated printed list of electronic databases, health agencies online, medical libraries online, discussion groups, Freenets, online journals, newsletters, etc. Some entries include descriptive information.

Hancock:1995:PGI

[Han95]  Lee Hancock.  Physi-

cians’ guide to the Internet. Lippincott-Raven, Philadelphia, PA, USA, 1995. ISBN 0-397-51634-7. xvii + 212 pp. LCCN R118 .H25 1995.

Hancock:1996:IIS

[Han96] Scott Hancock. Introduction to the Internet and the social science database archive. Thesis (B.A.), California Polytechnic State University, San Luis Obispo, CA, USA, 1996. 25 pp.

Hart:1992:BIIa

[Har92] Jeffrey A. Hart. The building of the Internet: implications for the future of broadband networks. Technical Report 90, Indiana Center for Global Business, Graduate School of Business Indiana University, Bloomington, IN, USA, 1992. 50 pp.

Harris:1993:MIN

[Har93a] Judi Harris. Mining the Internet: Networked information location tools: Gophers, Veronica, Archie, and Jughead. Computing Teacher, 21(1):16–19, August 1, 1993. ISSN 0278-9175.

Hart:1993:OHS

[Har93b] Michael Hart. Opinion: How shared ARE our shared resources, anyway? Internet research, 3(2):4–??, Summer 1993. CODEN IRESEF. ISSN 1066-2243.

Harnedy:1994:TS

[Har94a]  Sean J. Harnedy.  Total

SNMP: Exploring the Simple Network Management Protocol. CBM Books, Horsham, PA, USA, 1994. ISBN 1-878956-33-7. xvii + 626 pp. LCCN TK5105.55 .H37 1994. US$45.00.

Harrington:1994:CDP

[Har94b] John Harrington. Current developments in the provision of aerospace information: the role of electronic sources indlucing Internet. College of Aeronautics report; 9403. College of Aeronautics, Cranfield University, Cranfield, Bedford, England, November 1994. ISBN 1-871564-73-5 (paperback). ii + 34 pp. LCCN ????

Harris:1994:IRD

[Har94c] Judi Harris. Internet resource directory. Technical report, ????, ????, 1994. various pp.

Harris:1994:WFFb

[Har94d] Judi Harris. Way of the Ferret — Finding Educational Resources on the Internet. International Society for Technology in Education, Eugene, OR, USA, 1994. ISBN 1-56484-055-7. xiii + 191 pp. LCCN ???? US$24.95 (ISTE members US$22.45). Based on Judi Harris’ popular Mining the Internet column in The Computing Teacher journal, Way of the Ferret is designed to help you acquire the knowledge, skills, and experience necessary to locate and use resources found on the Internet. Carefully developed analogies help users grasp basics of Telnet sessions, FTP file transfers, information location tools, file encryption, and discussion groups. Finally, 15 different types of educational telecomputing activities help you plan for integrating Internet resources into your students’ academic explorations.

Harris:1994:WFFa

[Har94e] Judi Harris. Way of the ferret: finding educational resources on the Internet. International Society for Technology in Education, Eugene, OR, USA, revised edition, 1994. ISBN 1-56484-069-7. xiii + 191 pp. LCCN TK5105.875 .I57 H38 1994 Curriculum Lab.

Harrison:1994:UH

[Har94f] Mark Harrison. The Usenet Handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1994. ISBN 1-56592-101-1. 250 pp. US$24.95. A user’s guide to netnews.

Harrison:1994:IDC

[Har94g] Peter John Harrison. Internet Direct Connection Kit. Programmers Press, Foster City, CA, USA, 1994. ISBN 1-56884-135-3. xiv + 283 pp. LCCN TK5105.875.I57 H374 1994. US$29.95.

Harris:1995:WFF

[Har95a] Judi Harris. Way of the ferret: finding and using educational resources on the Internet. International Society for Technology in Education, Eugene, OR, USA, second edition, 1995. ISBN 1-56484-085-9. xiii + 290 pp. LCCN ????

Harris:1995:ISG

[Har95b] Stuart (Stuart H.) Harris. The IRC Survival Guide: Talk to the World with Internet Relay Chat. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-41000-1. viii + 213 pp. LCCN TK5105.875.I57 H37 1995. US$17.95, CDN$22.95.

Harrison:1995:UHU

[Har95c] Mark Harrison. The USE-NET Handbook: a User’s Guide to Netnews. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1995. ISBN 1-56592-101-1. 388 pp. LCCN ???? US$24.95.

Harmon:1996:UIO

[Har96a]  Charles Harmon.  Using

the Internet, online services and CD-ROMS for writing research and term papers. Neal-Schuman NetGuide series Neal-Schuman net-guide series. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-238-4. viii + 167 pp. LCCN LB1047.3.U75 1996.

Harris:1996:IEG

[Har96b] Cheryl Harris. An Internet education: a guide to doing research on the Internet. Integrated Media Group, Belmont, CA, USA, 1996. ISBN 0-534-25851-4. x + 165 pp. LCCN TK5105.875.I57 H366 1996.

Harnedy:1999:WBI

[Har99a] Sean Harnedy. Web-based Information Management: For the Enterprise. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-096018-7. xxiv + 389 pp. LCCN T58.6.H347

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130960187.html.

Harrington:1999:ENC

[Har99b]  Jan Harrington.  Ether-

net Networking Clearly Explained. Academic Press, New York, NY, USA, 1999. ISBN 0-12-326427-8. xv + 287 pp. LCCN TK5105.8.E83 H27 1999. US$41.95.

Harper:1995:BWN

[HAS95] Eric Harper, Matthew Flint Arnett, and R. Paul (Ravinder Paul) Singh. Building a Windows NT Internet server. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-484-8. xi + 222 pp. LCCN QA76.76.O63H3488 1995.

Haskin:1996:NJI

[Has96] David Haskin. News junkies Internet 500. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-461-8. ???? pp. LCCN PN4784.E5H37 1996.

Hawgood:1996:IG

[Haw96a] David Hawgood. Internet for genealogy. D. Hawgood, London, UK, 1996. ISBN 0-948151-12-9 (paperback). 24 pp. LCCN ????

Hawkes:1996:HTI

[Haw96b]  Lory Hawkes.  Hyperspa-

tial travel into the Internet: a guidebook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-229329-3 (paper). viii + 168 pp. LCCN TK5105.875.I57H39 1996.

HaydenBooks:1994:IC

[Hay94a]  Hayden Books.  Internet

Chicago. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-038-7. xi + 116 pp. LCCN TK5105.875.I57I543 1994. US$12.00, CDN$16.95.

HaydenBooks:1994:INY

[Hay94b]  Hayden Books.  Internet

New York. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-039-5. xi + 101 pp. LCCN TK5105.875.I57I563 1994. US$12.00, CDN$16.95.

HaydenBooks:1994:ISF

[Hay94c]  Hayden Books.  Internet

San Francisco. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-040-9. x + 118 pp. LCCN TK5105.875.I57I58 1994. US$12.00, CDN$16.95.

Haynes:1995:TIC

[Hay95] Phyllis Haynes. TPR-NE Internet conference, 1995. 1 videocassette (61 min.).

Heslop:1995:HPIb

[HB95] Brent D. Heslop and Larry Budnick. HTML publishing on the Internet for Windows: create great-looking documents online: home pages, newsletters, catalogs, ads and forums. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-229-1. xxii + 480 pp. LCCN QA76.76.H94H483 1995.

Hurley:1996:SBG

[HB96] Brian Hurley and Peter Birkwood. A small business guide to doing big business on the Internet. Self-Counsel business series. Self-Counsel Press, North Vancouver, BC, Canada, 1996. ISBN 1-55180-029-2. xvii + 174 pp. LCCN ????

Hart:1998:MWS

[HB98a] David L. Hart and Philip E. Bourne. Mac OS 8 Web Server Cookbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-520016-4. xxviii + 435 pp. LCCN QA76.76.O63H355

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0135200164.html;   http:
//www.sdsc.edu/Cookbook/
Mac.

Harvey:1998:DGO

[HB98b] Dennis Harvey and Steve Beitler. The Developer’s Guide to Oracle Web Application Server 3. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30960-2. 736 pp. LCCN QA76.76.D47H397 1998. US$44.95.

Hammond:1994:IPG

[HC94] Drew Hammond and Computer Channel, Inc. Internet practical guidelines for use, 1994. 1 videocassette.

Handley:1995:WWW

[HC95] Mark Handley and Jon Crowcroft. The World Wide Web: Beneath the Surf. UCL Press, London, UK, 1995. ISBN 1-85728-435-6. ix + 198 pp. LCCN TK5105.888 .H354 1995.

Handley:1999:VLC

[HCBO99] M. Handley, J. Crowcroft, C. Bormann, and J. Ott. Very large conferences on the Internet: the Internet multimedia conferencing architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):191–204, February 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/3/2068.pdf.

Hare:1995:BLI

[HE95a] Chris Hare and George Eckel.

Building a Linux Internet server. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-525-9. xiv + 340 pp. LCCN QA76.9.C55H36 1995.

Hassinger:1995:MOI

[HE95b] Sebastian Hassinger and Mike Erwin. Mecklermedia’s official Internet world 60 minute guide to VRML. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-710-6. xx + 218 pp. LCCN TK 5105.888 H38 1995.

He:1998:IRE

[He98]  Jimin He.  Internet Re-

sources for engineers: a Practical Handbook for Students. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 0-7506-8949-8. vi + 298 pp. LCCN TA158.7 .H4 1998. US$34.95.

Anonymous:1994:HGI

[Hea94] HealthLink: global Internet resources and databases, page various, 1994. ISSN 1077-6036. Marcus P. Zillman, Marco Island, FL, USA.

Anonymous:1995:HID

[Hea95] Healthcare Internet directory, page various, 1995. Stewart Publishing, Inc., Alexandria, VA, USA.

Hedrick:1987:IIPa

[Hed87a] Charles L. Hedrick. Introduction to the Internet protocols. Technical report, Rutgers University, Computer Science Facilities Group, New Brunswick, NJ?, 1987. i + 27 pp.

Hedrick:1987:IIPb

[Hed87b] Charles L. Hedrick. Introduction to the Internet protocols. Technical report, Rutgers, Center for Computers and Information Services Laboratory for Computer Science Research, New Brunswick, NJ, USA, 1987. i + 18 pp.

Heels:1993:LLI

[Hee93a] The legal list: Internet desk reference: law-related resources on the Internet and elsewhere, page various, 1993. Lawyers Cooperative Pub., Rochester, NY, USA.

Heels:1993:LLL

[Hee93b] Erik J. Heels. The legal list: law-related resources on the Internet and elsewhere. Technical report, University of Maine School of Law, Portland, ME, USA, 1993. 75 pp.

Heels:1994:LLLa

[Hee94a] Erik J. Heels. The legal list: law-related resources on the Internet and elsewhere. Technical report, University of Maine School of Law, Portland, ME, USA, 1994. vii + 88 pp.

Heels:1994:LLLb

[Hee94b] Erik J. Heels. The legal list: law-related resources on the Internet and elsewhere. Technical report, University of Maine School of Law, Portland, ME, USA, 1994.

Heels:1994:LLLc

[Hee94c] Erik J. Heels. The legal list: law-related resources on the Internet and elsewhere. Technical report, University of Maine School of Law, Portland, ME, USA, 1994.

Heels:1994:LLLd

[Hee94d] Erik J. Heels. The legal list:

law-related resources on the Internet and elsewhere. E. J. Heels, Eliot, ME, USA, v5.1 edition, September 17, 1994. ISBN 0-9643637-0-4. 274 pp. LCCN KF242.A1H443 1994.

Heels:1995:LLIa

[Hee95a]  Erik J. Heels.  The le-

gal list: Internet desk reference: law-related resources on the Internet and elsewhere. Lawyers Cooperative Publishing, Rochester, NY, USA, version 6.0 edition, 1995. ISBN 0-7620-0000-7. xx + 337 pp. LCCN ????

Heels:1995:LLIb

[Hee95b]  Erik J. Heels.  The le-

gal list: Internet desk reference: law-related resources on the Internet and elsewhere. Lawyers Cooperative Publishing, Rochester, NY, USA, October 1995. ISBN 0-7620-0012-0. xxxv + 341 pp. LCCN KF242.A1H443 1995.

Hegering:1999:IMN

[Heg99] Heinz-Gerd Hegering. Integrated Management of Networked Systems: Concepts, Architectures and their Operational Application. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-571-1. xviii + 651 pp. LCCN TK5105.5 .H43313 1999. US$66.95.

Heim:1995:IC

[Hei95a] Judy Heim. Internet for cats. No Starch Press, Daly City, CA, USA, 1995. ISBN 1-886411-07-7. ???? pp. LCCN M6231.C23H45 1995.

Heise:1995:ANA

[Hei95b] Donalyn M. Heise. An investigation of Nebraska art teachers’ perceptions and usage of Internet technology. Thesis (M.A.), University of Nebraska at Omaha, Omaha, NE, USA, 1995. 92 pp.

Helfrick:1991:ESN

[Hel91a] Albert Helfrick. Electrical Spectrum and Network Analyzers: a Practical Approach. Academic Press, New York, NY, USA, 1991. ISBN 0-12-338250-5. ix + 212 pp. LCCN TK7879.4 .H45 1991. US$101.00.

HL:1991:II

[Hel91b] Helix Limited. Implementing Internet. Helix Software Consultants Ltd, London, UK, 1991. ISBN 1-873984-00-6 (paperback). ii + 26 pp. LCCN ????

Held:1994:CCR

[Hel94] Gilbert Held. The complete cyberspace reference and directory: an addressing and utilization guide to the Internet, electronic mail systems, and bulletin board systems. VNR communications library. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01913-0. ix + 765 pp. LCCN TK5105.5 .H437 1994.

Held:1998:ENG

[Hel98] Gilbert Held. Ethernet Networks: From 10Base-T to Gigabit. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, third edition, 1998. ISBN 0-471-25310-3. 528 pp. LCCN TK5105.8.E83H45 1998. US$39.95.

Held:1999:GEP

[Hel99] Gilbert Held. Gigabit Ethernet — the potential illusion of speed. Sys Admin: The Journal for UNIX Systems Administrators, 8(6): 51–55, June 1999. CO-DEN SYADE7. ISSN 1061-

                                        2688. URL http://www.
                                        samag.com/.

Hempel:1993:EI

[Hem93] Paul. Hempel. Exploring the Internet. Media and methods, 30(2):26–??, November 1, 1993. ISSN 0025-6897.

Henderson:1982:LMO

[Hen82] Cecelia E. Henderson. Locating migratory objects in an Internet. Thesis (M.S.), M.I.T., Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1982. i + 75 pp. Supervised by Barbara Liskov.

Henderson:1994:WGI

[Hen94]  Harry Henderson.  The

Waite Group’s Internet Howto. Waite Group Press, Corte Madera, CA, USA, 1994. ISBN 1-878739-68-9. xviii + 468 pp. LCCN TK5105.875.I57 H46 1994. US$34.95, CDN$48.95.

Henderson:1995:ACI

[Hen95a] Danna K. Henderson. Airlines in cyberspace: the Internet and the World Wide Web provide new, broad electronic channels for carriers to sell tickets and reduce distribution costs. Air transport world, 32(11):26–32, November 1995. ISSN 0002-2543.

Henderson:1995:CTI

[Hen95b] Harry Henderson. Ching tung Internet = Internet how-to. Sung kang tien nao tu shu tzu liao ku fen yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-22-1849-2 (paperback). various pp. LCCN ????

Henderson:1999:DII

[Hen99] D. B. Henderson. Distributing 3-D information over the Internet using VRML. In Anonymous [Ano99b], pages 561–566. ISBN ???? LCCN ????

Herman:1993:USV

[Her93] Debra Herman, editor. UNIX System V NFS administration. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-016411-9. xiv + 164 pp. LCCN QA76.76.O63 U5521125 1993.

HWU:1994:IRN

[Her94] Heriot-Watt University. Internet resources newsletter, 1994.

Herz:1995:SI

[Her95a] J.-C Herz. I surfisti di Internet. I canguri. Feltrinelli, Milano, Italy, 1995. ISBN 88-07-70070-0. 239 pp. LCCN ????

Herz:1995:SINa

[Her95b] J. C. (Jessie Cameron) Herz. Surfing on the Internet: a nethead’s adventure on-line. Back Bay Books, Boston, MA, USA, paperback edition, 1995. ISBN 0-316-36009-0 (paperback). 321 pp. LCCN TK5105.875.I57 H47 1995.

Herz:1995:SINb

[Her95c] J. C. (Jessie Cameron) Herz. Surfing on the Internet: a nethead’s adventures online. Little, Brown and Co., Boston, MA, USA, 1995. ISBN 0-349-10614-2 (paperback), 0-316-35958-0. 321 pp. LCCN TK5105.875.I57 H47 1995.

Hernandez:1996:BTI

[Her96a] Tamsen M. Hernandez. Business, technology, and the Internet: a council report. Report 1143-96-CR, Conference Board, New York, NY, USA, 1996. ISBN 0-8237-0593-5. 39 pp.

Hersch:1996:FCP

[Her96b] Lisa J. Hersch. Freedom of choice: pornography and censorship on the Internet. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1996. 77 pp.

Heslop:1995:PIW

[Hes95] Brent Heslop. Publishing on the Internet for Windows: Creating Great-looking Documents Online: Home Pages, Newsletters, Catalogs, Ads and Forums. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-229-1. 350 pp. LCCN Z286.E43 H47 1995. US$49.95.

HP:1986:LCA

[Hew86] Hewlett–Packard Co. LAN Cable and Accessories Installation Manual, HP part no. 5955-7680 edition, January 1986. 204 pp. US$43.00.

HP:1988:STC

[Hew88] Hewlett–Packard Co. HP SiteWire Twisted-pair Cabling Installation Guide, HP manual part number 5959-2208 edition, January 1988. 100 pp. US$40.00.

Hayward:1996:HUI

[Hf96] Michael Hayward and Eudora (Computer file). How to use Internet Email: installing and using Eudora to get the most out of electronic mail.

Go!Tools. Motion Works Pub., San Francisco, CA, USA, 1996. ISBN 1-57712-002-7 (paperback). ???? pp. LCCN HE6239.E54H39 1996.

Hart:1995:BHI

[HG95] Michael S. Hart and Project Gutenberg. A brief history of the Internet, 1995.

Horn:1997:SVT

[HG97] Uwe Horn and Bernd Girod. Scalable video transmission for the Internet. Computer Networks and ISDN Systems, 29(15):1833–1842, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=15&aid=1768.

Hubaux:1999:IIT

[HGKB99] Jean-Pierre Hubaux, Constant Gbaguidi, Shawn Koppenhoefer, and Jean-Yves Le Boudec. The impact of the Internet on telecommunication architectures. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):257–273, February 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/3/2069.pdf.

Hunt:1998:NDC

[HGKM98] Guerney D. H. Hunt, Germán S.

Goldszmidt, Richard P. King, and Rajat Mukherjee. Network Dispatcher: a connection router for scalable Internet services. Computer Networks and ISDN Systems, 30(1–7):347–357, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/1-7/1899.pdf.

HBSPC:1995:DBI

[HH95a] Harvard Business School Publishing Corporation and Harvard University. Graduate School of Business Administration. Doing business on the Internet an interview with Shikhar Ghosh, CEO of Open Market, 1995. 1 videocassette.

Heslop:1995:HPIa

[HH95b] Brent D. Heslop and David Holzgang. HTML publishing on the Internet for Macintosh: create great-looking documents online: home pages, newsletters, catalogs, ads & forms. Ventana Press, Chapel Hill, NC, USA, Macintosh version edition, 1995. ISBN 1-56604-232-1 (kit), 1-56604-228-3 (book). xxvi + 539 pp. LCCN QA76.76.H94H48 1995. Includes 2 CD-ROMs.

Hogarth:1996:IGH

[HH96] Michael Hogarth and David Hutchinson. An Internet guide for the health professional. New Wind Pub., ????, second edition, 1996. various pp. LCCN W26.5 H63 1996, TK 5105.875 I57 H715i 1996.

Hauben:1997:NHI

[HH97] Michael Hauben and Ronda Hauben. Netizens: on the history and impact of Usenet and the Internet. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7706-6. xvi + 344 pp. LCCN TK5105.875.I57 H386

1997. URL http://www.
columbia.edu/~hauben/netbook/
;   http://www.columbia.
edu/~rh120/;   http://www.
loc.gov/catdir/bios/wiley046/
97005920.html;   http://
www.loc.gov/catdir/description/
wiley039/97005920.html;
 http://www.loc.gov/catdir/
toc/wiley031/97005920.html.

Heslop:1998:HPI

[HH98] Brent D. Heslop and David A. Holzgang. HTML Publishing on the Internet: everything you need to create professional-looking Web pages. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, second edition, May 1998. ISBN 1-56604-625-4. xxvi + 721 pp. LCCN QA76.76.H94H48 1998. US$49.95.

Handel:1994:ANC

[HHS94] Rainer Händel, Manfred N.

Huber, and Stefan Schröder. ATM Networks: Concepts, Protocols, Applications. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-42274-3. xvii + 287 pp. LCCN TK5105.35.H36 1994. US$43.25.

Hickox:1994:TIS

[Hic94] Charles R. Hickox. Training for Internet stages of concern among academic library staff in the AMIGOS consortium. Thesis (Ph.D.), East Texas State University, Commerce, TX, USA, 1994. xiv + 131 pp.

Hill:1996:UIS

[Hil96] Regina Brown Hill. Utilizing the Internet in schools. Technical report, University of North Carolina at Wilmington, Wilmington, NC, USA, 1996. iv + 50 pp. Field study.

Hillis:1997:HCT

[Hil97] Danny Hillis. The human connection: a time of transition. Communications of the ACM, 40(2):37–39, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
1997-40-2/p37-hillis/.

Hinden:1991:IRP

[Hin91] R. Hinden. Internet routing protocol standardization criteria, 1991. Includes floppy disk.

Hintikka:1993:IKT

[Hin93] Kari A. Hintikka. Internet: kalastusta tietoverkoilla. Silta. Painatuskeskus, Helsinki, Finland, 1993. ISBN 951-37-1471-3. viii + 184 pp. LCCN TK5105.875.I57H56 1993.

Hirschfeld:1997:FCF

[Hir97] Rafael Hirschfeld, editor. Financial cryptography: First International Conference, FC ’97, Anguilla, British West Indies, February 24– 28, 1997: proceedings, volume 1318 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CO-DEN LNCSD9. ISBN 3-540-63594-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1997.

Hirschfeld:1998:FCS

[Hir98] Rafael Hirschfeld, editor. Financial Cryptography: Second International Conference, FC ’98, Anguilla, British West Indies, February 23–25, 19987: proceedings, volume 1465 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64951-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1998.

Hiura:1996:UIU

[Hiu96] H. Hiura. Unicode and internationalization with UNIX and X-Window’s system. In Unicode Consortium [Uni96a], pages C3–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.]

– B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Hart:1993:NSI

[HJ93] Richard Hart and Eric James.

[NASA select internet], 1993. 1 videocassette (11 min.).

Hardcastle-Kille:1991:RDO

[HK91a] S. E. Hardcastle-Kille. Replication and distributed operations extensions to provide an Internet directory using X.500, November 1991. 1 data file 5 1/4 in. Available via FTP.

Hardcastle-Kille:1991:RRP

[HK91b] S. E. Hardcastle-Kille. Replication requirements to provide an Internet directory using X.500, November 1991. 1 data file 5 1/4 in. Available via FTP.

Harris:1995:NQTc

[HK95a] Stuart Harris and Gayle Kidder. Netscape Quick Tour for Windows: Accessing and Navigating the Internet’s World Wide Web. Ventana Press, Chapel Hill, NC, USA, special edition, 1995. ISBN 1-56604-266-6. ???? pp. LCCN TK5105.882 .H373 1995b. US$14.00.

Harris:1995:NQTd

[HK95b] Stuart Harris and Gayle Kidder. Netscape Quick Tour, Special Edition. Ventana Press, Chapel Hill, NC, USA, special edition, 1995. ISBN 1-56604-267-4 (Macintosh), 1-56604-266-6 (Windows). 184 (Macintosh), 192 (Windows) pp. LCCN TK5105.882 .H373 1995b. US$24.95.

Harris:1995:NQTa

[HKB95a] Stuart Harris, Gayle Kidder, and Gareth Branwyn. Netscape Quick Tour for the Macintosh: Accessing and Navigating the Internet’s World Wide Web. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-248-8. xx + 163 pp. LCCN TK5105.882 .H37 1995. US$14.00.

Harris:1995:NQTb

[HKB95b] Stuart Harris, Gayle Kidder, and Gareth Branwyn. Netscape Quick Tour for Windows: Accessing and Navigating the Internet’s World Wide Web. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-244-5. xx + 163 pp. LCCN TK5105.882 .H373 1995. US$14.00.

Hoffman:1996:IWU

[HKN96] Donna L. Hoffman, William D.

Kalsbeek, and Thomas P. Novak. Internet and Web use in the U.S. Communications of the ACM, 39(12): 36–46, December 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henry:1991:SSO

[HKR91] Marcia Klinger Henry, Linda Keenan, and Michael Reagan. Search sheets for OPACs on the Internet: a selective guide to U.S. OPACs utilizing VT100 emulation. Supplements to Computers in libraries. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1991. ISBN 0-88736-767-4. vii + 195 pp. LCCN Z699.35.C38 H46 1991.

Helman:1989:HNI

[HL89a] Daniel R. Helman and Darrell Long. High-speed networks and the Internet. Technical report UCSC-CRL-89-20, University of California, Santa Cruz, Computer Research Laboratory, Santa Cruz, CA 95064, USA, September 1989. 9 pp.

Helman:1989:HSN

[HL89b] Daniel R. Helman and Darrell Long. High-speed networks and the Internet. Technical report, Computer Research Laboratory, University of California, Santa Cruz, CA, USA, 1989. 9 pp.

Hegering:1993:EBC

[HL93] Heinz-Gerd Hegering and Alfred Läpple. Ethernet: Building a Communications Infrastructure. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-62405-2. vii + 389 pp. LCCN TK5105.8.E83 H45 1993.

Hoffman:1994:Ia

[HL94a] Paul Hoffman and John R. Levine. The Internet. Compaq Press: IDG Books, Foster City, CA, USA, deluxe edition, 1994. ISBN 1-56884-247-3. xiv + 241 pp. LCCN ????

Hoffman:1994:Ib

[HL94b] Paul Hoffman and John R. Levine. The Internet. Compaq Press: IDG Books, Foster City, CA, USA, 1994. ISBN 1-56884-461-1. xiv + 210 pp. LCCN TK5105.875.I57H637 1994b.

Hafner:1996:WWS

[HL96] Katie Hafner and Matthew Lyon. Where Wizards Stay Up Late: the Origins of the Internet. Simon and Schuster, 1230 Ave. of the Americas, New York, NY 10020, USA, 1996. ISBN 0-684-81201-0 (hardcover), 0-684-83267-4 (paperback), 0-684-87216-1 (e-book), 0-7567-9221-5 (paperback). 304 pp. LCCN TK5105.875.I57H338 1996.

Hui:1999:IAI

[HL99] Lucas Chi-Kwong Hui and Dik Lun Lee, editors. Internet applications: 5th International Computer Science Conference, ICSC ’99, Hong Kong, China, December 13–15, 1999: proceedings, volume 1749 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66903-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1749.

Hart:1995:UIT

[HLB^{+}95] Richard Hart, Dorothy Lucey, Thom Bird, Matthew Barzun, Amory Gabel, and Frank Voci. The ultimate Internet tour, 1995. 1 videocassette (59 min.).

Liu:1995:TSI

[hLcL95] Mei hui Liu and An chu Lin. Tsou! shang Internet chu! Po shuo ku wen “ching tien hsi lieh”. Po shuo ku wen yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-8948-62-X. 537 pp. LCCN ????

Hafner:1991:C

[HM91] Katie Hafner and John Markoff. Cyberpunk — Outlaws and Hackers on the Computer Frontier. Simon and Schuster, 1230 Ave. of the Americas, New York, NY 10020, USA, 1991. ISBN 0-671-68322-5. 368 pp. LCCN QA76.9.A25 H34 1991. US$22.95. Interviews with some of the crackers who have appeared conspicuously in the press in the past few years. One of the co-authors is the New York Times reporter who broke the Stoll story to the public.

Hurych:1996:EIH

[HM96] Jitka Hurych and Elaine Martin. Exploring Internet 5 health and medical resources, 1996. 1 videocassette (120 min.) guide.

Henderson:1997:SVE

[HM97] D. B. Henderson and H. L. McKague. Spatio-temporal visualization of earthquake hypocenters over the Internet using VRML. In Anonymous [Ano97a], pages 369– 373. ISBN 1-57083-043-6 (set), 1-57083-044-4 (vol. 1), 1-57083-045-2 (vol. 2), 1-57083-047-9 (vol. 4), 1-57083-048-7 (vol. 5). LCCN ???? Five volumes.

Hacking:1995:INW

[HMG^{+}95] Edwin Hacking, Denise Marcucci, John Gage, Arthur Van Hoff, Bill Joy, Geoffrey Baehr, and Karl Jacob. Internet: the next wave, 1995. 1 videocassette (60 min.).

Hardie:1992:IMLa

[HN92a] Edward T. L. Hardie and Vivian Neou. Internet: Mailing Lists. Internet information series. SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028, 1992. ISBN ???? i + 311 pp. LCCN TK5105.8.I57 I55 1992.

Hardie:1992:IMLb

[HN92b] Edward T. L. Hardie and Vivian Neou. Internet: mailing lists. Internet information series. SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028, September 1992. ISBN 0-944604-16-1. i + 337 pp. LCCN ????

Hardie:1993:IML

[HN93] Edward T. L. Hardie and Vivian Neou. Internet: Mailing Lists. SRI Internet information series Internet information series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993 edition, 1993. ISBN 0-13-289661-3, 0-13-327941-3. x + 356 pp. LCCN TK5105.875.I57I56 1993. US$26.00.

Hayward:1996:HPI

[HN96] Michael Hayward and Maureen Nicholson. How to publish on the Internet: build your own home page on the World Wide Web. Go!Tools. Motion Works Pub., San Francisco, CA, USA, 1996. ISBN 1-57712-003-5 (softcover). ???? pp. LCCN QA76.76.H94H39 1996.

Hoben:1996:GHC

[Hob96] John W. Hoben. 1996 guide to health care resources on the Internet. Faulkner and Gray’s Healthcare Information Center, New York, NY, USA, 1996. ISBN 1-881393-61-5. xvii + 335 + 47 + 113 + 31 pp. LCCN ????

Hoeltgen:1995:IPT

[Hoe95] Dominique Hoeltgen. Internet pour tous. Les ed. du Telephone, Paris, France, 1995. ISBN 2-909879-12-7. 238 pp. LCCN ????

Hoffman:1994:IKR

[Hof94a] Paul Hoffman. Intanetto kuikku rifarensu. Jasuto Shisutemu, Tokushima-ken, Japan, 1994. ISBN 4-88309-063-9. xxiii + 354 pp. LCCN ????

Hoffman:1994:IIRa

[Hof94b] Paul E. Hoffman. Internet Instant Reference. The Sybex instant reference series. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 0-7821-1512-8 (paperback). xvi + 317 pp. LCCN TK5105.875.I57 H64 1994. US$12.99.

Hoffman:1995:FWI

[Hof95a]  Allan Hoffman.  50 fun

ways to Internet. Career Press, Franklin Lakes, NJ, USA, 1995. ISBN 1-56414-176-4. 191 pp. LCCN TK5105.875.I57H63 1995.

Hoffman:1995:GDW

[Hof95b] Gary B. Hoffman. Gen-web a distributed world wide genealogy database on the Internet, 1995. 1 sound cassette.

Hoffman:1995:SI

[Hof95c] Gary B. Hoffman. Surfing the Internet, 1995. 1 sound cassette.

Hoffman:1995:DIWa

[Hof95d]  Paul Hoffman.  Destina-

tion INTERNET and world wide web. Compaq Press: IDG Books Worldwide, Inc., Foster City, CA, USA, 1995. ISBN 1-56884-469-7. xiv + 224 pp. LCCN TK5105.875.I57H636 1995.

Hoffman:1995:DIWb

[Hof95e] Paul Hoffman. Destination Internet and World Wide Web. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-727-0. xiv + 224 pp. LCCN TK5105.875.I57H636 1995b.

Hoffman:1995:IIR

[Hof95f] Paul Hoffman. The Internet instant reference. The Sybex instant reference series. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1995. ISBN 0-7821-1719-8. xvi + 346 pp. LCCN TK5105.875.I57H64 1995.

Hoffman:1995:IMB

[Hof95g] Paul Hoffman. Internet manual de bolsillo. Serie McGraw-Hill de informatica. McGraw-Hill, New York, NY, USA, 1a en espanol. edition, 1995. ISBN 970-10-0759-X (paperback). xiv + 378 pp. LCCN ????

Hoffman:1996:WHE

[Hof96a] David S. Hoffman. The Web of hate: extremists exploit the Internet. Technical report, Anti-Defamation League, New York, NY, USA, 1996. 60 pp.

Hoffman:1996:HIB

[Hof96b] Kurt Hoffman. Holes in the Internet: business has barely tapped the system’s potential, but some major security problems remain unsolved. Distribution, 95(6): 29–34, March 1996. ISSN 1066-8489.

Holledge:1994:ANI

[Hol94] Simon Holledge. Archaeology on the net: an Internet resource list. Wessex International Archaeology, Cambridge, England, 1994. ISBN 0-9524404-0-7. 79 pp. LCCN ????

Homer:1999:XIP

[Hom99]  Alex Homer.  XML in

IE5: programmer’s reference. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-157-6. xi + 479 pp. LCCN ????

Honeycutt:1995:UIW

[Hon95]  Jerry Honeycutt.  Using

the Internet with Windows 95. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0402-1 (paperback). xxvii + 376 pp. LCCN TK5105.875.I57H66 1995.

Harnad:1995:SJA

[HOO95] Stevan R. Harnad, Ann Okerson, and James Joseph O’Donnell. Scholarly journals at the crossroads: a subversive proposal for electronic publishing: an Internet discussion about scientific and scholarly journals and their future. Technical report, Office of Scientific and Academic Publishing, Association of Research Libraries, Washington, DC, USA, 1995. 242 pp.

Hopkins:1995:ILB

[Hop95] Gerald L. Hopkins. ISDN Literacy Book. Addison-Wes-ley, Reading, MA, USA, 1995. ISBN 0-201-62979-8. xvi + 367 pp. LCCN TK 5103.75 H67 1995. US$31.95.

Horan:1995:CEE

[Hor95] Thomas F. Horan. The converging elements of electronic commerce: the internet, multimedia messaging, and groupware. Technical report, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, 1995. 11 pp.

Horneffer:1998:MPA

[Hor98] Martin Horneffer. Methods for performance-analysis of Internet access points. Computer Networks and ISDN Systems, 30(16–18):1607– 1615, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/16-18/1977.pdf.

Howard:1995:IIS

[How95a] Garry Howard. Introduction to Internet security: from basics to beyond. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 1-55958-747-4. xxiv + 414 pp. LCCN TK5105.875.I57H69 1995.

Howard:1995:IV

[How95b] Jim Howard. The Internet voyeur. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1655-8. xxiii + 388 pp. LCCN T385.H7 1995. US$19.19.

Howard:1995:IVG

[How95c]  Jim Howard.  Internet

voyeur: le guide du chasseur d’images. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1555-4. xix + 355 pp. LCCN ????

Haskin:1996:QI

[HP96] David Haskin and Gail Perry.

Quicken 5 on the Internet. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-448-0. xxx + 432 pp. LCCN HG179.H319 1996.

Huynh:1995:CCM

[HPS95] Minh Huynh, Laird Popkin, and Matthew Stecker. Constructing a corporate memory infrastructure from Internet discovery technologies. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

????  URL http://www.
ncsa.uiuc.edu/SDG/IT94/
Proceedings/CorInfSys/huynh/

cmi.html. Includes CD-ROM.

Hart:1992:BIIb

[HRB92] Jeffrey A. Hart, Robert R. Reed, and François Bar. The building of the Internet: implications for the future of broadband networks. BRIE working paper 60, Berkeley Roundtable on the International Economy, University of California, Berkeley, Berkeley, CA, USA, 1992. 666–689 pp.

Hahn:1994:ICR

[HS94a] Harley Hahn and Rick Stout. The Internet Complete Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-881980-6. xxvi + 818 pp. LCCN TK5105.875.I57 H3 1994. US$29.95. Highly recommended. This is by far the most complete, illuminating, useful, exciting, readable :-) book on the Internet …. Covering mail, telnet,

finger, usenet, ftp, talk,

IRC, gopher, Wais, WWW, etc., …. And there is also a Catalog of Internet Resources (over 140 pages) ….

Hahn:1994:IGD

[HS94b] Harley Hahn and Rick Stout. The Internet golden directory: published in the United States as The Internet yellow pages. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882082-0. 447 pp. LCCN ????

Hahn:1994:IMR

[HS94c] Harley Hahn and Rick Stout.

Internet, manual de referencia. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 84-481-1882-0. xxv + 692 pp. LCCN ????

Hahn:1994:IYP

[HS94d] Harley Hahn and Rick Stout. The Internet Yellow Pages. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882023-5. xlv + 447 pp. LCCN TK 5105.875 I57 H343 1994. US$27.95. A catalog of hundreds of Internet resources in phone-book yellow-page format …. The Internet Yellow Pages strives to be an unmatched subject-oriented index to the Internet’s resources, a lofty goal given the number of competing indexes available. Hahn and Stout do it right — this book is a resource that long-time Internet addicts and new users will enjoy equally well. Organization is intuitive and consistent. Although it does not point you to as many online resources as some other directories, the ones it points to tend to be high-quality and truly interesting sites to visit. Well-researched and well-indexed.

Hahn:1995:IGD

[HS95a] Harley Hahn and Rick Stout. The Internet golden directory. Osborne, Berkeley, CA, USA, second edition, 1995. ISBN 0-07-882107-X (paperback). xxxvi + 812 pp. LCCN ????

Hahn:1995:IYP

[HS95b] Harley Hahn and Rick Stout. The Internet yellow pages. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1995. ISBN 0-07-882098-7. xxxvi + 812 pp. LCCN TK5105.875 I57 H35 1995.

Hahn:1995:YPL

[HS95c] Harley Hahn and Rick Stout. Yellow pages: l’annuaire Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1538-4. xxxv + 627 pp. LCCN ????

Harris:1998:LOS

[HS98] Jacob Harris and Vivek Sarkar. Lightweight objectoriented shared variables for distributed applications on the Internet. ACM SIG-PLAN Notices, 33(10):296– 309, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hahn:1994:IMK

[HSF94] Harley Hahn, Rick Stout, and Michael Frasse. Internet Membership Kit: Access and Tools: Your On-ramp to the Information Superhighway: Macintosh version. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-171-6. LCCN TK5105.875.I57 F72 1994. US$69.95. Bundle of Internet Yellow Pages and Internet Tour Guide. Includes application software for Mac or Windows.

Hansen:1995:HRI

[HSHF95] Stephen Hansen, Daniel Salcedo, Kenneth W. Hunter, and Eben Friedman. Human rights information sources on the Internet. Technical report, American Association for the Advancement of Science, Washington, DC, USA?, 1995. 21 pp.

Hardman:1998:SMA

[HSK98] Vicky Hardman, Martina Angela Sasse, and Isidor Kouvelas. Successful multiparty audio communication over the Internet. Communications of the ACM, 41(5): 74–80, May 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-5/
p74-hardman/.

Hahn:1995:IWC

[HSskw95] Harley Hahn, Rick Stout, and Po shuo ku wen. Internet wan cheng tsan kao shou tse. Po shuo ku wen “ching tien hsi lieh”. Maiko-lo-hsi-erh chu pan, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-8967-98-5, 957-8967-99-3. LCCN ????

Henri:1995:IGA

[HT95] Philip Tsang James Henri and Sandy Tse. Internet growth in Australia and Asia’s Four Dragons. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

                                        ????  URL http://www.
                                        ncsa.uiuc.edu/SDG/IT94/
                                        Proceedings/Int/tsang/tsang.

html. Includes CD-ROM.

Hampton:1994:IIS

[HTW94] Tim Hampton, Tiffany Taylor, and Linda Watkins. Inside the Internet starter guide. The Cobb Group, Louisville, KY, USA, 1994. ISBN 0-936767-22-7. 40 pp. LCCN ????

Harris:1995:FIE

[HU95] Ron Harris and David Ullendorff. Futurekids: the Internet expedition. Ziff-Davis Press, Emeryville, CA, USA, 1995. ISBN 1-56276-271-0. various pp. LCCN ????

Hubbel:1993:FLC

[Hub93] Hubbel Premise Wiring Inc., 14 Lord’s Hill Road, PO Box 901, Stonington, CT 06378-0901, USA, Tel: (203) 535-3809 or (800) 626-0005. Full Line Catalog No. 1100R, 1993.

Hubbard:1996:BDM

[Hub96] Jordan K. Hubbard, editor.

BSD Docs: the most complete collection of BSD-UNIX related articles and documents from the Internet. Walnut Creek CDROM, Walnut Creek, CA, USA, 1996. ISBN 1-57176-126-8. 4 pp. LCCN ???? Includes CD-ROM.

Hudson:1994:PNP

[Hud94] Hoyt Hudson. Plug n’ play Internet: release 2.0 for Microsoft Windows. InterAcess Co., Chicago, IL (9400 W. Foster Ave., Suite 111, Chicago 60656), 1994. various pp. LCCN TK5105.875.I57 H76 1994.

Hughes:1994:IPA

[Hug94] Phil Hughes. Internet public access guide: a guide for accessing the Internet through Unix-based computer systems. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1994. ISBN 0-916151-70-0. 63 pp. LCCN ???? Ideal for new users of UNIX-BASED Internet providers ….

Hughes:1995:AUI

[Hug95] Larry J. Hughes. Actually useful Internet security techniques. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-508-9. xv + 378 pp. LCCN TK5101.59 .H83 1995. “Covers SATAN and spoofing”.

Huitema:1995:RI

[Hui95] Christian Huitema. Routing in the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-132192-7. xii + 319 pp. LCCN TK5105.875.I57 H85 1995. US$42.00.

Huitema:1996:INI

[Hui96] Christian Huitema. IPv6: the new Internet protocol. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-241936-X. vii + 188 pp. LCCN TK5105.875.I57 H78 1996.

Huitema:1999:INI

[Hui99] Christian Huitema. IPv6: the new Internet protocol. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-850505-5. viii + 247 pp. LCCN TK5105.585 .H85 1998.

HRI:1976:HRI

[Hum76] Human Rights Internet newsletter, 1976. ISSN 0163-9048. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada.

HRI:1980:HRI

[Hum80] Human Rights Internet reporter, page various, 1980. ISSN 0275-049X; 0163-9048. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada.

HRI:1989:AHR

[Hum89] Human Rights Internet reporter. Africa: human rights directory and bibliography. Technical report, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1989. 308 pp.

HRI:1994:LOC

[Hum94] Human Rights Internet. A listing of organizations concerned with human rights and social justice worldwide: masterlist. Technical report, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1994. 278 pp.

Hume:1998:BRS

[Hum98] Andrew Hume. Book review: Sun Performance and Tuning: Java and the Internet. ;login: the USENIX Association newsletter, 23 (6):??, October 1998. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/1998-10/sun.html.

Hunt:1992:TIN

[Hun92] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-82-X (paperback). xxii + 471 pp. LCCN TK5105.9 .H86.

Hunter:1993:ICT

[Hun93]  Beverly Hunter.  Inter-

networking: coordinating technology for systemic reform. Communications of the ACM, 36(5):42–46, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
155057.html.

Hunt:1998:TIN

[Hun98] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1998. ISBN 1-56592-322-7. xv + 612 pp. LCCN TK5105.55 .H86 1998. US$32.95. URL http://www.

ora.com/catalog/tcp2/;
http://www.oreilly.com/
catalog/tcp2.

Huston:1994:DSD

[Hus94] G. Huston. Desperately seeking default: Internet policy update: A perspective from the Pacific. Computer Networks and ISDN Systems, 26 (Supplement 1):S3–??, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Hayworth:1995:LVS

[HV95] Learning Visions subject guide to the Internet, page various, 1995. Learning Visions, Boulder, CO, USA.

HISC:1995:HOA

[HW95] Hobbes Internet Software Collection and Walnut Creek CDROM. Hobbes OS/ 2 archived thousands of OS/ 2 programs direct from the Hobbes Internet Software Collection, 1995. ISBN 1-57176-037-7. Includes CD-ROM.

Hamalainen:1996:EML

[HWV96] Matti Hämäläinen, Andrew B. Whinston, and Svetlana Vishik. Electronic markets for learning: Education brokerages on the Internet. Communications of the ACM, 39(6):51–58, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228513.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228513.html.

Hoffman:1994:IIRb

[HY94] Paul Hoffman and Noam Yoran. Internet instant reference: [madrikh bazak]. Bag multisistem, Bene Berak, Israel, 1994. ISBN ???? 277 pp. LCCN ????

Ibe:1997:EAN

[Ibe97] Oliver C. (Chukwudi) Ibe. Essentials of ATM Networks and Services. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-18461-3. xv + 224 pp. LCCN TK5105.35.I24 1997. US$22.95.

IBM:1988:ITC

[IBM88] IBM Corporation, San Jose, CA, USA. IBM transmission control protocol/Internet protocol for VM: command reference manual, third edition, September 1988. various pp.

IBM:1989:ITCa

[IBM89a]  IBM Corporation.  IBM

Transmission Control Protocol/Internet Protocol for VM: network file system and remote procedure call manual. Technical Report SC09-1274-01, IBM Corporation, Armonk, NY, USA, January 1989. viii + 58 pp.

IBM:1989:ITCb

[IBM89b]  IBM Corporation.  IBM

Transmission Control Protocol/Internet Protocol for VM: programmer’s manual. Technical Report GC09-1206-03, IBM Corporation, Armonk, NY, USA (??), January 1989. various pp.

IBM:1989:IIT

[IBM89c] IBM Corporation, San Jose, CA, USA. Introducing IBM’s Transmission Control Protocol/Internet Protocol products, 1989. various pp.

IBM:1989:TCPb

[IBM89d] IBM Corporation, Mechanicsburg, PA, USA. Transmission Control Protocol/ Internet Protocol for MVS: installation and maintenance, June 1989. various pp.

IBM:1989:TCPc

[IBM89e] IBM Corporation, Mechanicsburg, PA, USA. Transmission Control Protocol/ Internet Protocol for MVS: programmer’s reference, June 1989. various pp.

IBM:1989:TCPa

[IBM89f] IBM Corporation, Mechanicsburg, PA, USA. Transmission Control Protocol/ Internet Protocol for MVS: user’s guide, June 1989. various pp.

IBM:1990:ITC

[IBM90] IBM Corporation, San Jose, CA, USA. IBM transmission control protocol/Internet protocol version 2 for programmer’s reference, 1990. xxii + 552 pp.

IBM:1991:ITCd

[IBM91a] IBM Corporation, Research Triangle Park, NC, USA.

IBM transmission control protocol/Internet protocol version 2 release 2 for VM: messages and codes, second edition, 1991. xiv + 174 pp.

IBM:1991:ITCb

[IBM91b] IBM Corporation, Research Triangle Park, NC, USA. IBM transmission control protocol/Internet protocol version 2 release 2 for VM: programmer’s reference, second edition, 1991. xxii + 560 pp.

IBM:1992:ITC

[IBM92a] IBM Corporation, San Jose, CA, USA. IBM transmission control protocol/Internet protocol: TCP/IP version 2 release 2.1 for MVS: user’s guide, third edition, 1992. xx + 301 pp.

IBM:1992:TCP

[IBM92b] IBM Corporation. Transmission control protocol/ Internet protocol: TCP/IP version 2 release 2.1 for MVS: planning and customization. Technical Report SC31-6085-02, IBM Corporation, San Jose, CA, USA, September 1992. xxii + 409 pp.

IBM:1993:TIP

[IBM93] IBM Corporation, San Jose, CA, USA. TCP/IP programmer’s toolkit, version 2.0 for OS/2 edition, 1993. xiv + 361 pp. 2 computer disks + 1 kit installation booklet (11 p.) + 1 programmer’s reference.

IETF:1990:PNI

[IC90] Internet Engineering Task Force and Corporation for National Research Initiatives, editors. Proceedings of the Ninteenth Internet Engineering Task Force, University of Colorado, December 3–7, 1990. Corporation for National Research Initiatives, Reston, VA, USA, 1990. ISBN ???? LCCN ????

IETF:1992:PIE

[IC92] Internet Engineering Task Force and Corporation for National Research Initiatives, editors. Proceedings of the … Internet Engineering Task Force. Corporation for National Research Initiatives, Reston, VA, USA, 1992. ISBN ???? LCCN ????

IMO:1994:IEI

[IC94] Information Market Observatory and Commission of European Communities. Directorate General, Telecommunications, Information Market and Exploitation of Research. The Internet and the European information industry. IMO Working Paper 94/3, European Commission Directorate General Telecommunications Information Market and Exploitation of Research, Luxembourg, Luxembourg, September 1994. 15 pp.

Ikeda:1998:SAT

[IDK98] Kazuyuki Ikeda, Matsuyuki Doi, and Tomiei Kazama, editors. State-of-the-art technology in anesthesia and intensive care: proceedings of the 18th International Symposium on Computing in Anesthesia and Intensive Care held in Hamamatsu city, Japan on 18–21 March 1998, volume 1168 of International congress series. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1998. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

Ihlenfeldt:1998:GID

[IE98] Wolf-D. Ihlenfeldt and Klaus Engel. Graphics in/for digital libraries — visualizing chemical data in the Internet — data-driven and interactive graphics. Computers and Graphics, 22(6): 703–714, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/6/625.pdf.

IEEE:1990:ISC

[IEE90] IEEE. 802.3i-1990 IEEE Supplement to Carrier Sense Multiple Access with Collision Detection CSMA/ CD Access Method and Physical Layer Specifications: System Considerations for Multisegment 10 Mb/s Baseband Networks (Section 13) and Twisted-Pair Medium Attachment Unit (MAU) and Baseband Medium, Type 10BASE-T (Section 14). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 1-55937-070-X. US$26.00. IEEE product number SH13763.

IEEE:1994:PCL

[IEE94] IEEE Computer Society.

Technical Committee on Computer Communications, editor. Proceedings / 19th Conference on Local Computer Networks, October 2– 5, 1994, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

IEEE:1995:PCL

[IEE95] IEEE Computer Society.

Technical Committee on Computer Communications, editor. Proceedings: 20th Conference on Local Computer Networks, October 16– 19, 1995, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

IEEE:1997:PAC

[IEE97] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings, 22nd annual Conference on Local Computer Networks: LCN ’97: November 2– 5, 1997, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

IEEE:1998:LPA

[IEE98] IEEE Computer Society. Technical Committee on Computer Communications, editor. LCN’98: proceedings: 23rd Annual Conference on Local Computer Networks: October 11–14, 1998, Lowell, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Ivler:1998:CDR

[IH98] J. M. Ivler and Kamrah Husain. CGI Developer’s Resource: Web programming in Tcl and Perl. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-727751-2. xx + 597 pp. LCCN TK5105.565.I95

1997. URL http://www.
phptr.com/ptrbooks/ptr_
0137277512.html.

Ishida:1993:ISI

[IL93] Haruhisa Ishida and Lawrence H.

Landweber. Introduction to the special issue on internetworking. Communications of the ACM, 36(8): 28–30, August 1993. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
163382.html.

ILAA:1991:UIH

[Ill91] Illinois Library Association, Academic, College, and Research and Libraries Forum. Using the Internet: a handson workshop. University of Illinois, Graduate School of Library and Information Science, Urbana, IL, USA, September 27, 1991. 74 pp.

IETF:1994:PTI

[ILSB94] Internet Engineering Task Force, Debra J. Legare, John W. Stewart, III, and Marcia J. Beaulieu, editors. Proceedings of the thirtieth Internet Engineering Task Force: University of Toronto, ON, Canada, July 25–29, 1994. Corporation for National Research Initiatives, Reston, VA, USA, 1994. ISBN ???? LCCN ????

Iwata:1995:ACT

[IMI^{+}95] A. Iwata, N. Mori, C. Ikeda, H. Suzuki, and M. Ott. ATM connection and traffic management schemes for multimedia Internetworking. Communications of the ACM, 38(2):72–89, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        204846.html.

InfoWorld:1980:I

[Inf80] InfoWorld, 1980. CODEN INWODU. ISSN 0199-6649. Infoworld Publishing Co., 155 Bovet Road Suite 800, San Mateo, CA 94402, USA, Tel:

(415) 572-7341 or (800) 227-8365.

IAC:1993:UGI

[Inf93] Information Access Company, Foster City, CA, USA.

User guide Infotrac Central 2000: Internet access to the Information Access Center, 1993. various pp.

IARI:1995:IGB

[Inf95a] InfoAdvantage guide: the best of strategic and competitive intelligence on Internet, 1995. InfoAdvantage Research International, Bellevue, WA, USA.

Anonymous:1995:IMI

[Inf95b] Infobahn: the magazine of Internet culture, page various, 1995. ISSN 1079-4069. Postmodern Communications, Inc., Foster City, CA, USA.

IS:1995:SOI

[Inf95c] InforSource, Inc. Seminar ona-disk Internet fundamentals, 1995. 3 computer disks.

IARI:1996:IIG

[Inf96] InfoAdvantage Internet guide, page various, 1996. InfoAdvantage Research International, Bellevue, WA, USA.

IGI:1996:EIU

[Inn96] Innovation Groups, Inc. Email and Internet use policies. Research packet (innovation groups), Innovation Groups, Tampa, FL, USA, 1996.

InPuT:1995:ISI

[INP95] Internet security: the impact of firewalls on Client/ Server applications, page various, 1995. INPUT, Mountain View, CA, USA.

IRC:1996:CCM

[Ins96] Institute for Research. Career as a computer multimedia developer, designer: (CD-ROM, internet). Careers research monographs 321, Institute for Research, Chicago, IL, USA, 1996. 24 pp.

Anonymous:1976:IPF

[Int76a] Internet 76: [proceedings of the] Fifth Internet World Congress, Project Implementation and Management: Bridging the Gap, [held at the] Metropole Hotel at the National Exhibition Centre, Birmingham, UK, September 13th to 17th 1976. NCC Publications, Manchester, UK, 1976. ISBN 0-85012-182-5. LCCN T56.8.I57 1976.

INTERNET:1976:IPF

[Int76b] Internet 76: Proceedings of the Fifth Internet World Congress, Project Implementation and Management: Bridging the Gap, held at the Metropole Hotel at the National Exhibition Centre, Birmingham, UK, September 13th to 17th, 1976. NCC Publications, Manchester, UK, 1976. ISBN 0-85012-182-5. LCCN T56.8 .I57 1976. Five volumes.

HRI:1981:IAF

[Int81] Human Rights Internet. The Internet at five: a five year report of the human rights Internet. Technical report, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1981. 16 pp.

IRG:1984:IMR

[Int84] Internet Research Group. Internet monthly reports, 1984.

IICHG:1985:IHG

[Int85] Internet on the Holocaust and genocide: an international information resource exchange towards understanding, intervention and prevention of genocide, page various, 1985. Institute of the International Conference on the Holocaust and Genocide; Denison University, Jerusalem, Israel.

HRI:1986:HRI

[Int86] Human Rights Internet reporter, page various, 1986. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada.

HRI:1987:HRI

[Int87] Human Rights Internet directory, page various, 1987. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada.

Internet:1988:IPI

[INT88] INTERNET … profiles of international development contractors and grantees, page various, 1988. ISSN 1051-5283. Network for International Technical Assistance, Chapel Hill, NC, USA.

IALCD:1989:ILC

[Int89a]  Internet-accessible  library

catalogs and databases, page various, 1989. Art St. George and Ron Larsen, ????

Internet:1989:ISP

[INT89b] INTERNET … summary profiles of international development contractors and grantees, page various, 1989. Network for International Technical Assistance, Inc., Chapel Hill, NC, USA.

IETF:1991:IEP

[Int91] Internet Engineering Task Force. Internet Engineering Steering and Group, ???? An Internet evolution plan for the IETF, 1991. 87 pp.

IBM:1992:TIN

[Int92a] International Business Machines Corporation. TCP/IP network file system kit, 1992. 2 computer disks and booklet (6 p.).

Anonymous:1992:I

[Int92b] Internet, page various, 1992. SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028.

LSSI:1992:I

[Int92c] Internet, 1992. Library Systems Services, Inc (LSSI), Gaithersburg, MD, USA.

HRI:1992:PIA

[Int92d] Human Rights Internet.

Peru: impact of the April 1992 coup. Occasional papers = document hors-serie occasional papers (human rights internet), Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, August 1992. ii + 21 pp.

SOLINET:1992:IN

[Int92e] Internet news, page various, 1992. Southeastern Library Network (SOLINET), Atlanta, GA, USA.

Anonymous:1992:IWb

[Int92f] Internet world, page various, 1992. ISSN 1064-3923. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA.

WETC:1993:IKC

[Int93a] Integrating the K–12 curriculum with the Internet: stepby-step, lesson plans and lesson ideas, project plans and project ideas for elementary, middle and high school. Wisconsin Educational Technology Conference, Appleton, WI, USA, 1993. ISBN ???? LCCN ????

IBM:1993:TIA

[Int93b] International Business Machines Corporation. TCP/IP applications kit, 1993. 7 computer disks + 1 kit installation booklet (11 p.).

IBM:1993:TIDb

[Int93c] International Business Machines Corporation. TCP/ IP domain name server, 1993. Includes 2 computer disks and installation booklet (11 p.).

IBM:1993:TIDa

[Int93d] International Business Machines Corporation. TCP/ IP DOS/Windows access kit, 1993. Includes 2 computer disks and installation booklet (11 p.).

IBM:1993:TIT

[Int93e] International Business Machines Corporation. TCP/ IP total kit, 1993. 17 computer disks + 1 kit installation booklet (11 p.) + 1 installation and administration manual.

IBM:1993:TIX

[Int93f] International Business Machines Corporation. TCP/IP X window system server kit, 1993. 7 computer disks + 1 kit installation booklet (11 p.).

IBR:1993:IBR

[Int93g] Internet business report, 1993. CMP Publications, Manhasset, NY, USA.

Anonymous:1993:ILa

[Int93h] Internet librarian, 1993. ISSN 1069-4463. Meekler Pub., Westport, CT, USA.

Anonymous:1993:IR

[Int93i] Internet research, page various, 1993. ISSN 1066-2243. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA.

HRI:1994:ERR

[Int94a] Human Rights Internet. Ethnic, religious and racial conflict: bibliographic abstracts, 1992–93. Human Rights Internet reporter; v. 15, no. 1, Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1994. 212 pp.

HRI:1994:GUF

[Int94b] Human Rights Internet.

Ghana: update on the Fourth Republic. Occasional papers = document horsserie / Human Rights Internet (HRI), Documentation, Information and Research Branch, Immigration and Refugee Board, Canada, Ottawa, ON, Canada, September 1994. iii + 23 pp.

WCS:1994:IY

[Int94c] The Internet and you, 1994.

2 videocassettes.

Anonymous:1994:IBC

[Int94d] Internet bulletin for CPAs, page various, 1994. ISSN 1078-2176. Kent Information Services, Kent, OH, USA.

Anonymous:1994:IBA

[Int94e] Internet business advantage: online solutions for business success, page various, 1994. ISSN 1078-6422. Wentworth Worldwide Media, Inc., Lancaster, PA, USA.

IBA:1994:IBD

[Int94f] Internet business directory, page various, 1994. ISSN 1078-9936. Internet Business Association, Cambridge, MA, USA.

IBA:1994:IIB

[Int94g] Internet international business directory, page various, 1994. ISSN 1078-9928. Internet Business Association, Cambridge, MA, USA.

Anonymous:1994:IBM

[Int94h] Internet business magazine, page various, 1994. ISSN 1079-0705. World Tech Press, Cambridge, MA, USA.

IES:1994:IYT

[Int94i] Internet Educational Services. The Internet and you: training workbook. Technical report, Winnebago Software Co., Caledonia, MN, USA, 1994. xv + 29 pp.

Anonymous:1994:IM

[Int94j] Internet marketing, page various, 1994. ISSN 1070-1028. Robert L. Nidever, Los Angeles, CA, USA.

Anonymous:1994:IV

[Int94k] Internet voyager, page various, 1994. ISSN 1078-540X. Blue Dolphin Communications, Sudbury, MA, USA.

IWI:1994:IWI

[Int94l] Internet world’s on Internet, 1994. ISSN 1066-9973. Mecklermedia, Westport, CT, USA.

IQPCM:1995:HMY

[Int95a] International Quality and Productivity Center Meeting. How to market your healthcare products and services on the Internet, 1995. 14 sound cassettes.

IREB:1995:FFA

[Int95b] International Research and Exchanges Board, Washington, DC, USA. FAQ: frequently asked questions: email and Internet in the NIS and Baltics, 1995. 24 pp.

Anonymous:1995:ICT

[Int95c] Internet and comms today, page various, 1995. ISSN 1355-5219. Paragon Publishing, Ltd., Bournemouth, UK.

Anonymous:1995:IC

[Int95d] The Internet connection, page various, 1995. ISSN 1080-8493. Bernan Press, Lanham, MD, USA.

Anonymous:1995:ILR

[Int95e] Internet law report, page various, 1995. Zacharia Publishing, McKeesport, PA, USA.

Anonymous:1995:ILb

[Int95f] The Internet lawyer, page various, 1995. ISSN 1087-7223. GoAhead Publications, Gainesville, FL, USA.

InternetMagazine:1995:NPH

[Int95g]  Internet Magazine.  .net

presents: How to make light of HTML: the plain English guide to creating your own Web site. Technical report, Future Publishing, Bath, England, 1995. 98 pp.

Anonymous:1995:IMT

[Int95h] Internet marketing and technology report, page various, 1995. ISSN 1082-1945. Computer Economics, Carlsbad, CA, USA.

Anonymous:1995:INb

[Int95i] Internet news, 1995. ISSN 1081-0048. William G. Behr, Arlington, TX, USA.

Anonymous:1995:INa

[Int95j] The Internet newsroom, 1995. Editors’ Service, Glen Echo, MD, USA.

IP:1995:LB

[Int95k] Internet Productions. Learn the basics, 1995. 1 videocassette (ca. 71 min.) disks.

Anonymous:1995:IP

[Int95l] Internet professional, page various, 1995. ISSN 1071-278X. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA.

Anonymous:1995:ISEb

[Int95m] Internet strategies for education markets: the Heller report, page various, 1995. ISSN 1083-1428; 1047-5230. N.B. Heller and Associates, Highland Park, IL, USA.

Anonymous:1995:IUc

[Int95n] Internet underground, page various, 1995. ISSN 1084-4805. Sendai Pub. Group, Lombard, IL, USA.

Anonymous:1995:IWW

[Int95o] Internet: what’s working for business, page various, 1995. ISSN 1082-1880. Computer Economics, Carlsbad, CA, USA.

InterNIC:1995:SRS

[Int95p] InterNIC. The scout report:

a service to the Internet community provided by the Inter-NIC, 1995.

IntraTrade:1995:PGI

[Int95q] Intra Trade Corporation. PC guide for Internet and online services: the easiest way to learn to use Internet and on-line services. Inter Trade Corp., Norcross, GA, USA, 1995. ISBN 1-881979-21-0. xii + 304 pp. LCCN ????

IPLA:1996:TCI

[Int96a] Intellectual Property Law Association of Chicago Domestic-Foreign Trademark Committee. Trademark and copyright issues on the Internet, 1996.

Anonymous:1996:IA

[Int96b] Internet advisor, page various, 1996. ISSN 1084-158X. Advisor Publications, Inc., San Diego, CA, USA.

Anonymous:1996:IDJ

[Int96c] Internet developer’s journal, page various, 1996. ISSN 1086-8569. Sys-Con Publications, Jersey City, NJ, USA.

Anonymous:1996:II

[Int96d] Internet InfoScavenger, page various, 1996. ISSN 1088-0666. InfoScavenger Communications, Evergreen, CO, USA.

IJAS:1996:IJA

[Int96e] Internet journal of African studies, 1996. ISSN 1363-2914. University of Bradford. Dept. of Social and Economic Studies, West Yorkshire, UK.

Anonymous:1996:ILR

[Int96f] Internet law researcher, page various, 1996. ISSN 1087-7703. Glasser LegalWorks, Little Falls, NJ, USA.

Anonymous:1996:ILN

[Int96g] Internet litigation news, page various, 1996. Litigation Reporting Service, Chalfont, PA, USA.

Anonymous:1996:IMa

[Int96h] Internet magazine, page various, 1996. ISSN 1082-7560. Internet Magazine, Torrance, CA, USA.

Anonymous:1996:IMb

[Int96i] Internet medicine, page various, 1996. ISSN 1086-5691. Lippincott-Raven Publishers, Philadelphia, PA, USA.

Anonymous:1996:IRSa

[Int96j] Internet reference services quarterly, page various, 1996. ISSN 1087-5301. Haworth Press, New York, NY, USA.

Anonymous:1996:ISP

[Int96k] Internet service providers:

quarterly directory, page various, 1996. Boardwatch Magazine, Littleton, CO, USA.

Anonymous:1996:ITa

[Int96l] Internet today, page various, 1996. ISSN 1355-5219. Paragon Publishing, Ltd., Bournemouth, Eng.

Anonymous:1996:I

[Int96m] InternetUser, page various, 1996. ISSN 1072-4168. Meckler Pub., Westport, CT, USA.

IDGS:1994:ICN

[Iow94] Iowa. Dept. of General Services, Division of Communication, Des Moines, IA, USA. The Iowa Communications Network: commonly asked questions about Part III sites, 1994. 10 pp.

IRCAM:1995:IMA

[IRC95] IRCAM (Research institute). Internet MIDI archives, 1995.

InterAction:1995:EIB

[IV95] InterAction and VITA. The essential Internet: basics for NGOs. Technical report, InterAction, Washington, DC, USA, 1995. 44 pp.

Ivens:1995:IWB

[Ive95] Kathy Ivens. The Internet Warp book: your complete guide to getting Online with OS/2 Warp. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 0-7615-0213-0. xv + 320 pp. LCCN TK5105.875.I57I94 1995.

JLearning:1995:IGC

[J3 95] J3 Learning Corporation. The Internet global communications, 1995. 1 videocassette (50 min.) (3 1/2 in.) + 1 course book (v, 98 p.).

Jacobson:1994:MCI

[Jac94] Van Jacobson. Multimedia conferencing on the Internet. Technical report, University College London, London, UK, 1994. 74 pp.

Jaffe:1994:IIP

[Jaf94] Lee David Jaffe. Introducing the Internet plus: a trainer’s workshop. Internet workshop series; no. 1. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, second printing, minor revisions edition, 1994. ISBN 1-882208-05-6. viii + 92 pp. LCCN TK5105.875.I57J33 1994.

Jain:1994:FH

[Jai94] Raj Jain. FDDI Handbook:

High-speed Networking using Fiber and Other Media. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-56376-2. xxvii + 528 pp. LCCN TK5105.72 .J35 199. US$49.50.

Jamsa:1994:IIA

[Jam94] Kris Jamsa. Instant Internet Access. Jamsa Press, Las Vegas, NV, USA, 1994. ISBN ???? LCCN ???? US$49.95.

James:1996:ONNa

[Jam96a] Phil James. Official Netscape Navigator 2.0 book: the definitive guide to the world’s most popular Internet navigator. Ventana Press, Chapel Hill, NC, USA, Macintosh edition, 1996. ISBN 1-56604-413-8. various pp. LCCN TK5105.882.J35 1996.

James:1996:ONNb

[Jam96b] Phil James. Official Netscape Navigator 2.0 book: Windows edition: the definitive guide to the world’s most popular Internet navigator. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-347-6. xxvi + 657 pp. LCCN TK5105.882.J35 1995.

Jander:1989:NN

[Jan89] Mary Jander. The naked network. Computer decisions (Teaneck, NJ: 1988), 21(4): 39–??, April 1, 1989. CODEN CODCB8, COMDEP. ISSN 0010-4558, 0898-1825.

Janal:1995:OMH

[Jan95] Daniel S. Janal. Online marketing handbook: how to sell, advertise, publicize, and promote your products and services on the Internet and commercial online systems. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-02058-9. xxviii + 370 pp. LCCN HF6146.I58 J36 1995.

Janal:1996:BYC

[Jan96] Daniel S. Janal. 101 businesses you can start on the Internet. Van Nostrand Reinhold, New York, NY, USA, 1996. ISBN 0-442-02202-6, 0-442-01938-6. xxi + 504 pp. LCCN HD62.5 .J36 1996.

Jarillo:1995:SNC

[Jar95] José Carlos Jarillo. Strategic Networks: Creating the borderless organization. Butterworth-Heinemann, Boston, MA, USA, 1995. ISBN 0-7506-2327-6. xii + 178 pp. LCCN ???? US$49.99.

Jamsa:1995:IP

[JC95] Kris A. Jamsa and Ken Cope.

Internet programming. Jamsa Press, Las Vegas, NV, USA, 1995. ISBN 1-884133-12-6 (paperback). xi + 588 pp. LCCN ACQUISITION.

Jamsa:1996:PIC

[JCD96] Kris Jamsa, Ken Cope, and Eric. Trad Dupas. Programmation Internet en C et C++. International Thomson Publishing, London, New York, Boston, Madrid, ldots, 1996. ISBN 2-84180-034-2. x + 645 pp. LCCN ????

Junor:1995:IUGa

[JD95a] Bill Junor and Chris DeMontravel. Internet: the user’s guide for everyone. Branden Pub. Co., Boston, MA, USA, 1995. ISBN 0-8283-2007-1. 144 pp. LCCN TK5105.875.I57 J86 1995b General Coll.

Junor:1995:IUGb

[JD95b] Bill Junor and Chris DeMontravel. Internet: the user’s guide for everyone. Branden, Boston, MA, USA, second, enlarged and updated. edition, 1995. ISBN 0-8283-2013-6. 218 + 4 pp. LCCN TK5105.875.I57J86 1995b.

Jarabo:1995:ICD

[JE95] Francisco Jarabo and Nicolas Elortegui. Internet: conexion desde el PC Domestico a ordenadores de todo el mundo. Paraninfo, Madrid, Spain, 1995. ISBN 84-283-2141-8 (paperback). 115 pp. LCCN ????

Johnson:1996:SIB

[JE96] Peggy Johnson and Lee English. The searchable Internet bibliography an on-disk annotated guide to timely materials about Internet, 1996. ISBN 0-8389-2189-2. 3 computer disks.

Jensen:1992:EIJ

[JEK92] Mary Brandt Jensen, Carol Ebbinghouse, and Brendan P. Kehoe. E-mail isn’t just a substitute for snail mail: a hitchhiker’s guide to the nets. American Association of Law Libraries, Chicago, IL, USA, 1992. 184 pp.

James:1996:FGI

[JFJ96] Henry James, David Freshwater, and Farm Journal. The Farmer’s guide to the Internet. TVA Rural Studies, Lexington, Ky., second edition, 1996. ISBN 0-9649746-1-4. xi + 334 pp. LCCN ????

Jackson:1991:PSA

[JG91] Lenore A. Jackson and J. Patrick Gary, editors. Proceedings of the Second Annual NASA Science Internet User Working Group Conference: proceedings of a conference held in San Mateo, California, February 11–14, 1991, number 3117 in NASA conference publication. National Aeronautics and Space Administration, Washington, DC, USA, 1991. LCCN TK5105.875.I57 N37 1991.

Jacobs:1996:FHF

[JGKR96] Stephan Jacobs, Michael Gebhardt, Stefanie Kethers, and Wojtek Rzasa. Filling HTML forms simultaneously: CoWeb-architecture and functionality. Computer Networks and ISDN Systems, 28(7–11):1385–1395, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1607.

Jarman:1994:USI

[JL94] D. Jarman and R. Y. Lee.

A user solution to Internetworking. In IEEE Computer Society. Technical Committee on Computer Communications [IEE94], pages 186–

192. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Junion-Metz:1995:IKR

[JM95] Gail Junion-Metz. Internet 101: K–12 resources. Technical report, ????, Chicago, IL, USA, 1995. various pp.

Junion-Metz:1996:IKR

[JM96a] Gail Junion-Metz. Internet 101: K–12 resources. Technical report, ????, ????, 1996. various pp.

Junion-Metz:1996:KRI

[JM96b] Gail Junion-Metz. K–12 resources on the Internet: an instructional guide. Internet workshop series; Number 5. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 1996. ISBN 1-882208-14-5 (paper). xii + 224 pp. LCCN ????

Jandt:1995:UIY

[JN95] Fred Edmund Jandt and Mary B. Nemnich. Using the Internet in your job search. JIST Works, Indianapolis, IN, USA, 1995. ISBN 1-56370-173-1. 225 pp. LCCN HF5382.7.J36 1995.

Jul:1995:IPE

[JO95] Erik Jul and OCLC Internet Cataloging Project. Intercat a public email list dedicated to facilitating communication concerning the OCLC Internet cataloging project, 1995.

Johnson:1994:IIT

[Joh94]  Diana L. Johnson.  An

introduction to the Internet: a training program. Thesis (M.A.), College of St. Scholastica, Duluth, MN, USA, 1994. 138 pp.

Johnson:1995:NIP

[Joh95] Ned Johnson. Navigating the Internet with Prodigy. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 0-672-30740-5. xvii + 330 pp. LCCN TK5105.875.I57J69 1995.

Johnson:1996:CIS

[Joh96a] Michael J. Johnson. Choosing an Internet service provider. Linux Journal, 24:??, April 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue24/choose.
html.

Johnson:1996:EPC

[Joh96b] Scott (J. Scott) Johnson.

Electronic publishing construction kit: creating multimedia for disk, CD-ROM, and the Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-12854-6 (paperback/CD-ROM). various pp. LCCN QA76.575.J65 1996.

Jones:1990:EI

[Jon90] Paul Jones. Exploring the Internet. ????, ????, 1990. ISBN ???? various pp. LCCN ????

Jones:1994:SCM

[Jon94a] Brant W. (Brant Wood) Jones. Surveying for creative market needs to predict product commoditization using the Internet as a survey mechanism. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1994. 71 pp.

Jones:1994:LRS

[Jon94b] Yolanda P. Jones. Legal reference services and the Internet: the promise and the peril. Technical report, The Author, Miami, FL, USA, 1994. 29 pp.

Jordan:1994:KSA

[Jor94] Sara Jordan. Kids’ songs about the Internet, 1994. ISBN 1-895523-48-6 (book). 1 sound cassette.

Joseph:1995:WLIa

[Jos95a] Linda C. Joseph. World link: an Internet guide for educators, parents, and students. Original Works/Greyden Press, Columbus, OH, USA, 1995. ISBN 1-57074-244-8. x + 260 pp. LCCN ????

Joseph:1995:WLIb

[Jos95b] Linda C. Joseph. World link: an Internet guide for educators, parents, and students. Original Works/Greyden Press, Columbus, OH, USA, revised edition, 1995. ISBN ???? xi + 274 pp. LCCN ????

Anonymous:1997:JIC

[Jou97] Journal of Internet cataloging, page various, 1997. Haworth Press, New York, NY, USA.

Juhnke:1994:ESP

[JP94] Joanne Juhnke and Chris Powell. Emotional support: physical loss, chronic illness, and bereavement: a guide to Internet resources. Technical report, University of Michigan, Ann Arbor, MI, USA, 1994. various pp.

Jaffe:1994:IITa

[JT94a] Lee David Jaffe and Roy Tennant. Introducing the Internet: a Trainer’s Workshop. Internet workshop series; no. 1. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 1994. ISBN 1-882208-03-X (book), 1-882208-05-6 (book and diskettes). viii + 92 pp. LCCN TK5105.875.I57 J23 1994 Ref 1 (text), pt. 2–3 (computer disks). US$30.00 (book), US$45.00 (book and diskettes). “Supplement to Crossing the Internet threshold.” Title on computer disk label: Introducing the Internet plus. On computer disk label: Windows. On computer disk label: Macintosh.

Jaffe:1994:IITb

[JT94b] Lee David Jaffe and Roy Tennant. Introducing the Internet: a trainer’s workshop. Internet workshop series; no.

1. Library Solutions Press, Berkeley, CA, USA, plus edition, 1994. ISBN 1-882208-05-6. viii + 92 pp. LCCN TK5105.875.I57J33 1994.

Jaffe:1994:IITc

[JT94c] Lee David Jaffe and Roy Tennant. Introducing the Internet: a trainer’s workshop. Internet workshop series; no. 1. Library Solutions Press, Berkeley, CA, USA, 2nd print., minor rev. edition, 1994. ISBN 1-882208-03-X. viii + 92 pp. LCCN ????

Junion:1996:KRI

[JT96] Gail Junion and Roy Tennant. K–12 resources on the Internet plus: an instructional guide. Internet workshop series; no. 5. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, trainer’s edition, 1996. ISBN 1-882208-15-3. xviii + 266 pp. LCCN ????

Jungling:1995:IBE

[Jun95]  Helmut Jungling.  Inter-

net und Bibliotheken: Entwicklung, Praxis, Herausforderungen. Kolner Arbeiten zum Bibliotheks- und Dokumentationswesen; Heft 21. Greven, Köln, Germany, 1995. ISBN 3-7743-0575-7.

185 pp. LCCN TK 5105.875 I57 I585 1995.

John:1994:ITH

[JV94] Nancy R. John and Edward J. Valauskas. The Internet troubleshooter: help for the logged-on and lost. American Library Association, Chicago, IL, USA, 1994. ISBN 0-8389-0633-8. xi + 145 pp. LCCN TK5105.875.I57 J65 1994.

James:1995:IGW

[JW95] Phil James and Jan Weingarten. Internet guide for Windows 95. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-260-

7. xxii + 522 pp. LCCN TK5105.875.I57J35 1995.

Kac:1996:IHN

[Kac96] Eduardo Kac. Internet hybrids and the new aesthetic of worldwide interactive events. In Blau et al. [B^{+}96b], pages 29–31. ISBN 0-89791-784-7. ISSN 1069-5419. LCCN T385 .S54 1996b.

URL http://www.acm.org:
80/pubs/citations/proceedings/
graph/253607/p29-kac/.

Kahin:1990:TPI

[Kah90] Brian Kahin. Toward a public information infrastructure: information policy and the Internet. Technical Report Contract no. L3-5445.0, Science, Technology and Public Policy Program, Kennedy School of Government, Harvard University, Cambridge, MA, USA, May 15, 1990. 42 pp.

Kahin:1992:BII

[Kah92] Brian Kahin, ed. Building Information Infrastructure: Issues in the Development of the National Research and Education Network. Mc-Graw-Hill, New York, NY, USA, 1992. ISBN 0-390-03083-X. iii + 443 pp. LCCN T58.5 .B85 1992. US$34.95. Includes the entire text of the High Performance Computing Act (HPCA) of 1991 that authorized the forthcoming National Research and Education Network (NREN), as well as numerous examinations of what it does and should mean. Papers from a workshop at Harvard’s Kennedy School of Government. Tough sledding in spots, but very rewarding in others.

Kahin:1994:IIS

[Kah94a] Brian Kahin. Information Infrastructure Sourcebook version 4. JFK School of Government/Harvard University, Cambridge, MA, USA, April 1994. ISBN none listed. 1700 pp. US$85. The fourth update of this compendium of official policy documents on information infrastructure, including the Internet. Includes historical policy documents, vision statements and position papers, program and project descriptions, reports and pending legislation.

Kahn:1994:VRG

[Kah94b] Robert E. Kahn. Viewpoint:

The role of government in the evolution of the Internet. Communications of the ACM, 37(8):15–19, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
179729.html.

Kallen:1994:PSR

[Kal94a] Ian Kallen. Post-soviet study resources on the Internet. Technical report, ????, ????, 1994. ii + 113 pp.

Kalnin:1994:IWI

[Kal94b] Mary Kalnin. Internet? what Internet?, 1994. 1 sound cassette.

Kaman:1993:BPC

[Kam93] Geradine Kaman. Broadband-ISDN: Personal connections to global resources. Internet research, 3(2):8–??, Summer 1993. CODEN IRESEF. ISSN 1066-2243.

Kane:1994:HGE

[Kan94a] Pamela Kane. The Hitchhiker’s Guide to the Electronic Highway. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel:

(503) 282-5215, 1994. ISBN 1-55828-352-8. xxv + 366 pp. LCCN TK5105.5 .K353 1994.

Kane:1994:ASS

[Kan94b] Pat Kane. Arts in society: Soft machines, hard battles. New statesman society, 7(318):31–??, September 2, 1994. CODEN NESSEF. ISSN 0954-2361.

Kanter:1995:DIW

[Kan95a] Beth Kanter. Dance on the Internet: Web resources. Afterimages, 2(2–3):10–17, Spring–Summer 1995. ISSN 1073-2101.

Kantor:1995:MOIa

[Kan95b] Andrew Kantor. Mecklermedia’s official Internet World: 60 minute guide to the Internet, including the World-Wide Web. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-342-9. xvii + 185 pp. LCCN TK5105.875.I57K36 1995.

Kantor:1995:MOIb

[Kan95c] Andrew Kantor. Mecklermedia’s official Internet World: 60 minute guide to the Internet, including the World-Wide Web. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-709252-X. xvii + 171 pp. LCCN ????

Kaplan:1996:IEP

[Kap96] Daniel Kaplan. Internet: Les enjeux pour la France: Livre blanc. Technical report, AF-TEL, Puteaux, France, 1996. 255 pp.

Karkare:1996:SEA

[Kar96] Sonia Karkare. Secure email architecture using PGP at the Timken Company. Thesis (M.S.), Kent State University, Kent, OH, USA, 1996. xv + 169 pp.

Kashipathi:1993:GUI

[Kas93] Arathi Kashipathi. A graphical user interface for IN-TERNET file transfer protocol. Thesis (M.S.), University of South Carolina, Columbia, SC, USA, 1993. vi + 66 pp.

Kochmer:1992:NUSa

[KB92] Jonathan Kochmer and Alison Brown. NorthWest-Net user services Internet resource guide (NUSIRG), adapted for use by OARnet / Jonanthan [sic] Kochmer; adaptations by Alison Brown. Zip Printing and Publishing, Columbus, OH, USA, OARnet, ver. 1.0., [adapted from] NUSIRG 3rd ed., ver. 3.1 edition, 1992. ix + 264 pp.

Krol:1995:MIG

[KCG95] Ed Krol, Pierre Cubaud, and Jacques Guidon. Le Monde Internet: guide and ressources. Ed. O’Reilly International Thomson, Paris, France, 2e edition, 1995. ISBN 2-84177-000-1. xxix + 543 pp. LCCN ????

Kretschmer:1996:SI

[KEdT96] Bernd Kretschmer, Jean-Philippe Elie, and Serge de Tavernier. Surfez sur Internet. Communication; 1541. Micro Application, Paris, France, 1996. ISBN 2-7429-0541-3. 412 pp. LCCN ????

Kehoe:1992:ZAI

[Keh92] Brendan P. Kehoe. Zen and the art of the Internet: a beginner’s guide to the Internet. Widener University, Chester, PA, USA, January 1992. iv + 96 pp.

Kehoe:1993:ZAI

[Keh93] Brendan P. Kehoe. Zen and the Art of the Internet: a Beginner’s Guide. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1993. ISBN 0-13-010778-6. xv + 112 pp. LCCN TK5105.875.I57 K44 1993. The first edition of this book was and is available only online, from various anonymous FTP servers. This second edition is somewhat enlarged and updated, yet is still the shortest introductory Internet book.

Kehoe:1994:ZAI

[Keh94] Brendan P. Kehoe. Zen and the Art of the Internet: a Beginner’s Guide. Prentice Hall series in innovative technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1994. ISBN 0-13-121492-6. xx + 193 pp. LCCN TK5105.875.I57 K44 1994. US$23.95. This guide should give you a reference to consult if you’re curious about what can be done with the Internet. It also presents the fundamental topics that are all too often assumed and considered trivial by many network users. It covers the basic utilities and information reaching other networks. An earlier, much less comprehensive version is available via FTP. Matisse Enzer says: Very friendly general overview book. Not really a how-to book, rather a Cultural Companion that explain what each type of Internet resource is, as well as the cultural and traditional uses and issues. Includes a chapter on “Things you will hear about… ” this chapter explains many famous (and infamous) Internet events and entities.

Kehoe:1996:ZAI

[Keh96] Brendan P. Kehoe. Zen and the art of the Internet: a beginner’s guide. Prentice Hall series in innovative technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 1996. ISBN 0-13-452914-6. xxxi + 255 pp. LCCN Tk5105.875.I51 K44 1996.

Kelleher:1992:NIM

[Kel92a] L. A. Kelleher. Navigating the Internet — a merit network cruise. Computer Networks and ISDN Systems, 25(4–5):460–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kelleher:1992:NI

[Kel92b] Laura A. Kelleher. Navigating the Internet. A merit information services cruise, Merit Network, Inc., Ann Arbor, MI, USA, 1992. 31 pp.

Kellner:1996:GI

[Kel96] Mark A. Kellner. God on the Internet. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-843-9 (paperback). xix + 308 pp. LCCN ????

Kent:1990:EPI

[Ken90] Stephen T. Kent. E-mail privacy for the Internet: New software and strict registration procedures will be implemented this year. Business communications review, 20(1):55–??, January 1, 1990. CODEN BCORBD. ISSN 0162-3885.

Kent:1993:IPE

[Ken93] Stephen T. Kent. Internet privacy enhanced mail. Communications of the ACM, 36(8):48–60, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
163390.html.

Kent:1994:MGI

[Ken94a]  Peter Kent.  10 Minute

Guide to the Internet. Alpha Books, Indianapolis, IN, USA, 1994. ISBN 1-56761-428-0 (paperback). xiii + 162 pp. LCCN TK5105.875.I57 K46. US$12.99, CDN$16.99.

Kent:1994:CIGa

[Ken94b] Peter Kent. The Complete Idiot’s Guide to the Internet. Alpha Books, Indianapolis, IN, USA, 1994. ISBN 1-56761-414-0. xv + 386 pp. LCCN TK5105.875 .I57 K46 1994. US$19.95, CDN$26.95. System requirements for computer disk: IBM-compatible PC; MS-DOS.

Kent:1994:CIGb

[Ken94c] Peter Kent. The complete idiot’s guide to the Internet. Alpha Books, Indianapolis, IN, USA, second edition, 1994. ISBN 1-56761-535-X. xx + 367 pp. LCCN TK5105.875.I57K47 1994b.

Kent:1994:CIGc

[Ken94d] Peter Kent. The complete idiot’s guide to the Internet, 1994. Includes CD-ROM.

Kent:1994:CIN

[Ken94e] Peter Kent. The complete idiot’s next step with the Internet. Alpha Books, Indianapolis, IN, USA, 1994. ISBN 1-56761-524-4 (paperback). xx + 373 pp. LCCN TK5105.875.I57K472 1994. US$19.95.

Kennedy:1995:IWW

[Ken95a] Angus J. Kennedy. The Internet and world wide web: the rough guide. Rough Guides, Ltd.: dist. by the Penguin Group, New York, NY, USA, 1995. ISBN 1-85828-198-9. 315 pp. LCCN TK5105.875.I57K45 1995.

Kennedy:1995:HGH

[Ken95b] Joyce Lain Kennedy. Hook up, get hired!: the Internet job search revolution. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11629-7, 0-471-11630-0 (paperback). xxii + 250 pp. LCCN HF5382.7.K458 1995.

Kent:1995:BS

[Ken95c]  Peter Kent.  Browser

shootout. Internet World, 6 (4):46–??, April 1, 1995. CO-DEN IERNE8. ISSN 1064-3923.

Kent:1995:CIGa

[Ken95d] Peter Kent. The complete idiot’s guide to the Internet for Windows. Que Corporation, Indianapolis, IN, USA, 1995. ISBN ???? xviii + 281 pp. LCCN TK5105.875.I57 K4715 1995.

Kent:1995:CIGb

[Ken95e] Peter Kent. The complete idiot’s guide to the Internet with Windows 95. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0629-6. xix + 342 pp. LCCN TK5105.875.I57K4716 1996.

Kent:1995:IF

[Ken95f] Peter Kent. Internet facil!

Prentice-Hall Hispanoamericana, Mexico, DF, Mexico, 1995. ISBN 968-880-470-3 (paperback). xviii + 387 pp. LCCN ????

Kent:1995:PCW

[Ken95g] Peter Kent. PGP companion for Windows: easy point-&- click encryption for your electronic information. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-304-2. xx + 172 pp. LCCN QA76.9.A25 K46 1995.

Kent:1995:UMI

[Ken95h] Peter Kent. Using Microsoft Internet Explorer. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0551-6. xxi + 335 + 1 pp. LCCN TK5105.875.I57K473 1995.

Keogh:1999:CMN

[Keo99] James Edward Keogh. Core MCSE: Networking Essentials. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010733-6. xviii + 358 pp. LCCN QA76.3.K46

1998. URL http://www.
phptr.com/bookseri/mcse.

html. Includes CD-ROM.

Kerwin:1995:LII

[Ker95] Robert J. Kerwin. Lawyers and the Internet: an intensive course on marketing and legal research. Technical Report 96-18.02, Massachusetts Continuing Legal Education, Inc., Boston, MA, USA, 1995. xii + 250 pp.

Kercheval:1998:TIA

[Ker98] Berry Kercheval. TCP/IP over ATM: a no-nonsense internetworking guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-768599-8. xv + 202 pp. LCCN TK5105.585 .K47 1998. URL

http://www.prenhall.com/
allbooks/ptr_0137685998.
html.

Kessler:1990:ICF

[Kes90] Gary C. Kessler. ISDN: concepts, facilities, and services. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-034242-3. xvi + 302 pp. LCCN TK5103.7 .K47 1990.

Kessler:1993:ICF

[Kes93] Gary C. Kessler. ISDN: concepts, facilities, and services. McGraw-Hill, New York, NY, USA, second edition, 1993. ISBN 0-07-034247-4. xviii + 454 pp. LCCN TK5103.7 .K47 1993. US$49.95.

Keshav:1997:EAC

[Kes97] Srinivsan Keshav. An engineering approach to computer networking: ATM networks, the Internet, and the telephone network. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63442-2. xviii + 660 pp. LCCN TK5105.5.K48 1997. US$47.95.

Katz:1993:TRI

[KF93] Dave Katz and Peter S. Ford.

TUBA: Replacing IP with CLNP. IEEE network, 7(3): 38–??, May 1, 1993. ISSN 0890-8044.

Klinger:1994:I

[KF94] Carol Klinger and Chris Fenwick. Internet, 1994. 1 videocassette (28 min.).

Kapor:1994:BDG

[KG94] Mitchell Kapor and Adam Gaffin. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. 195 pp.

Kong:1997:PSU

[KG97] Keith Kong and Dipak Ghosal. Pseudo-serving: a user-responsible paradigm for Internet access. Computer Networks and ISDN Systems, 29(8–13):1053–1064, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=08-13&aid=1710.

Kambil:1998:PAW

[KG98] Ajit Kambil and Mark Ginsberg. Public access Web information systems: Lessons from the Internet EDGAR project. Communications of the ACM, 41 (7):91–97, July 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-7/
p91-kambil/.

Kong:1999:MSS

[KG99a] Keith Kong and Dipak Ghosal. Mitigating serverside congestion in the Internet through pseudoserving. IEEE/ACM Transactions on Networking, 7 (4):530–544, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-4/p530-kong/.

Kotz:1999:MAF

[KG99b] David Kotz and Robert S. Gray. Mobile agents and the future of the Internet. Operating Systems Review, 33(3):7–13, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kidder:1995:HPI

[KH95] Gayle Kidder and Stuart (Stuart H.) Harris. HTML publishing with Internet Assistant: your guide to using Microsoft’s HTML addon. Ventana Press, Inc., Research Triangle, NC, USA, 1995. ISBN 1-56604-273-9. xx + 250 pp. LCCN QA76.76.H94H373 1995.

Kochmer:1991:NUSb

[KHJ91] Jonathan Kochmer, Eric S.

Hood, and Dan Jordt. NorthWestNet user services Internet resource guide (NUSIRG). Technical report, Northwest Academic Computing Consortium, Bellevue, WA, USA, December 1991. xi + 274 pp.

Kochmer:1992:NUSb

[KHJ92] Jonathan Kochmer, Eric S.

Hood, and Dan Jordt. NorthWestNet user services Internet resource guide (NUSIRG). Northwest Academic Computing Consortium, Bellevue, WA, USA, third edition, March 1992. xi + 279 pp.

Kochmer:1991:NUSa

[KHJG91] Jonathan Kochmer, Eric S.

Hood, Dan Jordt, and Larry Gales. NorthWestNet user services Internet resource guide (NUSIRG). Technical report, Northwest Academic Computing Consortium, Bellevue, WA, USA, October 1991. xi + 272 pp.

Kichen:1994:IIG

[Kic94] Steve Kichen. Investor’s Internet guide. Forbes, 154(7): 158–??, September 26, 1994. CODEN FORBA5. ISSN 0015-6914.

Kientzle:1995:IFF

[Kie95] Tim Kientzle. Internet file formats. Coriolis Group Books, Scottsdale, AZ, USA, 1995. ISBN 1-883577-56-X. xviii + 397 pp. LCCN TK5105.875.I57 K53 1995.

Killelea:1998:WPT

[Kil98] Patrick Killelea. Web Performance Tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-379-0. xix + 351 pp. LCCN QA76.9.C63 K55

1998.  US$32.95.  URL

                                        http://www.oreilly.com/
                                        catalog/webpt.

Kinder:1994:LII

[Kin94] Robin Kinder. Librarians on the Internet: impact on reference services. Haworth Press, New York, NY, USA, 1994. ISBN 1-56024-672-3. 410 pp. LCCN Z711.L487 1994.

King:1995:UDC

[Kin95a] Graham King. Understanding and Designing Computer Networks. Edward Arnold, London, UK and New York, NY, USA, 1995. ISBN 0-340-61419-6. vii + 208 pp. LCCN TK5105.5 .K56 1995. US$39.99.

Kinkoph:1995:CIG

[Kin95b] Sherry Kinkoph. The Complete Idiot’s Guide to Online Chat. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-595-3. xxv + 273 pp. LCCN TK5105.875.I57K56 1995. US$16.99.

Kalyanakrishnan:1999:RIH

[KIP99] M. Kalyanakrishnan, R. K.

Iyer, and J. U. Patel. Reliability of Internet hosts: a case study from the end user’s perspective. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):47–57, January 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/1-2/2035.pdf.

Kern:1998:BNE

[KJGH98] Harris Kern, Randy Johnson, Stuart D. Galup, and Dennis Horgan. Building the New Enterprise: People, Processes and Technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-079671-9. xvi + 382 pp. LCCN HD30.37.B843

1998. URL http://www.

                                        phptr.com/ptrbooks/ptr_
                                        0130796719.html.

Knierriem-Jasnoch:1996:RWF

[KJTS96] A. Knierriem-Jasnoch, B. Tritsch, and U. Schroeder. Reflection on WWW functionalities for educational purposes. Computers and Graphics, 20 (3):435–443, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.cgi?
                                        year=1996&volume=20&issue=
                                        3&aid=9600013.

Kahin:1995:PAI

[KK95a] Brian Kahin and James Keller. Public access to the Internet. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-11207-8, 0-262-61118-X (paperback). viii + 390 pp. LCCN TK5105.875.I57 P83 1995. “A publication of the Harvard Information Infrastructure Project.”.

Klau:1995:IA

[KK95b] Peter Klau and Michele Klau. Das Internet Adressbuch. IWT, Vaterstetten, Germany, 1995. ISBN 3-8266-2653-2. 285 pp. LCCN ????

Knel:1996:BIC

[KK96a] Linda Knel, compiler and Jean M. Kaleda, editor. The book of Internet career addresses: a list of career, educational, and financial aid sites. Technical report, Printed in cooperation with the Suffolk Cooperative Library System and Suffolk County Library Association, Reference and Adult Services Division, Bellport, NY, USA, 1996. xvi + 270 pp.

Krol:1996:WIU

[KK96b] Ed Krol and Bruce Klopfenstein. The whole Internet user’s guide and catalog. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, academic edition, 1996. ISBN 0-534-50674-7. xxix + 609 pp. LCCN TK5105.875.I57K76 1996.

Klatt:1995:WWU

[KKB95] Edward C. Klatt, Carolyn A.

Klatt, and Jay L. Burton. Windows to the world: Utah Library Network Internet training manual. Technical report, Dept. of Community and Economic Development, Utah State Library Division, Salt Lake City, UT, USA, March 1995. various pp.

Kadambi:1998:GEM

[KKC98] Jayant Kadambi, Mohan Kalkunte, and Ian Crayford. Gigabit Ethernet: migrating to high-bandwidth LANs. Prentice Hall series in computer networking and distributed systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-913286-4. xi + 365 pp. LCCN TK5105.8.E83K33

                                        1998. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0139132864.html.

Klopfer:1995:ISF

[KKF95] Fred Klopfer, Susan Klopfer, and Michael Fraley. Internet success with Fred. Vanatech Press, Kendallville, IN, USA, 1995. ISBN 0-9638715-4-4. xi + 193 pp. LCCN ????

Kraut:1998:SSI

[KKM^{+}98] Robert Kraut, Sara Kiesler, Tridas Mukhopadhyay, William Scherlies, and Michael Patterson. On site: social impact of the Internet: what does it mean? Communications of the ACM, 41(12): 21–22, December 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-12/
                                        p21-kraut/.

King:1994:DEJ

[KKO94] Lisabeth King, Diane Kovacs, and Ann Okerson. Directory of Electronic Journals, Newsletters and Academic Discussion Lists. Association of Research Libraries, Washington, DC 20036, USA, fourth edition, May 1994. ISBN none listed. v + 575 pp. LCCN Z 6951 D625 1994. US$54 (US$36 to members of the Association of Research Libraries). This directory is broken into two main sections: a catalog of 1,800 academic discussion lists and interest groups, and listings of 440 electronic journals, newsletters and digests — along with notes on how to access each publication. Scholarly lists are grouped by broad subject areas and journals and newsletters are listed in alphabetical order. Also includes five articles and bibliographies on electronic publishing.

Kovacs:1994:ESI

[KL94] Diane Kovacs and Kent State University. Libraries. Earth sciences Internet guide. Technical report, Kent State University Libraries, Canton, OH, USA, June 15, 1994. 68 pp.

Kong:1995:GAI

[KL95] Ling Kong and Sheila Lacroix. A guide to addictions information on the Internet and other public information sites. Technical report, Addiction Research Foundation Library, Toronto, ON, Canada, 1995. 85 pp.

Kuo:1998:NDC

[KL98] Geng-Sheng Kuo and Jing-Pei Lin. New design concepts for an intelligent Internet. Communications of the ACM, 41(11):93–98, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-11/
                                        p93-kuo/.

Klau:1994:IWV

[Kla94]  Peter Klau.  Das Inter-

net — Weltweit vernetzt. IWT-Verlag GmbH, Vaterstetten bei Muenchen, Germany, 1994. ISBN 3-88322-467-7. 455 pp. DM 59. This introductory text discusses how to connect to the Internet and covers the basic features of e-mail, data transfer, searching, discussion groups, network tools and more. The book is in German, but the examples, illustrations and all the standard terminology are in English.

Kleinrock:1994:RIF

[Kle94] Leonard Kleinrock. Realizing the information future: the Internet and beyond. National Academy Press, Washington, DC, USA, 1994. ISBN 0-309-05044-8. xv + 301 pp. LCCN TK5105.875.N37 N37 1994b. US$24.95. “Support for this project was provided by the National Science Foundation (under grant no. NCR-9223810)” — p. ii. “Leonard Kleinrock, chair”.

Klee:1995:CIR

[Kle95] Peter Klee. Cuba Internet resources, 1995.

Kluis:1995:STI

[Klu95] Kevin C. Kluis. Students, teachers, and the Internet. Technical report, ????, ????, 1995. iii + 38 pp.

Karlovsky:1995:MNI

[KLW95] Steve Karlovsky, Kevin E.

Leininger, and Linda Winkler. Mosaic and the new Internet: serving business and personal needs. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-912097-0 (paperback). xiv + 354 pp. LCCN TK5105.882.K37 1995.

Krol:1995:IYG

[KM95] Ed Krol and Jun Murai. Intanetto yuzazu gaido. Intanashonaru Tomuson Paburisshingu Japan, Tokyo, Japan, shohan. edition, 1995. ISBN 4-900718-12-2 (paperback). xxxvi + 783 pp. LCCN ????

Kardas:1996:UIS

[KM96a] Edward P. Kardas and Tommy M. Milford. Using the Internet for social science research and practice. Wadsworth Publishing Company, Pacific Grove, CA, USA, 1996. ISBN 0-534-51388-3. xv + 221 + 2 pp. LCCN H 61.95 K37 1996.

Kievit:1996:SWG

[KM96b] James Kievit and Steven Metz. The strategist and the web: a guide to Internet resources. Technical report, Strategic Studies Institute, U.S. Army War College, Carlisle, PA, USA, February 20, 1996. vi + 31 pp.

Kehoe:1995:IDA

[KMM95] Brendan P. Kehoe, Miguel A. Martinez Sarmiento, tr, and Javier Matuk Kanan, rev. Internet: del Arte al Zen. Prentice-Hall Hispanoamericana, Mexico, DF, Mexico, 1995. ISBN 968-880-489-4. XX + 193 pp. LCCN ????

Knoble:1996:AIS

[Kno96] Jim Knoble. Almost Internet with SLiRP and PPP. Linux Journal, 24:??, April 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue24/1174.
html.

Kochmer:1993:IPN

[Koc93] Jonathan Kochmer. Internet passport: NorthWestNet’s guide to our world online. NorthWestNet: Northwest Academic Computing Consortium, Bellevue, WA, USA, fourth edition, 1993. ISBN 0-9635281-0-6. xxxii + 515 pp. LCCN TK5105.875.I57K63 1993.

Koehler:1995:I

[Koe95]  Lora Koehler.  Internet.

A new true book. Children’s Press, Chicago, IL, USA, 1995. ISBN 0-516-01079-4. 45 pp. LCCN TK5105.875.I57K64 1995.

Kohn:1997:PGB

[Koh97] Daniel M. Kohn. Providing global broadband Internet access using low-earth-orbit satellites. Computer Networks and ISDN Systems, 29(15):1763–1768, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=15&aid=1783.

Kolletzki:1996:SIB

[Kol96] Stephan Kolletzki. Secure Internet banking with Privacy Enhanced Mail — A protocol for reliable exchange of secured order forms. Computer Networks and ISDN Systems, 28(14):1891–1899 (or 1891–1900??), November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1996/
28/14/1642.pdf.

Konopelski:1982:IIR

[Kon82] Louis J. Konopelski. Implemented Internet remote login on a personal computer. Technical report MIT/LCS/TM-233, Massachusetts Institute of Technology, Laboratory for Computer Science, Cambridge, MA, USA, December 1982. 80 pp.

Kongsli:1991:TII

[Kon91] T. A. Kongsli. TCP/IP internetwork communication through LANs interconnected by Dikon Meganet. Computer Networks and ISDN Systems, 23(1–3):57–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kotlas:1993:IB

[Kot93] Carolyn Kotlas. Internet bibliography. Technical report, Institute for Academic Technology, Research Triangle Park, NC, USA, 1993. 5 pp.

Kovacs:1995:ITG

[Kov95] Diane Kovacs. The Internet trainer’s guide. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01978-5. xiv + 240 pp. LCCN TK5105.875.I57 K68 1995.

Kappe:1993:AMD

[KP93] Frank Kappe and Gerald Pani. The architecture of a massively distributed hypermedia system. Internet research, 3(1):10–??, Spring 1993. CODEN IRESEF. ISSN 1066-2243.

Krol:1994:IBC

[KP94] Ed Krol and Ronald Petrusha. Internet in a box: the complete Internet solution with instant point-and-click access, 1994. 3 computer disks.

Kessler:1995:DLP

[KP95] Jack Kessler and Kermit M.

Patton. Digital libraries: putting the Internet to work. Technical Report D95-1947, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, August 1995. 23 pp.

Kaufman:1995:NSP

[KPS95] Charles Kaufman, Radia Perlman, and Michael Speciner. Network Security: Private Communication in a Public World. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-061466-1. xx + 504 pp. LCCN QA76.9.A25K345 1995. US$46.00.

Kragen:1994:CGC

[Kra94] Marshall Kragen. Cansearch, a guide to cancer resources a guide to cancer resources on the Internet, 1994.

Kraut:1996:I

[Kra96]  Robert Kraut.  The In-

ternet@Home. Communications of the ACM, 39(12):32– 35, December 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kirsh:1995:BLA

[KRK95] Ellen M. Kirsh, Lance Rose, and Bruce P. Keller. Business and legal aspects of the Internet and online services. Technical report, Law Journal Seminars-Press, New York, NY, USA, 1995. 505 pp.

Krol:1987:HGI

[Kro87] Ed Krol. The hitchhikers guide to the Internet. Technical report, Board of Trustees of The University of Illinois, Urbana, IL, USA, August 25, 1987. 23 pp.

Krol:1989:HGI

[Kro89] Ed Krol. The hitchhikers guide to the Internet. Technical report, University of Illinois Urbana, Urbana, IL, USA, 1989. 24 pp.

Krol:1992:WIU

[Kro92] Ed Krol. The Whole Internet User’s Guide & Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K86 1992. US$24.95. Perhaps the most ambitious of the new crop of books, this one is both a user’s guide and a catalog of resources in one. It is aimed at graduate students who want to use the Internet for research.

Krol:1993:WIU

[Kro93] Ed Krol. The Whole Internet: User’s Guide and Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K86 1992. US$24.95.

Krol:1994:WIU

[Kro94] Ed Krol. The Whole Internet User’s Guide and Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, April 1994. ISBN 1-56592-063-5. xxiv + 543 pp. LCCN TK5105.875.I57 K86 1994. US$24.95. Highly recommended. Definitely a MUST for any new and not-so-new Internet users …. It is a complete user’s guide to the Internet, covering everything from the basics, like electronic mail and newsgroups, to the newest developments. A large part of the book tells you how to find the resources you want. The second edition covers the latest developments on this rapidly-changing Internet and many new Internet tools like Mosaic, pine, xarchie, etc., ….

Krol:1995:CAM

[Kro95a] Ed Krol. Conectate al mundo de Internet. A Nutshell handbook. McGraw-Hill, New York, NY, USA, 2a. edition, 1995. ISBN 970-10-0637-2 (paperback). xxvi + 597 pp. LCCN ????

Krol:1995:WIW

[Kro95b] Ed Krol. The whole Internet for Windows 95: user’s guide and catalog. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-155-0. xix + 625 pp. LCCN TK5105.875.I57 K862 1995.

Kronick:1995:NNF

[Kro95c] Scott Kronick. Netscape Navigator: a friendly handbook on the coolest way to explore the Internet. Netscape Communications Corp., Mountain View, CA, USA, April 2, 1995. 218 pp.

Kirsh:1994:BLA

[KRS94] Ellen M. Kirsh, Lance Rose, and Shari Steele. Business and legal aspects of the Internet and online services. Law Journal Seminars-Press, New York, NY, USA, 1994. ISBN ???? 881 pp. LCCN ????

Kretschmer:1995:SI

[KS95] Brend Kretschmer and Gotz Schuchart. Surfez sur Internet. Collection Grand livre, communication, 1439. Micro Application, Paris, France, 1995. ISBN 2-7429-0439-5. 328 pp. LCCN ????

Korth:1997:SFT

[KS97] Henry F. Korth and Abraham Silberschatz. The science of future technology: Database research faces the information explosion. Communications of the ACM, 40 (2):139–142, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
1997-40-2/p139-korth/.

Kraut:1996:HFT

[KSM^{+}96] Robert Kraut, William Scherlis, Tridas Mukhopadhyay, Jane Manning, and Sara Kiesler. The HomeNet field trial of residential Internet services. Communications of the ACM, 39(12):55–63, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koga:1998:SAU

[KSM98] Y. Koga, K. Shiomi, and Y. Matsushita. Service architecture for utilizing TINAbased public network services from the Internet environment. In IEEE Computer Society. Technical Committee on Computer Communications [IEE98], pages 162–171. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Kidambi:1995:IMA

[KSO95] K. R. Kidambi, R. Sankar, and J. Ottensmeyer. Internetworking MANs to ATM for broadband services support. In IEEE Computer Society. Technical Committee on Computer Communications [IEE95], pages 102– 113. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Klusch:1999:CIA

[KSW99] Matthias Klusch, Onn M. Shehory, and Gerhard Weiss, editors. Cooperative information agents III: Third International Workshop, CIA’99, Uppsala, Sweden, July 31-August 2, 1999: proceedings, volume 1652 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66325-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 C53 1999.

Kessler:1992:MAN

[KT92] Gary C. Kessler and David A. Train. Metropolitan area networks. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-034243-1. xii + 436 pp. LCCN TK5105.85 .K47 1992. US$39.95.

Koop:1995:TIC

[KT95] C. Everett (Charles Everett)

Koop and Reed Tuckson. Teleconference and Internet conference partnerships for networked health information for the public, May 15, 1995. 1 videocassette (180 min.).

Tu:1994:TWK

[kTlC94] Li kang Tu and Sung lin Chen. Tsung Windows kan InterNet ti shih chieh. Sung kang tien nao tu shu tzu liao ku fen yu hsien kung ssu, Taipei shih, Taiwan, chu pan. edition, 1994. ISBN 957-22-1725-9 (paperback). various pp. LCCN ????

Tu:1995:ISY

[kTlC95] Li kang Tu and Sung lin Chen. Internet shih yung chih lu. Ju lin tu shu yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1995. ISBN 957-652-700-7 (paperback). various pp. LCCN ????

Kuhn:1994:LBM

[KTS94] Eva Kühn, Thomas Tschernko, and Konrad Schwarz. A language based multidatabase system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):509, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kugel:1995:FNI

[Kug95] Karin Beth Kugel. Feminism newsgroups on the Internet: have women found a safe space in cyberspace? Senior honors thesis, Brandeis University, Waltham, MA, USA, 1995. 53 pp.

Kumar:1996:MMI

[Kum96] Vinay Kumar. MBone — interactive multimedia on the Internet. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-397-3. xv + 252 pp. LCCN QA76.76.I59K85 1996.

Kuo:1997:CIA

[Kuo97]  Benjamin Kuo.  Create

Internet-aware apps with the Internet SDK. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Kurland:1996:NWY

[Kur96a] Daniel J. Kurland. The Net, the Web, and you: all you really need to know about the Internet– and a little bit more. Wadsworth Publishing Company, Pacific Grove, CA, USA, 1996. ISBN 0-534-51281-X (paperback). x + 213 pp. LCCN TK5105.875.I57K87 1996.

Kurz:1996:ILL

[Kur96b] Raymond A. Kurz. Internet and the law: legal fundamentals for the Internet user. Government Institutes, Rockville, MD, USA, 1996. ISBN 0-86587-506-5. xv + 248 pp. LCCN KF242.A1 K87 1996.

Kaashoek:1992:FIP

[KvRvST92] M. Frans Kaashoek, Robbert van Renesse, Hans van Staveren, and Andrew S. Tanenbaum. FLIP; an Internetwork Protocol for supporting distributed systems. Operating Systems Review, 26 (2):29, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaashoek:1993:FIP

[KvRvST93] M. Frans Kaashoek, Robbert van Renesse, Hans van Staveren, and Andrew S. Tanenbaum. FLIP: an internetwork protocol for supporting distributed systems. ACM Transactions on Computer Systems, 11(1):73– 106, February 1993. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1993-11-1/
p73-kaashoek/.

Kochan:1989:UN

[KW89] Stephen G. Kochan and Patrick H. Wood, editors. UNIX Networking. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48440-4. viii + 400 pp. LCCN QA76.76.O63 U546 1989.

Kapor:1993:DNC

[KW93] Mitchell Kapor and Daniel Weitzner. Developing the national communications and information infrastructure. Internet research, 3(2):21–??, Summer 1993. CODEN IRE-SEF. ISSN 1066-2243.

Kessler:1994:INI

[KW94] Jack Kessler and Dell Wilkinson. Information networks: Internet lessons from France’s Minitel? Technical Report D94-1883, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, November/December 1994. 12 pp.

Kantor:1995:III

[Kw95] Andrew Kantor and Internet world. An introduction to the Internet: including email, Usenet News, and the World-Wide Web. Mecklermedia, Westport, CT, USA, 1995. ISBN 1-57207-000-5. iii + 64 pp. LCCN ????

Klusch:1998:CIA

[KW98] Matthias Klusch and Gerhard Weiss, editors. Cooperative information agents II: learning, mobility, and electronic commerce for information discovery on the Internet: Second International Workshop, CIA ’98, Paris, France, July 1998: proceedings, volume 1435 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64676-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1435.

Lambert:1995:BM

[Lam95] Nancy Lambert. The better mousetrap. Searcher: the magazine for database professionals, 3(5):34–??, May 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Lam:1996:AHM

[Lam96] K. C. Lam. An application of hypertext markup language (HTML) for institute of textiles and clothing in the Internet. In Anonymous [Ano96x], pages 313–314. ISBN 951-722-545-8. LCCN ????

LaMacchia:1997:IFC

[LaM97] Brian A. LaMacchia. The Internet Fish Construction Kit. Computer Networks and ISDN Systems, 29(8–13): 1237–1248, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.

                                        cgi?year=1997&volume=29&
                                        issue=08-13&aid=1698.

LAN:1986:LLA

[LAN86] LAN—The Local Area Network Magazine, 1986. CO-DEN LANNER. ISSN 0898-0012. LAN, Miller Freeman Publications, P.O. Box 41904, Nashville, TN 37204, USA, Tel: (800) 933-3321 or (615) 377-3322. Monthly: US$20 per year.

Lang:1993:EIa

[Lan93] Mary Beth Lang. Exploring Internet, 1993. 1 videocassette (2 hr. + 4 min.).

Landwehr:1994:LCI

[Lan94] Brian R. Landwehr. Is linking a company to the Internet worth the costs? Technical report, ????, ????, 1994. ii + 22 pp.

Laporte:1996:I

[Lap96] Leo Laporte. Internet!, 1996. videocassettes.

LaQuey:1989:UDC

[LaQ89] Tracy Lynn LaQuey. Users’ Directory of Computer Networks. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Englewood Cliffs, NJ 07632, USA, 1989. ISBN 1-55558-047-5 (DP) and 0-13-950262-9 (PH). xv + 630 pp. LCCN TK5105.5 .U74 1990. US$34.95. Digital Part Number EY-C200E-DP.

A directory of users, domains, etc. of the Internet, BITNET, UUCP, etc.

LaQuey:1994:ICB

[LaQ94a] Tracy L. LaQuey. The Internet companion: a beginner’s guide to global networking. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-40766-3. viii + 262 pp. LCCN TK5105.875.I57 L37 1994. US$12.95, CDN$16.95.

LaQuey:1994:ICP

[LaQ94b] Tracy L. LaQuey. The Internet Companion Plus: a Beginner’s Start-up Kit for Global Networking: Includes Wordlink Software to Connect to the Internet. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-40837-6. viii + 262 pp. LCCN TK5105.875.I57 L39 1994. System requirements for computer disk: PC; 1 MB RAM; DOS; Windows 3.1; hard disk with 1.2 MB free space; modem. “An Editorial Inc. book.”.

Larkin:1994:AGP

[Lar94a] Nancy Larkin. The ACS guide: PC/DOS software, Macintosh software, CUN-YVM, electronic mail, the Internet. Technical report, Academic Computing Services, Hunter College, City University of New York, New York, NY, USA, 1994. 164 pp.

Larson:1994:CYA

[Lar94b] Dale L. Larson. Connect your Amiga!: a guide to the Internet, LANs, BBSs, and online services. Intangible Assets Manufacturing, Drexel Hill, PA, USA, 1994. ISBN 1-885876-02-5. 256 pp. LCCN QA76.8.A177L353 1994.

LaSalle:1995:LCI

[LaS95a] Kim LaSalle. Life in cyberspace: Internet, public relations and marketing. Technical report, LaSalle Communications, Indianapolis, IN, USA, 1995. 93 pp.

LaSalle:1995:LCP

[LaS95b] Kim LaSalle. Life in cyberspace: the professional’s guide to Internet public relations and marketing. Technical report, LaSalle Communications, Indianapolis, IN, USA, 1995. 103 pp.

Lawrence:1995:PBR

[Law95] Andy Lawrence. Publish and be robbed? New Scientist, 145(1965):32–??, February 18, 1995. CODEN NWS-CAL. ISSN 0262-4079, 0028-6664.

Lazenby:1997:BRN

[Laz97] Daniel Lazenby. Book review: Netizens: On the History and Impact of Usenet and the Internet. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
1997-12/hauben.html.

Levine:1993:ID

[LB93] John R. Levine and Carol Baroudi. Internet for Dummies. — For dummies. IDG Books, San Mateo, CA, USA, 1993. ISBN 1-56884-024-1. xxiv + 355 pp. LCCN TK 5105.875 I57 L48 1993. US$19.95. “Covers DOS, Windows Mac, and UNIX platforms!” — Contents: Pt. I. Getting onto the Internet.

— Pt.II. Mail and gossip. — Pt. III. Instant gratification.

— Pt. IV. Finding stuff on the net. — Pt. V. The part of tens. — Pt. VI. Resource reference.

Levine:1994:ID

[LB94a] John R. Levine and Carol Baroudi. The Internet for dummies. –For dummies. IDG Books, San Mateo, CA, USA, second edition, 1994. ISBN 1-56884-222-8 (paperback). xxviii + 427 pp. LCCN TK 5105.875 I57 L48 1994b. US$19.99, CDN$26.99. Contents: Pt. I. Getting onto the Internet.– Pt.II. Mail and gossip.– Pt. III. Instant gratification.– Pt. IV. Finding stuff on the net.– Pt. V. The part of tens.– Pt. VI. Resource reference. Intended as a beginners guide to the Internet includes useful sections on mail, gopher, news, ftp, etc. All in a fairly lighthearted style intended to set novices at ease. At over 350 pages, there is much information and the book is well indexed. Most of the material is oriented toward UNIX shell users, with some discussion of Winsock applications. Second edition notes: “We’ve updated and expanded the book, clarifying the first few chapters, expanding the coverage of WWW and other new services, and generally fixing stuff.”.

Levine:1994:IDA

[LB94b] John R. Levine and Carol Baroudi. Internet fur Dumme Anfanger. –For dummies. IWT Verlag, Bonn, Germany, aufl. edition, 1994. ISBN 3-88322-531-2. 460 pp. LCCN ????

Levine:1994:IPN

[LB94c] John R. Levine and Carol Baroudi. Internet pour les nuls. Pour les nuls. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 2-7361-1298-9. xix + 352 pp. LCCN ????

Levine:1994:UIS

[LB94d] John R. Levine and Carol Baroudi. Usare Internet senza fatica. McGraw-Hill Libri Italia, Milano, Italy, 1994. ISBN 88-386-0313-8. 315 pp. LCCN ????

Levine:1995:IPK

[LB95a] John R. Levine and Carol Baroudi. Internet i pyolgon kayo = The Internet for dummies, 2nd ed. –Pyolgon kayo;

8. Hongik Midio, Soul Tukpyolsi, Korea, chopan. edition, 1995. ISBN 89-8066-708-6. 455 pp. LCCN ????

Levine:1995:IPI

[LB95b] John R. Levine and Carol Baroudi. Internet para inexpertos. Limusa/Noriega, Mexico, DF, Mexico, 1995. ISBN 968-18-5204-4. 438 pp. LCCN ????

Levine:1995:IS

[LB95c] John R. Levine and Carol Baroudi. Internet secrets. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-452-2. xlii + 990 pp. LCCN TK5105.875.I57L482 1995.

Levine:1995:ISK

[LB95d] John R. Levine and Carol Baroudi. INTERNET sha kua shu. Tien nao shen tung sha kua tsung shu. Ching hua ta hsueh chu pan she, Pei-ching, China, ti 1 pan. edition, 1995. ISBN 7-302-01693-3. XXII + 306 pp. LCCN ????

Levine:1995:SI

[LBP95] John R. Levine, Carol Baroudi, and Beatriz Parra Perez, tr. Los secretos de Internet. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-790-2. 696 pp. LCCN ????

Levine:1995:ID

[LBY95] John R. Levine, Carol Baroudi, and Margy Levine Young. The Internet for dummies. –For dummies. IDG Books, San Mateo, CA, USA, third edition, 1995. ISBN 1-56884-620-7. xxxii + 400 pp. LCCN TK5105.875.I57L48 1995.

Levine:1996:ID

[LBY96] John R. Levine, Carol Baroudi, and Margy Levine Young. The Internet for dummies, 1996. ISBN 0-694-51667-8. 1 sound cassette (90 min.).

Li:1993:ESG

[LC93] Xia Li and Nancy Crane. Electronic Style: a Guide to Citing Electronic Information. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-909-X. xi + 65 pp. LCCN PN171.D37 L5 1993. US$15. Here is where to find out how to cite in your bibliographies references found on the Internet, on CD-ROMs, and online database searches.

Li:1996:ESG

[LC96] Xia Li and Nancy Crane. Electronic style: a guide to citing electronic information. Mecklermedia’s official Internet world. Mecklermedia, Westport, CT, USA, second edition, 1996. ISBN 0-88736-998-7. LCCN ????

Lysohir:1998:UWA

[LC98] Steve Lysohir and Computer Channel, Inc. Understanding Web-based applications HTML, dynamic HTML, SSL and more, 1998. 1 videocassette (47 min.).

Leiner:1997:SFT

[LCC^{+}97] Barry M. Leiner, Vinton G. Cerf, David D. Clark, Robert E. Kahn, Leonard Kleinrock, Daniel C. Lynch, Jon Postel, Lawrence G. Roberts, and Stephen S. Wolff. The science of future technology: The past and future history of the Internet. Communications of the ACM, 40(2):102–108, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
1997-40-2/p102-leiner/.

Long:1990:SRI

[LCP90] Darrell Long, J. L. Carroll, and C. J. Park. A study of the reliability of Internet sites. Technical report UCSC-CRL-90-46, University of California, Santa Cruz, Computer Research Laboratory, Santa Cruz, CA 95064, USA, September 1990. 17 pp.

Long:1991:ERH

[LCP91] Darrell Long, J. L. Carroll, and C. J. Park. Estimating the reliability of hosts using the Internet. Technical report UCSC-CRL-91-06, University of California, Santa Cruz, Computer Research Laboratory, Santa Cruz, CA, USA, 1991. 17 pp.

Lumpkin:1995:GSI

[LD95] Joan Brady Lumpkin and Susan B. Durnbaugh. Getting started with the Internet. Getting started with Windows series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-12418-4. x + 84 pp. LCCN TK5105.875.I57L86 1995.

Li:1999:DVC

[LD99] J. Li and E. De Doncker. Dynamic visualization of computations on the Internet. Lecture Notes in Computer Science, 1593:360–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lancki:1996:MIT

[LDG96] Ben Lancki, Abhijit Dixit, and Vipul Gupta. Mobile IP: Transparent host migration on the Internet. Linux Journal, 28:??, August 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue28/1271.
html.

LaQuey:1995:EIC

[LDH95] Tracy L. LaQuey, Stephan Deutsch, and Richard Harris. The European Internet companion: a beginner’s guide to global networking. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-42778-8. xiv + 272 pp. LCCN TK5105.875.I57L36 1995.

LeBlanc:1996:RPI

[Le 96] Dee-Ann Le Blanc. Running a perfect Internet site with Linux. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0514-1. xvii + 388 pp. LCCN IN PROCESS.

Levitus:1995:WMH

[LE95] Bob Levitus and Jeff Evans.

WebMaster Macintosh: How to Build Your Own Worldwide Server Without Really Trying. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-445574-3 (book), 0-12-445575-1 (book + CD ROM). ???? pp. LCCN TK5105.888 .L48 1995.

LeVitus:1996:CEI

[LE96] Bob LeVitus and Jeff Evans.

Cheap and easy Internet access for windows. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-445597-2. ???? pp. LCCN TK5105.875.I57L485 1996.

Leach:1995:RIA

[Lea95a] Martin David Leach. Registry of Internet addresses of biotech and biomedical related companies/organisations, 1995.

VPI:1995:LMI

[Lea95b] Learn to master the Internet, 1995. ISBN 1-57142-030-4. 1 videocassette (50 min.).

LeBlanc:1993:TFM

[LeB93]  James D. LeBlanc.  To-

wards finding more catalog copy: The possibility of using OCLC and the Internet to supplement RLIN searching. Cataloging & classification quarterly, 16(1):71–??, 1993. CODEN CCQUDB. ISSN 0163-9374.

Lee:1998:IBF

[Lee98] Matthew K. O. Lee. Internetbased financial EDI: towards a theory of its organizational adoption. Computer Networks and ISDN Systems, 30(16–18):1579– 1588, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/comnet/sub/1998/
                                        30/16-18/2016.pdf.

Legi-Slate:1993:LGS

[Leg93] Legi-Slate. LEGI-SLATE gopher service, 1993.

Anonymous:1994:LRI

[Leg94a] The Legal researcher’s Internet directory, page various, 1994. Legal Research of New York, Inc., Brooklyn, NY, USA.

LCBSC:1994:GAC

[Leg94b] Legislative Counsel Bureau, State of California. A Guide for Accessing California Legislative Information over Internet. Sacramento, CA, USA, 1994. 30 pp. This pamphlet tells how you can find California legislative information online. Explains what legislation info is available, what assistance is available and how the information is organized. The majority of the book is spent explaining the Internet, how to get access, how to use electronic mail and where to go for more detailed information. There’s also a simple glossary of legislative terms. This pamphlet is also available online, in PostScript format. Send e-

mail to: ftpmail@leginfo.
public.ca.gov with the

body

connect leginfo.public.
ca.gov
get   README_public_access_
guide_ps
quit.

Anonymous:1995:LRI

[Leg95a] Legal resources via the Internet, page various, 1995. ISSN 1083-0960. LawSource Legal Research Service, Albany, NY, USA.

LeGates:1995:IIB

[LeG95b] John C. LeGates. The Internet: is it a bird?: is it a plane?: will it fly? Incidental paper I-95-5, Program on Information Resources Policy, Harvard University Center for Information Policy Research, Cambridge, MA, USA, December 1995. ISBN 1-879716-27-5. 13 pp.

Leiner:1985:DIP

[Lei85] B. Leiner. The DARPA Internet protocol suite. ISI reprint series ISI/RS-85-153, University of Southern California, Information Sciences Institute, Marina del Rey, CA, USA, 1985. 15 pp.

Leiner:1993:PII

[Lei93] Barry Leiner, editor. Proceedings of INET 93: International Networking Conference, San Francisco, California, USA, 17–20 August 1993: the annual conference of the Internet Society. Internet Society, Reston, VA, USA, 1993. LCCN TK5105.875.I57 I56 1993.

Leiner:1994:IT

[Lei94] Barry M. Leiner. Internet technology. Communications of the ACM, 37(8): 32–33, August 1994. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/

toc/Abstracts/0001-0782/
179611.html.

Leisher:1996:AIUa

[Lei96]  M. Leisher.  An adven-

ture in implementing Unicode support on UNIX platforms. In Unicode Consortium [Uni96a], pages C6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Lemay:1995:TYW

[Lem95] Laura Lemay. Teach Yourself Web Publishing with HTML in a Week. Howard W. Sams, Indianapolis, IN 46268, USA, 1995. ISBN 0-672-30667-0. 397 pp. US$25. This book describes how to write design and publish information on the World Wide Web. In addition to describing the HTML language itself, it provides information on using images, sounds, video, interactivity, gateway programs (CGI), forms, and imagemaps. Through the use of dozens of real-life examples, the book helps you not only learn the technical details of writing Web pages, but also teaches you how to communicate information effectively through the Web.

Leon:1994:IWP

[Leo94a] R. E. Leon. The Internet world: its protocols and mainstream services. Online & CDROM review: the international journal of online & optical information systems, 18(4):229–??, August 1, 1994. CODEN ONCDEW. ISSN 0309-314X.

Leong:1994:WWW

[Leo94b] David Leong. World wide web: accessing information over the Internet. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1994. 17 pp.

Leondes:1998:NNSa

[Leo98a] Cornelius T. Leondes. Neural network systems, techniques, and applications: Algorithms and Architectures, volume 1. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443861-X. 459 pp. LCCN QA76.87 .N4774 1998. US$86.95.

Leondes:1998:NNSg

[Leo98b] Cornelius T. Leondes. Neural Network Systems Techniques and Applications: Control and Dynamic Systems, volume 7. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443867-9. xx + 438 pp. LCCN ???? US$81.95. Seven volumes.

Leondes:1998:NNSf

[Leo98c] Cornelius T. Leondes. Neural network systems, techniques, and applications: Fuzzy Logic and Expert Systems Applications, volume 6. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443866-0. xix + 416 pp. LCCN ???? US$81.95.

Leondes:1998:NNSc

[Leo98d] Cornelius T. Leondes. Neural network systems, techniques, and applications: Implementation Techniques, volume 3. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443863-6. xviii + 401 pp. LCCN ???? US$81.95.

Leondes:1998:NNSd

[Leo98e] Cornelius T. Leondes. Neural network systems, techniques, and applications: Industrial and Manufacturing Systems, volume 4. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443864-4. xx + 395 pp. LCCN ???? US$79.95.

Leondes:1998:NNSb

[Leo98f] Cornelius T. Leondes. Neural Network Systems Techniques and Applications: Optimization Techniques, volume 2.

Academic Press, New York, NY, USA, 1998. ISBN 0-12-443862-8. xxii + 398 pp. LCCN ???? US$86.95.

Leondes:1998:NNS

[Leo98g] Cornelius T. Leondes. Neural Network Systems Techniques and Applications: Seven-Volume Set. Academic Press, New York, NY, USA, 1998. ISBN 0-12-443860-1. ???? pp. LCCN ???? US$457.95.

Lescher:1995:OMR

[Les95] John F. Lescher. Online Market Research: Cost-Effective Searching of the Internet and Online Databases. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-48929-5. xvii + 269 pp. LCCN HF5415.2 .L45 1995. US$19.95. URL http://www.

aw.com/devpress/titles/
48929.html.

Levy:1984:HHC

[Lev84] Steven Levy. Hackers: Heroes of the Computer Revolution. Anchor Press/Doubleday, Garden City, NY, USA, 1984. ISBN 0-385-19195-2 (hardcover), 0-440-13405-6 (paperback). xv + 458 pp. LCCN QA76.6 .L469 1984. US$17.95 (hardcover), US$4.95 (paperback). Tales of the real hackers who invented the modern computer industry. Some of these people are still quite active on the nets today.

Levin:1995:FISa

[Lev95] Jayne Levin. The federal Internet source: a directory of nearly 500 federal, state and political Internet sites, and how to find and use them. National Journal Inc.: NetWeek L.L.C., Washington, DC, USA, 1995. ISBN 0-89234-064-9. viii + 96 pp. LCCN Z1223.Z7 F43 1995.

Levine:1996:IED

[Lev96]  John R. Levine.  Inter-

net E-mail for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-235-X. xxvi + 382 pp. LCCN IN PROCESS.

Long:1995:ENI

[LF95] H. Morrow Long and Sidnie Feit. Enabling networks for Internet access, 1995. 1 videocassette (136 min.)+ 1 computer disk.

Lev:1993:TAN

[LG93] Brian S. Lev and J. Patrick Gary, editors. The Third Annual NASA Science Internet User Working Group Conference: proceedings of a conference sponsored by the National Aeronautics and Space Administration and held in Greenbelt, Maryland, March 30–April 3, 1992, number 3170 in NASA conference publication. National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program; National Technical Information Service, distributor, Washington, DC, USA, 1993. ISBN ???? LCCN ????

Love:1995:DIC

[LG95] P. K. Love and M. V. Gosper. Developing interactive course materials: Using HTML to integrate conventional and Internet resources. In Pearce and Ellis [PE95], pages 359– 367. ISBN 0-7325-1220-4. LCCN ????

Lawrence:1998:SWW

[LG98] Steve Lawrence and C. Lee Giles. Searching the World Wide Web. Science, 280(??): 98–100, April 1998. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lawrence:1999:CAI

[LG99] Steve Lawrence and C. Lee Giles. Commentary: Accessibility of information on the web. Nature, 400(6740): 107–109, July 8, 1999. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

                                        tronic). URL http://www.
                                        nature.com/server-java/
                                        Propub/nature/v400n6740.

LaQuey:1993:IBG

[LGRS93] Tracy L. LaQuey, Al Gore, Jeanne C. Ryer, and Setsu Suzuki. Internet biginazu gaido = The Internet companion. Addison-Wesley Toppan information science series. Toppan Publishing, Tokyo, Japan, 1993. ISBN ???? xiii + 234 pp. LCCN ????

Lambert:1993:IBY

[LH93a] Steve Lambert and Walt Howe. Internet basics: your online access to the global electronic superhighway. Random House Electronic Pub., New York, NY, USA, 1993. ISBN 0-679-75023-1. xxx + 495 pp. LCCN TK5105.875.I57L34 1993.

Lambert:1993:IB

[LH93b] Steve Lambert and Walt Howe. Internet Basics— Your Online Access to the Global Electronic Superhighway. Random House, New York, NY, USA, 1993. ISBN 0-679-75023-1. xxx + 495 pp. LCCN TK 5105.875 I57L348 1993. US$27.00.

Li:1990:SIM

[Li90] Sai Lun Li. A study on Internet management. Thesis (M.S.), University of Washington, Seattle, WA, USA, 1990. vi + 99 pp.

BPL:1995:HM

[Lib95] Berkeley Public Library.

Health and medicine, 1995.

Lichty:1994:AOIa

[Lic94a] Tom Lichty. America Online’s Internet: Easy, Graphical Access — the AOL Way, Macintosh edition. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-175-9. xxxiv + 282 pp. LCCN QA76.57.A43 L5217 1994. US$24.90. The author says: “this book provides a complete description of America Online’s Internet Connection and access to the Internet via AOL. In a jaunty style well-known to readers of Ventana’s America Online Membership Kit, author Tom Lichty points out top Internet attractions and leads his audience through the process of accessing them with America Online. Updates to the book — text and graphics — are available online in AOL’s Internet Connection.”.

Lichty:1994:AOIb

[Lic94b] Tom Lichty. America Online’s Internet: Easy, Graphical Access — the AOL Way, Windows edition. Ventana Press, Chapel Hill, NC, USA, 1994. ISBN 1-56604-176-7. xxxiv + 272 pp. LCCN QA76.57.A43 L5217 1994b. US$24.90. The author says: “this book provides a complete description of America Online’s Internet Connection and access to the Internet via AOL. In a jaunty style well-known to readers of Ventana’s America Online Membership Kit, author Tom Lichty points out top Internet attractions and leads his audience through the process of accessing them with America Online. Updates to the book — text and graphics — are available online in AOL’s Internet Connection.”.

Lichty:1995:AOIa

[Lic95a] Tom Lichty. America Online’s Internet for Macintosh: easy graphical access, the AOL way. Ventana Press, Research Triangle Park, NC, USA, second edition, 1995. ISBN 1-56604-284-4. xxxiv + 360 pp. LCCN QA76.57.A43L522 1995.

Lichty:1995:AOIb

[Lic95b] Tom Lichty. America Online’s Internet for Windows: easy graphical access — the AOL way. Ventana Press, Research Triangle Park, NC, USA, second edition, 1995. ISBN 1-56604-283-6. xxxiv + 360 pp. LCCN QA76.57.A43L5217 1995.

Lin:1993:DIT

[Lin93a] Ning Lin. Diffusion of information technology: a case study of computer network and the role of government, industry, and academia in developing the Internet/NREN. Thesis (Ph.D.), University of Texas at Austin, Austin, TX, USA, 1993. xii + 192 pp.

Lindner:1993:IGU

[Lin93b] P. (Paul) Lindner. Internet gopher user’s guide. Technical report, University of Minnesota, Minneapolis, MN, USA, 1993. vi + 103 pp. RFC 1436.

Lindner:1993:IIS

[Lin93c] Paul Lindner. Internet information services and how to implement: Princeton Marriott Forrestal Village, Plainsboro, NJ, September 9, 1993. Jvncnet seminar series, GES Inc., ????, 1993. various pp.

Littman:1993:CID

[Lit93] Jonathan Littman. Commerce on the Internet: the Digital Gold Rush. UNIX/world, 10(12):42–??, December 1, 1993. ISSN 0739-5922.

Liu:1995:ILI

[Liu95] Lewis-Guodo Liu. The Internet and library and information services: a review, analysis, and annotated bibliography. Occasional papers 202, Graduate School of Library and Information Science, University of Illinois at Urbana-Champaign, Champaign, IL, USA, December 1995. 90 pp.

Liu:1996:ILI

[Liu96] Lewis-Guodo Liu. The Internet and library and information services: issues, trends, and annotated bibliography, 1994–1995. Bibliographies and indexes in library and information science, no.

10. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1996. ISBN 0-313-30019-4. ISSN 0742-6879. ???? pp. LCCN Z674.75.I58L58 1996.

Lamb:1995:CIH

[LJ95] Annette Lamb and Larry Johnson. Cruisin’ the information highway: Internet in the K–12 classroom. Vision to Action, Evansville, IN, USA, 1995. ISBN 0-9641581-1-6. ii + 234 pp. LCCN ????

Levin:1994:FIS

[LJN94] Jayne Levin, National Journal, and NetWeek. The Federal Internet source: a directory of nearly 300 sources of government information on the Internet, and how to find and use them / [Jayne Levin, editor]. National Journal Inc.: NetWeek Inc., Washington, DC, USA, 1994. ISBN 0-89234-062-2. viii + 43 pp. LCCN ????

LeBlanc:1995:ICT

[LL95a] Dee-Ann LeBlanc and Robert LeBlanc. Internet CD tutor, 1995. ISBN 0-7897-0285-1. Includes CD-ROM.

Lynch:1995:DMN

[LL95b] Daniel C. Lynch and Leslie Lundquist. Digital money: the new era of Internet commerce. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-14178-X (paperback). xvi + 285 pp. LCCN HF5415.1265 .L96 1996.

Lynch:1996:DMN

[LL96] Daniel C. Lynch and Leslie Lundquist. Digital money: the new era of Internet commerce. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-14178-X (paperback). xvi + 285 pp. LCCN HF5415.1265 .L96 1996.

Lim:1999:HDQ

[LL99] Ee-Peng Lim and Ying Lu. Harp: a distributed query system for legacy public libraries and structured databases. ACM Transactions on Information Systems, 17(3):294, July 1999. CODEN ATISET. ISSN 1046-8188. URL

http://www.acm.org/pubs/
citations/journals/tois/
1999-17-3/p294-lim/.

Lamersdorf:1998:TDS

[LM98] Winfried Lamersdorf and Michael Merz, editors. Trends in distributed systems for electronic commerce: international IFIP/GI working conference, TREC’98, Hamburg, Germany, June 3–5, 1998, proceedings, volume 1402 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64564-0 (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1402.

Long:1995:LSI

[LMG95] Darrell Long, Andrew Muir, and Richard Andrew Golding. A longitudinal survey of Internet host reliability. Technical report UCSC-CRL-95-16, University of California, Santa Cruz, Computer Research Laboratory, Santa Cruz, CA 95064, USA, 1995. various pp.

Labovitz:1998:IRI

[LMJ98] Craig Labovitz, G. Robert Malan, and Farnam Jahanian. Internet routing instability. IEEE/ACM Transactions on Networking, 6 (5):515–528, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1998-6-5/p515-labovitz/.

Little:1995:SIP

[LMLJ95] David L. Little, Kevin P.

Mullin, Annette C. Lamb, and Larry Johnson. Surfin’ the Internet practical ideas for K–12 classrooms, 1995. 1 videocassette (32 min. + 29 sec.).

Levin:1994:ISG

[LN94] Jayne Levin and NetWeek Inc. Internet software guide — for the PC. Technical report, NetWeek Inc., Washington, DC, USA, 1994. 28 pp.

Levin:1995:FISb

[LN95] Jayne Levin and National Journal, Inc. The federal Internet source. National Journal Inc., Washington, DC, USA, third edition, 1995. ISBN 0-89234-070-3. 143 pp. LCCN ????

Lang:1993:EIb

[LNH^{+}93] Mary Beth Lang, Jim Nocera, Elaine Honomichl, Micheal Maddox, and Tracy Smith. Exploring Internet, 1993. 1 videocassette (135 min.).

Ladd:1999:PIH

[LO99] Eric Ladd and Jim O’Donnell. Programmation Internet, HTML 4, XML et Java R 2: ressources d’experts. CampusPress France, Paris, France, 1999. ISBN 2-7440-0628-9. xvii + 1175 pp. LCCN ????

Loftis:1995:FTU

[Lof95] Karen K. Loftis. Free-nets: a tool for using the Internet in community development. Thesis (M.C.R.P.), University of Nebraska–Lincoln, Lincoln, NE, USA, 1995. vii + 196 pp.

Lathrop:1994:IGI

[LoI94] Lori Lathrop and American Society of Indexers. An indexer’s guide to the Internet. Technical report, The American Society of Indexers, Port Aransa, TX, USA, 1994. 38 pp.

Loney:1994:ODH

[Lon94] Kevin Loney. Oracle DBA handbook. Oracle Press, Berkeley, CA, USA, 1994. ISBN 0-07-881182-1 (hardcover). xxii + 672 pp. LCCN QA76.73.S35 L6 1994. US$34.95. Subtitle on cover: What every system administrator needs to know for effective and efficient database management.

Loshin:1995:TIE

[Los95]  Peter Loshin.  TCP/IP

for everyone. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-455827-5. xv + 425 pp. LCCN TK5105.585.L67 1995.

Lottor:1992:IG

[Lot92] M. Lottor. Internet growth:

1981–1991, January 1992. 1 data file 5 1/4 in. Available via FTP.

Loveman:1999:BCA

[Lov99] Jason Loveman. Building COM Applications with Internet Explorer. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-081909-3. xxii + 476 pp. LCCN QA76.76.A65L68

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0130819093.html. Includes

CD-ROM.

Lowery:1990:PPG

[Low90a] Carlyn M. Lowery. “protocols for providing performance guarantees in a packet-switching Internet”: research project. Master of sciences, plan ii, Department of Electrical Engineering, University of California, Berkeley, Berkeley, CA, USA, 1990. 19 pp.

Lowery:1990:PPP

[Low90b] Carlyn M. Lowery. “protocols for providing performance guarantees in a packet-switching internet”: research project. Master of sciences, plan ii., Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1990. 19 pp.

Lowery:1991:PPP

[Low91] Carlyn M. Lowery. Protocols for providing performance guarantees in a packet switching Internet. Technical report (International Computer Science Institute) TR-91-002, International Computer Science Institute, Berkeley, CA, USA, January 1991. 17 pp.

Lamb:1995:UEE

[LP95] Linda Lamb and Jerry Peek. Using Email Effectively. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1995. ISBN 1-56592-103-8. 160 pp. LCCN ???? US$14.95.

Liu:1994:MII

[LPJ^{+}94] Cricket Liu, Jerry Peek, Russ Jones, Bryan Buus, and Adrian Nye. Managing Internet Information Services: World Wide Web, Gopher, FTP, and more. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1994. ISBN 1-56592-062-7, 1-56592-051-1. xxxvi + 630 pp. LCCN TK5105.875.I57 M36 1994. US$29.95. Written specifically for database maintainers and system administrators, this Unix-based technical guide covers installing, setting up, and running Internet applications such as gopher holes, FTP and telnet sites, mailing lists, WAIS, and World Wide Web pages, and other sites, as well as keeping them secure.

LaQuey:1993:ICBa

[LR93a] Tracy LaQuey and Jeanne C.

Ryer. The Internet companion: a beginner’s guide to global networking. Voyager expanded book. The Voyager Company, New York, NY, USA, first electronic edition, 1993. ISBN ???? LCCN ???? 1 computer disk.

LaQuey:1993:ICP

[LR93b] Tracy LaQuey and Jeanne C.

Ryer. The Internet companion plus: a beginner’s startup kit for global networking: includes Wordlink software to connect to the Internet. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62719-1. x + 197 pp. LCCN TK5105.875.I57 L39 1993. US$19.95 (includes IBM PC diskette). System requirements for computer disk: IBM-compatible PC; 512K RAM; MS-DOS Computer disk contains WorldLink.

LaQuey:1993:ICBb

[LR93c] Tracy L. LaQuey and Jeanne C. Ryer. The Internet Companion: a Beginner’s Guide to Global Networking. Addison-Wesley, Reading, MA, USA, October 1993. ISBN 0-201-62224-6. x + 196 pp. LCCN TK5105.875.I57 L37 1993. US$10.95. This is the least expensive introductory guide for new users of the Internet, and perhaps the only one aimed at the general public. It is also being made available online by anonymous FTP from world.std. com, two chapters a month.

LaQuey:1993:WP

[LR93d] Tracy L. LaQuey and Jeanne C. Ryer. WorldLink for PC-Compatibles, 1993. 1 computer disk.

Lynch:1993:ISH

[LR93e] Daniel C. Lynch and Marshall T. Rose, editors. Internet system handbook. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56741-5. xxxii + 790 pp. LCCN TK5105.875.I57I58 1993. US$40. Reprinted with corrections, January 1993. A description of the process that produces Internet Standards, and some other user level material in addition to protocol material. Essays by people involved with the Internet. Strong on protocols, weak on direction. The book is divided into four parts. The first section deals with an historical and organizational background to the Internet. The second section, written by the people who built the Internet, deals with technical aspects of the major protocols and applications of the Internet. Part three, “Infrastructure,” covers issues not centrally relevant to the operation of the Internet, but supporting its use: network performance and management, backbone and node tools, directory services and operational security. The bibliographic information contained in the book overall might be worth the price alone. Quarterman’s contribution is carefully and fully researched and well organized. It includes not only texts, but periodicals and online sources as well.

LaQuey:1994:QIG

[LR94] Tracy L. LaQuey and Jeanne C. Ryer. Que es Internet guia de principiante a las redes mundiales. Addison-Wesley Iberoamericana, Wilmington, DE, USA, 1994. ISBN 0-201-63110-5. ix + 196 pp. LCCN ????

LaQuey:1995:QI

[LR95a] Tracy L. LaQuey and Jeanne C. Ryer. Que es Internet. Addison-Wesley Iberoamericana, Wilmington, DE, USA, 1995. ISBN 0-201-63110-5. ix + 175 pp. LCCN ????

Levinson:1995:GMI

[LR95b] Jay Conrad Levinson and Charles Rubin. Guerrilla marketing on the Internet: the complete guide to making money on-line. Piatkus, London, UK, 1995. ISBN 0-7499-1542-0 (hardcover), 0-7499-1517-X (paperback). xiii + 303 pp. LCCN ????

Levinson:1995:GMO

[LR95c] Jay Conrad Levinson and Charles Rubin. Guerrilla marketing online: the entrepreneur’s guide to earning profits on the Internet. Houghton Mifflin, Boston, MA, USA, 1995. ISBN 0-395-72859-2. 303 pp. LCCN HF5415.125 .L48 1995.

Lively:1996:JNN

[LR96] G. Martin Lively and Judy A. Reardon. Justice on the net: the National Institute of Justice promotes Internet services. Research in action Research in action (Washington, DC) 96-0186-P, U.S. Dept. of Justice, Office of Justice Programs, National Institute of Justice, Washington, DC, USA, March 1996. 7 + 1 pp.

Lu:1989:SCI

[LS89] W. P. Lu and M. K. Sundareshan. Secure communication in Internet environments: a hierarchical key management scheme for End-to-End encryption. IEEE Transactions on Communications, 37(10): 1014–??, October 1, 1989. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Lane:1993:IPI

[LS93a] Elizabeth S. Lane and Craig A. Summerhill. An Internet Primer for Information Professionals: A Basic Guide to Networking Technology. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-831-X. x + 182 pp. LCCN TK5105.875.I57 L35 1992. US$37.50. This book, not yet seen, is apparently aimed at information professionals, presumably meaning librarians and others.

Lawley:1993:IPI

[LS93b] Elizabeth Lane Lawley and Craig Summerhill. Internet primer for information professionals: a basic guide to Internet networking technology. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-831-X. x + 182 pp. LCCN TK5105.875.I57L35 1993.

Leshin:1995:IAS

[LS95] Cynthia B. Leshin and Dick Strandberg. Internet adventures: step-by-step guide for finding and using educational resources. Xplora Publishing, Paradise Valley, AZ, USA, 1995. ISBN 0-9645588-0-7 (paperback). xiii + 321 pp. LCCN TK5105.875.I57 L47 1995. Educator’s guide for putting classrooms on the internet.

Leshin:1996:IAV

[LS96] Cynthia B. Leshin and Dick Strandberg. Internet adventures, verion 1.2: step-by-step guide for finding and using educational resources. Allyn and Bacon, Needham Heights, MA, USA, 1996. ISBN ???? xv + 332 pp. LCCN ????

Lin:1998:SUD

[LS98] George Lin and Gayla Stewart. AS/400 System, Utilities, Database, and Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1998. ISBN 0-13-083067-4. 696 pp. LCCN ???? URL http:/

/www.phptr.com/ptrbooks/
ptr_0130830674.html.

Levine:1995:ISD

[LSBM95] John R. Levine, Gina Smith, Phillip Byrd, and John L. Meek. The Internet show drivers’ education for the information superhighway, 1995. ISBN 0-7806-0881-X. 1 videocassette (70 min.).

Laquey:1994:SPI

[LSLS94] Tracy L. LaQuey, Olivier Saint-Leger, and Baptiste Su. Sesame pour Internet: initiation au réseau planetaire. Addison-Wesley, Reading, MA, USA, October 1994. ISBN 2-87908-083-5. xi + 236 pp. LCCN ???? Translated by Olivier Saint-Leger and Jean-Baptiste Su. French version of Internet Companion, Second edition.

Ladner:1992:HSL

[LT92] Sharyn J. Ladner and Hope N.

Tillman. How special librarians really use the Internet: summary of findings and implications for the library of the future. ERIC reports ED 345 751, ERIC Document Reproduction Service, operated by CBIS Federal, Inc., Alexandria, VA, USA, 1992. 11 pp.

Ladner:1993:ISL

[LT93] Sharyn J. Ladner and Hope N.

Tillman. The Internet and Special Librarians: Use, Training and the Future. SLA research series; no. 10. Special Libraries Association, Washington, DC, USA, 1993. ISBN 0-87111-413-5. vi + 187 pp. LCCN Z674.82.I59 L33 1993. US$34.95.

Lucas:1994:GIR

[Luc94] Jane Lucas. Guide to Internet resources in criminal law and criminal justice. Technical report, International Centre for Criminal Law Reform and Criminal Justice Policy, Vancouver, BC, Canada, 1994. 44 pp.

Luker:1996:VNN

[Luk96]  Mark Luker.  Viewpoint:

NSF’s new program for highperformance Internet connections. Communications of the ACM, 39(10):27–28, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/pubs/toc/
Abstracts/cacm/236160.html.

Lunzer:1994:PNI

[Lun94] Thomas J. Lunzer. Public networks and the Internet: a reevaluation of security issues. Technical report, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, 1994. 18 pp.

Levine:1994:IDO

[LY94a] John R. Levine and Margaret Levine Young. Internet dla opornych: ciag dalszy. Read me, Warszawa, Poland, 1994. ISBN 83-85769-80-3. 378 pp. LCCN ????

Levine:1994:IDQ

[LY94b] John R. Levine and Margaret Levine Young. The Internet for dummies: quick reference. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1994. ISBN 1-56884-168-X. viii + 165 pp. LCCN TK5105.875.I57L4813 1994.

Levine:1994:MID

[LY94c] John R. Levine and Margaret Levine Young. More Internet for dummies. I D G Books Worldwide, Indianapolis, IN, USA, July 1994. ISBN 1-56884-164-7 (paperback). 390 pp. LCCN TK 5105.875 I57 L483 1994. US$19.95. Contents: Pt. I. Welcome back — Pt.II. Public Internet providers — Pt. III. Windows on the Internet — Pt. IV. Ten thousand interesting things to do on the net. Blurb from the author: “Intended as a follow-on to the best selling ‘Internet for Dummies’, includes in-depth looks at commercial on-line providers including Delphi, AOL and UNIX shell providers. Also SLIP/PPP access including how to get and install, Trumpet Winsock, Trumpet, WinVN, Eudora, WinQVTnet, WSFTP, etc., and an expanded resource list featuring newer gopher and WWW servers.”.

Levine:1995:IDQ

[LY95a] John R. Levine and Margaret Levine Young. The Internet for dummies: quick reference. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1995. ISBN 1-56884-977-X. viii + 210 pp. LCCN TK5105.875.I57L4813 1995.

Levine:1995:ISC

[LY95b] John R. Levine and Margaret Levine Young. Internet su cha shou tse. Meikuo tsui chang hsiao chi suan chi “sha kua” tsung shu. Tien tzu kung yeh chu pan she, Pei-ching, China, ti 1 pan. edition, 1995. ISBN 7-5053-2768-2. III + 215 pp. LCCN ????

Levine:1996:MID

[LY96] John R. Levine and Margaret Levine Young. More Internet for dummies: 2nd edition. I D G Books Worldwide, Indianapolis, IN, USA, second. edition, 1996. ISBN 1-56884-606-1 (paperback). xxvi + 406 pp. LCCN TK5105.875.I57.L483 1996.

Lindholm:1997:JVM

[LY97] Tim Lindholm and Frank Yellin. The Java Virtual Machine Specification. The Java Series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63452-X. xvi + 475 pp. LCCN QA76.73.J38L56 1997. US$36.53. URL http://www.

aw.com/cp/javaseries.html.

Marine:1993:IGS

[M^{+}93] April Marine et al. Internet — getting started. SRI Internet information series Internet information series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, update edition, 1993. ISBN 0-13-327933-2. xxii + 360 pp. LCCN TK5105.875.I57I55 1993.

Menasce:1998:CPW

[MA98] Daniel A. Menascè and Virgilio A. F. Almeida. Capacity Planning for Web Performance: metrics, models, and methods. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-693822-1. xv + 321 pp. LCCN QA76.9.C63M463

1998. URL http://webdevelopersjournal.
com/books/webperformance.
html;   http://www.phptr.
com/ptrbooks/ptr_0136938221.
html.

MacGregor:1993:TMB

[Mac93a] Ronald J. MacGregor. Theoretical Mechanics of Biological Neural Networks. Academic Press, New York, NY, USA, 1993. ISBN 0-12-464255-1. xii + 377 pp. LCCN QP363.3 .M33 1993. US$84.95.

Machovec:1993:TNI

[Mac93b] George Machovec. Telecommunications, networking, and Internet glossary. LITA monographs; 4. Library and Information Technology Association, 50 East Huron Street, Chicago, IL 60602, USA, 1993. ISBN 0-8389-7697-2. viii + 124 pp. LCCN TK5102 .M32 1993, Z680.5 .M35 1993. US$18.00 (LITA members US$16.20). “Greatly revised and expanded edition of the Telecommunications and networking glossary (LITA guides 3) which was published in 1990”.

MacDonald:1995:WII

[Mac95a] Randall M. MacDonald. What is the Internet? and Internet resources for Florida educators. Technical report, Florida Southern College, Lakeland, FL, USA, 1995. 8 pp.

MacLeod:1995:IGL

[Mac95b] Don MacLeod. The Internet guide for the legal researcher: a how-to guide to locating and retrieving free and feebased information on the Internet. Infosources Pub., Teaneck, NJ, USA, 1995. ISBN 0-939486-34-2. 305 pp. LCCN KF242.A1 M3 1995.

Madron:1989:LAN

[Mad89] Thomas W. Madron. LANs— Applications of IEEE/ANSI 802 Standards. Wiley-In-terscience, New York, NY, USA, 1989. ISBN 0-471-62049-1. xi + 308 pp. LCCN TK5105.7.M325 1989. US$34.95.

Madron:1996:CYL

[Mad96] Thomas William Madron.

Connect your LAN to the Internet: cost-effective access for small businesses and other organizations. Wiley Computer Publishers, New York, NY, USA, 1996. ISBN 0-471-14054-6 (paperback). ix + 356 pp. LCCN TK5105.875.I57 M34 1996.

Madsen:1998:IMW

[Mad98] Wayne Madsen. Internet malcontents of the world — unite! Communications of the ACM, 41(6): 27–28, June 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-6/
p27-madsen/.

IWM:1994:OI

[Mag94] Internet World Magazine. On INTERNET 94: An International Title and Subject Guide to Electronic Journals, Newsletters, Texts, Discussion Lists, and Other Resources on the Internet. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-929-4. 500 pp. US$45.00.

“Your guide to the full range of Internet-accessible data files — from artificial intelligence to women’s studies, from space exploration to rock music, from environment studies to AIDS research.” Nearly 6,000 mailing lists, electronic journals, archives, etc.

Mahoney:1995:GBR

[Mah95] Patrick Mahoney. Guide to business resources on the Internet. Technical report, Chester Fritz Library, Grand Forks, ND, USA, 1995. 17 pp.

Mais:1995:EST

[Mai95] Lynn M. Mais. An exploratory study of technology directors and teachers to determine how the Internet is being used as a tool in education with grades five and six in the Minneapolis/St. Paul metro area. Technical report, ????, ????, 1995. 39 pp.

Maitra:1996:BCI

[Mai96] Amit Maitra. Building a corporate Internet strategy: the MIS manager’s guide. Van Nostrand Reinhold, New York, NY, USA, 1996. ISBN 0-442-02246-8. various pp. LCCN HD30.37.M35 1996. Makulowich:1993:INN

[Mak93] John S. Makulowich. Internet: the network of networks. The quill, 81(7):28– ??, September 1, 1993. ISSN 0033-6475.

Malamud:1990:ANN

[Mal90] Carl Malamud. Analyzing Novell Networks. Van Nostrand Reinhold, New York, NY, USA, 1990. ISBN 0-442-00364-1. xv + 343 pp. LCCN TK5105.7 .M35 1990. US$37.95. Chapter 1 quickly reviews the ISO seven-layer model, explains the overall Novell architecture, and introduces the Network General Sniffer network analysis tool. Chapters 2 and 3 explain the Ethernet, Token Ring and ARCNET physical media. Chapter 4 discusses the IPX, SPX and network maintenance protocols; the author unfortunately does not explain why you DON’T want to use Novell’s default Ethernet encapsulation. He also notes that IPX checksumming is disabled on the example he shows, but doesn’t mention that this is usual Novell practice, or explain the implications. This chapter was written before packets greater than 512 bytes and block mode were implemented. Chapter 5 describes the NetWare operating system; the Bindery; communications utilities; file attributes, access rights, locking and synchronization; print services; job servers; and accounting. Chapter 6 describes the Message Handling Service (MHS) and some gateways between it and other type of messaging system. Chapter 7 briefly describes data organization, Btrieve, XQL and NetWare SQL. Chapter 8 describes the Netwise RPC product and its relationship to some other forms of RPC from other vendors. Chapter 9 describes the AT&T STREAMS protocol mechanism and its Novell implementation. Chapter 10 briefly describes the Net-Ware support for AppleTalk. Chapter 11 describes Net-Ware for VMS, Portable Net-Ware (which supersedes Net-Ware for VMS), and how Novell and DECnet networks can be gatewayed to each other. Chapter 12 briefly describes Novell support for WANS. Chapter 13 discusses gateways between Novell and TCP/IP, SNA (described as “not quite as nonproprietary as TCP/IP” in what is surely the understatement of the year!) NetBIOS and OSI. Much of the discussion is fairly superficial. There is, however, a good bibliography which provides references to more detailed documents. The glossary and index look serviceable. All chapters are illustrated with reproductions of Sniffer displays. Table of Contents: 1 Analyzing Networks 2 Data Links and LLC 3 Token Ring and ARC-NET 4 IPX and SPX 5 Net-Ware Core Protocols 6 Message Handling Service 7 Data Access 8 Remote Procedure Calls 9 STREAMS and Protocol Stacks 10 AppleTalk 11 DECnet and Portable Net-Ware 12 X.25 and Wide Area Networks 13 Gateways to Other Networks.

Malamud:1992:EIT

[Mal92] Carl Malamud. Exploring the Internet: a Technical Travelogue. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1992. ISBN 0-13-296898-3. xi + 379 + 5 + 16 pp. LCCN TK5105.875.I57 M36 1992. US$26.95. Many fine lunches and dinners with users, administrators, and developers of the Internet in many countries around the world, with accurate technical background. Certainly the most wickedly funny of all the books about the Internet. Contains a subplot about how and why ITU and ISO do not publish their standards online.

Maloni:1994:N

[Mal94] Kelly Maloni. netchat. Random House, New York, NY, USA, 1994. ISBN 0-679-75814-3. 281 pp. US$19.00, CDN$25.00. Focuses on interacting with others through newsgroups mailing lists, IRC chat groups, forums and support groups — in a word, contact. Coverage is not limited to the Internet, but also the major on-line services such as AOL, CompuServe, Prodigy, and BBSs. Divided into chapters titled Cybersex, Fans, On the Fringe, Net Op-Ed, Coffee House, Identity and Support. There are several listings on where to discuss sexual matters of all kind and even perhaps meet that “special someone” waiting out there just for you.

Malakhi:1995:IKV

[Mal95a]  Oded Malakhi.  Internet

Kompyuserv ve-rishtot meda la-mishtamesh ha-Yisreeli. Opus, Tel Aviv, Israel, 1995. ISBN ???? 366 pp. LCCN ????

Maloff:1995:NEY

[Mal95b] Joel Maloff. Net.profit: expanding your business using the Internet. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-701-7. xxii + 375 pp. LCCN HD30.37.

Maloni:1995:NGY

[Mal95c] Kelly Maloni. Net guide: your complete guide to the Internet and online services. Random House Electronic Publishing: Michael Wolff and Company, New York, NY, USA, second edition, 1995. ISBN 0-679-76456-9 (paperback). xii + 713 pp. LCCN TK5105.875.I57 N48 1995. US$27.95. “A Michael Wolff book.”.

Mamer:1996:THC

[Mam96] Karl Mamer. The truth is in here: conspiracies, mysteries, superstitions, kooks, and cults on the Internet. Go!Guides Go!Guides series. Motion Works Pub., San Francisco, CA, USA, 1996. ISBN 1-57712-011-6 (paperback). ???? pp. LCCN BF1024.65.M36 1996.

Manger:1995:EII

[Man95a] Jason J. Manger. The Essential Internet Information Guide. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-707905-1. ix + 515 pp. LCCN TK5105.875.I57 M364 1995. UK£22.95.

Manger:1995:WWW

[Man95b]  Jason J. Manger.  The

World Wide Web, Mosaic, and More. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-709132-9. vii + 292 pp. LCCN TK5105.875.I57 M367 1995. US$34.95, UK£23.95. The author says: A book on Mosaic and the Web covering many new areas such as publishing on the Net, detailed Common Gateway Interface scripting, fill-out-forms, HTML tutorials, software lists, Web-sites to visit, and much more besides. In summary, this is a novices guide to help people publish their work on the Internet using Mosaic.

Mann:1995:PN

[Man95c] Bill Mann. Politics on the Net. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0286-X. xxvi + 382 pp. LCCN ????

Manber:1996:FDR

[Man96] Udi Manber. Future directions and research problems in the World Wide Web. In ACM [ACM96], pages 213–

215. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996.

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        pods/237661/p213-manber/

. ACM order number:

475960.

Markiv:1993:IRE

[Mar93] Michael Markiv. Internet for Russian and East European studies. College and Research Libraries News, 54(8):444–??, September 1, 1993. ISSN 0099-0086.

Marcos:1994:ILT

[Mar94] Kathleen Marcos. Internet for language teachers. ERIC digest ERIC digest (Washington, DC, USA) EDO-FL-04-95, ERIC Clearinghouse on Languages and Linguistics, Center for Applied Linguistics, Washington, DC, USA, 1994.

Maran:1995:IWW

[Mar95a] Ruth Maran. Internet and World Wide Web simplified. IDG’s 3-D visual series. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-658-4. v + 213 pp. LCCN ????

Marchuk:1995:BIA

[Mar95b] Michael Marchuk. Building Internet applications with Visual Basic. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0214-2. xvi + 483 pp. LCCN TK5105.875.I57 M365 1995.

Anonymous:1995:MLI

[Mar95c] Marketing lists on the Internet, page various, 1995. ISSN 1084-7391. WolfBayne Communications, Colorado Springs, Co.

Marshall:1996:SGI

[Mar96a] Elizabeth L. Marshall. A student’s guide to the Internet: exploring the World Wide Web, Gopherspace, electronic mail, and more! Millbrook Press, Brookfield, CT, USA, 1996. ISBN 1-56193-923-3 (??invalid checksum). 159 pp. LCCN TK5105.875.I57 M3675 1996.

Marulanda:1996:IPR

[Mar96b] Richard L. Marulanda. The Internet as a public relations tool: a guide for universities and colleges. Thesis (M.A.), Fairleigh Dickinson University, Florham-Madison Campus, Madison, NJ, USA, 1996. 37 pp.

Martin:1998:IAT

[Mar98] David M. Martin. Internet anonymizing techniques. ;login: the USENIX Association newsletter, 23(3): ??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
1998-5/martin.html. Spe-

cial issue on security.

Martin:1999:TSI

[Mar99] Fredrick Thomas Martin. Top secret intranet: how U.S. intelligence built Interlink — the world’s largest, most secure network. Charles F. Goldfarb series on open information management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080898-9. 380 pp. LCCN JK468 .I6M413 1999. US$34.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130808989.html.

HRI:1989:MLH

[Mas89] Master list of human rights organizations and serial publications, page various, 1989. Human Rights Internet, Harvard Law School, Cambridge, MA, USA.

Masters:1993:PNN

[Mas93] Timothy Masters. Practical Neural Network Recipes in C++. Academic Press, New York, NY, USA, 1993. ISBN 0-12-479040-2. xviii + 493 pp. US$65.95.

HRI:1994:M

[Mas94] Masterlist, page various, 1994. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada.

MVW:1995:IHP

[Mas95] Master Video Works. The Internet: how to plug in, 1995. 1 videocassette (32 min.).

MVW:1996:YSI

[Mas96] Master Video Works. Your school and the Internet, 1996. 1 videocassette (39 min.).

Masys:1997:AMI

[Mas97] Daniel R. Masys, editor. The emergence of ‘internetable’ health care: systems that really work: proceedings, 1997 AMIA Annual Fall Symposium, formerly SCAMC, October 25–29, 1997, Opryland Hotel, Nashville, TN, AMIA ANNUAL SYMPO-SIUM 1997. Hanley and Belfus, Philadelphia, PA, USA, 1997. ISBN ???? ISSN 1091-8280. LCCN W26.5 S978 1997.

MathWorks:1993:GST

[Mat93]  MathWorks, Inc.  Guide

to services: technical support, product updates, Internet services, newsletter, short courses: for UNIX and VMS users. Technical report, MathWorks, Inc., Natick, MA, USA, September 1993. 16 pp.

Mathiesen:1995:MIP

[Mat95a] Michael Mathiesen. Marketing on the Internet: a proven 12-step plan for promoting, selling, and delivering your products and services to millions over the information superhighway. Maximum Press [Distributed by Publishers Group West], Gulf Breeze, FL, USA, 1995. ISBN 1-885068-01-8 (paperback). xxii + 400 pp. LCCN TK5105.875.I57M3677 1996.

Matthews:1995:IW

[Mat95b] Judith Matthews. Internet world ‘94. Searcher: the magazine for database professionals, 3(3):34–??, March 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Mathiesen:1996:MI

[Mat96a] Michael Mathiesen. Marketing on the Internet. Maximum Press, Gulf Breeze, FL, USA, second edition, 1996. ISBN 1-885068-09-3 (paperback). ???? pp. LCCN HF6146.I58M38 1996.

Mathieu:1996:MI

[Mat96b] Richard G. Mathieu. Manufacturing and the Internet. Engineering and Management Press, Norcross, GA, USA, 1996. ISBN 0-89806-164-4. ???? pp. LCCN TS149.M313 1996.

Maufer:1999:DIM

[Mau99a] Thomas A. Maufer. Deploying IP Multicast in the Enterprise. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-897687-2. xvii + 275 pp. LCCN TK5105.887.M39

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0138976872.html.

Maufer:1999:IFW

[Mau99b] Thomas A. Maufer. IP Fundamentals: What Everyone Needs to Know About Addressing and Routing. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-975483-0. 450 pp. LCCN TK5105.585.M38

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0139754830.html.

Maxwell:1995:HAF

[Max95a] Bruce Maxwell. How to access the federal government on the Internet. Congressional Quarterly, Washington, DC, USA, 1995. ISBN 1-56802-034-1. xviii + 402 pp. LCCN Z1223.Z7 M52 1995.

Maxwell:1995:SBF

[Max95b] Bruce Maxwell. Some of the best federal government Internet sites for law librarians. Technical report, Bates Information Services, Washington, DC, USA, 1995. 20 pp.

Maxwell:1995:WO

[Max95c] Washington online, page various, 1995. ISSN 1088-7466. Congressional Quarterly, Washington, DC, USA.

Maxwell:1995:MID

[Max95d] Christine Maxwell. McKinley Internet directory. New Riders Publishing, Carmel, IN, USA, new edition, 1995. ISBN 1-56205-440-6. xii + 802 (or viii + 820??) pp. LCCN ????

Maxymuk:1995:FGI

[Max95e] John Maxymuk. Finding government information on the Internet: a how-to-do-it manual. How-to-do-it manuals for librarians; no. 52 Howto-do-it manuals for libraries; no. 52. Neal-Schuman Publishers, Inc., New York, NY, USA, 1995. ISBN 1-55570-228-7. xv + 264 pp. LCCN Z1223.Z7M26 1995 J83.

Mayer:1996:CJF

[May96] Doug Mayer. Cyber jokes: the funniest stuff on the Internet. Andrews and McMeel, Kansas City, KS, USA, 1996. ISBN 0-8362-1075-1. ???? pp. LCCN PN6231.I62C93 1996.

Mazraani:1990:HSI

[Maz90] Tony Y. Mazraani. High speed Internet protocols and resource management in the Ethernet. Thesis (M.S.), Department of Electrical Engineering, Washington University, St. Louis, MO, USA, 1990. x + 136 pp.

McKinney:1993:EMA

[MB93] W. McKinney and Glenn Bunton. Exploring the medical applications of the Internet: A guide for beginning users. The American journal of the medical sciences, 306(3):141–??, September 1, 1993. CODEN AJMSA9. ISSN 0002-9629.

McClure:1995:ICC

[MBB95] Charles R. McClure, John Carlo Bertot, and John C. Beachboard. Internet costs and cost models for public libraries: final report. U.S. National Commission on Libraries and Information Science: For sale by the U.S. G.P.O., Supt. of Docs., Washington, DC (1110 Vermont Avenue, N.W., Washington 20005-3522), June 1995. ISBN 0-16-048113-9. vii + 47 pp. LCCN Z674.75.I8M33 1995.

Mintzer:1996:TOW

[MBC^{+}96] F. C. Mintzer, L. E. Boyle, A. N. Cazes, B. S. Christian, S. C. Cox, F. P. Giordano, H. M. Gladney, M. L. Kelmanson, A. C. Lirani, K. A. Magerlein, A. M. B. Pavani, and F. Schiattarella. Toward on-line, worldwide access to Vatican Library materials. IBM Journal of Research and Development, 40(2):139–162, March 1996. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

http://www.almaden.ibm.

                                        com/journal/rd40-2.html#
                                        two.

McClure:1991:NRE

[MBDR91] Charles McClure, Ann Bishop, Philip Doty, and Howard Rosenbaum. The National Research And Education Network (NREN): Research and Policy Perspectives. Ablex Publishing Corp., 355 Chestnut St., Norwood, NJ 07648, USA, 1991. ISBN 0-89391-813-X. ii + 744 pp. LCCN Q179.94 .M33 1991. US$45 (personal), US$95 (institutional). Masses of information on NREN, and examination of its potential effects on research.

Machover:1995:GTI

[MBM^{+}95] Carl Machover, Gavin Bell, Tamara Munzner, Fabio Pettinati, and Val Watson. 3D graphics through the Internet — a “shoot-out” (panel session). Computer Graphics, 29(Annual Conference Series):508–509, November 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/graph/218380/
                                        p508-machover/.

Murray:1995:HGI

[MBS95] Laura K. Murray, Karen Bordonaro, and Nancy Steele. The homesteader’s guide to the Internet frontier. Technical report, State University of New York, SUNY/OCLC Network, Office of Library Services, Albany, NY, USA, 1995. 57 pp.

McClure:1994:PLI

[MBZ94] Charles R. McClure, John Carlo Bertot, and Dougles Zweizig. Public libraries and the Internet: study results, policy issues and accommodations: final report. Technical report, U.S. National Commission on Libraries and Information Science, Washington, DC, USA, June 1994. ix + 62 pp.

McAteer:1995:MIH

[McA95] J. Davitt McAteer. MSHA Internet home page. Program information bulletin; no. p 95-28, U.S. Dept. of Labor, Mine Safety and Health Administration, Arlington, Va. (4015 Wilson Blvd., Arlington 22203-1984), 1995. 2 pp.

McBride:1995:IMS

[McB95] P. K. McBride. The Internet made simple. Made simple computer books. Made Simple Books, Oxford, England, 1995. ISBN 0-7506-2311-X (paperback). ix + 149 pp. LCCN ????

McBride:1996:IW

[McB96] P. K. McBride. The Internet for Windows 95. Clear and simple. Butterworth-Heinemann, Boston, MA, USA, 1996. ISBN 0-7506-9801-2 (paperback). ix + 149 pp. LCCN TK5105.875.I57M37 1996.

McConnell:1988:ICS

[McC88] John McConnell. Internetworking computer systems: interconnecting networks and systems. Prentice Hall advanced reference series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-473976-0. xv + 318 pp. LCCN TK5105.5.M4 1988.

McClure:1992:PLI

[McC92] Charles R. McClure. Public libraries and the INTER-NET/NREN: new challenges, new opportunities. Technical report, School of Information Studies, Syracuse University, Syracuse, NY, USA, July 1, 1992. ii + 34 + 5 pp.

McClure:1994:PGRa

[McC94a] Charles R. McClure. The project GAIN report: connecting rural public libraries to the Internet. Technical report, NYSERNet, Liverpool, NY, USA, February 15, 1994. various pp.

McClure:1994:PGRb

[McC94b] Charles R. McClure. The Project GAIN report: connecting rural public libraries to the Internet: project evaluation report. Technical report, Information Management Consultant Services, Inc., Manlius, NY, USA, February 15, 1994. various pp.

McCall:1995:PUI

[McC95] Dolores E. McCall. Physician use of the Internet: its impact on clinical decision making and patient care. Thesis (M.S.), Florida State University, Tallahassee, FL, USA, 1995. ix + 76 pp.

McCormick:1997:LIK

[McC97] Kevin McCormick. Linux as an Internet kiosk: Cybercafé — a company that uses Linux to set up public Internet terminals. Linux Journal, 42:??, October 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

McFedries:1995:CIG

[McF95] Paul McFedries. The Complete Idiot’s Guide to Internet E-Mail. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-596-1. xxii + 312 pp. LCCN TK5105.875.I57M382 1995. US$16.99.

McKenna:1994:LI

[McK94a] Mary McKenna. Libraries and the Internet. Technical report, ERIC Clearinghouse on Information and Technology, Syracuse University, Syracuse, NY, USA, 1994.

McKie:1994:ITG

[McK94b] Craig McKie. The Internet Toolkit: Getting Ready for the 21st Century. Thompson Educational Publishing, Toronto, ON, Canada, December 1994. ISBN 1-55077-066-7. 360 pp. US$24.95.

McKie:1995:ITG

[McK95] C. (Craig) McKie. The Internet toolkit: getting ready for the 21st century. Thompson Educational Pub., Toronto, ON, Canada, 1995. ISBN 1-55077-066-7. x + 324 pp. LCCN ????

McKenzie:1996:MII

[McK96a] Bruce C. McKenzie. Medicine and the Internet: introducing online resources and terminology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-262705-8 (paperback). xiii + 209 pp. LCCN R119.95 .M37 1996.

McKim:1996:IRC

[McK96b] Geoffrey W. McKim. Internet research companion. Que E and T, Indianapolis, IN, USA, 1996. ISBN 1-57576-050-9 (paperback). vi + 266 pp. LCCN TK5105.875.I57M3828 1996.

McLaren:1996:UUI

[McL96a] Bruce J. McLaren. Understanding and using the Internet. West’s microcomputing series. West Publishing Company, St. Paul, MN, USA, 1996. ISBN 0-314-06411-7 (softcover). xvii + 169 pp. LCCN TK5105.875.I57M383 1996.

McLoughlin:1996:WIB

[McL96b] Henry McLoughlin. WEST:

An Internet based education delivery and support environment. Computer Networks and ISDN Systems, 28(14):1887–1890, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1996/
28/14/1648.pdf.

McMullen:1994:NII

[McM94] Melanie McMullen. Networks 2000: Internet, information superhighway, multimedia networks and beyond. LAN networking library. M. Freeman Books, San Francisco, CA, USA, 1994. ISBN 0-87930-335-2. 338 pp. LCCN TK5105.5.N49 1994. US$24.95. “The best articles on emerging network technology from … LAN Magazine”–P. [4] of cover.

McNamara:1985:LAN

[McN85] John E. McNamara. Local Area Networks—An introduction to the technology. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Englewood Cliffs, NJ 07632, USA, 1985. ISBN 0-932376-79-7 (DP), 0-13-539669-7 (PH). viii + 165 pp. LCCN TK5105.7.M36 1985. US$30.00.

McNall:1995:WWW

[McN95] Ralph I. McNall. W3D, World Wide Web 3D browser for “surfing the internet” final report … December 15, 1994 thru June 14, 1995. NASA contractor report NASA CR-199156, National Aeronautics and Space Administration, Washington, DC, USA, 1995.

McQueen:1994:GMB

[McQ94a] Howard McQueen. Getting the most out of basic Internet tools: course 230. Technical report, CD Consultants, Inc., Washington, DC, USA, April 21, 1994. 175 pp.

McQueen:1994:IPT

[McQ94b] Howard McQueen. Interfaces, power tools and connectivity strategies: course 240. Technical report, CD Consultants, Inc., Washington, DC, USA, April 21, 1994. 179 pp.

Muller:1991:GFR

[MD91] Nathan J. Muller and Robert P. Davidson. The guide to frame relay & fast packet networking. Technical report, Telecom Library Inc., 12 West 21 Street, New York NY 10010, USA, Tel: +1 212 691 8215, +1 800 LI-BRARY, FAX: +1 212 691 1191, 1991. Numerous new technologies are emerging to support wide area networking; among them are frame relay, cell relay (ATM), SMDS, SONET, SDH, ISDN and BISDN. This confusing jumble of acronyms and technologies deserves a careful, thorough explanation. Unfortunately the authors did not rise to the occasion; instead they produced a jumble of unexplained acronyms and confused descriptions. Not recommended. Table of Contents 1 Public Network Trends 2 WAN Technologies 3 Principles of Frame Relay 4 Principles of Frame Relay ISDN 5 Principles of Cell Relay/ATM 6 Frame Relay Enablers: LANs, Fiber and their Interconnection 7 Carrier Services 8 Products and Networks 9 Network Planning and Design 10 Planning for Frame Relay 11 Managing Technology Acquisitions in a Down Economy Bibliography List of Acronyms Index.

Mikita:1993:IHS

[MD93] Elizabeth G. Mikita and Louise A. Drusedum. Introducing health sciences librarians to the Internet. Medical reference services quarterly, 12(3):1–??, Fall 1993. CO-DEN MRSQDK. ISSN 0276-3869.

Mailer:1994:USI

[MD94] Nicholas Mailer and Bruce Dickson. UK school Internet primer. Koeksuster Publications, London, UK, 1994. ISBN 0-9524072-0-5. 168 pp. LCCN ????

Meeker:1996:IR

[MD96] Mary Meeker and Chris DePuy. The Internet report. HarperBusiness, New York, NY, USA, 1996. ISBN 0-88730-826-0. various pp. LCCN HE7568.M44 1996.

Metzler:1999:LSG

[MD99] James A. Metzler and Lynn A. DeNoia. Layer 3 Switching: a Guide for IT Professionals. Prentice Hall series in computer networking and distributed systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-919838-5. xv + 232 (or xvi + 232??) pp. LCCN TK5015.7 .M47 1999. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0139198385.html.

Malamud:1994:FIP

[MDB^{+}94] Carl Malamud, Steve Deering, Bob Braden, Christian Huitema, Bob Hinden, Peter Ford, Steve Casner, Craig Partridge, Noel Chiappa, and Martin Lucas. The future of the Internet protocol, 1994. ISBN 1-56592-996-9. 4 sound cassettes (240 min.). Originally “broadcast” across the Internet from Washington, DC, USA on the program Geek of the week. “Internet talk radio” Carl Malamud interviews key players in the development of Internet: Steve Deering, Bob Braden, Christian Huitema, Bob Hinden, Peter Ford, Steve Casner, Craig Partridge, and Noel Chiappa.

Milojicic:1999:MPC

[MDW99] Dejan Milojicic, Frederick Douglis, and Richard Wheeler, editors. Mobility: processes, computers, and agents. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37928-7. xii + 682 pp. LCCN QA76.76.I58M65 1999. US$39.95.

McKenna:1992:SNC

[MDZ92] Thomas McKenna, Joel Davis, and Steven F. Zornetzer. Single Neuron Computation. Academic Press, New York, NY, USA, 1992. ISBN 0-12-484815-X. xiv + 644 pp. LCCN QA76.87 .S57 1992. US$102.95.

Mirchandani:1995:AII

[ME95] V. Mirchandani and D. Everitt.

An architecture for internetworking integrated services LANs. In IEEE Computer Society. Technical Committee on Computer Communications [IEE95], pages 433– 442. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

MartinezValero:1996:IPT

[ME96] Julian Martinez Valero and Carlos Esebbag Benchimol. Internet para torpes. Informatica para torpes. Anaya Multimedia, Madrid, Spain, 1996. ISBN 84-7614-814-3. 367 pp. LCCN ????

Mirchandani:1997:PIT

[ME97] V. Mirchandani and D. Everitt. Performance of integrated telephony service over standalone LAN and ATM Internetworked LANs. In IEEE Computer Society. Technical Committee on Computer Communications [IEE97], pages 80–88. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Mecklermedia:1994:OD

[Mec94] Mecklermedia. OPAC Directory 1994. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-962-6. 500 pp. US$70. A detailed listing of dial-in online public access catalogs and databases. Includes Accessing Online Bibliographic Databases, the annotated list of 700+ Internetaccessible OPACs.

Mecklermedia:1995:IW

[Mec95a] Mecklermedia Corporation.

Internet world, 1995. Includes CD-ROM.

Anonymous:1995:MOI

[Mec95b] Mecklermedia’s official Internet world World Wide Web yellow pages, page various, 1995. I D G Books Worldwide, Indianapolis, IN, USA.

IW:1996:MOI

[Mec96] Mecklermedia’s official Internet world Internet yellow pages, page various, 1996. ISSN 1087-6111. IDG Books Worldwide, Inc., Foster City, CA, USA.

MorrisMedia:1994:IC

[Med94a] Morris Media. The Internet connection, 1994. 1 videocassette (60 min.).

MorrisMedia:1994:ICY

[Med94b] Morris Media. The Internet connection your on ramp to the information superhighway, 1994. ISBN 1-55746-328-X. 2 videocassettes diskettes (3 1/2 in.) and instructional booklet. Anonymous:1996:MTW

[Med96] Med trib web: a physician’s guide to the Internet, page various, 1996. Medical Tribune, New York, NY, USA.

Meeks:1995:EFT

[Mee95]  Brock N. Meeks.  Elec-

tronic frontier: Target: Internet. Communications of the ACM, 38(8):23–25, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
208375.html.

MeierzuSieker:1982:TGI

[Mei82] Friedrich Meier zu Sieker. A telex gateway for the Internet. Thesis (B.S.), M.I.T., Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1982. 1 + iii + 59 pp. Supervised by Jerome H. Saltzer.

Meizel:1992:HSE

[Mei92] Janet Meizel. High school education and the Internet: The Davis Senior High School experience. Resource sharing & information networks, 8(1): 127–??, 1992. ISSN 0737-7797.

Meinke:1997:CCS

[Mei97] J. G. Meinke, editor. Consortium for Computing in Small Colleges: Southeastern conference; 10th — November 1996, Johnson City, TN, volume 12(3) of JOURNAL OF COMPUTING IN SMALL COLLEGES 1997. Consortium for Computing in Small Colleges, ????, 1997. ISBN ???? LCCN ????

Mellendorf:1994:AAI

[Mel94] Scott A. Mellendorf. Automating access to Internet resources at the reference desk. Online, 18(5):69–??, September 1, 1994. CODEN ONLIDN. ISSN 0146-5422.

Merola:1987:VUR

[Mer87] Sandy Merola. View of U.S.

research networks. Computer Networks and ISDN Systems, 13(3):249–??, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

MNI:1992:MNI

[Mer92] Merit Network, Inc. Merit network, inc. presents an Internet cruise, 1992. 1 computer disk.

MNI:1993:MNI

[Mer93] Merit Network, Inc. Merit network, inc. presents an Internet cruise, 1993. 1 computer disk.

Messerschmitt:1999:NAG

[Mes99] David Messerschmitt. Networked Applications: a Guide to the New Computing Infrastructure. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-536-3. xx + 396 pp. LCCN QA76.625 .M47 1999. US$44.95.

Meyer:1995:TUI

[Mey95] Margaret Meyer. Teacher uses of the Internet. Graduate paper (m.s. in i.m.), St. Cloud State University, St. Cloud, MN, USA, 1995. vi + 96 pp.

Maxwell:1994:NROa

[MG94a] Christine Maxwell and Czeslaw Jan Grycz. New Rider’s Official Internet Yellow Pages. New Riders Publishing, Carmel, IN, USA, 1994. ISBN 1-56205-306-X. xii + 877 pp. LCCN TK5105.875.I57 M368 1994. US$29.95. Another phonebook collection (better?) of information/topics for Internet explorers ….

Maxwell:1994:NROb

[MG94b] Christine Maxwell and Czeslaw Jan Grycz. New Riders’ official Internet yellow pages. New Riders Publishing, Carmel, IN, USA, second edition, 1994. ISBN 1-56205-408-2. xii + 802 pp. LCCN TK 5105.875 I57 M368 1994. US$29.99.

McBride:1994:IWP

[MG94c] James S. McBride and Seth Godin. The Internet white pages. IDG Books, San Mateo, CA, USA, 1994. ISBN 1-56884-300-3. ISSN 1080-7675. xxiv + 812 pp. LCCN TK5105.875.I57M38 1994.

Marx:1996:EI

[MG96] Gretchen Marx and Robert T. Grauer. Exploring the Internet. Exploring Windows 95 series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-504051-5. xvi + 287 pp. LCCN TK5105.875.I57M3677 1996.

Michael:1994:AZ

[MH94] James Michael and Mark Hinnebusch. From A to Z39.50: a Network Primer. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-766-6. 225 pp. LCCN TK5105.5 .M534 1994. US$25.00.

Mills:1991:IAB

[MHR91] C. Mills, D. Hirsh, and G. Ruth. Internet accounting background, November 1991. 1 data file 5 1/4 in. Available via FTP.

Miller:1989:LTH

[Mil89]  Mark A. Miller.  LAN

Troubleshooting Handbook. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1989. ISBN 1-55851-054-0 (book), 1-55851-056-7 (book/disk), 1-55851-055-9 (disk). 309 pp. LCCN TK5105.7 .M55 1989. US$29.95 (book), US$39.95 (book/disk), US$20.00 (disk).

Mills:1990:MPN

[Mil90] David L. Mills. Measured performance of the Network Time Protocol in the DARPA/NSF Internet system. Computer Communication Review, 20(1):65–??, January 1, 1990. CODEN CCRED2. ISSN 0146-4833.

Miller:1991:IGN

[Mil91a] Mark A. Miller. Internetworking—A Guide to Network Communications—LAN to LAN, LAN to WAN. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1991. ISBN 1-55851-143-1. iii + 425 pp. LCCN TK5105.5 .M55 1991. US$34.95.

Miller:1991:TI

[Mil91b]  Mark A. Miller.  Trou-

bleshooting Internetworks —Tools, Techniques, and Protocols. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1991. ISBN 1-55851-236-5. LCCN TK5105.5.M56 1991. US$34.95.

Miller:1992:TTI

[Mil92a] Mark A. Miller. Troubleshooting TCP/IP : Analyzing the Protocols of the Internet. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-268-3, 3-88229-025-0, 0-13-953167-X. 588 pp. LCCN TK5105.5 .M52 1992. US$44.95.

Milles:1992:IUI

[Mil92b] An introduction to using the Internet at Saint Louis University, page various, 1992.

St. Louis University School of Law, St. Louis, MO, USA.

Miller:1993:LTH

[Mil93a]  Mark A. Miller.  LAN

Troubleshooting Handbook. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 1993. ISBN 1-55851-301-9. LCCN TK5105.7 .M55 1993. US$34.95.

Miller:1993:MIS

[Mil93b] Mark A. Miller. Managing Internetworks with SNMP. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1993. ISBN 1-55851-304-3. xx + 507 pp. LCCN TK5105.55 .M55 1993. US$44.95.

Milles:1993:IHL

[Mil93c] James Milles. Internet handbook for law librarians. A law library information report, Law library information reports. Glanville Publishers, New York, NY, USA, 1993. ISBN 0-87802-093-4. ISSN 0268-8336. 64 pp. LCCN Z674.82.I59M55 1993.

Milles:1993:IUI

[Mil93d] James Milles. An introduction to using the Internet. Technical report, St. Louis University School of Law, St. Louis, MO, USA, 1993. 54 pp.

Miller:1994:IRD

[Mil94a] Elizabeth B. Miller. The Internet resource directory for K–12 teachers and librarians. Libraries Unlimited, Englewood, CO, USA, 94/95 ed edition, 1994. ISBN 1-56308-337-X. xvii + 199 pp. LCCN TK5105.875.I57 M55 1994.

Miller:1994:IKS

[Mil94b] Elizabeth Bowden Miller. The Internet for K–12 schools: accessing learning resources on the information superhighway. Thesis (S.L.I.S), University of South Carolina, Columbia, SC, USA, 1994. iii + 199 pp.

Miller:1994:EI

[Mil94c] Michael Miller. Easy Internet. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 0-7897-0012-3. ix + 228 pp. LCCN TK5105.875.I57M554 1994.

Miller:1994:BGP

[Mil94d] Michael J. Miller. Bill Gates ponders the Internet. PC Magazine, 13(17):79–??, October 11, 1994. ISSN 0888-8507.

Milles:1994:IUI

[Mil94e] James Milles. An introduction to using the Internet at Saint Louis University School of Law. Technical report, St. Louis University School of Law, St. Louis, MO, USA, 1994.

Milner:1994:ILG

[Mil94f] Jason Milner. Internet logging on in Gist 215, 1994. 1 videocassette (4 min.).

Miller:1995:IRDa

[Mil95a] Elizabeth B. Miller. The Internet resource directory for K–12 teachers and librarians. Libraries Unlimited, Englewood, CO, USA, 1995. ISBN ???? ISSN 1084-5798. various pp. LCCN ????

Miller:1995:IRDb

[Mil95b] Elizabeth B. Miller. The Internet resource directory for K–12 teachers and librarians: 1995 spring update. Technical report, Libraries Unlimited, Englewood, CO, USA, 1995. iii + 11 pp.

Miller:1995:IAG

[Mil95c] Mark A. Miller. Internetworking—A Guide to Network Communications—LAN to LAN, LAN to WAN. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 1995. ISBN 1-55851-436-8. xxii + 530 pp. LCCN TK5105.5 .M55 1995. US$44.95.

Miller:1995:EI

[Mil95d] Michael Miller. Easy Internet. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0012-3. 246 pp. US$24.99. The publisher says: “A fully illustrated step-by-step guide of how to best access and use the powerful tools of the Internet. E- mail, Usenet, Gopher, FTP, Telnet, WWW are clearly explained enabling the beginner to master these basics in a time efficient manner.”.

Millett:1995:ISE

[Mil95e] Neal G. Millett. The integration of the space environment laboratory’s product delivery systems onto Internet. Tmaster of science in computer and information science. thesis (m.s.c.i.s.), Regis University, Denver, CO, USA, 1995. 64 pp.

Mills:1995:PPD

[Mil95f] David L. Mills. Performance and policy dimensions in Internet routing. NASA contractor report NASA CR-197772. Technical report 95-4-1, National Aeronautics and Space Administration, Washington, DC, USA, 1995.

Miller:1996:IRD

[Mil96a] Elizabeth B. Miller. The Internet resource directory for K–12 teachers and librarians. Libraries Unlimited, Englewood, CO, USA, 95/96 edition, 1996. ISBN 1-56308-366-3. xv + 227 pp. LCCN ????

Miller:1996:TTI

[Mil96b] Mark Miller. Troubleshooting TCP/IP: analyzing the protocols of the Internet. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 1996. ISBN 1-55851-450-3. ???? pp. LCCN TK5105.585.M56 1996.

Miller:1996:NBG

[Mil96c] Mary Frances Miller. Net-Practice: a beginner’s guide to healthcare networking on the Internet. Opus Communications, Marblehead, MA, USA, 1996. ISBN 1-885829-27-2. xiii + 234 pp. LCCN W 26.5 M649n 1996. US$37.00.

Miller:1997:TIE

[Mil97] Philip Miller. TCP/IP Explained. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-166-8. xviii + 518 pp. LCCN TK5105.585 .M564 1997. US$49.95.

Miletsky:1999:WPG

[Mil99a] Jason Miletsky. Web Photoshop 5 To Go. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-011848-6. 320 pp. LCCN T385.M536

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130118486.html.

Miller:1999:MNA

[Mil99b] C. Kenneth Miller. Multicast networking and applications. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30979-3. xii + 282 pp. LCCN TK5105.887.M56 1999. US$39.95.

Minchek:1994:SI

[Min94a] Helen Minchek. Surfing the Internet, 1994. 1 videocassette (120 min.).

MESRCGII:1994:CLT

[Min94b] Ministère de l’Enseignement Supérieur et de la Recherche and Groupement français de l’industrie et de l’information, Paris, France. La Circulation de l’information à travers les réseaux Renater et Internet: Actes du colloque organisé le 11 fėvrier 1994 à Paris, 1994. various pp.

Minatel:1995:EWW

[Min95] Jim Minatel. Easy World Wide Web with Netscape. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0279-7. 208 pp. LCCN ???? US$19.99.

Miser:1995:IQW

[Mis95]  Brad Miser.  Internet

QuicKIT for Windows 95. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-240-1. xii + 176 pp. LCCN ????

Mitchell:1991:CND

[Mit91] B. S. (Barry S.) Mitchell.

Communications network development. Project Report A-8779, Georgia Tech Research Institute, Georgia Institute of Technology, Atlanta, GA, USA, 1991. various pp. Four volumes.

Mitchell:1992:TTE

[Mit92] B. S. (Barry S.) Mitchell.

TERIS (test and evaluation range Internet system) network development planning. Technical report, Georgia Tech Research Institute, Georgia Institute of Technology, Atlanta, GA, USA, 1992. various pp. Four volumes. Issued as Project management plans [nos. 1–2], and Reports [nos. 1–2], Project A-9214.

Marine:1992:IGSa

[MK92] April Marine and Susan Kirkpatrick. Internet: getting started. Technical report, SRI International, Menlo Park, CA, USA, 1992. vii + 312 pp.

Miller:1995:IDC

[MK95] Robert (Robert William)

Miller and Elissa Keeler. Internet direct: connecting through SLIP and PPP. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1995. ISBN 1-55828-432-X. x + 422 pp. LCCN TK5105.875.I57 M556 1995. ????

Millrod:1996:SI

[MK96a] Leslie Millrod and Mary Kelly. Surfing the Internet. Technical report, Huntington Public Library, Huntington, NY, USA, March 1996. 35 pp.

Musciano:1996:HDG

[MK96b] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-175-5. xx + 385 pp. LCCN QA76.76.H94 M87 199. US$29.95, CDN$42.95.

                                        URL http://www.oreilly.
                                        com/catalog/9781565921757.

MacHale:1998:LH

[MK98a] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.x, 1998. 6 videocassettes.

MacHale:1998:LHLa

[MK98b] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 1, 1998. 1 videocassette (126 min.) disk (3 1/2 in.) in pocket attached to back of container.

MacHale:1998:LHLc

[MK98c] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 3, 1998. 1 videocassette (117 min.).

MacHale:1998:LHLd

[MK98d] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 4, 1998. 1 videocassette (106 min.) disk (3 1/2 in.) in pocket attached to back of container.

MacHale:1998:LHLe

[MK98e] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 5, 1998. 1 videocassette (99 min.).

MacHale:1998:LHLf

[MK98f] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 6, 1998. 1 videocassette (91 min.).

Mayer-Kress:1995:HMI

[MKBB95] Gottfried Mayer-Kress, William Bender, and John Bazik. Hyper-media on the Internet as a tool for approaching global problems: a teleconferencing experiment. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN

????  URL http://www.
ncsa.uiuc.edu/SDG/IT94/
Proceedings/CorInfSys/gmk/
Mayer-Kress/HungerConf.

html. Includes CD-ROM.

Morgan:1994:IRE

[MKM94] Keith Morgan and Deborah Kelly-Milburn. Internet resources for economics. College and Research Libraries News, 55(8):475–??, September 1, 1994. ISSN 0099-0086.

Marine:1992:IGSb

[MKNW92] April Marine, ed., Susan Kirkpatrick, Vivian Neou, and Carol Ward. Internet: Getting Started. SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028, September 1992. ISBN 0-944604-15-3. xxii + 360 pp. LCCN TK5105.875.I57 I55 1994. US$39.00. How to join the Internet, and some context so you will know why. This book is not a guide, and not a catalog, but it does contain numerous contact

listings. ftp.nisc.sri.com:
netinfo/Internet-getting-
started-contents.

Marine:1994:IGS

[MKNW94] April Marine, ed., Susan Kirkpatrick, Vivian Neou, and Carol Ward. Internet — getting started. SRI Internet information series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, updated 1993 edition, 1994. ISBN 0-13-289596-X. xxv + 390 pp. LCCN TK5105.875.I57 I55 1994. US$39.00.

Masunaga:1998:WCA

[MKT98] Y. (Yoshifumi) Masunaga, Takuya Katayama, and Michiharu Tsukamoto, editors. Worldwide computing and its applications — WWCA’98: Second International Conference, Tsukuba, Japan, March 4–5, 1998: proceedings, volume 1368 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNCSD9. ISBN 3-540-64216-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.W18 1998.

Martin:1994:TIN

[ML94] James Martin and Joe Leben.

TCP/IP Networking: Architecture, Administration, and Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-642232-2. xv + 333 pp. LCCN TK5105.55 .M37 1994.

Mauldin:1995:LCI

[ML95] Michael L. Mauldin and Lycos, Inc. Lycos –the catalog of the Internet, 1995.

Mlekodaj:1995:BSC

[Mle95] Nancy Lancaster Mlekodaj.

A baseline study of career information on the Internet at state universities and land-grant colleges. Thesis (M.S.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1995. vii + 63 pp.

Mullin:1995:GAO

[MLJ95] Kevin P. Mullin, Annette C.

Lamb, and Larry Johnson.

Gophers, armadillos, and other Internet critters getting started with Internet tools, 1995. 1 videocassette (23 min.).

Malkin:1991:FQA

[MM91] Gary Scott Malkin and April Marine. FYI on questions and answers: answers to commonly asked “new Internet user” questions. Technical Report RFC 1206, United States Working Group, ????, 1991. 26 pp.

Malkin:1992:FQA

[MM92] Gary Malkin and April Marine. FYI on questions and answers: answers to commonly asked “new Internet user” questions. Technical Report RFC 1325, FYI 4, United States Working Group, ????, May 1992. 42 pp.

Maddison:1995:TLD

[MM95] David R. Maddison and Wayne P. Maddison. The tree of life a distributed Internet project containing information about phylogeny and biodiversity, 1995.

Magid:1995:WSB

[MMJ95] Jonathan Magid, R. Douglas Matthews, and Paul Jones. The Web server book: tools and techniques for building your own Internet information site. Ventana Press, Chapel Hill, NC, USA, 1995.

ISBN 1-56604-234-8. xxiv + 640 pp. LCCN TK5105.888 .J66 1995. US$49.95. Includes Linux and Netscape Navigator 1.1.

Mostafa:1997:MAI

[MMLP97] J. Mostafa, S. Mukhopadhyay, W. Lam, and M. Palakal. A multilevel approach to intelligent information filtering: Model, system, and evaluation. ACM Transactions on Information Systems, 15(4): 368–399, October 1997. CO-DEN ATISET. ISSN 1046-8188.

McClure:1993:LIP

[MMR93] Charles R. McClure, William E.

Moen, and Joe Ryan. Libraries and the Internet/ NREN: perspectives, issues, and challenges. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-824-7. xiv + 516 pp. LCCN Z674.82.I59 L48 1993. US$35.00.

McClure:1994:LIN

[MMR94] Charles R. McClure, William E.

Moen, and Joe Ryan. Libraries and the Internet/ NREN: perspectives, issues, and challenges. Mecklermedia, Westport, CT, USA, March 1994. ISBN 0-88736-824-7. xiv + 516 pp. LCCN Z674.82.I59 L48 1994. US$35.00, CDN$43.00. This major study identifies key factors within the library and larger environments that will affect libraries’ involvement in national networking policies. MacKie-Mason:1993:SEI

[MMV93] Jeffrey K. MacKie-Mason and Hal R. Varian. Some economics of the Internet. CREST working paper 93-16, Dept. of Economics, University of Michigan, Ann Arbor, MI, USA, 1993. 29 pp.

MacKie-Mason:1994:EFA

[MMV94] Jeffrey K. MacKie-Mason and Hal R. Varian. Economic FAQs about the Internet. Technical Report 94-95, Dept. of Economics, University of Michigan, Ann Arbor, MI, USA, 1994. 18 pp.

Mostafa:1993:JIH

[MNT93] Javed Mostafa, Tom Newell, and Ritchie Trenthem. JTR’s Internet handbook. J. M., T. N. and R. T., Austin, TX, USA, 1993. 145 pp.

Mostafa:1994:EIH

[MNT94] Javed Mostafa, Thomas Newell, and Richard Trenthem. The Easy Internet Handbook. Hi Willow Research and Publishing, Castle Rock, CO, USA, 1994. ISBN 0-931510-50-3. iv + 140 pp. LCCN TK5105.875.I57 M68 1994. US$20.

Marovac:1992:HTC

[MO92] Nenad Marovac and Larry Osburn. HyperNet: a tool to choreograph worldwide distributed hypermedia documents. Computers and Graphics, 16(2):197– 202, Summer 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mockapetris:1994:ITS

[Moc94] Paul Mockapetris. Internet Technology Series: Domain Name System. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-106865-2.

MODTAP:1993:AM

[MOD93a] MOD-TAP, 285 Ayer Road, PO Box 706, Harvard, MA 01451-0706, USA, Tel: (508) 772-5630. MOD-TAP Applications Manual, 1993.

MODTAP:1993:CC

[MOD93b] MOD-TAP, 285 Ayer Road, PO Box 706, Harvard, MA 01451-0706, USA, Tel:

(508) 772-5630. MOD-TAP Components Catalog, 1993. US$5.00.

Mogul:1988:ELI

[Mog88] Jeffrey C. Mogul. The experimental literature of the Internet: an annotated bibliography. WRL research report 88-3, Digital, Western Research Laboratory, Palo Alto, CA, USA, August 1988. iii + 16 pp.

Mogul:1989:ELI

[Mog89] Jeffrey C. Mogul. The experimental literature of the Internet: An annotated bibliography. Computer Communication Review, 19(1):58–??, January 1, 1989. CODEN CCRED2. ISSN 0146-4833.

Mogul:1995:OSS

[Mog95] Jeffrey C. Mogul. Operating systems support for busy Internet servers. WRL technical note TN-49, Western Research Laboratory, Palo Alto, CA, USA, 1995.

Mogul:1999:SSC

[Mog99]  Jeffrey Mogul.  Speedier

squid: a case study of an internet server performance problem. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/1999-2/squid.html.

Mohn:1995:PIC

[Moh95] Robert C. Mohn. Problems with integrating computer technology into the K– 12 educational curriculum: a study of the use of the Internet and video-conferencing in a fifth grade classroom. Report (M.S.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1995. v + 63 pp.

Mohan:1999:WMI

[Moh99a] C. Mohan. Workflow management in the Internet age. Lecture Notes in Computer Science, 1649:237–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1999:UWS

[Moh99b]  Jim Mohr.  UNIX Web

Server Administrator’s Interactive Workbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-020065-4. xxi + 587 pp. LCCN TK5105.888 .M642 1999. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130200654.html.

Mojnidar:1993:FFI

[Moj93] Jayathirtha Mojnidar. FINFS:

a flexible Internet network file system. Thesis (M.S.), Oklahoma State University, Stillwater, OK, USA, 1993. vi + 117 pp.

Molloy:1993:IRL

[Mol93a] Molly Molloy. Internet resources for Latin American studies. College and Research Libraries News, 54(7):395–??, July 1, 1993. ISSN 0099-0086.

Molta:1993:MIM

[Mol93b] Dave Molta. The Multipurpose Internet Mail Extensions: a tutorial. Network Computing, 4(8):146– ??, August 1, 1993. CODEN NCOMEV. ISSN 1046-4468.

Molva:1999:ISA

[Mol99] Refik Molva. Internet security architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(8): 787–804, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/
                                        31/8/2121.pdf.

Moore:1995:EVC

[Moo95] Dinty W. Moore. The emperor’s virtual clothes: the naked truth about Internet culture. Algonquin Books, Chapel Hill, NC, USA, 1995. ISBN 1-56512-096-5. xvii + 219 pp. LCCN HE7568 .M66 1995.

Moody:1996:IW

[Moo96] Glyn Moody. The Internet with Windows. Butterworth-Heinemann, Oxford, UK, 1996. ISBN 0-7506-2099-4 (UK/Europe), 0-7506-9704-0 (USA). xv + 651 pp. LCCN TK5105.875.I57M65 1996.

Morin:1993:B

[Mor93] Rich Morin. 1993 bookshelf. UNIX review, 11(8):63– ??, August 1, 1993. CODEN UNRED5. ISSN 0742-3136.

Morgan:1994:WGS

[Mor94] Eric Lease Morgan. WAIS and Gopher Servers: a guide for librarians and Internet end-users. Mecklermedia, Westport, CT, USA, 1994. ISBN 0-88736-932-4. ix + 119 pp. LCCN QA76.9.C55 M67 1994. US$30.00.

Mora:1995:BI

[Mor95a] Frederic Mora. La Bible Internet. Addison-Wesley France, Paris, France, 1995. ISBN 2-87908-105-X. xii + 359 pp. LCCN ????

Morphet:1995:WIC

[Mor95b] John P. Morphet. Windows on the Internet: the complete toolchest on CD-ROM. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-912173-X (paperback). xx + 322 pp. LCCN TK5105.875.I57M66 1995.

Morris:1995:HFP

[Mor95c] Mary E. S. Morris. HTML for Fun and Profit. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-359290-1. 264 pp. LCCN ???? US$35.95. Includes CD ROM.

Morrison:1995:OWI

[Mor95d] Deborah Morrison. OS/2 Warp Internet connection: your key to cruising the Internet and the World Wide Web. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-465-4 (paperback). xxii + 233 pp. LCCN TK5105.875.I57M664 1995.

Motley:1994:MUL

[Mot94] Lynne Motley. Modem USA: low cost and free online sources for information, the Internet, databases, and electronic bulletin boards via personal computer and modem in 50 states and Washington, DC. Allium Press, Takoma Park, MD, USA, second revised edition, 1994. ISBN 0-9631233-6-X. 401 pp. LCCN Z674.3 .M68 1994.

Moura:1996:SCD

[Mou96]  Brian Moura.  San car-

los discovers the internet, and shares lessons that were learned. Public management, 78(1):31–37, January 1996. ISSN 0033-3611.

Moy:1998:OAI

[Moy98]  John T. Moy.  OSPF:

Anatomy of an Internet Routing Protocol. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-63472-4. xvii + 339 pp. LCCN TK5105.875.I57M69 1998. US$44.95.

Mazraani:1989:SMC

[MP89] Tony Y. Mazraani and Gurudatta M. Parulkar. Specification of a multipoint congramoriented high performance Internet protocol. Technical report WUCS-89-20, Washington University, Dept. of Computer Science, Saint Louis, MO, USA, 1989. 32 pp.

McLaughlin:1991:BWW

[MP91] Pamela McLaughlin and Andrew (Andrew H.) Perry. Beyond the walls: the world of networked information. Technical report, Syracuse University, Syracuse, NY, USA, 1991. 61 + 146 pp.

Mui:1992:XWS

[MP92] Linda Mui and Eric Pearce.

X Window System Administrator’s Guide for X11 Release 4 and Release 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-83-8 (paperback), 1-56592-052-X (with CD ROM). xxiv + 346 pp. LCCN QA76.76.W56 D44 v.8 1992. US$29.95 (without CD ROM), US$59.95 (with CD ROM). URL http://

www.oreilly.com/catalog/
9780937175835;   http://
www.oreilly.com/catalog/

v8. Contains CD ROM with X11R4 and X11R5 source code, plus compiled versions for Sun 3 SunOS 4.1.1, Sun 4 SunOS 4.1.1, DECstation ULTRIX 4.2, and IBM RS/6000 AIX 3.2.

Malkin:1993:IUG

[MP93] G. (Gary Scott) Malkin and Tracy LaQuey Parker. Internet users’ glossary. Technical Report RFC 1392. FYI 18, United States Area, Internet Engineering Task Force, ????, January 1993. 53 pp.

Mullen:1995:IAK

[MP95] Al Mullen and Michael C. Piper. Internet access in Kansas: a primer and list of providers. Technical report, Kansas Library Network, Topeka, KS, USA, April 18, 1995. 10 pp.

Mack:1998:H

[MP98] E. Stephen Mack and Janan Platt. HTML 4.0. Biblioteca Profesional. Anaya Multimedia, Madrid, Spain, 1998. ISBN 84-415-0464-4. 768 pp. LCCN ????

Malkin:1995:GBI

[MPF95] G. (Gary Scott) Malkin, Tracy LaQuey Parker, and Rafael Fernandez Calvo. Glosario basico ingles-espanol para usuarios de Internet, 1995.

Metcalfe:1996:CI

[MPF96] Robert Metcalfe, Wayne Johnston Pond, and Lilian R. Furst. Casting the Internet, 1996. 1 sound disc.

Miser:1994:GIMb

[MPG94a] Brad Miser, Marta Justak Partington, and Brian Gill. Get on the Internet in 5 minutes!: for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-135-9 (paperback). 88 pp. LCCN TK5105.875.I57 M56 1994. US$9.99. A tiny little book and disk that will get you hooked up with Pipeline, an Internet Service Provider.

Miser:1994:GIMa

[MPG94b] Brad Miser, Marta Justak Partington, and Brian Gill. Get on the Internet in 5 minutes!: for Windows. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1994. ISBN 1-56830-136-7 (paperback). 86 pp. LCCN TK5105.875.I57M564 1994.

Morville:1996:ISH

[MRJ96] Peter Morville, Louis B.

Rosenfeld, and Joseph Janes. The Internet searcher’s handbook: locating information, people and software. Neal-Schuman netguide series Neal-Schuman net-guide series. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-236-8. xi + 236 pp. LCCN TK5105.875.I57 M668 1996.

Maddox:1995:CLT

[MRTR95] Matt Maddox, Scott Reinhart, Jim Taylor, and Jim Rosaschi. Connecting libraries to technology connecting to the Internet, 1995. 1 videocassette (36 min.).

Mumm:1994:ITS

[MS94] James A. Mumm and Ann Sitkin. The Internet in technical services the impact for acquiring resources and providing bibliographic access on technical services, 1994. 1 sound cassette.

Merali:1995:HLI

[MS95] Zinat Merali and Margaret Shaw. Highways for learning: an introduction to the Internet for schools and colleges. NCET, Coventry, England, 1995. ISBN 1-85379-318-3. iv + 99 pp. LCCN ????

Maddox:1996:IO

[MS96] Matthew Maddox and John Stapleton. Internet overview, 1996. 1 videocassette (120 min.).

MacHale:1998:LHLb

[MS98] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 2, 1998. 1 videocassette (117 min.).

Moravec:1992:TRA

[MSB^{+}92] Hans P. Moravec, Bruce Sterling, David Brin, R. Bruce (Robert Bruce) Miller, and Milton T. Wolf. Thinking robots, an aware Internet, and cyberpunk librarians: the 1992 LITA president’s program: presentations by Hans Moravec, Bruce Sterling, and David Brin. LITA president’s series. Library and Information Technology Association, 50 East Huron Street, Chicago, IL 60602, USA, 1992. ISBN 0-8389-7625-5. iv + 200 pp. LCCN Z678.9.T46 1992.

Mills:1990:IAW

[MSD90] David L. Mills, Paul Schragger, and Michael Davis. Internet architecture workshop: Future of the Internet system architecture and TCP/ IP protocols. Computer Communication Review, 20(1):6– ??, January 1, 1990. CODEN CCRED2. ISSN 0146-4833.

Msen:1990:IBP

[Mse90] Msen, Inc. Internet business pages, 1990.

McClure:1999:HEN

[MSK99] Stuart McClure, Joel Scambray, and George Kurtz. Hacking Exposed: Network Security Secrets and Solutions. Osborne/McGraw-Hill, Berkeley, CA, USA, 1999. ISBN 0-07-212127-0. 484 pp. LCCN TK5105.59 .M33 1999. US$39.99.

Mueller:1999:CII

[Mue99a] Milton Mueller. Commentary: ICANN and Internet regulation. Communications of the ACM, 42(6): 41–43, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p41-mueller/.

Mueller:1999:EII

[Mue99b]  Milton Mueller.  Emerg-

ing Internet infrastructures worldwide. Communications of the ACM, 42(6): 28–30, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p28-mueller/.

Muir:1994:BI

[Mui94] James A. Muir. Browsing the Internet. Technical report, Computing Services, Southwest Texas State University, San Marcos, TX, USA, 1994. various pp.

Mui:1995:WYC

[Mui95] Linda Mui. When You Can’t Find Your UNIX System Administrator. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1995. ISBN 1-56592-104-6. 156 pp. LCCN ???? US$17.95.

Mulliner:1992:SAC

[Mul92] Kent Mulliner. Southeast Asia collections and resources on the Internet. COR-MOSEA bulletin, 21(2):1–??, December 1, 1992. ISSN 0734-449X.

Mull:1997:ISP

[Mul97a] Mike Mull. Internet servers in perl. Linux Journal, 37:??, May 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp://ftp.

ssc.com/pub/lj/listings/
issue37/2064.tgz.

Mull:1997:PS

[Mul97b] Mike Mull. Perl and sockets.

Linux Journal, 35:??, March 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mulligan:1999:RSM

[Mul99] Geoff Mulligan. Removing the Spam: E-mail Processing and Filtering. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-37957-0. 176 pp. LCCN TK5105.73.M85 1999. US$19.95.

Murhammer:1998:TIT

[Mur98] Martin Murhammer. TCP/ IP Tutorial and Technical Overview. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-020130-8. 500 pp. LCCN

????  URL http://www.
phptr.com/ptrbooks/ptr_
0130201308.html.

Miller:1992:TRA

[MW92] R. Bruce (Robert Bruce)

Miller and Milton T. Wolf, editors. Thinking robots, an aware Internet, and cyberpunk librarians: a collection of background essays prepared for the 1992 LITA President’s Program, Tools for knowing, environments for growing: visions of the potential of information technology for human development. Library and Information Technology Association, 50 East Huron Street, Chicago, IL 60602, USA, 1992. ISBN 0-8389-7625-5. LCCN Z678.9 .T46 1992.

Maier:1994:SWKa

[MW94a] Gunther Maier and Andreas Wildberger. In 8 Sekunden um die Welt: Kommunikation über das Internet. Addison-Wesley, Reading, MA, USA, 2. aufl. edition, 1994. ISBN 3-89319-701-X. vii + 160 pp. LCCN ????

Maier:1994:SWKb

[MW94b] Gunther Maier and Andreas Wildberger. In 8 Sekunden um die Welt: Kommunikation über das Internet. Addison-Wesley, Reading, MA, USA, 3. überarbeitete aufl. edition, 1994. ISBN 3-89319-775-3. viii + 96 pp. LCCN ????

Maier:1994:ASW

[MW94c] Gunther Maier and Andreas Wildberger. In acht Sekunden um die Welt (Around the world in eight seconds). Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-701-X. 160 pp. DM39.90. It is a quite comprehensive introduction to the Internet, its history and the services available (Mail, News, Gopher, FTP, Telnet, WWW and more). For more information about the book, e-mail one of the au-

thors: wildberg@nestroy.
wu.wien.ac.at (Andreas

Wildberger). Three chapters of this book are available via WWW: URL http:/

/rektorat.wu-wien.ac.at/
stuff/netzbuch.html.

Milles:1994:AII

[MW94d] James Milles and Sally Wambold. Advanced Internet for the initiated, 1994. 3 sound cassettes.

Milles:1994:INI

[MW94e] James Milles and Sally Wambold. Internet for the novice an introduction to Internet access and use, 1994. 3 sound cassettes.

Magavi:1995:DIH

[MWB95] Sunil Magavi, Johnny Wong, and Prakash Bodla. Design and implementation of heterogeneous distributed multimedia system using Mosaic GSQL. Software—Practice and Experience, 25(11): 1223–1241, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Myers:1993:SSS

[MWL93] Judy Myers, Thomas Wilson, and John Lienhard. Surfing the sea of stories: Riding the information revolution. Computers in Libraries, 13(6): 30A–??, June 1, 1993. CO-DEN CPLIE8. ISSN 1041-

Murai:1994:IYI

[MY94] Jun Murai and Shin Yoshimura. Intanetto yogoshu = Internet working. Kyoritsu Shuppan, Tokyo, Japan, shohan. edition, 1994. ISBN 4-320-02716-7. ii + 129 pp. LCCN TK5105.875.I57 I57 1994.

Miller:1993:SIR

[MZ93] Donna Miller and Michael Zeigler. Striking it rich with the Internet: there’s gold in them thar’ networks. Technical report, Lebanon Valley College, Annville, PA, USA, 1993. various pp.

Nagai:1994:IN

[Nag94] Takeshi Nagai. Intanetto nyumon. Fujitsu, Tokyo, Japan, 1994. ISBN 4-938711-17-6. 196 pp. LCCN ????

US-NASA:1990:NSP

[NAS90] NASA. NASA STI program coordinating council, eighth meeting, June 2, 1992: using the Internet. Technical Report NASA technical memorandum NASA-TM-108019 4004232378, U.S. National Aeronautics and Space Administration, Washington, DC, USA, 1990. Distributed to depository libraries in microfiche. Shipping list no.: 93-0354-M. Microfiche. Washington, DC: National Aeronautics and Space.

US-NASA:1992:NS

[NAS92] NASA. Networking for Science. Washington, DC, USA, 1992. 12 pp. Cover title. At head of title: NASA Science Internet. Shipping list no.: 92-0638-P.

US-NASA:1993:GQI

[NAS93] NASA. Central Operation of Resources for Educators. Global quest: the Internet in the classroom, 1993. Videocassette. “Produced by: Imaging Technology Branch Ames Research Center” — Videocassette label. “This tape may be freely copied for the use of education.” Credits: Producer, Eric James; creative consultants, Andrea McCurdy, Marc Siegel; script writer, Rich Yacco; editor, Jay Scheibe. Performers: Host/narrator, Richard Hart. Tells how tapping into the Internet, the world’s largest computer network, will benefit educators and students. Teachers and students describe experiences using the Internet, and educators and others discuss how the Internet fits into the changing classroom.

NASA:1992:NS

[Nat92] National Aeronautics and Space Administration, Washington, DC, USA. Networking for Science, 1992. 12 pp.

NASA:1993:NHI

[Nat93a] National Aeronautics and Space Administration, Washington, DC, USA. NASA Headquarters Internet electronic mail guide: a handbook for electronic mail on NASA Science Internet, October 1993. ii + 29 pp.

NCSA:1993:IGIb

[Nat93b] National Center for Supercomputing Applications. An incomplete guide to the Internet: especially for K–12 teachers and students. Technical report, NCSA, Champaign, IL, USA, December 1993.

NCIS:1994:INN

[Nat94a] National Center for Information Security (United States Dept. of Veterans Affairs). Internet, a network of networks. Technical report, U.S. Dept. of Veterans Affairs, Washington, DC, USA, 1994. 1 videocassette (17 min. + 15 sec.).

NLMUS:1994:NI

[Nat94b] National Library of Medicine (U. S.). NLM and the Internet, 1994. 1 videocassette (4 min.).

NLMUS:1995:IHR

[Nat95] National Network of Libraries of Medicine (U. S.). Internet healthcare resources and information management, 1995. 1 videocassette (9 min. + 30 sec.).

NEI:1996:UIW

[Nat96a] National Institute on the Education of At-risk Students (U. S.). Using the Internet: World Wide Web pages featuring education. Consumer guide; no. 15 Consumer guide (United States. Dept. of Education); no. 15. AR 96-7028, National Institute on the Education of At-risk Students, Washington, DC, USA, 1996. 4 pp.

NSFDCISE:1996:CIP

[Nat96b] National Science Foundation (U. S.). Directorate for Computer and Information Science and Engineering. Connections to the Internet: program announcement. National Science Foundation (U.S.) NSF 96-64. NSF 96-64, National Science Foundation, Arlington, VA, USA, 1996. 11 pp.

Nitzan:1990:ONS

[NC90] Rebecca Nitzan and Ames Research Center. OSI in the NASA science Internet an analysis. NASA contractor report NASA CR-177561, National Aeronautics and Space Administration, Ames Research Center; For sale by the National Technical Information Service, Moffett Field, CA, USA, 1990.

Naythons:1995:IHFa

[NC95a] Matthew Naythons and Anthony Catsimatides. The Internet health, fitness and medicine golden directory. Osborne series. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, special edition, 1995. ISBN 0-07-882195-9 (paperback). xx + 346 pp. LCCN RA776 .N39 1995.

Naythons:1995:IHFb

[NC95b] Matthew Naythons and Anthony Catsimatides. The Internet health, fitness and medicine yellow pages. Osborne IYP series. Osborne/ McGraw-Hill, Berkeley, CA, USA, special edition, 1995. ISBN 0-07-882188-6. xx + 346 pp. LCCN ????

Chang:1994:TSI

[nChCiY94] Min ni Chang, Chien hao Chen, and Chung i. Yeh. Tou shih Internet wang lu. Ho shuo ko chi tsung shu; 34. Ho shuo ko chi wen hua yu hsien kung ssu, Tai-pei shih, Taiwan, chu pan. edition, 1994. ISBN 957-8682-34-4. 4 + 389 pp. LCCN ????

NCSA:1993:IGIa

[NCS93] NCSA Education Group. An incomplete guide to the Internet: and other telecommunications opportunities, especially for teachers and students K–12. Technical report, NCSA, Champaign, IL, USA, 1993. various pp.

Nixon:1998:ILP

[NDM98] Carol Nixon, M. Heide Dengler, and Mare L. McHenry, editors. Internet librarian ’98: Proceedings of the second Internet Librarian Conference, Monterey, CA, November 1–5, 1998. Information Today, Medford, NJ, USA, 1998. ISBN 1-57387-067-6. LCCN Z 674.82 I59 I585 1998.

Neff:1994:GI

[Nef94] Dianah Neff. Government on the Internet, 1994. 1 videocassette (60 mins.).

Nejmeh:1994:IST

[Nej94] Brian A. Nejmeh. Internet: a strategic tool for the software enterprise. Communications of the ACM, 37 (11):23–27, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
188288.html.

Nelson:1995:FGI

[Nel95] Stephen L. Nelson. Field guide to the Internet. Microsoft Press, Bellevue, WA, USA, 1995. ISBN 1-55615-822-X. x + 194 pp. LCCN TK 5105.875 I57 N45 1995.

Nemes:1996:MGM

[Nem96] Gabrielle Nemes. 10 minute guide to Microsoft Exchange. 10 minute guides. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0677-6. ix + 192 pp. LCCN QA76.9.C55N46 1996.

Neou:1994:IC

[Neo94a] Vivian Neou. The Internet CD. SRI international. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-123852-3 (paperback). x + 260 pp. LCCN TK5105.875.I57 N46 1994. US$49.95. System requirements for disc: CD-ROM reader (ISO 9660compatible); modem or Ethernet card; Internet access. At head of title: SRI International.

Neou:1994:IDA

[Neo94b] Vivian Neou. Internet: Domain Administration. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-511180-3.

Neou:1995:IML

[Neo95] Vivian Neou. Internet mailing lists navigator: a guide for Windows users. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-193988-2 (paperback). xi + 466 pp. LCCN TK5105.73.N46 1995.

SRINIC:1983:IPI

[Net83] Network Information Center, SRI International, Menlo Park, CA, USA. Internet protocol implementation guide, 1983. 151 pp.

NetWorld:1986:NW

[Net86] Network World, 1986. ISSN 0887-7661 (print), 1944-7655 (electronic). 161 Worchester Road, Framingham, MA 01701, USA, Tel: (508) 875-6400. Weekly; available free to qualified subscribers.

NetComp:1990:NC

[Net90] Network Computing, 1990.

ISSN 1046-4468. 600 Community Drive, Manhasset NY 11030-9789, USA, Tel: (516) 562-5071. Available free to qualified subscribers.

NetworkElectronics:1993:LAN

[Net93] Network Electronics, 4801 N. Ravenswood Ave., Chicago, IL. 60640-4496, USA, Tel: (312) 784-5100. Network Electronics LAN Catalog, 1993.

NetCruiser:1994:SNG

[Net94a] NetCruiser (Computer program). Surfin’ the net a guide for the ride!, 1994. 1 videocassette.

Anonymous:1994:NGO

[Net94b] NetGuide: the guide to online services and the Internet, page various, 1994. ISSN 1078-4632; 1077-4173. CMP Media, Manhasset, NY, USA.

NCC:1994:NNF

[Net94c] Netscape Communications Corporation, Mountain View, CA, USA. Netscape Navigator: a friendly handbook on the coolest way to explore the Internet, 1994. 159 pp.

Anonymous:1994:NBN

[Net94d] Netwise: business news and support for the Internet data superhighway, page various, 1994. ISSN 1075-2609. Infinity Link Corp., Malvern, PA, USA.

Anonymous:1995:NIM

[Net95a] Net: the Internet magazine, page various, 1995. Future Publishing Ltd., Bath, Avon.

Netcom:1995:NIS

[Net95b] Netcom. Netcruiser Internet software, 1995. 1 computer disk.

NCC:1995:ICN

[Net95c] Netscape Communications Corporation. Internet coach for Netscape Navigator Windows, 1995. 3 computer disks.

Anonymous:1995:NIE

[Net95d] Netware/Internet exchange, page various, 1995. ISSN 1080-3971. IDG Newsletter Corp., Boston, MA, USA.

NetGenesis:1996:BWW

[Net96a] Net. Genesis. Build a world wide Web commerce center: plan, program, and manage Internet commerce for your company. John Wiley, New York, NY, USA, 1996. ISBN 0-471-14928-4 (paperback). various pp. LCCN HD30.37.B84 1996.

Anonymous:1996:NIR

[Net96b] Net.10 international reporter:

the business and legal profile of the Internet, page various, 1996. ISSN 1088-4890. Reldus International, Philadelphia, PA, USA.

Neufeld:1994:DIG

[Neu94]  Gerald Neufeld.  Driving

the Infohighway — A Guide for VAX/VMS Users. New Fields Publishing, Brandon, Manitoba, Canada, 1994. ISBN 0-920436-54-4. vi + 286 pp. US$18.95, CDN$24.95. The book is for users who will be accessing the Internet via a VAX/VMS host. It is designed for a reader who has some familiarity with personal computers but who has had little or no experience in using the VAX. It assumes that the user will be accessing the VAX using a VT-100 or better terminal or terminal emulation program. It explains how to access the VAX using a PC on the campuswide Ethernet LAN or by modem. The book can be used as a tutorial for beginners. The novice user will find the step by step instructions, practical examples, and glossary very helpful. However, the book can also be used as a valuable resource for the more experienced user.

Neugass:1998:SSH

[Neu98a] Jonah Neugass. Step by step hsueh hui HTML chih tso wang yeh. Beginners’; 3. Tien nao jen wen hua shih yeh ku fen yu hsien kung ssu, Taipei shih, tseng ting pan. edition, 1998. ISBN 957-8309-09-0 (paperback). 194 pp. LCCN ????

Neumann:1998:IRI

[Neu98b] Peter G. Neumann. Inside risks: Internet gambling. Communications of the ACM, 41(2):112, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-2/
                                        p112-neumann/.

BBN:1981:HAF

[New81] Bolt Beranek and Inc. Newman. A history of the Arpanet: The first decade. Report 4799, Defense Advanced Projects Agency, 1400 Wilson Boulevard, Arlington, VA 22209, USA, April 1, 1981. URL https://

                                        ipj.dreamhosters.com/wp-
                                        content/uploads/2015/07/
                                        A-History-of-the-ARPANet.
                                        pdf.

Newby:1993:DDI

[New93] Gregory B. Newby. Directory of Directories on the Internet: a Guide to Information Sources. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-768-2. xvii + 153 pp. LCCN TK5105.875.I57 N49 1994. US$29.50. Intended for those who need to identify Internet information resources that point to other resources.

Newby:1994:DDI

[New94a] Gregory B. Newby. Directory of directories on the Internet: a guide to information sources. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-768-2. xvii + 153 pp. LCCN TK5105.875.I57 N49 1994. US$29.50.

Newton:1994:NTD

[New94b] Harry Newton. Newton’s telecom dictionary: the official dictionary of computer telephony, telecommunications, networking, data communications, voice processing and the Internet. Flatiron Publishing, New York, NY, USA, eighth edition, 1994. ISBN 0-936648-60-0. iv + 1171 pp. LCCN TK5102 .N485 1994. Cover title. “November 1994.”.

NJITDCDS:1995:FJI

[New95a] New Jersey Institute of Technology. Division of Career and Development Services. Finding a job on the Internet. Technical report, New Jersey Institute of Technology, Newark, NJ, USA, 1995. 29 + 8 pp.

Newman:1995:PMI

[New95b] Nathan Newman. Prop 13 meets the Internet: how state and local government finances are becoming road kill on the information superhighway: a report. Technical report, Center for Community Economic Research, University of California, Berkeley, Berkeley, CA, USA, August 1995. 14 pp.

Newton:1995:NTD

[New95c] Harry Newton. Newton’s telecom dictionary: the official dictionary of telecommunications, computer telephony, data communications, voice processing, client/server telephony, networking and the Internet. Flatiron Pub. Co., New York, NY, USA, nineth edition, 1995. ISBN 0-936648-75-9. iv + 1288 + 2 pp. LCCN ????

Newton:1996:NTD

[New96] Harry Newton. Newton’s telecom dictionary: the official dictionary of telecommunications, computer telephony, data communications, voice processing, PC telephony, Windows 95 and NT communications, networking and the Internet. Flatiron Pub. Co., New York, NY, USA, tenth edition, 1996. ISBN 0-936648-78-3. v + 1320 + 3 pp. LCCN ????

Nicol:1999:HIH

[NGP^{+}99] John R. Nicol, Yechezkal S.

Gutfreund, Jim Paschetto, Kimberly S. Rush, and Christopher Martin. How the Internet helps build collaborative multimedia applications. Communications of the ACM, 42(1): 79–85, January 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-1/
p79-nicol/.

Nguyen:1995:PIS

[Ngu95] Triet Jason Nguyen. Publishing on the Internet: a study of the World Wide Web, the market and opportunities for printers and publishers. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1995. 55 + 4 pp.

NTU:1995:CI

[NI95] National Technological University and International Communications Corporation. Commerce on the Internet, 1995. 2 videocassettes (100 min.).

Nickerson:1993:VRL

[Nic93] Gord Nickerson. The virtual reference library. Computers in Libraries, 13(5):37– ??, May 1, 1993. CODEN CPLIE8. ISSN 1041-7915.

Nicholson:1995:CIC

[Nic95] Dennis Nicholson. Cataloging the Internet: CATRIONA feasibility study. Library and information research report; 105. British Library Research and Development Department [Distributed by Turpin Distribution Services], London, UK, 1995. ISBN 0-7123-3299-5 (paperback). iv + 87 pp. LCCN ????

Nielsen:1995:MHI

[Nie95] Jakob Nielsen. Multimedia and Hypertext: the Internet and beyond. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-518408-5. xiii + 480 pp. LCCN QA76.76.H94 N543 1995.

Niederst:1998:WDN

[Nie98] Jennifer Niederst. Web Design in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-515-7. xviii + 560 pp. LCCN TK5105.888 .N542

1999.  US$24.95.  URL

                                        http://www.oreilly.com/
                                        catalog/wdnut.

NNK:1995:NI

[Nih95] Nihon Noritsu Kyokai. Nyumon Intanetto. Nihon Noritsu Kyokai Manejimento Senta, Tokyo, Japan, shohan. edition, 1995. ISBN 4-8207-1087-7. 238 pp. LCCN ????

Nilan:1993:SIG

[Nil93] Michael S. Nilan. Speculations on the impact of global electronic networks on human cognition and human organization. Internet research, 3 (3):47–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

NTU:1996:CI

[NIP96] National Technological University, International Communications Corporation, and PBS Adult Learning Satellite Service. Commerce on the Internet, 1996. 1 videocassettes (100 min.).

Noonan:1994:LGL

[NM94] Dana Noonan and Metronet.

LIBCAT: a guide to library resources on the Internet. Technical report, Metropolitan State University, St. Paul, MN, USA, March 1994. various pp.

Novak:1998:IWG

[NM98] Jeannie Novak and Pete Markiewicz. Internet world guide to maintaining and updating dynamic Web sites. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-24273-X. 365 pp. LCCN TK5105.888.N693 1998. US$29.99.

NJI:1994:FIS

[NN94] The federal Internet source:

a directory of nearly 300 sources of government information on the Internet and how to find and use them, page various, 1994. US$18.95. National Journal, Inc. and NetWeek, Inc., Washington, DC, USA.

Noack:1996:HW

[Noa96] David Noack. How’s the weather? Internet World, 7(1):78–??, January 1, 1996. CODEN IERNE8. ISSN 1064-3923.

Nofal:1995:IB

[Nof95a] Buzz Nofal. Internet basics, 1995. 1 videocassette (66 min).

Nofal:1995:SIS

[Nof95b] Buzz Nofal. Setting up an Internet server, 1995. 1 videocassette (95 min).

Nofal:1995:UUF

[Nof95c] Buzz Nofal. Understanding and using FTP, 1995. 1 videocassette (75 min).

Nofal:1995:UUG

[Nof95d] Buzz Nofal. Understanding and using gopher, 1995. 1 videocassette (77 min).

Nofal:1995:UUT

[Nof95e] Buzz Nofal. Understanding and using telnet, 1995. 1 videocassette (65 min.).

Nofal:1995:UUW

[Nof95f] Buzz Nofal. Understanding and using WWW, 1995. 1 videocassette (75 min).

Nofal:1996:CFT

[Nof96a] Buzz Nofal. Creating forms, tables and image maps, 1996. 1 videocassette (94 min).

Nofal:1996:CPW

[Nof96b] Buzz Nofal. Creating pages for the world wide web, 1996. 1 videocassette (85 min).

Nofal:1996:PI

[Nof96c] Buzz Nofal. Publishing on the Internet, 1996. 1 videocassette (88 min).

Noll:1997:TOI

[Nol97] A. Michael Noll. Technical opinion: Internet pricing vs. reality. Communications of the ACM, 40(8): 118–121, August 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-8/
p118-noll/.

Noonan:1991:GIB

[Noo91a] Dana Noonan. A guide to internet/BITNET. A metro library user network guide, Metropolitan State University, St. Paul, MN, USA, October 1991. 55 pp.

Noonan:1991:IL

[Noo91b] Dana Noonan. Internet libraries. Technical report, Metropolitan State University, St. Paul, MN, USA, 1991. various pp.

Noonan:1992:GIBa

[Noo92a] Dana Noonan. A guide to internet/bitnet. Technical report, Metropolitan State University, St. Paul, MN, USA, January 1992. 59 pp.

Noonan:1992:GIBb

[Noo92b] Dana Noonan. A guide to internet/bitnet. Technical report, Metropolitan State University, St. Paul, MN, USA, June 1992. 74 pp.

Noonan:1994:GIB

[Noo94] Dana Noonan. A guide to internet/bitnet. Technical report, Metropolitan State University, St. Paul, MN, USA, March 1994. various pp.

NorthWestNet:1992:NUS

[Nor92] NorthWestNet. NorthWest-Net user services Internet resource guide. Technical report, NorthWestNet, Bellevue, WA, USA, 1992. xi + 279 pp. US$20. NorthWest-Net’s Resource Guide.

Norris:1993:PSC

[Nor93] Frank Norris. The pit: a story of Chicago, 1993.

NI:1995:NDN

[Nor95] Northcoast digital news community and computer news from Northcoast Internet, 1995. ISSN 1086-9115. Northcoast Internet, Eureka, CA, USA.

Notess:1993:NLD

[Not93a] Greg R. Notess. On the nets — LC’s debut on the Internet. Database, 16(5):84–??, October 1, 1993. CODEN DTBSDQ. ISSN 0162-4105.

Notess:1993:NLP

[Not93b] Greg R. Notess. On the nets — the low-end power user: Using the Internet via a dialup account. Online, 17(6):87– ??, November 1, 1993. CO-DEN ONLIDN. ISSN 0146-5422.

Notess:1993:NID

[Not93c] Greg R. Notess. On the nets: Internet database review: UnCover. Database, 16(4):99–??, August 1, 1993. CODEN DTBSDQ. ISSN 0162-4105.

Notess:1994:IAP

[Not94a]  Greg R. Notess.  Inter-

net Access Providers: An International Resource Directory. Mecklermedia, Westport, CT, USA, 1994. ISBN 0-88736-933-2 (paperback). xxviii + 309 pp. LCCN TK5105.875.I57 N68 1994. US$30.00, CDN$37.50, UK£21.00.

Notess:1994:ICU

[Not94b] Greg R. Notess. Internet connections: Understanding your access options. Online, 18(5):41–??, September 1, 1994. CODEN ONLIDN. ISSN 0146-5422.

Notess:1994:N

[Not94c] Greg R. Notess. On the nets. Online, 18(5):110–??, September 1, 1994. CODEN ONLIDN. ISSN 0146-5422.

Nova:1994:CST

[Nov94] Craig Nova. Commentary: Slouching toward the Internet: a world of information can seem just out of reach. BYTE Magazine, 19 (10):282–??, October 1, 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Neou:1996:HCI

[NR96] Vivian Neou and Mimi Recker. HTML CD: an Internet publishing toolkit: for Windows. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-232331-1. xxxviii + 311 pp. LCCN QA76.76.H94N46 1996.

Needham:1994:IMD

[Nram94] Col Needham and rec. arts. movies. The Internet movie database, 1994.

Nedungadi:1994:HDI

[NS94] Kishnan Nedungadi and Michael Schneider. History and development of the Internet routing protocols. Technical report, Dept. of Computer Science, Macalester College, St. Paul, MN, 1994. 185 pp. Honors paper.

NNSC:1989:IRG

[NSF89] NSF Network Service Center. Internet resource guide. Technical report, NSF Network Service Center, Cambridge, MA, 1989.

NNSC:1991:IRG

[NSF91] NSF Network Service Center (NNSC). Internet resource guide. Technical report, Bolt Beranek and Newman, Inc., 10 Moulton Street, Cambridge, MA 02138, USA, 1991. 240 pp. US$15.00. The original Internet Resource Guide.

Available as nnsc.nsf.net:
resource-guide/README.

NNSC:1992:IRG

[NSF92] NSF Network Service Center. Internet resource guide. Technical report, NSF Network Service Center, Cambridge, MA, 1992.

Newby:1996:IWI

[NST96] Gregory B. Newby and Michael L. Sullivan-Trainor. Internet world Internet yellow pages. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-343-7. xliii + 908 pp. LCCN ????

Nugent:1995:IID

[Nug95] Michael Nugent. Ireland on the Internet: the definitive guide. Blackwater Press, Dublin, Ireland, 1995. ISBN 0-86121-647-4. 192 pp. LCCN ????

NYSERNet:1992:NNU

[NYS92] NYSERNet. K–12 Networking Interest Group. NYSER-Net: new user’s guide to useful and unique resources on the Internet. Technical report, NYSERNet, Liverpool, NY, USA, June 1992. 146 pp. Version 2.3, minor revisions.

NYSERNet:1994:PGC

[NYS94] NYSERNet, Inc. Project GAIN: connecting rural public libraries to the Internet, 1994. 1 videocassette (26 min. + 31 sec.).

Oblinger:1992:UI

[Obl92] Diana Oblinger. Understanding the Internet. Technical report, Institute for Academic Technology, Research Triangle Park, NC, USA, 1992. 8 pp. 1 sound cassette (20 min.).

Ochocki:1995:CYL

[Och95] Russel Ochocki. Connecting your Linux box to the Internet. Linux Journal, 9:??, January 1995. CODEN LIJOFX.

ISSN 1075-3583 (print), 1938-3827 (electronic).

Ormes:1995:LIC

[OD95] Sarah Ormes and Lorcan Dempsey. Library and Information Commission public library Internet survey: a report prepared by UKOLN. UKOLN, University of Bath, Bath, UK, 1995. ISBN 0-9516856-3-5 (paperback). 13 pp. LCCN ????

Omidvar:1998:NNP

[OD98] Omid Omidvar and Judith Dayhoff. Neural Networks and Pattern Recognition. Academic Press, New York, NY, USA, 1998. ISBN 0-12-526420-8. xvi + 351 pp. LCCN QA76.87 .O45 1998. US$59.95.

Obraczka:1993:IRD

[ODL93] Katia Obraczka, Peter B.

Danzig, and Shih-Hao Li. Internet resource discovery services. Computer, 26(9):8– ??, September 1, 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Owen:1995:LHS

[ODO95] Trevor Owen, Cheryl Dickie, and Ronald Davis Owston. The learning highway: a student’s guide to using the Internet in high school and college. Key Porter Books, Toronto, ON, Canada, 1995.

ISBN 1-55013-713-1. vi + 263 pp. LCCN ????

Anonymous:1995:UIS

[Of 95] Of Course!ware, Inc., Chicago, IL, USA. Using the Internet: student manual, 1995. various pp.

Ogbuji:1998:LMB

[Ogb98] Uche Ogbuji. Linux means business: Linux for Internet business applications. Linux Journal, 55:42–45, November 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

OHaver:1993:IRM

[O’H93] T. C. O’Haver. Internet resources for mathematics and science education. Technical report, University of Michigan Library, Ann Arbor, MI, USA, 1993. 14 pp.

GDIT:1994:TGI

[oIT94] Georgia. Division of Instructional Technology. A teacher’s guide to the Internet. Technical report, Georgia Dept. of Education, Division of Instructional Technology, Atlanta, GA, USA, 1994. various pp.

OLeary:1997:AIV

[OKP97] Daniel E. O’Leary, Daniel Kuokka, and Robert Plant. Artificial intelligence and virtual organizations. Communications of the ACM, 40(1):

52–59, January 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-1/
p52-o_leary/.

Olson:1995:CIR

[Ols95]  Nancy B. Olson.  Cata-

loging Internet resources: a manual and practical guide. OCLC, Dublin, OH, USA, 1995. ISBN 1-55653-189-3. 35 pp. LCCN Z 695.1 I6 C38 1995. Caption title. Computer printout. Los Angeles, Calif.: University of California, Los Angeles, 1995. 28 cm.

Olsen:1996:CRI

[Ols96] Dan R. Olsen, Jr. Computational resources and the internet. ACM Computing Surveys, 28(4es):142, December 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/surveys/1996-28-
4es/a142-olsen/.

Ogan:1995:OUI

[OM95] Christine L. Ogan and Merrill Morris. Opportunities unexplored: the Internet as mass medium. Technical report, ????, ????, 1995. 29 pp.

Omae:1995:IK

[Oma95] Kenichi Omae. Intanetto kakumei. Purejidentosha, Tokyo, Japan, 1995. ISBN 4-8334-1554-2. 250 pp. LCCN ????

Okerson:1995:WAI

[OO95] Ann Okerson and James O’Donnell. Whose article is it anyway? issues in intellectual property for academe in the 90s, 1995. 1 videocassette (119 min.).

OLeary:1996:I

[OO96] T. J. (Timothy J.) O’Leary and Linda I. O’Leary. Internet. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-049093-7. 144 pp. LCCN ????

Oppliger:1995:ISE

[Opp95] R. Oppliger. Internet security enters the Middle Ages. Computer, 28(10):100–101, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:1996:ASS

[Opp96] Rolf Oppliger. Authentication systems for secure networks. The Artech House computer science library. Artech House Inc., Boston, MA, USA, 1996. ISBN 0-89006-510-1. xvii + 186 pp. LCCN TK5105.59 .O77 1996. URL

                                        http://www.artechhouse.
                                        com/Detail.aspx?strIsbn=
                                        978-0-89006-510-5.

Oppliger:1997:ISF

[Opp97] Rolf Oppliger. Internet security: Firewalls and beyond. Communications of the ACM, 40(5):92–102, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
1997-40-5/p92-oppliger/.

Oppliger:1998:IIS

[Opp98]  Rolf Oppliger.  Inter-

net and Intranet security. The Artech House computer science library. Artech House Inc., Boston, MA, USA, 1998. ISBN 0-89006-829-1. xxi + 348 pp. LCCN TK5105.875.I57 O67

1998. URL http://www.
artechhouse.com/Detail.
aspx?strIsbn=978-0-89006-
829-8.

OReilly:1996:PMI

[O’R96] Tim O’Reilly. Publishing models for Internet commerce. Communications of the ACM, 39(6):79–86, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228522.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228522.html.

Steele:1994:UIG

[oS94] Jerri org Steele. Using Internet for genealogy, 1994. 1 sound cassette W-32; Repeat Performance.

Osheroff:1993:CVO

[Osh93] Shiela Keil Osheroff. Communications vendors and organizations providing Internet access. Serials review, 19(4):43– ??, 1993. ISSN 0098-7913.

OReilly:1988:MUU

[OT88] Tim O’Reilly and Grace Todino. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1988. ISBN 0-937175-09-9. xvi + 256 pp. LCCN QA76.8.U65 O64 1988.

Otis:1995:UGI

[Oti95] Yves Otis. User’s guide to Internet services for historians. Technical report, Canadian Historical Association, Ottawa, ON, Canada, 1995. 25 pp.

Otte:1994:ISB

[Ott94] Peter Otte. The information superhighway: beyond the Internet. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 1-56529-825-X (paperback). xiii + 241 pp. LCCN TK5105.87.O87 1994. US$19.99. “Written for anyone curious about the network of the future this book is a conceptual overview of the future of computer communications.”.

Overoye:1994:VGI

[Ove94] Dave Overoye. The video guide to the Internet, 1994. 1 videocassette (ca. 45 min.).

Owens:1993:IAA

[Owe93] Stephen C. Owens. Internet access at Frostburg State University: a hands-on user’s manual for those with network accounts on the FSU Digital VAX. Frostburg State University (FSU), Computer Science Dept., Frostburg, MD, USA, 1993. 45 (or 55??) pp.

PBFV:1994:GIE

[PA94] Phoenix/BFA Films and Video and Aames Teleproductions. Getting on the Internet easy, 1994. 1 videocassette (35 min.).

Parker:1995:BAI

[PA95] Andrew Parker and Ray Attwood. Business applications of the Internet: commercial opportunity or media hype? Financial times management reports. Financial Times Telecoms and Media Publishing, London, UK, 1995. ISBN 1-85334-346-3 (spiral). x + 206 pp. LCCN ????

Pabrai:1996:UI

[Pab96] Uday O. Pabrai. UNIX internetworking. The Artech House telecommunications library. Artech House Inc., Boston, MA, USA, second edition, 1996. ISBN 0-89006-778-3. xix + 454 pp. LCCN TK5105.5 .P32 1996.

Padlipsky:1985:ENS

[Pad85] M. A. Padlipsky. The Elements of Networking Style and Other Essays and Animadversions on the Art of Intercomputer Networking. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1985. ISBN 0-13-268129-3 (hardcover), 0-13-268111-0 (paperback). 236 pp. LCCN TK5105.5 .P34 1985. US$34.95, CDN$44.95. Mike was the Multics systemprogrammer at Project Mac and is personally responsible for putting two systems on the ARPAnet back when that meant hacking the OS designed and implemented a number of protocols including being working-group chair, and has been a commentator on Internet RFCs since the beginning. (He was also Dennis Ritchie’s officemate before Bell Labs pulled out of Multics and Dennis had to invent UNIX, and has a personal exemption from Dijkstra to use a GOTO when, in his professional opinion as a protocol implementor, it is necessary.) This is a book on style in protocol design and the great Bromdinagian TCP vs. ISO debates. It’s a nice piece of history and explains where TCP, TELNET, FTP came from.

Padovano:1993:NAU

[Pad93] Michael Padovano. Networking Applications on UNIX System V. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-613555-2. xvi + 544 pp. LCCN QA76.76.O63 P33 1993. US$44.

Padowicz:1995:IES

[Pad95] Julian Padowicz. Internet explained, short and sweet, 1995. ISBN 1-881288-10-2. 1 sound cassette.

Panturu:1995:UIU

[Pan95] Daniel Ion Panturu. Use of Internet by U.S. radio stations. Thesis (M.A.), Texas Tech University, Lubbock, TX, USA, 1995. vi + 99 pp.

Pantuso:1996:CIG

[Pan96]  Joe Pantuso.  The com-

plete Internet gamer. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-13787-1 (paperback). various pp. LCCN GV1469.15.C62 1996.

Paoli:1996:EWT

[Pao96] Jean Paoli. Extending the Web’s tag set using SGML: authoring new tags with Grif Symposia. Computer Networks and ISDN Systems, 28(7–11):1095–1104, May 1, 1996. CODEN CNISE9.

ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1625.

Parkhurst:1990:LPN

[Par90] Carol A. Parkhurst, ed. Library Perspectives on NREN: The National Research and Education Network. Library and Information Technology Association, 50 East Huron Street, Chicago, IL 60602, USA, 1990. ISBN 0-8389-7477-5. xi + 75 pp. LCCN Z674.7 .L52 1990. US$10.50.

Parker:1993:IGU

[Par93] Richard Parker. Interactive 3-D graphics using an Internet client and server. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1993. various pp.

Parhamovich:1994:FGI

[Par94a] Maggie Parhamovich. Federal government information on the Internet. Technical report, ????, ????, July 1994. 79 pp.

Parker:1994:TYT

[Par94b] Timothy Parker. Teach Yourself TCP/IP in 14 Days. Howard W. Sams, Indianapolis, IN 46268, USA, 1994. ISBN 0-672-30549-6. xxiv + 438 pp. LCCN TK5105.55 .P37 1994. US$29.99, CDN$39.99.

Partridge:1994:GN

[Par94c]  Craig Partridge.  Gigabit

Networking. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-56333-9. xiv + 396 pp. LCCN TK5105.5 .P3745 1994. US$46.25.

Parks:1995:ISG

[Par95] Dayna Parks. Internet survival guide: a beginner’s guide to mastering shareware for SLIP and PPP connections for Windows. MCp2s Cyberspace Research, Columbus, OH, USA, 1995. ISBN 0-9646746-0-2. 218 pp. LCCN TK5105.875.I57P37 1995.

Pareras:1996:IM

[Par96a] Luis G. Pareras. Internet y medicina. Guia de referencia. Masson, Masson, France, 1996. ISBN 84-458-0450-2. 709 pp. LCCN ????

ParraPerez:1996:IE

[Par96b] Beatriz Parra Perez. Internet en Espana. Informatica personal-profesional. Anaya Multimedia, Madrid, Spain, 1996. ISBN 84-7614-850-X. 252 pp. LCCN ????

Pascuzzi:1995:CGA

[Pas95]  Giovanni Pascuzzi.  Cy-

berdiritto: guida alle banche dati italiane e straniere, alla rete Internet e all’apprendimento assistito da calcolatore. Zanichelli, Bologna, Italy, 1995. ISBN 88-08-09106-6. 245 pp. LCCN KKH47.5.P37 1995.

Pasquale:1996:TIC

[Pas96] Joseph Pasquale. Towards Internet computing. ACM Computing Surveys, 28(4es): 215, December 1996. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/surveys/1996-28-
4es/a215-pasquale/.

Paulson:1999:IAIa

[Pau99a] L. C. Paulson. Inductive analysis of the Internet Protocol TLS. Lecture Notes in Computer Science, 1550:1– 12, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1999:IAIb

[Pau99b] Lawrence C. Paulson. Inductive analysis of the Internet protocol TLS. ACM Transactions on Information and System Security, 2(3): 332–351, August 1999. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tissec/1999-2-
3/p332-paulson/.

Paxson:1997:ERB

[Pax97] Vern Paxson. End-to-end routing behavior in the Internet. IEEE/ACM Transactions on Networking, 5 (5):601–615, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1997-5-5/p601-paxson/.

Paxson:1999:EEI

[Pax99]  Vern Paxson.  End-to-

end Internet packet dynamics. IEEE/ACM Transactions on Networking, 7(3): 277–292, June 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-3/p277-paxson/.

Pike:1995:UIY

[PB95] Mary Ann Pike and Scott Berkun. Using the Internet with your Mac. Que Corporation, Indianapolis, IN, USA, special edition, 1995. ISBN 0-7897-0212-6. xli + 927 pp. LCCN TK5105.875.I57P58 1995.

Press:1998:IDF

[PBF^{+}98] Larry Press, Grey Burkhart, Will Foster, Sy Goodman, Peter Wolcott, and Jon Woodard. An Internet diffusion framework. Communications of the ACM, 41(10): 21–26, October 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-10/
                                        p21-press/.

Piscitello:1993:OSN

[PC93] David M. Piscitello and A. Lyman Chapin. Open Systems Networking: TCP/ IP and OSI. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-56334-7. xxii + 600 pp. LCCN TK5105.55.P57 1993. US$49.50.

Polly:1995:I

[PC95] Jean Armour Polly and Steve Cisler. Internet@LJ. Library journal, 120(11):26–??, June 15, 1995. CODEN LIBJA7. ISSN 0363-0277.

Pausch:1992:LLS

[PCD92] Randy Pausch, Matthew Conway, and Robert De-Line. Lessons learned from SUIT, the Simple User Interface Toolkit. ACM Transactions on Information Systems, 10(4):320–344, October 1992. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Pethia:1991:GSO

[PCF91] R. Pethia, S. Crocker, and B. Fraser. Guidelines for the secure operation of the Internet, 1991. 1 data file 5 1/4 in.

Place:1996:EIY

[PDP96] Ron Place, Klaus Dimmler, and Thomas (Thomas R.)

Powell. Educator’s Internet yellow pages. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-232356-7. xiv + 386 pp. LCCN LB1044.87 .P53 1996. “A Web surfer book.”–Cover.

PG:1994:BDG

[PE94] Project Gutenberg and Electronic Frontier Foundation. Big dummy’s guide to the Internet, 1994.

Pearce:1995:ASC

[PE95] J. M. Pearce and A. Ellis, editors. Australian Society for Computers in Learning in Tertiary Education: AS-CILITE 95: Annual conference; 12th — 1995 Dec, ASCILITE -CONFERENCE PROCEEDINGS- 1995; 12th. The University of Melbourne, Faculty of Science, Melbourne, Victoria, Australia, 1995. ISBN 0-7325-1220-4. LCCN ????

Palmer:1999:DNE

[PE99] Jonathan W. Palmer and Lars Bo Eriksen. Digital newspapers explore marketing on the Internet. Communications of the ACM, 42(9): 32–40, September 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-9/
p32-palmer/.

Peal:1994:AI

[Pea94] David Peal. Access the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 0-7821-1529-2. xxv + 241 pp. LCCN TK5105.875.I57P42 1994. US$19.99. Includes diskette.

Peal:1995:AI

[Pea95] David Peal. Access the Internet! Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1995. ISBN 0-7821-1744-9. xxvii + 349 pp. LCCN TK5105.883.N43P43 1995.

Pearl:1998:PRI

[Pea98] Judea Pearl. Probabilistic Reasoning in Intelligent Systems: Networks of Plausible Inference. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-479-0. xix + 552 pp. LCCN Q335. .P383 1988. US$69.95.

PCnovice:1996:PNG

[Pee96a] Peed Corp., Lincoln, NE, USA. PC novice guide to the Internet, 1996. 191 pp.

Peete:1996:BRI

[Pee96b] Gary R. Peete. Business resources on the Internet PLUS: a hands-on workshop. Internet workshop series; no.

3. Library Solutions Press, Berkeley, CA, USA, 2nd printing, January 1996. edition, 1996. ISBN 1-882208-11-0. various pp. LCCN ????

Pfaffenberger:1995:HHL

[PEH95] Bryan Pfaffenberger, Eliezer Erez, and Shimrit Hilel. ha-Leksikon he-hadash lamahshev ha-ishi. Opus, R.G. z.o. Ramat Gan, Israel, mahadurah 1 be-ivrit. edition, 1995. ISBN ???? 326 pp. LCCN ????

Peha:1999:LHI

[Peh99]  Jon A. Peha.  Lessons

from Haiti’s Internet development. Communications of the ACM, 42(6): 67–72, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p67-peha/.

Peliks:1995:WWC

[Pel95a] Gerard Peliks. Le World-Wide Web: Creation de serveurs sur Internet. Addison-Wesley France, Paris, France, 1995. ISBN 2-87908-102-5. xii + 226 pp. LCCN ????

Pellegrini:1995:VED

[Pel95b] Steven Pellegrini. Virtual economics: the development of electronic commerce on the Internet. Thesis (B.A. Honors-Economics), Bucknell University, Lewisburg, PA, USA, 1995. ix + 70 pp. Supervised by: Peter K. Kresl.

Perry:1991:NUGa

[Per91a] Andrew Perry. New user’s guide to useful and unique resources on the Internet. Technical report, NYSER-Net, Syracuse, NY, USA, May 23, 1991. 71 pp.

Perry:1991:NUGb

[Per91b] Andrew Perry. New user’s guide to useful and unique resources on the Internet. Technical report, NYSER-Net, Syracuse, NY, USA, 1991. 146 pp.

Perry:1991:NNUb

[Per91c] Andrew Perry. NYSERNet: new user’s guide to useful and unique resources on the Internet. Technical report, NYSERNet, Liverpool, NY, USA, 1991. 146 pp.

Perlman:1992:IBR

[Per92a] Radia Perlman. Interconnections: Bridges and Routers. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56332-0. x + 389 pp. LCCN TK5105.5 .P474 199.

Perry:1992:NUGa

[Per92b] Andrew Perry. New user’s guide to useful and unique resources on the Internet. Technical report, NYSER-Net, Liverpool, NY, USA, 1992. 146 pp. Cover title.

At head of title: NYSER-Net. “A Project of the NY-SERNet K-12 Networking Interest Group and the NY-SERNet/NYS Library Networking Interest Group for Libraries.” Compiled by Andrew Perry. Contents: Library catalogs and campus information systems — Databases — Electronic discussion groups/forums — Directories — Information resources — FTP archives — Fee-based information services — Software/freeware — Bulletin board services — Miscellaneous.

Perry:1992:NUGb

[Per92c] Andrew Perry. New user’s guide to useful and unique resources on the Internet. In beyond the walls: the world of networked information / Pamela W. McLaughlin … [et al.]. — Syracuse, NY: Syracuse University, 1991 [i.e. 1992], NYSERNet, Liverpool, NY, USA, April 1992. 146 pp.

Perry:1992:NNU

[Per92d] Andrew Perry. NYSERNet:

new user’s guide to useful and unique resources on the Internet. Technical report, NYSERNet, Liverpool, NY, USA, June 1992. 146 pp.

Perry:1992:NSN

[Per92e] Andrew Perry. NYSERNet (SM): new user’s guide to useful and unique resources on the internet, version 2.2.

Technical report, NYSER-Net, Liverpool, NY, USA, January 1992. 158 pp.

Persson:1993:NRG

[Per93] Eric Persson. NetPower: Resource Guide to Online Computer Services. Fox Chapel Publishing, Lancaster, PA, USA, 1993. ISBN 1-56523-031-0. xiii + 774 pp. LCCN TK5105.875.I57 P47 1993. US$39.95. The most exciting section of this guide is devoted to the Internet. Netpower includes a primer and tutorial on using the network, information on getting started and navigating with Internet tools, and hundreds of Internet accessible resources with contact information and descriptions. The guide will point you in the direction of millions of megabytes of information all yours free for the downloading around the Internet.

Perlman:1995:PIH

[Per95a] Brett A. Perlman. Pricing the Internet: how to pay the toll for the electronic superhighway. CSIA discussion paper 95-01, Center for Science and International Affairs, John F. Kennedy School of Government, Harvard University, Cambridge, MA, USA, 1995. 29 pp.

Perry:1995:EMG

[Per95b] Bruce W. Perry. Environmental manager’s guide to the Internet and information superhighway. Business and Legal Reports, Madison, CT, USA, 1995. ISBN 1-55645-030-3. various pp. LCCN ????

Perkins:1998:MID

[Per98] Charles E. Perkins. Mobile IP: Design Principles and Practices. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-63469-4. xx + 275 pp. LCCN QA76.P425 1998. US$48.95.

Peterson:1986:CAI

[Pet86a] Larry L. Peterson. Clusterbased addressing in an Internet. Technical report TR 86-24, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, 1986. 18 pp.

Peterson:1986:NUH

[Pet86b] Larry L. Peterson. Naming users in a heterogeneous Internet: framework for a new approach. Technical report TR 85-28a, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, 1986. 20 pp.

Peterson:1987:PSN

[Pet87] Larry L. Peterson. Profile:

a system for naming Internet resources. Technical report TR 87-7a, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, June 10, 1987. 20 pp.

Peterson:1988:PNS

[Pet88] Larry L. Peterson. The Profile naming service. ACM Transactions on Computer Systems, 6(4):341–364, November 1988. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tocs/1988-6-4/
                                        p341-peterson/.

Peterson:1989:XPA

[Pet89] Larry L. Peterson. The xkernel: a platform for accessing Internet resources. Technical report TR 89-23, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, October 12, 1989. 19 pp.

Peters:1993:WGW

[Pet93a] Paul Evan Peters. Are we gardeners or are we farmers? guiding Internet information resource and service planners and developers. Educom review, 28(5):60–??, September 1, 1993. CODEN EDREEW. ISSN 1045-9146.

Peterson:1993:LIP

[Pet93b] Mary Peterson. Life on the Internet: Portrait of a collaboration. SIGUCCS newsletter, 23(3):3–??, November 1, 1993. CODEN SSNRD3. ISSN 0736-6892.

Petrucco:1995:IGP

[Pet95] Corrado Petrucco. Internet:

guida per i comuni mortali. Scenari. Il cardo, Venezia, Italy, 1995. ISBN 88-8079-044-7. 367 pp. LCCN ????

Pfaffenberger:1994:IPE

[Pfa94a] Bryan Pfaffenberger. Internet in Plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1994. ISBN 1-55828-385-4. xii + 463 pp. LCCN TK 5105.875 I57 P48 1994. US$19.95, CDN$25.95. System requirements for computer disk (NetManage Chameleon sampler. Version 3.11): PC; Microsoft Windows; floppy disk drive.

Pfaffenberger:1994:MUG

[Pfa94b] Bryan Pfaffenberger. Mosaic User’s Guide. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1994. ISBN 1-55828-409-5. x + 274 pp. LCCN TK5105.875.I57 P483 1994. US$24.95.

Pfaffenberger:1995:MWI

[Pfa95a] Bryan Pfaffenberger. Microsoft Windows 95 Internet kit. Microsoft Press, Bellevue, WA, USA, 1995. ISBN 1-55615-885-8, 1-55615-818-1. xix + 219 pp. LCCN QA76.76.O63 P5387 1995.

Pfaffenberger:1995:NNSa

[Pfa95b] Bryan Pfaffenberger. Netscape Navigator: surfing the Web and exploring the Internet. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-553132-X (paperback), 0-12-553133-8 (CD-ROM). xx + 362 pp. LCCN TK5105.882.P43 1995.

Pfaffenberger:1995:NNSb

[Pfa95c] Bryan Pfaffenberger. Netscape Navigator surfing the Web and exploring the Internet. AP Professional, Boston, MA, USA, 1995. ISBN ???? xx + 362 pp. LCCN ????

Pfaffenberger:1995:NNSc

[Pfa95d] Bryan Pfaffenberger. Netscape navigator: surfing the Web and exploring the Internet: Macintosh Version. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-553130-3, 0-12-553131-1 (CD-ROM). xx + 314 pp. LCCN TK 5105.882 P432 1995.

Pfaffenberger:1995:NNSd

[Pfa95e] Bryan Pfaffenberger. Netscape Navigator: surfing the Web and exploring the Internet: Windows Version. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-553132-X, 0-12-553133-8 (CD-ROM). xx + 362 pp. LCCN TK5105.882.P43 1995.

Pfaffenberger:1995:UBF

[Pfa95f] Bryan Pfaffenberger. The Usenet Book: Finding, Using, and Surviving Newsgroups on the Internet. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40978-X (paperback). 468 pp. LCCN TK 5105.875 I57 P485 1995. US$26.95, CDN$34.95.

Pfaffenberger:1995:WWW

[Pfa95g] Bryan Pfaffenberger. World Wide Web Bible. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, February 1995. ISBN 1-55828-410-9. 45 pp. LCCN TK5105.888 .P43 1995. US$27.95.

Pfaffenberger:1996:IPE

[Pfa96a] Bryan Pfaffenberger. Internet in plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1996. ISBN 1-55828-439-7. xxiv + 519 pp. LCCN TK5105.875.I57P48 1996.

Pfaffenberger:1996:MIS

[Pfa96b] Bryan Pfaffenberger. Microsoft Internet starter kit 96. Microsoft Press, Bellevue, WA, USA, 1996. ISBN 1-57251-309-9 (??invalid checksum??). various pp. LCCN TK5105.875.I57 P482 1996.

Pfaffenberger:1996:NNS

[Pfa96c] Bryan Pfaffenberger. Netscape Navigator 2.0: surfing the Web and exploring the Internet. AP Professional, Boston, MA, USA, Macintosh version edition, 1996. ISBN 0-12-553145-1, 0-12-553146-X (CD-ROM), 0-12-553147-8. xxiii + 444 pp. LCCN TK5105.882.P42 1996.

Powell:1993:II

[PG93] Thomas Powell and Barbara Glennan. Introduction to Internet, 1993. 1 videocassette.

Prall:1994:BDG

[PG94] Adam Prall and Adam Gaffin. Big dummy’s guide to the Internet. Technical report, The Electronic Frontier Foundation, Washington, DC, USA, 1994. v + 195 pp.

Pedersen:1995:PI

[PG95] Ted Pedersen and Mel Gilden. Pirates on the Internet. Cybersurfers; #1. Price Stern Sloan, Los Angeles, CA, USA, 1995. ISBN 0-8431-3933-1. 144 pp. LCCN PZ7.P3426Pi 1995.

Pabrai:1996:ITI

[PG96] Uday I. Pabrai and Vijay K. Gurbani. Internet and TCP/ IP network security: securing protocols and applications. J. Ranade workstation series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-048215-2 (paperback). various pp. LCCN TK5105.59.P33 1996.

Polly:1993:EI

[PGS93] Jean Armour Polly, Rick Gates, and Richard J. Smith. Exploring Internet, 1993. 1 videocassette (ca. 135 min.).

Polly:1994:EI

[PGS94] Jean Armour Polly, Rick Gates, and Richard J. Smith. Exploring Internet, 1994. 1 videocassette (VHS) (60 min.).

Pieper:1997:BDT

[PH97] Michael Pieper and Dirk Hermsdorf. BSCW for disabled teleworkers: usability evaluation and interface adaptation of an Internetbased cooperation environment. Computer Networks and ISDN Systems, 29(8–13): 1479–1487, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=08-13&aid=1689.

Anonymous:1995:PBI

[Phi95] Phillips Business Information’s Internet week, page various, 1995. ISSN 1081-2474. Phillips Business Information, Inc., Potomac, MD, USA.

Phinney:1993:GCU

[PHW^{+}93] Marianne Phinney, Henry Hartman, Terry Walch, Daniel Matta, Nancy Palmer, and Mario Ramos. The global connection using the Internet in the classroom, starlink teleconference, 1993. 1 videocassette (80 min.).

Pierret:1995:ICR

[Pie95] James A. Pierret, Jr. Internet curriculum and resource guide for middle school geography teachers. Technical report, ????, ????, 1995. 72 pp.

Pigot:1995:CVP

[Pig95] Thierry Pigot. Créez vos pages Web sur Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1614-3. xix + 207 pp. LCCN ????

Pike:1994:UM

[Pik94]  Mary Anne Pike.  Us-

ing Mosaic. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 0-7897-0021-2. xix + 390 pp. LCCN TK5105.875.I57 U85 1994. US$24.95.

Pike:1995:UI

[Pik95] Mary Ann Pike. Using the Internet. Que Corporation, Indianapolis, IN, USA, second edition, 1995. ISBN 0-7897-0077-8. xxxviii + 1241 pp. LCCN ???? Includes CD-ROM.

Pike:1996:UIW

[Pik96]  Mary Ann Pike.  Using

the Internet with Windows 95. Que Corporation, Indianapolis, IN, USA, special edition, 1996. ISBN 0-7897-0646-6 (paperback). xxxix + 1296 pp. LCCN TK5105.875.I57P579 1996.

Pillay:1990:IDD

[Pil90] Harish Pillay. An implementation of the Department of Defense’s transmission control protocol Internet protocol (TCP/IP) for the Microsoft OS/2 operating system. Thesis (M.S.), Oregon State University, Corvallis, OR, USA, 1990. 49 pp.

Pillman:1991:NRM

[Pil91] Hope Pillman. Networking I the road map of the Internet, 1991. 1 sound cassette (90 min.).

Pitter:1995:ESGa

[Pit95a] Keiko M. Pitter. Every student’s guide to the Internet. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-051773-8. viii + 183 pp. LCCN TK5105.875.I57 E96 1994.

Pitter:1995:ESGb

[Pit95b] Keiko M. Pitter. Every student’s guide to the Internet: Macintosh version. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-052109-3. xi + 196 pp. LCCN ????

Pitter:1995:ESGc

[Pit95c] Keiko M. Pitter. Every student’s guide to the Internet:

Windows version. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-052107-7 (paperback). xi + 196 pp. LCCN ????

Pitkow:1995:TIP

[PJ95] James E. Pitkow and R. K. Jones. Towards an intelligent publishing environment. Computer Networks and ISDN Systems, 27(6): 729–737, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1995/
27/6/1454.pdf.

Pentakalos:1994:ISI

[PK94a] Themos Pentakalos and Scott Kecken. Information superhighway or the Internet, 1994. 1 videocassette (VHS) (ca. 80 min.) quick reference booklet.

Pentakalos:1994:II

[PK94b] Themos Pentakalos and Scott Kecken. Introduction to the Internet, 1994. 1 videocassette (VHS) (90 min.). Title on reference booklet and in advertising: Internet, the video tape. Credits: Producer, Scott Kecken; writer/presenter, Themos Pentakalos. An introduction to the characteristics and capabilities of the Internet. Demonstrates the meaning and use of E-mail, lists, newsgroups, telnet, ftp, archie, gophers, veronica, WAIS, etc.

Petrazzini:1999:IDC

[PK99] Ben Petrazzini and Mugo Kibati. The Internet in developing countries. Communications of the ACM, 42 (6):31–36, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p31-petrazzini/.

Pu:1990:MMA

[PKL90] Calton Pu, Frederick Korz, and Robert C. Lehman. A methodology for measuring applications over the Internet. Technical report CUCS-044-90, Dept. of Computer Science, Columbia University, New York, NY, USA, September 28, 1990. i + 21 pp.

Place:1994:ERI

[Pla94a] Ron Place. Educational resources on the Internet: a book by a teacher for teachers. Internet Express, Colorado Springs, CO (1155 Kelly Johnson Blvd., Suite 400, Colorado Springs, CO 80920), 1994. ISBN ???? 454 pp. LCCN ????

Plattner:1994:SPA

[Pla94b] B. (Bernhard) Plattner, editor. Selected papers of the Annual Conference of the Internet Society/5th Joint European Networking Conference: June 3–17, 1994, Prague, Czech Republic, Computer networks and ISDN systems, v. 27, no. 3 (Dec. 1994) 0169-7552. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1994. ISBN ???? LCCN ????

Platon:1995:LPI

[Pla95] Thierry Platon. L’indispensable pour Internet, autres réseaux et BBS. Marabout informatique; 1098. Marabout, Alleur, Belgique, 1995. ISBN 2-501-02245-9. 571 pp. LCCN ????

ProjectGAIN:1994:CRP

[PM94] Project GAIN and Charles R. McClure. Connecting rural public libraries to the Internet: Project Gain, Global Access Information Network: project evaluation report prepared for NYSERNet, Inc. Technical report, Information Management Consultant Services, Manlius, NY, USA, 1994. various pp.

Peck:1995:W

[PM95a] Susan Peck and Linda Mui. WebSite. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1995. ISBN 1-56592-143-7. 342 pp. LCCN ???? US$499. Includes two diskettes.

Pedersen:1995:IKB

[PM95b] Ted Pedersen and Francis Moss. Internet for kids: a beginner’s guide to surfing the Net. Price Stern Sloan, New York, NY, USA, 1995. ISBN 0-8431-3957-9. 206 pp. LCCN TK5105.875.I57P43 1995.

Polly:1992:NIY

[PMB^{+}92] Jean Armour Polly, George Machovec, Dennis Brunning, Dennis J. Reynolds, George Brett, and Paul Evan Peters. Navigating the Internet: And you thought Columbus had it rough, 1992. 1 sound cassette.

Perkins:1994:IMH

[PMJ94] Charles Perkins, Andrew Myles, and David B. Johnson. IMHP: a mobile host protocol for the Internet. Computer Networks and ISDN Systems, 27 (3):479–491, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1994&volume=27&
issue=3&aid=1368.

Pitts-Moultis:1998:HSS

[PMTJ98] Natanya Pitts-Moultis, Ed Tittel, and Stephen N. James. HTML style sheets design guide. Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-211-4 (paperback). xv + 232 pp. LCCN QA76.76.H94P59 1998.

Perry:1991:NNUa

[PN91] Andrew Perry and NYSER-Net. K–12 Networking Interest Group. NYSERNet: new user’s guide to useful and unique resources on the Internet. Technical report, NYSERNet, Liverpool, NY, USA, October 1991. 146 pp.

Prater:1995:LLA

[PN95] Scott Prater and Tom Newell. LAIR legal automation and Internet review, 1995.

Podnar:1995:IEM

[Pod95] Thomas G. Podnar. Internet electronic mail. Thesis (M.S.), Dept. of Computer and Science, West Chester University, West Chester, PA, USA, 1995. 59 pp.

Polly:1992:SII

[Pol92] Jean Armour Polly. Surfing the INTERNET: an introduction. Technical report, The Author, ????, December 13, 1992. 26 pp.

Polly:1996:IKY

[Pol96] Jean Armour Polly. The Internet kids yellow pages. Osborne/McGraw-Hill, Berkeley, CA, USA, special edition, 1996. ISBN 0-07-882197-5. xxvii + 355 pp. LCCN ????

Pope:1994:IU

[Pop94] Ivan Pope. Internet UK.

Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-190950-9. ???? pp. LCCN TK5105.875.I57 P66 1994.

Postel:1980:DSI

[Pos80] Jon Postel. DOD standard INTERNET protocol. Technical report, University of Southern California, Information Sciences Institute, Marina del Rey, CA, USA, 1980. iii + 42 pp.

Postel:1983:ITP

[Pos83] Jonathan Bruce Postel. Internet Telenet protocol and options. Technical report, Network Information Center, SRI International, Menlo Park, CA, USA, 1983. ???? pp.

Post:1995:IER

[Pos95] Joyce A. Post. Internet and E-mail resources on aging. Technical report, Philadelphia Geriatric Center, Philadelphia, PA, USA, 1995. 44 pp.

Potter:1993:BIG

[Pot93] James E. Potter. Bridging the Internet gap: how to make the world your on-line oyster! Bridge Learning Systems, Inc., American Canyon, CA, USA, csuh version. edition, 1993. ISBN 0-9632069-8-2. 100 pp. LCCN ????

Potter:1994:BIGa

[Pot94a] James E. Potter. Bridging the Internet gap: includes 7 on-line keystroke-by-keystroke tutorials to speed you on your way into the world of the Internet. Norbry Publishing, Terra Cotta, ON, Canada, 1994. ISBN 0-921282-48-6. ???? pp. LCCN ????

Potter:1994:BIGb

[Pot94b] James E. Potter. Bridging the Internet gap: make the world your on-line oyster! Bridge Learning Systems, Inc., American Canyon, CA, USA, second edition, 1994. ISBN 0-9632069-9-0. 117 pp. LCCN ????

Potter:1995:INS

[Pot95] James E. Potter. Internet navigation simulator version 2.2t, 1995. ISBN 1-56087-081-8. 2 computer disks.

Powell:1993:CEC

[Pow93] Jill H. Powell. Choosing engineering CD-ROM titles: Results of an Internet survey; study identifies which titles are hot today — and tomorrow. CD-ROM Professional, 6(6):34–??, November 1, 1993. CODEN CRPFEX. ISSN 1049-0833.

Powell:1998:WSE

[Pow98] Thomas A. Powell. Web Site Engineering: Beyond Web Page Design. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-650920-7. x + 324 pp. LCCN TK5105.888.P62

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0136509207.html.

Partridge:1991:IRI

[PP91] Craig Partridge and Stephen Pink. An implementation of the revised Internet stream protocol (ST-2). Research report SICS/R-91/14-SE, Swedish Institute of Computer Science, Kista, Sweden, 1991. 26 pp.

Pike:1994:IQ

[PP94] Mary Ann Pike and Tod G.

Pike. The Internet quickstart. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 1-56529-658-3. xv + 387 pp. LCCN TK 5105.875 I57 P55 1994. US$21.99. A series of quick tutorials, this book explains Internet basics to absolute beginners. Lots of extras — like buzzword definitions, tips, and warnings — help users get more from this premier online service. Task oriented skill sessions cover topics such as log-on, E-Mail, database searches, Internet news and more.

Peterson:1991:AFIa

[PR91a] Larry L. Peterson and Herman C. Rao. Accessing files in an Internet: the Jade file system. Technical Report 90-30, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, 1991. 32 pp.

Peterson:1991:AFIb

[PR91b] Larry L. Peterson and Herman C. Rao. Accessing files in an Internet: the Jade file system. Technical Report NASA CR-189930, National Aeronautics and Space Administration, Washington, DC, USA, 1991. Distributed to depository libraries in microfiche. Shipping list no.: 92-1583-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1992] 1 microfiche.

Perry:1995:IBC

[PR95] Kathy Perry and Charlie Roney. Internet boot camp, 1995. ISBN 0-9642739-1-8. 1 videocassette (105 min.) disks (3 1/2 in.) + 1 index card.

Press:1993:PCI

[Pre93] Larry Press. Personal computing: The Internet and interactive television. Communications of the ACM, 36 (12):19–23, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        163352.html.

Press:1994:CI

[Pre94] Larry Press. Commercialization of the Internet. Communications of the ACM, 37 (11):17–21, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
188286.html.

Press:1995:PCI

[Pre95a] Larry Press. Personal computing: The Internet is not TV: Web publishing. Communications of the ACM, 38(3):17–23, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/info.cern.ch/hypertext/
Web/Daemon/Overview.html;
 http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
203332.html.

Prevost:1995:II

[Pre95b]  Ruffin Prevost.  Internet

insider. Osborne/McGraw-Hill, Berkeley, CA, USA, 1995. ISBN 0-07-882084-7. xii + 284 pp. LCCN TK5105.875.I57 P747 1995. US$14.95. Collection of articles, messages, etc. originally posted on the Internet. Wild wacky and way out there. That’s the only way to describe this toast to Internet’s more bizarre newsgroups, mailing lists, e-mags, and WWW pages. Produced in a tabloid format, the “Insider” explores haiku written by dogs, the SNAG (Sensitive New Age Guy) movement, and the threat of America being overrun by space-lizard overlords, with lots of observations by Socks the Clinton Cat. It’s all a bit much, but this is the first book on the Internet able to be understood without owning a computer.

Press:1997:PCTc

[Pre97a]  Larry Press.  Personal

computing — tracking the global diffusion of the Internet. Communications of the ACM, 40(11):11–17, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1997-40-11/
                                        p11-press/.

PGP:1997:PGPa

[Pre97b] Pretty Good Privacy, Inc. Pretty Good Privacy PGPmail 4.5 program: quick guide. Pretty Good Privacy, Inc., San Mateo, CA, USA, 1997. ISBN 0-9649654-3-7. iv + 34 pp. LCCN ????

PGP:1997:PGPb

[Pre97c] Pretty Good Privacy, Inc. Pretty Good Privacy PGPmail 4.5 program: reference manual. Pretty Good Privacy, Inc., San Mateo, CA, USA, 1997. ISBN 0-9649654-4-5. xiv + 202 pp. LCCN ????

Price:1995:SNI

[Pri95] Gary David Price. IN SITE new Internet site additions, a current awareness fortnightly for the information professional read by information professionals around the world, 1995.

Probert:1996:IS

[Pro96] Ian Probert. Internet spy.

Classified. Kingfisher, New York, NY, USA, American edition, 1996. ISBN 0-7534-5007-0. ???? pp. LCCN HV6773.P76 1996.

Pancerella:1999:DCC

[PRY99] Carmen M. Pancerella, Larry Rahn, and Christine Yang. The Diesel Combustion Collaboratory: Combustion researchers collaborating over the Internet. In ACM [ACM99a], page ?? LCCN ????

Pinse:1992:EMG

[PS92] D. Pinse and P. Sylvester.

An electronic mail gateway between EARN/INTERNET and the IBM internal network. Computer Networks and ISDN Systems, 25(4–5): 526–??, November 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pierce:1995:CAP

[PS95a] Julie Pierce and Elizabeth Shaw. Child abuse prevention Internet resources, 1995.

Preston:1995:IPR

[PS95b] Larry D. Preston and J. David Shabat. Injury prevention resources on the Internet: a guide for educators, students and public health professionals. Technical report, Center for the Advanced Study of Public Safety and Injury Prevention, School of Public Health, University at Albany, Albany, NY (Executive Park South, 12203), 1995. 10 pp.

Postel:1981:AIP

[PSC81] Jonthan B. Postel, Carl A. Sunshine, and Danny Cohen. ARPA Internet Protocol. Computer Networks: The International Journal of Distributed Informatique, 5(4): 261–271, July 1981. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Peete:1995:BRIa

[PT95a] Gary R. Peete and Roy Tennant. Business resources on the Internet: a handson workshop. Internet workshop series; no. 3. Library Solutions Press, Berkeley, CA, USA, 1995. ISBN 1-882208-10-2. x + 138 pp. LCCN HF54.5 .P44 1995. Supplement to Crossing the Internet threshold.

Peete:1995:BRIb

[PT95b] Gary R. Peete and Roy Tennant. Business resources on the Internet plus: a handson workshop. Internet workshop series; no. 3. Library Solutions Press, Berkeley, CA, USA, 1995. ISBN 1-882208-11-0. xvi + 252 pp. LCCN HF54.5 .P44 1995. “Supplement to Crossing the Internet threshold.” “Includes Windows and Macintosh diskettes of presentation slides.” Computer disks contain “compressed files: PowerPoint View and all slide modules, including narratives”– Disk labels.

Piliouras:1999:NDM

[PT99] Teresa C. Piliouras and Kounel Terplan. Network Design: Management and Technical Perspectives. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-3404-7. 405 pp. LCCN TK5101 .R73 1999. US$79.95.

Phillips:1994:ASW

[PU94] David Phillips and Yale University. American studies web: a guide to American studies resources on the Internet, 1994.

PLA:1994:ILM

[Pub94] Public Library Association.

The Internet a library management perspective, 1994. 1 sound cassette.

PTI:1995:SLG

[Pub95] Public Technology, Inc., Washington, DC, USA. Surfing the ’net: a local government guide to Internet connection, 1995. 56 pp.

Pung:1991:DIF

[Pun91] H. K. Pung. Design and implementation of a facsimile service for Internet. Technical report TRF6/91, National University of Singapore, Dept. of Information Systems and Computer Science, Kent Ridge, Singapore, June 1991. 22 pp.

Pfaffenberger:1995:QCI

[PW95a] Bryan Pfaffenberger and Dave Wall. Que’s computer and Internet dictionary. Que Corporation, Indianapolis, IN, USA, sixth edition, 1995. ISBN 0-7897-0356-4. x + 574 pp. LCCN QA76.15P387 1995.

Protheroe:1995:IMC

[PW95b] Nancy Protheroe and Elizabeth Wilson. The Internet manual for classroom use. Technical report, Educational Research Service, Arlington, VA, USA, 1995. ii + 36 pp.

Protheroe:1994:IHS

[PWK94] Nancy Protheroe, Elizabeth Wilson, and Lorene Kluge. The Internet handbook for school users. Technical report, Educational Research Service, Arlington, VA, USA, 1994. iv + 108 pp.

Pyle:1996:ECI

[Pyl96] Raymond Pyle. Electronic commerce and the Internet. Communications of the ACM, 39(6):22–23, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
228507.html;   http://www.
acm.org/pubs/toc/Abstracts/
cacm/228507.html.

Pyra:1995:UIR

[Pyr95] Marianne Pyra. Using Internet relay chat. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0020-4. xviii + 217 pp. LCCN TK5105.886.P97 1995. Quarterman:1994:MCC

[QCM94a] John S. Quarterman and Smoot Carl-Mitchell. The E-mail companion: communicating effectively via the Internet and other global networks. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-40658-6. xvii + 318 pp. LCCN HE6239.E54 C36 1994, TK5105.875.I57 Q37 1994. US$19.95, CDN$25.95. This introduction assists the e-mail user through sending and replying to messages, sending mail to other systems (such as from America Online to Compuserve), using mailing lists, finding and downloading files, locating email addresses, and encoding non-text data with MIME. The material, however, is quite short-sighted in several places.

Quarterman:1994:ICS

[QCM94b] John S. Quarterman and Smoot Carl-Mitchell. The Internet Connection—System Connectivity and Configuration. Addison-Wesley UNIX and open systems series. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54237-4. xv + 271 pp. LCCN TK5105.875.I57 Q37 1994. US$32.25.

QDG:1995:BI

[QP95] Que Development Group and Beatriz Parra Perez, tr. La biblia de Internet. Informatica personal-profesional. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-809-7. 672 pp. LCCN ????

Quint:1995:SAE

[QRV95] Vincent Quint, Cécile Roisin, and Irène Vatton. A structured authoring environment for the World-Wide Web. Computer Networks and ISDN Systems, 27(6): 831–840, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/comnet/sub/1995/
                                        27/6/1462.pdf.

Quarterman:1990:M

[Qua90] John S. Quarterman. The Matrix. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Englewood Cliffs, NJ 07632, USA, 1990. ISBN 1-55558-033-5 (DP) and 0-13-565607-9 (PH). xxvii + 719 pp. LCCN TK5105.5 .Q37 1990. US$50.00. Digital order number EY-C176E-DP-SS. A context book about all computer networks worldwide, including UUCP, FidoNet, BITNET, USENET, and the Internet. The longest of all the books listed, and the most comprehensive.

Quarterman:1993:RIB

[Qua93] John S. Quarterman. Recent Internet Books, March 1993. This is a list of books related to using the Internet. The document (rfc1432.txt) can be obtained by anonymous

ftp from /rfc at nic.ddn.

mil. See [Bee93].

Que:1994:IRQ

[Que94] Que Development Group.

The Internet Resource: Quick Reference. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 1-56529-748-2. 492 pp. LCCN TK5105.875.I57 I57 199. US$20.95.

QBPLISMWT:1995:SI

[Que95] Queens Borough Public Library. Internet Self-Managed Work Team. Surfing the Internet. Technical report, Queens Borough Public Library, Jamaica, NY, USA, 1995. various pp.

Quarterman:1993:UPO

[QW93] John S. Quarterman and Susanne Wilhelm. UNIX, POSIX, and Open Systems: The Open Standards Puzzle. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52772-3. xxix + 416 pp. LCCN QA76.76.O63 Q37 1993. US$42.50. Context about open systems standards. Includes a very brief history of the Internet, comments on the effects of networks on standards and the reverse, commentary on the benefits of publishing standards online, and an indepth examination of the IETF standards process used to produce Internet Standards. Also includes an Internet growth graph.

Resnick:1996:PTS

[RA96a] Rosalind Resnick and Heidi Anderson. A pocket tour of shopping on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1808-9. xi + 179 pp. LCCN TX335.R45 1996.

Ruh:1996:IBL

[RA96b] Joseph F. Ruh and Computer Law Association. The Internet and business: a lawyer’s guide to the emerging legal issues. Current issues publications series. Computer Law Association, Washington, DC, USA?, 1996. ISBN 1-885169-05-1. vii + 165 + 8 pp. LCCN ????

Reed:1990:DTA

[RAD90] Sidney G. Reed, Richard H. Van Atta, and Seymour J. Deitchman. DARPA technical accomplishments: An historical review of selected DARPA projects. IDA Paper P-2192, Institute for Defense Analysis, 1801 N. Beauregard Street, Alexandria, VA 22311-1772, February 1990.

URL https://apps.dtic.
mil/sti/pdfs/ADA239925.
pdf.

Raeder:1995:IEa

[Rae95a] Aggi Raeder. The Internet express. Searcher: the magazine for database professionals, 3(1):26–??, January 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Raeder:1995:IEb

[Rae95b] Aggi Raeder. The Internet express. Searcher: the magazine for database professionals, 3(2):18–??, February 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Raeder:1995:IEc

[Rae95c] Aggi Raeder. The Internet express. Searcher: the magazine for database professionals, 3(3):16–??, March 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Raeder:1995:IEd

[Rae95d] Aggi Raeder. The Internet express. Searcher: the magazine for database professionals, 3(4):8–??, April 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Raeder:1995:IEe

[Rae95e] Aggi Raeder. The Internet express. Searcher: the magazine for database professionals, 3(5):28–??, May 1, 1995. CODEN SMDPE8. ISSN 1070-4795.

Rago:1993:USV

[Rag93]  Steven A. Rago.  UNIX

System V Network Programming. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-56318-5. xv + 784 pp. LCCN 92-45276, QA76.76.O63 R34 1993. US$45.50.

Rahardjo:1994:BPS

[Rah94] Budi Rahardjo. Buku pegangan sistem Unix dan Internet. Open Pathways, Bellingham, WA, USA, 1994. ISBN 1-885130-11-2. xiv + 251 pp. LCCN ????

Rakov:1995:ISM

[Rak95] Simon A. Rakov. The Internet sampler: a Mosaicbased museum kiosk about the Internet. In Anonymous [Ano95-64], page ???? ISBN ???? LCCN ???? URL http:

//www.ncsa.uiuc.edu/SDG/
IT94/Proceedings/Museum/
rakov/rakov.html. Includes

CD-ROM.

Ramsoy:1995:GSI

[Ram95] Tor Jakob Ramsoy. The governing structure of the Internet: coordination without central control. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1995. 135 pp.

Randall:1994:TYI

[Ran94] Neil Randall. Teach Yourself the Internet: Around the World in 21 Days. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30519-4 (paperback). xiv + 676 pp. LCCN TK5105.875.I57R36 1994. US$27.00. This wellorganized tutorial can be used by individuals, in seminars, training sessions, and classrooms. It takes readers on a global learning expedition of the Internet in just 21 fun-filled lessons.

Randall:1995:AID

[Ran95a] Neil Randall. Aprendiendo Internet en 21 dias. Prentice Hall Hispanoamericana, Mexico, DF, Mexico, 1995. ISBN 968-880-576-9 (paperback). xiv + 676 pp. LCCN ????

Randall:1995:TYI

[Ran95b] Neil Randall. Teach yourself the Internet in a week. Sams.net Pub., Indianapolis, IN, USA, second edition, 1995. ISBN 0-672-30735-9. xxi + 622 pp. LCCN TK5105.875.I57T43 1995.

Rankin:1996:DBP

[Ran96a] Bob Rankin. Dr. Bob’s painless guide to the Internet. No Starch Press, San Francisco, CA, USA, 1996. ISBN 1-886411-09-3. ???? pp. LCCN TK5105.875.I57R38 1996.

Rannenberg:1996:SUU

[Ran96b] W. Rannenberg. Supporting Unicode on a UNIX system. In Unicode Consortium [Uni96a], pages C5–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Rasmussen:1994:IEI

[Ras94] M. Rasmussen. Interactive edutainment on the Internet.

Computer Graphics, 28(2): 139–??, ???? 1994. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Ratzan:1993:ICa

[Rat93a] Lee Ratzan. The Internet Cafe. Wilson library bulletin, 67(9):79–??, May 1, 1993. ISSN 0043-5651.

Ratzan:1993:ICb

[Rat93b] Lee Ratzan. The Internet Cafe. Wilson library bulletin, 67(10):91–??, June 1, 1993. ISSN 0043-5651.

Ratzan:1993:ICc

[Rat93c] Lee Ratzan. The Internet Cafe. Wilson library bulletin, 68(2):75–??, October 1, 1993. ISSN 0043-5651.

Raucci:1995:MW

[Rau95] Richard Raucci. Mosaic for Windows: a Hands-on Configuration and Set-up Guide to Popular Web Browsers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-97996-4. 192 pp. LCCN TK5105.882 .R36 1995. US$19.95.

Ravipati:1995:SI

[Rav95] Venkata R. Ravipati. Security on the Internet. Technical report, ????, ????, 1995. 37 pp.

Rotenberg:1995:CRD

[RAWS95] Marc Rotenberg, Ram Avrahami, Sanford Wallace, and Robert Sherman. Consumer rights with direct marketing on and off the Internet: does junk e-mail really bite?, 1995. 1 videocassette (VHS) (70 min.).

Reiner:1996:PPI

[RB96] Diane Reiner and Keith A. Blanton. Person to person on the Internet. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-104245-6. ???? pp. LCCN TK5105.73.R45 1996.

Rubin:1999:UD

[RB99] William Rubin and Marshall Brain. Understanding DCOM. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-095966-9. xv + 298 pp. LCCN QA76.9.D5R83 1999.

URL http://www.phptr.
com/bookseri/microsoft.
html;   http://www.phptr.
com/ptrbooks/ptr_0130959669.

html. Includes CD-ROM.

Reinwand:1994:PUL

[RBK^{+}94] Louis Reinwand, Jay Lance Burton, Carolyn Klatt, Sandi Long, and Jane Smith. Plug in to the Utah Library Network, reach out to the world: Utah Library Network and Internet training handbook. Information forum publication; #7, Utah State Library Division, Dept. of Community and Economic Development, Salt Lake City, UT, USA, 1994. various pp.

Rutten:1994:NG

[RBM94a] Peter Rutten, Albert Bayers III, and Kelly Moloni. Net Guide: Your Map to the Services, Information and Entertainment on the Electronic Highway. Random House, New York, NY, USA, 1994. ISBN 0-679-75106-8. xix + 356 pp. LCCN TK 5105.5 R982n 1994. US$19.00. As the ‘TV Guide of Cyberspace,’ this book provides pointers by subjects to Internet sites, USENET newsgroups, and commercial resources (CompuServe, American Online, etc.). It proves that there’s something for everyone somewhere out in the electronic world.

Rutten:1994:NGY

[RBM94b] Peter Rutten, Albert F. Bayers III, and Kelly Maloni. NET GUIDE: Your Map to the Services, Information and Entertainment on the Electronic Highway. Random House, New York, NY, USA, 1994. ISBN 0-679-75106-8. xix + 356 pp. LCCN TK 5105.5 R88 199. US$19.00.

Rodgers:1996:EML

[RCL96] Richard T. Rodgers, Jane E.

Colwin, and Paul D. Leopoulos. Effective modern legal research: using cd-rom, online services, and the Internet. Technical report, State Bar of Wisconsin CLE, Madison, WI, USA, 1996. various pp.

Reed:1995:DCL

[RD95] Chris Reed and Lars Davies. Digital cash — the legal implications. Technical report, IT Law Unit, Centre for Commercial Law Studies, Queen Mary and Westfield College, London, UK, 1995. 30 pp.

Rimmer:1995:IU

[RdB95] Steve Rimmer and Laurent de Blois. Internet Underground. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN ???? 305 pp. LCCN ????

Riche:1995:GMS

[RdG95] Morgan Riche and dir Guillermont. Guide des 1000 meilleurs sites Internet. F1RST, Paris, France, 1995. ISBN 2-87691-293-7. 234 pp. LCCN ????

Read:1998:EC

[Rea98]  Richard Read.  Essence

of Communications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-521022-4. ???? pp. LCCN TK5102.9.R43

                                        1998. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0135210224.html.

Reece:1993:IRG

[Ree93] Peter A. (Peter Andrew)

Reece. Internet: a reference guide for taming the dragon. LM Publications, Duncan, BC, Canada, 1993. ISBN 1-895524-01-6. 150 pp.

Regan:1993:IJT

[Reg93] Tom Regan. Internet as a journalist’s tool. Nieman reports, 47(3):26–??, Fall 1993. ISSN 0028-9817.

Reh:1996:BPI

[Reh96] Jane Phoenix Reh. Blackbird publishing on the Internet. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-358-1. ???? pp. LCCN ????

Reinalda:1992:CIP

[Rei92] Johannes K. Reinalda. A controller for Internet protocol routing of AX.25 packets. Thesis (M.S.), Oregon State University, Corvallis, OR, USA, 1992. 56 pp.

Reid:1993:ECS

[Rei93] Elizabeth Reid. Electronic chat: Social issues on Internet Relay Chat. Media information Australia, 67:62– ??, February 1, 1993. ISSN 0312-9616.

Anonymous:1992:REN

[Res92] Research and education networking, 1992. ISSN 1051-4791. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA.

Anonymous:1995:RBM

[Res95a] Resources for building and managing Internet services, page various, 1995. ISSN 1082-474X. Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA.

Ressler:1995:ITI

[Res95b] Michael P. Ressler. Introduction to TCP/IP and Internet security, 1995. ISBN 1-57305-053-9. 1 videocassette (45 min.) (various pagination: ill.; 28 cm.).

Ressler:1997:AEP

[Res97] Sandy Ressler. The Art of Electronic Publishing: the Internet and beyond. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-488172-9. xvii + 450 pp. LCCN Z286.E43R47 1997. US$39.95.

Reynolds:1991:HI

[Rey91]  J. K. Reynolds.  The

helminthiasis of the Internet. Computer Networks and ISDN Systems, 22(5):347– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Reynolds:1993:EID

[Rey93] Dennis Reynolds. Evaluating Internet Dial-Up access options. Computers in Libraries, 13(8):86–??, September 1, 1993. CODEN CPLIE8. ISSN 1041-7915.

Rekhter:1991:ABG

[RG91] Y. Rekhter and P. Gross. Application of the border gateway protocol in the Internet, October 1991. 1 data file 5 1/4 in. Obsoletes RFC 1164. Available via FTP.

Ruth:1993:MIC

[RG93] Stephen R. Ruth and Raul Gouet. Must invisible colleges be invisible? an approach to examining large communities of network users. Internet research, 3(1):36–??, Spring 1993. CODEN IRE-SEF. ISSN 1066-2243.

Ryder:1997:IE

[RH97] Randall J. Ryder and Tom Hughes. Internet for educators. Merrill, Upper Saddle River, NJ, USA, 1997. ISBN 0-13-239187-2 (paper). ???? pp. LCCN LB1044.87.R93 1997.

Rheingold:1994:VCHa

[Rhe94a]  Howard Rheingold.  The

Virtual Community: Homesteading on the Electronic Frontier. HarperPerennial, New York, NY, USA, 1994. ISBN 0-06-097641-1. LCCN TK5105.875.I57 R45 1994. US$13.00.

Rheingold:1994:VCHb

[Rhe94b]  Howard Rheingold.  The

Virtual Community: Homesteading on the Electronic Frontier. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-60870-7. viii + 325 pp. LCCN TK5105.5 .R48 1993. US$22.95.

Rosenbaum:1995:ISI

[RHK95] Howard Rosenbaum, Carol A. Hert, and Jeffrey Kosokoff. Internet skills for information networking. Technical report, School of Library and Information Science, Indiana University, Bloomington, IN, USA, 1995. 172 pp.

Ross:1995:QTI

[RHS95] C. P. Ross, S. Hassanzadeh, and B. R. Spies. A quick tour of Internet and the Information Superhighway, part 1: On-ramps, pit stops, and exits. Leading edge (Tulsa, OK), 14(3):179–??, March 1, 1995. CODEN LEEDFF. ISSN 0732-989X.

Ribeiro:1995:IEC

[Rib95] Gustavo Lins Ribeiro. Internet e a emergencia da comunidade imaginada transnacional. Serie antropologia; 181 serie antropologia (Brasilia, Brazil); no. 181, Departamento de Antropologia, Universidade de Brasilia, Brasilia, Brazil, 1995. 8 pp.

RURST:1994:INT

[Ric94a] Rice University. RiceInfo Support Team. Internet navigation tools, 1994.

Richardson:1994:NI

[Ric94b] Robert Richardson. Navigating the Internet. LAN: the network solutions magazine, pages 47–??, October 1, 1994. CODEN LANNER. ISSN 0898-0012.

Richards:1995:CBG

[Ric95a] Linda Richards. The Canadian business guide to using the Internet: what you can get if you’re on the net. Selfcounsel business series. Self-Counsel Press Ltd., North Vancouver, BC, Canada, 1995. ISBN 0-88908-849-7. xvii + 133 pp. LCCN ????

Richardson:1995:NI

[Ric95b] Brooke Richardson. News on the Internet. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1995. various pp.

Richardson:1995:WWS

[Ric95c] Eric Richardson. A walk on the weird side. Internet World, 6(4):80–??, April 1, 1995. CODEN IERNE8. ISSN 1064-3923.

Riis:1982:PMV

[Rii82] Jens Ove Riis, editor. Project management–tools and visions: proceedings of the 7. Internet World Congress, 1982. Projektplan, Copenhagen, Denmark, 1982. ISBN 87-980516-9-5. LCCN HD69.P75 I57 1982. Four volumes. “The World Congress on Project Management in the Copenhagen Congress Center from Sept. 12th to 17th, 1982 is … sponsored by the International Management Systems Association [and is hosted by] PROJEK-TPLAN, the Danish Project Management Society.” — P. [iii] “Inspirations from Copenhagen” — P. [4] of cover.

Rikitake:1994:ICK

[Rik94] Kenji Rikitake. Intanetto comyunithi: kokusai nettowaku saizensen = Internet community. Omusha, Tokyo, Japan, dai 1-han edition, 1994. ISBN 4-274-06090-X. xi + 276 pp. LCCN ????

Riley:1995:EOJ

[Ril95] Margaret F. Riley. Employment opportunities and job resources on the Internet: sometimes called the Riley guide [as adapted by Career Resources of the Academic Skills Center, Asnuntuck Community-Technical College]. Technical report, Asnuntuck Community-Technical College, Academic Skills Center Career Resources, Enfield, CT, USA, 1995. 86 pp.

Rimmer:1994:MI

[Rim94] Steve Rimmer. Mondo Internet. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1994. ISBN 0-07-053014-9, 0-07-053015-7 (paperback). ???? pp. LCCN TK5105.875.I57 R46 1994. US$36.95, US$24.95 (paperback).

Rimmer:1995:PI

[Rim95a] Steve Rimmer. Planet Internet. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1995. ISBN 0-07-053014-9 (recycled paper), 0-07-053015-7 (paperback recycled paper). xviii + 315 pp. LCCN TK5105.875.I57R46 1995.

Rimmer:1995:RPI

[Rim95b] Steve Rimmer. Return to planet Internet. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-053021-1. xxx + 300 pp. LCCN TK5105.875.I57R48 1995.

Rimmer:1996:IGT

[Rim96] Steve Rimmer. The Internet graphics toolkit. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912183-7 (hardcover), 0-07-912184-5 (paperback). xvii + 238 pp. LCCN T385.R5525 1996. Includes CD-ROM.

Rittner:1992:EEG

[Rit92] Don Rittner. EcoLinking:

Everyone’s Guide to Online Environmental Information. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1992. ISBN 0-938151-35-5. xi + 352 pp. LCCN GE30 .R57 1992.

US$18.95. Directed at concerned citizens, environmentalists and scientists interested in sharing ideas and research on environmental issues. Covers resources on FI-DONET, BITNET, INTER-NET, USENET, local bulletin boards, America Online, CompuServe, EcoNet, GE-NIE, WELL.

Rittner:1993:WEO

[Rit93] Don Rittner. The Whole Earth Online Almanac: Info from A to Z. Robert J. Brady Co., Bowie, MD 20715, USA, 1993. ISBN 1-56686-090-3. xiv + 545 pp. LCCN QA76.55 .R58 1993. US$32.95.

Rose:1995:GRO

[RJ95] Marshall Rose and Bruno Joachim. Gestion des réseaux ouverts: SNMPv2. Informatique intelligence artificielle. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 2-7296-0540-1. ix + 319 pp. LCCN ????

Rosenfeld:1994:ICG

[RJV94] Louis B. Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: guides to resources by subject. Neal-Schuman Publishers, New York, NY, USA, 1994. ISBN 1-55570-188-4. ???? pp. LCCN ????

Rosenfeld:1995:ICSb

[RJV95a] Louis Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: subject guides to health and science resources. Mansell, London, UK, 1995. ISBN 0-7201-2303-8 (paperback). ix + 529 pp. LCCN R859.7.D36 R67 1995.

Rosenfeld:1995:ICSd

[RJV95b] Louis Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: subject guides to humanities resources. Mansell, London, UK, 1995. ISBN 0-7201-2304-6 (paperback). ix + 368 pp. LCCN ????

Rosenfeld:1995:ICSe

[RJV95c] Louis Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: subject guides to social sciences, business, and law resources. Neal-Schuman Publishers, New York, NY, USA, 1995. ISBN 1-55570-220-1. ix + 424 pp. LCCN H61.95 .I58 1995.

Rosenfeld:1995:ICSf

[RJV95d] Louis Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: subject guides to social sciences, business and law resources. Mansell, London, UK, 1995. ISBN 0-7201-2305-4 (paperback). ix + 424 pp. LCCN ????

Rosenfeld:1995:ICSa

[RJV95e] Louis B. Rosenfeld, Joseph Janes, and Martha Vander Kolk. Internet compendium: subject guides to health and science resources. Neal-Schuman Publishers, New York, NY, USA, 1995. ISBN 1-55570-219-8. ix + 529 pp. LCCN R859.7.D36 R67 1995.

Rosenfeld:1995:ICSc

[RJV95f] Louis B. Rosenfeld, Joseph Janes, and Martha Vander Kolk. The Internet compendium: subject guides to humanities resources. Neal-Schuman Pub., New York, NY, USA, 1995. ISBN 1-55570-218-X (paper). ix + 368 pp. LCCN AZ221 .I58 1995.

Reddick:1995:OJU

[RK95a] Randy Reddick and Elliot King. The online journalist: using the Internet and other electronic resources. Harcourt Brace College, Fort Worth, TX, USA, 1995. ISBN 0-15-502018-8 (paperback). xi + 251 pp. LCCN PN4784.E5 R43 1995.

Robison:1995:IPN

[RK95b] David F. W. Robison and Jonathan Kochmer. The Internet passport: NorthWest-Net’s guide to our world online. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 1995. ISBN 0-13-194200-X. xxxviii + 667 pp. LCCN TK5105.875.I57 R63 1994. US$29.95.

Rowland:1995:RIC

[RK95c] Robin Rowland and Dave Kinnaman. Researching on the Internet: the complete guide to finding, evaluating, and organizing information effectively. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 0-7615-0063-4. xvi + 384 pp. LCCN TK5105.875.I57R73 1995.

Reddick:1996:OSM

[RK96] Randy Reddick and Elliot King. The online student: making the grade on the Internet. Harcourt Brace College Publishers, Fort Worth, TX, USA, 1996. ISBN 0-15-503189-9 (paperback). xiii + 317 pp. LCCN ????

Randall:1995:PI

[RL95a] Neil Randall and Celine Latulipe. Plug-n-play Internet. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30669-7. xiii + 383 pp. LCCN TK515.875.I57R34 1995.

Robbins:1995:DBI

[RL95b] Judd Robbins and LearnKey, Inc. Doing business on the Internet, 1995. 1 videocassette (88 min.).

Rowe:1994:IIT

[RM94] Lawrence A. Rowe and Paul Manley. Internet informational tools for classes, 1994. 1 videocassette (87 min.).

Resnick:1996:PIA

[RM96] Paul Resnick and James Miller. PICS: Internet access controls without censorship. Communications of the ACM, 39(10):87–93, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/pubs/toc/
Abstracts/cacm/234989.html;
 http://www.acm.org/pubs/
toc/Abstracts/cacm/236175.
html.

Robert:1994:IWL

[RML^{+}94] John Robert, John Mueller, Karen Long, Phred Tinampay, Sophia Marsh, and Serina Yang. Internet world live, 1994. 1 videocassette (60 min.).

Ruffin:1996:PHG

[RMU96] M. Holt Ruffin, Joan Mc-Carter, and Richard Upjohn. The post-Soviet handbook: a guide to grassroots organizations and Internet resources in the newly independent states. Center for Civil Society International, in association with University of Washington Press, Seattle, WA, USA, 1996. ISBN 0-295-97534-2. xvii + 393 pp. LCCN HS71.F6R84 1996.

Robison:1993:II

[RN93] David Robison and Michael Nanfito. An introduction to the Internet, 1993. Videocassette. “Part of larger project Internet Planning and Training for California Libraries sponsored in whole or part by U.S. Department of Education under provision of Library and Construction Act administered in California by State Librarian.” Performers: David F. W. Robison. Performers: Introduction by Gary Strong.

Ross:1997:ITA

[RN97] John Ross and Peter Norton. The Internet: a tutorial to accompany Peter Norton’s Introduction to Computers. Glencoe/McGraw Hill, New York, NY, USA, 1997. ISBN 0-02-802969-0. ???? pp. LCCN ????

Ryan:1994:MNU

[RNW^{+}94] George H. Ryan, Pat Norris, Jean Wilkins, Barb Allen, Sue Sueflow, Bill Erbes, and Beth McNeil. Mining “the net” using Internet for reference, 1994. 1 videocassette (2 hrs.).

Rodgers:1994:UIGa

[RO94a] William J. Rodgers and Melissa Oleen. USD Internet guide: navigating Internet services. Technical report, I. D. Weeks Library, University of South Dakota, Vermillion, SD, USA, Fall 1994. various pp.

Rodgers:1994:UIGb

[RO94b] William J. Rodgers and Melissa Oleen. USD Internet guide: navigating Internet services. Technical report, I. D. Weeks Library, University of South Dakota, Vermillion, SD, USA, Summer 1994. 141 pp.

Roach:1995:MGI

[Roa95a] J. Michael Roach. 10 minute guide to Internet Assistant for Word. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-616-X (paperback). vi + 168 pp. LCCN QA76.76.H94R63 1995.

Roach:1995:MGM

[Roa95b] J. Michael Roach. 10 minute guide to Microsoft Internet Explorer. 10 minute guides. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0628-8. xii + 191 pp. LCCN TK5105.875.I57R62 1995.

Robison:1994:AAI

[Rob94] David F. W. Robison. All about Internet FTP: Learning and Teaching to Transfer Files on the Internet. Internet workshop series; no.

2. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, second printing, minor revisions edition, 1994. ISBN 1-882208-04-8 (book), 1-882208-06-4 (book with diskettes). x + 90 pp. LCCN TK5105.875.I57 R63 1994 Ref pt. 1 (text), pt. 2– 3 (computer disks). US$30. “Supplement to Crossing the Internet threshold.” Title on computer disk label: All about Internet FTP plus. On computer disk label: Windows. On computer disk label: Macintosh.

Roberts:1995:DIW

[Rob95] Dave Roberts. Developing for the Internet with Winsock. Coriolis: Distributed to the book trade by IDG Books Worldwide, Scottsdale, AZ, USA, 1995. ISBN 1-883577-42-X (paperback). xvii + 620 pp. LCCN TK5105.55.R64 1995.

Rockman:1993:TAI

[Roc93] Ilene F. Rockman. Teaching about the Internet: The formal course option. The reference librarian, 39:65–??, 1993. CODEN RELBD6. ISSN 0276-3877.

Rogers:1997:OON

[Rog97] Joey Rogers. Object-Oriented Neural Networks in C++. Academic Press, New York, NY, USA, 1997. ISBN 0-12-593115-8. x + 310 pp. LCCN QA76.87 .R64 1997. US$41.95.

Romm:1996:JGI

[Rom96] Diane Romm. The Jewish guide to the Internet. Jason Aronson, Northvale, NJ, USA, 1996. ISBN 1-56821-914-8. xiv + 219 pp. LCCN BM67 .R66 1996 Ref.

Rooker:1997:BRI

[Roo97] Terry Rooker. Book review: Internet Economics. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
1997-10/mcknight.html.

Rose:1984:LTC

[Ros84]  Marshall T. Rose.  Low

tech connections into the ARPA Internet: the Raw-Packet split-gateway. Technical report 216, Department of Information and Computer Science, University of California, Irvine, Irvine, CA, USA, 1984. 14 pp.

Ross:1986:FT

[Ros86] Floyd E. Ross. FDDI — a tutorial. IEEE Communications Magazine, 24(5):10– 17, May 1986. CODEN ICOMD9. ISSN 0163-6804. A good general introduction to FDDI techniques and terminology. You will find it easier to read if you have a fair amount of prior experience and vocabulary in network technology. The proposed applications now seem a little dated, but the basic ideas remain valid. See [Ros89].

Ross:1989:OFF

[Ros89] Floyd E. Ross. An overview of FDDI: The fiber distributed data interface. IEEE Journal on Selected Areas in Communications, 7(7):1043– 1051, September 1989. CO-DEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic). An updated version of the same author’s 1986 paper [Ros86]. This version reflects three years’ more experience and goes into more detail in some areas. For this reason you may find it a little less readable as an introduction.

Rose:1991:SB

[Ros91] Marshall T. Rose. The Simple Book. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-812611-9. xxix + 347 pp. LCCN TK5105.5.R68 1991. US$54.00.

Rose:1993:IMC

[Ros93] Marshall T. Rose. The Internet message: closing the book with electronic mail. Prentice Hall series in innovative technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-092941-7. xxvi + 370 pp. LCCN TK5105.73 .R66 1993. US$44.00.

Rose:1994:MYC

[Ros94a] Donald Rose. Minding Your Cybermanners on the Internet. Alpha Books, Indianapolis, IN, USA, 1994. ISBN 1-56761-521-X. 194 pp. LCCN TK 5105.875 I57 R66 1994. US$12.99. From the publisher: “An entertaining and thorough guide to the critical do’s and don’ts of Internet behavior and communication for newcomers. Smileys, emoticons, and many helpful tips helps the newbie blend into the net culture with grace and ease.”.

Rose:1994:SBI

[Ros94b] Marshall T. Rose. The simple book: an introduction to Internet management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1994. ISBN 0-13-177254-6. xxiv + 456 pp. LCCN TK5105.5 .R68 1994. US$55.00.

Rose:1995:ICQ

[Ros95a] Donald Rose. Internet Chat quick tour: real-time conversations and communications online. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-223-2. xvi + 174 pp. LCCN TK5105.886 .R68 1995. US$14.00.

Ross:1995:IPT

[Ros95b] John Ross. Internet power tools. Random House Electronic Pub., New York, NY, USA, 1995. ISBN 0-679-75815-1, 0-679-76531-X (disk). x + 374 pp. LCCN ????

Ross:1995:PRC

[Ros95c] Steven S. Ross. Public relations in cyberspace. The Public relations journal, 51 (1):36–??, May 1, 1995. CO-DEN PREJAR. ISSN 0033-3670.

Ross:1996:UGM

[Ros96] John Ross. The underground guide to Microsoft Internet assistant: slightly askew advice on mastering the Web with WinWord. The underground guide series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48944-9. xvii + 174 pp. LCCN QA76.76.H94R67 1996.

Roth:1995:UIM

[Rot95a] Clark E. Roth. The use of Internet in the marketing and management classrooms in postsecondary institutions. Thesis (M.S.), Emporia State University, Emporia, KS, USA, 1995. 21 pp.

Rothenberg:1995:IEC

[Rot95b] Dianne Rothenberg. The Internet and early childhood educators some frequently asked questions. ERIC digest EDO-PS-95-5, ERIC Clearinghouse on Elementary and Early Childhood Education University of Illinois, Urbana, IL, USA, 1995.

Rothman:1996:NWP

[Rot96] David H. Rothman. Networld!: what people are really doing on the Internet, and what it means to you. Prima Publishing, Rocklin, CA, USA, 1996. ISBN 0-7615-0013-8 (paperback), 0-7615-0018-9 (paperback). vii + 344 pp. LCCN TK5105.875.I57R69 1996.

Roudebush:1994:DSS

[Rou94] Clair Roudebush. Directory of safety science educators: institutions, addresses, degree data, personnel, Internet addresses. Technical report, ????, ????, 1994. vi + 38 pp.

Rowe:1996:BID

[Row96] Jeff Rowe. Building Internet database servers with CGI. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-573-9, 1-56205-575-5. ???? pp. LCCN TK5105.888.R68 1996.

Rao:1993:AFI

[RP93] H. C. Rao and L. L. Peterson. Accessing files in an Internet: The Jade file system. IEEE Transactions on Software Engineering, 19(6):613–??, June 1, 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Riley:1994:PBI

[RR94a] Mark Riley and Joe Rizzo. Putting business on the Internet, 1994. 1 videocassette (46 min.).

Rousseau:1994:WDP

[RR94b] Bertrand Rousseau and Mario Ruggier. Writing documents for paper and WWW. A strategy based on FrameMaker and WebMaker. Computer Networks and ISDN Systems, 27(2):205–214, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1994&volume=27&
issue=2&aid=1383.

Ram:1995:IKS

[RR95a] P. Ram and D. K. Rand. Internet kiosk: Satan: doubleedged sword. Computer, 28 (6):82–83, June 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ram:1995:SDS

[RR95b] P. Ram and D. K. Rand.

Satan: double-edged sword. Computer, 28(6):82–83, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Randall:1995:ISK

[RRF95] Lisa Randall, Maurizio Riccio, and R. Toy Fowler. Internet starter kit. Technical report, Prepared by the Bureau of Transportation Statistics, Information Technology Center, Washington, DC, USA, 1995. vi + 73 pp.

Riley:1996:GIJ

[RRO96] Margaret Riley, Frances E. Roehm, and Steve Oserman. The guide to Internet job searching. VGM Career Horizons, Lincolnwood, IL, USA, 1996. ISBN 0-8442-8197-2 (paperback). ix + 213 pp. LCCN HF5382.7.R557 1996.

Rai:1998:HAI

[RRS98] Arun Rai, T. Ravichandran, and Subhashish Samaddar. How to anticipate the Internet’s global diffusion. Communications of the ACM, 41(10):97–106, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-10/
                                        p97-rai/.

Raymond:1991:NHD

[RS91] Eric S. Raymond, ed. and Guy Steele. The New Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-18145-2 (hardcover), 0-262-68069-6 (paperback). xx + 433 pp. LCCN PN6231.E4 H3 1991. US$10.95 (hardcover). The authority on hacker jargon, and a very amusing book. Look it up in here when you doubt a definition in the press.

Raymond:1993:NHD

[RS93a] Eric Raymond and Guy L.

Steele, editors. The New Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, second edition, 1993. ISBN 0-262-68079-3. xxi + 505 pp. LCCN PN 6231 E4H3 1993. US$14.95. The New Hacker’s Dictionary is a great book for learning about the various slang, jargon and customs and folklore of the ’net (as well as other lairs of the hacker.) Very silly and highly recommended. You can get it for free online (The Jar-

gon File) at http://www.
ccil.org/jargon/jargon.

html, but the bound book makes great bathroom reading and contains silly cartoons and stuff.

Rost:1993:DHR

[RS93b] Martin Rost and Michael Schack. DFUe — Ein Handbuch. Recherchen in weltweiten Netzen (DFUe — A Handbook. Research in Worldwide Nets). Heise-Verlag, Hannover, Germany, 1993. ISBN 3-88229-026-9. 389 pp. DM58.00. An introduction to Internet (e-mail, FTP, telnet, IRC, archie, gopher, hytelnet, wais, WWW), uucp-Net, Bitnet (Netserver, Listservs, Trickle-Server, Bitftp, Astra), WIN, FidoNet, Maus-Net and Z-Net with many examples for each.

Rus:1997:CIC

[RS97] Daniela Rus and Devika Subramanian. Customizing information capture and access. ACM Transactions on Information Systems, 15(1):67– 101, January 1997. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/rus.html.

Rose:1998:IMD

[RS98] Marshall T. Rose and David Strom. Internet Messaging: From the Desktop to the Enterprise. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-978610-4. xxiv + 307 pp. LCCN TK5105.73.R6623

                                        1998. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0139786104.html.

Rao:1998:MI

[RSD98] H. Raghav Rao, A. F. Salam, and Brian DosSantos. Marketing and the Internet. Communications of the ACM, 41 (3):32–34, March 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-3/
                                        p32-rao/.

Resnick:1994:IBGb

[RT94a] Rosalind Resnick and Dave Taylor. Internet business guide. Sams.net Pub., Indianapolis, IN, USA, second edition, 1994. ISBN 1-57521-004-5. xxxi + 470 pp. LCCN HD30.335 .R47 1995.

Resnick:1994:IBGa

[RT94b] Rosalind Resnick and Dave Taylor. The Internet business guide: riding the information superhighway to profit. SAMS Publishing, Indianapolis, IN, USA, September 1994. ISBN 0-672-30530-5. xxxix + 418 pp. LCCN HD30.335 .R47 1994. US$25.00, CDN$34.95. Of the various books on doing business on the Internet this is the best one I’ve seen. By reading this book, you’ll find out how to use the Internet to build market share, track down business leads, communicate with colleagues and search online databases, all through extensive examples and detailed case studies. The authors also explain how you can offer cost-effective customer support, and access critical information using the net. Explores the many opportunities available on the Internet, and get tips on shipping for the best deal on Internet access. For leading edge marketing strategies and bottom-line information.

Rose:1995:NS

[RT95] Candi Rose and Dirk Thomas. net.sex. Howard W. Sams, Indianapolis, IN 46268, USA, January 1995. ISBN 0-672-30702-2. 250 pp. US$19.99. Subtitled “Discreet Guide to the Adult Side of the Internet” it gives “Practical advice on how to access the ‘adult side of the Internet”’. Indeed, it does. It’s a surprisingly wellwritten, thoughtful and detailed guide to finding adultoriented stuff on the ’net. Lots of insightful text and not one sleazy picture. Includes info on discussion lists, anonymity, finding dirty pictures and so on.

Raghavan:1998:NMS

[RT98] S. V. Raghavan and Satish K. Tripathi. Networked multimedia systems: concepts, architecture and design. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-210642-6. xvii + 397 pp. LCCN QA76.575.R34

                                        1998. URL http://www.
                                        prenhall.com/allbooks/esm_
                                        0132106426.html.

Rexford:1999:SVB

[RT99] Jennifer Rexford and Don Towsley. Smoothing variablebit-rate video in an Internetwork. IEEE/ACM Transactions on Networking, 7(2):202–215, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1999-7-2/p202-rexford/.

Rubinstein:1994:IE

[Rub94] Dick Rubinstein. The Internet for everybody, 1994. Includes CD-ROM and videocassette (123 min.).

Rubinstein:1995:RI

[Rub95] Ed Rubinstein. Retailing on the Internet. A discount store news special report, Lebhar-Friedman, New York, NY, USA, 1995. 83 pp.

Rudkin:1995:SSO

[Rud95]  Kristine M. Rudkin.  A

system for search and organization of Internet resources using information retrieval methodologies. Thesis (M.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1995. ix + 135 pp.

Russett:1994:TPT

[Rus94] James A. Russett. Telecommunications and pre-service teachers: the effects of using electronic mail and a directed exploration of Internet on attitudes. Thesis (Ed. D.), University of Nebraska–Lincoln, Lincoln, NE, USA, 1994. v + 113 pp.

Ruskin:1998:NIT

[Rus98] K. J. Ruskin. New Internet technology: HTML tools, Java, teleconferencing, and streaming multimedia. In Ikeda et al. [IDK98], pages 229–234. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

Rutter:1996:CI

[Rut96] Michael Rutter. Conquering the Internet. South-Western Educational Publishing, Cincinnati, OH, USA, 1996. ISBN 0-538-65873-8 (paperback). xii + 196 pp. LCCN TK5105.875.I57C66 1996.

Rossi:1999:IBC

[RV99] D. Rossi and F. Vitali. Internet-based coordination environments and documentbased applications: a case study. Lecture Notes in Computer Science, 1594:259–274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1995:EBB

[RWD95] Chris Reed, Ian Walden, and Lars Davies. Electronic bulletin boards and the law. Internet law research project, Centre for Commercial Law Studies, Queen Mary and Westfield College, London, UK, 1995. 24 pp.

Ryan:1993:RRP

[Rya93] Joe Ryan. Resource review: a pathfinder to core resources for network users. Internet research, 3(2):69–??, Summer 1993. CODEN IRE-SEF. ISSN 1066-2243.

Ryan:1994:GGI

[Rya94] Joe Ryan. A guide to government information available on the Internet. Technical report, Syracuse University, Syracuse, NY, USA, September 1, 1994. 58 pp.

Ryan:1995:GGIa

[Rya95a] Joe Ryan. Guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, January 1995. 186 pp.

Ryan:1995:GGIb

[Rya95b] Joe Ryan. Guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, January 1995. 196 pp.

Ryan:1995:GGIc

[Rya95c] Joe Ryan. Guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, April 1995. 212 pp.

Ryan:1995:GGId

[Rya95d] Joe Ryan. Guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, August 1995.

Ryan:1995:GGIe

[Rya95e] Joe Ryan. Guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, September 1995. 454 pp.

Ryan:1995:RBMa

[Rya95f] Joe Ryan. Resources for building and managing Internet services. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, January 1995. 101 pp.

Ryan:1995:RBMb

[Rya95g] Joe Ryan. Resources for building and managing Internet services. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, April 1995. 108 pp.

Ryan:1995:RBMc

[Rya95h] Joe Ryan. Resources for building and managing Internet services. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, 1995. 120 pp.

Ryan:1995:UGS

[Rya95i] John Ryan. A uses and gratifications study of the Internet social interaction site LambdaMOO: talking with “dinos”. Thesis (M.A.), Ball State University, Muncie, IN, USA, 1995. 60 pp.

Ryan:1996:GGI

[Rya96a] Joe Ryan. A guide to government information available on the Internet. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, March 1996. 339 pp.

Ryan:1996:RBM

[Rya96b] Joe Ryan. Resources for building and managing Internet services. Technical report, Ryan Information Management Press, 1049 Ackerman Ave., Syracuse, NY 13210-3035, USA, March 1996. 165 pp.

Ryer:1995:PTH

[Rye95] Jeanne C. Ryer. A pocket tour of health and fitness on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1711-2 (paperback). xiv + 176 pp. LCCN R859.7.D36R94 1995.

Raysman:1996:CLI

[RZ96] Richard Raysman and Joseph P.

Zammit. Computer law: Internet and other complex technologies. Technical report, Law Journal Seminars-Press, New York, NY, USA, 1996. 704 pp.

Slingers:1997:SFT

[Sac97] Sachem Web Slingers. The science of future technology: Spinning webs into the 21st century. Communications of the ACM, 40(2):125– 128, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/cacm/1997-40-2/
p125-sachem_web_slingers/

.

Salkind:1995:HI

[Sal95a] Neil J. Salkind. Hands-on Internet. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1995. ISBN 0-7895-0161-9. ix + 198 pp. LCCN TK5105.875.I57S34 1995.

Salomon:1995:WIC

[Sal95b] Norbert Salomon. Warping to the Internet: connecting to the Internet with OS/2 Warp. Abacus, Grand Rapids, MI, USA, Data Becker edition, 1995. ISBN 1-55755-284-3. vii + 275 pp. LCCN ????

Salus:1995:CNA

[Sal95c] Peter H. (Peter Henry) Salus. Casting the Net: From Arpanet to Internet and Beyond. UNIX and open systems series. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-87674-4 (paperback). xviii + 299 pp. LCCN TK5105.87 .S25 1995.

US$26.85. Map on liner pages.

Salza:1995:CCC

[Sal95d] Giuseppe Salza. Che cosa ci faccio in Internet?: [viaggio nel mondo delle reti: il manifesto della next generation]. Ritmi; 15. Theoria, Roma, Italy, 1995. ISBN 88-241-0412-6. 158 pp. LCCN ????

Salkind:1996:HIW

[Sal96] Neil J. Salkind. Handson Internet for Windows. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1996. ISBN 0-7895-0170-8. xiv + 178 pp. LCCN TK5105.875.I57S35 1996.

Sams:1994:IES

[Sam94] Sams Development Group.

The Internet Explorer’s Software Kit. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30521-6. ???? pp. LCCN ???? US$82.50.

Sams:1995:IU

[Sam95] Sams.net. The Internet Unleashed. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30714-6. 1300 pp. LCCN ???? US$35.00. Includes CD ROM.

Santifaller:1991:TIN

[San91] Michael Santifaller. TCP/IP and NFS: internetworking in a UNIX environment. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54432-6. x + 235 pp. LCCN TK5105.7 .S27 1991. Translated by Stephen S. Wilson from the German edition, “TCP/IP und NFS in Theorie und Praxis”.

Sanghi:1992:EAE

[San92]  Dheeraj Sanghi.  Experi-

mental assessment of endto-end behavior on Internet. Computer science technical report series / University of Maryland; CS-TR-2909 UMIACS-TR-92-62, University of Maryland, College Park, MD, USA, June 1992. 9 + 8 pp.

Sandore:1994:AAI

[San94a] Beth Sandore. AGRICOLA across the Internet: end user needs. Technical report, ????, Urbana, IL, USA, May 6, 1994. iii + 37 + 4 pp.

Santifaller:1994:TIO

[San94b] Michael Santifaller. TCP/ IP and ONC/NFS: internetworking in a UNIX environment. Addison-Wes-ley, Reading, MA, USA, second edition, 1994. ISBN 0-201-42275-1. xii + 288 pp. LCCN TK5105.7 .S2713 1994. US$39.76. Translated by Stephen S. Wilson from the German edition, “TCP/IP und NFS in Theorie und Praxis”.

Sanford:1995:EI

[San95] Clive C. Sanford. Exploring the Internet. Irwin advantage series for computer education. Irwin, Chicago, IL, USA, 1995. ISBN 0-256-18185-3. xi + 335 pp. LCCN TK5105.875.I57S36 1995.

Saroughian:1994:RAC

[Sar94] Andy Saroughian. Restructuring aero 320 course notes for Internet access. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1994. various pp.

Sattler:1995:ITC

[Sat95] Michael Sattler. Internet TV with CU-SeeMe. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 1-57521-006-1. xiii + 300 pp. LCCN TK5105.875.I57S37 1995.

Saustrup:1992:IPA

[Sau92] Arne Saustrup. Internet presentation by the ACCD information systems: presented by arne saustrup: October, 28 1992, 1992. 1 videocassette (VHS) (ca. 120 min.).

Savetz:1993:ISF

[Sav93] Kevin M. Savetz. Internet services frequently asked questions and answers: Version 1.0, 2 July 1993. Technical report, ????, ????, 1993. various pp.

Savetz:1994:YIC

[Sav94] Kevin M. Savetz. Your Internet Consultant: The FAQs of Life Online. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30520-8. xxxvii + 550 pp. LCCN TK5105.875 .I57 S38 1994. US$25.00. Subtitle on cover: Answers to all your Internet questions.

Savetz:1995:MM

[Sav95] Kevin Savetz. The medium is the matrix. Internet World, 6 (4):70–??, April 1, 1995. CO-DEN IERNE8. ISSN 1064-3923.

Savage:1996:FGM

[Sav96] Charles Savage. Fifth Generation Management: Dynamic Teaming, Virtual Enterprising and Knowledge Networking. Butterworth-Heinemann, Boston, MA, USA, second edition, 1996. ISBN 0-7506-9701-6. xxv + 341 pp. LCCN HD31 .S322 1996. US$21.95.

Sawyer:1993:IFM

[Saw93] Carl C. Sawyer. The inverted file — A matter of confidence: Asking reference questions over the Internet. Online, 17(4):8–??, July 1, 1993. CODEN ONLIDN. ISSN 0146-5422.

Strangelove:1994:HAI

[SB94] Michael Strangelove and Aneurin Bosley. How to advertise on the Internet: an introduction to Internetfacilitated marketing and advertising. The Internet Business Journal; Strangelove Internet Enterprises, Ottawa, ON, Canada, October 1994. ISSN 1201-0758. x + 211 pp. LCCN HF6146.I58 S77 1994. US$49.50.

Sadler:1995:UIE

[SB95] Will Sadler and Andrew Berendts. Using Internet email. Que Corporation, Indianapolis, IN, USA, special edition, 1995. ISBN 0-7897-0237-1 (paperback). xxv + 692 pp. LCCN TK5105.73.S24 1995.

Scheller:1994:IWD

[SBGK94] M. Scheller, K.-P. Boden, A. Geenen, and J. Kampermann. Internet: Werkzeuge und Dienste — von Archie bis World Wide Web (English: Internet: Tools and servers. From Archie to World Wide Web). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., July 1994. ISBN 3-540-57968-0. xix + 368 pp. LCCN TK5105.875.I57I59 1994. DM 49. A basic Internet primer for beginners in German. First the Internet’s history is handled in a short introduction and some essential terms are pointed out. Subsequently all of the major tools are explained. The book is finished with ‘the sweets‘, a list of different Internet services Parts of the book are available via the World Wide Web at: http:

                                        //www.ask.uni-karlsruhe.
                                        de/books/inetwd.html.

Salvail:1996:WGI

[SBM96] Leon Salvail, Steven K. Besler, and Sacha Michel Mallais. Webmaster’s guide to Internet connectivity. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-574-7. ???? pp. LCCN TK5105.875.I57S355 1996.

Siewiorek:1982:CSP

[SBN82] Daniel P. Siewiorek, C. Gordon Bell, and Allen Newell. Computer Structures: Principles and Examples. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-057302-6. xvi + 926 pp. LCCN QA76.9.A73

                                        C65.  URL https://
                                        archive.computerhistory.
                                        org/resources/text/bell_
                                        gordon/bell.computer_structures_
                                        principles_and_examples.
                                        1982.102630397.pdf.

Schroeder:1984:EGG

[SBN84] Michael D. Schroeder, Andrew D. Birrell, and Roger M. Needham. Experience with Grapevine: The growth of a distributed system. ACM Transactions on Computer Systems, 2(1):3–23, February 1984.

Schwartz:1995:QII

[SBP^{+}95] Candy Schwartz, Marian Bremer, Robin Peek, Hope N. Tillman, and Clifford A. Lynch. Quality of information on the Internet, 1995. 3 sound cassettes.

Chiou:1986:DAV

[sC86] Ian Yiing shyang Chiou. Design and analysis of a voice/ data Internet transport system. Thesis (Ph.D.), Ohio State University, Columbus, OH, USA, 1986. xi + 221 pp.

Su:1988:IIE

[SC88] Jim Su and Dan Christofferson. ISDN Internet environment and standards analysis. Technical Report AD-A206583, University of Arizona, Computer Engineering Research Laboratory, Tucson, AZ, USA, 1988. ix + 179 pp.

Schmidt:1992:IPT

[SC92] Andrew Schmidt and Roy Harold Campbell. Internet protocol traffic analysis with applications for ATM switch design. Report UIUCDCS-R-92-1735, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, May 1992. 22 pp.

Schmidt:1993:IPT

[SC93] A. Schmidt and R. Campbell. Internet protocol traffic analysis with applications for ATM switch design. Computer Communication Review, 23(2):39–??, April 1, 1993. CODEN CCRED2. ISSN 0146-4833.

Samsel:1994:II

[SC94a] Michael Samsel and Jim Cissel. An introduction to Internet, 1994. 1 videocassette (VHS) (ca. 30 min.).

Schlotman:1994:SUG

[SC94b] Brian Schlotman and Mary J. Clark. SAN users guide to the Internet: a virtual world traveler’s companion, for users of VMS and Ultrix. University of Cincinnati, Center for Information Technology Services, Academic Information Technology Services, Cincinnati, OH, USA, 1994. iii + 87 pp.

Samsel:1995:II

[SC95a] Michael Samsel and Jim Cissel. An introduction to Internet, 1995. 1 videocassette (30 min.).

Sternburg:1995:IGS

[SC95b] Annie Sternburg and Nebraska Library Commission. The Internet: getting started. Technical report, Nebraska Library Commission, Lincoln, NE, USA, 1995. 30 pp.

Sternburg:1995:IPDa

[SC95c] Annie Sternburg and Nebraska Library Commission. Internet providers: dial-up and SLIP/PPP access. Technical report, Nebraska Library Commission, Lincoln, NE, USA, 1995. 5 pp.

Sternburg:1995:IPDb

[SC95d] Annie Sternburg and Nebraska Library Commission. Internet providers: direct leased line connections. Technical report, Nebraska Library Commission, Lincoln, NE, USA, 1995. 5 pp.

Schwartz:1990:SRD

[Sch90] Michael F. Schwartz. Supporting resource discovery among public Internet archives using a spectrum of information quality. Technical report CU-CS-487-90, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, September 1990. 15 pp.

Schwartz:1991:MSCa

[Sch91a] Michael F. Schwartz. A measurement study of changes in service-level reachability in the global TCP/IP Internet: goals, experimental design, implementation, and policy considerations. Technical report CU-CS-551-91, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, October 1991. 5 pp.

Schwartz:1991:MSCb

[Sch91b] Michael F. Schwartz. A measurement study of changes in service-level reachability in the global TCP/IP Internet: Goals, experimental design, implementation, and policy considerations, 1991. 1 data file 5 1/4 in.

Schwartz:1991:RDG

[Sch91c] Michael F. Schwartz. Resource discovery in the global Internet. Technical report CU-CS-555-91, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, November 1991. 18 pp.

Schulzrinne:1992:VCA

[Sch92a] Henning Schulzrinne. Voice communication across the Internet: a network voice terminal. COINS technical report 92–50, University of Massachusetts at Amherst, Dept. of Computer and Information Science, Amherst, MA, USA, June 1992. 34 pp.

Schwartz:1992:CIR

[Sch92b] Michael F. Schwartz. A comparison of Internet resource discovery approaches. Technical report CU-CS-601-92, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, July 1992. 15 pp.

Schuyler:1993:TWI

[Sch93a] Michael Schuyler. Train8@WLN.COM invades Internet. Computers in Libraries, 13(5):30– ??, May 1, 1993. CODEN CPLIE8. ISSN 1041-7915.

Schwartz:1993:IRD

[Sch93b] Michael F. Schwartz. Internet resource discovery at the University of Colorado. Computer, 26(9):25– ??, September 1, 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaefer:1994:CIS

[Sch94] Marc Schaefer. Courte information sur le système UNIX et les réseaux. ALPHANET NF, Battieux 6c 2013 Colombier, CH, 1994. ISBN none. 60 pp. 10 SFr. Beginner introduction to UNIX and introduction to networking tools (uucp, mail, news, www, ftp, etc.).

Schneider:1995:IL

[Sch95a] Karen Schneider. Internet liberian. American Libraries, 26(6):568–??, June 1, 1995. ISSN 0002-9769. Schnittgrund:1995:ICH

[Sch95b] Rita Therese Schnittgrund.

Inner city high school students’ attitudes, beliefs, and intentions toward using the Internet for information retrieval. Doctor of philosophy, administration, supervision, and curriculum development. thesis (Ph.D.), University of Colorado at Denver, Denver, CO, USA, 1995. ix + 140 pp.

Schofield:1995:UIBa

[Sch95c] Sue Schofield. The UK Internet book. Addison-Wes-ley, Reading, MA, USA, 1995. ISBN 0-201-42766-4. xix + 301 pp. LCCN TK5105.875.I57S39 1995.

Schofield:1995:UIBb

[Sch95d] Sue Schofield. The UK Internet book. Addison-Wesley, Reading, MA, USA, revised edition, 1995. ISBN 0-201-87731-7. xxi + 339 pp. LCCN TK5105.875.I57S392 1995.

Schultze:1995:ICE

[Sch95e] Quentin J. (Quentin James) Schultze. Internet for Christians: everything you need to start cruising the Net today. Gospel Communications Network, Gospel Films, Inc., 1995. ISBN 1-55568-155-7. 160 pp. LCCN ????

Schupp:1995:EGI

[Sch95f] Jonathan F. Schupp. Environmental guide to the Internet. Government Institutes, Inc., Rockville, Md., 1995. ISBN 0-86587-449-2. 252 pp. LCCN HC79.E5 S293 1995.

Schuyler:1995:VTL

[Sch95g] Michael Schuyler. The view from the top left corner. Computers in Libraries, 15 (5):25–??, May 1, 1995. CO-DEN CPLIE8. ISSN 1041-7915.

Schneider:1996:IAC

[Sch96a] Karen G. Schneider. The Internet access cookbook: a librarian’s commonsense guide to low-cost connections. Neal-Schuman NetGuide series. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-235-X (paperback). xvi + 316 pp. LCCN Z674.82.I59S36 1996.

Schwartau:1996:CIB

[Sch96b]  Winn Schwartau.  The

complete Internet business toolkit. Van Nostrand Reinhold, New York, NY, USA, 1996. ISBN 0-442-02222-0. xxiii + 262 pp. LCCN HD30.37 .S34 1996.

Shelly:1996:IIC

[SCJ96] Gary B. Shelly, Thomas J.

Cashman, and Kurt A. Jordan. The Internet: introductory concepts and techniques (UNIX). Shelly Cashman series Shelly, Gary B. Shelly Cashman series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1996. ISBN 0-7895-0003-5. viii + 112 pp. LCCN TK5105.875.I57S52 1996.

Starr:1999:UDS

[SCS99] Thomas J. J. Starr, John M.

Cioffi, and Peter Silverman. Understanding Digital Subscriber Line Technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-780545-4. 480 pp. LCCN TK6421.S85

                                        1999. URL http://www.
                                        phptr.com/ptrbooks/ptr_
                                        0137805454.html.

Shoch:1981:MEI

[SCT81] John F. Shoch, Danny Cohen, and Edward A. Taft. Mutual encapsulation of internetwork protocols. Computer Networks: The International Journal of Distributed Informatique, 5(4): 287–301, July 1981. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Stout:1996:ISR

[SD96a] Rick Stout and Morgan Davis. The Internet science, research, and technology yellow pages. Osborne IYP series. Osborne/McGraw-Hill, Berkeley, CA, USA, special edition, 1996. ISBN 0-07-882187-8 (paperback). xxxviii + 362 pp. LCCN TK5105.875.I57 S76 1996.

Summers:1996:IHG

[SD96b] Charles K. Summers and Bryant Dunetz. ISDN: how to get a high-speed connection to the Internet. John Wiley, New York, NY, USA, 1996. ISBN 0-471-13326-4 (paperback). xiii + 346 pp. LCCN TK5105.875.I57 S86 1996.

SRI:1985:DPH

[SDUU85] SRI International, DDN Network Information Center, United States. Dept. of Defense, and United States. Defense Communications Agency. DDN protocol handbook v.2. DARPA Internet protocols. Technical report, National Technical Information Service, Washington, DC, USA, 1985.

Sirkin:1996:SIT

[SE96] Arlene Farber Sirkin and Bill Erbes. Say Internet and they will come successful teambuilding, 1996. 1 videorecording (3 hrs.).

Seagopher:1990:SUI

[Sea90] Seagopher, Inc. Seattle USA an international city destination on the Internet, 1990.

Smith:1994:LNE

[SEDK94] Neil Smith, Andy Ewers, Lorcan Dempsey, and Jack Kessler. Libraries, networks and Europe: a European networking study. Library and information research reports; 101. British Library Research and Development Department, London, UK, 1994. ISBN 0-7123-3295-2 (paperback). ii + 91 pp. LCCN Z674.83.E85 L53 1994. Contents: The European Internet / Andy Ewers — Network resource delivery: a European library perspective / Lorcan Dempsey. — The French case: networked libraries, the Internet and the Minitel / Jack Kessler.

Segell:1995:GIH

[Seg95]  Glen Segell.  Guide to

IHR-info Hypertext Internet Server. IHR guides; no. 4. University of London, Institute of Historical Research, London, UK, 1995. ISBN 1-871348-19-6 (paperback). viii + 94 pp. LCCN ????

Seiter:1994:IMDc

[Sei94a] Charles Seiter. The Internet for Macs for dummies. –For dummies. IDG Books, San Mateo, CA, USA, 1994. ISBN 1-56884-184-1 (paperback). xxiv + 291 pp. LCCN TK5105.875.I57 S45 1994. US$19.95.

Seiter:1994:IMDb

[Sei94b] Charles Seiter. The Internet for Macs for dummies: quick reference. –For dummies. IDG Books, Foster City, CA, USA, 1994. ISBN 1-56884-967-2 (paperback). viii + 184 pp. LCCN ????

Seiter:1994:IMDa

[Sei94c] Charles Seiter. The Internet for Macs for dummies: starter kit. –For dummies. IDG Books, Foster City, CA, USA, 1994. ISBN 1-56884-244-9 (paperback). xxiv + 304 pp. LCCN TK 5105.875 I57 S45 1994. US$19.95.

Seifert:1998:GET

[Sei98] Rich Seifert. Gigabit Ethernet: Technology and Applications for High-Speed LANs.

Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-18553-9. xv + 411 pp. LCCN TK5105.8.E83S45 1998. US$49.95.

Severance:1995:SVF

[Sev95] C. Severance. Standards:

The value of the formal standards process. Computer, 28 (8):82–83, August 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sferrino:1982:MBM

[Sfe82] V. J. (Vincent John) Sferrino.

A multiport buffer memory for an Internet packet voice/ data gateway. Technical report (Lincoln Laboratory); 612 ESD/TR-82/056, Massachusetts Institute of Technology, Lincoln Laboratory, Lexington, MA, USA, 1982. viii + 59 pp.

Singh:1995:CNI

[SFL95] Paul Singh, Rick Fairweather, and Dan Ladermann. Connecting Net-Ware to the Internet. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-398-1. xiii + 294 pp. LCCN TK5105.9 .S46 1995. US$32.00.

Saustrup:1993:IUT

[SG93a] Arne Saustrup and Jim Garrison. Internet the use of Telnet and FTP, 1993. 1 videocassette (VHS) (ca. 60 min.).

Smith:1993:LGG

[SG93b] Richard J. (Richard James) Smith and Jim Gerland. Let’s go gopherin’ E-mail based course sessions, 1993.

Smith:1993:NI

[SG93c] Richard J. (Richard James) Smith and Mark Gibbs. Navigating the Internet. SAMS Publishing, Indianapolis, IN, USA, 1993. ISBN 0-672-30362-0 (paperback). xxiii + 500 pp. LCCN TK5105.875.I57S65 1993.

Smith:1994:NID

[SG94a] Richard Smith and Mark Gibbs. Navigating the Internet: Deluxe Edition. Howard W. Sams, Indianapolis, IN 46268, USA, April 1994. ISBN 0-672-30485-6. xxvi + 640 pp. LCCN TK5105.875.I57 S65 1994. US$29.95. Nearly identical to Navigating the Internet (above) except it includes a disk with Internet Software for Microsoft Windows and a few extra chapters: “A Better Window on the Web: Mosaic”, “Using the Internet for Business”, as well as versions of the online references Gopher Jewels and PDIAL. I suppose it’s worth the extra five bucks for this book versus the vanilla Navigating the Internet, especially if you use Windows and need the software.

Smith:1994:IUM

[SG94b] Richard J. (Richard James)

Smith and Mark Gibbs. Intonet ui modun kot: nugu na i chaek ul ilko namyon intonet sayong si piryohan modun kot ul alsu itta = Internet: international bestseller. Inpo Buk, Soul Tukpyolsi, Korea, chopan. edition, 1994. ISBN ???? 15 + 617 pp. LCCN ????

Smith:1994:NI

[SG94c] Richard J. (Richard James)

Smith and Mark Gibbs. Navigating the Internet. SAMS Publishing, Indianapolis, IN, USA, deluxe edition, 1994. ISBN 0-672-30485-6. xxvi + 640 pp. LCCN TK5105.875.I57 S65 1994. US$29.95.

Siemann:1995:WWI

[SG95] Robert Siemann and Shirley Ann Gartmann. Welcome to the world of the Internet: an Internet simulator for PC-DOS and Macintosh, UNIX (version 2.0). Technical report, ????, ????, 1995. 62 pp.

Schwartz:1998:SSS

[SG98] Alan Schwartz and Simson Garfinkel. Stopping Spam: Stamping Out Unwanted Email and News Postings. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-388-X. x + 191 pp. LCCN TK5105.73 .S38 1998. US$19.95. URL

                                        http://www.oreilly.com/
                                        catalog/spam.

Smith:1995:NI

[SGM95] Richard J. (Richard James) Smith, Mark Gibbs, and Paul McFedries. Navigating the Internet. Sams.net Pub., Indianapolis, IN, USA, third edition, 1995. ISBN 0-672-30718-9. xxii + 501 pp. LCCN ????

Siyan:1995:IFN

[SH95a] Karanjit Siyan and Chris Hare. Internet Firewalls and Network Security. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-437-6. xv + 410 pp. LCCN TK5105.875.I57 H36 1995. US$35.00, CDN$47.95, UK£32.49. “Master the complexities of network security”–Cover.

Stites:1995:IAF

[SH95b] Barbara J. Stites and Samantha K. Hastings. Internet access for Florida’s public libraries: a project report. Technical report, Tampa Bay Library Consortium, Tampa, FL, USA, 1995. vii + 72 pp.

Shah:1991:IEW

[Sha91a]  Devang S. Shah.  Inter-

net experiments with window mechanisms for replay detection. Thesis (M.S.), University of Maryland at College Park, College Park, MD, USA, 1991. vii + 115 pp.

Shaikh:1991:RTD

[Sha91b] Asad Ali Shaikh. Real time data acquisition protocol for Internet applications: a thesis. Thesis (M.S.), Department of Electrical and Computer Engineering, Clarkson University, Potsdam, NY, USA, 1991. ix + 58 pp.

Sha:1995:IRC

[Sha95a] Vianne Tang Sha. Internet resources for cataloging, 1995.

Sharma:1995:FI

[Sha95b] Shashank Sharma. Fundamentals of the Internet. Masters of science in computer information systems management. thesis (M.S.), Capstone Project Committee and Computer Information and Systems Dept., College of Business, Ferris State University, Big Rapids, MI, USA, June 1995. viii + 58 pp.

Stey:1995:IHR

[SHB95] John A. Stey, Bill Hengstenberg, and Frank Barton. Internet healthcare resources and information management, 1995. 1 videocassette (10 min.).

Shetler:1977:IIE

[She77] Stanwyn G. Shetler. Internet international environmental resources network: teacher guide. Strategic Environmental Planning, Threshold, Inc., Washington, DC, USA?, second edition, 1977. ISBN ???? various pp. LCCN ????

Sheltzer:1985:NTI

[She85] Alan Brian Sheltzer. Network transparency in an Internet[work] environment. Thesis (Ph.D.), UCLA, Computer Science Dept., Los Angeles, CA, USA, 1985. xi + 129 pp. Report CSD-850028.

Sherman:1986:NPM

[She86] Mark Steven Sherman. A network package for the Macintosh using the DoD Internet protocols. Technical report PCS-TR 86-124, Dartmouth College, Dept. of Mathematics and Computer Science, Hanover, NH, USA, 1986. 45 pp.

Shea:1994:N

[She94] Virginia Shea. Netiquette. Albion Books, San Francisco, CA, USA, 1994. ISBN 0-9637025-1-3. 154 pp. A unique book on network etiquette (the dos and DON’T s of online communications) …. With lots of valuable tips for beginners …. This slim book does an excellent job of helping those new to Cyberspace learn the basic tenants of network etiquette, or “netiquette”. It’s easy to read and loaded with interesting real-life examples. Chapters cover netiquette of email and of discussion groups as well as providing excellent netiquette guidelines for business, home and school use. Particularly useful sections cover Electronic Style (looking good online, tone of voice and signature files) and the “core rules” of the online world. One chapter covers copyright issues and another the etiquette of e-mail privacy. There’s even a tiny chapter on “Love and Sex in Cyberspace”. This book should be required reading for anyone collecting their Internet Learner’s Permit. (You can retrieve a brief excerpt from Netiquette: “The Core Rules of Netiquette” by sending e-mail to netiquetterequest@albion.com with the words “archive send core” in the subject line.).

Shefski:1995:III

[She95a] William J. Shefski. Interactive Internet: the insider’s guide to MUDs, MOOs and IRC. Prima Publishing, Rocklin, CA, USA, 1995. ISBN 1-55958-748-2. xvi + 202 pp. LCCN TK5105.875.I57 S518 1995.

Shenoy:1995:HEN

[She95b] Uday V. Shenoy. Heat Exchanger Network Synthesis: Process Optimization by Energy and Resource Analysis. Gulf Publishing, Houston, TX, USA, 1995. ISBN 0-88415-391-6. xviii + 642 pp. LCCN ???? US$105.00.

Shirky:1994:IE

[Shi94] Clay Shirky. The Internet by E-mail. Ziff-Davis Press, Emeryville, CA, USA, 1994. ISBN 1-56276-240-0 (paperback). xv + 220 pp. LCCN TK 5105.875 I57S55 1994. US$19.95.

Shields:1996:CIV

[Shi96] Rob Shields. Cultures of Internet: virtual spaces, real histories, living bodies. Sage Publications, London, UK, 1996. ISBN 0-8039-7518-X, 0-8039-7519-8 (paperback). viii + 196 pp. LCCN ????

Snyder:1997:TUS

[SHN97] Garth Snyder, Trent R. Hein, and Evi Nemeth. Tools for UNIX system administrators. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-665431-2. LCCN QA76.76.O63 T45 1997. Companion CD-ROM to UNIX system administration handbook, 2nd ed..

Shunk:1993:ERI

[Shu93] Dan L. Shunk. An example of a regional Internet that will amplify the area’s global competitiveness positioning. Creative manufacturing engineering program technical paper MS93-228, Society of Manufacturing Engineers, Dearborn, MI, USA, 1993. 13 pp.

Shurety:1999:BNC

[Shu99]  Samantha Shurety.  E-

Business with Net.Commerce. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-083808-X. xxx + 718 pp. LCCN HF5548.32 .S5

1998. URL http://www.
phptr.com/ptrbooks/ptr_
013083808X.html. Includes

CD-ROM.

SIDivision:1993:ISM

[SI 93] SI Division. Internet shouyouka ni mukete (CIX) (Feasibility Reserch for Commercial Internet in Japan). Toppan Publishing, Tokyo, Japan, 1993. ISBN 4-8101-8907-4. 142 pp. Yen 1600. This is the first and only book on commercial Internet service in Japan. Published as reserch report, its purpose is to clarify Internet status in the United States and business feasibility study on CIX in Japan. Coverage includes: What is Internet and its historical story in the US, Structure of commercial service in Internet, Image of commercial Internet in Japan. Report covers the feasibility of Internet within various regulated Japanese environments. It is must-read book for companies in Japan wishing enter the Internet business.

Silverberg:1995:AGA

[Sil95] Ira Silverberg. AirLinks: a guide to air pollution information on the Internet. Technical report, University of Tennessee, Knoxville, TN, USA, Spring 1995. ix + 106 + 96 pp.

SMP:1996:II

[Sil96] Silicon Mountain Publishing. Introduction to the Internet, 1996. 1 videocassette (VHS) (29 min.) computer disk (3 1/2 in.).

Simmonds:1993:PFE

[Sim93] Curtis Simmonds. Painless file extraction: The A(rc) — Z(oo) of Internet archive formats. Online, 17(6):60–??, November 1, 1993. CODEN ONLIDN. ISSN 0146-5422.

Simmons:1995:IPI

[Sim95a] Edlyn Simmons. Intellectual property and the Internet. Searcher: the magazine for database professionals, 3(1): 38–??, January 1, 1995. CO-DEN SMDPE8. ISSN 1070-4795.

Simpson:1995:ILM

[Sim95b] Carol Mann Simpson. Internet for library media specialists. Professional growth series. Linworth Pub., Worthington, OH, USA, 1995. ISBN 0-938865-39-0. xi + 144 + 18 pp. LCCN Z674.82.I59S56 1995.

Singh:1981:DRS

[Sin81] Vineet Singh. The design of a routing service for campuswide Internet transport. Thesis (M.S.), M.I.T., Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1981. 109 pp. Supervised by Jerry Saltzer.

Singh:1999:IDW

[Sin99a]  Harinder S. Singh.  In-

teractive data warehousing via the Web. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080371-5. 480 pp. LCCN QA76.9.D37S563 1999. US$49.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130803715.html.

Singh:1999:PDM

[Sin99b] Harinder S. Singh. Progressing to Distributed Multiprocessing. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-095683-X. ???? pp. LCCN QA76.9.D5S54435

1999. URL http://www.
phptr.com/ptrbooks/ptr_
013095683X.html.

Siyan:1994:NTG

[Siy94] Karenjit Siyan. NetWare training guide: NetWare TCP/IP and NetWare NFS. New Riders Publishing, Carmel, IN, USA, 1994. ISBN 1-56205-409-0. xiv + 1050 pp. LCCN TK5105.5 .S55 1994. Includes CD-ROM.

Sun:1998:ACC

[SJZ^{+}98] Chengzheng Sun, Xiaohua Jia, Yanchun Zhang, Yun Yang, and David Chen. Achieving convergence, causality preservation, and intention preservation in realtime cooperative editing systems. ACM Transactions on Computer-Human Interaction, 5(1):63–108, March 1998. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/tochi/1998-5-1/
                                        p63-sun/.

Schoenleber:1995:IH

[SK95] Claus Schoenleber and Cornelius Keck. Internet Handbuch. Franzis’-Verlag, Poing bei Muenchen, Germany, 1995. ISBN 3-7723-7292-9. 315 pp. DM 78.

Sugiura:1998:ISA

[SK98] Atsushi Sugiura and Yoshiyuki Koseki. Internet Scrapbook: automating Web browsing tasks by programming-by-demonstration. Computer Networks and ISDN Systems, 30(1–7):688–690, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/comnet/sub/1998/
                                        30/1-7/1886.pdf.

Schill:1996:IAE

[SKB96] Alexander Schill, Sabine Kühn, and Frank Breiter. Internetworking over ATM: Experiences with IP/IPng and RSVP. Computer Networks and ISDN Systems, 28(14):1915–1927, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1996/
28/14/1639.pdf.

Skyrme:1999:KNC

[Sky99] David J. Skyrme. Knowledge Networking: Creating the Collaborative Enterprise. Butterworth-Heinemann, Boston, MA, USA, 1999. ISBN 0-7506-3976-8. xvii + 311 pp. LCCN HF54.5 .S55 1999. US$29.95.

StGeorge:1991:IAL

[SL91] Art St. George and Ron Larsen. Internet: accessible library catalogs and databases. Technical report, University of New Mexico, Albuquerque, NM, USA, January 28, 1991. 67 pp.

StGeorge:1992:IAL

[SL92] Art St. George and Ron Larsen. Internet: accessible library catalogs and databases. Technical report, University of Maryland, College Park, MD, USA, January 6, 1992. vii + 120 pp.

Shipley:1995:EPR

[SL95] Buddy Shipley and Williams F. Lyons. Ethernet Pocket Reference Guide. Shipley Consulting International, ????, 1995. 85 pp. LCCN ???? US$24.95.

Smith:1994:EIB

[SLB94] Tracy Smith, Elizabeth Lane Lawley, and Karen Becker. Exploring Internet 2 back to basics. D and F Associates, DeKalb, IL, USA, 1994. ISBN ???? LCCN ???? 1 videocassette (120 min.).

Sha:1994:ILR

[SLJ94] Vianne Tang Sha and Lyonette Louis-Jacques. Internet as a library-wide resource tools for acquisitions, serials, cataloging, ILL and reference, and related legal and technical problems, 1994. 1 sound cassette.

Sloan:1993:RSN

[Slo93] Bernie Sloan. Resource sharing and the networks. Internet research, 3(2):2–??, Summer 1993. CODEN IRESEF. ISSN 1066-2243.

Sloman:1994:NDS

[Slo94] Morris Sloman, editor. Network and Distributed Systems Management. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-62745-0. xix + 666 pp. LCCN QA76.9.D5 N53 1994. US$50.50.

Saint-Leger:1994:CCP

[SLP94] Olivier Saint-Leger and Thierry Pigot. Clés de contact pour Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 2-7361-1445-0. 274 pp. LCCN ????

Saint-Leger:1995:CCP

[SLP95] Olivier Saint-Leger and Thierry Pigot. Clés de contact pour Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, nouv. edition, 1995. ISBN 2-7361-1553-8. xiv + 298 pp. LCCN ????

StGeorge:1991:ILC

[SLR91] Art St. George, Ron Larsen, and Carlos Robles. Internetaccessible library catalogs and databases. Technical report, University of Maryland, College Park, MD, USA, April 16, 1991. various pp.

Sluman:1995:IME

[Slu95] Roelf Sluman. Internet made easy: a guide to the Internet for CompuServe members. International Thompson Computer Press, London, UK, 1995. ISBN 1-85032-240-6. viii + 255 pp. LCCN ????

Singh:1993:ESL

[SM93] Jagtar Singh and Jack Meadows. Electronic serials for library and information specialists on Internet. ASLIB Proceedings, 45(9):

234–??, September 1, 1993. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Anonymous:1995:SBI

[Sma95] The small business Internet newsletter, page various, 1995. ISSN 1081-0560. Internet and More, Silver Spring, MD, USA.

Stone-Martin:1994:RHW

[SMB94] Martha Stone-Martin and Laura Breeden. 51 Reasons: How We Use the Internet and What it Says About the Information Superhighway. FAR-NET, Lexington, MA, USA, 1994. ISBN None. 124 pp. US$19.95. This is a unique book exploring how people and organizations in all 50 United States (plus the capitol) are using the Internet as the first stepping stone toward the Clinton administration’s “information superhighway”. Written as an oral narrative from the mouths of the people making these projects work, the book covers uses of the Internet in libraries, K-12 education, agriculture, economic development, government, health care and higher education. Although brief, 51 Reasons gives novel ideas for the use of the Internet and provides contact information for each of the individuals/projects mentioned.

You can get the text of the book online via gopher (gopher.cni.org:70/11/cniftp/miscdocs/farnet) and FTP (ftp.cni.org:/CNI/documents/farnet/storiesindex).

Smedinghoff:1996:OLS

[Sme96] Thomas J. Smedinghoff. Online law: the SPA’s legal guide to doing business on the Internet. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48980-

5. xxiv + 544 pp. LCCN KF390.5.C6O55 1996.

Smith:1993:FRP

[Smi93a] Philip Smith. Frame Relay: Principles and Applications. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62400-1. xii + 268 pp. LCCN TK5105.5.S627 1993. US$41.95.

Smith:1993:BGIa

[Smi93b] Una Smith. A biologist’s guide to Internet resources. Technical report, Dept. of Biology, Yale University, New Haven, CT, USA, 1993. various pp.

Smith:1993:BGIb

[Smi93c] Una Smith. A biologist’s guide to Internet resources: version 1.4, 26 may 1993. Technical report, Department of Biology. Yale University, New Haven, CT, USA, 1993. 36 pp.

Smith:1993:BGIc

[Smi93d] Una R. Smith. A Biologist’s Guide to Internet Resources. Yale University, Department of Biology, Osborn Memorial Laboratories, PO Box 6666, New Haven, CT, 06511-8155 USA, 1.7 edition, 10 November 1993. 45 pp. This 45-page document is available via gopher, anonymous FTP and e-mail from various archives. For a free copy via e-mail, send the text send

pub/usenet/sci.answers/
biology/guide/* to the e-
mail address mail-server@
rtfm.mit.edu.

Smith:1994:AIG

[Smi94a] John W. T. Smith. Accessing the Internet: a guide for the UK and Ireland. International Thomson, London, UK, 1994. ISBN 1-85032-147-7 (paperback). xii + 164 pp. LCCN ????

Smith:1994:IRP

[Smi94b] Lisa DiIorio Smith. Internet resources for pharmaceutical reference. Thesis (M.S.), University of North Carolina at Chapel Hill, Department of Library Science, Chapel Hill, NC, USA, 1994. 36 pp.

Smith:1995:CAA

[Smi95a] Brian Carl Smith. A content analysis of American online newspapers distributed on the Internet. Thesis (M.S.), Oklahoma State University, Stillwater, OK, USA, 1995. vi + 62 pp.

Smith:1995:IB

[Smi95b] Dana Smith. The Internet and beyond, 1995. 1 videocassette (1 hr. + 49 min.).

Smith:1997:IC

[Smi97] Richard E. Smith. Internet Cryptography. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-92480-3. xx + 356 pp. LCCN TK5102.94.S65 1997. US$27.95.

Savage:1995:GYN

[SMLD95] B. Savage, Myles McGregor-Lowndes, and Lester Deakin. Getting your nonprofit organisation started on the Internet: software, hardware, service providors and other issues. Program on Nonprofit Corporations Working paper series, no. 59 1037-1516. Program on Nonprofit Corporations, Queensland University of Technology, Brisbane, Qld., August 1995. ISBN 1-86435-007-5. 29 pp. LCCN ????

Swindell:1996:IMB

[SMM96] Simon R. Swindell, R. Russell Miller, and Garry S. A. Myers. Internet for the molecular biologist. Current innovations in molecular biology; v. 3. Horizon Scientific, Wymondham, Leicestershire, UK, 1996. ISBN 1-898486-02-6 (paperback). 187 pp. LCCN IN PROCESS.

Smythe:1995:IDR

[Smy95] Colin Smythe. Internetworking: Designing the Right Architectures. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-56536-6. xxiv + 473 pp. LCCN TK5105.7 .S58 1995. US$37.75.

Sneed:1995:BII

[Sne95a] Stephen F. Sneed. Browsing Internet information resources using a gopher client controlled by adaptive user profiles. Thesis (M.S.), University of Alabama at Birmingham, Birmingham, AL, USA, 1995. vi + 55 pp.

Snell:1995:CAI

[Sne95b] Ned Snell. Curious about the Internet? SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30459-7 (paperback). xv + 247 pp. LCCN TK5105.875.I57S66 1995.

Snell:1995:NIW

[Sne95c] Ned Snell. Navigating the Internet with Windows 95. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 0-672-30765-0. xx + 404 pp. LCCN QA76.76.O63S5931995.

Snyder:1995:IRR

[Sny95]  Joel M. Snyder.  Inside

risks: Research on the Internet. Communications of the ACM, 38(8):130, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
208377.html.

Snyder:1996:MCE

[Sny96] Joel Snyder. Money changes everything. Internet World, 7(1):44–??, January 1, 1996. CODEN IERNE8. ISSN 1064-3923.

Stevenson:1994:W

[SO94] Robert Louis Stevenson and Lloyd Osbourne. The wrecker, 1994.

IS:1992:Na

[Soc92a] News, 1992. ISSN 1060-7811.

Internet Society, Reston, VA, USA.

IS:1992:Nb

[Soc92b] News, 1992. ISSN 1060-7803. The Society, Reston, VA, USA.

IS:1995:IFI

[Soc95a] Internet Society. The ISOC forum international electronic publication of the Internet society, 1995.

IS:1995:OIP

[Soc95b] OnTheInternet: an international publication of the Internet Society, page various, 1995. ISSN 1081-3969. The Society, Hopewell, NJ, USA.

Sohier:1994:IGS

[Soh94]  Danny J. Sohier.  IN-

TERNET, guide de survie de l’internaute. Collection Réseau simplifié. Editions Logiques, Montréal, PQ, Canada, 1994. ISBN 2-89381-245-7. 181 pp. LCCN ????

Sohier:1995:I

[Soh95a] Danny J. Sohier. Internet. L’Informatique nouvelle vague. Editions logiques, Montréal, PQ, Canada, 1995. ISBN 2-89381-203-1. 126 pp. LCCN ????

Sohier:1995:IGD

[Soh95b] Danny J. Sohier. INTER-NET, guide d’exploration de l’internaute. Collection Réseau simplifié. Editions Logiques, Montréal, PQ, Canada, 1995. ISBN 2-89381-253-8. 232 pp. LCCN ????

Sohier:1996:GLG

[Soh96] Danny J. Sohier. Guide de l’internaute 1996: guide de survie, guide d’exploration. Collection réseau simplifié. Ed. de l’homme; Ed. logiques, Montréal, PQ, Canada, nouv. rev. corr. et augm. edition, 1996. ISBN 2-7619-1321-3. 477 pp. LCCN ????

Soman:1994:VCI

[Som94]  Sadhana Soman.  Video

conferencing over the Internet: an experimental study. Thesis (M.S.), University of South Carolina, Columbia, SC, USA, 1994. vi + 34 pp.

Sondhi:1992:TIT

[Son92] Amit Sondhi. TP4/IP: transition and coexistence strategy from Internet to DSI. Thesis (M.S.), Utah State University. Department of Computer Science, Logan, UT, USA, 1992. vii + 38 pp.

Souter:1993:ISB

[Sou93] Bill Souter. Internet Security: a Brief Survey. ??, ??, 1993. ISBN 0-9625476-2-X. US$8.95.

Schwartz:1994:AIG

[SP94a] Michael F. Schwartz and Calton Pu. Applying an information gathering architecture to Netfind: a white pages tool for a changing and growing Internet. IEEE/ ACM Transactions on Networking, 2(5):426–439, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1994-2-5/p426-schwartz/.

Shrikumar:1994:TLE

[SP94b] H. Shrikumar and Rehmi Post. Thinternet: life at the end of a tether. Computer Networks and ISDN Systems, 27(3):375–385, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1994&volume=27&
issue=3&aid=1357.

Seiter:1995:IMD

[SP95a] Charles Seiter and David Pogue. The Internet for Macs for dummies. –For dummies. IDG Books, San Mateo, CA, USA, second edition, 1995. ISBN 1-56884-371-2 (paperback). xxii + 316 pp. LCCN ????

Sosinsky:1995:AQT

[SP95b] Barrie Sosinsky and Elisabeth Parker. Acrobat Quick Tour. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-255-0. 150 pp. LCCN Z286.E43 S658 1995. US$14.95.

Spafford:1988:IWP

[Spa88] Eugene H. Spafford. The Internet worm program: an analysis. Purdue technical report; CSD-TR-823 CSD-TR-823, Department of Computer Sciences, Purdue University, West Lafayette, IN, USA, December 8, 1988. 40 pp.

Spafford:1989:IWC

[Spa89a] Eugene H. Spafford. The Internet worm: Crisis and aftermath. Communications of the ACM, 32(6):678–??, June 1, 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spafford:1989:IWP

[Spa89b] Eugene H. Spafford. The Internet worm program: An analysis. Computer Communication Review, 19(1):17–??, January 1, 1989. CODEN CCRED2. ISSN 0146-4833.

Spector:1987:CDT

[Spe87] Alfred Z. Spector. Camelot:

a distributed transaction facility for Mach and the Internet — an interim report. Research paper. CMU-CS-87-129, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 36 pp.

SLABC:1993:PVS

[Spe93] Special Library Association Boston Chapter. Plundering the virtual shelves: the Internet for special librarians, 1993. 4 videocassettes and 1 folder of accompanying handouts.

Spilbergs:1992:PVP

[Spi92] A. Spilbergs. Projektu vadisanas problemas un tendences parejas perioda uz tirgus ekonomiku: starptautiska simpozija Internet’ 91: referatu tezes. Technical report, Latvijas universitate, Riga, Latvia, 1992. 118 pp. English, Latvian, and Russian.

Spry:1994:IBC

[SPR94] SPRY, Inc. Internet In A Box: The Complete Internet Solution. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1994. ISBN ???? ???? pp. LCCN ???? US$149.00.

SPRY:1995:IB

[SPR95] SPRY, Inc. Internet in a box, 1995.

Segev:1998:ISC

[SPR98] Arie Segev, Jaana Porra, and Malu Roldan. Internet security and the case of Bank of America. Communications of the ACM, 41(10): 81–87, October 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-10/
p81-segev/.

Schwartz:1993:CGI

[SQ93] Michael F. Schwartz and John S. Quarterman. The changing global Internet service infrastructure. Internet research, 3(3):8–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

Shah:1995:PMI

[SR95a] Rawn Shah and Jim Romine. Playing MUDS on the Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11633-5. viii + 328 pp. LCCN GV1469.15.S52 1995.

Smoot:1995:AOG

[SR95b] Kelley Smoot and Edna O. F.

Reid. Asian oil and gas Internet resources, 1995.

Shah:1996:PWI

[SR96] Rawn Shah and Jim Romine.

Playing wargames on the Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-11634-3 (paperback). xi + 308 pp. LCCN U313.S53 1996.

Schulzrinne:1999:ITA

[SR99] Henning Schulzrinne and Jonathan Rosenberg. Internet Telephony: architecture and protocols — an IETF perspective. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):237–255, February 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1999/
31/3/2071.pdf.

SRI:1982:IPI

[SRI82a] SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028. Internet protocol implementation guide, 1982. 151 pp.

SRI:1982:IPT

[SRI82b] SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028. Internet Protocol Transition Workbook, March 1982. 521 pp. AD-A153607.

SRINIC:1982:IPT

[SRI82c] SRI International. Network Information Center. Internet protocol transition workbook. Network Information Center, SRI International, Menlo Park, CA, USA, 1982. 521 pp.

Sriram:1993:SRH

[Sri93] K. B. Sriram. A study of the reliability of hosts on the Internet. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1993. vii + 48 pp.

Savetz:1996:MMT

[SRL96] Kevin Savetz, Neil Randall, and Yves Lepage. MBone: multicasting tomorrow’s Internet. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-723-8 (paperback). 234 pp. LCCN QA76.76.I59S28 1996.

Seltzer:1997:ASR

[SRR97] Richard Seltzer, Deborah S. Ray, and Eric J. Ray. The AltaVista Search Revolution: How to Find Anything on the Internet. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882235-1. xxii + 274 pp. LCCN TK5105.875 .I57 S44 1997. US$16.99.

Steen:1995:TIP

[SRS^{+}95] Douglas R. Steen, Mark R.

Roddy, Derek Sheffield, Michael Bryan Stout, and D. J. Hura. Teaching with the Internet: putting teachers before technology. Resolution Business Press, Bellevue, WA, USA, 1995. ISBN 0-945264-19-4. 334 pp. LCCN ????

Sparks:1995:UIW

[SRV95] Kathryn Sparks, Virgil Richie, and ViaGrafix. Using Internet with Windows 95, 1995. 1 videocassette (49 min.) (3 1/2 in.) + 1 course index.

Sachs:1993:TLS

[SS93] David Sachs and Henry Stair.

Telix lite software, 1993. 1 computer disk.

Sachs:1994:HMT

[SS94a] David Sachs and Henry Stair.

Hands-on Mosaic: a Tutorial for Windows Users. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, December 1994. ISBN 0-13-172321-9. 358 pp. LCCN TK5105.875.I57 S32 1994. US$29.95. This book/disk package contains nine ready set go sessions which take the user through TCP/IP (which comes on the disk), downloading and installing Mosaic, and navigating the Web. “Hands-On Mosaic” is perfect for the novice, packed with scripts and sample screens. Not recommended for classroom use.

Sachs:1994:HIB

[SS94b] David Sachs and Henry H. Stair. Hands-on Internet— A Beginning Guide for PC Users. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-056392-7. xiv + 275 pp. LCCN TK5105.875.I57 S33 1994. US$27.95 (includes IBM PC diskette). System requirements for computer disk: IBM-compatible PC; DOS; Hayes-compatible modem; hard disk.

Salaun:1994:IKC

[SS94c] Christophe Salaun and Jean-Arthur Silve. Internet kit de connexion. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1994. ISBN 2-7361-1454-X. 35 pp. LCCN ????

Sachs:1995:IIW

[SS95a] David Sachs and Henry H. Stair. Instant Internet With Websurfer. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-210675-2. xxi + 217 pp. LCCN TK5105.875.I57 S34 1995. US$24.95. “Internet Chameleon software inside”– Cover.

Schumann:1995:EGI

[SS95b] Roland W. Schumann and Jon F. Schupp. Energy guide to the Internet. Utility Data Institute, Washington, DC, USA, August 1995. ISBN 1-56760-064-6. various pp. LCCN HD9502.A2S399 1995. UDI-2475-95.

Smith:1995:DTE

[SS95c] Brandon D. Smith and Michael L. Scott. Development of a technology education guide to the Internet. Thesis (M.A.), Ohio State University, Columbus, OH, USA, 1995. viii + 75 pp. Advisor: Michael L. Scott, College of Education.

Sobol:1995:IBG

[SS95d] Robert E. Sobol and Kevin J.

Scanlon. Internet book of gene therapy: cancer therapeutics. Appleton and Lange, Stamford, CT, USA, 1995. ISBN 0-8385-3101-6. ???? pp. LCCN RC271.G45I58 1995.

Stang:1995:IST

[SS95e] David J. Stang and Norman Data Defense Systems. Internet security threats and firewalls. Norman technical report 6, Norman Data Defense Systems, Fairfax, VA, USA, 1995. 28 pp.

Sutton:1995:IC

[SS95f] Rory Sutton and Tony T. Santana. Internet class, 1995. 1 videocassette (76 min.).

Steadman:1996:PIS

[SS96] Carl Steadman and Jason Snell. Providing Internet services via the Mac OS. Addison-Wesley Developers Press, Reading, MA, USA, 1996. ISBN 0-201-48998-8. xv + 409 pp. LCCN TK5105.875.I57S75 1996.

Solomonides:1999:REI

[SS99] C. Solomonides and M. Searle. Relevance of existing intelligent network infrastructure to the Internet. Lecture Notes in Computer Science, 1597:459–468, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strudwick:1995:IP

[SSA95] Karen Strudwick, John Spilker, and Jay Arney. Internet for parents. Resolution Business Press, Bellevue, WA, USA, 1995. ISBN 0-945264-17-8 (paperback). 352 (or 379??) pp. LCCN ????

Schnyder:1996:IBD

[SSC96] Sandy Eddy Schnyder, Michael M. Swertfager, and Margaret M. E. Cusik. The Internet business-to-business directory. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1751-1. xv + 689 pp. LCCN HF54.56 .E44 1996.

Seiler:1998:EIA

[SSI^{+}98] R. J. Seiler, A. M. Seiler, J. M. Ireland, A. M. Guy, and K. N. Woodward. Enhancing Internet access for people with disabilities. Computer Networks and ISDN Systems, 30(1–7):739–741, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/1-7/1914.pdf.

Schwartz:1989:ESP

[ST89] Michael F. Schwartz and Panagiotis G. Tsirigotis. Exploiting semantics to provide Internet white pages without global cooperation. Technical report DCR-8420944, CU-CS-444-89, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, October 1989. 9 pp.

CST:1995:TCI

[ST95] Casper Star-Tribune. Trib.com the Internet newspaper, 1995.

Stanford:1988:SUP

[Sta88] Stanford University. Networking and Communication Systems. The Stanford University personal computer/ Internet protocol: SU-PC/ IP. Technical report, Networking and Communication Systems, Stanford University, Stanford, CA, USA, 1988. A communication program for IBM personal computers: PC, PC/XT, PC/AT, PS/2 50, 60, and 80.

Stallings:1989:I

[Sta89] William Stallings. ISDN: An Introduction. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1989. ISBN 0-02-415471-7. xii + 418 pp. LCCN TK5103.7.S73 1989.

Stallings:1990:LN

[Sta90]  William Stallings.  Local

Networks. Collier Books, Macmillan Publishing Company, New York, NY, USA, third edition, 1990. ISBN 0-02-415531-4. xvi + 510 pp. LCCN TK5105.7 .S78 1990.

Stallings:1992:IBI

[Sta92] William Stallings. ISDN and Broadband ISDN. Collier Books, Macmillan Publishing Company, New York, NY, USA, second edition, 1992. ISBN 0-02-415475-X. xvii + 633 pp. LCCN TK5103.7.S73 1992.

Stallings:1993:LMA

[Sta93a] William Stallings. Local and Metropolitan Area Networks. Macmillan Publishing Company, New York, NY, USA, fourth edition, 1993. ISBN 0-02-415465-2. xvi + 550 pp. LCCN TK5105.7 .S77 1993.

Stallings:1993:NS

[Sta93b] William Stallings. Networking Standards: a Guide to OSI, ISDN, LAN, and MAN Standards. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56357-6. xv + 646 pp. LCCN TK5105.5 .S73 1993.

Stallings:1993:SSC

[Sta93c] William Stallings. SNMP, SNMPv2, and CMIP: the Practical Guide to Network-Management Standards. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63331-0. xvii + 625 pp. LCCN 92-38220, TK5105.5 .S763 1993. US$49.50.

Stamper:1994:LAN

[Sta94a]  David A. Stamper.  Lo-

cal Area Networks. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1994. ISBN 0-8053-7724-7. xi + 401 pp. LCCN TK5105.7.S784 1994.

Stankovic:1994:ISI

[Sta94b] Nikola D. Stankovic. Internet for Southeast: Internet, the information highway, for students, faculty and staff of Southeast Missouri State University. Technical report, Nik Stankovic, Cape Girardeau, MO, USA, 1994. 56 + 27 pp.

Stallings:1995:IBI

[Sta95a] William Stallings. ISDN and Broadband ISDN with Frame Relay and ATM. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-02-415513-6. 581 pp. LCCN TK5103.75 .S73 1994. US$56.00.

Stallings:1995:MOI

[Sta95b] William Stallings. Mecklermedia’s official Internet world Internet security handbook. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-700-9. xxx + 288 pp. LCCN TK5105.59.S72 1995.

Stallings:1995:NIS

[Sta95c] William Stallings. Network and Internetwork Security. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-02-415483-0. xiii + 462 pp. LCCN TK5105.5 .S728 1995. US$55.00. Provides an overview of security threats mechanisms and services. Details security principals such as conventional encryption and confidentiality, public-key cryptology, authentication and digital signatures, viruses and worms.

Stallings:1995:PYP

[Sta95d] William Stallings. Protect your privacy: the PGP user’s guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-185596-4. xvi + 302 pp.

LCCN TK5102.85.S73 1995. With a foreword by Phil Zimmermann.

Stamper:1995:BDC

[Sta95e] David A. Stamper. Business Data Communications. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1995. ISBN 0-8053-7715-8. xxviii + 608 pp. LCCN TK5105.S734 1994.

Stauffer:1995:UIY

[Sta95f] Todd Stauffer. Using the Internet with your Mac. Que Corp., Indianapolis, IN, USA, 1995. ISBN 0-7897-0665-2. xxii + 308 pp. LCCN ????

Standage:1998:VIR

[Sta98a] Tom Standage. The Victorian Internet: the Remarkable Story of the Telegraph and the Nineteenth Century’s On-line Pioneers. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1998. ISBN 0-8027-1342-4. ix + 227 pp. LCCN HE7631 .S677 1998. US$12.00, CAN$17.00.

Stanfield:1998:CCU

[Sta98b]  Allison Stanfield.  Cyber

courts: using the Internet to assist court processes. Computer Networks and ISDN Systems, 30(1– 7):559–566, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/

                                        tree/store/comnet/sub/1998/
                                        30/1-7/1878.pdf.

Stallings:1999:CNS

[Sta99a] William Stallings. Cryptography and network security: principles and practice. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-869017-0. xvii + 569 pp. LCCN TK5105.59.S713

                                        1999. URL http://www.
                                        prenhall.com/allbooks/esm_
                                        0138690170.html.

Stallings:1999:SSS

[Sta99b] William Stallings. SNMP, SNMPv2, SNMPv3, and RMON 1 and 2. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-48534-6. xv + 619 pp. LCCN TK5105.5.S732 1999. US$54.95.

Stevens:1990:UNP

[Ste90] W. Richard Stevens. UNIX Network Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-949876-1. xi + 772 pp. LCCN QA76.76.O63 S755 1990.

Steiner:1991:DPN

[Ste91a] Henry J. Steiner. Diverse partners: non-governmental organizations in the human rights movement: the report of a retreat of human rights activists, co-sponsored by Harvard Law School Human Rights Program and Human Rights Internet. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1991. ISBN 1-879875-00-4. vii + 90 pp. LCCN K3239.6 1989a.

Stern:1991:MNN

[Ste91b] Hal Stern. Managing NFS and NIS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-75-7. xxiv + 410 pp. LCCN TK5105.5 .S74 1991. URL

http://www.oreilly.com/
catalog/9780937175750;
http://www.oreilly.com/
catalog/nfs.

Sterling:1992:HCL

[Ste92] Bruce Sterling. The Hacker Crackdown: Law and Disorder on the electronic frontier. Bantam Doubleday Dell Publishing Group Inc., 666 Fifth Avenue, New York, NY 10130, USA, 1992. ISBN 0-553-08058-X. xiv + 328 pp. LCCN HV6773.2 .S74 1992, KF9350 .S73 1992. US$23.00. An in-depth examination of the forces of law who try to deal with computer crime, and of the issues involved, written by one of the science fiction writers who invented cyberpunk. The real story behind Operation Sundevil and the Legion of Doom. Readable, informative, amusing, and necessary.

Stevens:1994:TII

[Ste94] W. Richard Stevens. TCP/ IP Illustrated, Volume 1: The Protocols. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63346-9. xix + 576 pp. LCCN TK5105.55 .S74 1994. US$65.95.

Sterne:1995:WWW

[Ste95a] Jim Sterne. World Wide Web marketing: integrating the Internet into your marketing strategy. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-12843-0 (paper). xvii + 331 pp. LCCN HF5415.1265 .S742 1995.

Stewart:1995:HID

[Ste95b] Scott Alan Stewart. 1995 Healthcare Internet directory. Stewart Publishing, Inc., Alexandria, VA, USA, 1995. ISBN 0-936999-31-4. xvii + 121 pp. LCCN ????

Stefik:1996:IDA

[Ste96a]  Mark Stefik.  Internet

Dreams: Archetypes, Myths, and Metaphors. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-19373-6. xxiv + 412 pp. LCCN ZA3250.U6I58 1996. US$30.00.

Stevens:1996:TII

[Ste96b] W. Richard Stevens. TCP/IP Illustrated, Volume 3: TCP for Transactions, HTTP, NNTP, and the UNIX Domain Protocols. Addison-Wesley Professional Computing Series. Addison-Wes-ley, Reading, MA, USA, 1996. ISBN 0-201-63495-3. xix + 328 pp. LCCN TK5105.55.S74 1994. US$41.47.

Stein:1997:HSM

[Ste97] Lincoln D. Stein. How to Set Up and Maintain a Web Site. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 0-201-63462-7. xxi + 793 pp. LCCN TK5105.888.S74 1997. US$39.95.

Stein:1998:WSS

[Ste98a] Lincoln D. Stein. Web Security: a Step-by-Step Reference Guide. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-63489-9. x + 436 pp. LCCN TK5105.59.S74 1998. US$29.95.

Stevens:1998:UNP

[Ste98b] W. Richard Stevens. UNIX Network Programming: Networking APIs: Sockets and XTI, volume 1. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1998. ISBN 0-13-490012-X. xx + 1009 pp. LCCN QA76.76.O63S755 1998. US$59.00. URL http:/

/www.phptr.com/ptrbooks/
ptr_013490012X.html.

Stevens:1999:UNP

[Ste99a] W. Richard Stevens. UNIX Network Programming, Interprocess Communications, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-081081-9 (hardcover). xvii + 558 pp. LCCN QA76.76.O63S755 1998. US$54.00. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130810819.html.

Stewart:1999:BID

[Ste99b] John W. Stewart, III. BGP4: inter-domain routing in the Internet. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37951-1. xiii + 137 pp. LCCN TK5105.555.S74 1999. US$19.95.

Still:1994:IL

[Sti94a] Julie Still, editor. The Internet Library: Case Studies of Library Internet Management and Use. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-965-0. 200 pp. LCCN Z674.82.I59 I59 1994. US$37.50.

Still:1994:ILC

[Sti94b] Julie Still. The Internet library: case studies of library Internet management and use. Mecklermedia, Westport, CT, USA, 1994. ISBN 0-88736-965-0. ix + 185 pp. LCCN Z674.82.I59I59 1994.

Stoll:1989:CET

[Sto89] Clifford Stoll. The Cuckoo’s Egg: Tracking a Spy Through the Maze of Computer Espionage. Doubleday, New York, NY, USA, 1989. ISBN 0-385-24946-2, 0-307-81942-6 (ebook), 0-7434-1145-5, 0-7434-1146-3, 1-299-04734-3. vi + 326 pp. LCCN UB271.R92 H477 1989; UB271.R92 H4771 1989; UB271.R92 S47 1989. US$18.95. URL http:

//vxer.org/lib/pdf/The%
20Cuckoo%27s%20Egg.pdf.

Stockman:1993:GCG

[Sto93] Bernhard Stockman. Global connectivity — the Global Internet eXchange (GIX). Computer Networks and ISDN Systems, 26(3):297– 303, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1993&volume=26&
issue=3&aid=1258.

Stoll:1995:HAI

[Sto95a] Clifford Stoll. Hyper alert:

Internet? bah! Newsweek, 125(9):41–??, February 27, 1995. ISSN 0028-9604.

Stoll:1995:SSO

[Sto95b] Clifford Stoll. Silicon Snake Oil: Second Thoughts on the Information Superhighway. Doubleday, New York, NY, USA, 1995. ISBN 0-385-41993-7. 247 pp. LCCN QA76.9.C66 S88 1995. US$22.00.

Stout:1995:NGR

[Sto95c] Connie Stout. Networks for goals 2000 reform: bringing the Internet to K–12 schools, July 25–September 30, 1994. Technical report, Southwest Educational Development Laboratory, Austin, TX, USA, 1995. vi + 135 pp.

Stoll:1999:HTH

[Sto99] Clifford Stoll. High Tech Heretic: Why Computers Don’t Belong in the Classroom and Other Reflections by a Computer Contrarian. Doubleday, New York, NY, USA, 1999. ISBN 0-385-48975-7. xv + 221 pp. LCCN LB1028.5 .S77 1999. US$24.95.

Strangelove:1992:EMG

[Str92] Michael Strangelove. The electric mystic’s guide to the Internet: a complete bibliography of networked electronic documents, online conferences, serials, software and archives relevant to religious studies. Technical report, Research Centre for the Study of Religion, Department of Religious Studies, University of Ottawa, Ottawa, ON, Canada, 1992.

Strangelove:1993:EMG

[Str93a] Michael Strangelove. The electric mystic’s guide to the Internet: a complete directory of networked electronic documents, online conferences, serials, software and archives relevant to religious studies. Technical report, Research Centre for the Study of Religion, Department of Religious Studies, University of Ottawa, Ottawa, ON, Canada, 1993.

Strother:1993:S

[Str93b] Lesley Strother. smileys.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-041-4. 595 pp. A collection of 650 “smileys”. While not an Internet book per se, smileys are certainly used enough on the Internet to warrant an entry here. :-).

Stremikis:1994:FGI

[Str94] John R. Stremikis. Field guide to the Internet: essential words and phrases, essential tools, essential concepts, effective projects and effective practice: integrating the K– 12 curriculum with the Internet: materials to accompany workshops, April–June 1994. ????, ????, 1994. 409 pp.

WKAR:1995:SLI

[Str95] StreetWatch Lansing: #503 Internet: equal service for all, Sept. 20, 1995, 1995. 1 videocassette (060 min.).

SA:1994:BGI

[Sua94] Suarey Associates. The beginner’s guide to the Internet for windows, 1994. 2 computer disks.

Suarez:1995:BGI

[Sua95] Patrick J. Suarez. The beginner’s guide to the Internet. The Hanford Press, North Canton, OH, USA, deluxe hardcover edition, 1995. ISBN 1-884889-09-3. vi + 177 pp. LCCN ????

Summers:1999:ASI

[Sum99] Charles K. Summers. ADSL Standards, Implementation, and Architecture. The CRC Press advanced and emerging communications technologies series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-9595-X. 186 pp. LCCN TK5105 .S86 1999. US$69.95.

Sunday:1995:YBI

[Sun95] Michael Sunday. This is your brain on the Internet. Ziff-Davis, Emeryville, CA, USA, 1995. ISBN 1-56276-356-3. xi + 47 pp. LCCN PN6231.I62S86 1995.

Sun:1996:DIC

[Sun96] Y. Sun. Design and implementation of the CJK unified character set on UNIX system. In Unicode Consortium [Uni96a], pages B6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

SNIC:1990:SGS

[SUR90] SURAnet Network Information Center. SURAnet guide to selected Internet resources. Technical report, SURAnet Network Information Center, College Park, MD, USA, 1990. various pp.

SURAnet:1992:SIA

[SUR92] SURAnet Network Information Center, College Park, MD, USA. SURAnet: information available on the Internet: a guide to selected sources, June 29, 1992. 56 pp.

SURA:1993:SGS

[SUR93] SURAnet guide to selected Internet resources, page various, 1993. SURAnet Network Information Center, College Park, MD, USA.

Srinivasan:1999:FAL

[SV99] V. Srinivasan and G. Varghese. Fast address lookups using controlled prefix expansion. ACM Transactions on Computer Systems, 17(1): 1–40, February 1999. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tocs/1999-17-1/
                                        p1-srinivasan/.

Schwartz:1993:DSI

[SW93] Michael F. Schwartz and David C. M. Wood. Discovering shared interests using graph analysis. Communications of the ACM, 36(8):78–89, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        163402.html.

Soloway:1997:LED

[SW97] Elliot Soloway and Raven Wallace. Log on education: Does the Internet support student inquiry? don’t ask. Communications of the ACM, 40(5):11–16, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        1997-40-5/p11-soloway/.

Swanson:1991:IPB

[Swa91] Richard B. Swanson. Internet protocol broadcast storms. Thesis (M.S.), University of Colorado, Boulder, CO, USA, 1991. vi + 84 pp.

Scott:1998:VPNa

[SWE98] Charlie Scott, Paul Wolfe, and Mike Erwin. Virtual Private Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1998. ISBN 1-56592-319-7. xi + 177 pp. LCCN TK5105.875.I57 S36 1998. US$29.95. URL http://www.

oreilly.com/catalog/vpn;
http://www.oreilly.com/
catalog/vpn/.

Swingler:1996:ANN

[Swi96] Kevin Swingler. Applying Neural Networks: a Practical Guide. Academic Press, New York, NY, USA, 1996. ISBN 0-12-679170-8. xviii + 303 pp. LCCN ???? US$66.95.

Shinn:1995:LBA

[SY95] Betty Shinn and Jack Yoshi.

Lotus Buddhist art gallery striving to be the Buddhist art and communication center in the Internet, 1995.

Sydow:1996:IMD

[Syd96] Dan Parks Sydow. The Internet for Macs for dummies: quick reference. –For dummies. IDG Books, San Mateo, CA, USA, second edition, 1996. ISBN 1-56884-983-4 (paperback). xvi + 198 pp. LCCN IN PROCESS.

Shibayama:1998:VSC

[SYTM98] E. Shibayama, J. Yabe, S. Takahashi, and M. Matsuda. Visualizing semantic clusters in the Internet information space. Lecture Notes in Computer Science, 1532:409–410, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singhal:1999:NVE

[SZ99] Sandeep Singhal and Michael Zyda. Networked virtual environments: design and implementation. Addison-Wes-ley and ACM Press, Reading, MA, USA and New York, NY 10036, USA, 1999. ISBN 0-201-32557-8. 352 pp. LCCN QA76.76.I59S56 199. US$54.95.

Turner:1998:JPB

[T^{+}98] E. Shane Turner et al. Java programming basics for the Internet. South-Western Educational Publishing, Cincinnati, OH, 1998. ISBN 0-538-68012-1. ???? pp. LCCN QA76.73.J38 J379 1998.

Trager:1996:BEM

[TA96] James Trager and Robert Andrews. Bookshelf 1996–97 edition multimedia reference library, 1996. Includes CD-ROM.

Tsai:1999:CGP

[TAH99] Wei K. Tsai, John K. Antonio, and Garng M. Huang. Complexity of gradient projection method for optimal routing in data networks. IEEE/ACM Transactions on Networking, 7 (6):897–905, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-6/p897-tsai/.

Tanenbaum:1988:CN

[Tan88] Andrew S. Tanenbaum. Computer Networks. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1988. ISBN 0-13-162959-X. xv + 658 pp. LCCN TK5105.5 .T36 1988.

Tannenbaum:1993:HPU

[Tan93] Todd Tannenbaum. How to play the Unix network name services game: DNS. Network Computing, 4(3):156– ??, March 1, 1993. CODEN NCOMEV. ISSN 1046-4468.

Tanenbaum:1996:CN

[Tan96] Andrew S. Tanenbaum. Computer Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1996. ISBN 0-13-349945-6. xvii + 814 pp.

LCCN TK5105.5.T36 1996. US$77.85.

Tatters:1995:NIA

[Tat95a] Wes Tatters. Navigating the Internet with America Online. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 0-672-30763-4. 574 pp. LCCN QA76.57.A43T38 1995.

Tatters:1995:NIC

[Tat95b] Wes Tatters. Navigating the Internet with CompuServe. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 0-672-30761-8. xviii + 474 pp. LCCN TK5105.875.I57T37 1995.

Taves:1995:PTG

[Tav95]  Scott Taves.  A pocket

tour of games on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1694-9 (paperback). xi + 190 pp. LCCN GV1469.15.T38 1995.

Taylor:1995:ICG

[Tay95] Holly Taylor. Internet: a comprehensive guide. Jain Publishing Co., Fremont, CA, USA, 1995. ISBN 0-87573-060-4 (paper). ix + 431 pp. LCCN TK5105.875.I57T38 1995.

Triantafillou:1990:DNM

[TB90] P. Triantafillou and M. Bauer. Distributed name management in Internet systems: a study of design and performance issues. Journal of Parallel and Distributed Computing, 9(4):357–368, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tillman:1993:SNT

[TB93] Hope Tillman and Marian Bremer. Searching for network treasures a librarian’s perspective and guide to Internet training, 1993. 2 video cassettes (159 min.).

Todino:1991:UUU

[TD91] Grace Todino and Dale Dougherty. Using UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA76.76 O63 T63 1991. US$21.95. Well written on how to use uucp and Netnews.

Thomas:1999:DBC

[TE99] Jeffrey L. Thomas and Francis M. Edgington. Digital Basics for Cable TV Systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-743915-6. xv + 283 pp. LCCN TK6678.T46

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0137439156.html.

Teeter:1994:GUI

[Tee94] Robert J. Teeter. A guide to useful Internet resources for Santa Clara Valley Water District staff. Technical report, Santa Clara Valley Water District, San Jose, CA, USA, December 1994. 3 + 80 pp.

Templeton:1995:IJB

[Tem95] Brad Templeton. The Internet Joke Book. Peer to Peer Communications, San Jose, CA, USA, 1995. ISBN 1-57398-025-0. 220 pp. LCCN ???? US$14.95.

Tennant:1992:IBE

[Ten92] Roy Tennant. Internet basics: ERIC digest. Technical report, Syracuse University, ERIC Clearinghouse on Information Resources, Syracuse, NY, USA, October 1992. 8 pp.

Terplan:1999:WBS

[Ter99] Kornel Terplan. Web-based Systems and Network Management. Advanced and emerging communications technologies. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-9598-4. 259 pp. LCCN TK5105.5.T475 1999. US$59.95.

Testa:1996:GTI

[Tes96] Bridget Mintz Testa. Graphical treasures on the Internet. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-685375-4. ???? pp. LCCN T385.T454 1996.

TLA:1995:LIM

[Tex95] Texas Library Association.

Conference (Dallas). Libraries on the Internet moving beyond connectivity, 1995. 1 sound cassette.

Tan:1999:CSI

[TFG99] Zixiang (Alex) Tan, William Foster, and Seymour Goodman. China’s state-coordinated Internet infrastructure. Communications of the ACM, 42 (6):44–52, June 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-6/
p44-tan/.

Tittel:1995:MGJ

[TG95] Ed Tittel and Mark Gaither.

Mecklermedia’s official Internet world: 60 minute guide to java. I D G Books Worldwide, Foster City, CA, USA, 1995. ISBN 1-56884-711-4. xxiii + 253 pp. LCCN TK5105.888 .T58 1995. “Mecklermedia’s official Internet world.”.

Tomasic:1997:DSE

[TGL^{+}97] Anthony Tomasic, Luis Gravano, Calvin Lue, Peter Schwarz, and Laura Haas. Data structures for efficient broker implementation.

ACM Transactions on Information Systems, 15(3):223– 253, July 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/tomasic.
html.

Turletti:1996:VI

[TH96] Thierry Turletti and Christian Huitema. Videoconferencing on the Internet. IEEE/ACM Transactions on Networking, 4(3): 340–351, June 1996. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1996-4-3/p340-turletti/.

Thatcher:1995:EIO

[Tha95] James W. Thatcher. Exploring the Internet with OS/ 2 and screen reader/2. Research report RC. International Business Machines Corporation. Research Division; 19961 International Business Machines Corporation. Research Division. Research report; RC 19961. RC 19961 (88390), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, March 2, 1995. 6 pp.

Anonymous:1993:ILb

[The93] The Internet letter, 1993. ISSN 1070-9851. Net Week Inc., Cabin John, MD, USA.

SUNY:1994:IH

[The94a] The Internet homesteader, page various, 1994. ISSN 1076-4143. SUNY/OCLC Network and State University of New York. Office of Library Services, Albany, NY, USA.

Anonymous:1994:IIJ

[The94b] The Internet insider’s jobs review, page various, 1994. ISSN 1088-954X. Career Advancement Systems, Jamestown, NY, USA.

Theakston:1994:NLP

[The94c] Ian Theakston. NetWare LANs: Performance and Troubleshooting. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63175-X. 400 pp. LCCN TK5105.7 .T53 1994. US$34.50.

Anonymous:1996:IN

[The96a] The Internet newsletter, page various, 1996. ISSN 1088-0615. Leader Publications, New York, NY, USA.

Anonymous:1996:ITb

[The96b] The Internet teacher, page various, 1996. ISSN 1088-7032. Regulus Communications, Lincoln, NE, USA.

Anonymous:1996:LAI

[The96c] The legal automation and Internet review, page various, 1996. ISSN 1087-3651. Fred B. Rothman and Co., Littleton, CO, USA.

Thiel:1993:EYM

[Thi93] Eric Thiel. Everything you might want to know about the internet, but are afraid to ask!: a new users resource. Technical report, Lawrence Livermore National Laboratory, Livermore, CA, USA, 1993. various pp.

Thistle:1994:ITI

[Thi94] Dawn Thistle. Internet: traveling the information superhighway, manual. Technical report, College of the Holy Cross, Worcester, MA, USA, 1994. iv + 31 pp.

Thompson:1994:WWW

[Tho94] Bernie Thompson. World Wide Web. Linux Journal, 3:??, July 1994. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Thomas:1995:ISE

[Tho95a] Brian J. Thomas. The Internet for scientists and engineers: online tools and resources. SPIE Optical Engineering Press, Bellingham, WA, USA, January 1995. ISBN 0-7803-1162-0, 0-8194-1806-4. xix + 450 pp. LCCN TK5105.875.I57 T48 1995. US$30. From the author’s preface: “This book was written to serve three basic purposes: first, to introduce readers to the process of ‘getting online’; second, to provide a basic tutorial to primary Internet software ‘tools’; and last, to point to a variety of resources compiled specifically for the professional scientist or engineer.” Includes examples for Unix, Windows, and Macintosh computers. Includes excellent indices of online resources in: aerospace, agriculture, anthropology, archaeology, artificial intelligence, astronomy, biology, chemistry, computer science, electronics and electrical engineering, energy, engineering, geology, imaging, linguistics, mathematics, medicine, meteorology, oceanography, physics, security, statistics, and virtual reality.

Thompson:1995:IIB

[Tho95b] Brent Allen Thompson. An Internet information browser for use in a digital-audio/ video-interactive-decoder environment. Thesis (M.S.), Iowa State University, Ames, IA, USA, 1995. 52 pp.

Thompson:1995:IRS

[Tho95c] Hugh A. Thompson. Internet resources: a subject guide. Association of College and Research Libraries, Chicago, IL, USA, 1995. ISBN 0-8389-7785-5 (paperback). v + 111 pp. LCCN TK5105.875.I57 I54 1995.

Thomas:1996:ISE

[Tho96a] Brian J. Thomas. The Internet for scientists and engineers: online tools and resources. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996 edition, 1996. ISBN 0-8194-2148-0 (softcover), 0-8194-2017-4 (softcover). xxi + 495 pp. LCCN TK5150.875.I57T48 1996.

Thomas:1996:JPI

[Tho96b] Michael D. Thomas. Java programming for the Internet: a guide to creating dynamic, interactive Internet applications. Ventana Communications Group, Research Triangle Park, NC, USA, 1996. ISBN 1-56604-355-7. ???? pp. LCCN QA76.73.J38J38 1996.

Thomas:1996:ITI

[Tho96c] Stephen A. Thomas. IPng and the TCP/IP protocols: implementing the next generation Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-13088-5 (cloth). xiv + 481 pp. LCCN TK5105.585 .T46 1996.

Thomsen:1996:RCD

[Tho96d] Elizabeth Thomsen. Reference and collection development on the Internet. Howto-do-it manual; no. 66. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-243-0. ???? pp. LCCN Z711.T487 1996.

Thornburgh:1999:FCM

[Tho99] Ralph H. Thornburgh. Fibre Channel for Mass Storage. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010222-9. 170 pp. LCCN TK7895.B87T48 1999.

URL http://www.phptr.
com/ptrbooks/ptr_0130102229.
html.

Anonymous:1994:TI

[Thr94] Threading the Internet, page various, 1994. ISSN 1078-1838. Endless Horizons, Louisville, Ky.

Takemura:1995:IUY

[TI95] Mitsuhiro Takemura and Joichi Ito. Intonet 7-ilgan ui yohaeng. Maeil Kyongje Sinmunsa, Soul, Korea, chopan. edition, 1995. ISBN 89-7442-062-7. 236 pp. LCCN ????

Tillman:1995:ITP

[Til95] Hope N. Tillman. Internet tools of the profession: a guide for special librarians. Special Libraries Association, Washington, DC, USA, 1995. ISBN 0-87111-430-5. v + 200 pp. LCCN QA76.55.I68 1995.

TimeWarner:1994:PTW

[Tim94] Time Warner, Inc. Pathfinder Time Warner’s new home on the Internet, 1994.

Tsai:1997:ICD

[TJ97] I-Fei Tsai and Rong-Hong Jan. Internetting connectionless data networks with a wide area public ATM network. Computer Networks and ISDN Systems, 29(7):797–810, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=7&aid=1677.

Tauber:1995:MAI

[TK95a] Daniel A. Tauber and Brenda Kienan. Mosaic access to the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1656-6. xxiii + 256 pp. LCCN TK5105.882.T38 1995. US$26.99.

Tauber:1995:SIN

[TK95b] Daniel A. Tauber and Brenda Kienan. Surfing the Internet with Netscape. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1740-6. xxv + 332 pp. LCCN TK5105.882.T39 1995.

Tauber:1996:SIN

[TK96] Daniel A. Tauber and Brenda Kienan. Surfing the Internet with Netscape Navigator 2. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1996. ISBN 0-7821-1872-0. xxviii + 403 pp. LCCN ????

Taylor:1999:CJW

[TK99] Christopher Taylor and Tim Kimmet. Core Java Web Server. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080559-9. xxii + 593 pp. LCCN TK5105.888.T39

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130805599.html. Includes

CD-ROM.

Todd:1993:AII

[TL93] Leslie Todd and Barry Leiner, editors. Abstracts of INET 93: International Networking Conference, San Francisco, California, USA, 17– 20 August 1993: the annual conference of the Internet Society. The Society, Reston, VA, USA, 1993. ISBN ???? LCCN ????

Tate:1996:UGI

[TL96] Thomas G. Tate and Glynis D. Long. U.S. government information: Internet and other electronic resources. ACE, ????, second edition, 1996. ISBN ???? 363 pp. LCCN ????

Tan:1997:CNI

[TMF97] Zixiang Alex Tan, Milton Mueller, and Will Foster.

China’s new Internet regulations: two steps forward, one step back. Communications of the ACM, 40(12): 11–16, December 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1997-40-12/
p11-tan/.

Todino:1992:MUU

[TO92] Grace Todino and Tim O’Reilly. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, tenth edition, January 1992. ISBN 0-937175-93-5. 368 pp. LCCN QA76.76 O63 T63. US$27.95. Well written on the setting-up and the maintenance of UUCP and Netnews.

Tennant:1994:CIT

[TOL94a] Roy Tennant, John Ober, and Anne Grodzins Lipow. Crossing the Internet threshold: an instructional handbook. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, second edition, 1994. ISBN 1-882208-07-2. viii + 169 pp. LCCN TK5105.875.I57 T46 1994.

Tennant:1994:CUI

[TOL94b] Roy Tennant, John Ober, and Anne Grodzins Lipow. Cruzando el umbral de la Internet: un manual instructivo. OAS Press, Washington, DC, USA, 1994. ISBN 0-8270-3420-2. ???? pp. LCCN TK5105.875.I57 T4618 1994. Spanish translation of [TOLF93], by Jose Arias Ordonez, Absalon Escobar, and Nohora C. Diaz.

Tolhurst:1994:IRQ

[Tol94c] William A. Tolhurst. The Internet Resource: Quick Reference. Que quick reference series. Que Corporation, Indianapolis, IN, USA, 1994. ISBN 1-56529-748-2 (paperback). 492 pp. LCCN TK5105.875.I57 I57 1994. US$17.99.

Tennant:1993:CIT

[TOLF93] Roy Tennant, John Ober, Anne G. Lipow, and Foreword by Clifford Lynch. Crossing the Internet Threshold: an Instructional Handbook. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 1993. ISBN 1-882208-01-3. viii + 134 pp. LCCN TK5105.875.I57 T46 1993, K662.I4 T46. US$45.00. See [TOL94b].

Tomaiuolo:1993:NDR

[Tom93] Nicholas G. Tomaiuolo. On the nets — Internet database review: The FDA BBS. Database, 16(6):82–??, December 1, 1993. CODEN DTBSDQ. ISSN 0162-4105.

Tools:1995:UIE

[Too95] Tools for the future (Television program). Use the Internet easily, 1995. 1 videocassette (37 min.).

Townsend:1995:SAD

[Tow95] Robert L. Townsend. SNMP Application Developer’s Guide. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01874-6. xvi + 238 pp. LCCN TK5105.5 .T69 1995. US$44.95. Includes two diskettes.

Tolhurst:1994:UI

[TPBH94] William A. Tolhurst, Mary Ann Pike, Keith A. Blanton, and John R. Harris. Using the Internet, Special Edition. Que Corporation, Indianapolis, IN, USA, special edition, 1994. ISBN 1-56529-353-3 (paperback). xx + 1188 pp. LCCN TK 5105.875 I57 T65 1994. US$39.95. System requirements for computer disk: IBM PC.

Tseng:1996:LIP

[TPH96] Gwyneth Tseng, Alan Poulter, and Debra Hiom. The library and information professional’s guide to the Internet. Library Association Publishing, London, UK, 1996. ISBN 1-85604-151-4 (paperback). viii + 199 pp. LCCN ????

Tittel:1994:ME

[TR94a] Ed Tittel and Margaret Robbins. E-mail Essentials. Academic Press, New York, NY, USA, 1994. ISBN 0-12-691397-8. xx + 298 pp. LCCN HE6239.E54 T57 1994. US$24.95. Slade writes: I find it very difficult to define who or what this book is for, but I would recommend it primarily to managers in charge of large networks and possibly Internetworks needing to evaluate email systems with extended functionality. The title would suggest this is a discussion of the basics of email. That is true only in the most simplistic and superficial way.

Tittel:1994:NDE

[TR94b] Ed Tittel and Margaret Robbins. Network Design Essentials. AP Professional, Boston, MA, USA, 1994. ISBN 0-12-691395-1. xx + 282 pp. LCCN TK5105.5.T58 1994.

Tittel:1995:IAE

[TR95] Ed Tittel and Margaret Robbins. Internet Access Essentials. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-691393-5 (paperback). xxiii + 374 pp. LCCN TK5105.875.I57 T58 1995. Adequately covers the basics: what TCP/IP is how to get connected, types of connections, Internet tools, netiquette, and case studies.

Tretter:1995:HUI

[Tre95] Marietta Tretter. How to use the Internet. Ziff-Davis Press, Emeryville, CA, USA, second edition, 1995. ISBN 1-56276-348-2 (paperback). ix + 254 pp. LCCN ????

Tabizel:1996:IIP

[TRHR96] David Tabizel, Nick Rosen, John Hardy, and Richard Rosen. The Internet in 1996: an investment perspective. Durlacher, London, UK, 1996. ISBN ???? xvi + 191 pp. LCCN ????

Triantafillou:1987:INM

[Tri87] Peter Triantafillou. Issues of name management in Internet systems. Thesis (M.Sc.), Dept. of Computer Science, University of Western Ontario, London, UK, 1987. viii + 117 pp.

Trier:1991:UMC

[Tri91] Stephen Trier. User’s manual: the Case Western Reserve University personal computer Internet protocol CWRU-PC/IP: network communications software for the MS-DOS operating system. Case Western Reserve University, Cleveland, OH, USA, version 1.1: March 4, 1991. edition, 1991. various pp.

Trier:1993:UMC

[Tri93]  Stephen Trier.  User’s

manual: the Case Western Reserve University personal computer/Internet protocol CWRU-PC/IP: network communications software for the MS-DOS operating system. Case Western Reserve University, Cleveland, OH, USA, version 2.10 edition, August 23, 1993. various pp.

Troutman:1994:IPM

[Tro94] Leslie Troutman. An Internet primer for music librarians: Tools, sources, current awareness. Notes — Music Library Association, 51(1):22– ??, September 1, 1994. ISSN 0027-4380.

Terry:1988:MSV

[TS88] Douglas B. Terry and Daniel C.

Swinehart. Managing stored voice in the Etherphone system. ACM Transactions on Computer Systems, 6(1):3– 27, February 1988. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1988-6-1/
p3-terry/.

Treese:1998:DSI

[TS98] G. Winfield Treese and Lawrence C. Stewart. Designing Systems for Internet Commerce. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-57167-6. xiii + 375 pp. LCCN QA76.9.S88T74 1998. US$39.95.

Tsudik:1989:DAI

[Tsu89] G. Tsudik. Datagram authentication in Internet gateways: Implications of fragmentation and dynamic routing. IEEE Journal on Selected Areas in Communications, 7(4):499–??, May 1, 1989. CODEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic).

TPI:1995:CSI

[Tsu95] Tsunami Productions, Inc. Cyber surfin’ Internet for the rest of us, 1995. 1 videocassette (ca. 30 min.).

ATT:1995:AD

[TT95a] American Telephone and Telegraph Company. AT&T 800 directory, 1995.

ATT:1995:ATD

[TT95b] American Telephone and Telegraph Company. AT&T toll-free 800 directory, 1995.

Tung:1999:KNA

[Tun99] Brian Tung. Kerberos: a Network Authentication System. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37924-4. 192 pp. LCCN TK5105.59.T86 1999. US$19.95.

Turkheimer:1994:PIN

[Tür94]  Frank Türkheimer.  Pri-

vacy and the Internet: The next step. Computer Networks and ISDN Systems, 27 (3):395–402, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1994&volume=27&
issue=3&aid=1359.

Turkle:1995:LSI

[Tur95a] Sherry Turkle. Life on the screen: identity in the age of the Internet. Simon and Schuster, 1230 Ave. of the Americas, New York, NY 10020, USA, 1995. ISBN 0-684-80353-4. 347 pp. LCCN QA76.9.C66 T87 1995.

Turlingon:1995:WWW

[Tur95b] Shannon Turlingon. Walking the World Wide Web. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-208-9. 350 pp. LCCN TK5105.888 .T87 1995. US$29.95.

Turlington:1995:KI

[Tur95c] Shannon R. Turlington. Kids Internet 500. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-324-

7. ???? pp. LCCN TK5105.875.I57T87 1995.

Turlington:1995:WWW

[Tur95d] Shannon R. Turlington.

Walking the World Wide Web: your personal guide to great Internet resources. Ventana Press, Chapel Hill, NC, USA, Netscape edition, 1995. ISBN 1-56604-269-

0. xxii + 322 pp. LCCN TK5105.888.T87 1995.

Turner:1995:IIH

[Tur95e] Rick Turner. The Internet: the information highway becomes reality, 1995. 1 videocassette (54 min.).

Tuss:1993:EOA

[Tus93] Joan Tuss. Easy online access to helpful Internet guides. Online, 17(5):60–??, September 1, 1993. CODEN ON-LIDN. ISSN 0146-5422.

Takase:1995:JMS

[TW95] Emi Takase and Rae Jean Wiggins. Japanese materials science and engineering information on the Internet: a case study. Mitjp; 95-05 mitjp (series); 95-05, MIT Japan Program, Massachusetts Institute of Technology, Cambridge, MA, USA, 1995. 20 + 8 pp.

Tweney:1994:TGI

[Twe94] Dylan Tweney. The Traveler’s Guide to the Information Highway. Ziff-Davis Press, Emeryville, CA, USA, 1994. ISBN 1-56276-206-0. 139 pp. US$24.95. A guide to the Internet, CompuServe, Prodigy, America Online, Delphi, GEnie, and ZiffNet for CompuServe, this book is the first of its kind to feature real maps of each service, making it exceptionally easy to navigate the on-line world. There are also easyto-follow instructions for using each one, plus tips, shortcuts, and pointers to the best and most interesting features. A comprehensive and easyto-use introduction to the Internet and the largest commercial on-line services.

Tyson:1995:NIO

[Tys95] Herbert L. Tyson. Navigating the Internet with OS/2 Warp. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30719-7. xvii + 437 pp. LCCN TK5105.875.I57T97 1995.

UbietoArtur:1995:DAM

[Ubi95] Antonio-Paulo Ubieto Artur. Documentacion automatizada: manual de uso de la red Internet. Ciencias de la documentacion; 1. Anubar Ediciones, Zaragoza, Spain, 1995. ISBN 84-7013-256-3. 334 pp. LCCN ????

Ubos:1995:WC

[Ubo95] Jeff Ubos. Wheels of commerce. Internet World, 6(4): 62–??, April 1, 1995. CODEN IERNE8. ISSN 1064-3923.

US-Congress:1989:CSV

[Uni89] United States. Congress.

House. Committee on Energy and Commerce. Subcommittee on Telecommunications and Finance. Computer security: virus highlights need for improved Internet management: report to the chairman, Subcommittee on Telecommunications and Finance, Committee on Energy and Commerce, House of Representatives. Washington, DC, USA, 1989. 48 pp.

US-EIA:1993:CCO

[Uni93a] United States. Energy Information Administration. COGIS, Comprehensive Oil and Gas Information Source: available on Internet. Washington, DC, USA, 1993. 1 pp. Shipping list no.: 93-0598-P.

USEIA:1993:CCO

[Uni93b] United States. Energy Information Administration, Washington, DC, USA? CO-GIS, Comprehensive Oil and Gas Information Source: available on Internet, 1993. 6 pp.

UWECCNS:1993:IRG

[Uni93c] University of Wisconsin–Eau Claire. Computing and Networking and Services, Eau Claire, Wis. Internet resource guide, 1993. 110 pp.

USBLS:1995:EBS

[Uni95a] United States. Bureau of Labor Statistics, Washington, DC, USA. Employee benefits survey: a BLS reader, 1995. vi + 254 pp. Shipping list no.: 90-442-P. “February 1995”.

USIM:1995:SIM

[Uni95b] United States Institute of Marketing, Oakmont, PA, USA. Successful Internet marketing, 1995. 139 pp.

UMSILS:1995:IIP

[Uni95c] University of Michigan. School of Information and Library Studies, Ann Arbor, MI, USA. IPL: the Internet public library, 1995.

UMCSL:1995:IL

[Uni95d] University of Missouri–Columbia.

School of Law, Columbia, MO, USA. Internet for lawyers, 1995. 115 pp.

UWVP:1995:IRC

[Uni95e] University of Washington.

Video Production. Internet roadside cafe, 1995. videocassettes (times vary).

UC:1996:PCT

[Uni96a] Unicode Consortium, editor. Pre-conference tutorials proceedings: Software development + the Internet: going global with Unicode: Ninth International Unicode Conference, San Jose, CA, September 4–6 1996. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, 1996. ISBN

???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.]

– B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

UC:1996:SDI

[Uni96b] Unicode Consortium, editor. Software Development + the Internet: Going Global with Unicode: Ninth International Unicode Conference, 4–6 September, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, 1996. ISBN

???? LCCN ???? URL

http://www.unicode.org/
unicode/iuc9conf/index.
html;   http://www.unicode.
org/unicode/uni2book/u2ord.
html.

USFDA:19xx:FDA

[Unixx] United States. Food and Drug Administration. Food and drug administration home page, 19xx.

Upjohn:1995:IRN

[Upj95] Richard Upjohn. Internet resources for the NIS. Technical report, Center for Civil Society International, Seattle, WA (2929 NE Blakeley St., Seattle 98105), 1995. 84 pp.

USWC:1992:AIG

[US 92] US West Communications.

Accessing the Internet the gateway to an information gold mine, 1992. 1 videocassette (17 min.).

USWEST:1993:AIG

[US 93] US WEST, Inc. Accessing the Internet the gateway to an information gold mine, March 15, 1993. 1 videocassette (11 min.).

USDOD:1983:IPM

[U.S83] U.S. Dept. of Defense, Washington, DC, USA. Internet protocol: military standard, August 12, 1983. viii + 72 pp. MIL-STD-1777.

USNTIA:1993:NII

[U.S93] U.S. Dept. of Commerce, National Telecommunications and Information Administration, Washington, DC, USA. The National information infrastructure: agenda for action, 1993. various pp.

USENIX:1997:PUS

[USE97] USENIX, editor. Proceedings of the USENIX Symposium on Internet Technologies and Systems: December 8–11, 1997, Monterey, California. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

USLD:1994:ULN

[Uta94] Utah State Library Division, Salt Lake City, UT, USA. Utah library network and Internet training handbook, 1994. various pp.

USGAO:1989:CSV

[UU89] United States. General Accounting Office and United States. Congress. House. Committee on Energy and Commerce. Subcommittee on Telecommunications and Finance. Computer security: virus highlights need for improved Internet management: report to the chairman, subcommittee on telecommunications and finance, committee on energy and commerce, house of representatives. Technical Report GAO/IMTEC-89-57, GAO, Washington, DC, USA, June 1989. 48 pp.

Vacca:1996:ISS

[Vac96a] John R. Vacca. Internet security secrets. IDG Books, San Mateo, CA, USA, 1996. ISBN 1-56884-457-3. xxxii + 758 pp. LCCN TK5105.59.V33 1996.

Vacca:1996:VBV

[Vac96b] John R. Vacca. VRML: bringing virtual reality to the Internet. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-709910-7, 0-12-709911-5 (CD-ROM). xxxv + 552 pp. LCCN T385.V322 1996.

Vacca:1998:CH

[Vac98]  John Vacca.  The Ca-

bling Handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-080531-9. ???? pp. LCCN TK5103.12.V33

1998. URL http://www.
phptr.com/ptrbooks/ptr_
0130805319.html.

Vacca:1999:MIS

[Vac99] John R. Vacca. MCSE:

Implementing and Supporting Microsoft System Management Server 2. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-020226-6. ???? pp. LCCN QA76.3.V33

1999. URL http://www.
phptr.com/bookseri/mcse.
html.

Valauskas:1993:MMI

[Val93a] Edward Valauskas. Mac monitor — One-Stop Internet shopping: NCSA Mosaic on the Macintosh. Online, 17 (5):99–??, September 1, 1993. CODEN ONLIDN. ISSN 0146-5422.

Valovic:1993:RCN

[Val93b] T. Valovic. The role of computer networking in the emerging virtual marketplace. Telecommunications (Americas ed.), 27(5):40–??, May 1, 1993. ISSN 0278-4831.

Vansickle:1994:TGI

[Van94] Sharon Vansickle. A teacher’s guide to the Internet. Technical report, Norcross High School, Norcross, GA, USA, 1994. various pp.

VanLeeuwen:1995:PBP

[Van95] Steven Van Leeuwen. Project bartleby the public library of the Internet, 1995.

Vandendorpe:1996:SYW

[Van96] Laura Vandendorpe. Scientist of the Year weaves Web over the world. Research & Development, 38(12):14– 16, 18, November 1996. CO-DEN REDEEA. ISSN 0746-9179. Profile of WWW inventor Tim Berners-Lee. Also includes discussion of the Jigsaw HTTP server, written entirely in Java.

Varadan:1995:PTT

[VB95a] Savitha Varadan and Kenyon Brown. A pocket tour of travel on the Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1760-0. xii + 211 pp. LCCN G153.4.V37 1995.

Varadan:1995:VI

[VB95b] Savitha Varadan and Kenyon Brown. Voyages sur Internet. Pocket tour Internet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1633-X. xx + 188 pp. LCCN ????

ViaGrafix:1994:CI

[Vc94] ViaGrafix and Internet communicating. Communicating with Internet, 1994. 1 videocassette (60 min.) (3 1/2 in.).

Wees:1995:IVJ

[vdW95] J. G. L. van der Wees. Internet voor juristen. Kluwer, Deventer, The Netherlands, 1995. ISBN 90-268-2793-8. x + 200 pp. LCCN ????

Valauskas:1996:ITS

[VE96] Edward Valauskas and Monica Ertel. The Internet for teachers and school library media specialists: today’s applications, tomorrow’s propects. Neal-Schuman NetGuide series Neal-Schuman net-guide series. Neal-Schuman Publishers, New York, NY, USA, 1996. ISBN 1-55570-239-2. x + 231 pp. LCCN LB1044.87.I58 1996.

Vince:1998:VWI

[VE98] John Vince and Rae Earnshaw, editors. Virtual worlds on the Internet. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8700-2. 360 pp. LCCN QA76.625.V57 1998. US$55.00.

Velocci:1995:STI

[Vel95] Anthony L. Velocci. Southwest to tap Internet for ticketless travelers. Aviation week & space technology, 142(26): 29, June 26, 1995. ISSN 0005-2175.

VanMeter:1998:VNVa

[VFH98a] Rodney Van Meter, Gregory G. Finn, and Steve Hotz. VISA: Netstation’s virtual Internet SCSI adapter. Operating Systems Review, 32 (5):71–80, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanMeter:1998:VNVb

[VFH98b] Rodney Van Meter, Gregory G. Finn, and Steve Hotz. VISA: Netstation’s Virtual Internet SCSI Adapter. ACM SIGPLAN Notices, 33 (11):71–80, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/291069/
p71-van_meter/.

Veljkov:1994:PGIa

[VH94a] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 1 — Telnetting. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-943-X. 64 pp.

LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGIb

[VH94b] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 2 — Transferring Files with File Transfer Protocol (FTP). Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-944-8. 64 pp. LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGIc

[VH94c] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 3 — Using and Navigating Usenet. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-945-6. 64 pp. LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGId

[VH94d] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 4 — The Internet E-Mail. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-946-4. 64 pp. LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGIe

[VH94e] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 5 — Basic Internet Utilities.

Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-947-2. 64 pp. LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGIf

[VH94f] Mark Veljkov and George Hartnell. Pocket Guides to the Internet: Volume 6 — Terminal Connections. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1994. ISBN 0-88736-948-0. 64 pp. LCCN TK5105.875.I57 V45 1994. US$9.95.

Veljkov:1994:PGI

[VH94g] Mark D. Veljkov and George Hartnell. Pocket guides to the Internet. Mecklermedia, Westport, CT, USA, 1994. ISBN 0-88736-943-X (vol. 1), 0-88736-944-8 (vol. 2), 0-88736-945-6 (vol. 3), 0-88736-946-4 (vol. 4), 0-88736-947-2 (vol. 5), 0-88736-948-0 (vol. 6). LCCN TK5105.875.I57 V45 1994. USS$7.00 (per volume).

Viagrafix:1994:TYI

[Via94] Viagrafix. Teach yourself Internet, 1994. 1 videocassette (52 min.) (3 1/2 in.).

Vickery:1988:IPI

[Vic88] William Edward Vickery. IN-TERNET 1988 profiles of international development contractors and grantees. Technical report, Network for International Technical Assistance, Chapel Hill, NC, USA, 1988. various pp.

Vickery:1989:IPE

[Vic89] William Edward Vickery. IN-TERNET 1989–1990 profiles electronic index, 1989. 2 computer disks.

Viehland:1993:NFM

[Vie93] Dennis W. Viehland. Notes from the Field — Dear Mr. President: a story of misinformation distribution in cyberspace. Internet research, 3 (3):57–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

Vincent:1994:FSI

[Vin94a] Patrick Vincent. Free Stuff from the Internet. Coriolis Group Books, Scottsdale, AZ, USA, 1994. ISBN 1-883577-11-X. 459 pp. US$19.99. Once you get on the Internet, everything is free so the title is somewhat of a misnomer. Predominately a guide to areas where users can download software, hear sound bites, or take a tour of Europe. The books are arranged by categories such as “Business,” “Around the House,” and “Entertaining,” and “Educational Ideas.”.

Vincent:1994:FTI

[Vin94b] Patrick Vincent. Free $tuff from the Internet. Coriolis Group Books, Scottsdale, AZ, USA, 1994. ISBN 1-883577-11-X (paperback). 459 pp. LCCN ????

Vincent:1995:BCC

[Vin95a] Patrick Vincent. Les bonnes combinés: carnet d’adresses. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1521-X. xxii + 188 pp. LCCN ????

Vincent:1995:WMF

[Vin95b] Patrick Vincent. Way more! free $tuff from the Internet. Coriolis Group Books, Scottsdale, AZ, USA, 1995. ISBN 1-883577-50-0 (paperback). xxx + 418 pp. LCCN TK5105.875.I57V56 1995.

Vincent:1996:FTI

[Vin96] Patrick Vincent. Free $tuff from the Internet. Coriolis Group Books, Scottsdale, AZ, USA, second edition, 1996. ISBN 1-883577-79-9 (paperback). 590 pp. LCCN ????

VLBISSTHJR:1994:HI

[Vir94] Virginia. Library Board. Internet Staff Study Team and House Joint Resolution no.

76. HJR 76 Internet, 1994. 2 sound cassettes.

VSGC:1997:EGI

[Vir97] Virginia Space Grant Consortium. The educator’s guide to the Internet: a handbook with resources and activities. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-49609-7. xi + 236 pp. LCCN ????

Viterbi:1995:CPS

[Vit95] Andrew J. Viterbi. CDMA:

Principles of Spread Spectrum Communication. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-63374-4. xix + 245 pp. LCCN TK5103.45 .V57 1995. US$59.25.

Valauskas:1995:IIL

[VJ95] Edward Valauskas and Nancy (Nancy R.) John. The Internet initiative: libraries providing Internet services and how they plan, pay, and manage. American Library Association, Chicago, IL, USA, 1995. ISBN 0-8389-0668-0. xiv + 220 pp. LCCN Z674 .I59 1995. Vandenwauver:1998:SIE

[VJ98] M. Vandenwauver and F. Jorissen. Securing Internet electronic mail. Lecture Notes in Computer Science, 1528:209– 223, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitek:1999:SIP

[VJ99] Jan Vitek and Christian D.

Jensen. Secure Internet programming: security issues for mobile and distributed objects, volume 1603 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66130-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 501 pp. LCCN QA76.625 .S43 1999 Bar.

Viegnes:1995:GI

[VM95] Laurent Viegnes and Michel Marecaux. Guide Internet. Micro Application, Paris, France, 1995. ISBN 2-7429-0552-9. 258 pp. LCCN ????

Volkman:1997:CCSe

[Vol97a] Victor R. Volkman. C/C++ sources: Internet on the Internet. C/C++ Users Journal, 15(5):87–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1997:CCSc

[Vol97b] Victor R. Volkman. C/C++ sources: UNIX and database resources on the Internet. C/C++ Users Journal, 15 (3):89–??, March 1997. CO-DEN CCUJEX. ISSN 1075-2838.

VagaFernandez:1995:I

[VP95] Jose Maria Vaga Fernandez and Rafael Perez Munoz, aut. Internet. Al Dia en una Hora. Anaya Multimedia, Madrid, Spain, 1995. ISBN 84-7614-735-X. 128 pp. LCCN ????

Volkerding:1997:LIS

[VRFJ97] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. The Linux Internet Server. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel:

(503) 282-5215, 1997. ISBN 1-55828-545-8. xxi + 530 pp. LCCN QA76.9.C55R444 1997. US$39.95, CDN$55.95.

URL http://www.mispress.

com/. Includes CD-ROM.

Varakulsiripunth:1986:CPI

[VSN86] Ruttikorn Varakulsiripunth, Norio Shiratori, and Shoichi Noguchi. Congestion-control policy on the internetwork gateway. Computer Networks and ISDN Systems, 11(1): 43–58, January 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Vetter:1994:MWW

[VSW94] Ronald J. Vetter, Chris Spell, and Charles Ward. Mosaic and the World Wide Web. Computer, 27(10): 49–57, October 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vitek:1997:MOS

[VT97] Jan Vitek and Christian Tschudin, editors. Mobile object systems: towards the programmable Internet: second international workshop, MOS ’96, Linz, Austria, July 8–9, 1996: selected presentations and invited papers, volume 1222 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-62852-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.625 .M68 1996.

Vriethoff:1985:PMP

[VVB85] Willem Vriethoff, Co Visser, and Hanko Boerma, editors. Project management, “Clarity for the 90’s”: proceedings of the 8th INTERNET World Congress, May 19–24, 1985, De Doelen, Rotterdam, The Netherlands. North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN ???? LCCN ????

Wilkinson:1998:PPT

[WA98] Barry Wilkinson and Michael Allen. Parallel Programming: Techniques and Applications Using Networked Workstations and Parallel Computers. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-671710-1. xv + 431 pp. LCCN QA76.642.W54

1999. URL http://www.
phptr.com/allbooks/esm_
0136717101.html.

Wagner:1995:UIV

[Wag95a] Judith O. Wagner. Using the Internet in vocational education. ERIC digest; no. 160 ERIC digest (Columbus, Ohio); no. 160. EDO-CE-95-160, ERIC Clearinghouse on Adult, Career, and Vocational Education Center on Education and Training for Employment, College of Education, the Ohio State University, Columbus, OH, USA, 1995.

Wagner:1995:CIT

[Wag95b] Richard Wagner. The CompuServe Internet tour guide: easy graphical access– the CompuServe way. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-259-3. xxiv + 210 pp. LCCN TK5105.875.I57W34 1995.

Walsh:1993:DCI

[Wal93] R. Walsh. Development of a community information service: The National Capital Area Public Access Network (CapAccess) — A work in progress. Internet research, 3 (2):41–??, Summer 1993. CO-DEN IRESEF. ISSN 1066-2243.

Walker:1994:LIE

[Wal94a] Thomas D. Walker. Libraries and the Internet: education, practice, and policy. Library trends, 42(4); Library trends; vol. 42, no. 4. University of Illinois Graduate School of Library and Information Science, Champaign, IL, USA, Spring 1994. ISBN ???? ISSN 0024-2594. 585–758 pp. LCCN ????

WCC:1994:IIC

[Wal94b] Walnut Creek CDROM. Internet info CD-ROM, March 1994. Includes CD-ROM.

Waltz:1995:IID

[Wal95a] Mitzi Waltz. The Internet international directory. Ziff-Davis, Emeryville, CA, USA, 1995. ISBN 1-56276-329-6 (paperback). 526 pp. LCCN ????

Walzer:1995:RSI

[Wal95b]  Norman Walzer.  Rural

schools and the Internet: providing an ‘on/off ramp’ to the information superhighway of the 21st century. Rural research report; vol. 6, issue 4 (winter 1994–95), Western Illinois University, Macomb, IL, USA, 1995. 7 + 1 pp.

Wambold:1995:NGL

[Wam95] Sally Wambold. Net gain learning to bank on the Internet, 1995. 1 sound cassette.

Wang:1993:ISI

[Wan93] Tao Wang. Improvement and simulation of Internet routing information protocol. Thesis (M.S.), Southwest Texas State University, San Marcos, TX, USA, 1993. iii + 93 pp.

Wang:1994:SOL

[Wan94] Yuting Wang. A study of online library search methods through Internet at California University of Pennsylvania. Research project (m. ed.), California University of Pennsylvania, California, PA 15419-1394, USA, 1994. iii + 37 pp.

Wang:1996:IUX

[Wan96]  Paul S. Wang.  An in-

troduction to UNIX with X and the Internet. PWS Pub. Co., Boston, MA, USA, 1996. ISBN 0-534-94768-9. ???? pp. LCCN QA76.76.O63W365 1996.

Warschauer:1995:EET

[War95] Mark Warschauer. E-mail for English teaching: bringing the Internet and computer learning networks into the language classroom. Teachers of English to Speakers of Other Languages, Alexandria, VA, USA, 1995. ISBN 0-939791-62-5 (paperback). 120 pp. LCCN ????

Ward:1999:CIP

[War99] Andrew F. Ward. Connecting to the Internet: a Practical Guide about LAN-Internet Connectivity. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37956-2. xv + 291 pp. LCCN TK5105.875.I57W375 1999. US$19.95.

Watson:1993:GEI

[Wat93] Ian Watson. The great electronic information bazaar — a rough guide to exploring the Internet. ASLIB Proceedings, 45(6):153–??, June 1, 1993.

CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Watson:1996:PIA

[Wat96] Mark Watson. Programming intelligent agents for the Internet. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912206-X. xvi + 240 pp. LCCN TK5105.875.I57W38 1996.

Wack:1994:KYS

[WC94] John P. Wack and Lisa J. Carnahan. Keeping your site comfortably secure an introduction to Internet firewalls. NIST special publication Computer security 800-10, U.S. Dept. of Commerce, Technology Administration, National Institute of Standards and Technology; For sale by the Supt. of Docs., U.S. G.P.O., Gaithersburg, MD, USA, 1994. xii + 70 pp.

WMM:1995:AI

[WC95a] Wentworth Worldwide Media, Inc and Classroom Connect. The amazing Internet, 1995. 1 videocassette (20 min.).

WMM:1995:E

[WC95b] Wentworth Worldwide Media, Inc and Classroom Connect. Email, 1995. 1 videocassette (44 min.).

WMM:1995:ST

[WC95c] Wentworth Worldwide Media, Inc and Classroom Connect. Search tools, 1995. ISBN 0-932577-22-9. 1 videocassette (43 min.).

Woo:1995:WAI

[WC95d] Dennis Woo and Raye Cole.

Webmaster: an Introduction to Electronic Publishing on the Global Internet. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94433-8. 224 pp. LCCN ???? US$24.95.

Washburn:1993:TIR

[WE93] Kevin Washburn and Jim T.

Evans. TCP/IP—Running a Successful Network. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62765-5. xvii + 537 pp. LCCN TK5105.5 .W37 1993.

Wagner:1996:MIT

[WE96] Ronald L. Wagner and Eric Engelmann. The McGraw-Hill Internet training manual. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-066937-6 (paperback). xx + 332 pp. LCCN TK5105.875.I57 W37 1996.

Weber:1995:CWP

[Web95] Kate Weber. Chapter 6, in which Pooh proposes improvements to Web authoring tools, having seen said tools for the Unix platform. Computer Networks and ISDN Systems, 27(6): 823–829, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1995/
27/6/1465.pdf.

Webster:1999:YSG

[Web99] Bruce Webster. The Y2K Survival Guide: Getting To, Getting Through, and Getting Past the Year 2000. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-021496-5. xxvii + 544 pp. LCCN QA 76.76 S64 W43 1999. URL http:/

/www.phptr.com/year2000/
survial.html.

Weingarten:1993:NNI

[Wei93] Fred W. Weingarten. NREN and the national infrastructure: a personal vision. Internet research, 3(3):2–??, Fall 1993. CODEN IRESEF. ISSN 1066-2243.

Weiss:1995:CIG

[Wei95a] Aaron Weiss. The complete idiot’s guide to protecting yourself on the Internet. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-593-7. xx + 281 pp. LCCN TK5105.875.I57.W45 1995.

Weiss:1995:HSJ

[Wei95b] Aaron Weiss. Hop, skip, and jump. Internet World, 6(4): 41–??, April 1, 1995. CODEN IERNE8. ISSN 1064-3923.

Weiss:1995:HS

[Wei95c] Aaron Weiss. Hot sites. Internet World, 6(4):76–??, April 1, 1995. CODEN IERNE8. ISSN 1064-3923.

WMM:1994:AOI

[Wen94] Wentworth Worldwide Media. All in one Internet: the Internet graphical software pak and tourguide handbook. Wentworth Worldwide Media, Inc., Lancaster, PA, USA, 1994. ISBN 0-932577-01-6. LCCN ???? 3 computer disks.

WMM:1995:DWW

[Wen95] Wentworth Worldwide Media. Discovering the World Wide Web: Understanding NetScape, 1995. ISBN 0-932577-21-0. 1 videocassette (53 min.).

WLN:1994:WIT

[Wes94] Western Library Network, Lacey, WA, USA. WLN Internet training tip, 1994.

Wessels:1995:IDI

[Wes95] John A. Wessels. Information dissemination via the Internet: development of a homepage for the NIST time and frequency division. Thesis (M.S.C.I.S.), Regis University, Denver, CO, USA, 1995. ii + 49 pp.

West:1997:UIB

[Wes97]  E. J. W. West.  Using

the Internet for business — Web oriented routes to market and existing IT infrastructures. Computer Networks and ISDN Systems, 29(15):1769–1776, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1997&volume=29&
issue=15&aid=1771.

Wesel:1998:WMC

[Wes98] Ellen Kayat Wesel. Wireless Multimedia Communications. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-63394-9. xxxi + 298 pp. LCCN TK5103.2.W47 1998. US$62.95.

Walsh:1994:WI

[WG94a] Michael J. Walsh and Eric Gagnon. What’s on the Internet. InternetInfo, Fairfax, Va., U S A, 1994. ISBN 1-884640-16-8. 261 pp. LCCN ????

Wood:1994:GVF

[WG94b] Lisa R. Wood and Kristen Garlock. Guide to video and film resources on the Internet, 1994.

White:1995:GOS

[Whi95] Jim White. Get on-line in Southern California: your regional guide to the Internet and the World Wide Web. Cyberguides. Aaron Blake Publishers, Gardena, CA, USA, 1995. ISBN 0-937609-14-5. 90 pp. LCCN ????

ORA:1993:WIC

[Who93] The Whole Internet catalog, 1993. ISSN 1072-043X. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA.

Wiencko:1993:BEV

[Wie93] Joseph Wiencko. The Blacksburg electronic village. Internet research, 3(2):31–??, Summer 1993. CODEN IRE-SEF. ISSN 1066-2243.

Wiese:1996:KI

[Wie96] Michael Wiese. Kids on the Internet, 1996. ISBN 0-941188-53-1. 1 videocassette (30 min.).

Wiggins:1993:SCR

[Wig93] Gary Wiggins. Some chemistry resources on the Internet. Technical report, Indiana University Chemistry Library, Bloomington, IN, USA, 1993. 9 pp.

Wiggins:1995:WSW

[Wig95a] Richard Wiggins. What to say when MIS won’t connect your library to the Internet.

Searcher: the magazine for database professionals, 3(1): 18–??, January 1, 1995. CO-DEN SMDPE8. ISSN 1070-4795.

Wiggins:1995:IEG

[Wig95b] Richard W. Wiggins. The Internet for Everyone: a Guide for Users and Providers. J. Ranade workstation series. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-067019-6 (paperback), 0-07-067018-8 (hardcover). xxi + 655 pp. LCCN TK5105.875.I57 W54 1995. US$45.00 (hardcover), US$29.95 (paperback). The purpose of this book is to tell people how to use the Internet, but also how to provide information via the Internet — picking a kind of server technology, installing a Gopher, Web, or FreeWAIS server, an overview of HTML, pointers to more detailed online info, etc.

Wilder:1993:GTI

[Wil93] Floyd Wilder. A Guide to the TCP/IP Protocol Suite. Artech House Inc., Boston, MA, USA, 1993. ISBN 0-89006-693-0. xx + 313 pp. LCCN TK5105.5 .W55 1993. US$77. A technical reference guide describing all major Internet protocols.

Wilkes:1994:IPS

[Wil94a] T. Wilkes, editor. IPCC 94 Proceedings. Scaling New Heights in Technical Communication. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1936-2. LCCN T10.5.I54 1994. IEEE Catalog No. 94CH3442-1.

Williams:1994:IIK

[Wil94b] Katherine P. Williams. Integration of the Internet into the K–12 academic environment with emphasis on Minnesota education. Technical report, ????, ????, 1994. iv + 76 pp.

Willett:1995:TLI

[Wil95a] Edward F. Willett. THO-MAS legislative information on the Internet, 1995.

Williams:1995:IT

[Wil95b] Bard Williams. The Internet for teachers. –For dummies. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-600-2. xxviii + 350 pp. LCCN LB1044.87.W55 1995.

Wilson:1995:TUI

[Wil95c] Patricia E. Wilson. Training users on the Internet pitfalls, perils, and pleasures, June 1995. 2 sound cassettes.

Williams:1996:INE

[Wil96a] Constance D. Williams. Internet for newbies: an easy access guide. Libraries Unlimited, Englewood, CO, USA, 1996. ISBN 1-56308-483-X. ???? pp. LCCN TK5105.875.I57W555 1996.

Williams:1996:BOI

[Wil96b]  Joseph Williams.  BOTS

and other Internet beasties. Sams.net Pub., Indianapolis, IN, USA, 1996. ISBN 1-57521-016-9 (paperback). xxv + 505 pp. LCCN ????

Williamitis:1998:ICM

[Wil98] Anthony Williamitis. Internet connections with the 56Kbps modems. Linux Journal, 45:??, January 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Wilde:1999:WWT

[Wil99] Erik Wilde. Wilde’s WWW: technical foundations of the World Wide Web. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64285-4. xxv + 594 pp. LCCN TK5105.888.W545 1999.

Winslett:1993:SGO

[Win93] Marianne Winslett. SIG-MOD goes online: new member services via Internet. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):4– 6, September 1993. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

WinStruct:1996:LMW

[Win96a] WinStruct, Inc. Learn! Microsoft Windows 95 Internet, email, faxing and more!, 1996. 1 videocassette.

Winters:1996:GIF

[Win96b] Tracey Winters. Get it free on the Internet: the pennypincher’s guide to online bargains. Go!Guides Go!Guides series. Motion Works Pub., San Francisco, CA, USA, 1996. ISBN 1-57712-012-4 (paperback). ???? pp. LCCN TX335.W515 1996.

Winston:1998:MTS

[Win98] Brian Winston. Media technology and society: a history: from the telegraph to the Internet. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1998. ISBN 0-415-14229-6, 0-415-14230-X (paperback). xiv + 374 pp. LCCN P96.T42 W49 1998.

Wishon:1990:DSI

[Wis90] Gordon D. Wishon. Data security and integrity in open networks: a prototype implementation of Internet standard privacy-enhanced electronic mail. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1990. vi + 332 pp.

WTCSSMIC:1996:IPG

[Wis96] Wisconsin Technical College System. Statewide Marketing Internet and Committee, Madison, WI, USA. Internet policy guide, 1996. various pp.

Withers:1996:SPI

[Wit96] John P. Withers. SGML publising on the Internet: designing, building and distributing Internet documents. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-366-2. ???? pp. LCCN ????

Weeby:1995:IPN

[WK95] Steve Weeby and KKStudios. The Internet primer: navigating the Internet, 1995. 1 videocassette (VHS)(55 min.).

Willebeek-LeMair:1998:BAV

[WLKS98] M. H. Willebeek-LeMair, K. G. Kumar, and E. C. Snible. Bamba — Audio and video streaming over the Internet. IBM Journal of Research and Development, 42 (2):269–280, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
almaden.ibm.com/journal/
rd/422/willebeek.html.

Wiseberg:1990:DOD

[WLM90] Laurie S. Wiseberg, Guadalupe Lopez, and Sarah Meselson. Directorio de organizaciones de derechos humanos: America Latina y El Caribe = Human rights directory: Latin America and the Caribbean. Human Rights Internet Reporter; vol. 13, nos.

2 and 3, special issue (Jan. 1990). Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, January 1990. ISBN 0-939338-05-X. 528 pp. LCCN JC571.H7696a.

Wang:1998:CPC

[WLW98] Huaiqing Wang, Matthew K. O. Lee, and Chen Wang. Consumer privacy concerns about Internet marketing. Communications of the ACM, 41(3): 63–70, March 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-3/
p63-wang/.

Watkins:1995:IEB

[WM95] Christopher Watkins and Stephen Marenka. The Internet edge in business. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-737840-5, 0-12-737841-3 (CD-ROM). xv + 272 pp. LCCN HF54.56 .W38 1995.

Wiseberg:1987:HRI

[WMRS87] Laurie S. Wiseberg, Eileen Maloy, Laura Reiner, and Hazel Sirett. Human Rights Internet directory: Eastern Europe and the USSR. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1987. ISBN 0-939338-03-3. 304 pp. LCCN JC599.E92H86 1987.

Wiebe:1996:WWE

[WMS96] James H. Wiebe, Jeffrey M. Moreton, and Simeon P. (Simeon Peter) Slovacek. Works for Windows for educators: includes coverage of the Internet, CBI, and LinkWay. Franklin, Beedle, Wilsonville, OR, USA, 1996. ISBN 0-938661-74-4. viii + 391 pp. LCCN QA76.76.I57W49 1996.

Winder:1995:AYN

[Wn95] Davey Winder and net. All you need to know about the Internet. Future Publishing, Bath, UK, 1995. ISBN 1-85870-064-7 (paperback). xvi + 324 pp. LCCN ????

Williams:1994:ITV

[WNS^{+}94] Evan Williams, Erik Nielsen, Robert Siemann, Shirley Ann Gartmann, and Brendan P. Kehoe. Internet TV video guide to global connectivity, 1994. 1 videocassette (120 min.) binder, 1 booklet, 4 IBM compatible computer disks, 2 simulator computer disks and additional miscellaneous information.

Wolff:1994:N

[Wol94] Michael Wolff. netgames. Random House, New York, NY, USA, 1994. ISBN ???? 272 pp. LCCN ???? US$19.00. Games on the Internet.

Wolff:1995:N

[Wol95] Michael Wolff. netmoney.

Random House, New York, NY, USA, 1995. ISBN ???? LCCN ???? Financial information on the Internet.

Wolgemuth:1996:LOE

[Wol96] Amy Wolgemuth. Learning online: an educator’s easy guide to the Internet. IRI/Skylight Training and Pub., Palatine, IL, USA, 1996. ISBN 1-57517-009-4. xvii + 156 pp. LCCN LB1044.87.W65 1996.

Wolfe:1998:EFG

[Wol98] Rosalee Wolfe. EDUCA-TION: 3D freebies: a guide to high quality 3D software available via the internet. Computer Graphics, 32(2): 31–33, May 1998. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Wood:1994:NAD

[Woo94] Lamont Wood. The Net After Dark. John Wiley, New York, NY, USA, November 1994. ISBN 0-471-10347-0. xi + 356 pp. LCCN TK5105.5 .W64 1994. US$16.95. Ad says: The Underground Guide to the Coolest the Newest and the Most Bizarre Hangouts on the Internet, CompuServe, AOL, Delphi, and More.

Wood:1995:NAD

[Woo95a] Lamont Wood. The net after dark: the underground guide to the coolest, newest, and the most bizzare hangouts on the Internet, CompuServe, AOL, Delphi, and more. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-10347-0 (paperback). xi + 356 pp. LCCN TK5105.5.W64 1994.

Woodul:1995:CEI

[Woo95b] Charles E. Woodul. Characteristics of an early Internet adoption in a middle school. Thesis (Ed.S.), Department of Educational Leadership, East Carolina University, Greenville, NC, USA, 1995. 169 pp.

Woodward:1996:ITG

[Woo96] Meredith Bain Woodward. Internet travel guide: practical business and adventure resources. Go!Guides series. Motion Works, San Francisco, CA, USA, 1996. ISBN 1-57712-007-8. ???? pp. LCCN G155.A1W64 1996.

WHVPCI:1994:HWH

[Wor94] Workshop on Handling Vernacular on Personal Computers and over the Internet (Boston, MA). Handouts of Workshop on Handling Vernacular on Personal Computers and over the Internet. ????, 1994. various pp.

Woods:1995:DFG

[WoRIGSoO95] Wendy L. (Wendy Lynn)

Woods and University of Rhode Island. Graduate School of Oceanography. Drinking from the firehose: global change datasets available through Internet. GSO technical report 95-1 NML/GSO, University of Rhode Island, Graduate School of Oceanography, Narragansett, RI, USA, 1995. 42 pp.

Wiseberg:1989:AHR

[WR89] Laurie S. Wiseberg and Laura Reiner. Africa, human rights directory and bibliography. Human Rights Internet reporter, vol. 12, no. 4, special issue (winter 1988/1989); Human Rights Internet reporter; vol. 12, no. 4. Human Rights Internet, University of Ottawa, Ottawa, ON, Canada, 1989. ISBN 0-939338-04-1. ISSN 0275-049X. 308 pp. LCCN JC571.H7696a.

Wright:1993:VA

[Wri93] Robert Wright. Voice of America. New republic, 209 (11):20–??, September 13, 1993. ISSN 0028-6583.

Wright:1995:LEC

[Wri95a] Benjamin Wright. The law of electronic commerce: EDI, Email, and Internet: technology, proof, and liability. Little, Brown and Co., Boston, MA, USA, second edition, 1995. ISBN 0-316-95620-1. various pp. LCCN KF1357.5 .W753 1995. Loose-leaf for updating.

Wright:1995:GYC

[Wri95b] Chris Wright. Getting your church onto the Internet. Technical report, C. Wright, Keighley, UK, September 24, 1995. 28 pp.

Wiseberg:1980:NAH

[WS80] Laurie S. Wiseberg and Harry M. Scoble. North American human rights directory 1980: human rights Internet. Garrett Park Press, Garrett Park, MD, USA, 1980. ISBN ???? 181 pp. LCCN ????

Wolf:1994:AMO

[WS94] Gary Wolf and Michael Stein. Aether Madness: an Offbeat Guide to the Online World. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1994. ISBN 1-56609-020-2. 297 pp. US$21.95. Culture counterculture and no-culture — with an emphasis on the latter two labels — this lively tour showcases the diversity found on the Internet and Bulletin Boards, from space aliens and sex mania, to Limbaugh and recreational drug use. Full text available

at http://www.aether.com/
Aether/.

Wright:1995:TII

[WS95] Gary Wright and W. Richard Stevens. TCP/IP Illustrated: Volume 2. The Implementation. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-63354-X. xxii + 1174 pp. LCCN TK5105.55 .S74 1994. US$52.75.

Wilkins:1994:MNI

[WSA^{+}94] Jean E. Wilkins, Sara Suelflow, Barbara McFadden Allen, Beth McNeil, and Bill Erbes. Mining “the net” for information using Internet for reference, 1994. 1 videocassette (120 min.).

Wu:1997:ATN

[Wu97] Theodore Wu. ATM Transport and Network Integrity. Academic Press, New York, NY, USA, 1997. ISBN 0-12-765558-1. xviii + 347 pp. LCCN TK5105.35 .W8 1997. US$74.95.

Watters:1996:IPP

[WvRA96] Aaron Watters, Guido van Rossum, and James C. Ahlstrom. Internet programming with Python. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55851-484-8. xviii + 477 pp. LCCN QA76.73.P98 W38 1996.

Welsh:1994:II

[WWBD94] David Welsh, Sally T. Wood, Laura Blobaum, and Robert Dunlap. Interacting with Internet, 1994. 1 videocassette (30 min.) guide.

Wyatt:1994:NI

[Wya94a] Allen Wyatt. Navigating the Internet. Jamsa Press, Las Vegas, NV, USA, 1994. ISBN ???? iv + 60 pp. LCCN ????

Wyatt:1994:SI

[Wya94b] Allen L. Wyatt. Success with Internet. Jamsa Press, Las Vegas, NV, USA, 1994. ISBN 1-884133-01-0 (paperback). xv + 429 pp. LCCN ???? US$29.95.

Wyatt:1995:MI

[Wya95a]  Allen Wyatt.  La ma-

gia de Internet. McGraw-Hill/Interamericana de Mexico, Naucalpan de Juarez, Mexico, 1995. ISBN 1-884133-01-0. xv + 457 pp. LCCN ????

Wyatt:1995:SI

[Wya95b] Allen Wyatt. Success with Internet. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1995. ISBN 0-7895-0087-6. xvi + 443 pp. LCCN TK5105.875.I57W93 1995.

Wyatt:1995:YWI

[Wya95c] Allen Wyatt. Your Windows 95 Internet surfboard. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-721-1. xx + 420 pp. LCCN QA76.76.O63W94 1995.

Wang:1998:SIG

[WYLO98] X. F. Wang, X. Yi, K. Y.

Lam, and E. Okamoto. Secure information gathering agent for Internet trading. Lecture Notes in Computer Science, 1544:183–193, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:SAM

[WYLZ99] X. F. Wang, X. Yi, K. Y.

Lam, and C. Q. Zhang. Secure agent-mediated auctionlike negotiation protocol for Internet retail commerce. Lecture Notes in Computer Science, 1652:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xerox:1981:ITPa

[Xer81a] Xerox Corporation. Internet transport protocols. Technical Report XSIS 028112, Xerox Corp., Stamford, CT, USA, 1981. viii + 61 pp.

Xerox:1981:ITPb

[Xer81b] Xerox Corporation. Internet transport protocols. Technical Report XNSS 028112, Xerox Corp., Stamford, CT, USA, December 1981. viii + 64 pp.

Xie:1994:MCI

[Xie94] Weihong Xie. Mobile connectivity with Internet: a jacketed forward approach.

Thesis (M.S.), University of North Carolina at Charlotte, Charlotte, NC, USA, 1994. vii + 90 pp.

Xopen:1988:XPG-7

[X/O88] X/Open Company, Ltd. X/ Open Portability Guide, Networking Services, volume 7. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-685892-9.

Yourison:1995:IS

[YA95] Karola Yourison and Special Libraries Association. Internet security, June 1995. 1 sound cassette.

Anonymous:1996:YIL

[Yah96] Yahoo! Internet life, 1996. ISSN 1088-0070. Ziff-Davis Pub. Co., Boulder, CO, USA. Includes CD-ROM.

Yan:1988:DEC

[Yan88] Xiong Yan. Design of an experimental computer LAN interconnection in university environment. Computer Networks and ISDN Systems, 15 (5):369–373, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yanoff:1993:SIC

[Yan93] Scott Yanoff. Special Internet connections. Technical report, University of Wisconsin-Milwaukee, Milwaukee, WI, USA, 1993. 17 pp.

Yanoff:1994:SIC

[Yan94] Scott Yanoff. Special Internet connections. Technical report, University of Wisconsin-Milwaukee, Milwaukee, WI, USA, 1994. 25 pp.

Yasin:1995:MHN

[Yas95] Sa’iyd Yasin. A manual on how to navigate the Internet for fourth and fifth grade science educators. Thesis (M. of Ed.), Central State University, Wilberforce, OH, USA, 1995. 77 pp.

Young:1995:IF

[YL95] Margaret Levine Young and John R. Levine. Internet FAQs. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-476-X. xxxii + 412 pp. LCCN TK5105.875.I57Y68 1995.

Yeager:1996:WST

[YM96] Nancy J. Yeager and Robert E.

McGrath. Web server technology: the advanced guide for World Wide Web information providers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-376-X. xv + 407 pp. LCCN TK5105.888.Y43 1996. US$52.95.

Yee:1999:WFD

[YmtEC99] Michael Townsend Yee and Cal Poly master’s thesis Engineering College. Web forms and dynamic HTML: simplifying data transfer on the internet. Thesis (M.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1999. xi, 68 pp.

Yeung:1997:OCR

[YN97] K. H. Yeung and K. W. Ng. An optimal cache replacement algorithm for Internet systems. In IEEE Computer Society. Technical Committee on Computer Communications [IEE97], pages 189–194. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Young:1994:IWD

[You94] Margaret Levine Young. The Internet for Windows for dummies: starter kit. – For dummies. IDG Books, San Mateo, CA, USA, 1994. ISBN 1-56884-237-6 (paperback). xxxii + 416 pp. LCCN ????

Young:1995:IRW

[You95] Terry Leland Young. The Internet: the role of the Webmaster. Thesis (M.A.), University of Texas at Austin, Austin, TX, USA, 1995. xii + 132 pp.

Tseng:1994:ISW

[yTZ94] Jui yuan Tseng and Ruiyuan Zeng. Internet shih wu shou tse. Ti 3 po tung hsun wang lu hsi lieh tsung shu; TJ14. Ti san po wen hua shih yeh ku fen yu hsien kung ssu: Tsung ching hsiao chih tao chu pan yu hsien kung ssu, Tai-pei shih, Taiwan, tsai pan, hsin pan. edition, 1994. ISBN 957-23-0274-4. various pp. LCCN ????

Yu:1993:IVM

[Yu93] Bernard Yu. An Internet voice mail protocol. Project (M.S.), Electrical and Computer Engineering, University of Kansas, Lawrence, KS, USA, 1993. v + 70 pp.

Yourdon:1998:TBW

[YY98] Ed Yourdon and Jennifer Yourdon. Time Bomb 2000: What the Year 2000 Software Crisis Means to YOU! P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-095284-2. xxv + 416 pp. LCCN QA76.76.S64 Y68 1998. URL http:/

/www.phptr.com/year2000;
http://www.y2kinfo.com.

Zazulak:1995:IGI

[Zaz95] Eleanor Zazulak. The Internet: a guide to Internet use in the federal government. Planning and Communications Directorate, Treasury Board of Canada Secretariat, 1995. 2 + iii + 45 pp. “July 17, 1995.” “Author/information: Eleanor Zazulak”–Prel. leaf [1]. “Developed by the Information, Communications and Security Policy Division, Financial and Information Management Branch, Treasury Board Secretariat”–Prel. leaf [2]. Includes bibliographical references.

Zakalik:1995:GGI

[ZB95] Joanna M. Zakalik and Sara Burak. Gale guide to Internet databases. Gale Research, New York, NY, USA, 1995. ISBN 0-7876-0198-5 (paperback). xli + 478 pp. LCCN TK5105.875.I57 G34 1995.

Zegura:1997:QCG

[ZCD97] Ellen W. Zegura, Kenneth L. Calvert, and Michael J. Donahoo. A quantitative comparison of graph-based models for Internet topology. IEEE/ACM Transactions on Networking, 5 (6):770–783, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1997-5-6/p770-zegura/.

Anonymous:1995:ZIL

[ZD 95] ZD Internet life, 1995. ISSN 1088-9914. Ziff-Davis Pub., Boulder, CO?

Zhang:1993:RNR

[ZDE93] Lixia Zhang, Stephen Deering, and Deborah Estrin. RSVP: a new resource ReSer-Vation Protocol. novel design features lead to an Internet protocol that is flexible and scalable. IEEE network, 7 (5):8–??, September 1, 1993. ISSN 0890-8044.

Zeff:1996:NGI

[Zef96] Robbin Lee Zeff. The nonprofit guide to the Internet. Nonprofit law, finance, and management series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-15359-1 (paperback). various pp. LCCN HD62.6.Z43 1996.

Zelwietro:1995:PAI

[Zel95] Joseph Zelwietro. Progressive activism on the Internet. Alternatives, 21(3):16–??, June 21, 1995. ISSN 0300-9734.

Zeltserman:1999:PGS

[Zel99] David Zeltserman. A Practical Guide to SNMPv3 and Network Management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-021453-1. xiv + 337 pp. LCCN TK5105.583.Z45

1999. URL http://www.
phptr.com/ptrbooks/ptr_
0130214531.html.

Zhang:1994:SEU

[Zha94] Haiyan Zhang. The study of the educational use of Internet by the faculty at California University of Pennsylvania. Research project (m. ed.), California University of Pennsylvania, California, PA 15419-1394, USA, 1994. iv + 42 pp.

Zhang:1996:FFI

[Zha96] Allison B. Zhang. File formats on the Internet: a guide for PC users. Special Libraries Association, Washington, DC, USA, 1996. ISBN 0-87111-441-0. xiii + 161 pp. LCCN QA76.575.Z47 1995.

Zhou:1984:DIB

[Zho84] Songnian Zhou. The design and implementation of the Berkeley Internet name domain (BIND) servers: research project. Thesis (M.S. in Electrical Engineering), Department of Electrical Engineering, University of California, Berkeley, 1984. 47 pp.

Zillman:1995:KGI

[Zil95a] K12Link: … global Internet resources and databases, page various, 1995. ISSN 1078-1447. Marcus P. Zillman, Marco Island, FL, USA.

Zillman:1995:LGI

[Zil95b] LawLink: … global Internet resources and databases, page various, 1995. ISSN 1077-9639. Marcus P. Zillman, Marco Island, FL, USA.

Zillman:1995:RGI

[Zil95c] RefLink: … global Internet resources and databases, page various, 1995. ISSN 1078-1455. Marcus P. Zillman, Marco Island, FL, USA.

Zillman:1995:TGI

[Zil95d] TheoLink: global Internet resources and databases, page various, 1995. ISSN 1078-0556. Marcus P. Zillman, Marco Island, FL, USA.

Zimmermann:1995:OPU

[Zim95a] Philip R. Zimmermann. The Official PGP User’s Guide. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-74017-6. xviii + 127 pp. LCCN TK5102.85 .Z56 1995. US$14.95. URL http:

//www-mitpress.mit.edu/
mitp/recent-books/comp/
pgp-user.html.

Zimmermann:1995:PSC

[Zim95b] Philip R. Zimmermann.

PGP: Source Code and Internals. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-24039-4. xxi + 907 pp. LCCN TK5102.85 .Z58 1995. US$55.00. URL http:

//www-mitpress.mit.edu/
mitp/recent-books/comp/
pgp-source.html.

Zeeman:1993:SID

[ZM93] Johan Zeeman and Dennis MacKinnon. Spreading the ILLINET: a distributed network for library resource sharing in Illinois. Internet research, 3(2):60–??, Summer 1993. CODEN IRESEF. ISSN 1066-2243.

Znati:1994:HOT

[Zna94] Taieb Znati. Hands-on TCP/ IP. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30149-0. 550 pp. LCCN ???? US$51.95.

Zornetzer:1995:INE

[Zor95] Steven Zornetzer. An Introduction to Neural and Electronic Networks. Academic Press, New York, NY, USA, second edition, 1995. ISBN 0-12-781882-0. xxiv + 500 pp. LCCN QA76.87 .I5895 1995. US$133.95.

Zeltserman:1998:BNM

[ZP98] Dave Zeltserman and Gerard Puoplo. Building network management tools with Tcl/Tk. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-080727-3. xiv + 429 pp. LCCN QA76.73.T44 Z45 1998. URL

http://www.prenhall.com/
allbooks/ptr_0130807273.
html.

Zhu:1998:SUP

[ZTLV98] Jinsong Zhu, Maria Toeroe, Victor Leung, and Son Vuong. Supporting universal personal computing on Internet with Java and CORBA. In ACM [ACM98], page ?? ISBN ???? LCCN

????  URL http://www.
cs.ucsb.edu/conferences/
java98/papers/upc.pdf;
http://www.cs.ucsb.edu/
conferences/java98/papers/

upc.ps. Possibly unpublished, except electronically.

ZuSeiker:1982:TGI

[Zu 82] Frederich Meier Zu Seiker.

A Telex gateway for the Internet. Technical Report MIT/MAC/TM-222, Massachusetts Institute of Technology, Laboratory for Computer Science, Cambridge, MA, USA, May 1982. 59 pp.